A Bibliography of Accuracy and Stability of Numerical Algorithms

Nicholas John Higham University of Manchester Department of Mathematics

Manchester M13 9PL England

Tel: +44 (0)161 275 5800 FAX: +44 (0)161 275 5819

E-mail: higham@ma.man.ac.uk (Internet)

19 August 2024 Version 0.75

Abstract

This is a bibliography of publications in the book Accuracy and Stability of Numerical Algorithms.

Title word cross-reference

(0, 1) [Chi93]. (A, B) [Kåg93, KP94]. (A − λB) [Kåg84]. (AR − LB, DR − LE) = (C, F) [Kåg94]. (Y A − DZ, Y C − BZ) = (E, F) [Chu87]. 16 × 16 [Tod53]. 27 [Gol67]. 2 × 2 [Lev80, Uhl81, Uhl85, Win71]. 3 × 3 [BL91, Fra89, Lad76]. 8 [Gol67]. ^{3} [HW75]. A [Dem79, PW69, SS91]. A − λB [DK93a, DK93b]. A^{−1} [AYB90, Alb77]. A^{−1}A^{∗} [DJ74]. A^{1/2} [Alb77]. A^{T} Q + QA = −C [Bar79a]. {a_{ij}} [Asp59]. a_{ij} = 0 [Asp59]. A′Q + QA = −C [HMW77a]. Au = λBu [Wil77a]. AX + X^{∗}A^{∗} = C [LR83]. AX + XB = C [BS72, GNV79]. AX + Y B = C [Ziȩ85, Ziȩ87]. AX − XB = C [Hig93d, SN91]. AX − XB = Y [BR96]. AX − XB^{T} = C [Bye84]. AX − Y B = C [BK79]. Ax = b [Har80]. Ax = λBx [SS91, PW69, Wil76]. Ax = λCx [SE87]. AX^{2} + BX + C = 0 [Dav83]. AX^{2} − I = 0 [Laa58]. AXA^{∗} = B [Bak84]. AXB + CY D = E [BK80]. AXB − CXD = E [Chu87]. AXB^{T} + CXD^{T} = E [GWL^{+}92, GLAM92]. AXD − BXC = E [Ept80]. AY + Y B = −Q [Kre72]. B [PW69]. B^{∗} [SS91]. B_{n} [Ebe71a]. Bx = Ax [Wil77a]. C_{n}(x) [Hig96b]. c_{p} [McC67]. cos(iπ/n) [Oli75]. ΔH [Gab87]. e [Bai88a]. E(X) = E{E(x|Y)} [Eni73]. ℓ^{p} [Boy74]. ℓ_{1} [SW87, Ziȩ87]. ℓ_{2} [GV74, Wat91c]. f(x) = 0 [Kah79]. f(X) = A [EU92]. GL(n, C) [DJ74]. H [ABO88, Bea79, Har86a, Kol95, Neu86, NP84, Pol87]. H = W [MS64]. H_{∞} [PKPL91, BBK89]. I + A + ⋯ + A^{N−1} [LN92]. I − 2ww^{H} [Wil65c]. ∫_{0}^{∞} exp(−ct)dt/t(1 + t^{2}) [Act74]. ∫_{0}^{∞} \frac{sin x}{x} dx [Har09]. j > i + p [Asp59]. k [Sor92, Wat92g]. L_{1} [RR73, OW85, Rot84]. l_{2} [CP76, Sha85]. l_{i}nfty [Rot84, BJ89, Cad73, Cad74, CL92]. L_{p} [Fis83, Li91, MS74b, Wat77, GS83, FGH71]. λ [SW81]. LDL^{H} [Bar91b]. LDL^{T} [FP74b, Sun92b]. LU [AO86, ABO88, Bar91b, Cha84, Cry73, DD89, DHS95, ES92, FP81, FNP82, FP86b, Har86a, Mat84a, Par67, VC81, DD90, GMSW87b]. M [AO86, Bar86b, Bea79, DL92b, Fan60, FP81, Joh82, JE95, Neu86, Peñ95, VC81, FT88, TW92b]. n [Smi68a, TSY88, TST86]. n × n [Kah73a, Chi93]. p [Bax91, Hig92b, HW79b, Yip86]. π [Bai88b, Bai88a, Bec71]. P ≤ A ≤ Q [Har80]. p ≤ b ≤ q [Har80]. p → 1 [Fis83]. px = xq [Syl84]. QR [BJ74, Dem89a, Dem90a, DV92a, Fos90, GP93, Gen75, Gul95, Hig90h, Hig91c, Kar74, Kni95, KT76, Mas94, Par67, RS88a, SK77, Ste77b, Sun91, Wat92a, Wat95, Zha93a]. QZ [Wil79]. S_{n}(x) [Hig96b]. sep^{−1} [KP92]. T A − BT = C [Hea77]. we^{w} = χ [FSC73]. x [Kah80b]. X(0) = C [HMW77b]. x ∗ (1/x) ≠ 1 [Ede94b]. x ∗ [1/x] < x/x = 1 [Kah94c]. X = A_{1}X + XA_{2} + D [HMW77b]. x^{(i+1)} = P x^{(i)} + q [AR92]. XA + AY = F [HMW77c]. XA + BX + C = 0 [Mil88b]. XA + BX = C [Smi68b]. XA − BX = R [Dat88]. XA = A^{T} X [DD88]. XY = A [Ziȩ83]. Z [Kah73a]. Z^{0} [A^{+}90]. ∥A^{−1}∥_{∞} [Dem79, Var76a]. |X′X| [Bat83]. ∥Z − Z^{∗}∥ ≤ ∥Z + Z^{∗}∥(log_{2} n + 0.038) [Kah73a].

-Approximation [Wat77, MS74b]. -Decomposition [Kar74].

-Decompositions [Mat84a, Har86a]. -Digit [Gol67]. -Factorization [Cry73, BJ74]. -Factorizations [Cha84]. -form [FT88]. -Matrices [AO86, ABO88, DL92b, Fan60, FP81, Joh82, JE95, Neu86, NP84, Peñ95, Pol87, VC81, Gab87, Har86a]. -Matrix [Bar86b, Kol95, SW81]. -Norm [Bax91, Cad74, Hig92b]. -operators [Bea79]. -problem [Kåg84]. -Scaling [GV74, Sha85]. -Solutions [CP76, Ziȩ87]. -step [Sor92]. -th [Smi68a].

1 [BH87, Gar73a]. 1-Structure [RB89]. 1.0b [Ken93b]. 10967-1 [LIA93]. 10P [DD89]. 13th [GW90]. 14th [GW92a]. 15C [Hew82, Kah83a]. 15th [GW94]. 16 [For60]. 16th [Gol77, Sal81]. 1900-1920 [Bel31]. 1930s [Anoxxb]. 1940s [Gar77]. 1950s [Anoxxb]. 1993 [BCEP94, LIA93]. 19th [Gol77, Sal81]. 1A [Gar90b]. 1B [Gar90c].

2 [Bus90, DD89, Gar73b, Gar90e]. 2-by-2 [Mol74].

3 [AHB87, DCDH90b, DCDH90a]. 3-D [AHB87]. 3090 [DD89]. 3090/VF [DD89].

48G [Hew93].

50th [Gau94]. 52 [Kah77]. 524 [Bre78b]. 5th [Maz76].

600J [DEKV92]. 60th [Hig92c]. 64 [PW85]. 65-1 [Wah65]. 6600/7600 [DM71].

722 [CC93]. 754 [Hou81]. 754-1985

[IEE85]. 77 [BM79, Wor89].

8 [Duf89a]. 8087/80287 [Fri85].

8087/80287/80387 [Sta88]. ’80s [Gid86]. 854-1987 [IEE87]. 860 [HGD90]. 8X [MR87].

90 [Bai94b, BGA90, DCH^{+}95, MR90].

A. [BDDC88, Bod49, Bri92, CD86, Inc92]. Abscissa [Bag94]. Absolute [BSW61, Bha93]. Abstract [Ben87a, Gal94, Jac53]. Abusage [Par73]. Abuse [GH70]. Abuses [Fin94, Rat85]. Academic [HW79a, Lue87, Wol85]. Accelerate [BLS91]. Accelerated [Bal80, LPS92]. Accelerating [Fox69]. Acceleration [Dax90c, Dax90b, Pan84b, Saa84]. Accent [Hon91]. accept [Ste67]. Acceptability [BRB76]. Acceptance [Coo90]. Access [BRK^{+}87, BDCS88, Day83a]. Accidental [Led87]. Account [Wat80b]. Accumulated [Bar81a, Lin76]. Accumulation [Rad48]. Accumulators [Ros65]. Accuracy [BRB76, BRR85, Boh77, CGR93, DK85, DC77, DS95, Don82, Don83, GS90b, GAKK93, Gol67, Hig89a, Hig93a, Hig95b, Hig96a, Hof89a, JW86, JI92, Kah90c, Kah94a, Kos93, KC73, Lin74, Mal92a, Mal92b, PW74, ST86, Tsa91, Tsa92, TW92b, DMW83, CL79, DSH95]. Accurate [AGL95, AS85, BD90, BPE92, DK88, DK90, DV92a, DG93, EKNX93, FP92a, GB91, Gou86a, Hig90f, JW85a, Kah66a, FB91, Lin75, Lin70, Mal71, Mas94, Mat93a, Mat95a, MNP84, MT86, Pri92, Sla92, Tan90a, UvG89, AL75]. Accurately [Cli90, SW90, GP90]. ACE [Tur45, Asp81, CK81, CD86, Wil54a, Wil54b, Wil75c, Wil80]. Achievements [Par85]. Acid [WC53]. Acids [WC53]. ACM [ACM87]. Acoustic [LPBW87]. ACRITH [BRR85, BRK^{+}87, JW86, KL85]. Acrosin [Har76]. Acta [act94, act95, Hig96c, Ise92]. Action [Dix76, Wag91]. Active [CH88b]. Activity [CK80a, CK80b, CK81]. Acute [Inv93]. Acyclic [DG93]. ad [JBNP91]. ADA [Geh89, CH80, FG84, FKR86]. Adam [Rol75b]. Adams’ [JS77]. Adaptive [Bis88a, FGP91, PP92, Ric75, SB92, BL90, Bur90]. Adding [Vit68]. Addition [FG82, FT86, Møl65b, RK75, RS88b, Swe65]. Additions [Syl50]. additive [DK84].

Addressing [RG92a]. Adds [Kah91b]. ADI [Wac92]. Adjectives [Rol67a]. adjoint [Putxx]. Adjustment [WH16]. Admissible [OPW65]. Admits [Mar94]. Ado [Kah87a]. Advanced [And90a, BRR85, Geh89, Hew82, KD93, KD95, Oxf89b, Und80, Wil95]. Advances [Grä94, Wil77b]. Advantages [Kah94b, Rei92]. Adventures [Ula76, Fey85]. Advice [Mau91]. Advisory [Boi90]. Aeronautical [BCD95]. After [Fox90b, Ruh69]. After-Dinner [Fox90b]. Again [Mac95]. Against [Jes93]. Age [Hwa95, Kel88, Mer90]. aggregated [Ste83a]. Aggregating [Pan78, Pan84b]. Agile [Pas85]. AGM [BB87]. Agonizing [She94]. Agreement [tBK84]. ahead [CH92a]. Aid [GP59, Ver49, BDCS88, DBKF90]. Aided [BB92, MS91]. Aids [Gou73, Mac77b, MFGK82, CBB54]. Aims [Col77, SE62]. Airline [Bro89]. Aitken [Min74]. AIX [IBM93]. Alan [CK84, Hof85c, Hod83, Tur59, Whi87]. Albers [Alb91]. Album [Gri89]. Alexander [Min74]. Algebra [Ant94, AR94, Bab72, Bas83, BF93a, BL77, BP92b, Bro91b, Bun87, CJLP92, CJLP93, CDPW92, CDW94, Cia89, ILA93, Dat85, DD86, DST88, DDC^{+}87, Dem92d, DHvdV92, Dem93a, DHvdV93, DGL91b, DGL91a, DGK84, DH90, DvdGW92, DW93, DPW94, DPW95, DEK95, DW95, DCHH88b, DCDH90b, DCHH88a, DCDH90a, Ede93b, Ede94a, FF63, FF64, Fad59, For67, Fox64, Gal94, GJM87, GJMS88, GPS90, GMW91, GAKK93, GL91b, Gol73b, GHMP74, Gra87, Hag88, HK87, HWH91, Hel78, Hig85, HS87, Hig94d, Hig95d, Hil88, Ing95, Jac53, Kah66b, Kah90d, KS88, KS92, Kos93, LS82, LS83, Lau85, LHKK79, Lay94, Leo93, LJ93, Mal92a, Mal92b, Mal93, MM65, Mir61, Mol78, MGM93, Nas79, Nas90a, Nob69, Riv90, Sto91, Str88, Str93, Tod77, Tre90, Van91, Wil54a, WR71, Wil74b]. Algebra [Wil77b, Wil78a, ABB^{+}90, ABD^{+}90, BD89b, Bis90a, DS86a, DH86, DS86b, DMdB90, Gla82, Gro87, Nas85, Sch89a, SR90, Str89, Tho92]. Algebraic [AL84, Ata40, BM75, Bye85, Bye87, Dah78, DST88, EMN88, FKF68a, For53, FM67, FHW48, GL86, GL88, GL91a, GL95a, Ham82a, HV94, Inc83, Inc84, KH90a, Kup71, Lau79, Lev58, MD92, Mos71a, Rob80, SH83, vdS70a, vdS70b, Spe91, Sto75b, Sto77, Wil63b, Wil65a, Wil68b, Wil70, AL75, Hil29, KLW90, Pet82, Sch91, Ioh82]. Algebraicznych [Wil67a]. algébriques [FKF68b]. ALGOL [Lin72, Nic70, BFBS64]. Algorithm [AA94, Ant85, AL85, BD91, Bai88b, BF88, BLS91, Bai93a, BHH93, BIMF76, Bar91a, BS72, BB71, BP92a, BGD^{+}89, Bre91, Brü95b, Buc94a, Cad88, CH92a, CI95a, CV89, Cod93, CR76, CH88b, CT65, Cox75, CH89, Dat87, Dav83, DH89, DK85, DGL91b, Don82, DS88, DS92, DS93a, DCHH88b, DCDH90b, DHSS94, EK89, EH91a, EP87c, Fan93, FM85, Fra89, Fre93, GL94, GJMS88, GWL^{+}92, GM73, GS71, Gol73b, GB86, GE94, GE95a, GE95b, Gul92, HK87, HW95, Hen92, HV94, Hig91b, HP94a, HP94b, HR90, JP91, JS92, Jon62, KF63a, KF63b, KF63c, Kar74, Kas95, KLP93, KR78, Lad76, Lan93, LPW83, LN92, LZS91, LZ92, LL93, LZ91, LPS87, MS93a, McK62, MS78b, Mol2b, MS78c, ML85, NRT92b, Pan78, PKL90, PR81, PSS82, Par90d, Par94, PPY93, Phi87]. Algorithm [PA92, RR73, Rob90, Roh92, RG92b, Ruh94, SB80a, SS91, Smi62, Smi65, Smi91, Ste71a, Ste85a, Swa93, Tha61, Tom92, Tor89, Tor91, Tsa81b, Tur91a, War76, War77, Wat92a, Wat95, Wat91c, Wil79, Win68, DW87, APP88, AL73, AD90a, Bar65, Bis89a, Bis91, Bog74, BBdH91, Cad73, Cli72, CM89, Cyb80, Cyb84a, DD88, DE84, DS87, Dub91b, Fil81, FG86, Gra86, JP90, KR80a, KR80b, Kie73, LM85, Van87b, Mar63, OW85, Paa71, PV86, Par90b, Ruh70, SK77, Ste79a, Ste90h, Ste91d, Str87, Wat82, Wee89, Bre78b, Cod88a, CC93, Cra86, KR80a, For60, Hig88c, Hig91a]. Algorithmen [Spi76]. Algorithmic [BBD^{+}94, CdB80, FS83, Knu85a, Cad74]. Algorithmics [BB88b]. Algorithms [AHU74, AG92, AL86, AL84, BDG94a, Ban76, BR77, BZY93, BG76, BS89a, BL80, BP94b, Bis88b, BE73a, Bjö91b, BBdHS95, BKO94, Bre70a, Bre70b, Bre73a, Bro91c, Bun87, BDV89, Car89, CGL83, CF92, CLR90, Csa76, CW85, DGKS76, DST88, DHP94, DH92, DL93, DL94, DGK84, EHHR88, ES89, Fei82, FP92a, FP74a, Fis74, FS76, Fox69, FP92b, GPS90, GHN^{+}90, Gan90, GL91a, GM74a, Geo78, Goe94, GO89, Gra93a, Gre94, GK69, GC91, GA94, HHKR93, Hel74, Hel78, Hig86b, Hig87b, Hig90k, Hig95b, Hig96a, HJ81, HJ88, Hof89b, Hul82, JáJ92, KP92, KP93, KP94, Kah73b, KLP92a, KLP92b, Knuta, Knu81b, Kos93, Koz92, Kre76, Kro87, KW92, LPS92, LS80, Le85b, LO82, Lu95, Mal72, Mal93, MC90, Mat92a, Mat95b, MBI77, Mil88a, Mil78, MW80, Mod88]. Algorithms [MW86, Noc92, OVR90, Osb85, Pan78, Pap92, Par71, Pow75, Pri91a, Pry85, Qui87, Row90, Ruh73, Sch88a, Sch88b, SHW86, Sed88, Sed90, Sma85, Smo95, SW94, SW82, Spe91, Stp93, Stu80, Stu81, Tan91, Tem75, Tor93, Tra73, TSY88, Tsa81a, Tsa83, UvG89, Vav92, Vit68, WE91, Wat92b, Wil89, Wil68c, Woź80, BPT90, BC90, Bye86, Cal86, Cyb78, Cyb79, Cyb81, CB90b, Den84, DS86a, DH86, DO90, DBKF90, DGR79, Duf89b, Gro90, HLW83, Kub62, LHPW87, OW90, Pai81, Par67, Pow87a, RS88a, Wri90, DK93a]. Alliant [Duf89a, DD90]. Allied [Gow84]. Almost [Ber71a, BG90, Kah70, Kah88, PS79, Wil60a, Wil59, Wil68a, Ano78b, Ros69, Ruh75, Wed73a]. Almost-Commuting [Kah70]. Along [Gor93a]. alphabet [Hil29]. Alphabets [Mur76]. Also [Rei93]. Alternating [Deu83, GB86]. Alternative [CH89, Dig91, ST80, Fre80]. Always [Ogr88]. Amdahl [Gus88]. America [Bry89, DAM88, SD86, Tru85]. American [Bal86, Caj25, Hou82, Neu88, ‘Vi52, Wil72a]. Ames [Bai93c]. Among [Par76]. Amplitude [LM90]. Anagrams [Ber73]. Analog [CH87, HM90a]. analysed [Wed78]. Analyses [Mei83, Mil78, SB88, Wil74a]. Analysing [Bra87]. Analysis [Abd71, Abe88, AHU74, APP88, AM94, And71, Ano78a, Ano93, AL85, AL86, ACD^{+}93, Atk89, Atk93, Bab69, Bar81b, BB85b, Bar88a, BZ90, Bar91a, Bar93a, Bar91d, Bel70, Bev69, Bjö87, Bjö91a, Bjö91b, BBG92, dBP77, BB92, Bre70b, BF90, BC86, BT92, Bun71a, BF89, BF93b, But87, Car59, CCKT83, CI95a, CF92, CH94, CL90a, Cle78, Cod81, CdB80, Cyb83, DK84, Dei86, Dem87d, DH95, DR91, Eir93, EWK90, EP94b, Ell94, For59b, FP68, Fox78, Fox90a, Fro69, GL94, Gar72, Gas83, GS90a, Gen69, Gen75, GK90a, Gol77, GO78, Gol84, GZ93, Gow75, Gow84, Gre52, Gri84, GW86, GW90, GW92a, GW94, Gul92, Gul95, Han69, Han92, HN81, Har58, Hen82, Hen86, Hen64, Hig87b, Hig90a, Hig90k, HK93a, Hil74, Hir65, HJ85, HJ91]. Analysis [Hou53, Hou64, Mül83, IK66, IP87b, Jac77, JP91, JO74, JR82, Kåg94, Kah72b, Kah91a, Kah93a, Kie87, KC91, Kni93, Kör88b, Koz92, Kri80, KPS77, Lan56, LS80, LPW83, LS82, LS83, LM90, LS91, Lin74, Van84b, Lus89, Man81, Mat92a, Mat95b, MS91, Mik91, Mil71a, Mil75d, Mil75a, Mil76, MW80, MS78b, MS78a, MM83b, Moo66, Moo79, Neu85, NP84, New73, New74, Nic77, O’C93, Oli77, Olv83a, Ort63, Ort72, Osb85, Pai73, Par71, Par90c, PEM81, Pry81, Pry85, RR78, Ram71, RJ93, Ric83, Ric92, Row90, Sal89, Sal81, Sau78, Sca50, Sch68a, SRS73, Seb77, SB94, Sma85, Sma90, Smi75, Söd93, Sor85, Spe80, Ste73b, Ste71a, Ste92a, SB80b, Sto77, Stp93, Stu80, Stu85b, Stu85c, Swe65, SY78, Tan91, Tau50, Tay82, Ter85]. Analysis [Tod62, Tre92a, Tsa81a, Tsa83, Tur89, Uhl92b, Van87a, VV91, Var62, VA86, War77, Wat92b, Wen66, vW66, Wil60b, Wil61a, Wil62, Wil65c, Wil68b, Wil68c, Wil71a, Wil85, Wil86b, Won90, Woź74, Woź80, Wun48, YC95, Zha93a, Ziv82, AP81, BD76, BDCS88, Bul80, CM89, Cyb78, Cyb81, Dem88b, Dwy67, FNO87, GMSW86a, GK89, GV80, HK65, JP90, LT88, Lin88, Van83a, Luk84, Luk85, MFGK82, Man82, OW85, Pai80, Par78, Ple86, Rao80, Shr84, Woź78, CGL83]. Analyst [Pap93a, Wil71b]. Analytic [BB87, VG87, BGBMN90]. Analytical [Lan38, Sto75b]. Anatomy [Rol75b]. ancient [Jen77a]. Anderson [Ano78b]. Andrew [CK84]. Andy [Lee88]. Anecdotes [Tro84a]. Angle [Kah89a]. Angles [BG73, BL91]. Anguished [Led87]. Anniversary [Gau94]. Annotated [Col77, GS76a, NR76]. Annual [Ede91b]. Anomalies [KL85]. Anomalous [Ben38]. Anonymous [Hig92d]. ANSI [IEE85]. ANSI/IEEE [IEE85]. Answers [Fox71]. Antenna [LPBW87]. Anthology [Led87]. Anti [GS84]. Anti-Hadamard [GS84]. Anton [Hig95d]. Any [Kah79]. Anybody [Byr88]. Anything [Gea81]. Apology [Har40]. Appl. [Hig96b]. Applicable [Gra87]. Application [Bar79b, Boy94, Ept80, Gar79, GC91, Jon62, Kah93a, Kat60, Nab9x, PV74a, Sor92, Ste80b, Tur36, Wat92e, Ben24, Van77b, Tog86, Ben24, CB90a]. Applications [AB79, Bar86b, Bar93b, BS68, Bar90a, BL88, BBG92, Bon91, BF90, Car89, Cas79, CH92b, Cyb83, Dat88, DKW82, DK93a, DK93b, EAS94, ES89, Far87b, Fer94, Fie86, Gol91b, GB89, Gra69, Gro89, GDD94, Hal90a, HW78a, Hel74, Hig86a, Hig88c, Hig89d, Hig90j, Hod92, Hou81, JE95, KF87, KL80, Fer95, Lay94, Leo93, MO79, MD64, MP77, Mil71a, MGM93, Moo79, NR76, Nas76, PS91, Pap93b, Ral81, RM71, SV77, STW84, Smi65, Str88, TG81, Ter85, Tyr91, WH16, AG88, Bar75b, Dah83, Dwy67, FU69, Gla82, GS58, ITH88, Kie73, LHPW87, MG91, MM88, AR94, DK93b, Hig95d]. Applied [BI77, Bar89, Bas83, CH94, Gar73b, Gar82b, GO78, HY81, Hag88, HR90, Lan56, Nob69, Par64, RJ93, Str86a, Ple86, Str87]. Applying [PR68, PR69b, Woe92]. Appraisal [Lon67]. Approach [ADdR89, BBD^{+}94, BZ93, CdB80, Gru70, HP90, HW91b, Jes93, KM86a, KM83, KM86b, LS80, LS82, LS83, Lau88b, Man81, Olv78, Spe80, Tsa83, Van81, Woo85, Chu86, Chu91a, Gay84, Kub70, VV87]. Approaches [BN87a, GMSW86b]. approaching [DH86]. Approche [Bru89]. approximant [Bou73a]. Approximants [BS92a, Hop82, BHK91, Bou73b, Hal72, Hal74b]. Approximate [CDG91, GL95b, KPJ82, Kov76, Kup71, OP64, Ste71b, Ziv95, AL75, How83a]. Approximating [DGR79, Mat91a, Van84a, Vas90]. Approximation [ADW84, ASS73, And85, Che82a, Che86, CSW83, CH89, Dun82, EY36, GHS87, Gre52, HW88, Hol74b, Lás94, MC90, Pow81, Rei91, Riv69, Riv90, SV77, Sch83a, Sny66, SH93, Tan94, Tre90, VP93, Wat77, Wat80a, Wat83b, Wat92f, Wat92g, Chu91b, Fis83, GS71, Gro90, Van77b, MS74b, OW85, Phi77, RR73, Sch91, SW87]. Approximations [BP67, BBB89, Fik67, FHM67, Gol68, HW80, Has55, FB91, Pow67, SR80, DK83, FGH71, Hal86, LR81, Osb65, Rao80, Ste67]. Approximators [Far85]. Arabia [Uni92]. arbitrarily [Rum91]. Arbitrary [DK93a, DK93b, HS90, Lei89, Pri91a, Sol80, GLW89, Hua75, Mas91, Rad62, Ruh68]. Architecture [HP90, HJ88, IT89, Roj94, HJ81]. Architectures [And90a, DLN^{+}94, Got92, JP94, Rob90, Sch88b, BPT90, DvdG92]. Area [Kah89a]. Argument [Fer94, Fer95].

Arguments [Var76a]. Arise [Nab92]. Arising [CH89, FGS95, Gau69, GMPS92, HW80, Mat84a, PW71, Rut47, Rut52, Var72, BJW85, Sha82a]. Arithmetic [AS92, AB79, AS85, BSBF89, Bar81a, BB85b, BB85a, Bar88b, BZ90, BRR85, BB84, Bre74, BHY80, Bre78b, Bre78a, Bro81, Car59, COT89, Cod73, CCG^{+}84, CC93, CHH83, CW87, Dek82, Dem87d, Dem91, Gar76b, GM74a, Gol91a, Gre94, IEE81, Hen80, Hig91e, HK95, Hou81, IEE85, IEE87, JW85a, JW86, Kah80a, Kah81, Kah83b, Kah86a, Kah86b, Kah87b, Kah89b, Kah89c, Kah90b, KT91, Kah95, KL73, Kie81, KPS77, KM81, KM86b, LIA93, Mal72, MK85, Met77, Mil88a, Olv78, Olv83a, Olv86, Pla88, Pri91a, Ric55, Sch81a, Smi91, Stu81, Stu82, Swa80b, Tan89, Tan90b, Tan92, Tie70, Tur91b, WF82, Wil63a, Yoh73, Yoh79b, Yoh80, Kah90a, Olv83b, VS88a]. Arithmetic-Geometric [BB84]. Arithmetics [Bar81b, Cra79, Pri92, SS84, Yoh79a]. Arithmétique [Pic72, PV74a, Mul89]. Armchair [Mil78]. Arnoldi [Bro91c, LS94, LSV95, Saa80, Sor92, TT96]. ARPACK [LSV95]. Arranging [Mil43]. Array [Luk86, Sch86c, BV86, Van88]. Arrays [BI87, LPBW87, SHW86, Sch83c]. arrival [AGS91, Van87b]. Arrow [AG92]. Arrowhead [OS90, Zha92]. Art [Anh94, BSB92, BR93, EBR87, Fle46, IP87b, Jac77, Knu81a, Knuta, Knu81b, Kuy91, Lan89a, Lit78, Nei82, Nei84, Per91, PFTV86, PFTV88, PTVF92a, PTVF92b, SWD76, Sha87, She93, Spe91, Sta91, Wil85]. Article [Goo92, Hof85a]. Articles [Gar77, Syl50]. Artificial [JMV94]. Artist [Mac77b]. Artistic [Knu93]. Arts [Com83]. Asked [Ola93]. Aspect [Pol57]. Aspects [Bar88a, DD86, Eir93, Gau67, Lau85, Mau91, Mui82, Ric88a, Ruh83, VV91, Wil78b, GS74, Pai87, Ric88b, Sor90]. Assaults [Led87]. Assembling [PH85]. Assertions [BGM^{+}94]. Assessing [EP87c, EP87b, FG84, NS90]. Assessment [For92, HD80, Sch90a]. Assign [Dat87]. assignment [BN87a, CM89]. Assistants [Swa79]. assisted [Vrs86]. Associated [BG77, dBK86, Jor68, Kni93, NP84, Ste73d, Tra66, DD88, HM90b, Vrs86, Wri82]. Asymptotic [BW93, BD74a, CW82, Gau53a, Gau53b, Pan84b, Won90, Wri82, Ste84a]. Asynchronous [Bul91, BF92, Lei89]. Athens [HPP88]. Atom [HW75]. Attainable [Bre73b, Lás94]. Attention [Mad88]. Attitude [Wah65]. attractor [Hén76]. attributes [Lef78]. Attributions [BG89]. AUGMENT [BHY80]. Augmented [ADdR89, Bjö92, BP94c, MT95]. Augmenting [KT91]. August [Gau94, PA88]. Auld [Hal88]. Ausgleichung [Läu61]. Austria [PA88]. Authenticity [DC87]. Author [Pre78, Tho83]. Authors [But92a, CBB54, Gil87, Har76, Lue87, Sta77a, Swa79, Fla71]. Authorship [Rol75a]. Autobiography [Lem61, wQT89]. Automata [KO63, Loe74]. Automathography [Hal85]. Automatic [Ano46, Bai93b, BBG92, Cha66, CR72, Fit93, Gil51, GC91, ITH88, Iri91, Lyn69, LK76, Lyn83, MS76, PdDKÜK83, Ral81, Sal89, SD90, Sto77, Gri88, Tur45, Wil48a, Wil48b]. Autour [Bar79b]. Availability [Jac78]. Available [Boi90, Dek71, Lus89, SH83, TW91b]. Average [TS90, WWWS86]. Average-Case [TS90]. Awaken [Gar90a]. Awakening [vdW75]. Award [ACM87]. Awards [Gar85]. Awful [Kod79, Led68]. AWK [AKW88].

B [dB72, Cox72, Cox78]. B-Spline [Cox78]. B-Splines [dB72, Cox72]. Babble [Cha74a]. Babuskasche [Nic70]. Babylonian [vdW75]. Backgammon [Ber80, Hol73, JC70]. Backward [AM94, ADD89b, BGM^{+}94, dBP77, Bra95, CI94, CM89, Eir93, Fox78, GL95a, Gul95, HH92a, Hig93d, KF63b, Kas95, Lem76, NP84, Spe80, Sun95a, Sun95c, Var91, WKS95, Yal91, Bjö88]. Bad [Hal91a, Pow86a, Tyr94]. Bag [Sta77a, Sta77b]. Balanced [Bus69b]. Balancing [PR69a, Gra71, LHPW87]. Ball [Tao84b, Rot84]. Baltimore [Hig90i]. Banach [Ost73]. Band [Boh75, CKS78, Dem77, DMS84, Hal80, Lin92a, Róz87, Róz88, YI79, BLR90, CMR90, PW69]. Banded [BF81, Bar84, BW93, Dem79, JP93, Lan93, Mac89a, MR89, Vas90]. Bareiss [BBdH91, BBdHS95]. Barnett [Hig90j]. Barrier [Lar83, GMS^{+}86]. Bartels [KR78, Pow87a]. Base [BR69, Mat70, Ack89]. Based [Bai94b, HN81, KLV91, Lin92b, Par92c, Pug92, RR82, TW92a, Wil62, Wil65c, Bar65, Bye86, Cal86, DGK80, KV91, KO91, Vas90]. Bases [TL81, Hil89]. Basic [Ash93, CCC82, DGL91b, DGL91a, DCHH88b, DCDH90b, DCHH88a, DCDH90a, EL69, HK87, KW83, LHKK79, Mac90, Par92c, Tod77, Wil78b, KK80, HHKR93, Hig86c]. Basis [BHR94, CS84, Dun82, GMS^{+}85, Moo36, SB95]. Basis-Kernel [SB95]. Battling [Hom92]. bcc [HW75]. Be [Bar92a, Bus68, Dav92, Dav93a, For53, Gar90b, HR80, Hal85, Kah87c, Kah90b, Kah93b, Kur92, Mad83, Sch86d, Ste81b]. beamforming [Cyb86]. Beats [DC87, Ber80]. Beauty [Pic90, Pic91]. Became [Tau88]. Beeblebrox [Gra92]. Beer [Sch92b]. Before [Cip89, Coo94, Hal90a, Sig86]. Beginner [LR93, LaQ94, Sam83]. Beginners [KD93, KD95]. Beginning [Hol73]. Behaved [Ogr88]. Behavior [Gre89, GS92, GT93, HK93b, Trexx, Ste84a]. Behaviour [Ch85, Gau53a, Gau53b, HW76b, vdSvdV87, vdV90]. Behind [Hal95]. Being [Mas94]. Beitrag [Hil94]. Bell [Rei93]. Bellcore [Lan89b]. Bemerkungen [Alb77]. Benchmark [Don88a, Kar85, Pre93]. Benchmarks [BBB^{+}91, Bro91a, Kah94a]. Benjamin [Lem61]. berechnung [Alb77, Sch33]. Berkeley [Wan88]. Bernstein [FR87]. Best [Ano76, BB71, Byr88, FS55, GV74, Hol74b, Lás94, Mat85, RS88b, Sau78, Sch90c, Sto62, Wat92f, Fis83, FGH71, LR81]. Bestimmung [Rut66]. BETA [Ken93b]. Better [Bar78, Bar82, Bol88, GA94, Kah90a, Pir82, Woe92, Ors69]. Between [AR85, BG73, GS90a, KP93, Syl51b, TZ59, TL81, Gea77, Han90b]. Beyond [CO84, Moh85, TTRD92]. BFGS [Pow86a, Pow87b]. Bi [vdV92]. Bi-CG [vdV92]. Bi-CGSTAB [vdV92]. Bible [PM87, Knu91]. bibliographic [Hil89]. Bibliography [Bee93, Col77, Ein79, GS76a, HH93, NR76, OVR90]. BibNet [Gro95a]. Bidiagonal [CDW94, DDLT91, DK90, GE95a, Cha74b, Chu86]. Big [Pou83]. Bigger [Pou86]. Bilinear [Kre72]. Billion [BBB89]. Binary [Buc59, CC93, Gay90, IEE81, IEE85, Kah86a, Kah95, MA87, Pax91, SS84, Wal71]. Binary-Decimal [Gay90]. Bini [Tsa81b]. binormal [BHK91]. Biography [Gro94, Wil65d]. Biology [Cul85, Dud77, JS83, Mur93, Pec87]. Biometric [Fin94]. Biometrika [Goo92]. Bird [Rus72, Gid87]. Birthday [Hig92c]. Bisection [Bye88, DDR94, BBK89]. Bisectors [Par71]. Bit [Kah87a, Ziv91, Frö90, Tro84b]. Bitmap [Cas93]. Bitmap-Free [Cas93]. Bits [Gol67]. Björck [Hig87b]. Björck-Pereyra [Hig87b]. black [PNO89]. Blacklist [Bie37]. BLAS [ADD89a, CMR90, DCHP90, DD89, DD90, DH92, DHSS94, DMR92, HK87, Hig90e, Ken93a, KL89a, KV91, KLV91, MR89]. BLAS3 [GJM87]. Bledy [Wil67a].

Bletchley [HS93a]. Block

[ADNR92, BD89a, Ban76, BG90, Bun76, Cal86, Cas89, CGM85, DEKV92, Dem83, DH92, DHS95, DL92b, DHS89, Dub91a, EP94b, ES92, FV62, GLO81, GP90, JP91, KP92, Kha72, LJ93, Van86a, MS93a, vMS94, Mat84a, Mat84b, Meu92, Pol87, RG92b, Sch88a, SP88, SHW86, Sha85, SS89, TG81, Var72, AD90a, Bis89a, CM86, DO90, Geo72, GU77, JP90, Van88, Ple86, Vas90]. Block- [ES92]. Block-Cyclic [LJ93]. Block-oriented [Cal86]. Block-Tridiagonal [Var72]. blocking [Bis88a, BL90, SD90]. Blocks [BD93c, BBC^{+}94]. Blockwise [Pol87]. Bloomsbury [D^{+}93, Man90]. Bob [Nei82, Nei84]. Body [Gol83, LS82, LS83, SW91, SW94, App85]. Boeing [DGL92]. Boil [Ark86]. Bombardment [BBD^{+}94]. Boojums [Mer90]. Book [Ano79, Ano83, Ano85d, Ano88, Ano8x, Anoxxc, Anoxxd, BM84, BG89, B^{+}81, EE84, Edm85, For70, Gra90, Hal67b, Hal82a, Har86b, Hin77, HW79a, Hol74a, Hol75, JC70, Kau78, Knu86e, L^{+}64, Led68, Lev89, LL74, Man88, Nei82, Nei84, Rol75f, Swa88, Und80, Won88, Sie87, Knu86c]. Books [Hal88, O’C78, O’C86]. bordered [GP90]. Borweins’ [Bai88b]. Boson [A^{+}90]. Both [‘Vi52]. Bottleneck [RG92a]. Bottom [PS94]. Bottom-Up [PS94]. Bound [HP92b, Jon84, Lás94, Lee95, Tao84a, Var75]. Boundary [AMR88, BR77, BG77, BD86, CH94, Eva71, Kel93, FU69, Sch91]. Boundary-Value [Kel93]. Bounding [Dem79, Hig90b, Var76a]. Boundless [Ske89]. Bounds [AK75, Bar93b, Bar90b, Bar91b, Bar92a, Bar94, Bha93, Bjö91a, Boh75, Bro73, CP95, CG67, CK75, CS89b, CH88b, CR79, DEG72, DKW82, Dei90, DK93a, DK93b, DH93, Des63a, FL74, GI88, GL95c, Gol62, Grc92, Hen62a, Hig90c, Jen82, Jen83, Kåg77a, KPJ82, KH95, Kit95, Kol95, Lás91, Lem75a, Li95, Mat93b, Neu92, O’L90, OP64, OW82, Olv86, Peñ94, Löt83, Löt84, RK95, Sch92a, Sha91, Ste71b, Ste73d, Ste77b, Sun91, Sun92a, Sun92b, Sun95c, Sun95b, WKS95, Wed72, Wil61b, Ziv95, AL75, FHM67, Hig83b, KLW90, Olv83b, Ste83a, Ste91e, Tsa74a, Wat87]. box [PNO89]. Brain [Gra92]. Brains [Ber49]. Branch [Kah87a]. Breaking [KP81, Whi87]. Breakthroughs [Par85]. Brent [BHY80]. Brian [Fre87a]. Bridges [LM90]. Brief [Car58, Jac86, Pét66, AM87, KY88]. Bringing [Kuy91]. Britain [Cro90]. British [Gau94, WCK89, Lav80]. Broadway [Ano8x]. Browne [Cra74]. Broyden [GS78, Gay79, O’L93b]. Bruaset [Hig96d]. Buffon [WT94]. Bug [Hal95, Hop81]. Bugbears [Ber71b]. Bugs [Gla95]. Building [Cha74a, BBC^{+}94]. Built [Lav80]. Bunch [JP93]. Bur. [Hig96b]. Business [CK89, TK89]. buys [Gay91]. BVP [Mat84a]. BVPs [dBK86]. Byte [Buc81].

C [BDDC88, Hig90i, Hig90j, Hig96b, KKL^{+}93, Min74, AS92, BM84, DCHP90, DLN^{+}94, FGMS90, FG84, Gla93, IBM93, KKL^{+}93, PFTV88, PTVF92a, RJLK78, Sed90]. C-XSC [KKL^{+}93]. CADAC [CHH83]. Calculating [Ano48, Bab37, BS74, dB72, BK77, Els76, GK65, Har84b, Kah89a, Leh94a, Rad48, Ver49, Ste75c]. Calculation [AM65, CT65, GK83, Hot43, KE82, LS78, TT96, Wil65b, Bri91, FGH71, KG83, PR69a, Ste90c, Wil58, KF63a]. Calculations [Cha66, EAS94, Fos86, Gol69, Str86b, HH91a, Ste81c]. Calculator [Ano46, Blo48, Kah79, Kah80c, Sch83a, Smi75, KP77]. calculators [Ste81c]. Calculus [Bod56, WR44, Dav73, Pol85]. Calibration [For92]. Call [DK94, Rol76c]. Cambridge [Hig96c, But81, But92a, CK80a, Cry87, Wil75a]. Can [Bar92a, Boa81, Bus68, CEN91, DGGW89, For53, Fos94, Gar73c, Kac66, KP77, Pan84c]. Canceling [Pan78]. Cancellation [Fle86, Gur92]. Cancer [Ber91]. Cannot [GWW92]. Canonical [FS83, GW76, GZ93, TA32, Wil79]. capital [Ste87b]. Capitalization [Boa66]. Careful [Ber65, Ber71b]. Carnival [Gar75]. Carolina [BCEP94]. Carpenters [Ano85a, Ano85b]. carrés [Ben24]. Cartoons [Har91b]. Cascades [SS84]. Cascading [Ros65]. Case [BR77, Buc94b, Dat87, Har76, Jes93, KR65, Tho83, TS90, Uhl81, Ziȩ85, PW85, Uhl85, Wed73a]. Casebook [SE62]. Catalog [Kro92, Kro94]. catalogue [Gro90]. Catch [Mad88]. Cauchy [Tyr91, GK89, GK90a, GK90b, HM90a, HM90b, IIM87]. Cauchy-Toeplitz [Tyr91]. Causes [Mar94]. Cawdrey [SN46]. CCI [DK94]. CDC [DM71]. Celebrating [Gid87, EG91]. Celebration [Red93]. CELEFUNT [Cod93]. celui [Ben24]. Centenary [BCEP94]. Central [SC70]. Century [DAM88, Ewi94, Gol77, Sal81, Gau94, MHR80]. CERFACS [HHK93]. Certain [And90b, CH94, Gra93a, HB59, JW85a, Mil71a, Peñ94, Ste73d, Sto62, Tod49, Tod50, Tod54a, Var72, Buc77, Kub63, dO69, Phi77, WY64]. CESTAC [BC86]. CFD [SS90a]. CG [vdV90, vdV92]. CG-S [vdV90]. CGSTAB [vdV92]. Chain [Tsa81a, Bur90, FM86]. Chains [Bar86b, Bar93b, Gra93a, Hey87, IM94, Mey89, O’C93, O’L93a, Ren72, Ste93a, Ste94c, GM86, HP84, Mey75, Ste83a, SZ91b]. challenge [GMS88]. Challenges [Kah90d]. Challenging [Sch87]. Chamber [‘Vi52]. Chambers [Cha90, Cha91, Cha93a]. champion [Ber80]. Chandler [CR94]. Change [FM85, MP83, DS79a, Ip87a]. Changes [Hym86]. Changing [Lod75]. Chaos [CEN91, Cor92a, Dav93b, PJS92a, Pic90, SSL93, Ste89]. Chaotic [BBP87, CM69, Cor92b, HYG88, SU84].

Chapter [HL86]. Character [Fey85]. Characteristic [FF51, Giv54, HB59, Tau49, Uhl81, Tau70]. Characteristics [Cod73]. Characterization [GV74, LR81, Wat92c, All88]. Characters [Geo78]. Charles [Uhl92b]. Charlie [Gri89]. Charts [Mac77c, Mac77a]. Chasing [BZY93, WE91, Zha92]. Chauvenet [Abb78]. Chebychev [ST75, CR92]. Chebychev-Vandermonde [CR92]. Chebyshev [FP68, Hig96b, Riv90, Sny66, ADW84, Bog74, BKO94, Cle55, Coo67, Cox86, DS95, Dun82, Fun88, GO94, GS71, GO88, Oli77, RO91, Saa84, Woź77b]. Check [Gau83b]. Checking [Bee93, Kah91c, PR93]. Checksum [Par92a]. Chemical [ABG48, Shr84]. Chemistry [EBR87, EAS94, Gar77, GL91b, Rut47, Rut52]. Chemotherapy [Ber91]. Child [Frö90]. China [LYK89]. Chip [Gla95, Mar94]. Chips [SW89]. Choice [BR69, Buc59, Dan76, Dun82, Hul82, SC70, DK83]. Choleski [Gre85]. Cholesky [Ano22, Ben24, BPT90, BBVdH87, Car91, CP95, CDO^{+}94, Dem89b, DOV94, EP94b, FGM95, Gay91, GHL86, GSS96, Hig90a, Kie87, Man79, MR89, Mei83, Pow86b, RS88a, Sau72, Sch93, SE90, Ste79a, Ste93c, Sun91, Sun92b]. Choosing [Ker83]. Chord [Hym86]. Chordal [PPY92, PPY93]. Chords [Pre91]. Chris [Hig95d]. Christoffel [Gal71]. Chronology [Par85]. Circle [CGR90, FV62, SU84]. Circles [Eve69, Eve88]. Circuit [Mar94]. Circulant [Cha88, CNP94, Dav79, CS89a, SE87]. Circulants [Par82]. Circular [Lin92b]. Citation [Bus90, Gar64, Gar72, Gar79, Gar82b, Gar84, Gar90a, Gar90e, Gar91a, Gar90d]. Cite [Gar82b]. Cited [Bus90, Gar73a, Gar73b, Gar77, Gar82a, Gar84, Gar90b, Gar90c, Gar90d, Gar90e, Gar91b]. Clarity [You89b]. Class [ABS84, Cho69, Cle59, DGGW89, Dub83, EP94a, Fai71, KKL^{+}93, MP74, Nab92, Syl50, VG87, Wat91d, Wat92f, Woź80, Rum91, WG78]. Classes [Dav85, Mil71a]. Classical [Mar89, Wil74a]. Classics [Gar84, Sch82, Gar90d, Gar90e]. Classroom [PJS92a, PJS92b, Pic91]. Clean [Kah89c]. Clearer [Woo67]. Clenshaw [ES88, Oli77]. Clique [PPY93]. clone [PH85]. Close [vdSvdV87, Wil68a]. Closed [Ste71b]. closeness [Ruh75]. Closest [Cau64, Kel75, Ruh87]. Cobuild [Col87]. Code [Buc94a, DR93, DS93b, DR95, Whi87, BGD^{+}89, Bur90, Duf89a]. Codebreakers [HS93a]. Coder [Ben85, Ben88]. Codes [BGS87, BG90, Hig88c]. codiagonal [Wil58]. Coefficient [BR77, Doy57, Grc89, Fre90]. Coefficients [Coo67, Cro41, Gen69, Hop82, OP64, OPW65, Oet65, Uhl92a, Fre79, Kog55]. Coincidence [Syl51a]. colleagues [Har84a]. Collected [Ano04, ACD^{+}93, Bri92, Inc92, Tau63]. Collection [Abb78, Bai94a, Dic92, DGL92, GK69, Hig89b, Hig91b, MHR80, UvG89, Wri93, Mor90]. College [Ran88]. Collegiate [Mer83, Mer93]. Collinearity [BKW80, Ste87a, VV87]. Collins [Col87, Har91a]. Collocation [CH94, DS95, Fun88]. colloquium [Gla82]. Colon [Dil81]. Coloring [Kau78]. Colossal [Edm85]. Columbia [Edm85, Gau94]. Column [tBK84, Fos86, Ste84b, Ste90c]. Combination [Gau95, CM83, Cra86]. Combinatorial [BL86, HK65, Hal67a, Wil89]. Combinatorics [Bor94, Bru92, PW79a, WSW72]. Combining [Fis74, Ing95]. Comes [For68]. Commandant [Ano22]. Commandments [Hig92a]. Commemorative [Fox78]. Comment [Bjö78, EE84]. Commentary [Hig96b]. Comments [Bar79a, IEE81, LK76, Ode90, Wil71b, Hof85a]. commercial [GAG88]. committee [JBNP91].

Commodore [PW85]. Common [HT87, Van87d, Wol85, BKM86]. Commonly [Ker69]. Communicable [Kel85]. Communicate [QCM94]. Communicating [Boo84, Boo93, HH93, Rat85, SG91, Tur85, Mer90]. Communication [Bru82, DEK95, Lar83, Mat90d, She93, Ste90b, TK89, Don88b, Kel88]. Commuting [Kah70, Ber71a, PS79, Ros69]. Compact [Nas79, Nas90a, Oxf89a, Pug92]. Companion [Ber90, BF87, CM87, EM94, EM95b, GMS93, KL88, Kit95, LR93, LaQ94, NM90, QCM94, TT94]. Comparative [Ton92]. Comparing [FG84, Inv93]. Comparison [BG76, Box65, Box66, Bro91c, Gre72, HP84, MBI77, Neu86, SS62, Ste78, BJW85, EH79]. Comparisons [Hey87, Pap92]. compartmental [All83]. Compatibility [Kov76, OP64, RG67]. Compatible [Kov77, PH86]. Compatibles [SoIPUG86, SoIPUG87, Sta88]. Compendium [Ash86b]. Compensated [Kah93a]. Compiler [Far88b, PW77]. Compilers [Bro79, Pre93, Bur87]. Complement [Cot74, GMSW90]. Complementation [Mey89]. Complete [Ano85a, Ano85b, Ros90, BZY93, Bra83, Ede92a, EM95a, Gou91a, Gow73, Gow86, NM90, Ste81a, Tho42, Vav89, Ano78b, Gay91, Kub62]. completely [Fan51, Kah77]. Complètement [GK37]. Completeness [GJ79]. Completing [DM86]. Completion [EHM94, Mat69]. Complex [BBP87, BMPW66, BGG88, CH87, Cod93, CT65, Cro41, DLR85, Duf81, Ehr70, Fam88, For92, Fra58, Hig88c, Hig92f, Hir65, HFT94, Ips94a, Joh78, Kah87a, KT91, Kah92c, Ker89, Lew92, Moh85, Olv83a, Olv86, PJS92b, Ste85b, TSY88, TST86, Zie79, AL73, CL74, Ebe71b, FH60, Fre90, KR80a, KR80b, MN79, Uhl85, Smi62]. Complexity [Boj84, BM75, BB87, Bre76, CL91, CW82, Knu84a, Kro87, Mil75b, Pan92, Par92c, Tra73, TW92a, Tsa81a, BS83, Gen78, MS86, Vav91]. Component [Bjö91a, Grc92]. Component-Wise [Bjö91a, Grc92]. Components [CI95b, Lar66]. Componentwise [Bjö88, Dem92c, GK93b, HH92b, HK93a, Hig94d, RK95, Smo95, Sun92a, Zha93a]. Comportement [Bru89]. composed [Ano78b]. Composite [Cad88]. Composition [Rub88, BKW90]. Compositors [CBB54, Har83]. Compression [OP83, BLR90]. Computable [AL75, CG67, Ste83a, Tur36]. Computation [AS92, Aik85, Bai88b, BKP93, Bar93b, BL95, BT89, BRR85, BRK^{+}87, Boh77, BL91, BB84, BLV85, BHM94, Cha84, CF93, CFB93, CS84, CH90c, DST88, DLR85, DH95, DL92b, Dub83, DG94, EAS94, EK92, ET73, Far88b, Fer94, FF87b, For70, Fox71, FK77, GL95b, Giv54, Giv58, Gla93, GW76, Gro89, GDD94, Hof89a, HS93b, Jen77b, JM92, Kah90b, KL80, Fer95, KBS94, KM83, LHPW87, Leh94b, Lem75b, Lui96, Mai84, MT94, MT95, MNP84, MA87, Mod88, Ng84, Nod71, Par74, PA88, Ral65, Ral91, SP88, Ske89, SA56, Ste74a, Sun92d, Thi88, Var90d, Wil60b, Wun48, Boj84, BGBMN90, FT88, Gau94, ITH88, Kåg77b, KR80a, KR80b, WG78]. Computational [AG90, Bau74, BRB76, BD73, BM75, BB87, Bur74, BL86, CF92, CKP95, DD86, DSH95, DC77, FF63, Fad59, For67, Gau67, GHMP74, GW73, GK69, Gus70, Hil88, Kah90c, Ker90, Van92, MS74b, Mil75b, PW79a, Tod61, Won90, CDM79, Gau94, JBNP91, VV91]. Computations [AH83, Ber92, BP94b, BR76, CCF96, Cha92, CV88, CH80, CW85, DB76, DHS89, DW95, Duf84a, Dwy51, Dwy53, EKNX93, Far88a, FMM77, GPS90, GHN^{+}90, Gen73, GV83, GV89, GS92, HHK93, Hem67, Hig86c, Hig89c, Hig90i, Hig93c, LF80, Lau85, Lin75, Lin81, Van90, Lon84, Mat93a, Mat95a, MT86, Mol2a, Mol90b, New62, Pan92, Pri92, Ric81, Ste73e, Van83b, Wat91b, YS89, Bis90a, Dat89, Gen78, GS74, Van89, Pai79, Pan90, Ple86, Ste90b, Tsa74a]. Compute [Bar91a, BBB89, Fra89, MV78, Dav83, GM86]. Computed [Dei90, Don82, Don83, GL95a, PW74, Wil61b, DMW83, ST91]. Computer [AwJS85, AHU74, AM65, Bar86c, BN71, Boaxx, Bom93, BB92, BM92, Bro91b, Cip88, DST88, DM86, Eva81, FvDFH90, For57a, FM67, For68, For69a, FMM77, Fra92b, Frö90, Gar76b, GL81a, GHR67, Gol91a, Gol72, GJ65, HVZ90, HWH91, Har81b, Hem67, HP90, Hof87, Kah88, Kah94a, Knu81a, Knuta, Knu81b, KM81, KM86b, Lon67, MMR78b, MS91, Mic80, Mil87a, Mil75c, Mit87, Nei82, PW78, Par90a, Par93, PH94, Pet80a, Pet80b, Roj94, Rub88, Sal89, SB77, Sch81a, Sch88b, Spe91, Ste80a, Sto75b, Sto77, Sto91, Swa80b, Wan91, War60, Web92b, WWG51, Wil75a, WCK89, Wol91, Ber80, Knu62, Kol78, Maz76, Vrs86, CGS83, Knu86d, MFGK82]. Computer-Drawn [Cip88]. Computerized [Kah90d]. Computerizing [Wei85]. Computerland [Sch74]. Computerorientierte [KS88]. Computers [AS85, Bai91, Bai93c, BMSV92, BB49, Bul91, CKW85, CDPW92, DDC^{+}87, DKH86, DDSv91, DV92b, DW93, Don95, DW95, DM71, GJ79, Has55, HJ81, HJ88, Hom92, IT89, KR65, Lav80, Nas79, Nas84, Nas90a, OV85, Pic90, Qui87, RW67, Ran75a, Ran75b, Ric55, Sas93, Vit68, War76, Wil74b, Wil75d, Woo58, ABB^{+}90, ABD^{+}90, BD89b, Bis90a, DS86a, DS86b, DS88, Gay91, HK65, Olv83b, DF87, Hoe87]. Computes [LSV95]. Computing [AG89, AK75, AL86, Asp90, Ata40, BD91, BDM93, BD93a, Bai93e, BD90, BV86, Bis89b, BB71, BG73, Bjö90a, BH95a, Brü95b, BC86, CL79, CGL83, CH90a, CI95a, CF93, CFB93, CH90b, CCC82, Cro90, CV86, DK87, DG93, Día89, DS95, DDGM89, DT93, DT94, Dow93, Ede93b, EH91a, ES89, FTxx, FGM95, FM68b, Fra58, Fre87a, Fre87b, Fun88, GH93, Gen69, Gil51, GLO81, GO96, GL80, Gro89, HHKR93, HW95, HLPW86, Hey87, Hig86a, Hig86b, Hig87a, HS87, Hig88a, Hig90c, HP94a, HP94b, HW79b, HRV92, HCSW88, Inc85, JS92, Jun95, KP94, KG80, KLP93, KKL^{+}93, Lab57, LZ91, Van78, Van85b, Van86a, Luk80, Luk86, MB93, MS93b, MHR80, Mio90, Moo75, Moo88, MS83, Nas90b, OS90, Pai86, PKPL91, PA88, PFTV86, PFTV88, PTVF92a, PTVF92b, RK95, Saa80, Sch83c]. Computing [Sch81c, SB80a, SA73, Sha87, SK93, Ste83b, Ste85a, Swa93, Tao84a, Tao84b, Thi88, TW91b, TSY88, TST86, Tur45, Tur50, Tur91a, VS88a, War77, Wat90, Wat92d, Wat96, Wil48a, Wil48b, Wil55a, Wou87, BBK89, CM86, DK81, DHW92, Fil81, FH60, Fre79, FP86b, FP86a, GU77, Gow85, HP84, HW78b, Pai81, PV86, RS88a, SS74, Tsi90]. comrade [Bar75b]. Concavity [Mat90a]. concept [Hof85a]. Concepts [AS83, Kah90c, Ste75d]. Concerning [Har80, Tau49, Tod61]. Concise [Mer91, Cha91, Lon93, Oxf90, Woo62]. Concrete [GKP89, Knu89b]. Concurrent [CDPW92, AS83]. Condensed [DHS89, DvdG92]. Condition [And91a, AR85, BDM93, Bau63a, Bis90b, BLP90, dBK86, Bro73, Bye84, Bye85, CR86, CF92, CH94, CMSW79, CCV82, CR83, Dem83, Dem87c, Dix83, Ede88, Ede89, Ede92b, EHM94, Erh93, FL74, FGP91, FM52, Gau69, Gau78b, Gau83a, GI88, Geu82, GK92, GK93a, GK93b, GL81b, GEJ95, Hag84, Hig83a, Hig86b, Hig87c, Hig88c, HH92a, Hig95a, KW89, KP94, Kah72a, Kat60, KL89b, KL91a, KL94, Kur92, Lem75a, Van87c, Loi69, Mat95c, O’L80, PP92, Ric66b, Roh89a, SB92, vdS69, vdS70a, SW91, SW94, Ste80b, SZ91a, Tan94, Tau50, Tod49, Tod50, Tod54a, Tod54b, Tod68, Wed85, WWWS86, Zie88, ZWS86, Bjö88, DCHS88, FR87, Han90a, Hig83b, Van82a, MZB87, Ste90c]. Condition-Number [Bro73]. Conditionally [Bax91]. Conditioned [CF88, FGS95, FS55, Gau75b, GW76, Men56, Uhl92a, PW79b, Rum91, Ste81b, Wil67b, Wil72c]. conditionées [FKF68b]. Conditioning [CGR90, AR92, Dor71, Fle85a, Fle94, MS74a, Rhe66, Ric65, Sha82a, BDM89, Cyb86, Dem90b, Osb60]. Conditions [BG77, Eva71, Sha82c]. conducted [MMR78b]. Conduction [Gil86, HLW83]. Conference [AK90, BCEP94, GW90, GW92a, GW94, HW76c, HPP88, Ack89]. Conferences [WCK89]. Confession [McC78]. Confessions [Ben88, Ben85]. Configuration [SW91]. Confirming [Kah91c]. Conflicts [Gea77]. Confluence [Kah72a]. Confluent [BE73a, Dem90a, GP71, Gau62, Gau63, Gau78a, HG80, Hig90k, KF63a, Lu94, Lu95, Lu96]. conformations [GHR93]. Conformity [Wic93]. Confusion [BM92]. Congress [AAR87, Jam74, Pet68]. Congruence [Cli66]. conical [All88]. Conjecture [Ede92a, EM95a, PC83]. Conjectures [Dem87b, Var90d]. Conjugate [AK90, CGO76, CGM85, EAS94, Fre90, GO89, Gre89, GS92, Gre94, Mag60, Not93, PS92, Pow87b, She94, Ste73c, Str91, Woź80, CS89a, CM86, Ste75a]. Conjugate-Gradient [Gre89, Woź80]. Conley [Ros90]. Connected [BLV85]. Connection [Ant85, Har95, Wed72, Gab87, Vas90, BHH93, Sch90a]. Connections [LM90]. Connoisseur [Dic92]. Conquer [AA94, BDG94a, BS92b, GE95a, GE95b, Jes93, KM86a, LL93, GR90, Sor90, ST91]. Consensus [Lef78]. Conserving [Kah72a]. Considerations [AGG88, GMSW86a, DHW92]. Consistent [NP78, Cad73, Cad74, Fle84]. Constant [Bai88a, BR77, BR71b]. Constants [Bai88a, Bri91]. Constellations [Orr74]. Constrained [AE93, Aru92, BH88, Bar88a, BMO92, BV92a, Box65, Box66, Fle81b, FGS95, GL94, GM74b, GM74c, GMSW87a, GMSW88, Gul92, GW92b, Gul94, Gul95, Van85c, Ste81b, Swa74, Wed85, Wri92, You79, Dem87a, GGM89, Gay84, GHM^{+}86, Van83a, Woo89]. Constraints [Eld80, Löt83, Fle85b]. Constructing [Chu95, Pan78]. Construction [BP67, BB49, Bue77, Wil80, LN62]. Containment [SSSU87]. Contemporary [Gal94, Lon95, Sum87, AAR90]. contents [WY63, BDDC88]. contiguous [DGK80]. Continent [Bry89]. Continuant [Rut47, Rut52]. Continuation [BH95a, LL93, Lui96]. Continued [Ant85, Bla64, Cor92a, JT74, Kah87d, KBS94, STY83]. Continuity [Dav85]. Continuous [JW85a, Chu91a, Fan51, GS71]. Contrôle [PV74a]. Contracted [Kah91b]. Contribution [Ano78a, Bru89, Geu82, Par90c]. Contributions [Aga93, Caj25, Gar76b, Ips94a, Ng84, SG83c, Bel31]. Control [Bar75a, Bol85, BGFB94, CGR93, CC91, DD86, DK88, Enr89, Gar90a, Grä94, Gus70, Ham82c, Hod92, Hul82, Lau85, Van82b, SR80, SW82, DK83, Dat89]. Controllability [HG80, KL88, Pai81]. Controlled [Ano46, Cha66, CHH83, Cor92b, Bis91]. Controlled-Precision [CHH83]. Controlling [GMSW91]. Convection [RTP93, RT94]. Convection-Diffusion [RT94]. Converge [DL92a]. Convergence [AN63, AR85, AR92, Art86, BIM76, BI77, Boaxx, Dax90a, Dem77, Deu83, JR75, Lei89, Lyn69, NR72, Not93, O’C73, SS89, SB88, vdSvdV87, Ste94d, Str91, Tao84a, Tor91, Tor93, vdV90, AL73, Bal80, Cli72, Fis83, Gay79, GO88, Mas91, PV86, Ruh68, Ste75a]. Convergent [Bai88b, Fan93, FP63, Le85b, MP77, NP78, CL92, Li91, Ove83].

Converging [vdV92]. Conversations [AAR90]. Conversion [Buc94b, DK94, Mat70, Pax91, TL81]. Conversions [Gay90, Mat68]. converter [FGMS90]. Converting [Ziv95]. Convex [BD85, GHM^{+}86]. Convolution [GF74, MZB87]. Cookbook [Bak85a, Tea91]. Cookery [Ark86]. Cooking [Dav88]. coordinate [JS77]. coordinates [Lef78]. Coordinex [PC83]. Coping [Tru85]. Copy [But81, But92a, Swa79]. Copy-Editing [But81, But92a]. Copyedit [O’C86]. Copyediting [Jud90, Sta91]. core [Gay91]. Cornelius [BCEP94]. Corner [Mol90a]. Correct [Sch84]. Correcting [Kah91c, Pet80a]. Correction [Bar88a, BV92a, Car58, GHR67, Pet80b, Pic72, DLS83, RS81]. corrections [Kie73]. Correctly [Gay90, Ziv91]. Correctness [Dek82, DDR94, Hul79]. Correlate [Dil81]. Correlation [MO84, SBKZ93]. Correlations [GZ93]. Correspondence [Wil71c, Wil71d]. Corresponding [GLO81]. Cosine [SB80a, BIF69]. Cost [Gli89, Pri92, PW77]. Couette [TTR92]. count [KP77]. COUNTDOWN [NM90]. Counter [CR83]. Counter-Examples [CR83]. Counterexample [Dem87b]. Counterexamples [KP74]. coupled [Van88]. Courant [Rei86, IIM87]. Course [Bar82, Bis88b, CJLP93, DH95, EL69, Har67, Mon89, Ort72, Ort87, RR78, SR90]. Course-grained [Bis88b]. Courses [ILA93]. Covariance [Pai85a]. Covariances [Bur74]. Craft [Lan89a]. CRAIG [Sau94]. Cramer [Mol74]. Crash [Mon89]. CRAY [DD89, DE84, Bai88c, Cal86, DH86, Kah90b, Pap92]. CRAY-2 [DD89, Bai88c, Cal86]. Crazy [Led90]. CRD [Bar81b]. Crest [Jos91]. Crimes [Sto91]. Criteria [ADR92, BB94, JR75, Lyn69, Pow67]. Criterion [Ada67, Bra95, Kar74, NR72]. Critical [Wic92]. critique [Lin88].

Critiquing [Kuy91]. Crooked [Rol75d]. Cross [GA94]. Crout [For60, McK62, Tha61]. Cruel [Hau91]. Crummy [TS80]. Cryptograms [MM83b]. Cryptography [Hil29, Lev58]. CS [Van84b, Van85b, PW94]. Cubic [FK77, Kah86d, Mag60, McK84, Nam85, Squ87]. Curbs [Kah72a]. Curious [Fey85, III84]. Current [Box66, Oxf89b, Oxf90, Rat85, Und80, Woo62, WFF89]. Curriculum [CJLP93]. Curvature [EAS94, MS79b]. Curve [Brü95b, Smi65, Fri85]. Curve-Fitting [Smi65]. Cuts [Kah87a]. cweb [Thi86]. Cyclic [AM94, BG94, EHHR88, LJ93, SS89, Tem75, Zaw91, FH60].

D [BDDC88, Hoe87, AHB87]. D. [Kah73b]. d’Algorithmes [PV74a]. d’Analyse [Cau41]. Danilewski [Fra58]. Dans [PV74a]. Data [Ack89, AS85, BR93, BKW80, Ben87a, Bev69, CCKT83, CR79, DW80, Erh93, For57a, HK87, Hen92, KLV91, LS82, LS83, MS76, Osb85, RG67, Roh90, RK95, Hil89]. Data-Fitting [For57a]. Database [War89]. David [IEE81]. Davidon [Ste67]. Davis [CR94]. day [Gla82]. Days [Ros90, Pre91]. Dazzle [Anh94]. Dealing [Kol78]. Debate [Roj94]. Decade [Mos71b]. Decay [DMS84]. Decays [A^{+}90]. December [BCEP94, Syl51a]. Decimal [AS85, Bai88b, Buc59, CHH83, Gay90, Kah86a, Pax91]. Decimal-Binary [Gay90]. Decimals [Leh43]. Deciphered [Rej81]. Decomposing [MS78c]. Decomposition [BD93a, Bar86a, Bar90b, Bre70b, BLV85, BKP76, Bun82, CH87, DDLT91, DK93a, DK93b, Dub94, EP94b, Fos86, FP81, Gan90, Gen74, GDD94, GE95c, GW92b, HN81, Hig86a, Hig90a, HS90, HP93, Hig94c, HP94a, HP94b, HP91, JS92, Kar74, KL91a, KL92, KL80, Kni95, Liu87, Van82b, Van84b, Van86a, MS93a, Mat93b, Mat95c, PS81, PW94, PS94, PE95, Pow86b, Rei91, Ste83b, Ste85a, Ste93b, Ste93d, TG81, Uhl81, Wed72, Wil78b, Zha93a, AGS91, AG88, BV86, Bis89b, BGBMN90, CH90a, Chu86, DGK80, DLR81, GR70, Ham85, Han90c, Hau67, HLPW86, Kåg84, Van76a, Luk80, Man82, Mao86, MG91, MZ91, PV81, Pai85b, Pai86, PV86, Ste90f, Wri90]. Decompositions [AMS74, BZY93, Bar91b, BGG88, Cai84, CI94, FNP82, Gen75, Gol69, Van85b, Mat84a, Mat84b, Ste94b, Sun92a, AP76, Chu89, Har86a, MG89, Pai80, Ste84a, Wed78]. decompositon [Tog86]. Deconvolution [Lin92b]. Defect [CC91, Cor92b]. Defect-Controlled [Cor92b]. Defective [Dem90b, SB88]. Defense [Smi81, Uni92]. Deferred [Bar88a, BV92a, vdG93]. deficiency [Wat88b]. deficient [DS80, Wed73a]. Defined [GL80, Kah87c, OPW65, Pow67]. défini [Ben24]. Defining [Hop82, SW91]. Definite [AR85, Bak84, BR71a, CV86, Doy57, EHM94, Fra89, GL81a, GSS96, GV79, Ham82b, HW79b, MPW66, Mat92b, SW81, SS91, Van95, Bel68, CM83, Cra86, DI86, Fle85b, FG86, Hig90a, KW83, KM76, MM88, OSV79, Pul66, Ste81b, Tau67, Tau70]. Definiteness [Ker90]. Definiteness/Semidefiniteness [Ker90]. Definition [dJ77, Tre92a]. definitions [Rin55]. Deflation [LS94, Ste81d]. Degeneracy [Fle88, Fle93, Fle94, GKS76, Ste84c]. Degree [Tal91]. Deivations [CL79]. Delight [Saw43]. Demonstration [Mol82a]. Denormalized [Coo81]. Denoted [Ebe71a]. Dense [CDW94, DGK84, DH90, DvdGW92, Ede91b, Ede93b, Ede94a, GPS90, LJ93, PR89, DH86, Pan90]. Deoxyribose [WC53]. Department [MMR78b]. Departure [Lee95]. d’équations [Ben24]. Derivation [FS83, Nam85, KT85]. Derivative [Bat83, DS95, Le85a, Fun88, Mat91b, Rin56]. Derivative-Free [Le85a]. Derivatives [Bre73a, Fle65, Kub91, Pow75, Smi65, Woź74, BS83, Dwy67, ITH88, Ste67]. Derive [Bar92a]. Derived [STY83]. Deriving [HS81a, Vas90]. Descent [Bol84, FP63]. describe [Kah90a]. Describing [Bro68, Mac90]. description [Uhl85]. Design [AHU74, BD92, BD93b, BD94, CKW85, CDO^{+}94, CDW94, DW93, DPW94, GJMS88, Koz92, LPBW87, Lau85, Mil87a, Nas85, PH94, PW85, Pet80b, Pla93, Rub88, BCJR87, DK84, DCHS88, DS88, DBKF90, GMS92]. Designers [WF82]. Designing [DDK92, Har85, Qui87]. Designs [Gla95]. Desktop [Har86b, Mil87a, Bud87]. Destabilizing [MH87]. Detecting [Fos79a, Pet80a]. Detection [BV92b, Par92a, PV74b]. Determinant [Chi93, Fra58, HG80, Hel74, LZ92]. Determinants [Caf63, Cro41, FH93, Mak91, Rut47, Rut52, Syl51b, Tur29, VG87, Bel31]. Determination [Gou86a, Joh78, Man81, Sch84, Pul66, Ruh70]. Determine [Cod88a, GT93]. determined [Han90c]. Determining [SW94]. DEUCE [Wil80]. Develop [Lor74]. Developing [FS76, BD89b]. Development [Cow84, DDC^{+}87, GS90a, Har81b, McI82, Ost83, Ste77c, Swa88, Tur45, Gow85]. Developments [Gil74, GMSW88, Har84b, KF87, Wil75a]. Deviating [Mee80]. Devil [Bie11]. Devilish [Ede95]. DF [SF86]. DFP [Pow86a]. DFT [BH95b]. Dhahran [Uni92]. Diagnostic [Gay88]. Diagnostics [BKW80, KC82]. Diagonal [BD93c, BG90, Bun71a, Bus69b, Chu95, Cle59, FP62, Fos90, Hig95f, Kah66a, Mir59, Sha85, Sha91, SW62, Str72, Var76a, Var76b, FG86, Wil60a, Wil68a]. Diagonale [Gab79]. Diagonalizable [LTT92]. Diagonalization [BGBM93, GH59, Ebe71b, Kog55, LHPW87]. Diagonalize [Dem83]. diagonalizes [AL73]. Diagonally [BD90, FV62, FNP82, BS74]. Diagonals [Lás91, MRK76]. Dice [Ste89].

Dick [Hym86]. Dictionaries [Bai87, Lan89a, Ors69, SE62, Sta80]. Dictionary [Bie11, Bor94, Bry87, Cha90, Cha91, Cha93a, Col87, DN89, Eis92, FF92, Fow68, Har91a, HT87, Hig94e, Hou82, 87, Lee81, Lon91, Lon95, Mer83, Mer93, Neu88, Nix81, Par92d, Pét66, Oxf89a, Oxf89b, Oxf89c, Oxf90, Oxf93, Ran88, Ray91, III84, Sch87, SE62, SN46, Sum87, Tul91, Und80, War89, Woo62, Wei85]. Did [TE87]. Diet [Ros90]. Difference [Fer94, FW60, HW80, Hen63, Fer95, RM67, Var72, You54, Osb65, Ste67]. Differences [BR93, KF63a, KF63b, KF63c, MNP84]. Different [tBK84, FS76, New81, TL81, Boj84, Kah77]. Differentiable [Gou86a]. Differential [Ame77, AMR88, BP67, BD86, Bur95, But87, CGO76, CKP95, Cor92b, Ept80, FP92a, FW60, Gea81, Gil51, GO96, HW91a, HW76a, HW80, Hen62b, HW91b, Hul67, JS83, KH90a, Lam91, OV85, Pet82, SWD76, SW84, Sha94, Ste73b, Vit69, AP81, BIM90, BJW85, Chu86, Dah83, Lax65, Osb65, Pol85, Wil76, Wil77a]. Differential/algebraic [Pet82]. differentiate [GM86]. Differentiation [GP70, GC91, Iri91, LM66, Ral81, Gri88]. difficult [Dem88b]. Diffusion [RTP93, RT94]. Digest [NA xx]. Digit [Boy94, Gol67, Hil95, Rai76]. Digital [Ano48, BB49, CKW85, For57a, GHR67, Gil51, Has55, Hem67, Knu87a, KM86b, OP83, Ral65, RW67, Ran75a, Ran75b, Ric55, Rub88, Ste80a, Van91, WF82, WWG51, Wil74b, Woo58, AP76, MS79a]. Digits [Bai88b, Bar88b, BBB89, New81, Rai69b, Tsa74b, Tur82, Olv83b]. Dilation [SC70]. Dilations [DKW82]. Dimension [Gar64]. Dimensional [Hal74c, Hén76, LR81]. dimensionality [Rao80]. Dimensions [Mat85, KT76]. Dinner [Fox90b]. Dinosaur [Nor91]. Direct [ABS84, BD91, BH88, Ber91, BP71, BGN70, CG67, CH88b, DM86, DT91, DRMN79, DER86, DR93, DR95, Hey87, Hig93c, JO74, Kåg93, Osb65, Sch90b, Swa72, Swa74, Tor89, Wil61a, Dor70, HP84, SZ91b, Tor92]. Direction [AGS91, TTRD92, BIF69, Van87b, NS90, Ste73c]. Direction-of-arrival [AGS91, Van87b]. Directional [Tor89]. Directions [Duf84a, FGM95, Gou86a, MS79b, Pow87b]. DIS [LIA93]. Discourse [Dav72]. discovered [TE87]. Discovery [BGH62, Coo94, Cri89, Wat80b]. Discrete [ADW84, BP67, Bar88b, BJ89, BF93a, BH95b, GLKH90, GL95c, Han92, Hen62b, O’C73, Ple76, Wat77, Dor70, Fis83, Han90a, Han90b, Han90c, MS74b, RR73]. Discrete-Time [GLKH90, GL95c]. Discretization [Ste73b]. Discretized [dBK86]. discrimination [BMW86]. Discursive [Sch77]. Discussion [Ste87a, Eis92]. Disease [Kel85, Ken85]. disks [Tod65]. Displacement [KKM79, LS82]. Display [Tuf83]. displays [BK86a]. Dissertation [Ste81a]. Dissertations [Tur87]. Distance [Bax91, Bye88, Dem87c, Dem92c, EH91a, FTxx, Far87b, Mat85, Ziv82, GHR93]. Distributed [BT89, Bis88b, CDPW92, DvdG92, EHHR88, HR88, KP92, LC89, Bis89b]. Distributed-Memory [EHHR88, HR88, LC89]. Distribution [Bar81a, Bar88b, DG87, DRW93, Ede91a, Ede92b, Gay85, Par86, Rai69a, Rai69b, Tur82, Ura68, FM86, Mac88]. Distributions [BB85a, BG79, Hey87, Tsa74b, FP86b, FP86a, HP84, WY63]. Disturbances [Far85]. Divergence [Boaxx]. Diversions [Gar61]. Divide [AA94, BDG94a, BS92b, GE95a, GE95b, Jes93, KM86a, GR90, Sor90, ST91, LL93]. Divide-and-Conquer [BS92b, GE95a, GE95b, KM86a, ST91]. Divided [KF63a, KF63b, KF63c, MNP84]. Division [Smi62, Ste71a, Ste85b, Tsa83, Tur45]. Division-Free [Tsa83]. divisors [Wil84b]. DNA [Wat80b]. Do [Asl82, For68, Fox71, Gea81, GT93, GA94, Kah81, Loc85, Par94, Pow75, Hol75, KS92, Cow89a]. Doctoral [Ste81a]. Document [Dil93, KD93, KD95, Lam86b, Lam87a, Lam94, van86b]. Does [Cip89, Hon91, Ste89]. Doing [Web92b, Wol91]. Domain [Gli89]. Dominance [SSSU87, Var76a, Var76b]. Dominant [BD90, FV62, FNP82, BS74]. Donald [Alb91, Fre87b, Hof85a]. Done [Kah90b]. Donkin [Syl51a]. Dont [Ber77a]. DON’Ts [Jac76]. Doolittle [Far87a]. Doomed [Gor84, Gor93b]. Double [Cri74, KLV91, Møl65a, Møl65b, Wat80b, Mol90b]. Double-Precision [Møl65a, Møl65b]. Double-rounding [Mol90b]. Doubled [Kah87b, Lin81, SS84]. Doubled-Mantissa [SS84]. Doubled-Precision [Kah87b, Lin81]. Doubly [PL82]. Down [Hal90b]. Downdates [Ste94e]. Downdating [BPE92, EP94b, GE95c, PE95, BPT90, BBVdH87, Ste79a]. Dr. [Ano78a]. Draft [LIA93, IEE81]. Drawing [Gur93]. Drawn [Cip88]. Drift [RK75]. Driven [Tan89, Tan90b, Tan92]. Drum [GWW92, Kac66]. Dual [BIMF76, KLJ89]. Dubious [MV78]. d’un [Ben24]. Dundee [Gri84, GW86, GW90, GW92a, GW94]. d’Une [PV74b, Bar79b, Pic72]. Durbin [Cyb79, Cyb80]. durch [Spi76]. Dust [Asl82]. Dynamic [AP94, Cod73, Tan94]. Dynamical [Ch85, HW91b]. Dynamically [Cod88a, HH88]. Dynamics [BBP87, CHQZ88, Or94, Har65].

E-Mail [QCM94]. E. [Rei93]. E2 [GS71]. Eager [Gor84, Gor93b]. Early [Cro90, Fox90a, HGD90, Lav80, McC84, Sch81c, Ste93b, Wil75a, Wil75d, WCK89, CK80a, CK80b, CK81]. Earth [QTD91]. Easy [Dav88]. Eckart [GHS87]. Eclectic [Mil87b]. Edge [Fle94]. Editing [BSB92, But81, But92a, Eis92, Lev83, O’C78, Swa79, Tar92, Har84a]. Edition [Hig90i, Tho42, Web92b, Hig95d]. Editor [Ano91, Pas85, Rol75e]. Editorial [EE84, Rod92, Rol76d, Swa79]. Editors [But92a, CBB54, Gar90a, Lon93, MS89, Sta77b, Tar92, HP86b]. eds [Hig90j]. EDSAC [CK80a, WWG51, Wil75a]. educational [Fle81a]. Effect [Att57, DM90, Lyn69, Ske81]. Effective [CH94, Mat90d, Nas84, Tur85, TK89, Tur74, WPG92]. Effectively [CF88, QCM94]. Effectiveness [Gre79]. Effects [HW75, Kar91, LS78, Ste79a]. Efficiency [Kro87, PH72, Sma85]. Efficient [Ben82, BMSV92, Fam88, FP74a, GB86, GDD94, GE94, Hig86b, HS93b, LS78, LPBW87, Le85a, Nix81, PR85, PR89, Qui87, SV89, Ste80b, Tan90a, TW92b, VA86, App85, Cad74, Enr79, MS79a]. Egg [Ark86]. Egyptian [vdW75]. Eigendecompositions [DK87]. Eigenelements [Saa80]. Eigenproblem [AL84, BDM93, DM86, GE94, GE95b, Kah93b, Kha72, Lau88b, LZ94, BDM89, GD89, Wil72c]. Eigenproblems [BDG94a, BG77, BCD95, GW73, GE93, LL93, PR81, GR90]. Eigenreduction [Sla92]. Eigenroutine [BD92, BD93b, BD94]. eigenroutines [DCHS88]. Eigensolver [LTT92]. Eigensolvers [BB94]. Eigenspaces [KP94]. eigenstructure [Cyb84b]. Eigensystem [GBDM77, Mat93a, Mat95a, SBD^{+}76, WG78]. Eigensystems [GW76, KPJ82, PW74, Wil61b]. Eigensytems [BD90]. Eigenvalue [AA94, AGG88, Bai94a, Bar93a, Bar93c, BS89a, CI94, CB90a, CW85, CW86, CV86, DKH86, DHS89, DS92, DvdG92, DS93a, Ede91a, Ede93a, EAS94, EH91b, Fan93, Gea69, HV94, HS93b, Ips94a, Jen83, Jes93, Jun95, Ker89, KW92, LZS91, LZ92, LZC92, LPS87, Mac95, Nod71, Ort63, Par64, Par65, Par80, PSS82, Roe86, Rut58, Saa84, Saa92, Sch86c, Sco81, SHW86, Ste73d, SZ91a, Sun95a, Van81, WE91, Wat93a, Wil62, Wil65a, Wil70, Wil87a, Yau93, BG87, DM84, DS87, FNO87, GGM89, Gol73a, Kub62, OSV79, Osb78, Paa71, Ruh73, Sor90, SE87, Wat87, Wil77a]. Eigenvalues [And90b, BD91, BD73, BS92b, Cha93b, Cho69, Chu95, CJ70, Dat87, Dei90, DG93, Dix83, Don82, Ede88, Ede89, EKS94, EM94, EM95b, Fai71, Fra58, HS59, How83b, Hym57, Kåg93, KP94, Kah66a, KM86a, LTT92, LSV95, LZ91, Van87c, OS90, Par84, PW69, RT92b, Ruh94, vdS75a, vdSvdV87, SZ91a, Sti91, Swa93, TM62, TTRD92, TTRD93, Wil86c, DM78, DK81, DMW83, Fan51, FHM67, GU77, Hor54, Van82a, Van84a, Mat90b, PR69a, Ruh75, Sch83c, Ste91e, Tod65, Wil68a]. Eigenvector [EH92, Cyb86]. Eigenvectors [BD73, CJ70, DK70, DR94, Fai71, Hym57, LSV95, Van87c, Mor80, OS90, Ber71a, BS74, DMW83, PR69a, ST91, Ste75c, Wil58]. Eigenwerte [Rut66]. Eighty [Har81a]. Einführung [KS88]. einiger [Neu74]. Einstein [Reg87]. EISPACK [DM84, GBDM77, SBD^{+}76]. Either [Tor92]. Electromagnetics [Wan91]. Electronic [Ano93, CKW85, DG87, Gay85, Lon67, Mad87, Ste80a, WWG51, Wil55a]. Element [Gil86, HLW83, LNOP82]. Element-by-element [HLW83]. Elementary [Ant94, AR94, Atk93, BB84, BD86, Bre76, Bur80, CJLP92, CW80, Cod93, CdB80, Eve66, GB91, Hig95d, HFT94, Kah87a, Leh94b, Men56, SK93, Stu80, Tan91, Wil59, Ziv91, Wil84b]. Elements [ABG48, CF92, FP62, Fos90, Hen64, Kat85, KP78, Mir59, Moo75, Ode90, Par76, Ste74b, SW79, Thi88, LR81, CF93, CFB93]. Elimination [Ban76, BB85b, Boh75, Bro73, CG85, Cli73, Coh74a, Cry68, CR72, DK77, Dax82, DP88, DEGR88, Ede92a, Fos94, FP81, GP92, GP93, GKO94, Gou91a, HH89, Hig90b, Hig90f, Mac89b, Mil71b, ON93, OW82, Pap92, PW75, Ple74, Rei71, Rob90, Sau78, Ske79, Ske80, Sor85, Ste74b, Ste93a, Ste94a, Str69, Stu82, Stu85b, Stu85c, Stu85a, Syl40, Tre85, TS90, Vav89, van77d, Wri93, Yal91, YC95, Bus71, Geo72, GM77, GP90, GN82, Mac88, Ker92, Läu61]. Elliptic [BR77, BL84, CGO76, HW80, RB85, Sch81b, You54, FHM67]. ELLPACK [RB85]. Eloquence [DC87]. ELR [DK81]. Emacs [CR91, Sta86, HH88, Sta86]. Embrace [dP80]. Emergence [Dil81]. Emitter [Sch86b]. emphasis [All83]. Emphasizing [Fra92a]. Empirical [Dil81, Knu71, Lan38]. Enchirdia [Pas75]. Enclosing [Roh93]. Encyclopedia [Cry87]. endlicher [Neu74]. Enduring [Anoxxb]. Engine [Bab37, Tur45, Wil48a, Wil48b]. Engineering [BRK^{+}87, Bro75, IBM88, MS88, MD64, Mic90, Mil84, Moo36, Pit90, Rat85, Hil89]. Engineering/Scientific [BRK^{+}87]. Engineers [Bar78, FM68b, Ham73, Jen77b, Ken82]. English [Bor94, BR67, BR68, Bru82, Wei85, Bai87, Ber65, Ber71b, Ber77a, Bry90, Bur89, Cha90, Col87, Cry88, Day92, EL69, FF31, Fow68, GH43, GH70, Har91a, Her65, 87, Led87, Led90, Lon91, Lon95, Mac90, Neu88, OW77, Ock75, Ors69, Par73, Oxf89a, Oxf89b, Oxf89c, Oxf90, Oxf93, Qui68, QS90, Rob88, SN46, Sum87, Swa71, Swa88, Swa80a, Trz93, Und80, ‘Vi52, Wat91a, Woo62, WFF89]. Enhancement [Cad88]. Enhances [Hig90h, Hig91c]. Enhancing [KLW89]. Enigma [CK84, Har95, Hod83, Hof85c, Rej81]. enormous [Kol78]. Enough [For70, Gol67]. Entertains [Bei66]. Entries [Chu95]. Entrywise [JE95, O’C93]. Entscheidungsproblem [Tur36]. Environment [For78, HH88, Ost83, DD90]. Environments [DT93, DT94, KC82, MS93b, Wou87]. Envisioning [Tuf90]. Epidemiology [Ken85]. Epilogue [Fox90b]. Episodes [Swa88]. Eponymy [HR80, Smi81]. Equal [Lew92]. equalities [Ste90d]. Equality [BH88, Bar88a, BV92a, Eld80, GL94, Van85c, Van83a]. Equality-Constrained [Van85c]. equally [TW91a]. Equation [AGL95, BK79, BK80, Bak84, BS68, BS72, BD74a, BR96, Buz92, Bye84, Bye85, Bye87, Dat88, DSH95, Dou59, Duf81, EU92, For69a, For69b, FSC73, GLKH90, GL88, GWL^{+}92, GLAM92, Ham82a, Ham82b, HK88, Hod92, HMW77c, HR92, Inc83, Inc84, Jon84, KW89, KP92, KP93, Kåg94, Kah79, Kah86d, Kre72, LR83, Low60, MT94, Mil88b, Mol2b, MD92, Nam85, Rob80, Ste92b, Syl84, Wac88, Ziȩ85, All88, BIM90, BJW85, Cal86, Chu86, DD88, Dav81, Dav83, Don87, Dor70, Laa58, SH83, Smi68b, Ziȩ83, Ziȩ87, Bar79b]. Equations [AG90, Ame77, AL84, AMR88, Ata40, Auc92, BM82, BSD60, BG76, BD74b, BP70, Bjö87, Bol84, BBB89, BMPW66, BD86, BP71, Bun85, Bur95, But87, BGN70, CI95b, CV89, CG67, Chu87, CGO76, CKP95, Cor92b, Cro41, CNRT81, DW80, Dax90a, DS89, DK85, DS83, DV92b, Don95, DRMN79, Duf81, DR82, DR83, DR93, EMN88, Ept80, ED83, FM52, For53, FW60, Fos91, FHW48, GL86, GL91a, Gea81, GL95a, GL95c, Gil51, GF74, GO96, GS76b, Grc89, Grc90, Grc92, HW91a, Hal77, Hal80, HW76a, HW80, Hen62b, Hig90h, Hig91c, Hig91d, HS95, HP92a, Hou50, HW91b, Hul67, JS83, KLJ89, KH90a, KLP92a, KLP92b, KR78, Kov76, Kov77, Kup71, Lan70, Lau79, LYK89, Le85a, LN72, Lem75b, MPW66, McK84, Men56, Nab92, O’L93b, OP64, OV85, Ost73]. Equations [Pai73, PW71, Rhe66, Roh89b, Roh93, Roh94, RO88, Sal60, SWD76, SW84, Sha94, Sny55, Sol80, Spe91, Ste73b, Ste81d, Usp48, Van81, Var72, Van77a, Ver49, Vit69, Wac92, Wan91, WWWS86, Wes68, Woź77a, Wri95, You54, Zie88, Ben24, AP81, ADD89a, Bal80, BMO92, Bar65, BJW85, Cad73, Cad74, CS89a, Chu89, Cyb80, DEG72, Dah83, DNR87, Fel75, Fre80, GS78, HL86, Hua90, ITH88, KKM79, KLW90, KM76, Kog55, Lax65, dO69, Osb65, Per69, PW79b, Pet82, Ste73c, Vet75, Wil67b, Wil76, Wri82]. Equilibration [Bun71b, Lem76, McK62, Ske81, vdS69, vdS70a]. Equilibrium [Vav92]. Equipment [Sch90c]. Equivalence [Dem83, GMS^{+}86, Rad62, Rin55]. Equivalent [Lem75b, Syl51b]. Ergodic [Bar86b, Pet81, FM86]. Errata [RW92]. Erreurs [PV74a]. Error [Abd71, AB79, AM94, AL85, AL86, ADD89b, AM65, Att57, Auc92, Bar81a, Bar81b, BB85b, BB85a, Bar88a, BZ90, Bar91a, Bar93a, Bar93b, BR71b, Bev69, Bjö91a, Bjö91b, BBG92, BGM^{+}94, dBP77, Bra95, Bre70b, BF90, BC86, Bul80, Car59, CG67, CB90a, CH94, CG85, Cle78, Cyb78, Cyb81, DKW82, Dei90, Dem87d, DK93a, DK93b, DH93, DM90, DM71, Dou59, Eir93, Enr89, Fox78, Ful87, GL94, Gen69, Gen75, GL95a, GK89, GK90a, GF75, GFB85, Grc92, Gul92, Gul95, Hen63, Hen80, Hig87b, Hig90b, Hig90c, HH92a, HK93a, Hig93d, Hou58a, Mül83, Iri91, JO74, Jor68, Kah72b, Kas95, KLW90, Kie87, Kni93, Kub91, LS80, LPW83, Lin76, Lyn64, Mei83, Mik91, Neu85, NP84, New73, New74, O’C93, OP64, Oli77, Oli79, Olv78, OW82]. Error [Olv83a, Olv83b, Olv86, Ort63, Pai73, Pai80, Par92a, Pry81, Pry84, Pry85, Qui94, QT90, Ral65, Ram71, Sau78, SR80, Ste71a, Ste71b, Ste73d, Ste92a, Sto77, Stp93, Stu80, Stu82, Stu85b, Stu85c, Stu85a, Sun92b, Tan91, Tay82, Tie70, Tsa81a, Tsa83, Ura68, VA86, War77, Wil60b, Wil61a, Wil61b, Wil62, Wil65c, Wil68b, Wil68c, Wil71a, Wil74a, Wil85, Wil86b, Wol65, Woź74, Woź80, Wun48, Zie88, Ziv82, Ziv95, AL75, CM89, Cyb79, DEG72, ITH88, KL89c, Lau75, Pow87a, Ste79a, Ste83a, Var91, Woź78, Stu85b]. Error-Control [Enr89]. Errors [Bau74, Bod49, Boh75, CI94, Dem89b, Des63b, Dwy53, Fle86, Fle88, For59a, Gar90a, Gau95, Gol62, Gra93a, Gur92, Gus70, HT87, Hen66, MM70, HS66, Kah65, KL73, Knu89a, LS78, Lem75b, Lin75, Low60, vMS94, Not93, Peñ94, Pet80a, Pet86, Pow67, Rad48, Ros65, SZ80, Ske89, Tsa74b, Tur48, Var90a, Wil63b, Wol64, Wol85, vdV90, Bjö88, Fos79a, Ste90e, VV87, Yal94]. Esimation [Van87a]. ESP [Swa88]. ESPRIT [Van87b]. Essays [Bro75, Gar85, MHR80, Pla93]. Essence [Hof85b]. Essentials [Hoo80]. Essex [Hig96d]. ESSL [DK94]. Estimate [BR71b, BB71, CMSW79, GEJ95, Sol80, Wah65, GM86, Lau88a]. Estimates [Auc92, Dem87e, Gau75a, GKL95, Hag84, KL89b, KL91a, KL94, Kub91, Qi84, Stu82, Stu85b, Stu85a, Zie88, Bjö88, ITH88, KL89c, Var91]. Estimating [CG85, Dix83, He90, Hig88c, Hig92b, KP93, KW92, Van87c, OSV79, O’L80, PSS82, Wat92e]. Estimation [And91a, AM65, Bar74, Bar91d, Bis90b, BLP90, CL72, FGP91, GL81b, Gus70, Hig87c, Hig88c, Iri91, KP94, Kat60, Lar66, Lin75, Mat95c, Mei69, PP92, SV79, Sch86b, SB92, Tan94, AGS91, All83, DCHS88, Van82a, Van87b, Mar63]. Estimator [Bye84, CCV82, Hig90d]. Estimators [CR83, KW89, KP92, Ste80b, ZWS86]. ETA [DD89]. ETA-10P [DD89]. etc [Kah83a]. Ethics [Har76]. Etude [PV74a]. Euclidean [Blu78, Far87b, Ost73, Ove83]. Euclidian [Mat85]. Euler [Bai88a, BL91, Nab92]. Europe [PA88, Bry91]. European [Jos91]. Evading [RK75]. Evaluates [Kah80c]. Evaluating [AD90a, Cro41, Fik67, Fle65, Fri84, JT74, LN92, Mat91b, Oli75, Oli77]. Evaluation [Bla64, Bre74, Bre76, CGR93, CJ70, Coo67, Cox72, Cox78, CH80, EGL^{+}87, Fik67, Fos79b, Fra58, Gar72, Gus70, HPW90, Kah86c, Knu62, New73, New74, NT58, Oli79, Olv86, RR82, Stu81, Woź74, Ziv83, Ziv91, Ein79, Van79, KW83]. Event [SGB91, BGS87]. Events [GSB88]. Ever [Byr88, Gar84, Gar90b, Qui83]. Every [Gol91a, Hym86, Kah73a, Bud87].

Everything [Ash91]. Everywhere

[Bar93d]. Evolution [DH90, TTR92]. Exact [EP94a, Fer94, Fer95, LO82, RK95, Ziv95]. exactly [VS88a]. Examinations [Bar82]. Example [Bag94, Dor71, HR90]. Examples [BN71, CR83, KKN^{+}92, KP74]. Exception [DL93, DL94, Hul82, HCSW88, HFT94, Kah92a]. Exception-Handling [Kah92a]. Exchange [HW75, Qui83]. Exclusion [BF60]. Execution [Bai93a]. Exercices [Cau41]. Exercise [Ros90]. Exercises [Eis92]. Existence [BB95, Cha84, Cul66]. Expanding [HB59]. Expansion [Coo67, KBS94, Lin76, OP83, SBKZ93, Tan75]. Expansions [DGM86, Smi68a]. Expected [Fle85a]. Experience [Bai93c, HGD90, Hig90d, Nix81]. Experiences [Bom93, Thi86, KL89a, MS74b, SR90]. Experiencing [SD86]. Experiment [Pet80b]. Experimental [Ost83, SF86, KBS90]. Experiments [CGT92b, Erh93, GD89, Hil88, Jor68, Ric66a, Tod53, CDM79, CGS83, GM77, JBNP91, Pic91]. Expert [HRV92, BCJR87]. Explained [MR87, MR90, Sch89b]. Explaining [Bro78]. Explanation [Don88a, Syl51a]. Explicit [Dav73, FM84, Ker69, Lan70, Rot84, Uhl81, Uhl85]. Exploiting [Hig90e]. Explorations [Mic92a]. Expm1 [Tan92]. Exponential [Bag94, MNP84, MB93, MV78, Ng84, Rin64, Ste92a, Tan89, Tan90a, War77, BMW86, How83a, Kåg77a, Van75, Van77b, Van77c, Van78, Mat91b, Roc81, Sch91]. Exposition [CJLP92, Mar89]. Expository [Abb78, GS76a, Gil91b, SG83a, SG83b]. Expression [Eis68, WY64]. Expressions [AB79, BP67, Bre74, McK84, Stu81]. Extended [DCHH88b, DCHH88a, KKL^{+}93]. Extending [Dek71, Moh85]. Extensible [HH88]. Extension [BRK^{+}87, FM52, GBDM77, Inc83, Sol80, ABD^{+}90]. Extensions [DGL91a, DPW94, Hal86, KR78]. External [Bai89]. Extra [Bai88c]. Extraction [Par84, Gow58]. Extrapolation [SB88]. Extremal [Bus69b, Dix83, Pin85, Li88]. extreme [Ste75a]. Eye [Mac77d]. Eyes [Ewi94].

F [Hig90i, BMW86]. F90 [Ola93]. Fête [Hig92c]. Faber [ES88]. facility [MS88]. Factor [BK69, Gar94a, Gar94b, Gre52, Li95, Mag60, RJ93, Wat92b, Luk84, Luk85, Mac88]. factored [Ip87a]. Factoring [JP94, CB90b, Ste69]. Factorisation [EO79]. Factorisations [Bha94, CI91]. Factorization [AO86, ABO88, Bea79, Bis88b, BBdHS95, BH74, Car91, CH92b, CP95, CDO^{+}94, Cry73, DGKS76, Dem89a, Dem90a, DHS95, DOV94, DGR^{+}91, ER74, FGM95, Fos90, GP93, GHL86, GSS96, GK90a, Gul95, Hea67, Hig90h, Hig91c, JP93, Kel62, Kie87, Mat92d, Mei83, Sau72, SE90, SW81, Sco81, Ste77b, Sun95b, VC81, Bis89a, Bis91, BC90, BBVdH87, BJ74, DD89, DD90, DO90, GM86, Van88, Man79, MR89, MS79a, OW90, PW77, Ste79a, Ste91d]. Factorizations [BP94c, Cha84, Cha87, CMR90, DEKV92, FP74b, HP91, Pol87, Sch93, SB92, Ste93c, Sun91, Sun92b, BL90, BH91, CH90a, FP86b, FP86a, Gay91, Pai87, PP89, Pow87a, Vas90]. Factorizing [Fle76]. Factors [BL95, CS89b, ETA90, FM84, FM85, HH89, GMSW87b, Gre85, RS88a]. Fail [Fos94]. Failure [Uni92]. Fake [Ano79, Ano83, Ano85d, Ano8x, Anoxxc, BG89, B^{+}81, Edm85, Won88]. Fallacies [Ker90]. False [Ede92a]. Fame [Gar90d]. Familiar [Bar92b]. Family [Dig91, Ost54]. Fan [Wat92g]. FAN’s [Ola93]. Far [Ano24]. Faraday [Jef60, Jon70, Wil65d, Wil71c, Wil71d]. Fast [AP94, BD87, BL80, BKO94, BB84, BH74, CR93, CR76, Csa76, Dem89a, DS89, DH92, DL93, DFW90, EK89, GS66, GO94, Hig88b, Hig90e, Hig90g, HS90, Kni95, LN92, LNOP82, Lu94, Lu95, LQ87, MP74, Mat93a, NRT92a, Pai79, Pan78, PR89, PS92, PP92, PA92, Ram71, Rat82, Rei91, RS88a, Sch73, Swe84, TSY88, TST86, Ziv91, vdV92, Van92]. Faster [DL94, HW79b, Inc85, Pan84a, dP80, HW78b]. Fastest [AJ88]. Fault [BGM^{+}94]. Faults [Bie37]. Favorite [Ker84]. Favourite [bGBB87]. Feedback [PEM81, DK84]. Feigenbaum [Bri91]. Feingold [Wil87b]. Few [EK92]. Fewer [Pry81, Lin72]. FFT [Coo90, CT93, Coo94]. FFTs [Bai89]. Fiabilité [Bru89]. Fibonacci [GL80]. Fidelity [Dav72]. Field [BH95a, Joh78, Fle84, Uhl85]. Fields [EBR87, Hen62a]. Fifth [HW76c]. Fifty [Gra92]. Fifty-Ninth [Gra92]. Figures [Qui83, Rai69a, Tom88]. File [BM92, Ste91a]. Fill [LNOP82]. Fill-in [LNOP82]. Filtering [SW82, Pai85a]. Filters [Sor92]. Final [Eis68, LS82]. Finally [Ruh87]. Find [Blu78, Cip89, Goe94, LL86, PM88, Pem69, Sch90c, Cra86]. Finder [Rod78]. Finding [Ada67, BD85, Bre76, CM83, Dek69, Le85b, O’L93a, Roh92, Wat94, Wil55b]. Fine [Per91, Sta91]. Fingers [Buc59]. Finite [Bar88b, CCF96, CF93, CFB93, FW60, GS92, Gre94, Hal74c, HW80, HK93b, HK95, LNOP82, Lyn87, MK85, Mil88a, Ode90, Osb85, SL54, Tau49, TM62, Tod54b, Var72, Cad73, LR81, Mat90a, Mey75, Osb65, Pot76b]. Finite-Difference [FW60, Var72, Osb65]. Finite-Dimensional [Hal74c, LR81]. Finite-element [LNOP82]. First [ACM87, Asl82, Bha93, CJLP93, DH95, Ede91b, Hop81, HPP88, Rai69a, Rai69b, Rai76, RR78, Sam83, Ste80a]. Fischer [IIM87]. Fists [Buc59]. Fit [AN63, Mat85]. Fitting [AHB87, DW80, For57a, SC70, Smi65]. Five [GA94]. Fixed [Ura68, GK83]. Fixed-Point [Ura68]. Flaw [Ede95, Mar94, SB94].

Fletcher [Pow87a]. Floating [Ano86a, BSBF89, Bar81a, Bar81b, BB85b, BB85a, BZ90, Bar89, Boh77, BGM^{+}94, Bre73b, Bro81, Car59, Car91, CO84, Cli90, Cod73, Cod81, CCG^{+}84, Cod88b, CC93, Dek71, Dem89b, Dem92e, DDR94, Esp78, Far88b, Fat82, FG82, FT86, GB91, GM74a, Geo78, Gol91a, GFB85, Gre72, IEE81, Hen80, Hig91e, Hig93a, Hou81, IEE85, IEE87, JSW83, KP79, Kah80a, Kah80b, Kah81, Kah83b, Kah86a, Kah86b, Kah87b, Kah92a, Kah95, KL73, Kar85, KPS77, KC73, LIA93, LO82, Lin74, Lin75, Lin81, Lin70, Mal71, Mal72, MI81, Mat70, MS76, Mol67, Møl65b, Pla88, Pri91a, Pri92, RK75, RS88b, Sch81a, Sch84, SB94, Smi91, SW90, Ste74a, Stu82, Swe65, Tan89, Tan90b, Tan92, Tie70, VS93, Vit68, Wal90, Wic89, Wic92, Wil60b, Wil63a, Wil88b, Yoh73, Rum91, Tsa74a]. Floating-Point [Ano86a, Bar89, Bre73b, Bro81, Cod73, Cod81, CCG^{+}84, Cod88b, CC93, Dek71, Esp78, Far88b, Fat82, FT86, Geo78, Gol91a, GFB85, Hou81, IEE85, IEE87, JSW83, KP79, Kah80b, Kah81, Kah83b, Kah86a, Kah86b, Kah87b, Kah92a, Kah95, KL73, Kar85, KPS77, Lin74, Lin75, Lin81, Lin70, Mal71, Mal72, MI81, Mat70, MS76, Pla88, RK75, RS88b, Sch81a, Sch84, Smi91, SW90, Ste74a, Stu82, Swe65, Tan89, Tan90b, Tan92, Tie70, VS93, Wal90, Wil60b, Wil63a, Yoh73, Boh77, Wil88b, Rum91, Tsa74a]. Floquet [FJ91]. Flottante [Pic72]. Flow [HR90]. Flows [TTR92]. Fluid [CHQZ88, Har65]. Flushes [Gla95]. Following [HR90]. Fonction [Bar79b]. Font [Hof85a]. Fonts [Knu87a]. Fool [Bai91]. Foot [SD86]. Football [Kee93]. Footer [Grä94]. Footnote [Ken85]. Footnote-in-mouth [Ken85]. Forays [Mal89]. Forecasting [Ing95]. Forecasts [Ing95]. Foreign [BR67, BR68, Lar83]. Foreign-Language [Lar83]. Form [Aas71, AR92, BD93c, CDW94, CH89, DvdG92, Doy57, FS83, For92, Gei91, GM73, Giv58, GW76, Kåg93, KL88, Lan93, Lei89, Mee80, MH57, PL82, Par92b, RR82, Tal91, Tao84b, Tho42, Wil65c, Wil79, Zha92, Bus69a, Dub91a, FT88, FG86, GLW89, Han90b, Ip87a, KR80a, KR80b]. Formal [Bar89, dJ77, Wic89]. Formalization [Mat70]. Formats [BM92]. Forms [DHS89, Ioh82, Ric65, Wil59, GS58]. Formula [Cas89, EH92, Har72, Pow87b, Yip86]. Formulae [Gau83b, Gus70, Fle89, Lev83]. Formulas [AS64, McC87]. Formulation [Inc83, Inc84, FNO87, Wan91]. Forsythe [Var90b]. Fortran [Bur87, EPL94, MR87, MGH81a, Bai93a, BRK^{+}87, DGL91a, Don82, DH79, DE84, EP87b, GWL^{+}92, HK87, Hig88c, Kau78, Knu71, McC84, PFTV86, PTVF92b, Smi91, Sun92d, Sun92c, Tro84a, Bai93b, Bai94b, BM79, Blu78, BGA90, Bre78b, Bre78a, BDCS88, Buc94a, Buc94b, CGT92a, DCH^{+}95, DCHH88b, DCHH88a, DS94, DR82, DR93, DR95, FGMS90, GHM^{+}86, GMSW86c, HL86, IBM93, Ken91a, Ker75, Ker93, KO91, KLS^{+}94, Kub91, LHKK79, MD64, MR90, Mol2a, Mon89, NAG90, Ola93, Pre93, Rei92, Wil95, Wor89]. FORTRAN-77 [GWL^{+}92]. Fortran-90 [Bai94b, DS94]. Forward [KF63c, PL93, Stu85b, Stu85c, Wat95]. Found [Ruh87]. Foundations [Sma90, Maz76]. Four [HW75, Inv93]. Four-Atom [HW75]. Fourier [Kör88b, LT88, Van92, Boy94, BH95b, CT65, GS66, Gen69, New73, Oli77, Par82, Ram71]. Fox [MS78c]. FPV [Ano86a]. Fractal [BDM^{+}88]. Fractals [Bar93d, PJS92a, PJS92b]. Fraction [KBS94]. fractional [Coh74b, WA67]. Fractions [Ant85, Bla64, Cor92a, JT74, Kah87d, Knu85b, STY83]. Frameworks [Van92]. Frank [Ano84, Var86a]. Franklin [Lem61]. Fraud [Bra87]. Frechet [Mat91b].

Free [BDG94a, Cas93, Fre93, Le85a, LNOP82, MI81, Tsa83, WSW72, GK83]. Freedom [Pas85]. freien [Zie75]. Frequency [New81]. Frequently [Ola93]. Fresnel [Act74]. Friends [KB92]. Frobenius [Dub94, Kee93]. Frontal [DS93b]. Fry [ACD^{+}93]. FTN90 [NAG92, Sal93]. FTP [Hig92d, Ste91a]. FTPing [Ste91b]. Fucntions [STY83]. Full [CNRT81, ADD89a]. Fully [Par94, DS87]. Fumblerules [Saf90]. Fun [GS66]. Function [BFP88, Bre76, Bye87, BHM94, Cod82, Cod87, CH89, Cri75, DB76, Fik67, Fle65, GL86, KLJ89, KL91a, KL91b, KL92, KL93, KL95, KBS94, Le85b, MNP84, Nas79, Nas90a, NM65, PKL90, Pow86a, Rob80, Sch83a, Sti91, Tan89, Tan90b, Tan92, TSY88, Ziv83, Bal80, BGBMN90, Bye86, CH90b, CGS83, DK83, Rin55, Rin56]. Functional [HG80, Row90, Dav73]. Functions [AS64, Bax91, Boh77, BB84, CGR93, CK57, CW80, Cod93, Dav85, GK93a, Gou86a, Gre93, GL80, Hew82, HFT94, Kah87a, KL89b, KL94, Kit85, FB91, Lan38, Mat95c, Mir60, Par74, Par76, Riv69, STW84, Sta78, Syl51b, Tan90a, Tan91, Uhl92a, Ziv91, BMW86, Des63a, Fil81, GS71, Kåg77b, Lag74, LM85, Mat90a, Mat91a, Vrs86, CC93]. Fundamental [Bis90a, Knuta, BP94b]. Fundamentals [Wat91b]. Further [Fis74, Hey87, How83a, Kah65, Tur84]. Futility [Kah89b]. Future [BSBF89]. FX [DD90, Duf89a]. FX/8 [Duf89a]. FX/80 [DD90].

G [Hig90i]. Gained [Coo90]. Gains [PEM81]. Galerkin [Wat87]. Gambling [DF87]. Game [CH68]. Games [Aug84]. Gatlinburg [Var90c]. Gauge [Mir60]. Gauss [For51, Ste94a, EJ86, FM52, AM87, BR71a, DH89, Dub91a, GK83, PW75, Sau78, Yal91]. Gaussian [Fos94, GKO94, Str69, Vav89, Ban76, BB85b, Boh75, Bro73, Bus71, CG85, Coh74a, Cry68, CR72, DK77, Dax82, DP88, DEGR88, Ede92a, Gau83b, GLW89, GM77, Gou91a, GN82, HH89, Hig90b, Hig90f, Ker92, Kos92, Mac88, Mac89b, Mil71b, ON93, OW82, Pap92, PD92, Rei71, Rob90, Ske79, Ske80, Sor85, Ste74b, Ste93a, Ste94a, Stu82, Stu85b, Stu85c, Stu85a, Tor65, Tre85, TS90, van77d, Wri93, YC95]. Gdansk [Maz76]. GEMM [KV91, KLV91]. GEMM-Based [KLV91, KV91]. GEMMW [DHSS94]. Gems [CJLP92]. Genauigkeitsfragen [Bau66]. General [Bre74, But87, Cyb83, Gei91, GMSW91, Giv58, Gru70, Hym57, Joh78, KL94, Tau72, Wat91d, DK81, GMSW87b, Han90b, Kåg84, Pai85b, Per69, Ruh70, Tau70]. Generalisation [ES88]. Generalised [EO79, Kas95]. Generality [PW94]. Generalization [GL86, GHS87, Var86a, Wil42, LS90, Ruh68]. Generalizations [FV62, Wat93b, MZ91]. Generalized [AL84, BD93a, CM79, CGO76, DK93a, DK93b, FH93, FNP82, Gow75, KW89, Kåg93, KP93, Kåg94, Kau87, Van85b, Lu95, MN89, MG89, Nab91, Nab9x, NR76, Nas76, PS81, PS91, RM71, Sch66, Sco81, Ste83b, Ste92b, Van81, Van77a, VS88b, Wan91, Wed85, Zie86, DLR81, Kåg84, Kub70, Van83a, Mey75, Pai79, Pai85b, Pai86, Pai87, Sch83c, SS74, Tan75, Wil77a, Zie78, Zie84]. Generalizing [CCV82, Van76a]. Generated [BR71b, Hou58a]. Generating [BJ89, MO84, Gre85]. Generation [AOU87, BD87, DM89, For57a, MS76, Ste80b, TS80]. Generators [And90a, PM88]. Genius [Dun90]. Geodesy [Kol78]. Geodetic [WH16]. Geometric [BB84, Hof89a, Mil88a, SSSU87]. Geometry [Dem90b, KL93, BC90]. George [Ano76, Dys93, Var90b]. Gerling [For51]. Gerschgorin [FV62, Tod65]. Gershwin [Ano76]. Get [Fox71, Hol74a, PP94].

Getting [Woo92]. Giant [Ber49]. gigaflop [DH86]. Give [Har91b, JHL93]. Given [Mat85, OP64, RG67, Syl51a, GR89, Kel75]. Givens [BI87, Gen73, Gen75, Ham74, Rat82]. Giving [Bai91]. Gleanings [Ano24]. Gleichungssysteme [Bau66]. gleitendem [SB53]. Glenn [Mil43]. Global [DS75, LR93, LaQ94, QCM94, CL92, LM85]. globally [Li91]. Glossary [‘Vi52]. GMBACK [Kas95]. GMRES [BHR94, Bro91c, Kar91, NRT92b, SRD94, TW92b]. GMRES-Method [Kar91]. GNU [CR91, HH88, Sta86]. go [Knu74a]. God [Ste89]. Goertzel [Gen69]. Goes [Fit93, Lew78]. Golden [L^{+}64]. Goldstine [Sal81]. Golub [Hig90i, Hig92c, Moh85, Pow87a]. Good [EE84, Hig92a, JHL93, Kah83b, Kir92, Man90, MH80, PM88, Par73, Saf90, She93, Hal91a, Lau88a]. Gooder [Hal82b]. Got [Bry90, Reg87]. Gourmet [Spi90]. Gover [Hig90j]. GP [Gou91b]. grade [Sam83]. Graded [SZ91a, Ste94b]. Gradient [AK90, BB88a, CGO76, CGM85, EAS94, GO89, Gre89, GS92, Gre94, PS92, PKPL91, She94, Str91, Woź80, CM86, Fre90]. Gradients [Not93, CS89a, Ste75a]. Graduate [ILA93]. grained [Bis88b]. Gram [JP90, Abd71, Bjö67b, BP92a, Bjö94, DGKS76, Hof89b, JP91, OW90, Ric66a, Ruh83]. Grammar [Gor84, Led68, Saf90]. GRAP [BK86a, BK86b]. Graph [Mil76, PPY92, PPY93]. Graphic [Com83]. Graphical [CCKT83]. Graphics [BM92, BH87, FvDFH90, Mac77a, Mac77b, Mac77c, Tuf88, Pic90]. Graphs [AS64, Bau74, BK86b, DG93, Mac77a, Mac77c, Mac90, Yal94]. Great [Ano78b, Dun90, Par85, Wil72a]. Greats [Anoxxa]. Greek [vdW75]. Green [Dys93]. Greenbaum [BDDC88]. Greps [Hum88]. Grooks [Hei92a, Hei92b]. Group [MMR78b, GM86, Mey75, CJLP93].

Groupes [Aut02]. Growth

[DP88, FS80, Gou91a, HH89, Jor68, Tho42, Mac88, MZB87, Pow87a]. GSVD [Han89, Han90b]. Guaranteed [CC91, GAKK93, Kos93, Mal92a, Mal92b, ST86]. guard [Olv83b]. Guerrilla [Bro89]. Guide [Ame91, ABB^{+}92, ABB^{+}95, Ano86a, Aug84, Bak61, Bar86c, Bar78, Bar82, Bee92, Ber65, Ber71b, Boi90, BGA90, CB81, Car58, CM93, CS87, Day92, Dig91, DBMS79, DPW95, DGL92, Eis92, FF87a, GBDM77, GJ79, GMSW86c, Hew93, HWH91, IBM88, IBM93, Jud90, Ken93a, Ken93b, Kat85, KD93, KD95, Kro92, Kro94, LR93, LaQ94, Lam86b, Lam94, Lin84b, Lon93, Mac90, Man90, Mat92c, MH80, Mil87a, MLB87a, MLB87b, MW93, Mos71a, Par90a, Par93, Par73, Pec87, Pre78, Rat85, Rob88, Saf90, Sch79, SBD^{+}76, Spi90, Sun92d, Sun92c, Tar92, Trz93, Woe92, Zin88a, Zin90, AD90b, Den84, GHM^{+}86, Mol82b]. Guidelines [DCHS88, JBNP91, Mat90d]. Guinness [Man88].

H [Ano78a, Far87a, Hig90i, Par90c, Eva76, Nab91, Sal81]. H-Matrices [Nab91]. Hacker [Cor88, Ray91]. Hadamard [HW78a, Wal93, BGH62, Bir75, EM95a, GB63, GS84, HM90a, WSW72, Wat92e]. Hadronic [A^{+}90]. half [Gau94]. half-century [Gau94]. Halftones [Knu87a]. Hall [Gar90d, Pre78]. Halley [Gan85]. Halmos [Alb91, EG91, Gil91b, Pri91b, SG83c, SG83b, SG83a]. Hamilton [Mac95]. Hamiltonian [DDLT91, Van84a, PV81, SS92, SSC94]. Hammarling [BDDC88]. Hand [HH92b, OP64, SG92, Sun95c, Ste81c]. hand-held [Ste81c]. Handbook [AS64, Mer91, Bow90, Bro89, But81, But92a, CL90a, CV88, Com83, Cor88, Gor84, Gor93b, GH43, Hew82, Hig93b, Hig94b, Ken82, KLS^{+}94, vL78, vL92, Lue87, Mic92a, MS89, PP94, SG91, Wes68, You89b, Zwi92]. Handheld [Kah80c]. Handling [DL93, DL94, Hul82, HCSW88, HFT94, Kah92a]. Hankel [Ioh82, Tyr94]. Hard [PM88, PR93, Roh93, RK95]. Harder [Sch86d]. Hardness [Roh95]. Hardware [GS90a, PH94]. Hardware/Software [PH94]. Hardy [Fox87a]. Harmonic [Ter85]. Harmonics [FK77]. Harry [Sto80]. Hartley [BF93a]. Harvest [Mac77d]. Harwell [DGL92]. Haugeland [Kah73b]. Head [Buz82]. Header [Grä94]. Hear [DGGW89, GWW92, Kac66]. Heat [Dou59, Gil86, Low60, HLW83]. Hebetization [Rol76c]. Held [AK90, Ste81c]. Helix [Cri74, Wat80b]. Helmert [Lan65]. Helmut [Ips94a]. Here [Bry91]. Heritage [Hou82]. Herman [Sal81]. Hermite [ET73, GP71]. Hermitian [AG88, KM76, CES93, Chu95, CM83, Fre93, Kah75, Kel75, Mat92b, Ste85a, Ste91e, SH93, Sun95a, VS93]. Hessenberg [BD89a, BGH92, Bus69a, CS86, Chi93, CDW94, Cho69, Dat87, Dub91a, Fai71, GNV79, Gra86, Ike79, Van82b, vMS94]. Hessenberg-Like [CS86]. Hessians [Ste81b]. Hierarchical [Bai89, GJM87, GJMS88, RG92a]. High [Bai88c, BF88, Bai93d, BMvN46, BB49, DDC^{+}87, DDK92, DS86a, DKH86, DW93, DPW94, DLN^{+}94, DW95, Dow93, Fat82, GS90b, GvN51, IT89, JW86, JP94, KLV91, KLS^{+}94, MS93b, vNG47, Rad48, Tal91, Tho92, TW92b, ABB^{+}90, ABD^{+}90, BD89b, Bis90a, DS86b, DS88, PH85, BRR85, DPW95]. High-Accuracy [JW86, BRR85]. High-Level [Fat82]. High-Performance [DDC^{+}87, DKH86, JP94, ABB^{+}90, ABD^{+}90, BD89b, Bis90a, DS86b, DS88, PH85]. High-Speed [BF88, BB49, Rad48]. Higher [Caj29, Kah89b, Vrs86]. Higher-Precision [Kah89b]. Highly [AG89, APS93, BCD95, Gar73a, Gar73b, Gar77]. Hilbert [Hal67b, Hal82a, Rei86, AY81, Att57, BD85, Cho83, Pow86b, SL54, Tau49, Tod54b, Tod61, You88]. Hilbert-Courant [Rei86].

Hill [Lon93]. Hille [Bag94]. Hints [Har75]. Historical [Ode90, You89a, You90]. History [AAR87, AM87, Bec71, Bel31, Boy68, Caj29, CK89, Eva81, Gol77, GO89, Hal90a, Har76, HPS83, Ips94b, Iri91, Jac86, KR65, McC84, MHR80, Nas90b, PW94, Sal81, Sig86, Ste93b, Tho83, Wex81, Woo58, PW85]. Hobgoblins [Ber71b]. hoc [JBNP91]. Hodges [CK84]. Hollywood [Pre91]. Home [CM87, Mic92b]. Homology [DGGW89]. Homotopies [LZC92]. Homotopy [LZS91, LZ92, LL93]. Homotopy-Determinant [LZ92]. Hopkins [Hig90i]. Horn [Uhl92b]. Horner [Mül83]. Horrors [‘Vi52]. Hosts [HHK93]. Hough [Kah73b, Sch89b]. House [Bod86, 87]. Householder [BV87, BG67, BG65, Cup84, CB90b, Dan76, Kau87, MS78c, OW90, Ort63, PR68, PR69b, Pug92, RS86, SV89, Tsa75]. Householder-Fox [MS78c]. Howard [Hig95d]. HP [Hew82, Hew93, Kah83a]. HP-15C [Hew82, Kah83a]. Huang [Bro85]. Hue [Rol75c]. hulls [All88]. Humanities [Gar79]. Hurts [Kah90b]. Hurwitz [Asn70, Kem82]. Hybrid [Dax90c, NRT92b, Neu92]. Hydrodynamic [TTRD92, TTRD93]. Hydrodynamical [HR90]. Hyman [Hym86, War76]. Hyperbolic [CB90b, KL93, RS86, APP88, DI86]. Hypercube [GL88]. hyperplane [Rot84]. hyperpower [Tan75]. hyphenate [Ock75]. Hypothesis [Roe86].

IBM [CS87, DEKV92, DD89, DMdB90, Fri84, HP86b, KL89a, KL85, MS88, PA88, PH85, PH86, SoIPUG86, SoIPUG87, Sta88]. ICCG [RG92a]. ICL [CK89]. ideas [Dat89]. idempotent [HC63]. Identification [BS92a]. Identifying [BKW80]. Identity [Bre70b, KL93, Pot76b]. Idiom [Cok75]. Idioms [FF92]. IEC [LIA93]. IEEE [CC93, Dem91, Fat82, GB91, Hou81, IEE85, IEE87, Kah80a, Kah80b, Kah83b, Kah87b, Kah95, Pax91, Tan89, Tan90b, Tan92]. II [DK93b, BD94, Bjö68, Caj29, Gau53b, Gau63, GvN51, Goo92, HW91a, Hei92b, Mal92b, MC90, Mil75a, MS78a, Rut52, Shi70, Stu85c, Tha61, Tod54a, Tsa92, Wil86c]. IIB [SoIPUG87]. III [Cor88, DK70, Dav93a, Gau78a, KvSR90, Ric77, Shi72]. Ill [Dem87c, DK88, Dem90b, Dor71, FGS95, GW76, Han92, Kah72a, Men56, Rhe66, Han90a, Han90c, PW79b, Rum91, Wil67b, Wil72c, Har91b]. Ill-Condition [Kah72a]. Ill-Conditioned [FGS95, GW76, PW79b, Wil67b, Wil72c]. Ill-Conditioning [Dor71, Rhe66, Dem90b]. ill-determined [Han90c]. Ill-Posed [Dem87c, DK88, Han92, Han90a, Han90c]. ILLIAC [Luk80, Sny55]. Illuminated [Knu91]. Illuminating [Moo36]. Illustrated [AAR87]. IMA [Hig92c]. Image [OP83, AP76, MS79a, SV79]. Images [BDM^{+}88]. Impact [GJMS88, Gar90a, Gar94a, Gar94b, Har84b, Rob90, She93, HH91a]. Impenetrable [Mad83]. implement [NOP84]. Implementation [AD90b, BD89a, BHR94, Bar88a, BH87, Cap71, CDO^{+}94, Cod82, CR76, DGL91b, DCHH88b, DCDH90b, Gal71, GMS92, GC91, Kah73b, LSV95, Sha82b, Tan89, Tan90b, Tan92, DBKF90, Nas85]. Implementations [KL89a, Mil88a]. Implementing [DGK84, DH86, HFT94, Sed78, Tsa75, Yoh79a]. Implications [Fat82, Mol90b, PP89]. Implicit [Dem92a, Ept80, Sor92, Ste81d, HLW83]. Implicitly [LS94, LSV95]. Implies [JW77, Ske80]. Importance [Bau63a]. Impossible [Mol90a]. Improve [Byr87, Phi87]. Improved [DH93, PS91]. Improvement [DW84, JP91, Sny55, JP90]. Improving [Don82, Don83, DMW83, FM52, Gre79, JI92, Kov70, Lin74, Wal90, Har81a, Har84a, Kub65, TK89]. In-and-Out [Mat68]. in-core [Gay91]. Inaccuracy [Dem92a, DW84]. Inaccurate [CR79, Kov76, Oet65]. Inadequate [Lyn69]. Inch [Ros90]. Including [Gar82a, Mac90]. incoming [Bul80]. Incompatibilities [Sch74]. Incomplete [Pol87, Man79, Vas90]. inconnues [Ben24]. Inconsistencies [O’C73]. Inconsistent [Ziȩ85]. Increasing [Hen92]. Incremental [Bis90b, BLP90, Ste90c]. Indefinite [AGL95, BG76, Bro91b, BP71, CH92a, DRMN79, DR83, DGR^{+}91, DR95, Fle76, GMPS92, JP93, JP94]. Independent [CCG^{+}84, IEE87, LIA93, vW66]. Index [CH88a, COT89, Gar91a, Lam87b, Bus90, Gar64, Gar90e]. Indexes [BK88]. Indexing [Gar64, Gar79]. Inductive [PS94]. industrial [Tog86]. Industry [BCD95]. Inequalities [BB61, BGFB94, FH55, HLP52, MM64, MO79, Shi67, Shi70, Shi72, Bel68, Fan51, HM90b, KZ91, Mat92b]. Inequality [AY81, MT93, Löt83, HM90a]. Inertia [BK77, GMSW91, DD88]. Inertia-Controlling [GMSW91]. Inertial [CL72]. inexact [GO88]. Infarction [Inv93]. Infectiousness [Gre92]. inférieur [Ben24]. Infinite [Art86, DL92a, Smi68a]. Influence [Bra95, Cha92, Ede93b, Ede94a, Sch77, Var90b, PW77]. Influential [BKW80]. INFO [Gal91]. Informal [Zin88a, Zin90]. Information [HH93, HP77, Kah84, Kel85, Kel88, LL86, PW78, Par92c, Rat85, Sal89, TW92a, Tuf83, Tuf90, WY63, LIA93]. Information-Based [Par92c, TW92a]. ing [Pat88]. Inherent [Dem92a, Lem75b]. Initial [Cas89, Enr89, EP87b, Lam91, LS82, LS91, RM67, AD89, Lau75]. Initial-Value [RM67]. Inner [Neu85, Win68, Wee89]. Innocent [Gor84, Gor93b]. Innovation [Gid86]. Innovators [Wil72a]. Input [Dat87]. Instability [CH89, Mac89a, Mat95d, Mil71b, Mil75c, PL93, TT87, TTR92, Tsa81b, Bye86].

Installation [Ken93a, Ken93b]. Institute [PA88]. Instructional [Har85, Har81a]. Instrumentation [BBG92]. Instruments [Har84b, Kah77]. Integer [LIA93]. Integers [Geo78]. Integral [BM82, Wan91, Wou65, Act74, Cha74b, Har09, Wri82]. Integrals [Bar92a, Bro91b, Dav86, Kah80c, Lyn87, Act74, KW83, Van78]. Integration [DR84, Gil51, Hul67, KF87, Low60, Mos71b, PdDKÜK83, QTD91, Rad48, Zwi92, Bur90]. Integrations [Qui94, QT90]. Integrators [SS92]. Integrity [SF87]. Intel [DO90, HK87, HGD90, Kah83a, Mar94, Tho88]. Intelligence [Inc92, JMV94, Tur50]. Intelligent [HRV90]. Intelligible [Boa81]. Intensity [MB93]. Interactive [Bro79, Woo85]. Interchanges [Fos86]. Interest [MMR78b]. Interesting [For53, III84]. Interface [BHY80, DK94, DCH^{+}95, DEK95, PH94]. Interfaces [Sch74, DLS83]. Interfacing [But92b]. Interior [FGS95, GMS88, Wri92, Wri95]. Interior-Point [Wri95, GMS88]. Interlace [EH91b, IIM87]. International [AAR87, Bra93, BCEP94, Bru82, HPP88, Inv93, Jam74, KvSR90, LIA93, Pet68]. Internet [Kro92, Kro94, LR93, LaQ94, QCM94]. Interpolating [ADW84]. Interpolation [Cox75, Dek69, EK89, Gus70, KF63b, KF63c, Lag74, Lan38, Pow67, RR82, Rei90, Smi65, Tal91, Wer84, TW91a]. Interpolatory [KE82]. Interpretation [AwJS85]. Interpreters [Bro79]. Intersection [BD85]. intersections [Rot84]. Interval [AH83, DK85, Gay88, Han69, Kah80a, Man81, Moo66, Moo79, Moo88, Nic75, Nic77, Nic80, Nic85, Roh89b, Roh92, Roh93, RK95, Yoh79b, Yoh80, RR73]. Interval-Analysis [Nic77]. Interview [Eva76]. Interviews [AA85, Alb91]. Intractability [GJ79]. Introducing [Har66]. Introduction [AH83, Atk89, Bar75a, BFBS64, Bel70, Che82a, Cia89, CLR90, EWK90, Fou81, Fox64, Fro69, Geh89, GO96, Grä85, Gra69, HC92, Hig96b, Hil74, JáJ92, Kit71, MM65, Mir61, Ort88, PW79a, PJS92a, Riv69, Rub88, SA73, She94, Ste73e, SB80b, Str86a, Str93, Tay82, TA32, Van83b, Wan88, WF82, YS89, You88, Pal86, Col91]. Introductory [Bru92, COT89, Goo92, IEE81]. INV [CM86]. Invariance [DR94, Ste84b]. Invariant [BD91, BHM94, Dem87e, DL92b, KLP93, Mir60, PW74, Ste71b, Ste87b, Wil75b, DHW92, HM90a, Van82a]. Invariants [Tur29]. Invented [Roj94, Ste80a]. Inverse [Att57, BDG94a, Coh75, CDG91, Els76, ETA90, GK65, Hag89, HS81a, HP94c, Ips94b, JI92, Joh82, Kah93b, Kol95, Lau88b, Meu92, Osb78, PS91, PW79b, Pow69, RM71, Róz87, SL54, Wil72b, AP81, BG87, DGK80, DGR79, FNO87, Fun88, Mey75, PP89, SS74, Ste75b, Wed78]. Inverses [Abd74, Asp59, BF81, Bar84, Caf63, CM79, CS86, Cho69, Dem77, DMS84, Dia78, Gau62, Gau63, Gau75a, Gau78a, Hen62a, Ike79, Ker69, Mee80, NR76, Nas76, NP84, Peñ95, PW70, Ste77a, VG87, VS88b, Wed73b, Zie86, BLR90, KT85, Sch83c, Tan75, Vas90, Zie78, Zie84]. Inversion [BF88, BHH93, BR71a, BH74, CR93, CH92c, Csa76, Doy57, Ehr70, EP94a, Gau69, GK90b, GO94, Gre55, Lew82, Mit48, NT58, Pei62, SC90, Tod53, Wes68, Wil61a, YI79, Zie79, GM86]. Invertible [Sch65]. Inverting [BE73b, GvN51, MH57, vNG47, ST75]. Investigation [PH72]. Investigations [Bar79b]. involutary [LN62]. involutory [HC63]. Involving [Bai88a, CR93, Gau83a, Gra93a, Hig88b, Van78]. IPA [EV85]. iPSC [DO90, HGD90]. iPSC/860 [HGD90, DO90]. Irreducible [Els76, Nod71, Har86a]. Irreproducible [Sch86a]. ISI [Gar90a]. Isn’t [For70]. ISO [LIA93, JW85b]. ISO/IEC [LIA93]. isolated [Ste75a, Tod65]. Isotropic [BG79]. Issues [Vav91]. Iterated [Kat60]. Iterates [Hen62a]. Iteration [BD89a, BF92, CJ70, Dia78, Dub94, DG94, Els76, HK93b, Ips94b, JI92, Kah92b, Kel65, KLJ89, Kni93, LS94, LSV95, Lei89, LZ94, McK62, Pai94, PS91, SB88, TT96, Zaw91, CGS83, Gan85, Kub63, LT88, Osb78, PW79b, Ste75c, Vrs86, Wil72b]. Iterations [Bul91, CNP94, KL93, NRT92a, Ste94d, Ste69, Woź78]. iterationsverfahren [Alb77]. Iterative [AR92, ADR92, Axe94, BI75, BIM76, BIMF76, BI77, BBD^{+}94, BBC^{+}94, BR71b, Bjö67a, BG67, Bjö68, BB71, Bjö78, Bjö90a, Bru95a, Dax90a, Dax90c, Dax90b, Des63b, DEK95, FGN92, Gul94, HY81, HW76b, Hig90h, Hig91c, HK93a, Hig95c, Hig96d, Hof89b, JW77, KL91b, Kie81, KH90b, Kov70, MPW66, MP77, MBI77, Mil88b, Mol67, NP78, O’L93a, dP80, SG92, Ske80, SS84, Spe80, Var62, Wac88, Wac92, Wil55b, You54, You71, You89a, You90, Zie78, Zla82, Bea79, BD76, DCHS88, DFW90, DGR79, GW66, GO88, GP90, Gow58, Laa58, Ple86, Pul66, SS74, Ste90d, Tan75, Sch33]. ITPACK [KY88]. IV [CSW83, EP87c, Grä94, Luk80].

J [Alb91, Ano78a, BDDC88, Hig90j, Eva76]. J. [Par90c]. Jacobi [DV92a, Hac93, AR85, BF92, Chu91a, FH60, Gab87, Hua75, KG83, Van86a, Mac95, Mas91, Mas94, Mat93a, Mat95a, Ruh68, SHW86, SS89, Sla92, Ste85a]. Jacobi-Like [Ste85a, Chu91a, Gab87]. Jacobi-type [Hua75]. James [Ano04, Ano87, Fox87a]. Jargon [‘Vi52]. Jazz [Anoxxd, Bal86, Cok75, Lee88, Lev89, Sto75a, Won88, Gid86]. Jellied [Orr74]. Jill [Knu87b]. Jim [Fox78, Fox90b, Gol90, Lee88]. JNF [KR80a]. John [Asp90, Rei93, Tau63]. Johns [Hig90i]. Johnson [Bos86, SN46, Uhl92b]. Joking [Fey85]. Jones [Ano84]. Jordan [AM87, BR71a, DH89, FS83, GW76, KR80a, KR80b, Loi69, MP83, Läu61, PW75, RB89, Yal91]. Jordan-Elimination [Läu61]. Joseph [Ano04]. Journal [Ano93, EV85, Gar72, Gar82b, Gar90a, Gea91, Mad74, Mad87, Sch86a, SD86, Syl51a]. Journalists [FDR86]. Journals [Gar73c, Gar76a, Gar82b, O’C78, O’C86]. Journey [Dun90, Far87c]. Joy [Hig92d, Spi90]. Julia [Vrs86]. July [MMR78b]. June [AK90, Syl51a].

Kac [EK94, TT91]. Kahan [Nic70].

Kahan-Babuskasche [Nic70]. Karmakar [Tur91a]. Karmarkar [GMS^{+}86, Str87]. Kaufman [JP93]. keep [Hil89]. Kent [Ano93]. Kernel [SB95]. kernels [ADD89a]. Key [Kah79]. King [FF31]. Kingdom [Fox90a]. kit [Wil88b]. kleinsten [Läu61]. knots [GK83]. Know [Gol91a, Hym86, Bud87]. Known [Cho69, Lin74, Rei93]. Knuth [Fre87b, Hoe87, Hof85a]. Kogbetliantz [PV86]. Komma [SB53]. Konditionszahl [Zie74]. konvergente [Rut60]. Kreiss [WT94]. Kronecker [Wil79, HS81b, HPS83, VP93]. Krylov [HR92, Bra95, HS95, Pai94, Ruh94]. KSR [Ken91a, Ken91b]. KSR1 [Bom93, Pap93a]. KSRlib [Ken93a, Ken93b]. KSRlib/BLAS [Ken93a]. KSRlib/LAPACK [Ken93b]. kubisch [Rut60]. Kutta [BGD^{+}89, But87, Cas89]. Ky [Wat92g].

L.A. [Ano84]. L.S.D. [Tur84]. Laboratory [BZ93, Ash86a, Mol82a, CK81, HGD90, Wil48a, Wil80, Woo58]. Labotatory [Wil75c]. Lady [Ano84]. Lagrange [RR82, Wer84]. Lagrangean [Mer88]. Laguerre [Kah92b, LZ94, Par64, Wat94]. lambda [Kub70]. lambda-matrices [Kub70]. Lamé [Wil65b]. LAMS [Leh92, Spi91]. LAMS-TEX [Leh92, Spi91]. L’analyse [Bru89]. Lancaster [Tur89]. LANCELOT [CGT92a, CGT92b].

Lanczos [BCEP94, Hig96b, Ste91c, BH95a, Coo94, CW85, FGP91, Gol73b, GU77, GLO81, GO89, Gre89, GS92, Gre94, KW92, MT94, MT95, PR81, PSS82, PNO89, Par90d, Par94, Ton92]. Landen [Squ87]. Lang [Hal88]. Language [AKW88, Bai87, BK86b, Boa66, BR67, BR68, Bur89, Col87, CH80, Cry87, Cry88, DM91, EJ86, Fat82, Fin94, For93, GH70, HC88, 87, Ive62, KT91, Ker84, LIA93, Lar83, Led87, Led91, Lon91, Pot76a, Sav67, BK86a, Fox87b, Gou86b, KO91, RJLK78, KKN^{+}92]. Languages [Bar86c, But92b, FG84, Loe74, Wex81, Ros72]. LAPACK [AD89, AD90a, AD90b, ABB^{+}90, ABD^{+}90, ABB^{+}92, ABB^{+}95, DO90, DK94, DPW94, DPW95, DCH^{+}95, Hig92e, Hig95c, Ken93b, KP93]. LAPACK-Style [KP93]. LAPACK/ESSL [DK94]. Laplace [Gau69, ST75]. Large [Ano48, Ata40, Bai94a, Ber92, BS92c, CES93, Col84, CH88b, CL90b, Col91, CGT92a, CGT92b, CNRT81, CW85, CW86, DHP94, Día89, Ede91b, Ede93b, ES92, Fos90, Gem86, GL81a, Hea84, HH89, HP92a, Ker89, KH90b, LM90, LSV95, MH57, MGM93, PR81, Rie56, Ruh94, Saa80, Saa92, Sau72, ST86, Sch73, Wou87, Woź77b, You71, Dat89, DFW90, Enr79, Gay87, ITH88, Ste90b, Woź78]. Large-Amplitude [LM90]. Large-Scale [Ano48, Ber92, CL90b, Col91, CGT92a, CGT92b, ES92, Ker89, Sau72, Wou87, Dat89, ITH88]. Largest [Fan93, KW92, PSS82, OSV79]. Last [Ziv91]. Latent [Wil55b, Kub70]. lattice [Cyb84a]. Launches [Ano93]. Law [Ben38, Gus88]. Lawson [Cli72]. Layout [Mil87a]. LCAS [Kah91a]. Leading [Bar88b, Peñ94, Tur82]. Learn [DM91, Zin88b]. Learner [Oxf89b, Und80]. Learning [CR91]. Least [Abd71, Abd74, AK75, ADdR89, AHB87, Aru92, BI87, BH88, Bar88a, BV92a, Bjö67a, BG67, Bjö67b, Bjö68, Bjö78, Bjö87, Bjö90b, Bjö91a, Bjö91b, Bjö91c, BPE92, Bjö96, BG65, CNP94, Chu91b, Cli73, DS89, DS79a, Eld80, Far88a, GL94, Gau95, Gay88, Gen73, GS90b, Gol65, Gol68, GKS76, GB86, GDD94, Gul92, GW92b, Gul94, Gul95, He90, Hea84, JO74, Jor68, Kar74, Kel62, Lar66, LH74, LN72, Van85c, Lon67, Lon84, Man81, MD92, PP89, Löt83, Löt84, PW70, Ple74, Pow67, PR68, PR69b, Rei91, Sau78, vdS75b, Ste77a, Van87a, VV91, Wah65, WKS95, Wed85, Wei90, WH16, APP88, All88, Bjö88, CH90a, Cyb84a, Dem87a, GHM^{+}86, GW66, GR70, GV80, Ip87a, Van76b, Van83a, Mar63, Pai79, Ste75b, Ste87a, Ste90e, VV87, Wed73a, Wed78]. least-change [Ip87a]. Least-Squares [ADdR89, AHB87, BH88, Bjö78, GB86, Jor68, Kel62, LN72, Van85c, All88, GHM^{+}86, Mar63]. Lecture [Asp81, Bar93c, Kah73b, Kah95, Kod79]. Lectures [ACM87, Bli72, CKW85, CCF96, Jac53, Jef60, Van76b, Rut90, Wed34]. Lecturing [Bro78]. Led [Uni92]. Left [Gea81, Hol75]. Legal [Ano79, Anoxxc]. Legendre’schen [Hil94]. Legible [Gou73]. Legit [Ano83]. Leja [Rei90]. Length [CCG^{+}84, Tie70]. Leroy [Ano78b]. Let [Kod79, Rol76c]. Letter [Rol75e]. Letter-to-the-Editor [Rol75e]. Letters [Jon70, Min74]. Leur [Bru89]. Level [ADD89a, COT89, DD89, DD90, DH92, DHSS94, DMR92, Fat82, Hig90e, KL89a, KLV91, Tho83, CMR90, KV91, MR89, DCDH90b, DCDH90a]. Level-3 [DH92, KLV91, KV91]. Level-Index [COT89]. Levinson [Cyb79, CH92a, Cyb80]. Lewis [ACD^{+}93]. Lexicographers [SE62]. Lexicography [Lan89a, Sim87]. Libraries [DDK92, DvdGW92, DW93, DW95]. Library [Anoxxb, CGG^{+}91, CDPW92, CDW94, DDC^{+}87, DLN^{+}94, FS76, FHS78, GB91, Ham94, HP88, IBM88, Cra89, Ken93a, Ken93b, KKL^{+}93, NAG90, PC-87, WWG51, ABB^{+}90, ABD^{+}90, BD89b, CH90d, Gay87, HL86, MS88]. Lie [Huf54]. Life [Ano85c, Bos86, Far87a, Jon70]. Lighthearted [Saf90]. Like [Bar91d, CR93, CS86, Gau83a, Kah89a, Lu96, Rei91, Ste85a, AG92, BH92, Chu91a, Gab87, Hig88b, Hig90k, Lau75, Vrs86]. Limit [Gem80]. limitation [Van77b]. Linda [Ano85c, Ano86b]. Line [Anoxxa, Dax90b, Dil93]. Linéaires [Aut02, Ben24, FKF68b]. Linear [ABS84, Abd71, Abd74, Ant94, AR94, AL85, AL86, ADD89b, AR92, AGL95, Ata40, Auc92, Bab72, BLS91, BP67, BMvN46, BI87, BBD^{+}94, BBC^{+}94, BSD60, BJ89, BG76, Bjö67a, BG67, Bjö67b, Bjö68, BG73, Bjö87, Bjö91a, BP94c, Bol85, Bol84, dBP77, dBK86, BMPW66, BGFB94, BS92c, BP92b, Bro68, BP71, Bun76, BK77, Bun87, BDV89, BG65, But87, CM79, CJLP92, CJLP93, CF88, CI95b, CG67, CK75, CDPW92, CDW94, Cia89, Cli73, CP76, CL91, ILA93, CR79, CDG91, Cro41, CNRT81, Cul85, Dah78, Dan84, Dat85, DD86, Dax90a, Dax90c, Dei86, Dek69, DHP94, DK85, DDC^{+}87, Dem92d, DHvdV92, Dem93a, DHvdV93, DGL91b, DGL91a, DGK84, DS84, DS86b, DH90, DDSv91, DvdGW92, DV92b, DW93, DPW94, DPW95, Don95, DEK95, DW95, DCHH88b, DCDH90b, DCHH88a]. Linear [DCDH90a, DRMN79, Duf81, DR82, DR83, DR93, DR95, Dwy51, Ede91b, Ede93b, Ede94a, EMN88, Eld80, ES92, FF63, FKF68a, Fad59, Far85, Far88a, FM52, For53, FM67, For67, FHW48, Fox64, FGN92, Fre93, Frö53, GL94, GJM87, GJMS88, GPS90, Gas83, Gay85, Gil74, GMW91, GS90b, GAKK93, GK92, GL91b, Gol65, Gol73b, GV79, GHMP74, Gra87, Grc89, Grc90, Grc92, GEJ95, Gul92, GW92b, Gul94, Gul95, Hag88, HW76b, HW80, HK87, Hea84, Hel78, Hig85, HS87, Hig90h, Hig91c, Hig91d, HH92a, HH92b, Hig94d, Hig95d, HP86a, Hof87, Hou50, MM70, Ing95, Jon62, Jor68, Kah66b, Kar74, Kas95, Kat76, Kel65, Kie81, KH90b, KR78, Kos93, Kov76, Kov77, Kup71, Lan70, LS82, LS83, Lau85, LHKK79, LYK89, Lay94, Lem75b, Leo93, LJ93]. Linear [Lu94, Lu95, Lus89, MP74, Mal92a, Mal92b, Mal93, Man81, MM65, MH87, Mat92b, MN89, Men56, MP77, Mir61, Mol78, MGM93, MD92, NRT92b, Nas79, Neu92, NP78, NP84, Nob69, O’L93b, OP64, OPW65, Oet65, Ort88, PR89, Peñ94, Löt83, Ple74, PEM81, PR68, PR69b, Rie56, RG67, Rob80, Roh89a, Roh89b, Roh93, Roh94, RK95, Roh95, RG92b, Sau72, ST86, Seb77, vdS70a, vdS70b, vdS75b, Smo95, Sny55, Sol80, Spe91, Spe80, Ste71b, Ste77a, Ste81d, Stp93, Str88, Str93, Sun95c, Ton92, Tra66, Tre90, TTR92, Van91, Ver49, vdVD89, WKS95, Wat92f, Wed85, WWWS86, Wes68, Wil54a, Wil60c, WR71, Wil74a, Wil74b, Wil76, Wil77b, Wil77c, Wil78a, Wol65, Woź77b, Wri95, Yip86, You71, Zie88, Ziȩ85, Zla82, vdV92, AL75, ADD89a, ABB^{+}90]. linear [ABD^{+}90, Bar65, BD89b, Bis90a, Bjö88, Bog74, Boj84, Buc77, Cad73, Cad74, Cal86, CF91, Chu89, CL92, CM83, Cra86, Cyb81, Cyb84a, DEG72, DK84, DCHS88, DFW90, Don87, DS86a, DH86, DMdB90, DNR87, Enr79, Fel75, Fis83, FGH71, FH90, Fre90, Geo72, GMSW86b, GMS^{+}86, GHM^{+}86, GMS88, Gla82, GS74, GO88, GP90, Gro87, Hua90, KKM79, KM76, Kog55, LR81, Loo72, Nas85, Osb78, Pai79, Pai85a, Pai85b, PR85, Per69, RR73, Sch89a, SR90, SW87, Ste73c, Ste75b, Ste90d, Ste91c, Str89, Tho92, Vet75, Wed78, Wil67b, Woź78, Ziȩ83, Ziȩ87, CJLP93, Mol2b, Nas90a]. Lineare [KS88]. Linearen [FF64]. linearer [Bau66]. Linearly [GM74c, Syl51b, Gay84]. Lines [KP81, RT92a]. Lingo [DM91]. Liniowa [KS92]. LINPACK [Bye84, CCV82, DBMS79, DS84, Don88a, Ste77c]. LINPACK-Style [Bye84]. Liouville [Eva71]. Lipschitz [Dav85, Kit85]. LISP [Wan84]. LISPer [FF87c]. List [Gar82a, Jef60, McI82]. Literary [Ber90, Bie37]. Literate [BKM86, Gur93, Knu84b, RM91, Thi86, Van87d]. Literature [GP59, Sch79, SF87, Hil89]. Little [Bie37, FF87c, Gay91]. Littlewood [Bol86].

Lives [Rus72]. Loan [Hig90i]. Local [Bul91, Dem77, Fan93, KL73, Bis91, Cal86, DW84]. local-memory-based [Cal86]. Localization [CK57]. Locating [GSB88]. Location [Mil71a, Sch86b, SGB91, BGS87]. Log [GL80]. Logarithm [SA75, Tan90a, Tan90b, Wou65, Cul66, KL89c]. Logarithmic [BB85b, BB85a, Str72, Str75]. Logarithms [Hel68, Putxx]. Logic [Gla95, KU68]. logical [Lam87a]. Logiciels [Bru89]. Long [Qui94, QT90]. Long-Term [Qui94, QT90]. Longman [HT87, Hig96d, Lon91, Lon95, Sum87]. Look [BRB76, CH92a, Cle78, DvdGW92, TT91, Gou91b]. Look-ahead [CH92a]. Lookup [Tan91]. Loop [Cow89a]. loops [DH79, SD90]. loosely [Van88]. Loss [Ros90, BP92a, FG82, FT86]. Lost [Bry89]. Lösung [Bau66]. Low [Cap71, Fle90, Gli89, AG88, Tho92]. Low-Cost [Gli89]. Lower [Chi93, EY36, GI88, HP92b, Lás94, Var75, AL75]. Lowers [Wat88a, Dem87a]. Lowry [Gar90b]. LP [Fle94, Fox69]. LR [Rut58, Rut60]. LR-Transformation [Rut58, Rut60]. LS [Sun95c]. LSQR [Sau94]. LSSOL [GHM^{+}86]. LU [CDO^{+}94, FM84, FP86a, Mat84b, Ste93c]. LU-Decompositions [Mat84b]. Lush [Ano85c]. Lyapunov [Bal80, BS68, BN87b, CV89, GLKH90, GL95b, GL95c, Ham82b, HK88, HS95, HP92a, Hod92, Jon84, KH95, Wac88, Wac92, Woo89]. Lyapunov-type [Woo89].

M [Bod49, Bri92, CD86, Hig90j, Inc92, Tur59, AS82, Goo92, Jun95, Nab9x]. M-Matrices [Nab9x, AS82]. M-Matrix [Jun95]. M. [Far87a]. MA27 [DR82]. MA42 [DS93b]. MA47 [DR95]. MA48 [DR93]. MACHAR [Cod88a]. Machine [Ata40, Bow53, Cod88a, CT65, Dek82, DGK84, FS76, Frö53, Gil51, Gay87, BHH93, Sch90a]. Machinery [Ano48, Tur50].

Machines [Ber49, Har84b, McC79, Rad48, Sch88a, Tor89, Tor92, Wil55a, DT91, Sch89a]. Macintosh [Com94]. Macro [Spi90]. Macros [Knu87b, Leh92]. Mad [Cri89]. Madison [MMR78b]. Magical [Mol93]. Magnetic [BSD60]. Magnetic-Tape [BSD60]. Magnus [Hig96d]. Mahler [Leh60]. Mail [DG87, Gay85, QCM94]. Mainly [Caj29]. Maintaining [GMSW87b]. majorants [Dah83]. Majorization [MO79]. Make [Boa81, Gou73]. MakeIndex [Lam87b]. Making [Eva81, Hur90, SW89]. mal [FKF68b]. mal-conditionées [FKF68b]. Man [Bro75]. Man-Month [Bro75]. Management [HS83, Ack89]. Manchester [Red93, CK80b, Wil75d]. Mandelbrot [PJS92b, Vrs86]. Mandelbrot-like [Vrs86]. Manifestations [Cot74]. manifesto [Sie86]. Manifold [AL86]. manipulability [Tog86]. Manipulation [BT64, Dav86, PW78, Sto75b, Sto77, van86b]. Manitoba [HW76c]. Mantissa [SS84]. Manual [Ano46, BH95b, CGG^{+}91, CW80, Com88, EJ86, Fla71, Gil87, Cra89, JW74, JW85b, Lam86b, Lam94, Lon93, NAG90, NAG92, Ame65, Uni69, Uni82, Sal93, Sta86, Tur87, HL86, Sam83]. Manuscript [Hay88, WPG92]. Many [EK93, EKS94, JSW83, Leh43, App85]. many-body [App85]. Maple [CGG^{+}91, The93, GH93, HC92]. Mapping [Cad88, Hén76, Ruh94]. Maps [SU84]. Marcel [Gil91a]. March [Hig92e]. Marching [Ban76, BR77]. Marcus [Gro94, ST80]. Mark [NAG90, Blo48, CK80b, Cor94, TT91]. Markov [Bar86b, Bar93b, FM86, GM86, Gra93a, HP84, Hey87, IM94, Mey75, Mey89, O’C93, O’L93a, SS76, Ste78, Ste83a, SZ91b, Ste93a, Ste94c]. Marvin [Gro94]. Maryland [Hig90i]. Mass [DGGW89, Wat87]. Masses [Bai91, Sea89].

Massive [Gay91]. Massively [GA94, HS93b]. Master [Dav92, Dav93a]. Mastering [Low84]. Masters [AFJS87]. Mat [Sch92b]. Materials [EAS94]. Math [Cau41, CM93, For70, Grä85, HW76c, Cra89]. Math. [Hig96b]. Mathematica [Her88, Wag91, Wol91, SK93, Web92b]. Mathematical [AS64, AA85, AAR87, AAR90, Ano04, Bab37, Bab69, Bar75a, BP79, BP94a, Boi90, Caj25, Caj29, Cow84, Dav72, DG87, Ede95, Ell94, Eve69, Eve88, FMM77, FHS78, GS76a, GB91, Gar61, Gar75, Gla95, Har84b, HR80, Hig93b, Hig94b, Hig94e, HRV90, JS83, KW83, Knu79, Knu85a, KLR89, Ler83, MMR78b, Maz76, MM82, Moo75, Mur93, Pét66, Pol57, Ral91, RW67, Ric71, Ric77, Ric81, Ric88a, Ric88b, Sca50, Sch79, Trz93, Var90d, Ziv91, CDM79, Lev83, PW77, LW86]. Mathematician [Asl93, Hal85, Har40, Lit78, Saw43, Sch86d, Ula76]. Mathematicians [Gar82a, Hal91b, Jam74, Pet68, Rej81, Tho83]. Mathematics [Abb78, Aga93, Bei66, Boa81, Bon91, Boy68, BZ93, Bri92, Caj29, CBB54, CR41, DN89, Día89, Dun90, DAM88, Eva74, Ewi94, Fie86, Fit93, For57b, Fra92a, Gar73a, Gar73b, Gar77, Gar82b, Gar90e, Gil87, GO78, GKP89, HH91b, Hal70, Hal74d, Hal90b, Har67, Hay88, Hig96d, JMV94, Jos91, KU68, Kah83a, Kel91, KC75, Kra93, Mau91, Mur76, Nic75, Nic80, Nic85, Par85, Pem69, Pic91, Ren72, Rut90, Saw55, Ste73a, SHSD73, Ste90a, Ste75d, Ste89, Str86a, Swa79, Tod77, vdW75, Web92b, WR44, Wol91, YG72, YG73, Day83a, EG91, Gau94, Pet91, Str87, Bus90, Gau94, Knu89b, Tur45]. Mathematik [Var90c]. MATLAB [GMS92, Hig89b, Hig91b, Mat92c, Mat92d, Mol82b, Mol93, Sig92, Sig94, But92b, GH93, Han92, Hig93e, Hig95g, MLB87a, MLB87b, SR95, SR90]. matric [Rin55, Rin56]. Matrice [Bar79b, PV74b]. Matrices [AO86, ABO88, AOU87, And87, And90b, CGR90, AG92, AR85, ADNR92, Art83, Art86, Asp59, Bai94a, Bar86a, BD90, BV92b, BL95, Bar90a, BF81, Bar84, Bau63b, Bau63a, Bau69, BR71a, Bax91, BW93, tBK84, BP79, BP94a, BT64, Bir75, BG79, BV87, BLP90, BS92b, BG77, BM86, Bro68, BE73b, Bun71b, Bun74, Bun82, BGG88, BGH92, Bus68, Bus69b, Bye88, CR93, CS86, CES93, Cau64, CR86, Cha93b, Cho69, Cle59, CJ70, Cry73, Cry76, CR72, Dan76, DL92a, Dav79, Dem89a, Dem90a, Dem77, DMS84, DK90, Dem90b, DG93, Den81, DLR85, DL92b, Dix83, DGK84, DHS89, DER86, Duf88, DGR^{+}91, Ebe71a, Ede88, Ede89, Ede93a, EP94a, EH92, EGL^{+}87, FKF68a, Fai71, FH55, Fan60, FTxx, Far87b, FV62, FL74, FP62, Fie86, Fle76, FS55, Fra58]. Matrices [FP81, FNP82, Gan59a, Gan59b, GK37, Gau53a, Gau53b, Gau62, Gau63, Gau75a, Gau75b, Gau78a, Gau83a, GI88, Gea69, Gem80, Gem86, GMS92, GKK90, GK90b, GK93a, GO94, GKO94, GS83, GvN51, GH59, GB63, GS84, Gra87, Gra69, Gre93, GK69, GL81b, Hac93, Han93, Han95, HW78a, Hel68, HS81a, Hen62a, Hig87c, Hig89b, Hig91b, Hig92f, Hol92, HW79b, Hou64, Hym57, Ike79, Ioh82, Jen77a, Joh82, JP93, JP94, JE95, Kah70, KPJ82, Kat60, Kau87, Kel62, Ker69, Kha72, Kit95, Kos92, Lad76, Lan65, Lan69, LT85, Lan93, Law72, LTT92, Leh60, Leh94b, Lem75a, Lem76, Lew82, Li88, LZC92, Loi69, Lot55, LQ87, Mac43, MO84, Mat92a, Mat92b, Mat95b, Mat84a, Mat84b, MS74a, Mee80, Men56, Meu92, MH57, Min88, Mir56a, Mir59, Nab91]. Matrices [Nab92, Nab9x, Neu86, vNG47, NP78, NP84, OS90, Ost54, Pan84a, Pan92, Par74, Par76, Par84, Par86, Peñ95, Per52, Pin85, Pol87, RM71, Rei91, RT92b, Rei87, Rin64, RB78, Roh89a, Róz87, Rut68, Saa80, Sch73, SRS73, Sha82c, Sha85, Sha91, SB88, vdS69, vdS75a, Ste80b, SZ91a, Ste94b, SW62, SB95, Syl84, Tao84a, Tau50, TM62, Tau72, Tod49, Tod54a, TT94, Tom92, Tre92b, Trexx, TST86, Tur29, TA32, Tyr91, Tyr94, VG87, Van95, Var79, VC81, VS88b, VS93, WSW72, Wal93, Wat91c, Wed34, Wil55b, Wil65c, Wil42, Win71, Wou65, YI79, You79, Zie79, Zie86, AS82, All88, And85, AG88, Ber71a, BLR90, BS74, BJ74, Chu86, Chu91a, Chu91b, CM83, Cra86, CL74, CMR90, Cyb81, Cyb84b, CB90b, DK83, Des63a, Ebe65, Ebe71b]. matrices [EP87a, FKF68b, Fil81, FG86, Fre90, FU69, Gab87, GHR93, GK89, GK90a, Gra86, GJSW87, GN82, Hal91a, Har86a, HW88, HLPW86, HM90b, HC63, JS77, Kah75, KKM79, KG83, KvW74, Kub70, LR81, LN62, MN79, MSS84, MZB87, MM88, Osb60, Paa71, PV81, Pan90, PS79, Rad62, Rin60, Ros69, Róz88, Ruh68, Ruh75, Rum91, SK77, Sch83c, Sha82a, Ste75c, Tau67, Tau70, Vas90, WG78, Wid65, Wil60a, Wil58, Wil68a, Wil72c, Wil84b, Wri82]. Matrix [Aas71, AMS74, AK75, Asn70, Att57, Bai88c, BF88, BK79, BK80, Bak84, BB95, BHH93, BI75, BIMF76, BI77, Bar86b, BS68, BFP88, BJ92, Bar92a, Bar94, BS72, Bar91d, Bas83, BS92a, BD73, BD74a, BD74b, BL88, Bel70, BR71b, Bha94, BM94, BL80, BF93a, BP94b, BB71, BH83, BMSV92, Bod49, Bod56, BL91, BGFB94, BP92b, Bre70a, Bre70b, BK69, BH74, BKP76, BR76, Bye87, BHM94, Caf63, CGR93, Cha92, Chi93, Cho83, Chu87, Chu95, CMSW79, Coh75, CS84, CV88, CW82, CW87, CH92c, Csa76, DEKV92, Dat87, Dat88, DR94, Dem83, DK87, DM89, Dem92c, DB76, Don82, DLN^{+}94, DHSS94, DS94, Doy57, Dub94, DRMN79, DS79b, Duf84a, Duf84b, DGL89, DGL92, Dwy53, EY36, Ede91a, EKS94, EM94, EK94, EM95a, EM95b, Ehr70, Els76]. Matrix [EHM94, ETA90, ER74, EU92, Eve66, Fam88, Fan93, FTxx, FF51, FP74a, Fis74, Fos86, Fra89, FS80, Fri91, GHN^{+}90, GBDM77, GL86, GWL^{+}92, GLAM92, Gau69, Gei91, GMSW84, Giv54, Giv58, GK92, Gol91b, GK65, Gol68, Gol69, GV74, GLO81, GV83, GHS87, GV89, GZ93, GW73, GB89, Gra71, Grc90, GT93, Gre55, GKL95, HG80, Hag89, Hal77, Hal80, He90, Hea67, HS81b, Hig83a, Hig86b, Hig86c, Hig86d, Hig87a, Hig88a, Hig88c, Hig89c, Hig89d, Hig90a, Hig90e, Hig90d, Hig90g, HS90, Hig90i, Hig90j, Hig92b, Hig92f, Hig93c, Hig93e, Hig94c, HK95, Hig95g, Hil88, HS95, HK71, HJ85, HJ91, HMW77c, Hot43, Hou58b, How83b, Inc85, Jen77b, Jen82, JM92, JS92, Joh78, Jun95, KR83, Kåg93, KP93, KP94, Kah66a, Kah73a, Kau83, Kem82, KL89b, KLJ89]. Matrix [KL91a, KL91b, KLP92a, KLP92b, KL92, KLP93, KL94, KL95, Ker90, KR78, Kni95, KBS94, Kol95, Kre72, KM86a, LPS92, Lás91, Lan70, LR83, Lar66, LN72, LSV95, LN92, LPP93, Lev58, Lev80, LZ91, Liu87, Van84b, Van89, Van90, MRK76, MM64, Mar89, MP83, MT95, Mat85, Mat90c, Mat95c, Mat95d, MB93, MP77, MBI77, MW80, Mil87b, Mil88b, Mir56b, Mir58, Mit48, Mod88, Mol2a, MV78, Mor80, NRT92a, NT58, New62, Ng84, Nod71, O’L80, Ort87, Pai73, Pan78, Pan84c, Pan84b, PS91, PKL90, Par92a, Par64, Par65, PD92, Pei62, PV94, Pow69, Pry85, Qi84, Rao80, Ric81, RB89, Roh92, Ruh87, SL54, Sch92a, Sch77, SC70, SB80a, SC90, STY83, STW84, Smi68b, SBD^{+}76, Sta78, Ste73e, Ste77b, Ste81c, Ste85a, SS90b, Ste94f, Sti91, Sto62, Str86b, SH93]. Matrix [Sun92a, Swa93, TG81, Tau49, TZ59, Tau88, TT91, Tod50, Tod53, Tod54b, Tod61, TSY88, Tsa81a, Tur48, Uhl92b, Var75, Var86a, Var93, Var62, Wac88, War77, Wat91b, Wat88a, Wat91d, Wat92c, Wat92g, WT94, Wes68, Wil59, Wil61a, Wil68c, Yau93, Zha92, Zie88, Ziȩ85, ZWS86, All88, AL73, Ash86a, Bal80, BIF69, BIM90, BMO92, Bar75b, BGH62, Bel68, BL90, BG87, BBK89, BGBMN90, Bus69a, BJW85, Bye86, Bye88, Cha74b, Chu89, Coh74b, Cul66, CV86, Dah83, DK83, DD88, Dat89, DM78, Dav81, Dav83, DK81, DGK80, DLR81, Don87, DM84, DBKF90, DGR79, Duf77, Duf89a, Duf89b, Dwy67, Fel75, Fle85b, FH60, GLW89, Gen78, GMSW87b, Gol73a, HH91a, Hig83b, Hor54, HS59, HW78b, How83a, Hua90, Hua75, Kåg77a, Kåg77b, KR80a, KR80b, KL89c, KM76, Kog55]. matrix [Kub63, KZ91, Laa58, LS90, Van75, Van77b, Van77c, Van78, Van79, Van84a, Van85a, Mao86, Mat90a, Mat91a, Mat91b, Mat69, MS79a, Mol82a, OSV79, dO69, Pai85a, PR69a, Pol85, Pot76b, Pow87a, Pul66, Roc81, Ruh69, Ruh70, SV79, Sch91, SW81, ST80, Smi68a, Ste90b, Ste91e, Tan75, Uhl85, VS88a, Vet75, Wat87, WA67, Wed78, WY64, Zie78, Zie84, Ziȩ83, Ziȩ87, Sch33, Spi76]. Matrix-Matrix [DS94]. Matrix-Multiplikation [Spi76]. Matrix-Sign-Function [GL86]. Matrizen [Gab79, Rut66]. Matter [Hon91]. Matthews [Pow87a]. Max [Bun71b, Kah87c]. Max-Norm [Bun71b]. Maximal [tBK84, Nod71]. maximaler [Gab79, Zie74]. maximization [MS86]. Maximum [Boh77, Chi93, Fan51, Pow67, Tao84b, Tor65]. Maybes [Ber77a]. McGraw [Lon93]. McGraw-Hill [Lon93]. Me [Lee88]. ME28 [Duf81]. Mead [DW87]. Mean [BR71b, BB84]. Meaning [BT64]. Meaningless [Fox71]. Means [Dek69]. Measure [Pry84, Ziv82]. Measurement [BF90, Ful87, Kla63, Sig86]. Measurements [HN81, Tay82]. Measures [Hig91e, Rhe66, Ebe65, EP87a]. Measuring [Bye88, KF90]. Mechanical [Inc92]. Mechanics [DDLT91, KR65]. Media [Mac77b]. Medical [CB81, Cri75, Kin67a, Kin67b, Kin67c, Kin67d, Rol67a, Rol67b, Rol75b, Rol75c, Rol75a, Rol75d, Rol75e, Rol75f, Rol76a, Rol76b, Rol76c, Rol76d, Rol76e]. Medicine [Dig91, Dud77]. Meet [Lee88, Mac95]. Meetings [Boo93, CB81]. Melodies [Ano78b]. Melody [Far87c]. Memoir [Far87a, Mic92b]. Memory [Bai89, Bul91, CDPW92, DEKV92, DDSv91, DvdG92, EHHR88, GJM87, GJMS88, GHL86, Hal87, HR88, KP92, Kre76, LC89, Lor74, LL74, RG92a, BDCS88, Cal86, Gay91, HH91a, Wri90]. Merit [Tom88]. Merriam [Mer93, Mer91]. Merriam-Webster [Mer93, Mer91]. Mertens’ [BFP88]. Mesh [BLV85, LZ91]. Mesh-Connected [BLV85]. Message [LC89, Ste90b]. Message-Passing [LC89]. Meta [Hof85a]. Meta-Font [Hof85a]. Metafont [Hof85a, Knu86e]. Metalgorithm [Ric75]. Metallic [Sch92b]. Metamagical [Hof85b]. Metamathematics [Hof85a]. Metaphysics [Hof85a]. Meteorology [ACD^{+}93]. Method [Abd71, AO86, AR85, ADNR92, Bar91d, BR71a, BD73, BD74b, BBP87, BS92b, BH83, Bjö87, Bjö92, Box65, Boy74, BD85, BH95a, Bun71a, Bye88, BHM94, Cap71, Cli73, CGO76, CGM85, Cro41, DM86, DV92a, Deu83, DM90, ES88, EO79, FTxx, FP63, Fos91, FM91, Gar90b, Gen69, GMSW90, Gol62, GNV79, GLO81, GS76b, Hac93, Ham82a, Hig86d, Hig92f, HP94c, Hig95f, HW79b, Inc85, Kåg93, Kar91, KL92, Lau79, Le85a, LN72, LC89, Van85c, Van86a, Lyn64, Mag60, MP74, MT94, MT95, MH57, Mil43, Moh85, MS79b, NM65, O’L93b, Ogr88, Oli77, Ort63, Pai73, PKPL91, PH72, Par64, Pol57, Pug92, RT92a, Saa80, SU84, SC90, She94, SS89, SG92, Sla92, Sor92, Ste83b, Str91, Tao84a, TST86, VA86, War76, Wat94, Woź77b, WH16, Ziv83, Bal80, Bar65]. method [BBK89, CL92, CM86, CGS83, Dan87, DK81, DGK80, FH60, Fre79, Gan85, GS78, Gay79, GMSW86a, GMS^{+}86, GMS88, GU77, Gow58, GR90, Gre85, Han87, HW78b, Hua75, Li91, Van84a, Van87b, Luk85, Mas91, MS84, NS90, Osb78, Ove83, PW79b, Pul66, Ruh68, SS74, Sor82, Ste67, Ste75a, Ste90d, SZ91b, Tan75, Wri82]. méthode [Ben24, PV74b]. Methoden [FF64]. Methodologies [MM82]. methodology [DSB88]. Methods [ABS84, Act70, AYB90, Ame77, AR92, AK90, Axe94, BM82, BV92b, Bar93a, Bar90a, Bar89, BBC^{+}94, BB88a, Ber91, BG73, Bjö90b, Bjö91c, Bjö96, Bol84, Box65, Box66, Bra95, BS92c, Bre76, Bro85, Bru95a, BT92, BE73b, BP71, Bun76, BK77, Bun85, BGBM93, Bur95, But87, BGN70, BL86, CHQZ88, Car89, CCKT83, CH94, Col77, CEN91, Cor92b, CH92c, DBA74, DR84, Dax90b, Dem87e, DS83, DW84, DEK95, Dub83, DER86, EMN88, EP87b, Ept80, ES89, FF63, Fad59, FGP91, Fik67, Fle80, Fle81b, Fle86, Fle87, FGS95, FW60, For67, FMM77, FM68b, Fra58, GS90a, vdG93, GM74b, GM74c, Gil74, GMSW84, GMSW91, Gil86, GF74, Gol65, GO96, GHMP74, GW73, Gra90, Grc89, Gre55, GR88, Gre72, Gur92, HY81, HW76a, HW76b, Ham73, Han93, Han95]. Methods [Hea84, Hen62b, Hen63, Hey87, Hig90h, Hig91c, HK93a, Hig96d, HS95, HP88, Hot43, Hou50, HB59, HR92, Hul67, IK66, KW89, KvSR90, KMN89, Kau83, Kel93, KL91b, KH90b, Kov70, KO68, KPS77, Lab57, Lam91, LYK89, Lin74, Lin92a, Lon84, Loo72, Mat93a, Mat95a, MD64, Met77, MP77, Moo79, Moo88, Mur72, Nas79, Nas90a, NP78, O’L93a, Oli75, Pai94, Par92a, PL82, Pow70, Pow81, Pow86a, Pre75, Qui94, RW67, Ric83, Ric92, RM67, Saa92, SB88, Sny66, Spe80, Ste73b, Ste75c, Swa72, Tao84b, Ton92, Tor92, Tra66, TT87, Wan91, Wat77, Wat80a, Wil55b, Wil60c, Wil61a, Wri92, Wri95, You54, You89a, You90, Bea79, BD76, Cyb86, DS79a, DT91, DFW90, Fle90, Fre90, FNO87, Gab87, GMS^{+}86, GMS88, GO88, GD89, HP84, JS77, Van83a]. methods [Osb65, Sch83b, Sha82b, Sor90, ST75, Ste73c]. Metric [Ash93, DM90, FH55]. Mexico [Hen82, Hen86]. MGS [Ple74]. Michael [Jef60, Jon70, Wil65d, Wil71c, Wil71d]. Michigan [Var90c]. Micro [Eva81]. Microcomputer [Hig86c].

Microcomputers [DW87, Wou87]. Microprocessors [Bai93e]. MicroTEX [Var86b]. Middle [Tho83]. Might [Kah90b]. Migrating [Ker93]. Mill [Woo92]. Miller [Mil43]. Million [QTD91]. Min [Kah87c]. Mind [Car58, Gol83, Hof85b]. Minimal [BIM90, Bau63a, Fre93, Kov77, Par92b, Löt84]. minimax [Ziȩ83]. Minimisation [Nas79, Nas90a]. Minimization [Bre73a, Buc94a, FM68a, FP63, Fle65, Mir58, NM65, Pow75, Str72, LM85, Ste67]. Minimizing [EHM94, Fan93, HK71, Ove83]. Minimum [CV86, Kas95, Wat92b, Cad73, Cad74]. Minor [Pre91, Syl51b]. minorants [Dah83]. Minors [FP62]. MINPACK [MSGH84]. minres [Luk85]. Miracle [Led91]. Mirroring [Gro95b]. Mirsky [GHS87]. Miscalculate [Mar94]. Miscellany [Bol86]. Misconvergence [Par90d]. Misdemeanours [Sto91]. Misleading [BG89]. Misquotes [BG89]. Miss [Ber71b]. Missile [Uni92]. Misteaks [Cip89]. Mixed [GK93b]. M’Lady [And91b]. Mode [SBKZ93]. Model [AB79, Bro81, CB90a, Far85, Hen80, Rob80, Sch84, SBKZ93, Tie70, BMW86, Mor90, Pai85b, Sor82, DGL91b, DCHH88b, DCDH90b]. Modeling [Ste78]. Models [BF90, Ful87, Hen66, HS66, MN89, SS76, All83, Boj84, GS74]. moderately [PH85]. moderately-priced [PH85]. Modern [Asp90, Ber65, BL77, Fow68, HW76a, Lab57, Lee81, Sch88b, Ste75d, Wil71a, Woe92, Wor89, Jen77a, Knu86d]. Modes [Cul85]. Modification [AGG88, FP74b, FM84, GE94, Pug92, Yip86, Sor82, Ste67]. Modifications [Fos91, Ham74, Pow69, PP89]. Modified [BP92a, EHHR88, FGM95, Hal77, MS79b, Oli77, RSD77, Sch93, SE90, AG88, Gol73a, OW90, VS88a]. Modifying [BZY93, GW92b, Ste74b]. Modular [BBB89]. modulus [VS88a]. moindres [Ben24]. Molecular [GHR93, WC53].

Moment [Wan91, Bul80]. Moments [Ede91a, DEG72]. Monde [LM66]. Monitoring [Bar86a, Bus71, ER74]. Monotone [FB91, Mat90a]. Monotonic [BSW61]. Monotonicity [IIM87]. Monotony [Kin67a]. Month [Bro75]. Monthly [Ewi94, Fla71]. Moore [CKW85, SS74]. Morrison [Yip86]. Moscow [Pet68]. Most [Boy94, Bus90, DKH86, Gar82a, Gar84, Gar90b, Gar90c, Gar90d, Gar90e, Gar91b, DE84]. Most-Cited [Bus90, Gar82a, Gar84, Gar90b, Gar90c, Gar90d, Gar90e, Gar91b]. Mother [Bry90]. Motion [SU84, SC70]. mouth [Ken85]. Movement [SW91, SW94]. Moyls [ST80]. MP [Bre78b, Car91, DH86, Pap92]. Mr [Fey85, Syl51a]. MS [AFJS87]. MS-DOS [AFJS87]. MTNS [KvSR90]. MTNS-89 [KvSR90]. Much [Kah87a]. Muir [Bel31]. Multi [Tor89, Woo85, Day83a]. multi-access [Day83a]. Multi-Directional [Tor89]. Multi-objective [Woo85]. Multicomputer [Ede94a]. Multidigit [KO63]. Multidimensional [GP70, Gow84, GB86]. Multidirectional [Tor91]. Multifrontal [BL95, DR83]. Multigrid [Bri87]. Multiple [BKP93, Bre76, BHY80, Bre78b, Bre78a, HH92b, Par92a, PV94, Sch86b, SF86, SG92, Smi91, SZ91a, Sun95c, GK83, Sch83b, VS88a, Wil68a]. Multiple-Precision [Bre76, Bre78b, Bre78a, Smi91]. Multiplication [Bai88c, BL80, BMSV92, Bre70a, Bre70b, BH74, CR76, CW82, CW87, DS94, Fam88, GFB85, Hig90e, Hig90g, HK71, Kah92c, KO63, Kni95, LPS92, MRK76, Mat95d, Moh85, Pan78, Pan84c, Tsa81a, Ura68, Win71]. Multiplications [GA94, Hig92f, HK71, Lad76, Yau93]. Multiplicative [Pry85]. Multiplicativity [GS83]. Multipliers [FJ91]. Multiplikation [Spi76]. Multiply [DHSS94, Kah91b, Pan84a].

Multiply-Adds [Kah91b]. Multiplying [Hig92f, Lad76, Vit68]. Multipoint [Ste94d]. Multiprecision [Bai93b, Bai93d, Bai94b, Bai93a]. Multiprocessing [Duf89a, DD90]. Multiprocessor [DEKV92, GL88, GHL86, LPS87, RG92a, Wri90]. Multiprocessors [EHHR88, HR88, LC89, DD90]. Multishift [BD89a, DG94, Dub91b]. Multistep [Qui94]. multitasking [DH86]. Multivariable [PEM81, DK83, DK84]. Multivariate [Che86, Gow84, Kri80, Mui82, Dwy67, Ham85, Rao80]. Music [Fox87b, Hur90, Gou86b, Lan89b]. Musician [Anoxxc, Hym86]. Musicians [Bal86, Lee88]. Must [Bar78, Bar82, Mad83]. mutation [KO91]. mutation-based [KO91]. My [Ano84, Ker84, Man80, Mic92b, Pre91]. Myocardial [Inv93]. Mysteries [Tre85]. Mysterious [Gea91]. Mystery [Mol93]. Mythical [Bro75]. Myths [HV94].

NA-NET [DR91]. nach [Läu61]. NAG [CH90d, Ham94, HP88, NAG90]. NAGWare [NAG92]. Naive [Hal74a]. Name [HR80]. NAS [BBB^{+}91]. NASA [Bai93c]. Nat. [Hig96b]. National [HGD90, CK81, Wil48a, Wil75c, Wil80, Woo58]. Natural [New81, RJ93]. Nature [LK76, WY64]. Navier [TTR92]. Navigation [BI77]. Near [Ano24, For59b, Kah70, KC82, Uhl81, Bou73a, Van85a, Ros69]. near-approximant [Bou73a]. Near-Characteristic [Uhl81]. Nearest [Dem87c, Dem90b, Dem92c, Hig88a, Phi77]. Nearly [CR86, Mey89, Phi87, Sha91, Ste81d, FG86, Kah75, MSS84, SZ91b]. Nearness [Hig85, Hig89d, Wat91d]. Necessary [HK71, Sch86d, Sha82c]. Necessity [Wun48]. Need [Kah81]. Needle [Kah89a, WT94]. Needle-Like [Kah89a]. Needs [Jac78]. Negative [Cha74b, Els76, Ham82b, KLJ89, KLP93, MS79b, Nod71]. Négatives [GK37].

Neighborhoods [Mos86]. neighbouring [Wil84b]. Neill [Nei82, Nei84]. Neither [Bry91]. Nelder [DW87]. Nelson [Ano86b]. nested [SD90]. NET [DR91]. Netherlands [AK90]. Netlib [BDG^{+}94b, Gro95a]. Networking [LR93, LaQ94]. Networks [Ch85, Ell94, QCM94]. neue [Spi76]. Neumann [Asp90, Deu83, Gol72, Kat60, O’C73, Ple76, Tan75, Tau63]. Neumann-type [Tan75]. Neural [Ch85, Ell94]. Never [BG89]. Neville [GP92, GP93]. newly [TE87]. newly-discovered [TE87]. News [FDR86, Ola93]. Newspaper [Wat89]. Newton [CGS83, GM74c, Ham82a, MS84, Rei90, SU84, Sor82, BHR94, BF88, BBP87, DS79a, DW84, Fle86, Fle89, FGM95, GMS^{+}86, Gre85, Gur92, Hig86d, KF63b, KF63c, KLW90, KL92, KLP93, KT85, MS79b, NS90, Osb78, PS91, PW79b, Sch83b, SC90, Tal91, Wer84]. Newton-Squaring [KLP93]. Newton-type [GM74c]. Next [Ano78c]. Nijmegen [AK90]. Nineteen [MV78]. ning [Gid86]. Ninth [Gra92, Mer83]. No [Kah88, Pre91]. nombre [Ben24]. Nomenclature [HR80, MG89]. Non [AS85, CES93, DS92, Els76, FP62, Fre93, GK37, Ham82b, Hen62a, Loo72, MS89, Nod71, vdS75a, Ste85a, Trexx, Bar65, Ebe65, Osb78, Wil86a, Ziȩ83, Jos91]. Non-Decimal [AS85]. Non-European [Jos91]. Non-Hermitian [Fre93, Ste85a]. Non-linear [Loo72, Bar65, Osb78, Ziȩ83]. Non-Negative [Els76, Ham82b, Nod71]. Non-Normal [Hen62a, vdS75a, Trexx]. non-normality [Ebe65]. Non-Positive [FP62]. non-review [Wil86a]. Non-Sexist [MS89]. Non-symmetric [DS92]. Noncommutative [Gil91a, Lad76]. Nonegative [BJW85]. Nonfiction [Zin88a, Zin90]. Nonlinear [AG90, BP81, Bar74, Buc94a, CGT92a, CGT92b, DS83, Dix72, Dra92, FM68a, GMSW86c, GS76b, LM90, Le85a, MMR78a, MMR78b, Man94, Rhe66, Spe80, TS75, Vav91, Woź77a, All83, BD76, Den84, DNR87, Fle81a, GS78, GMSW86b, GMSW87a, ITH88, Mar63, Mor90, Ruh73]. Nonnegative [Bak84, BP79, BP94a, FS80, Gra87, Mat90c, Min88, NP78, KM76]. Nonnegativity [Asn70]. Nonnormal [BCD95, KPJ82, KvW74]. Nonnormality [Cha92, Loi69, EP87a]. Nonorthogonal [DGM86]. Nonsingular [Hea77, JE95, Wil42]. Nonsingularity [PR93, Roh90]. Nonstandard [Cra79, Yoh79a]. Nonstiff [Cas89, SWD76]. Nonsymmetric [AA94, BD92, BDM93, BD93b, Bai94a, BD94, BDG94a, BS92c, DS93a, Hou58b, Jes93, Kas95, LZ92, LZC92, NRT92a, NRT92b, O’L93b, Ruh94, Saa84, SG92, Ton92, vdV92, BDM89, DFW90]. Nor [Bry91]. Norm [AY81, Bar91d, Bax91, Blu78, Bun71b, DKW82, Gau75a, Gem80, Hig90d, KH95, KZ91, Mat90c, PKPL91, Tao84a, Tao84b, And85, BBK89, Cad74, Des63a, Hig88c, Hig92b, MS86, Ruh69]. Norm-Preserving [DKW82]. Normal [BHK91, Cau64, Fos91, GH59, GJSW87, Hen62a, Hol92, Hol74b, Kit85, Lás94, MT94, Mir56a, Ruh87, vdS75a, Trexx, Chu91b, DK83, Gab87, Hal74b, KR80a, KR80b, Phi77, Put57, Rog76, Ruh75]. normales [Ben24]. Normality [Lee95, Ebe65]. Normed [SW62, LR81]. Norms [BF60, BSW61, BL88, Boy74, GS83, Gre93, GKL95, Mir58, Mir60, Neu92, Ruh94, SS62, Sta78, Sto62, Str72, Str75, Wat92c, Wat92e, Wat92f, You79, Zie88, Bau75, HM90a, OW85, Ove83, Rot84, Tsi90, Wat92g]. North [BCEP94]. Notation [Knu92, SZ80]. Notations [Caj29, AS83]. Note [Ano91, Bar86a, BV92a, Bar94, Ben24, Bun82, CS86, Car58, CH68, CG85, Cle55, Coh74a, CS84, Des63b, Ebe71a, Fan60, FH93, For59a, GW66, HG80, Ham74, Hor79, JSW83, Kie87, Mas94, Mil71b, Mir56a, Mol90b, New81, Nod71, Møl65a, Pet86, Rei71, Rie56, Sal60, Smo95, Ste85b, Ste92b, Sun95a, Syl40, Tau50, Tsa75, van77d, Wat94, Wic92, Wil72c, Yip86, BBVdH87, GMSW86b, Gow58, Van79, Ste79b]. Notes [Bar93c, FHW48, Ham82c, Hig96d, Ken93a, Ken93b, Kah73b, Kah92b, Kah95, Knu92, Tau50, Gle86]. Nothing [Kah87a]. Nouns [Rol67a]. Novel [SC90]. NP [GJ79, PR93, Roh93, RK95, Roh95]. NP-Completeness [GJ79]. NP-Hard [PR93, Roh93, RK95]. NP-Hardness [Roh95]. NPSOL [GMSW86c]. NSDTST [EP87c]. Nucleic [WC53]. Null [Bar93b, CS84, Fos86, GL94, GL95b, GMS^{+}85, Ste84b]. Number [And90a, Bar89, Boa79, BB87, Bre73b, Bro73, Bur80, CH94, CMSW79, CR83, Dem83, Ede92b, EHM94, GI88, GL81b, GEJ95, Ham87, Hig86b, Hig87c, HK71, Kin67a, Kin67b, Kin67c, Kin67d, KC73, MK85, PM88, Riv90, Rol67a, Rol67b, SW91, SA75, Syl51a, WWWS86, ZWS86, Hig83b]. Numbers [AR85, BDM93, Bei66, Ben38, Bir75, Cli90, Coo81, Dem87c, Dix83, Ede88, Ede89, Erh93, FL74, Geo78, GK92, GK93a, GK93b, GL80, Hig83a, Hig95a, Kah92d, KO63, Knu74b, Lem75a, LO82, Loi69, Mac90, MI81, Mol95, New81, O’L80, Pri91b, Roh89a, Ske89, vdS69, SW94, SW90, SZ91a, Tod68, Tur36, Wed85, Zie88, MZB87]. Numeric [BM75, Dav86, HK87, Mat70, Mol90b]. Numerica [act94, act95, Hig96c, Ise92]. Numerical [Abe88, Act70, Aga93, Ame77, Ano78a, Ano93, AMR88, ACD^{+}93, Atk89, Atk93, Bab69, Bab72, Bai88a, BP81, BM82, Bar79a, BL84, BG73, Bjö96, Bla64, BRR85, Bol84, BZ93, Bro85, BT92, BF92, Bun87, BGBMN90, BGBM93, BF89, BF93b, But87, Bye85, Bye86, Cas79, Cia89, CL90a, Cod73, Col84, CL90b, Col91, CGO76, CGT92b, CdB80, CEN91, Cor92b, Cow77, Cox72, Cox78, CH80, CH90c, DBA74, Dav81, DR84, Dem84, DDK92, Dem92d, DHvdV92, DL93, Dem93a, DHvdV93, Dem93b, DL94, DS83, DH95, DH90, DHW92, Dor71, Dou59, Dub83, Ede93b, Ede94a, Eir93, EWK90, Ept80, Erh93, Eva71, Eva74, Fei82, Fie86, FS76, For78, For59b, Fos79b, Fox64, FP68, Fox90a, FP92b, Fro69, GP70, Gas83, Gau69, Gau78b, Gea81, GS90a, GM74b, GMW91, Giv54, Gla93, GAKK93]. Numerical [GvN51, Gol77, Gol65, Gol73b, GO78, Gol84, GO96, Gra90, GR88, Gri84, GW86, GW90, GW92a, GW94, HH91b, Hag88, HW76a, Ham82b, HYG88, HHKR93, Ham73, Har65, HW76c, Har58, HW95, Hea84, Hel78, Hen82, Hen86, HD80, Hen64, Hig85, HS87, Hig94d, Hig95b, Hig96a, Hil74, HP88, HMW77c, HMW77a, HMW77b, Hou50, Hou53, Hou64, HR90, Hul67, Hul79, Hul82, HAC^{+}85, Ips94a, IK66, IP87b, Jac77, Jac78, JW77, Joh78, JR82, Jon62, JT74, dJ77, Kåg77b, KvSR90, Kah66b, KC82, Kah91c, KMN89, KF87, Kel93, KC91, Lam91, Lau85, Lau88b, Lax65, Lem76, LS91, Loo72, Low60, LM66, MD64, Mik91, Mil71a, Mil75b, Mil75c, Mil75a, Mil84, MW86, Mol78, Mur72, Nab92, Nas79, Nas90a, vNG47, NR72, Ort72, OVR90, Pap93a, Par90c, Pow70]. Numerical [PFTV86, PFTV88, PTVF92a, PTVF92b, RR82, RR78, Ric83, Ric92, Row90, Ruh83, Rut90, Saa92, Sal81, SSC94, Sca50, Sch68a, Sch88b, SRS73, SA73, Sha87, Sha94, Ske79, Ske80, Ske89, SK93, Sma90, Sny66, Sor90, Ste78, SB80b, SRD94, Stu80, SY78, Tau50, Tod62, Tod77, Tra73, Tre90, Tre92a, Tsa81b, Tsa83, Tur89, UvG89, Van83b, Vav92, Van77a, Vit69, Wat80a, Wat96, Wen66, Wes68, WR44, vW66, Wil71b, Wil74b, Wil77b, Wil78a, Woź77b, Woź77a, Wun48, YS89, YG72, YG73, Zaw91, Bur90, Bus71, Cyb80, Cyb84a, Dem88b, Ein79, Enr79, FT88, FR87, FNO87, GMSW86a, Gla82, GS74, GM77, Gro87, Han90c, KR80a, KR80b, LS90, Lin88, Pai81, Par78, Ruh70, ST80, SS74, DR91, Pri92, Sun92d, Thi88].

Numerically [Fos91, GM73, Gil74, Kah89c, Pai79]. Numerics [Bjö94, Com88, Com94]. Numérique [PV74b]. Numériques [PV74a, Bru89]. Numerische [FF64, KS88, Var90c]. Numeryczna [KS92]. NUMVEC [HL86]. Nyström [BGD^{+}89].

O.K. [FJ91]. Oak [HGD90]. Oberlech [PA88]. Obfuscation [Cri75]. Object [DPW94]. Object-Oriented [DPW94]. Objective [Pow86a, Woo85]. Obliczeń [KS92]. Oblique [Bro67, BK69, Cra74, Gre52, Luk84]. Obliterated [Gar90b]. Observations [BB92, Gau95, Har81b, Ing95, WR44, WH16]. Obtaining [FF51, Cad74]. Occasional [Kah89b]. Occurring [Ker69]. ODE [CC91, Enr79, GS90a, Pet82, Sha82a, SR95]. ODEs [dBK86, BGS87, Eir93, GSB88, HT93, SGB91]. Odor [Dem91]. OED [Sch92b, Sim87]. Off [Abd71, AB79, Cap71, Dem93b, Des63b, Dia78, Dou59, FP62, For59a, Gol62, GF75, MM70, Low60, Lyn64, Pry81, Qui94, Qui83, Tur48, VA86, Ziv82, BC86, CB90a, Cyb79, Bod49]. Off-Diagonal [FP62]. Office [Reg87]. Official [‘Vi52]. Old [Dic92, Hal91b]. Olga [Sch77]. One [Ark86, BBB89, Dav88, Dav72, EY36, FM85, Gau95, GWW92, Gor93a, GE94, Hig88c, Kac66, Ost54, PJS92a, Pow69, SC70, SB92, Syl51a, Ziv83, Gla82, GP90, KBS90]. one-day [Gla82]. One-norm [Hig88c]. One-Parametric [Ost54]. Ones [PM88, Ziv95]. Only [JW85a]. Onto [BD85, AL86]. Open [Dem92d]. Opening [Kin67d]. Operation [Ano46, Ham87]. Operations [Dem92b, Pan84b, Ric55, DMdB90, Olv83b]. Operator [Bha93, BR96, BN87b, Goo79, Grc89, HS81b, Jon62, KvSR90, RTP93, RT94, TTR92, Wat92e, Kel75]. Operators [Bou73a, Dav85, GL95b, Hol74b, Kat76, KH95, Kit85, Sch65, Ste71b, Trexx, Wou65, ASS73, Bea79, BHK91, Fan51, FHM67, Hal72, Hal74b, Kel75, Phi77, Put57, Putxx, RS88a, Rog76]. Optical [Shr84]. Optimal [AS93, BB95, BI75, Bau63a, BS92a, Bro68, Cha88, DKW82, GR89, LPBW87, Mao86, MS74a, MA87, Pan78, PKPL91, Sha85, Sha91, Str69, Stu82, Stu85a, Sun95c, WKS95, Wat91c, EH79]. Optimale [SB53]. Optimality [Har72]. Optimally [Bau63b, Bau69, Bus68, FL74, Gau75b, Sha82c]. Optimisation [Ber91, Cia89, Dix72, Gre80]. Optimization [Box65, Box66, Col84, CL90b, Col91, CGT92a, CGT92b, DS83, DS75, Dix76, Fle80, Fle81b, Fle87, FGS95, Fou81, GM74b, GM74c, GMW81, GMSW84, GMW91, GMPS92, Hig93c, IBM93, KO68, Loo72, MGH81a, MGH81b, MW93, Mur72, Noc92, Osb85, Pow70, Sto75b, Swa74, Tor92, Vav91, Woo85, DW87, Wri92, Den84, Fle85b, Fle90, Gay84, Pow76, Rot84]. Optimize [Lem76]. Optimizing [Wol89]. Optimum [Hul67, Vit68]. Optimzation [GMSW88]. Options [Kah80a]. Oral [Anh94, Fra92a]. Orbit [QT90, QTD91]. Orbital [Qui94]. Orbits [CKP95, HYG88]. Orchestra [Ano84, Ano86b]. Orchestral [Mil43]. Order [BMvN46, Bha93, EM95a, GvN51, Gol62, Mit48, vNG47, BGH62, Mat90a, Vrs86]. Ordering [RS88b]. Orderings [EGL^{+}87, SS89, Mas91]. Orders [tBK84]. Ordinary [AMR88, Bur95, But87, CKP95, Ept80, Gea81, HW91a, HW76a, Hen62b, Hul67, Lam91, SWD76, SW84, Sha94, Ste73b, AP81, BJW85, MZ91, HW91b]. Ordinateurs [Mul89, PV74a]. Ordination [Gow84]. Organ [Hur90, Sto75a]. Organization [Bow53, HVZ90, PH94, HH91a]. Oriented [DPW94, Cal86]. Origin [ABG48, Buc81, CT93, Tro84b]. Original [Anoxxc, Dud77]. originally [Ano78b]. Origins [Asp90, Dan84, Dan87, FF92, Mal89, Mic80, Ran75a, Ran75b]. Orthogonal [AOU87, BL95, tBK84, BB71, BP94c, Bro72, CR93, CL72, CH87, Cli66, EK89, For57a, Gal71, Gau83a, Gre52, Ham82c, Hig88b, Luk83, LQ87, Mir59, Rat82, Sch66, Smi65, Söd93, Ste80b, SB95, Wat93b, Wil62, BIM90, FG86, GR90, Kel75, Pai80, Par90b, Sch68b, SW87]. Orthogonal-Symmetric [CH87]. orthogonaler [Rut66]. Orthogonalisation [MBI77]. Orthogonality [BP92a, Bra95, ES93, KC82, ST91]. Orthogonalization [BIF69, BI75, BIM76, BIMF76, Bjö67b, Bjö94, CK57, Cyb83, Hof89b, Lon84, Ric66a, Ruh83, Sch73, Kub65, Swe84]. Orthogonalizing [Mat92a, Mat95b]. Orthogonaux [Aut02]. Orthonormal [CS84, Phi87]. Orthonormality [Kov70]. orthonormalization [Mao86]. Oscillations [LM90]. Oscillatoires [GK37]. Other [And90a, Box65, But92b, CD86, Day92, Gar85, Lem61, QCM94, Bal80, LHPW87]. Our [Bai87, Kah90c, Led87, Pot76a, Wat91a]. Outer [OP83]. outgoing [Bul80]. Outlandish [Dic92]. Outmoded [Ber71b]. Over-Relaxation [Lyn64]. Overcoming [CH89]. Overdetermined [Neu92, Bog74]. Overflow [FT86, MI81]. Overflow/Underflow [MI81]. Overflow/Underflow-Free [MI81]. Overhead [Gou72]. Overrelaxation [JR75]. Overview [BS89a, Col91, DPW94, SS92, You89a]. Owner [BH95b]. Oxford [Aug84, CBB54, Hig90j, Har83, Par92d, Tul91, Und80, Wei85].

P [Vav89]. P-Complete [Vav89]. P. [SG83a]. p754 [Kah83b, IEE81]. Package [Ano86a, Bai93d, BHY80, Bre78b, Bre78a, Cod87, Cod93, CGT92a, CGT92b, DGL91b, DS84, GWL^{+}92, GMSW86c, Han92, KL85, PdDKÜK83, Smi91, Spi90, Yoh79b, Cas91, DM84, GHM^{+}86, Nas85]. Packages [EP87b]. Padé [SV77, Hop82, KL89c, Van77b, KL93, War77]. PADRE2 [Kub91]. Paged [CNRT81]. Paging [Mol2a]. Pain [She94]. Painless [DGM86]. Pair [Kåg93, KP94]. Pairs [GZ93, KP93]. Pairwise [Bar91a, Sor85]. Pal [Com83]. Palindromes [Ber73]. Paper [Boo93, CT93, Day77, Day83b, Day88, Gar80, HB86, Kat85, Par93, Pyk85, TS80, Woo92, Ker83, Ste87b]. Papers [Abb78, Ano04, ACD^{+}93, CD86, For57b, FGHK83, Gar77, Gar84, Gar90b, Gar90c, Gar90d, Gar91b, Mic90, OW77, Ran75a, Ran75b, Sch86a, TS80, Trz93, Tur87, UvG89, Woo92]. Paradoxes [Kah90c]. Paragraphs [KP81]. Paralanguage [She93]. Parallel [AG89, APS93, AS93, BD92, BD93b, BD94, BDG94a, BBB^{+}91, Bai91, Bai93c, Bar93c, BS89a, BT89, Bre74, Bre91, Bul91, Bur95, CGR93, Car89, CK75, CKS78, CDW94, CL91, Csa76, DEKV92, Dat89, DM86, DHP94, DI86, DHvdV92, Dem92e, Dem92b, DHvdV93, DDR94, DDGM89, DS92, DT93, DS93a, DT94, Duf89b, Ede93b, EK92, Eva82, ED83, ES89, Fei82, FF87b, FP92b, GJM87, GPS90, GHN^{+}90, GL91a, vdG93, GHL86, Got92, GA94, Ham94, HR88, Hel74, Hel78, HP93, HP94a, HP94b, HP94c, Hig95e, HJ81, HJ88, HP92a, HS93b, JáJ92, JS92, Ken91b, KF90, Kog74, LF80, Lan93, LO82, LZS91, LZ91, MRK76, Mal93, Mat95d, Mod88, ML85, OW90, OV85, Ort88, OVR90, PR89, PKL90, Pap93b, PA88, PPY92, PA92, Qui87, Rob90, RO88, RG92a, RG92b, SB77]. Parallel [Sch88a, Sch88b, SHW86, SS89, Stp93, Swa93, Tor89, Tor92, Tra73, TW91b, Tsa91, Tsa92, Tur89, Wat92a, Wri90, Bis91, DT91, DS87, DSB88, DBKF90, GAG88, Gen78, PR85, PP89, Par90b, Ple86, SK77]. Parallelism [CH90d, Dem93b, Duf88, Rol76a]. Parallelizable [BF88, LTT92]. Parallelizing [HV94]. Parameter [All83, Bar74, Sch86b, Van87a, BIM90]. Parameterization [For78]. Parametern [Zie75]. Parameters [Cod88a, Sch84, Mar63]. Parametric [Ost54]. Paranoia [Kar85]. PARI} [Gou91b]. Park [HS93a]. Parker [Gri89]. Part [BD92, BD94, BM94, DAM88, Gau95, HW91b, LIA93, Mat92b, PJS92a, PJS92b, HS59, BD93b, Bus90, DK93a, DK93b, Gar73a, Gar73b, Gar90b, Gar90c, Gar90d, Gar90e, Gar91b, Gau95, SoIPUG87, Stu85b, Stu85c]. Partial [Ame77, BZY93, Bun74, CGO76, Dax82, FGM95, FW60, Fos94, GKO94, Hig95f, Liu87, Mac89b, OV85, Ske81, Vav89, van77d, Wri93, You54, BS83, ITH88, Lax65, Mac88, Osb65]. Partially [Bro72, Luk83]. Participants [Hig92c]. participatory [Ste87b]. Particular [HW76b, Ziȩ85, PW77]. Partitioned [AYB90, HP94c, BJ74]. Partitioning [PPY92, PPY93]. Parts [Alb91, Hal85]. Pascal [JW74, JW85b, Law72, Syl50, KKN^{+}92, BP92b, Coh75, FG84, Gol72, Gra92, HHKR93, Ker84]. PASCAL-XSC [KKN^{+}92, HHKR93]. Passenger [Bro89]. Passing [LC89, Ste90b]. Past [TW91b]. Path [HR90]. Patriot [Uni92]. Pattern [BK69, DR94, Gru70, Hof85b, Pic90, Tor93, GR89]. Patterns [Str89, BDCS88]. Paul [Alb91, EG91, Gil91b, Pri91b, SG83c, SG83b]. PC [CS87, Fri84, HK87, Hig89c, HP86b, MLB87a, PC-87, PH85, SW88, SR90, SoIPUG86, SoIPUG87, Sta88]. PC-BLAS [HK87]. PC-Matlab [MLB87a, SR90]. PCTEX [Var86b]. PDFIND [Cra86]. PDS [Tor92]. Peacock [Jos91]. Pearls [Ben85, Ben86, BKM86, Ben87a, Ben87b, Ben88]. Peculiar [Dav75, Rai69b]. Peer [Gar80]. Penalty [Gou86a, CH90b]. Pencil [DK93a, DK93b]. Pencils [Dem83, DK87, KR83, Ste94f]. Penguin [Bry87, DN89, III84]. Pennine [Wai85]. Penrose [SS74]. Pentium [Ede95, Hal95, Mar94, SB94]. People [AA85, AAR90, Lav80]. Pereyra [Hig87b]. Perfidious [Coh94, Wil84c]. Perform [GB86]. Performance [Bai91, Bai93d, BBG92, BF92, DDC^{+}87, DDK92, DKH86, DV92b, DW93, DPW94, DLN^{+}94, DPW95, Don95, DW95, Dow93, EP87c, Fos79b, IT89, JP94, KLV91, KF90, KLS^{+}94, Lyn82, MS93b, Pre93, RSD77, ABB^{+}90, ABD^{+}90, BD89b, Bis90a, DS86a, DS86b, DS88, Fri84, Fri85, HH91a, PH85]. Performing [Mil78, Wat83a]. Periodic [BG77, Eva71, EO79, LM90]. Perish [Hil87, Ros75]. Permutation [HS81b]. permutations [GP74]. Perplexed [Mos71a]. Perron [EKNX93, Kee93]. Personal [Cri74, Cri89, Kah79, Kra93, Mil87a, Sch77, She93, Wat80b, Hil89]. Perspective [Kra93, Pap93a]. Perspectives [TW92a, Wat93a]. Perturbation [Bar90b, Bar91b, Bar92a, Bar94, Bha93, Bjö91a, Bol85, CP95, CS89b, DK70, DR94, Dor71, DOV94, EI93, Eld80, EP94b, EH91b, FM91, GZ93, GE93, HH92b, Hig93d, Hig94d, JE95, Kåg94, Kat76, LPP93, Li95, Mat93b, O’C93, Löt83, Roh94, Sch92a, Söd93, Ste73d, Ste77b, Ste77a, SS90b, Ste90e, Ste90f, Ste90g, Ste93a, Ste93c, Ste94f, Stu81, Sun91, Sun92a, Sun92b, Sun95c, Sun95b, WKS95, Wat88a, Wed72, Wed73b, Wed85, Wei90, Zha93a, Dem87a, Kåg77a, Ste75b, Ste79b, Ste91d]. Perturbations [Bha94, CR86, Erh93, MP83, MH87, vdS75a, Sun95a, Var90a, VS93, Wed78, AG88, DK84, Wat88b]. Perturbed [Gre89, BD76, Sch91, Ste84b]. PFORT [Ryd74]. Phases [PEM81]. PhD [PP94]. Phenomenon [Hil95, TT87]. Phillips [Bag94]. Photographic [Hal87]. Phrases [Pét66]. Phys [Cau41]. Physical [Bev69, Bus90, CK81, Gar90e, Tay82, Wil48a, Wil75c, Wil80, Woo58]. Physical-Sciences [Bus90, Gar90e]. Physics [Dys93, Har84b, Kit71, Rut47, Rut52, Gar77]. Pi [BB87, BBB89]. Piano [Ano78b, Lev89, Wat83a]. Picard [Han90a]. Pictures [Cip88]. Piecewise [Cox86]. Pilot [Asp81, CK81, Wil54a, Wil54b, Wil75c, Wil80]. Pinyin [Hol80]. Pipeline [DGK84]. Pitfalls [For70, SA56]. Pitman [Hig96d]. Pivot [Coh74a, Cry68, Ste74b, Tor65]. Pivoting [Bjö92, Bun71a, Bun74, CG85, DK77, Dax82, Ede92a, EM95a, For60, Fos94, FP81, GKO94, Gou91a, HH89, Hig95f, Liu87, Mac89b, ON93, Peñ94, PW75, Ske81, vdS70a, Sor85, Tha61, Vav89, van77d, Wri93, YC95, Bis89a, Bis91, GN82, Mac88]. Pivots [Cha84]. place [Str87]. Places [Leh43, Lod75]. Plagiarism [Mal89]. Plain [Cla92, Fle46, Gow73, Gow86, Rob88]. Plan [Ros90]. Plane [AP94, Giv58, Ham74, TTR92, Wil63a]. Plane-Rotations [Wil63a]. Planetary [QT90]. Play [Ste89, Sto75a]. Please [Ano78c, Tur91b]. Plus [Anoxxa]. Pocket [Com83, Leh94a, Smi75]. poem [TE87]. Poets [Lee81]. Point [Ano86a, AL86, BSBF89, Bar81a, Bar81b, BB85b, BB85a, BZ90, Bar89, BB88a, BGM^{+}94, Bre73b, Bro81, Car59, Car91, CO84, Cli90, Cod73, Cod81, CCG^{+}84, Cod88b, CC93, Dek71, Dem89b, Dem92e, DDR94, Esp78, Far88b, Fat82, FG82, FT86, GB91, GM74a, Geo78, Gol91a, GFB85, Gre72, IEE81, Hig91e, Hig93a, Hou81, IEE85, IEE87, JSW83, KP79, Kah80a, Kah80b, Kah81, Kah83b, Kah86a, Kah86b, Kah87b, Kah92a, Kah95, KL73, Kar85, Kel93, KPS77, KC73, LIA93, LO82, Lin74, Lin75, Lin81, Lin70, Lon67, Mal71, Mal72, MI81, Mat70, MS76, Mol67, Møl65b, Pla88, Pri91a, Pri92, RK75, RS88b, Sch81a, Sch84, SB94, Smi91, SW90, Ste74a, Stu82, Swe65, Tan89, Tan90b, Tan92, Tie70, Ura68, VS93, Vit68, Wal90, Wic89, Wic92, Wil60b, Wil63a, Wri95]. Point [Yoh73, Boh77, GMS88, Rum91, Tsa74a, Wil88b]. Points [AHB87, Oli75, Rei90, SW91, Ste75a, TW91a]. Poiseuille [TTR92]. Poisson [Dor70, BGN70, Buz92]. Polar [AMS74, Bar90b, Cai84, CS89b, CH87, Dub94, Gan90, Hea67, Hig86a, HS90, HP93, Hig94c, HP94b, KL91a, KL92, Li95, Mat93b, Uhl81, Wil42, Hau67]. pole [BN87a, CM89]. Polish [Rej81]. Poly [BBD^{+}94]. Poly-Iterative [BBD^{+}94]. Polyhedral [Wat92f]. Polynomial [Ada67, BP94b, Coh94, EK93, EM94, EM95b, Fik67, HPW90, HB59, JT75, Kah86c, Kur92, LN92, Mos86, New74, Oli79, Olv86, PW71, Pow67, Rei91, Ric65, SR80, Smi65, Sor92, Ste71a, Tal91, TL81, Uhl81, Uhl92a, Wer84, Wil84c, Woź74, Ziv83, Fre79, Ste69, TW91a]. Polynomials [BGH92, CR93, EK94, EK89, ET73, For57a, FP68, Gal71, Gau78b, Gau83a, Goe94, Hig88b, Hig96b, Kit95, Mil71a, Pai94, Pan92, RR82, Riv90, TL81, TT94, Tra66, Wat94, Wil65b, FR87, Knu62, Van79]. Polynoms [Hil94]. Pompous [Gre92]. Popular [Wil72a]. PORT [FHS78]. Portability [Cow77]. Portable [Bai93d, Blu78, Cod87, Cod93, DDK92, DHSS94, Kah93b, Wal82, Yoh79b, Yoh80, ABB^{+}90, ABD^{+}90]. Portrait [CES93, MT94, MT95]. Portraits [Bal86, GKK90, Sla87]. Posed [Dem87c, DK88, Han92, Han90a, Han90c]. Position [For92, LS82, LS83]. Positive [All88, AMS74, And87, AR85, Bak84, Bar86b, BR71a, Bax91, dBP77, Bou73b, Cry73, Cry76, CV86, Doy57, EHM94, FP62, Fra89, GL81a, GV79, Hal72, Hig88a, HW79b, Kem82, KLJ89, Ker90, Mak91, MPW66, Mat92b, Peñ95, Pin85, SS91, SH93, Tau67, Tau70, ASS73, And85, Bel68, Bou73a, CM83, Cra86, DI86, FG86, HW88, HM90b, MM88, OSV79, Pul66, WY64]. Positive-Definite [Doy57, SS91, Tau67, Tau70]. Positively [Gou73]. Positivity [Bau75, GP92, GP93, Kar68]. Possible [Sto62, MM88]. Poster [Mat90d]. Posteriori [Auc92, Dei90, Tsa74a, OW82].

PostScript [Wic93]. Power [Boy74, Hop82, KW92, Lor74, Tuf88]. PowerPC [Com94]. Powers [Bab37, FS80, Gau53a, Gau53b, HK95, Kor88a, MP77, You79, Cha74b, Coh74b, KZ91, WA67]. practicable [Zie78]. Practical [Bak85b, Bee92, CB81, CKS78, Fle80, Fle81b, Fle87, GMW81, HW76b, Hig90g, Jud90, LPS92, Lee95, MBI77, ML85, PW71, Sch79, SW91, Swa80a, Tar92, Trz93]. Practice [BB88b, Cod88b, Dav75, FvDFH90, Fos94, HP88, KM81, MH80, Und80, Wil72b]. pre [Osb60, Sie86]. pre-conditioning [Osb60]. pre-TEX [Sie86]. Precise [Abe88, AS92, Bri91]. Precision [BKP93, Bar88b, Bre73b, Bre76, BHY80, Bre78b, Bre78a, Bur74, CCF96, Cha66, CF93, CFB93, CHH83, Dek71, Dem87d, GS92, Gre94, Hig91e, HK93b, HK95, Hul82, JW85a, KLV91, Kah87b, Kah89b, Kie81, Lin81, MK85, Mil88a, Møl65a, Møl65b, Pri91a, Smi91, Wil77c]. Precompiler [Cra79, Kub91]. Preconditioned [AK90, Bru95a, CNP94, Hig96d, HS95, PS92, Ton92, vdV90]. Preconditioner [Cha88, CS89a]. Preconditioners [GMPS92, Sch93, GR89, Vas90]. Preconditioning [CGM85, EH79]. Preconditionings [CDG91, CM86]. Predicting [GS92]. prediction [Cyb81, Cyb84a]. Prefix [Dem92e, Dem92b, EK92, LF80, Mat95d, MA87]. Preliminary [TTR92]. Prelude [Saw55]. Prentice [Pre78]. Preparation [CH88a, KD93, KD95, Lam86b, Lam94, WWG51, Ack89]. Preparing [Mac77c]. Prepositions [Kin67b]. preprocessor [Ker75]. prescribe [Kah90a]. Prescribed [Mat85, Hor54]. Presence [Fle88, Neu85, Not93, SB88, vdSvdV87, vdV90, DK84, She93]. Present [Par93, TW91b, Woo58]. Presentation [Anh94, Dud77, Fra92a, GP59]. Presentations [Byr87, Woe92]. Presenting [FGHK83, Gar80]. Preserving [DKW82, BH91]. Prespecified [DK85]. Press [CBB54, Har83, Hig90i, Hig90j, Hig96c]. Presubstitution [Kah87d]. Prettyprinting [Bee93]. priced [PH85]. Primer [Cla92, Sig92, Sig94]. Principal [FP62, PEM81, TSY88, TST86, DGK80, FH60]. Principles [BB49, Hou53, JT75, SY78, FvDFH90]. Print [Hay88, SW90]. Printing [BK88, CBB54, Gou86b, Van87d]. Priori [Dei90, Wil68b]. Private [Don88b]. Prize [Abb78]. Prize-Winning [Abb78]. Pro [MLB87b]. Pro-Matlab [MLB87b]. Probabilistic [Bar81b, BB85b, CB90a, Hen66, HS66, YC95]. probabilities [GM86]. Probability [Bor94, Fos90, Hal90a, Dem88b]. Problem [AA94, AGG88, Aru92, Bar93a, Bar93c, Boy94, CB90a, Dem87c, DS92, DvdG92, DS93a, Eld80, EH91b, Eva71, GS90b, Gol83, GNV79, Gul95, Hal67b, Hal82a, HV94, Hig88d, Inc83, Inc84, JW86, Jes93, Kah71, Kar74, Lam91, LZS91, LPS87, Lu95, Mac95, Mio90, Mol90a, Nas84, O’C73, Ort63, Par64, Par80, Löt83, Löt84, PW70, Ple76, PC83, Rai76, Roe86, Sau78, Sch66, Sch81b, Söd93, Sun95a, Sun95c, Uni92, Van87a, VV91, Wah65, WKS95, WE91, Wat93a, Wat83b, WT94, Wei90, Wil65a, Wil70, Wil78a, Yau93, Zha93b, AP81, Bul80, Dem88b, DS87, Fle81a, Fle84, GGM89, GV80, Kåg84, Kub62, Kub70, MS86, OW85, Par90b, Roc81, Ruh73, Ste75b, SE87, Wed73a, Wed78, Wil77a, Woo89]. Problem-Solving [JW86]. Problems [Abd71, Abd74, AE93, ADdR89, AMR88, Bab72, Bai94a, BP81, BR77, BI87, BH88, Bar88a, BV92a, BJ89, BS89a, BL84, Bjö67b, Bjö87, Bjö96, Bol85, BM75, Box66, BD86, Cas89, Cli73, CW86, DD86, DK88, Dem92d, Dor71, DGL89, EI93, GL94, GH93, Gay85, Gen74, Gil86, Gol65, GKS76, GJ65, GDD94, Hal91b, HHKR93, Han92, Hea84, Hig85, Hig89d, Hig90c, HW79a, Hof89a, Ips94a, JW85a, Jen83, Kau83, Kel93, Ker89, Kni93, Kog74, KO68, LH74, LS91, LNOP82, LZ92, LZC92, Van84b, Van85c, Lus89, Mal92a, Mal92b, Mal93, Man81, MM82, Mol78, Pap93b, Par65, PW71, Pin85, PV94, PR68, PR69b, Qui83, RB85, RM67, Roh95, Rut58, Saa84, Saa92, SS92, SSC94, Sch86c, SW81, Sco81, SGB91, vdS75b, Ste73d, Ste77a, Tod61, Tra66, Van91]. Problems [Var90d, Wat91d, Wat92g, Wat93b, Wed85, Wil87a, Wol65, Woo85, Wri93, Bjö88, BG87, BBK89, BGS87, CL92, Cyb84a, Cyb86, DD88, Dem87a, DM84, FNO87, FU69, Gol73a, Han90a, Han90b, Han90c, Li91, Mor90, Osb78, Pai79, Rot84, Sch91, Sch68b, Sor90, VV87, Wat90, Lar83]. Proc [HW76c, Jam74, Pet68]. Procédé [Ben24]. Procedure [GH59, KLW89, Cad74]. Procedures [CK57, FP74a, Jor68, Pei62, SY78]. Proceedings [Ano48, AK90, DS79b, Gla82, GO78, GW90, GW92a, GW94, HPP88, KR83, KvSR90, MT86, PA88, MMR78b, Maz76, BCEP94]. Procesach [Wil67a]. Process [BKW90, Bro73, Dax90c, Gea91, Gil51, Rol76b, Kub65]. Processes [BIM76, BI77, BR71b, Dax90a, Des63b, HYG88, Mik91, Pai94, Rad48, SS76, Tur48, Wil63b, Wil68b, Kub63, DLP79, Bod49]. Processing [Aru92, BR87, CH88a, Cyb83, Dil93, Eva82, Hay89, KvSR90, ML85, Sal89, SF86, Tur89, Van91, van86b, APP88, AP76, Cyb78, Van89, MS79a]. Processor [BKP93, CK75, GJM87, HK87, KF90, Luk86, MRK76, Phi87, SB94, Zin83, Lam87b]. Processors [BLV85, EK92, IT89, LZ91, MA87, BV86, Bis89b, DSB88, DBKF90, DGR79, GAG88, Gen78, Van88]. Procrustes [AE93, Bro67, Cra74, Gow75, Gru70, Hig88d, Luk84, Par90b, Sch66, Sch68b, Söd93, Wat93b]. Product [GA94, HPS83, Neu85, OP83, Tsa81a, Win68, HLPW86, Smi68a, Wee89]. Production [Com83, GP59, HD80, MM82, Lam87a]. Products [Art83, Art86, BV87, Cup84, DL92a, GF75, HS81b, Mat92a, Mat95b, SV89, Ste94b, VP93, Hal91a, HM90a, Pol85]. Professional [Ano83, EBR87, Hym86]. Professionals [Day92, Kat85]. Professions [Bee92]. Professor [Syl51a]. Profilers [Ben87b]. Profiles [Fre87a, Fre87b, Lyn82, AA85]. Profit [GS66]. Program [Ros90, AJ88, BKM86, BBG92, Blu78, Dek82, Hor79, Kah88, Knu86b, Knu86a, Pax91, Pet80b, Ste91a, Ack89, App85, Ber80, Nas85, Ock75, PNO89, RS81]. Programme [Nei82]. Programmer [BGA90, Kah93b]. Programmers [Lam86a]. Programming [AKW88, BM79, Ben85, Ben86, BKM86, Ben87a, Ben87b, Ben88, CK80a, CK80b, CK81, Dan84, EJ86, ES92, EPL94, FG84, FM68a, Fle93, Gay85, Gil74, GMSW86c, GMSW90, GMSW91, Gra90, Gur93, HH88, HJ81, HJ88, HC88, MM70, Ive62, Ken91a, Ken91b, KC82, KT91, KK80, KP74, KP78, Ker84, Knu81a, Knuta, Knu74a, Knu81b, Knu84b, Knu93, Loe74, Lus89, MMR78a, Man94, MD64, MH80, Pla93, Ros72, Sau72, Thi86, Van87d, Wal82, Wal90, Wex81, Wol64, Wol65, AS83, DSB88, DNR87, Fle81a, FH90, GMSW86a, GMSW86b, GMS^{+}86, GHM^{+}86, GMSW87a, GMS88, MMR78b, RM91, RJLK78, RR73]. Programs [AwJS85, Bai93a, Bai93b, Ben82, CH88b, DCHH88b, DCDH90b, Gea69, HHKR93, JT75, Knu71, Lon67, NT58, Pet80a, Pet80b, WWG51, BDCS88, Fos79a, Har84a, DLP79, Sed78]. Progress [BB49, Hal90b, Pan84b, Par78, PR81, Wil48b]. Progressions [CW87]. Project [HS83, Ost83, Sim87, BD89b, KY88, MSGH84, RM91]. projected [GS78, GMS^{+}86]. Projection [AL86, ADNR92, BS92c, GF74, Hal80, MS78c].

Projection-Recurrence [Hal80]. Projections [BD85, Dem77, O’L90, Ste77a, Tur91a]. projective [GMS^{+}86]. Projector [Gou72]. Projects [Gra90]. Prolate [Var93]. Prologue [Gol90]. Proof [Car58, Har91b, Lam95, PS94, Yuv78]. Proof-Correction [Car58]. Proofreading [Har75, Swa79]. Proofs [Ler83, MS91, DLP79, ST80]. Propagated [Att57]. Propagation [CB90a, Hen63, Hen66, Hen80, HS66, Lau75, Low60, Oli79, Cyb79]. Properties [AN63, CGR90, BD87, BJ92, Bus69b, Cai84, Cry76, FU69, GM74a, GMS^{+}85, Mal72, Men56, Pai81, Pla88, Pri92, STY83, Ziȩ87, Fan51, Gay79, Li88, MM88, SS74, Wri82, Ziȩ83, MG91]. Property [Cad88]. Proposal [DCH^{+}95, DMR92, Str86b, Tur45, MG89]. Proposals [Cod81]. Proposed [CCG^{+}84, DCHP90, Fat82, IEE81, Hou81, KP79, Kah80a, Kah80b, Kah83b]. Prosaic [Mer90]. Prose [GH43, Gre92, Lan92, Rol75b]. Prospective [Har84b]. Prospects [KU68, Bai87]. Prospectus [ABD^{+}90, DDC^{+}87]. Proven [Woe92]. provenant [Ben24]. Provisional [BDDC88]. proximinal [Rog76]. Pseudo [Abd74, GK65, PW70, RT92b, Ste77a, Wed73b, KT85, Wed78]. Pseudo-Eigenvalues [RT92b]. Pseudo-Inverse [GK65, Wed78]. Pseudo-Inverses [Abd74, PW70, Ste77a, Wed73b, KT85]. Pseudoinverses [O’L90, DS80]. Pseudospectra [BH95a, GT93, Lui96, RT94, TT94, TT96, TTR92, Tre92b, Trexx]. Pseudospectrum [Brü95b, RTP93, Ruh94]. Pseudozeros [TT94]. Psychic [Kor88a]. Public [Bra83, Gli89, Ken82]. Publication [CT93, EV85, Gea91, Rod92, Rol76b]. Publications [Bus90, EBR87, Gar82a, Gar90e, McC84]. Publish [Day83b, Day88, Hal75, Hil87, Mic90, Rol75f, Ros75]. Published [Gar77, Jef60, Sch86d, Woo92]. publisher [Bud87]. Publishers [But92a, Man80]. Publishing [Har86b, Jac86, Mic92a, Mil87a, Sie87]. Punch [Ver49]. Punctuation [Ber77b, Car58, Gor93b]. punishment [Ste87b]. Puns [Hau91]. Punter [DF87]. Pure [Gar73a, Gar82a, Har67, Bri92, Gar82b]. Purpose [Pla93]. Purposes [Cle59]. Pursuit [Cri89]. Putting [Gre87]. Puzzle [Ede95]. Puzzles [Gar61]. Pythagorean [Ant85, Dub83, Inc85, MM83a, Sch92a].

qd [FP92a]. QL [GD89]. QR [AG92, BD89a, Bis88b, Bis89a, Bis91, BH91, Cha87, CH90a, CH92b, CI91, CI95a, CDO^{+}94, DGKS76, Dem92a, Dub91b, DG94, vdG93, Gen74, GM86, Gra86, GD89, GW92b, Hen92, HV94, HP91, Van88, MS93a, Mat92d, OW90, Pai87, PL93, Rei91, SB92, Ste84a, Ste92a, Ste93c, Ste94b, Sun95b, War76, Wat82, Wri90]. QR-Decomposition [GW92b]. QR-factorizations [BH91, CH90a]. QR-like [AG92]. QR/QL [GD89]. QUADPACK [PdDKÜK83]. Quadraten [Läu61]. Quadratic [BD74b, CH88b, Doy57, Fle93, For69a, For69b, GMSW90, GMSW91, LN72, Pow86a, Roh95, SW81, Syl51b, Tao84b, AL73, BJW85, Dav81, GMSW86a, GHM^{+}86, PV86, Ruh68, Wil84b]. Quadratically [Fan93, CL92, Ove83]. Quadrature [BD87, Gau83b, Gus70, Kah84, LK76, Lyn83, Lyn87, Ric75, Lau75, Lau88a, ST75, Wri82]. Quadratures [KE82, GK83]. Qualitative [CF93, CFB93, Lef78]. Quantitative [HP90, Tuf83, Lef78, Tho92]. Quantum [GL91b]. Quartic [Leh94a, Sal60]. Quartically [Bai88b]. Quasi [BW93, DW84, Fle86, Fre93, GSS96, Gur92, Møl65a, Møl65b, Sch83b, Van95, DS79a, Fle89, Gre85, KT85]. quasi-Choleski [Gre85]. Quasi-Definite [GSS96, Van95]. Quasi-Minimal [Fre93]. Quasi-Newton [DW84, Fle86, Gur92, Sch83b, DS79a, Fle89, KT85]. quasi-quasi-Newton [Gre85]. Quasi-Toeplitz [BW93]. quasiaccretive [MN79]. Quaternions [Mac95]. Queen [Bei66]. Quest [Ras86]. Questing [Hof85b]. Questions [Gau78b]. Queuing [Kau83]. Quicksort [Sed78]. Quiet [Mac77d]. Quincy [Ano84]. Quotations [Bar92b, Bra93, Mac77d, Par92d]. Quotes [BG89].

R [SG83a, Uhl92b]. Radii [HP86a, KL88, Lew92]. Radio [bGBB87]. Radius [BFP88, Els76, Gem86, HW95, Wat96]. Radix [CCG^{+}84, IEE87]. Radix- [CCG^{+}84]. Radix-Independent [IEE87]. Raleigh [BCEP94]. Ramanujan [BBB89]. ramp [Lag74]. Random [AOU87, And90a, And90b, Ede88, Ede89, Ede91a, Ede93a, EK93, EKS94, EK94, Gem80, Gem86, Hou89, KW92, PM88, Sch89b, Ste80b, Web73, Web82, 87]. Randomized [Inv93]. Range [AS92, DJ74, FT88, LS90, ST80]. Ranges [FS76]. Rank [AGG88, BV92b, BH91, CR86, Cha87, CH92b, CI91, EY36, FM85, Fos86, GKS76, GE94, HP91, Man81, PE95, Pow69, SB92, Ste84c, Ste93d, Wat88a, Yip86, AGS91, AG88, CH90a, Dem87a, DS80, Han90c, KBS90, Ste91d, Wat88b, Wed73a]. Rank- [Yip86]. Rank-1 [CR86]. rank-deficiency [Wat88b]. rank-deficient [DS80]. Rank-One [GE94, SB92]. Rank-Revealing [CI91, HP91, PE95, Ste93d, BH91, AGS91, Ste91d]. Ranked [Gar91b]. Ranking [Kee93]. ranks [KKM79]. Rapid [Shr85]. Rapidly [FP63, Le85b]. Rate [Cli72, Deu83, Not93, Str91, MZB87]. Rates [DMS84]. RATFOR [Ker75]. Rational [CGR93, CH89, EK89, Kah86b, KL91b, MK85, Ric65, Ruh94, SV77, Ziv83, CGS83, Ker75, Vrs86]. Rationale [CC91]. Rationals [ADW84]. Ratios [Sto62]. Ravages [Mal89]. Rayleigh [MSS84, Tan94]. Re [LSV95]. Re-started [LSV95]. Read [Cli90]. Readability [Kla63, McC87]. Reader [GH43]. Readers [CBB54, Har83]. Readings [BN71]. Real [AA94, Ano88, AR85, Ash93, BD93c, BD73, BMPW66, CJ70, Cro41, Den81, DLR85, EK93, EKS94, Ehr70, Giv54, Ham87, Hig87a, Hig88c, Hig92f, Hir65, JP93, Kåg93, Kah73a, Kah86d, Kur92, LTT92, Lew92, LZC92, MGM93, Str91, Tur91b, Chu91a, Chu91b, Cra86, Cul66, Fre79, Gea77]. Real-Time [MGM93, Gea77]. Realistic [Bro81, Dei90, Wat87]. Reality [HV94]. Realizations [Par92b]. Really [Dav72]. Reasonably [Yoh79b]. Reasons [Ano86b]. Recapture [BP92a]. Rechenanlagen [SB53]. Rechengenauigkeit [SB53]. Recipes [Knu85b, PFTV86, PFTV88, PTVF92a, PTVF92b, Sha87]. Recommendation [Wil84a]. Recommendations [CJLP93, CGL83]. Record [Bai87]. Recreational [Ede95]. Recreations [Bei66]. rectangle [Dor70]. rectangulaires [FKF68b]. Rectangular [FKF68a, Kni95, Sau94, Ste94f]. Recurrence [Act74, CK75, Gau67, Hal80, Kog74, Par76, Stp93, vdVD89]. Recurrences [Gre89]. Recurring [Var76b]. Recursions [Cas79]. Recursive [FGP91, APP88]. Redheffer [BJ92, RB89]. Reducible [Mey89]. Reducing [Bus69a, Kah65, Lan93, Ros65, Wol64, Yau93, Zha92]. Reduction [Aas71, AM87, AM94, Bev69, BG94, DHS89, DvdG92, Fer94, Gei91, Fer95, Par92b, Rob80, Tor65, Wil59, Chu91a, Dub91a, GLW89, Rao80, CDW94]. Réels [Aut02]. Reevaluating [Gus88]. Referee [For80, Man75, Pyk85, Smi90, Tho83]. Refereeing [CH68, Lin84a, Gle86]. Referees [Jon74, Par90a]. Reference [CGG^{+}91, HW76b, IBM88, Cra89, KKN^{+}92, Lam86b, Lam94, Mac90, NAG92, Sal93, Swa88, WWG51]. References [Boa66, Rol76e]. Refined [Mei83]. Refinement [Bjö67a, BG67, Bjö68, Bjö78, Bjö90a, GL95a, Gul94, Hig90h, Hig91c, Hig95c, JW77, Kie81, MPW66, Mol67, Ske80, SS84, Zla82, DCHS88, GW66, GP90, KLW90, MSS84, Zie78]. Refining [Dem87e, Ste91d]. Reflections [Cup84, Par71, Gol90]. Reflectors [SP88]. reflexive [Tan75]. Refutation [Kah91a]. region [CH90b, Gay84, MS83, Sor82]. Regions [LS91]. Regression [BRB76, BKW80, CR88, DC77, Gen74, Hig90c, Seb77, Man82, Ste87a]. Regular [Kåg93, KP93, KP94, Ple76]. Regularization [Han89, Han92, Han93, Han95, Han87, Han90b]. Rehabilitation [DH89]. Reid [Fre87a]. Rejections [Ber90]. Related [BFP88, BF93a, BBdHS95, Bro85, EBR87, Gau78b, Lyn87, Pap93b, RB78, BBK89, Pai81, Wat90]. Relation [Eni73, Hig94c, PC83, Syl51b]. Relations [AR85, Gau67, Han90b, vdVD89, Act74]. Relative [Dem87d, EI93, GS90b, GFB85, GE93, JE95, LS80, LPW83, LPP93, Pry84, Ziv82]. Relaxation [CM69, For51, Lyn64]. Relay [Frö53]. Release [CGT92a, DGL92, IBM88, Ken93a, Ken93b, AD89, Ano86a, CGT92b]. Released [Hig92e]. Reliability [Dem84, Fra92b, HHK93, Hor79, Moo88]. Reliable [Bjö90a, CH90c, SGB91, BGS87, Enr79, FH90]. Remark [Goe94, Hig91a, Tau49, Ura68]. Remarks [Bau69, Goo92, Kah65, Lyn82, Sma90, Zie88]. Reminiscences [Dan84, Var90c]. Removable [Kah89c]. Reordering [Kåg93, PA92]. Reorthogonalization [DGKS76]. Repeated [BS92b]. Replacing [MM83a]. Report [Bue77, CD86, ILA93, Gra70, JW74, JW85b, TTR92, Wil48b, Zie86, JBNP91].

Reporting

[BR93, FDR86, HP77, CDM79, JBNP91]. Reports [Coo64, EBR87, Mic90, Ste91b]. Repository [Gro95b]. Represent [HYG88]. Representation [BV87, Buc59, Cox78, Geo78, GHR67, GMS^{+}85, Ham87, MI81, PW78, Pug92, Rin64, SV89, SS76, SB95, Wil42, Wou65, Pai85a]. Reprint [For59a]. Require [Pow75]. Requirements [Kre76]. Research [Duf84a, Fra92a, Gre79, HH93, Hig96d, HS83, JHL93, Kuy91, Mol78, Ost83, Ste77c, UvG89, Duf77, SG83c]. Residual [Fre93, GL95c, KPJ82, Ske81, Stu85b, Wil77c, Ste91e]. residue [VS88a]. résolution [Ben24]. Resolvent [Ruh94]. Resolving [Fle93]. Resource [Gro95a]. Resources [BDG^{+}94b]. Respect [Hop82]. Responsibility [Pas85]. Restarted [LS94]. Restricted [AGG88, Bis89a, KT85, MG91]. Result [Bol85, Yuv78, Fle89]. Résultats [PV74a]. Results [AD89, AS85, Bai88a, Bai91, Eis68, LPP93, dP80, Roh95, Sch86a, SC90, SH83, Gen78, JBNP91, TW91a]. Retrieval [Sal89]. Retrospect [KU68]. Return [Eve88]. Reuse [Hen92]. Reveal [GM74a, Geo78, Mal72]. Revealing [Cha87, CH92b, CI91, HP91, PE95, Ste93d, AGS91, BH91, CH90a, Ste91d]. Revelations [Tur84]. Revenge [DF87]. Reverse [DEK95]. Review [Bel31, Bod49, CK84, EE84, Fle65, HS81b, Her88, Hig90i, Hig90j, Hig95d, Hig96d, Hig96c, Hoe87, Hof85c, Meu92, Sal81, Sha87, SH83, Uhl92b, WPG92, Web92b, You90, KY88, Wil86a]. Reviewers [SE62]. Reviewing [HW79a, Wol85]. Reviews [LW86, SoIPUG87]. Revised [Tho42, JW85b]. Revising [Lan92]. Revisited [Buz92, LZ94, Wil86b]. Revisiting [Fri91]. Rewriting [Rol67b]. Reziproken [Sch33]. Rhyming [Lee81]. Rhythm [Gid86]. Rhythm-a-ning [Gid86]. Rib [Rol75b]. Riccati [AL84, Bal80, Bar79b, BD74a, Bye85, Bye87, CV89, GL86, GL88, GL91a, GL95a, Ham82a, Inc83, Inc84, KLJ89, KLW89, KLW90, KH90a, KLP92a, KLP92b, Lau79, MD92, Rob80, SH83, Van81]. Richard [Ano83]. Richardson [ACD^{+}93, Gol62, GO88, Zaw91]. Rid [Hol74a]. Riddle [Ano86b]. Ridge [HGD90]. Riemann [Roe86]. Riesz [Gil91a]. Right [Bie37, HH92b, OP64, SG92, Sun95c]. Right-Hand [HH92b, OP64, SG92, Sun95c]. Rights [Jon74]. Rigid [LS82, LS83, SC70, SW91, SW94]. Rigorous [CKP95, Wil61b]. ring [BV86]. RISC [Bai93e, DMdB90, IBM93]. Ritz [MSS84, vdSvdV87, Tan94]. Road [EK94]. roads [Tho92]. robots [Tog86]. Robust [And91a, Dem92b, PR93, Shr85, BN87a, Wee89, DK93a, DK93b]. Robustness [Hof89a]. Roger [Uhl92b]. Role [Dat85, For80, Moo88, Tau72, Har84a, Mey75, Tau70]. Roll [BKP93]. Roll-Up [BKP93]. Romberg [Lau75]. Romberg-like [Lau75]. Ronstadt [Ano85c, Ano86b]. Room [WSW72]. Root [Ada67, BH83, EKNX93, Fra89, Hig86d, Lev80, Mos86, TSY88, Bel68, Gow58, HW78b, Pul66]. Rootfinder [Shr85]. Roots [AMS74, Den81, DLR85, EK93, EM94, EM95b, Gen73, Goe94, Hig87a, HW79b, Jos91, Kur92, McK84, MM83a, Nam85, Sch65, Sch92a, Tau49, TST86, Uhl92a, Wat94, Wil55b, AS82, CL74, MN79, Pot76b, Put57, Putxx, Rin60, Smi68a, Tau70]. Rorres [Hig95d]. Rose [Mau87]. Rosemary [Ros90]. Rosenbrock [Sha82b]. Rotation [Bro67, BK69, Bro72, Cli66, Cra74, DK70, Gru70, Ham74, Luk83, APP88]. Rotational [Hou58a]. Rotations [AP94, BI87, tBK84, Giv58, Rat82, Wil63a, DI86, Luk84]. Rothamsted [Gow85]. Rotten [Ber90]. Roulette [Ede93a]. Round [Abd71, AB79, BC86, Cap71, CB90a, Cyb79, Des63b, Dou59, Gol62, GF75, MM70, Low60, Lyn64, Pry81, Qui94, VA86, Ziv82]. Round-Off [AB79, Cap71, Des63b, Dou59, Gol62, GF75, MM70, Low60, Lyn64, Pry81, Qui94, VA86, Ziv82, BC86, CB90a, Cyb79]. Rounded [Dia78, Gay90, Ziv91, Lag74]. Rounding [Bau74, Boh75, DM90, For59a, Gra93a, Hen80, Mül83, Iri91, Kie87, KPS77, Lem75b, LO82, Lin75, Lin76, vMS94, Neu85, Not93, Oli79, Qui83, Roh90, SZ80, Stu80, Sun92b, Tur48, Ura68, Wil63b, Woź74, vdV90, ITH88, Lau75, Ste79a, Bod49, Mol90b]. Rounding-Error [Kie87, Sun92b]. Rounding-Off [For59a, Tur48, Bod49]. Roundings [Yoh73]. Roundoff [Bar81a, BB85a, BR71b, DM71, Fle88, Hen66, HS66, Kah86c, Kah92c, KL73, Kar91, LS78, LS80, Mil75d, Mil76, Mil78, MW80, MS78b, MS78a, QT90, Ram71, Tie70, Tsa74b, Ura68, War77, Woź78, Woź80, Yal94]. Roundoff-Error [Woź80, Woź78]. Routine [Lyn83, Cra86]. Routines [CDO^{+}94, DCHP90, EP87c, Fik67, GBDM77, JW86, KLV91, Lyn69, LK76, SBD^{+}76, KL89a]. Row [BS92c, Gra92]. Rule [Mol74, Ura68]. Rules [BD87, Ber71b, CBB54, Gus70, Har83, Lyn87]. Run [Woo92]. Rundungsfehler [Wil69]. Rundungsfehleranalyse [Neu74]. Runge [BGD^{+}89, But87, Cas89]. running [PH85]. Russian [Bor94].

S [BDDC88, Hig90j, Kah73b, vdV90].

S.O.R. [HW76b, O’C73]. Saddles [SSL93]. Safety [Ske89]. Said [BG89, Byr88]. Same [Syl51a]. Sample [Bar91a, CGL83, GKL95, KL94, KLV91]. Samuel [Bos86]. sand [Kah83a]. Satellite [Wah65]. Satisfactory [For69b]. Satisfies [Kah73a]. Satisfy [Asp59]. Saudi [Uni92]. Say [Pir82]. Saying [Mad90]. Scalability [Sch90b]. Scalable [CDPW92, DvdGW92, Ede94a]. ScaLAPACK [CDPW92, CDO^{+}94]. scalar [Gay91]. Scale [Ano48, Bai94a, Ber92, CL90b, Col91, CGT92a, CGT92b, CW86, Día89, ES92, Ker89, MGM93, Sau72, Wou87, Dat89, ITH88]. Scaled [BI87, BD90, Bau63b, Bau69, Bus68, Ede92b, FL74, O’L90, Sha82c, Ste84a]. Scaling [AP94, Bau63a, CR72, Dah78, ES93, GV74, Gow84, HPW90, KLW89, KL92, PL82, Sha85, Sha91, Ske79, Wat91c, EH79, Ste84b]. Scalings [BB95]. SCALP [FM91]. Scattering [KvSR90]. Scene [Par84]. Schatten [Dav85]. Scheme [Mül83, Zha92, Van88, Ple86]. Schemes [Fis74, vdG93, Oli79, RR82, Bul80]. Schmidt [DGKS76, JP90, Sch73, Abd71, AY81, Bjö67b, BP92a, Bjö94, Hof89b, JP91, OW90, Ric66a, Ruh83]. Scholarly [Jac86]. Scholars [vL78, vL92]. Scholarship [Dil81]. School [CKW85]. Schools [Var90c]. Schröder [Vrs86]. Schur [BD93c, BH83, Cot74, DK93a, DK93b, GMSW90, GNV79, KW89, Kåg93, KLW89, Lau79, Mat95c, Mor80, PV81, Ste85a]. Schur-Complement [GMSW90]. Schur-Riccati [KLW89]. Schwarz [HM90a, HM90b]. SCI [Gar91a, Gar90b, Gar90c, Gar90d, Gar91b]. Science [BDM^{+}88, Boo84, Boo93, BR67, BR68, EW90, FDR86, Gar76a, Gar79, Gar85, Gar91a, GS90c, Har81b, Kir92, LL86, Mad83, MD64, Mer90, O’C91, Par90a, Par93, Par85, Pit90, Sav67, SG91, Swa88, vdW75, Web73, Web82, Web92a, vW66, Maz76, Bus90, Gar64, Gar90e]. Sciences [Bas83, BP79, BP94a, Bev69, Bus90, GS76a, Gar90e, Hig93b, Hig94b, MMR78b, RJ93]. Scientific [Anh94, Bai93e, BRK^{+}87, Bol88, Boo93, Bra87, BC86, Cri89, Cro90, Day77, Day83b, Day88, Day92, DH95, DT93, DT94, EBR87, EL69, FKR86, Fox71, GH93, Gar73c, GO96, GS90c, Gro89, Hig96d, Hoe86, Hoo80, HRV92, HCSW88, IBM88, Cra89, Kah90b, Kat85, KKL^{+}93, KM83, Lar83, Lin84b, Mac77d, Man75, Mat90d, Mio90, MT86, Moo36, Moo75, Moo88, Nas84, Nas90b, OW77, O’C78, O’C86, PFTV86, PFTV88, PTVF92a, PTVF92b, Rat85, Ric88a, Sha87, Sie87, Smi75, SF87, Swa71, TS80, TK89, Var90d, Wil95, MS88, Ric88b]. Scientist [For68, Gol91a, Hig94e]. Scientists [Bar78, Day92, FM68b, FDR86, Ham73, Jen77b, Ken82, Kuy91]. Sciwrite [Dix73]. Scope [Moh85]. scores [Fox87b]. Scrabble [BF87, Hin77, AJ88]. Search [Ber91, Dax90b, GB63, Gou86a, Hig93c, Hul67, Mil75c, Rei93, Swa72, Swa74, Tor89, Tor91, Tor92, Tor93, DT91, NS90]. Searching [GA94, Pet91]. Secant [GS76b, Bar65, DS79a, DFW90, Ip87a, Sch83b]. Second [Bha93, Gol62, Hig90i, Kub91, Nei84, Ort72, Ort87, Web92b]. Secret [Bod86, Lew78, Sha79]. Secrets [Pou83, Pou86]. Sector [STW84, TSY88]. Segments [SL54, Tau49, Tod54b]. Select [Gar84]. Selecta [SG83c, SG83b, SG83a]. Selected [Hay89, Sch86a, Wil71c, Wil71d, Che86, Ran75a, Ran75b]. Selection [Mac77d, Ste90c, VV87]. self [Fle84, Putxx]. self-adjoint [Putxx]. Semantics [KC82]. Semi [Fle85b, Hig90a, Róz88]. Semi-definite [Fle85b, Hig90a]. semi-separable [Róz88]. Semidefinite [Hig88a, Kel65, SH93, All88, And85, HW88, HM90b, WY64]. Semidefiniteness [Ker90, Mak91]. Semiiterative [EMN88]. Seminar [Fra92a, PA88]. Seminormal [Bjö87]. Seminumerical [Knu81b]. Sensitivity [Bar91d, BH92, CI95b, DM78, Dei86, FM86, GLKH90, GL95a, HK88, Hop82, Jon84, KH90a, Wil86c, GM86, Van77c, Roc81, Tog86]. Sentence [Gor93b, Kin67d]. Sentimental [Ano86b]. Sentiments [Fox90b]. separable [Osb65, OW85, Róz88]. Separating [Sti91]. Separation [How83b, KP93, Var79]. September [Maz76]. Sequence [Ano46, BGH92]. Sequential [Bur95, FM68a, Ste94e, Tor92, Tra73, GMSW86a]. Series [And71, Boy94, Cle55, CT65, Cox86, Cyb83, ES88, Hew93, Hig96b, Hop82, New73, Oli77, Smi65]. Services [BDG^{+}94b, Sch90c]. Set [CR83, CH88b, DCHH88b, DCHH88a, DCDH90a, DR82, FTxx, Gea69, Hal74a, Har80, Lem75b, Lot55, Lus89, Oet65, PJS92b, Sny55, DCDH90b]. Sets [AHB87, BD85, CNRT81, DL92a, DRMN79, DR82, Phi87, WSW72, DNR87, Rog76, Vrs86]. setting [Jus72]. Seventh [Hig95d]. Several [Box66, FG86, Pic91]. Severe [FT86]. Sexist [MS89]. Shadowing [CKP95, Cor92b]. Shadows [SSL93]. Shakespeare [Har66, TE87]. Shape [Ros90, GWW92, Kac66, SBKZ93]. share [Sha79]. Shared [DEKV92, DDSv91, GHL86, KP92, Wri90]. Shared-Memory [DEKV92, GHL86]. Sharply [OPW65]. Sheet [BKP93]. shell [Bou78]. Sherman [Yip86]. Shifted [Man79]. Shifting [vdG93, Ste71a, Wat92a]. Shifts [DG94, Wat95]. Short [Cro41, Pec87, Pic91]. Shorter [Oxf93]. Shorthand [SZ80]. Should [Gol91a, HR80, Hym86, Kah87c, Bud87]. Shoulder [GH43]. SIAM [Gea91]. SICEDR [Don82]. Sided [Kol95, Sch68b]. Sides [HH92b, OP64, SG92, Sun95c]. SIG [PC-87]. Sign [Bye87, BHM94, DR94, DB76, GL86, Hig94c, How83b, Kah87a, KLJ89, KL91a, KL91b, KLP92a, KLP92b, KL92, KL93, KL95, KBS94, PKL90, Rob80, STY83, Sti91, SA75, Bal80, Bye86, DLR81]. Sign/Logarithm [SA75]. Signal [Aru92, Cad88, Cyb83, Hay89, KvSR90, Sch86b, SF86, Van91, APP88, BS89b, Cyb78, Van89, Kel88]. Signe [Bar79b]. Significance [FG82, FT86, Kah88, Mei69, Met77]. Significant [Bar88b, BR93, Boy94, Gar76a, Hil95, Rai69a, Tsa74b, Tur82]. Significant-Digit [Hil95]. Signs [Dan76]. Silicon [Sla87]. SIMD [BPT90, BMSV92, LZ91]. similar [Act74]. Similarität [Gab79]. Similarities [Str72]. Similarity [Ham82c, Rat82, TZ59, Wil59, Ruh69]. Similarly [GL80]. Simple [Bro81, CH88b, CH89, Gea69, Gou86a, Grä85, GEJ95, McK84, Nam85, Qi84, Tsa83, Yuv78, Ste91e, Ark86]. Simplest [Mol90a]. Simplex [FM84, GM73, NM65, Ogr88, PH72, RSD77, DW87, Dan87, GMS88]. Simplification [Mos71a]. Simplified [Mac77a, Mac77b, Mac77c]. Simply [Kah93a]. Simulation [Gea77, App85]. Simultaneous [BSD60, BGBM93, CJ70, FHW48, HL86, Sny55, Ver49, Chu91a, FG86, LHPW87, Ste75c]. Sinatra [Ano84]. Sing [Wat91a]. Single [FS76, JW85a, Wil77c, Dat87]. Single-Precision [Wil77c]. Singular [AK75, And90b, AP76, BD93a, Bar86b, BZY93, BS89a, Ber92, BLV85, BGG88, BN87b, CR86, CI94, CI95a, Dax90a, DDLT91, DK90, Dem92c, DG93, Don83, Dor71, EMN88, EI93, FP92a, GK65, Gol68, GR70, GLO81, GDD94, GE95c, HN81, He90, HK93b, HP93, HP94a, HP91, HP92b, Jen82, JS92, Kar74, Kel65, Kit95, KL80, LPP93, Van85b, Van86a, Luk86, MS93a, MP77, MM83b, PS81, PS94, Par86, Qi84, Roh92, Sch86c, Ste81d, Ste83b, Ste93b, Var75, Wat92d, Wed72, Wil78b, BV86, Bis89b, BGBMN90, CH90a, Chu86, Dem88a, FT88, Ham85, Han90c, HLPW86, Hor54, HC63, Kåg84, Van76a, Luk80, Man82, Mao86, Mat90b, MG89, MG91, MZ91, Pai85b, Pai86, PV86, Ruh75, SV79, Ste79b, Ste84a, Ste90f, Ste91d, Tog86, Tsi90, Wat90]. Singular-Value [Wil78b, Luk80]. Singularité [PV74b]. Singularity [For59b, Kah89c]. singularly [Sch91]. Sir [Bel31]. Six [Min74]. Size [BB88a, Coh74a, Cry68, Ske81]. Skew [Bun82, Hac93, Inc84, Rin60, Buc77, Paa71, WG78]. Skew-Symmetric [Bun82, Hac93, Inc84, Paa71, WG78]. Skewes [Boa79]. Sky [Moo93]. skyline [Fel75]. skyline-stored [Fel75]. SLAP [Sea89]. Slash [MK85]. Slide [Ano78c]. Slides [Kod79]. Slightly [Gre89, Mee80]. Slowed [Hal90b]. Sluggish [Ste94c]. Small [Bry89, Cha84, GKL95, KL94, Lod84, MP83, Nas84, Peñ94, PV94, Gay87, Nas85]. Small-Sample [GKL95, KL94]. Smallest [Bar86b, Ede91a, HP92b, Jun95, MH87, Var75, Wat88a, Dem87a, Tod65, Wat88b]. Smile [Web92a]. Smoothly [vdV92]. Snippet [Fox78]. Snoopy [Sch82]. Social [DLP79, SS76]. Society [BM86]. Software [AL84, Boi90, Bro75, But92b, CDW94, CW80, Cod82, Cow77, Cow84, Dem84, DK93a, DK93b, DG87, DH90, DRW93, Don95, DW95, Duf84b, Eva74, For92, For78, Fos79b, GWL^{+}92, Gli89, HD80, HRV90, Hul79, Jac78, JW86, KP93, KP94, Kah80b, Kah91c, Kah84, KMN89, KF87, KP76, Lin81, MM82, MS93b, Mil75d, MW80, Mil84, MS78a, MGH81a, MGH81b, MW93, Ost83, Par84, PH94, Pla93, Pol89, Ric71, Ric77, Ric81, Ric83, Ric88a, Ric92, SW84, Yoh79b, Yoh80, ZWS86, BDM89, CDM79, Ein79, Gea77, KW83, KO91, PW77, Ric88b, LPW83, MS78b, Uni92]. Solid [KR65, Kit71]. Solo [Ano78b]. Solution [Abd71, Abd74, AG90, APS93, AS93, AM94, AR92, AMR88, Ata40, Axe94, BLS91, BP81, BMvN46, BI87, BH88, Bar93c, Bar79a, BBD^{+}94, BBC^{+}94, BSD60, BD74b, BL84, BP70, BP94c, BMPW66, Cas79, CI95b, CG67, Chu87, Cli73, CL91, CGO76, Cra74, DS89, Dor71, Dou59, DRMN79, DR83, DR93, DR95, EMN88, Ept80, Eva71, EO79, ED83, FKF68a, Fel75, FM67, FHW48, Fre80, FGN92, FSC73, GP70, GL86, GLAM92, Gea81, GL81a, GF74, GJ65, Hal77, Hal80, Ham82b, Han92, Har80, HR88, Hig88b, HP94c, HP92a, HMW77c, HMW77a, HMW77b, Ips94a, JW85a, Jon84, Kel65, Kog55, Kog74, Kre72, LN72, Lem75b, Lin92a, Lu94, Lu96, Mac89a, Mag60, MPW66, Mil88b, ML85, Nab92, OP64, Oet65, OV85, Ort88, Ost73, PR89, Pap93b, PW71].

Solution

[PPY92, PA92, RG67, Rob80, RO88, Rut58, Sal60, Sau94, Sch66, SGB91, Sha94, Sol80, Squ87, Tem75, Tra66, Var72, Ver49, Vit69, Wac88, Wac92, Wan91, Wes68, Wil74a, Wil77c, Wol65, Woź77b, You71, Zla82, vdV92, All88, ADD89a, BIM90, Bog74, BGS87, Buc77, Cad73, Cad74, Cal86, Dav81, DI86, DFW90, Dor70, DNR87, Duf89b, Enr79, GW66, Kub62, Kub70, Laa58, Lax65, Osb65, PR85, Sha82a, SZ91b, Wil67b, Ziȩ83, BS72]. Solutions [Bak84, BRB76, BD74a, Bjö67a, BG67, Bjö68, Bjö78, Bjö91a, BPE92, BG65, CP76, CR79, DM86, DK88, Fra92b, Frö53, GL95a, Hea77, Hig89a, KLJ89, Kov76, Kov77, Lan70, OPW65, Roh93, RK95, vdS70b, vdS75b, Sny55, TW92b, AL75, BJW85, CH90a, Chu89, Don87, GR70, Han90c, Hua90, KLW90, KM76, Rot84, Vet75, Ziȩ87]. Solve [BR96, Kah79, Kah86d, Pol57, RG92a]. solvents [Dav83]. Solver [CR92, Duf81, For69b, LNOP82, Mol2b, PS92]. Solvers [ADR92, AGL95, CKS78, DH93, DKH86, Enr89, EP87c, Hig95e, KLW89, Sch90b, Sch81b, Wri95]. Solves [And91a, GP90]. Solving [AL85, ADD89b, BG76, Bjö67b, Bol84, BKO94, BP71, Bun76, BK77, Bun85, BDV89, BGN70, Bye87, Cro41, CNRT81, Dax90a, DHP94, DK85, DS84, DDSv91, DV92b, DR82, DS93b, EHHR88, For53, For69a, GP71, GH93, GL88, GWL^{+}92, GS78, GSB88, GS90b, Gol65, GS76b, HW91a, Ham82a, Hig87b, Hig90h, Hig90k, Hig91c, Hig91d, HK93b, Hof87, Hou50, JW86, KW89, KP93, Lau79, LH74, LYK89, Le85a, LC89, LZ92, LZC92, LZ94, MP74, Mal93, vMS94, Mio90, Nas84, Pai73, Löt84, RB85, RG92b, Saa84, SB77, ST86, Sch86c, SW81, Sco81, SWD76, Smo95, Spe91, Stp93, TS75, Tsa91, Tsa92, Van81, Wat93b, WWWS86, Wil60c, Woo85, Woź77a, Yip86, You54, Bal80, Bar65, Boj84, DD88, DM84, GO88, Ste73c, Ste90d].

Some [BD87, BS68, Bar75b, Bar94, Bel68, BB92, Bro73, BK77, Cai84, CH92b, Cry76, DD86, Dwy67, FH55, FL74, Gar76b, Gay79, Gen78, Gle86, Gol73a, Gol73b, GS74, GO89, Hal88, Ham82c, Hot43, Hou50, JW86, Jor68, Kar91, KL80, FB91, Kov70, LSV95, Lin74, Van82a, Van84b, Lyn87, Mac89b, Mal93, Mar89, Mee80, Men56, Mil87b, Pai87, Par92c, Pin85, Qi84, Roh95, Rut47, Rut52, SC90, STY83, SB88, SH83, Sma90, Ste69, Sun92a, Tao84b, Tsa74a, Tyr91, Wat93a, Wat92c, Wil71b, Wil77b, Woź74, Zie88, Cyb78, FU69, Gab87, Gay91, HP84, Kie73, Kub62, Li88, Van83a, MM88, Pai80, Vas90, Wat90, Dat89, Fox90b, HW79a, LM90, SR90]. something [Kah77]. Somme [Pic72]. Song [Wil72a]. Songbook [Cah89]. Songs [Knu84a]. Songwriters [Lee81]. SOR [LT88, SN91]. Sorensen [BDDC88]. Sorry [Mad90]. Sounder [Woo67]. Source [SF86, Swa88]. Sources [BKW80, Cow84, Kah84]. Soviet [War60]. Space [Ash93, CS84, FH55, Fos86, GL94, GMS^{+}85, Hal67b, Hal82a, Nix81, Pow86b, SW62, You88]. spaced [TW91a]. Spaces [BD85, Hal74c, Ost73, Ter85, LR81]. SPARCompiler [Sun92d, Sun92c]. Sparse [AYB90, APS93, AS93, ADdR89, ADD89b, ADNR92, Bar86a, BV92b, BL95, Ber92, BLP90, Bun76, BR76, CES93, Col84, CH88b, CDG91, DGL91b, DGL91a, DV92b, DLN^{+}94, DRMN79, DS79b, Duf81, DR82, DR83, Duf84a, Duf84b, DER86, Duf88, DGL89, DGR^{+}91, DMR92, DGL92, DR93, DS93b, DR95, ETA90, ER74, EGL^{+}87, EO79, GL81a, GMS92, GMSW84, GMSW90, GL81b, Hea84, HP94c, Kau87, LSV95, Liu87, Mat92d, Par84, PPY92, PA92, Rei87, Sau94, ST86, Sch93, Sch90b, Sco81, Zla82, ZWS86, ADD89a, Duf77, DNR87, Duf89a, Duf89b, Gay91, Geo72, GMSW87b]. Sparsity [DEGR88, GR89]. Speaker [Bra83, Par93]. Speakers [Mac90, MS89]. Speaking [Bee92, Boo84, Boo93, CB81, Ken82, Tur85, Tur74].

SPECFUN [Cod87]. Special

[BD87, BG94, Cod87, Fie86, GSB88, MMR78b, Mau91, MH57, Nei82, SGB91, Tom92, WWG51, BGS87]. Specification [DCHP90, Dem92e, For93, Wic89]. Specifications [Dem92b, Mac77b]. Specified [BD91, BK69, Bro72, KP94, Luk83, Chu91b]. Spectra [BW93, Kah75, Kos92, Trexx, You79, Shr84]. Spectral [Bag94, BDG94a, BFP88, BJ92, Cai84, CHQZ88, CES93, DLR81, Els76, Gem86, GKK90, Hal74b, Hen62a, Hol92, Mal92a, Mal92b, Mal93, MT94, MT95, Mat90c, TT87, AG88, Des63a, NOP84, CGR90]. Spectrum [Kah73a, Ost54, dP80, Chu91b, Ste75a]. Speech [Asl89, Tur85]. Speed [Bai88c, BF88, BB49, DS95, Pan84c, Rad48, Rod92, Gay91]. Spelling [DLS83, GHR67, Pet80a, Pet80b, McI82, RS81]. Sperry [PH86]. Spirals [Dav93b]. Spline [AN63, Cox75, Cox78, CH89, Dem77]. Splines [dB72, Cox72, Pre75]. Splitting [Lei89]. Splittings [Ple76]. Spread [Mir56b]. Square [AMS74, BR71b, BH83, CL74, Fra89, Gen73, Hig86d, Hig87a, Lev80, MM83a, Sch92a, AS82, Bel68, HW78b, MN79, Pot76b, Pul66, Put57, Putxx, Rin60, ST80]. Squares [Abd71, Abd74, ADdR89, AHB87, Aru92, BI87, BH88, Bar88a, BV92a, Bjö67a, BG67, Bjö67b, Bjö68, Bjö78, Bjö87, Bjö90b, Bjö91a, Bjö91b, Bjö91c, BPE92, Bjö96, BG65, CNP94, Cli73, DS89, Eld80, Far88a, GL94, Gay88, Gen73, GS90b, Gol65, Gol68, GKS76, GB86, GDD94, Gul92, GW92b, Gul94, Gul95, Hea84, JO74, Jor68, Kar74, Kel62, Lar66, LH74, LN72, Van85c, Lon67, Lon84, Man81, MD92, Löt83, Löt84, PW70, Ple74, Pow67, PR68, PR69b, Rei91, Sau78, vdS75b, Ste77a, Van87a, VV91, Wah65, WKS95, WSW72, Wed85, Wei90, WH16, APP88, All88, Bjö88, CH90a, Chu91b, Cyb84a, Dem87a, GHM^{+}86, GW66, GR70, GV80, Van76b, Van83a, Mar63, Pai79, PP89, Ste75b, Ste87a, Ste90e, VV87, Wed73a, Wed78]. Squaring [Fri91, KLP93]. Squeezing [DE84, DKH86]. SR [Kah77]. SR-52 [Kah77]. Stability [AR92, Bab69, Bab72, Bar86a, BL80, Bjö87, Bjö92, BBdHS95, Bol84, Bro85, Bun85, Bun87, BDV89, CH92c, DSH95, Dem87b, DH92, Dem93b, DHS95, Dia78, Erh93, ER74, FGS95, GSS96, GN82, Hig90h, Hig90k, Hig91c, Hig92f, HP94c, Hig95b, Hig95f, Hig95e, Hig96a, HP86a, IM94, JP90, JP91, JW77, JT74, dJ77, KL88, KLW89, Lem76, LS91, Lew92, Lin92a, Lin92b, Mat84a, Mat84b, Mil75b, Per69, PW75, Pri92, RR82, RT92a, Rei71, Row90, SB88, Ske79, Ske80, vdS70b, vdS75b, Smo95, Ste94e, SRD94, TS90, TTRD92, TTRD93, Vit69, Wat95, Wil59, Woź77b, Woź77a, Wri95, Yal91, Yip86, Zaw91, BD76, BBdH91, Bus71, Bye86, CF91, Cyb80, Cyb84a, DD88, Stu85c]. Stability-Enhancing [KLW89]. Stabilized [GLW89]. Stable [AO86, ABO88, BZY93, BPT90, BK77, Bun82, Bye88, Cas79, DGKS76, DK87, FTxx, FM84, FM85, Fos91, GLKH90, GM73, Gil74, GS76b, GE94, Ham82b, HK88, HW79b, KH95, Oli75, TSY88, Vav92, Ziv83, Gau90, Van85a, Pai79]. Stack [Kah94b]. Stalk [Cip88]. Stamp [Kod79]. Stand [Hig96b, Tur91b]. Standard [BM79, CL79, Cod81, CCG^{+}84, CC93, DV92b, Don95, Far85, Fat82, GB91, IEE81, Hou81, IEE85, IEE87, KP79, Kah80a, Kah80b, Kah81, Kah87b, LIA93, Han90b, JW85b, Kah83b, Kah95]. standardized [MG89]. Standards [Anoxxb, Anoxxd, Cod88b]. Star [Ash86b]. Start [KW92]. started [LSV95]. State [IP87b, Jac77, Kit71, SWD76, Spe91, Wil85, Ano93]. Stated [Syl51a]. Statements [Knu74a]. Static [Cod73]. Stationary [AR92, Dax90a, Dax90c, Hey87, HK93a, HK93b, Kni93, O’L93a, FM86, FP86b, FP86a, GM86, HP84]. Statistical [And71, Bas83, BF90, CCC82, FGHK83, Gol69, Goo92, GKL95, Hem67, HS87, KG80, KL94, KC73, Lin75, Mai84, Mil78, Mui82, SB94, Thi88, Tie70, War77, CF91, Gow85, Rao80]. Statistically [BR93]. Statistics [Bor94, Bur74, Gra69, Hal90a, Huf54, Llo84, Mac89b, Sig86, Ste94a, Fle81a, Ham85]. Statistique [Bru89, PV74a]. Status [Kah95]. STDTST [EP87c]. Steepest [Fle94]. Step [BB88a, Gil51, Kat85, CH90b, MS83, Sor92]. Step-By-Step [Gil51, Kat85]. Stephen [Web92b]. Stevenson [IEE81]. Stewart [KR78]. Stiff [Aik85, Cas79, Sha82a]. Stiffness [HT93]. Stochastic [BC86, Mey89, PL82, Ste90g, VA86, MSS84]. Stock [Qui83]. Stokes [TTR92]. Stolen [Mal89]. Stoneham [Sto80]. Stop [Car58]. Stopping [Ada67, ADR92, BB94, Bra95]. Storage [BLR90, SV89, Fle90]. Store [BSD60, CNRT81, Nix81]. stored [Fel75]. Stormy [Mos71b]. Story [HS93a, Hou89, Lav80, Lew78, Sch89b, Whi87]. strange [Hén76]. Strapdown [BI75, BIMF76, BI77, CL72, Mao86]. Strassen [Pan78, BF88, BLS91, BHH93, CR76, DHSS94, Kre76, Sol80, Yuv78]. Strassen-Type [BHH93]. Strategies [Bun74, Dax82, Inv93, Peñ94, Wat92a]. Strategy [Liu87, ON93, BL90]. Stretching [Grc90]. Strict [Stu85a]. Strictly [Pin85, MN79]. Strong [Bun87, BDV89, Smo95]. Structural [BS92a, Ple86]. Structure [AwJS85, BH91, DEGR88, Gre52, Her65, MP83, RB89, WC53, Wat80b, Ruh70, Cri75]. Structure-preserving [BH91]. Structured [BH92, GK92, GK93b, GKO94, HH92a, Knu74a, Van91, Wat92d, CB90b, Dem88a, FT88, Pan90, Tsi90, Wat90, Bar91d]. Structures [BN71, GK92, Vet75]. Structuring [Ler83]. Student [Bak85a, EBR87, HS83, Und80]. Students [Bar78, Bar82, BKW90, BR67, BR68, For57b, Kat85, Par93, PP94, Var90b]. Studies [Gar82b, Gol84, Hir65]. Study [A^{+}90, BRK^{+}87, BB87, Buc94b, CJLP93, Erh93, Gri89, Kar91, Knu71, KC73, Pre93, Tau72, Tay82, Ton92, Wic93, CF91, KBS90, Van75, Tau70, Vrs86]. Sturm [BGH92, Eva71]. Style [Anh94, Bye84, Gre87, KP93, KP74, KP78, Kir92, Lon93, Ame65, Uni69, Uni82, Ras86, Sto80, SW79, Wat89, Wil88a, You89b]. Stylist [Bak85b]. Subdifferential [Wat92c]. Subgradient [Bar91d, Tao84a]. Subgraphs [PPY92, PPY93]. Subject [Gau95, Löt83, Löt84]. submatrices [DGK80, Wat88b]. Submatrix [Wat88a, Dem87a]. Subprograms [DGL91b, DGL91a, DCHH88b, DCDH90b, DCHH88a, DCDH90a, HK87, LHKK79]. Subroutine [Cod88a, Don82, FHS78, IBM88, PdDKÜK83, MS88]. Subroutines [DR82, MGH81a, WWG51]. Subset [VV87]. Subspace [HS95, HR92, KLP93, Pai94, HW88, Ste90h]. Subspaces [BD91, BG73, BHM94, Dem87e, PW74, Ste71b, Ste73d, BS89b, DHW92, LR81, Van82a, Ste87b, Wil75b]. Substitution [Lev58]. Substitutions [Hym86]. Subtracted [MT93]. Subtraction [FG82, FT86]. Successfully [Ber75, O’C91]. Successive [Dek69, JR75, Lyn64]. Sufficient [Kur92, Sha82c]. Suggestions [For57b, HW79a]. suitable [Wed78]. Suite [DM89, SR95]. Sum [Fer94, HS81a, Fer95, WSW72, Ove83]. Sum-Free [WSW72]. Summary [SH83]. Summation [Bar91a, Cap71, Cle55, Esp78, Gre72, Hig93a, JSW83, Kah93a, Kie73, LO82, Lin70, Mal71, Neu74, Wal71]. Summen [Neu74]. Summer [Var90c]. Summierungsverfahren [Nic70]. Summing [Smi65]. Sums [Ant85, Dub83, Inc85, Lin74, MA87, MM83a, Sch92a].

Super [Lor74, RSD77]. Super-Power [Lor74]. Supercalculators [SW88]. Supercompilers [Wol89]. Supercomputational [EW90]. Supercomputers [And90a, BSBF89, Laz88, Pit90, Wol89]. Supercomputing [Car89, Got92, HPP88]. Supernatural [Kor88a]. Superstars [Gar90c]. Supervisors [PP94]. Supplement [Gau95]. Support [CC93, Far88b]. Suppress [CEN91]. Supremum [SS62]. Surely [Fey85]. Surreal [Knu74b]. Survey [BL77, Bro91a, Bru95a, CR94, COT89, Duf84b, Ede91b, Gar76b, Hel78, Hig87c, Hig94d, Hig96d, Kah72b, Van90, MM64, Pow70, RB78, Tod62, TW91b, Wil68c, YG72, YG73, BG87, Duf77, GAG88, Sch89a, Zie84]. survival [Wil88b]. Survive [Ste81a]. Survivor [FF87a]. Suspension [LM90]. SVD [DCHS88, GE95a, Han87, Han90b, Van83a, Pap93b, Shr84]. SVD-Related [Pap93b]. Swapping [BD93c]. Swedish [Mac90]. SWEEP [Goo79]. Switzerland [Sch81c]. Sylvester [Ano04, DSH95, BMO92, GWL^{+}92, GLAM92, GL95b, HR92, KW89, KP92, KP93, Kåg94, Ste92b, Syl51a]. Symbiosis [GS90a]. Symbolic [CM93, Dav86, Gro89, Kah90d, The93, Mos71b]. Symbolism [Caj25]. Symbols [Mur76]. Symmetric [Aas71, AGG88, AG92, AGL95, Bar93a, Bar93c, BG76, BS89a, BS92b, BL91, BP71, Bun71b, Bun74, BKP76, BK77, Bun82, BDV89, BGG88, CH87, Chu89, CJ70, CW85, CV86, DK77, Dax82, Doy57, DRMN79, DR82, DR83, DGR^{+}91, DR95, EH92, Fan93, Fle76, FGS95, GSS96, Giv54, GE94, GE95b, Hac93, Hig88a, Hig88d, Inc84, Jen83, JP93, JP94, Kah66a, KM86a, Lan93, Lar66, LZS91, LL93, LZ94, Liu87, LPS87, Mee80, Meu92, Mir60, OS90, Ort63, Par80, PR81, Par94, Pot76b, SRS73, SW81, Sco81, SS91, Sla92, Smo95, Swa93, Tau72, Ter85, Van95, Yau93, Zha92, AL73, Buc77, Cra86, DI86, DCHS88, Don87, DS87, DS92, Ebe71b, Fel75, FG86, Fre90, GD89, Hua90, KBS90, Mat69, Paa71, PW69, SK77, Sch83c, WG78]. Symmetric-Definite [SW81]. symmetrisation [Kub63]. symmetrizer [VS88a]. Symplectic [SS92, Van84a]. Symposia [GO78, Var90c]. Symposium [Ano48, Gau94, KvSR90, Maz76, Wil60c, MMR78b]. Syndrome [Hol80]. Syne [Hal88]. Synonym [Rod78]. Syntax [Bee93, KC82]. Synthetic [Ste71a]. System [AL85, ADdR89, Bai94b, Bar89, Bjö92, Boi90, BGFB94, CKS78, CL91, DH93, DMdB90, Ede91b, GEJ95, Hig95e, IBM93, KC75, Lam86b, Lam94, MPW66, MS93b, ML85, Oet65, PM87, RG67, Rub88, SF86, STW84, SA75, Uni92, Van77a, Web92b, Wol91, Yip86, Ack89, Bis89b, Cad73, Cad74, GS71, Gro87, KO91, Kub65, LHPW87, dO69, Wil77a]. System/6000 [DMdB90, IBM93]. Systematic [TS80]. système [Ben24]. systémes [FKF68b]. Systems [ABS84, AG90, AS93, AM94, ADD89b, AR92, Ata40, BLS91, BMvN46, BBD^{+}94, BBC^{+}94, Bar91d, BH92, BG76, Bis88b, BP70, BE73a, BP94c, Boh75, BG94, dBP77, dBK86, BKO94, BMPW66, BS92c, BG90, Bre73b, Bre91, Bro68, BP71, Bun76, BK77, Bun85, BDV89, BL86, CL72, Cas79, CF88, Cha88, CH92a, CI95b, CK75, CP76, CR79, CDG91, Cro41, Dah78, Dat85, Dax90a, Dei86, DHP94, DS89, DK85, DB76, DS84, DDSv91, Dra92, DR93, DS93b, DR95, EMN88, EHHR88, EO79, Eva82, ED83, FKF68a, FGS95, FM52, FM67, FGN92, Fre93, Frö53, GP70, GP71, GJMS88, Gau90, GL81a, GMPS92, GSS96, GV79, HW80, HWH91, HR88, Hig87b, Hig88b, Hig89a, Hig90b, Hig90k, HH92a, HH92b, HK93b, HP94c, HP86a, Hof87, Hou50, HRV90, HRV92, HW91b, Kah90d]. Systems [Kas95, Kel65, KL88, Kie81, KH90b, KC73, Lam91, LC89, Lin92a, Lu94, Lu95, Lu96, LM66, Mac89a, MP74, MH87, Mat92b, vMS94, MK85, MP77, Mey89, Mol74, ML85, NRT92b, Neu92, NP78, NP84, OPW65, Ort88, PR89, PA88, PJS92b, Peñ94, PEM81, RO91, Rie56, Roh89a, Roh89b, RK95, RO88, RG92b, SB77, Sau94, ST86, SG92, vdS70a, vdS70b, Smo95, Sol80, Ste81d, Stp93, Tem75, Ton92, TS75, TW91b, Tsa91, Tsa92, Var72, Var90a, Vav92, WF82, Wic92, Wil60c, Wil74a, Wil77c, Woź77b, You71, Zla82, vdV92, AL75, BCJR87, Bog74, Boj84, Buc77, CF91, Cyb80, DEG72, DK83, DK84, Day83a, DI86, DCHS88, DFW90, Enr79, Fre80, Fre90, GS78, Geo72, GO88, GP90, ITH88, Van88, PR85, Per69, Ros72, Ste73c, Ste90b, Ste91c, Var91]. systems [Woź78, DST88]. Systolic [BI87, CV89, Sch86c, Sch89a, SHW86, BPT90, Sch83c].

T [Rei93]. T.P.A. [Anoxxb]. Table [RW92, Tan89, Tan90b, Tan91, Tan92]. Table-Driven [Tan89, Tan90b, Tan92]. Table-Lookup [Tan91]. Tables [AS64, Hig96b, Leh43, SL54, Bue77]. Taboos [Ber71b]. Taine [Rei93]. Tale [Hum88, Mol95]. Talk [Fle46, Hal74d, JHL93]. Talking [For57b]. Tangent [KL93]. Tape [BSD60]. Target [Bro72, Luk83]. Targets [SE62]. Task [IEE81, Smi90, Kol78]. Taussky [Sch77]. Taussky-Todd [Sch77]. Taylor [Lin76]. Teach [Kra93, Ste90a]. Teacher [Cip89]. Teaching [Hal94, Low84, Ren72, Wat83a]. team [RM91]. Teams [Kee93]. Tears [Pry81, Lin72]. Technical [Bak61, Bol88, Byr87, CK89, Coo64, Eis92, GP59, Gra70, HH93, Her65, Hig96d, Hoo80, HP77, Mol90b, Per91, Rat85, Ste91b, Tar92, TK89, Tur74, Woe92, Woo92, You89b, Cor94]. Technique [Cyb83, Dek71, LQ87, Pan78, MSS84]. Techniques [CKW85, DK77, Dil93, DV92b, EI93, Ell94, FM68a, Fou81, Gli89, LS94, Low84, Pan84b, Saa84, Ste78, Wil62, Woe92, BIF69, Pai80, ST91, Mit87, Ral81]. Technobbable [Bar91c]. Technology [Bee92, Gar79, Kir92, LIA93, LL86, Swa88, War60]. Tell [Gar73c, Sch90c]. Tempered [Gor93b]. Templates [BBC^{+}94, DEK95]. Ten [Asl93, Wol85]. tend [Ste81b]. Tensor [Pol85]. Term [Gau67, MT93, Qui94, QT90, Tro84b, Tur87]. Terminates [O’L93b]. Termination [NR72]. Terminology [Mac90]. Terms [JSW83]. Terrible [Man80]. Test [Bai94a, BI77, BJ89, BG79, Caf63, Cle59, Cod93, DM89, DGL91b, DCHH88b, DCDH90b, DGL89, Ede93a, Gay85, Gea69, Gur92, Han93, Han95, Hen66, Hig89b, Hig91b, Lot55, Lus89, MS76, Pei62, Rut68, Sch81a, Zie86, BMW86, Hig93e, Hig95g]. Testing [Cod82, GK69, JT75, Kah91c, Ker90, MGH81a, MGH81b, Pax91, Tan90a, Fle81a, KO91, Kor88a]. Testmatrizen [Zie74, Zie75]. Tests [HS66, JW86]. Tetraflop [SS90a]. Texas [Kah77]. Text [BR87, Har85, Sal89, van86b, Har81a, Har84a, HP86b, MFGK82]. text-editing [Har84a]. Textbook [McC78]. Texts [Knu91]. th [Smi68a, TSY88, TST86, HW79b]. Their [BL88, Bha94, Fie86, GF74, Hal90a, Har84b, HW78a, Hig95a, Kro87, PP94, STW84, Sta78, Tan91, Wil55a, DKW82, FF92, FU69, GS58, Hop82, Laz88, Lee88, MM88, Smi65, Tau70, BT64]. Them [Cip89, Gar90a, Gou73, Lav80, Woo92]. Themas [Hof85b]. Theme [Ker92]. Themes [bGBB87]. Theodorus [Dav93b]. Theorem [Art83, FV62, Gal71, Gem80, GHS87, Hel74, IIM87, Kee93, Pow69, Roh94, Syl51a, WT94, ST80, Ste75b, Ste91d, Wil87b, Syl50]. Theorems [BF60, DW84, Dun90, EH91b, Far87b, Gil91a, Mar89, SS62, Stu85c, Var76b, BJW85, Mat90b, DLP79, Syl50]. Theoretical [Bro91c, DD86, Har81b, Par90a, Par93, Wen66, KBS90, AGG88]. Theorie [Hil94]. Theorists [Sch77]. Theory [Bar75a, Bei66, BB87, BGFB94, Bur80, BL86, CF92, CF93, CFB93, Che82a, Che86, CSW83, CW85, Dat85, DD86, DK88, DK93a, Eld80, Eve66, Gal71, Gan59a, Gan59b, GJ79, Gar79, Gau95, Geu82, Gol91b, GB89, GE93, Hal67a, Hal74a, HHKR93, Hig90j, HH92b, Hig93d, Hig94d, Hod92, Hou64, Ioh82, JE95, KvSR90, Kat76, KM81, Lan69, LT85, Van82b, MM64, Mey89, Mil87b, Mui82, Noc92, O’C93, Ort87, Par92c, Per52, Pet81, Pow81, Ric66b, Riv90, SV77, Sch77, STW84, SW82, SS90b, Ste90g, Ste93a, Ste94f, Stu81, Tau88, Tre90, Tur29, TA32, Usp48, Wat80a, Wed73b, Wed85, DEG72, Mey75, Ste90e, Ste90f, Wil72b, BDM89, BB88b, CKW85, Cod88b, GC91, KP94, MO79, SP88]. There [Asl89, Bry91, Gea81, Jus72]. Thesaurus [D^{+}93]. Theses [Tur87]. Thesis [Hol74a, Hol75]. Things [Byr88]. Think [Ber49, Hal82b, McC79]. Thinking [HB86, Knu85a, Rol75d, Woo67]. Third [Tor65]. Thistlebottom [Ber71b]. Thomas [Bel31]. Thought [Rhe85]. Thoughts [Rol75b, Rol75c, Rol75a, Rol75d, Rol75e, Rol75f, Rol76a, Rol76b, Rol76c, Rol76d, Rol76e, Van82a]. Three [Bur87, CR83, Dem87e, Far87b, Gau67, GM77, Hal85, HW75, Hig91e, Hig92f, Le85b, Mol78, QTD91, Tre85, DD90]. Three- [HW75]. Three-Term [Gau67]. Thrombolytic [Inv93]. thru [Anoxxb]. Till [For68]. Time [And71, Anoxxb, CK75, Coo94, Cyb83, GLKH90, Gar90b, Gar90c, Gar90d, Gar91b, GL95c, GL80, MGM93, Gea77]. Tippe [Or94]. Tips [Dil93, Gli89, Gra83]. title [Ker83]. titration [Shr84]. Today [For67]. Todd [Sch77]. Toeplitz [PD92, Bar84, BW93, BBdHS95, Bre91, Bun85, Cha88, CS89a, CH92a, CNP94, Cyb80, Cyb81, Cyb84b, CV86, GS58, GN82, Ioh82, Lau88b, Lin92a, LQ87, Mac89a, Mak91, Mee80, MM88, PS92, Par86, RT92b, RS88a, RB78, Str86b, SE87, SH93, Swe84, Tyr91, Var91, Wid65]. Toeplitz-circulant [SE87]. Tolerance [BGM^{+}94]. tomorrow [Cas91]. Tongue [Bry90]. Tool [BRR85, BC86, Cow89b, Gar72, Gay88, Hig94a, DBKF90]. Toolbox [BD92, BD93b, BD94, CM93, HHKR93, The93, Hig93e, Hig95g]. TOOLPACK [Ost83, BH87, Pol89]. Toolpack/1 [BH87]. Tools [BK88, BDCS88, Cow89a, DSB88, DT93, DT94, Han92, Kel88, KP76, MS93b, Mio90, Pol89, Ral91, Rhe85, Sch90c, Woe92, Che82b, Lan89b]. Top [Anoxxa, Ash91, Gar82a, Gar90b, Gar90c, Or94]. Topics [Bar91d, Gow84, Gra87, Han69, Hay89, HJ91, Mil75a, Che86, Uhl92b]. Topology [Ash93]. Torchbearer [Tau88]. Tori [DL92b]. TOs [Jac76]. Total [Aru92, Asn70, GP92, GP93, Kar68, MD92, Van87a, VV91, Wat83b, Dem87a, GV80, OW85, VV87]. Totally [And87, dBP77, Cry73, Cry76, Kem82, Peñ95, Pin85]. Tour [Mol93]. Tower [Cha74a]. Town [Bry89]. Trace [KH95, MT93, Wat92b, And85]. Tracing [Brü95b]. Tracking [PR81, Ste90h]. Trade [Sch90c, Sto91]. Trading [Dem93b]. Tradition [Gid86]. Traditional [LYK89]. Trajectory [Cip88]. Transcendence [Bai88a]. Transcendental [FSC73, FB91, McK84]. Transfer [Ste91a, BBK89, DK83]. transfer-function [DK83]. Transform [Ros90, BF93a, BH95b, Gau69, Lin92b, Par82, Ram71, Van92]. Transform-Based [Lin92b]. Transformation [BG67, CR88, FM52, Kau87, Pry85, Rut58, Rut60, Sal89, TZ59, Tsa75, FG86, NOP84, Ruh69]. Transformations [Box66, BG65, CM79, CL72, Dem83, FF51, Gen73, Gen75, Ham82c, Mil76, PR68, PR69b, Rat82, Sch73, SV89, SW62, Wil59, Wil62, Wil65c, JS77, LHPW87, RS86]. Transforming [Cow89a, Giv58]. Transforms [GS66, ST75]. Transients [Ste94c]. Transitive [Gor84, PPY92, PPY93]. Translation [Bai93a, Bai93b, HK87]. Transmission [Wat95]. Transportable [For78]. Transpose [Fre93, TZ59]. Transpose-Free [Fre93]. transputers [Bur90]. Trapezoid [Lem75a]. Travails [Kah93b]. Travels [Sch74, Bry89, Bry91]. Treasury [Dic92]. Treatise [WR44]. Treatment [BM82, Van77a]. Treats [Cho83]. Tree [MA87, PPY93, MZ91]. Trek [Ash86b]. Trench [Cyb79]. Tri [Bus69b, Kah66a]. Tri-Diagonal [Bus69b, Kah66a]. Trial [Inv93]. Triangle [Gra92, Kah89a]. Triangular [APS93, AS93, AK75, And91a, Bar86a, Bre70b, BH74, CKS78, Cup84, EHHR88, ER74, ED83, FM85, Giv58, GK90a, HR88, Hig87c, Hig89a, HP94c, Hig95e, KP92, Lem75a, LC89, Luk86, ML85, Par74, Par76, PPY92, PA92, RO88, RG92a, SB77, Tsa91, Tsa92, Wat92e, Wil59, DGK80, Fil81, Hig83b, PW77]. Triangularization [Hou58b, LQ87]. Triangularizations [Par92a]. triangularizing [Hua75]. Tribute [Sch77]. Tricks [AFJS87, Asl93, Cho83]. Tridiagonal [Aas71, AM94, BG94, BE73b, CDW94, Dem92a, EO79, Gei91, GE95b, Hig86b, Hig90b, Kah93b, Kha72, KM86a, Lan93, Lew82, LZS91, LL93, LZ94, LPS87, MP74, Mat84b, Mee80, Meu92, Par92b, PL93, RG92b, Swa93, Tem75, Var72, Zha92, FU69, GLW89, GD89, GN82, Har86a, SK77]. Tridiagonalization [GP74, Hou58a]. Tridiagonalizing [BS92b]. Trigonometric [Lan38, Lyn87, Bul80]. Trilinear [Pan84b, Pan78]. Triple [Cle59]. Triple-Diagonal [Cle59]. Triplex [Nic70]. Triplex-ALGOL [Nic70]. trouble [Dub91b]. Troublesome [Bry87]. True [HYG88]. Trummer [Lu95]. Truncated [Han90c, CH90a, Han87, Han89, NS90]. truncated-Newton [NS90]. Truncation [Kah65, Kar74, Ros65, Wat92e, Wol64]. trust [CH90b, Gay84, MS83, Sor82]. trust-region [Gay84]. Truth [Hal95].

Trying [Sch86d]. Tunes [Hym86]. Tuning [IBM93]. tunneling [LM85]. Turing [CK84, CD86, Goo92, Hof85c, Mic80, Wil80, HAC^{+}85, ACM87, Bod49, Bri92, Hod83, HC88, Inc92, Tur59, Whi87]. Tutorial [Bri87, Goo79]. Twain [Cor94]. Twelve [Bai91]. Twentieth [MHR80]. Twenty [ACM87, Kah90d, Gar90e]. Two [AE93, AHB87, Bar91d, BB88a, tBK84, Bir75, Dav88, Dav72, Dem87b, EP87b, Gau95, HW75, Hum88, Kel93, Ker69, Knu92, Kol95, Mat90b, Mol95, Pas75, PJS92b, Ste91e, Stp93, TW91a, Var79, Wat77, Zha92, CM83, Cra86, HLPW86, Hén76, KT76, Sch68b]. Two- [HW75]. two-dimensional [Hén76]. Two-Point [BB88a, Kel93]. Two-Sided [Kol95, Sch68b]. Two-Way [Zha92]. Type [Bag94, BHH93, Ede91a, GP71, Gra83, Rub88, SGo74, Swa79, You54, Fre90, GM74c, Hua75, Jus72, Tan75, Woo89]. Typefaces [Knu86d]. Types [Ben87a]. Typesetting [BR87, BK86b, Day83a, Hoe87, KC75, Knu89b, Leh92, Mur76, Spi90, BK86a, Cas91, Fox87b, Hoe86, Jus72, Sie86]. Typewriter [Nei82, Nei84]. Typical [Bow53]. Typing [Pet86]. typographic [BCJR87]. Typography [Knu79, Mil87a, Rub88, Sas93].

UK [Hig96d]. Ultimate [Ano85d, Ano8x, B^{+}81, Won88]. Ultra [Lew78]. ULV [Ste93d]. unbalanced [Par90b]. Unbiased [Far85]. Uncertainties [Eis68, Tay82]. Uncertainty [Sig86]. Unconstrained [Buc94a, DS83, FM68a, Fle80, GM74c, KO68, MGH81a, MGH81b, Mur72, Noc92, Pow70, Pow75, Tor92, Fle90, Pow76]. Uncontrollability [EH91a]. uncoupled [MSS84, SZ91b]. Uncoupling [Mey89]. Underdetermined [AL85, CP76, DH93]. Underflow [Coo81, Dem84, FT86, Neu85]. Underflow-Free [MI81]. Undergraduate [Fra92a]. Undergraduates [Mau91].

Undersized [Sch86c]. Understand [Par94]. Understanding [Wat82, DBKF90]. Undetected [Pet86]. Unfolding [GB86]. UNICOS [Cra89]. Unidimensional [BI77]. Unified [Mar89, KT85]. Uniform [IM94, Tra66]. uniprocessor [Cal86]. Uniprocessors [IT89]. unique [Bou73a]. Uniqueness [BB95, Hau67, Cul66]. Unit [CGR90, CHH83, Sch81a, Tao84b, Rot84]. unitärer [Gab79]. Unitarily [Aru92, Mir60, HM90a]. Unitary [BM94, BGH92, EH91b, EH92, Giv58, Hou58b, Leh94b, Li95, Mat92a, Mat95b, Rin64, Sch73, Gra86, GR90, HS59, Kel75, Van87b, Mat69, Rad62]. United [Fox90a]. Uniting [Pan78]. Units [GM74a]. University [BKW90, CBB54, Hig90i, Hig90j, Hig96c, MMR78b, Wil75d, CK80a, CK80b, Har83, Var90c]. UNIX [BR87, Bou78, PM87, Lan89b, Wan88]. Unlocking [Bur89]. Unnormalized [BZ90]. Unraveling [BM92]. Unrolling [DH79]. Unseen [Pic90]. Unstable [Bye88, FTxx, LQ87, Wri93, Van85a]. Unstructured [Dax90a]. Unsymmetric [Duf81, DR93, DS93b, EGL^{+}87, GV79, Hen92, HV94, LZ91, Saa80]. Untersuchungen [Spi76]. Unusual [Hau91]. Unzalässige [PN85]. up- [BPT90]. Update [Bar93a, ES92, Wil89, KBS90]. updated [Ano8x]. Updates [FM84, SB92, Ste94e, DS79a, GS78, Ip87a, KT85, Sch83b]. Updating [Cup84, DGKS76, FM85, FP86b, FP86a, Hag89, Pow87b, Ste92a, Ste93d, BS89b, Pai80, Pow87a, Ste90h]. Upon [Led87]. Upper [Hig83b, Lás91, Lee95, Sha91, AL75]. URV [AGS91, PE95, Ste91d]. Usage [Ber65, Ber71b, Ber77a, Fow68, HK87, LHKK79, Par73, Saf90, Swa80a, Woo62, WFF89]. Use [ADD89a, And91a, Ber75, Bli72, Box66, BH87, Buz82, DD89, DD90, For57a, GJM87, Gar91a, GH70, Ham82c, Hig90g, Laz88, LS91, Lyn83, Man82, MS79b, New81, Qui68, QS90, Rob80, Und80, Van87a, Wic92, WWG51, Wil55b, Wil65c, Wil77c, Woo58, Zla82, CH80, DK83, DGR79, DNR87, KT85, Sie86]. Used [Bar92a, Pét66, Tsi90]. Useful [Dic92, McC87]. useless [McC87]. User [Ame91, ABB^{+}92, ABB^{+}95, CS87, DBMS79, DPW95, DMR92, DGL92, GHM^{+}86, GMSW86c, Hew93, JW74, JW85b, Kro92, Kro94, Lam86b, Lam94, Lon67, Mat92c, MLB87a, MLB87b, Rob88, Den84, DLS83, Mol82b, Ano86a, CM93, Sun92c]. Users [Gro95a, Hig94a, KD93, KD95, Und80]. Uses [Bat83, Gol73b, Rat85, Sta80, Wil55a]. Using [AS92, Bai88b, BLS91, BSD60, BCJR87, BP94c, BH95a, Bre70b, BLV85, CH89, DV92b, DRW93, Don95, FP74a, FGM95, Fos86, GH93, Gar94b, Gay87, GHR67, Gla93, Gli89, GM86, Gul95, Hil89, HFT94, JW85a, KBS94, Lad76, Leh92, Van82b, Van84b, Lon84, Mac90, Mil88a, Par92a, Pol89, Qui94, Ros65, Sau72, Sch79, SHW86, Ste90a, Sti91, Sto75b, Sto77, Wat92f, AGS91, Bis89a, CMR90, DEG72, DLR81, DH86, Mao86, MR89, RB85, Sau94, Sch83b, Sch83c, Sie87, Sny55, Thi86, VV87, VS88a, HP88].

v [Hig96b, CGG^{+}91, Kri80, PM87]. V1.0 [DPW95]. Vacuum [Asl82]. Validating [Kah91c]. Validating/Testing/Confirming/Checking/Correcting [Kah91c]. Validation [Ano86a]. Value [AK75, AMR88, BD93a, BR77, Bar86b, BZY93, BS89a, Ber92, Bha93, BD86, BLV85, BGG88, Cas89, CI94, DDLT91, EI93, Enr89, EP87b, GDD94, GE95c, HN81, He90, HP93, HP94a, HP91, HP92b, JS92, Kel93, KL80, Lam91, LS91, Van85b, Van86a, MS93a, MM83b, PS81, PS94, RM67, Sch86c, Ste83b, Ste93b, Var75, Wat92d, Wed72, Wil78b, AP76, BV86, Bis89b, BGBMN90, CH90a, Chu86, FT88, FU69, GR70, Ham85, Han90c, HLPW86, Kåg84, Kub70, Van76a, Luk80, Man82, Mao86, MG89, MG91, MZ91, Pai85b, Pai86, PV86, Sch91, Ste84a, Ste90f, Tog86, Tsi90, Wat90]. valued [BGBMN90, Mat91a]. Values [And90b, BH95a, CI95a, DK90, DG93, Don83, FP92a, Giv54, GK65, Gol68, GLO81, Hen62a, Jen82, Joh78, Kit95, LPP93, Luk86, Par86, Qi84, vdSvdV87, Dem88a, FH60, Hor54, HC63, Mat90b, Ruh75, SV79, Ste79b, Ste91d, Uhl85]. Vampire [Gor84]. Vancouver [Gau94, Jam74]. Vandermonde [CGR90, Bar91d, BH92, BP70, BE73a, BKO94, CR92, CR93, Dem89a, DS89, Dem90a, EP94a, FH93, Fre80, GP70, GP71, Gau62, Gau63, Gau75a, Gau75b, Gau78a, Gau83a, GI88, Gau90, GK89, GK90a, GO94, HG80, Hig87b, Hig88b, Hig90k, JS77, Lu94, Lu95, Lu96, RO91, Rei91, TG81, Tom92, TS75, Var90a, Var91, Van77a, VS88b]. Vandermonde-Like [Bar91d, CR93, Gau83a, Lu96, Rei91, BH92, Hig88b, Hig90k]. Variable [DM90, Hen62b, Kie81, Lev58, Ziv83, MS86]. variable-complexity [MS86]. Variable-Precision [Kie81]. Variables [Ash93, Löt84, Ste90e, VV87]. Variance [Bar91a, CGL83]. Variances [Bur74]. Variant [DHSS94, GS76b, vdV92, Thi86]. Variante [Rut60]. Variants [DS94, PH72, AD90a]. Variation [BM94, Hen62a, Hol92]. Variational [Pre75, Fle89]. Variations [Gil91a, Ker92, Saa80]. Various [Bre73b, DV92b, Don95]. Varying [Dem87d, Tie70]. Vec [HS81b]. Vec-Permutation [HS81b]. Vector [And90a, Blu78, DGK84, DDGM89, DDSv91, GHR67, Hal74c, Hof87, IT89, MRK76, O’L93a, OV85, Ort88, OVR90, Phi87, Rob90, SSSU87, SB88, Vet75, War76, Bis89b, DD90, DGR79, FM86, Kub65, MS88]. Vector/Parallel [MRK76]. Vectorization [PV94, vdVD89]. Vectorizing [Wee89]. Vectors [Bar93b, BN87b, FF51, GL95b, GLO81, Mac43, Phi87, Pry84, Ruh83, Wil55b, Ste84b]. Vegetarian [Bow90]. Velocity [LS83]. Verbs [Kin67c]. Verfahren [Neu74]. Verifiable [Mil88a]. Verified [HHKR93]. Verifier [Ryd74]. Verify [Rol76e]. Vermutung [PN85]. Versa [Gar82b]. Versatile [Cra79, Hig94a]. Version [Ame91, AR94, Bai94a, GHM^{+}86, GMSW86c, Hig95d, Hig95g, IBM93, Ken93a, Ken93b, LIA93, MT94, MT95, PFTV86, Sta86]. Verstärkungen [PN85]. Versus [Bag94, Ehr70, Kah86a, Kah94a, Wer84]. Very [Dav75, GS90b, JSW83, Nas85, Wil72c]. Veteran [Hom92]. VF [DEKV92, DD89, KL89a]. VF/600J [DEKV92]. Via [DG87, AS92, All88, CW87, DL93, DL94, EK89, HP93, Mat95c, Par82, QCM94]. Vice [Gar82b]. View [Cri74, Cri89, Lon67, Pow75, Fre80, Pow76]. Viewpoint [Roj94]. Vintage [Anoxxd, Lav80]. Virgule [Pic72]. Virtual [CNRT81]. Virtue [Rol76a]. Visual [Gou73, Mat90d, Tuf83, Lam87a]. Visualization [SBKZ93]. Voice [Ber75]. Vol [Tod77]. Volume [Ano85a, Ano85b, Ano85d, Anoxxb, BP94b, Caj29, CW85, Fle80, Fle81b, IT89, KvSR90, Knuta, Knu81b, Wil71c, Wil71d, Hig96d, ACD^{+}93]. Volumes [Abb78]. Vortex [BKP93]. vs [Tho83, Tho88]. VTEX [Cas91].

w [Wil67a]. W. [Kah73b]. Wainwright [Wai85]. Walk [Web73]. Walker [SD86]. Walks [Web82]. Wannier [CK57]. Want [Hal85, Har91b, Rol75f]. Wanted [Ors69]. War [Goo92, Lew78]. Wars [Wil88a]. Watchers [Moo93]. WATFOR [CS87]. WATFOR-77 [CS87]. Watt [Gen69]. Way [Bry90, Gor93a, Mer90, Nix81, Zha92, Wai85]. Ways [Bai91, MV78, Har81a, How83a, Vas90]. Weak [Bun87]. Weather [Ing95]. WEB [Thi86, BKM86]. Webster [Mer83, Mer93, Mer91, Neu88]. Weighted [BH88, Gul92, GW92b, Gul94, Gul95]. Weighting [CR88, Van85c, Van83a]. Weights [KE82]. Weird [Dic92]. Weitek [Tho88]. Well [CF88, Gil87, Gor93b, Mit87, Ogr88, Uhl92a, Zin88a, Zin90, Ste81b]. Well-Conditioned [CF88, Uhl92a]. Well-Tempered [Gor93b]. Where [Rol75a, Har80]. Which [Asp59, Bus68, DL92a, Dav92, Dav93a, Hig94e, Nab92, Wat88a, Wri93, AL73, Dem87a]. Whips [Ren72]. Who [Lav80, McC79, Reg87, Roj94, Rol75a, Ste80a]. Whole [Kro92, Kro94]. Whose [DRMN79, Peñ95]. wider [Sie86]. Wielandt [Fra58, Ips94a]. Wild [Cip88]. Wiley [Hig95d]. Wilkinson [Asp81, Eva76, Ano78a, Ano87, Fox78, Fox87a, Fox90b, Gol90, Par90c, PC83, PN85]. Will [Gar90b, Tur91b, Ros90]. Windowing [Ste92a]. Winning [Abb78]. Winnipeg [HW76c]. Winograd [Par82, Bre70b, DHSS94, Har72]. Wisconsin [MMR78b]. Wisconsin-Madison [MMR78b]. Wise [Bjö91a, Grc92]. Wishart [Ede91a]. Within [EV85, Hig90e, NS90]. Without [Bre73a, CG85, DG94, Fle65, Fos86, FP81, Gen73, Sch86d, SW81, Sco81, She94, TTRD93, YC95, GN82, Olv83b, Sch65]. Wobbling [Gor93a]. Wolfe [Ano83]. Wolfram [Web92b]. Wonderful [Dic92]. Woodbury [Yip86]. Word [Aug84, Buc81, CCG^{+}84, Dic92, Man90, Zin83]. Word-Length-Independent [CCG^{+}84]. WordPerfect [FF87a]. Wordprocessors [SoIPUG86, SoIPUG87]. Words [Bry87, Dic92, Gor93a, Gow73, Gow86, Mal89, Man88, III84, Sch87, SGo74, Tul91, Van87d, BKM86, Ock75, Pri91b]. Wordstar [FF87a]. Work [Act70, Bar82, Dud77, Goo92, Ham94, Jac76, Lam86a, Lit78, SW89, Sim87, Wil80, WH16]. Workbench [MFGK82]. Workpieces [For92]. Works [Bri92, Gar73a, Gar73b, Inc92, Tau63]. Workshop [HHK93, Hig92c]. Workstation [BH87]. World [AJ88, DF87, Lee88, Lod84, Mic92b, Mol90a, Neu88, Pic90, Ber80, Goo92]. worth [Dub91b]. Would [Kah93b]. Would-be [Kah93b]. Wprowadzenie [KS92]. Write [Bar78, Bar82, Bie37, Day77, Day83b, Day88, Gra70, Hal70, Lam95, Mic90, Ste73a, SHSD73, TS80, Wil84a, TE87]. Writer [Ber65, Ber71b, CM87, Cor94, GP59, Hoo80, McC78, Mic92a, Mit87, Sch90c, You89b, MFGK82]. Writers [Mer91, GH43, Lon93, MS89, Tar92, Tur87, Sch90c]. Writing [Bak61, Bar78, Bar82, Bee92, Ben82, BKW90, Bol88, Bro79, Che82b, Coo64, Cow89b, EBR87, GS76a, Gil87, Gil91b, GS90c, Gre79, Hig92a, Hig93b, Hig94b, Hwa95, Jac76, Kel91, Kin67a, Kin67b, Kin67c, Kin67d, KLR89, L^{+}64, Lin84b, Mau91, Mic92a, MS89, Mit87, OW77, O’C91, Pec87, Per91, Pét66, Rat85, Rol67a, Rol67b, Rol75b, Rol75c, Rol75a, Rol75d, Rol75e, Rol75f, Rol76a, Rol76b, Rol76c, Rol76d, Rol76e, SG83b, SG83a, Ste90a, Swa71, Trz93, TK89, Tur74, WPG92, Woo67, Zin83, Zin88b, Zin88a, Zin90, Boo84, Boo93, Kir92, Rol76d, Woo92, You89b]. Writings [Jef60, Lem61]. Written [But92b, Kah83a]. Wrong [Mer88]. Wrongs [Jon74]. WY [BV87, Pug92, SV89].

X [DH86]. X-MP-4 [DH86]. XBibTEX [KB92]. Xnetlib [DRW93]. XSC [HHKR93, KKN^{+}92, KKL^{+}93]. XXIX [Rol75a]. xxvi [Hig96b]. XXVII [Rol75b]. XXVIII [Rol75c]. XXX [Buc94a, Rol75d]. XXXI [Rol75e]. XXXII [Rol75f]. XXXIII [Rol76a]. XXXIV [Rol76b]. XXXV [Rol76c]. XXXVI [Rol76d]. XXXVII [Rol76e].

Y-MP [Car91, Pap92]. Year [QTD91, Sim87]. Years [ACM87, Sch81c, EG91]. Yes [And91b]. Yielding [Kub91]. York [Hig95d]. Young [GHS87, Hal91b]. You’re [Fey85].

Zaokragleń [Wil67a]. Zaphod [Gra92]. Zautomatyzowanych [KS92]. Zeitgewinns [Spi76]. Zen [BR93]. Zero [Bre76, Dek69, Le85b]. Zero-Finding [Bre76]. Zerofinding [JT75]. Zeros [Kit95, Mil71a, Ste71a, Fre79]. zur [Alb77, Hil94, Neu74, Spi76]. Zwei [PN85].

References

Aarnio:1990:SHD

[A^{+}90] P. Aarnio et al. Study of hadronic decays of the Z^{0} boson. Physics Letters B, 240(1,2):271–282, 1990. CODEN PYL-BAJ. ISSN 0370-2693 (print), 1873-2445 (electronic). DELPHI Collaboration.

Albers:1985:MPP

[AA85] D. J. Albers and G. L. Alexanderson, editors. Mathematical People: Profiles and Interviews. Birkhäuser Boston Inc., Cambridge, MA, USA, 1985. ???? pp.

Adams:1994:TDC

[AA94] Loyce Adams and Peter Arbenz. Towards a divide and conquer algorithm for the real nonsymmetric eigenvalue problem. SIAM J. Matrix Anal. Appl., 15(4): 1333–1353, 1994.

Albers:1987:IMC

[AAR87] Donald J. Albers, G. L. Alexanderson, and Constance Reid, editors. International Mathematical Congress: An Illustrated History, 1893–1986. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96479-7. 64 pp.

Albers:1990:MMP

[AAR90] D. J. Albers, G. L. Alexanderson, and C. Reid, editors. More Mathematical People: Contemporary Conversations. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1990. ???? pp.

Aasen:1971:RSM

[Aas71] Jan Ole Aasen. On the reduction of a symmetric matrix to tridiagonal form. BIT, 11:233–242, 1971. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Aggarwal:1979:REM

[AB79] Vijay B. Aggarwal and James W.

Burgmeier. A round-off error model with applications to arithmetic expressions. SIAM Journal on Computing, 8(1): 60–72, 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abbott:1978:CPC

[Abb78] J. C. Abbott, editor. The Chauvenet Papers: A Collection of Prize-Winning Expository Papers in Mathematics, Volumes 1 and 2. Mathematical Association of America, Washington, DC, USA, 1978. ???? pp.

Anderson:1990:LPL

[ABB^{+}90] E. Anderson, Z. Bai, C. H. Bischof, J. W. Demmel, J. J. Dongarra, J. J. Du Croz, A. Greenbaum, S. J. Hammarling, A. McKenney, and D. C. Sorensen. LA-PACK: A portable linear algebra library for high-performance computers. LA-PACK Working Note #20, Department of Computer Science, University of Tennessee, ????, 1990. ???? pp.

Anderson:1992:LUG

[ABB^{+}92] E. Anderson, Z. Bai, C. H. Bischof, J. W. Demmel, J. J. Dongarra, J. J. Du Croz, A. Greenbaum, S. J. Hammarling, A. McKenney, S. Ostrouchov, and D. C. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-294-7. xv + 118 + listings pp.

Anderson:1995:LUG

[ABB^{+}95] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1995. ISBN 0-89871-345-5 (paperback). xix + 325 pp. LCCN QA76.73.F25 L36 1995.

Abdelmalek:1971:REA

[Abd71] Nabih N. Abdelmalek. Round off error analysis for Gram–Schmidt method and solution of linear least squares problems. BIT, 11:345–368, 1971. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Abdelmalek:1974:SLL

[Abd74] Nabih N. Abdelmalek. On the solution of the linear least squares problems and pseudo-inverses. Computing, 13:215–228, 1974. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Anderson:1990:PEL

[ABD^{+}90] E. Anderson, C. H. Bischof, J. W. Demmel, J. J. Dongarra, J. J. Du Croz, S. J. Hammarling, and W. Kahan.

Prospectus for an extension to LAPACK: A portable linear algebra library for highperformance computers. LAPACK Working Note #26, Department of Computer Science, University of Tennessee, ????, 1990. ???? pp.

Aberth:1988:PNA

[Abe88] Oliver Aberth. Precise Numerical Analysis. Wm. C. Brown Publishers, Dubuque, IA, USA, 1988. ISBN 0-697-06760-2. x + 225 pp.

Alpher:1948:OCE

[ABG48] R. A. Alpher, H. Bethe, and G. Gamow. The origin of chemical elements. Physical Review, 73(7):803–804, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Abikoff:1986:T

[Abi86] W. Abikoff. TEX. The Mathematical Intelligencer, 8:64–76, 1986. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Ahac:1988:SFM

[ABO88] Alan A. Ahac, John J. Buoni, and D. D. Olesky. Stable LU factorization of H-matrices. Linear Algebra and Appl., 99:97–110, 1988.

Abaffy:1984:CDM

[ABS84] J. Abaffy, C. G. Broyden, and E. Spedicato. A class of direct methods for linear systems. Numer. Math., 45:361–376, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ashford:1993:CPL

[ACD^{+}93] Oliver M. Ashford, H. Charnock, P. G. Drazin, J. C. R. Hunt, P. Smoker, and Ian Sutherland, editors. Collected Papers of Lewis Fry Richardson. Volume 1: Meteorology and Numerical Analysis. Cambridge University Press, ????, 1993. ISBN 0-521-38297-1. xiv + 1016 pp.

Ackenhusen:1989:DBM

[Ack89] J. G. Ackenhusen. Data base management system for conference program preparation. IEEE Trans. Prof. Commun., PC-32:194–199, 1989.

ACM:1987:ATA

[ACM87] ACM Turing Award Lectures: The First Twenty Years, 1966–1985. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-54885-2, 0-89791-183-0. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Acton:1970:NMW

[Act70] Forman S. Acton. Numerical Methods That Work. Harper & Row, New York, NY, USA, 1970. ISBN 0-88385-450-3. xviii + 541 pp. Reprinted by Mathematical Association of America, Washington, DC, with new preface and additional problems, 1990.

Acton:1974:RRF

[Act74] F. S. Acton. Recurrence relations for the fresnel integral ∫_{0}^{∞} exp(−ct)dt/\sqrt{t(1 + t^{2})} and similar integrals. Comm. ACM, 17: 480–481, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1994:AN

[act94] Acta Numerica 1994. Cambridge University Press, ????, 1994. ISBN 0-521-46181-2. 572 pp. LCCN QA297 .A21 1994.

Anonymous:1995:AN

[act95] Acta Numerica 1995. Cambridge University Press, ????, 1995. ISBN 0-521-48255-0. 491 pp.

Anderson:1989:RIR

[AD89] E. Anderson and J. J. Dongarra.

Results from the initial release of LA-PACK. Technical report, LAPACK Working Note #16, Department of Computer Science, University of Tennessee, Knoxville, ????, 1989. ???? pp.

Anderson:1990:EBA

[AD90a] E. Anderson and J. J. Dongarra.

Evaluating block algorithm variants in LAPACK. Technical report, LAPACK Working Note #19, Department of Computer Science, University of Tennessee, Knoxville, ????, 1990. ???? pp.

Anderson:1990:IGL

[AD90b] E. Anderson and J. J. Dongarra. Implementation guide for LAPACK. Technical report, LAPACK Working Note #18, Department of Computer Science, University of Tennessee, Knoxville, ????, 1990. ???? pp.

Adams:1967:SCP

[Ada67] Duane A. Adams. A stopping criterion for polynomial root finding. Comm. ACM, 10(10):655–658, 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Amestoy:1989:ULB

[ADD89a] P. R. Amestoy, M. J. Daydé, and I. S. Duff. Use of level 3 BLAS kernels in the solution of full and sparse linear equations. Technical report, Report TR 89/9, CERFACS, Toulouse, France, ????, 1989. ???? pp.

Arioli:1989:SSL

[ADD89b] M. Arioli, J. W. Demmel, and I. S. Duff. Solving sparse linear systems with sparse backward error. SIAM J. Matrix Anal. Appl., 10(2):165–190, 1989.

Arioli:1989:ASA

[ADdR89] M. Arioli, I. S. Duff, and P. P. M. de Rijk. On the augmented system approach to sparse least-squares problems. Numer. Math., 55:667–684, 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1992:BPM

[ADNR92] M. Arioli, I. S. Duff, J. Noailles, and D. Ruiz. A block projection method for sparse matrices. SIAM J. Sci. Stat. Comput., 13(1):47–70, 1992.

Arioli:1992:SCI

[ADR92] Mario Arioli, Iain S. Duff, and Daniel Ruiz. Stopping criteria for iterative solvers. SIAM J. Matrix Anal. Appl., 13(1):138–144, 1992.

Almacany:1984:DCA

[ADW84] M. Almacany, C. B. Dunham, and J. Williams. Discrete Chebyshev approximation by interpolating rationals. IMA J. Numer. Anal., 4:467–477, 1984.

Andersson:1993:TCP

[AE93] Lars-Erik Andersson and Tommy Elfving. Two constrained Procrustes problems. Report LiTH-MAT-R-1993-39, Department of Mathematics, Linköping University, Sweden, December 1993. ???? pp.

Angermeyer:1987:TMM

[AFJS87] J. Angermeyer, R. Fahringer, K. Jaeger, and D. Shafer. Tricks of the MS-DOS Masters. Howard W. Sams and Co., Inc., Indianapolis, IN 46268, USA, 1987. ???? pp.

Arbenz:1988:SDH

[AG88] P. Arbenz and G. H. Golub. On the spectral decomposition of hermitian matrices modified by low rank perturbations with applications. SIAM J. Matrix Anal. Appl., 9:40–58, 1988.

Almasi:1989:HPC

[AG89] G. S. Almasi and A. Gottlieb.

Highly Parallel Computing. Benjamin/ Cummings Pub. Co., Menlo Park, CA, USA, 1989. ???? pp.

Allgower:1990:CSN

[AG90] E. L. Allgower and K. Georg, editors.

Computational Solution of Nonlinear Systems of Equations, volume 26 of Lectures in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1990. ISBN 0-8218-1131-2. xix + 762 pp. LCCN QA372 .C6374 1990.

Arbenz:1992:QLA

[AG92] P. Arbenz and G. H. Golub. QR-like algorithms for symmetric arrow matrices. SIAM J. Matrix Anal. Appl., 13(2):655– 658, 1992.

Agarwal:1993:CNM

[Aga93] R. P. Agarwal, editor. Contributions in Numerical Mathematics. World Scientific, Singapore, 1993. ISBN 981-02-1437-5. 475 pp.

Arbenz:1988:RRM

[AGG88] P. Arbenz, W. Gander, and G. H.

Golub. Restricted rank modification of the symmetric eigenvalue problem: Theoretical considerations. SIAM J. Matrix Anal. Appl., 104:75–95, 1988.

Ashcraft:1995:ASI

[AGL95] Cleve Ashcraft, Roger G. Grimes, and John G. Lewis. Accurate symmetric indefinite linear equation solvers. Manuscript, September 1995.

Adams:1991:DAE

[AGS91] G. Adams, M. F. Griffin, and G. W. Stewart. Direction-of-arrival estimation using the rank-revealing URV decomposition. Technical report, Technical Report CS-TR 2640, Department of Computer Science, University of Maryland, ????, 1991. ???? pp.

Alefeld:1983:IIC

[AH83] Göltz Alefeld and Jürgen Herzberger. Introduction to Interval Computations. Academic Press, New York, 1983. ISBN 0-12-049820-0. xviii + 333 pp.

Arun:1987:LSF

[AHB87] K. S. Arun, T. S. Huang, and S. D. Blostein. Least-squares fitting of two 3-D points sets. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-9(5):698–700, 1987. CODEN IT-PIDJ. ISSN 0162-8828.

Aho:1974:DAC

[AHU74] A. V. Aho, J. E. Hopcroft, and J. D. Ullman. The Design and Analysis of Computer Algorithms. Addison-Wesley, Reading, MA, USA, 1974. ???? pp.

Aiken:1985:SC

[Aik85] Richard C. Aiken, editor. Stiff Computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1985. ISBN 0-19-503453-8. xiv + 462 pp.

Appel:1988:WFS

[AJ88] A. W. Appel and G. J. Jacobson. The world’s fastest Scrabble program. Comm. ACM, 31:572–578, 585, 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:1975:CBL

[AK75] Ned Anderson and Ilkka Karasalo. On computing bounds for the least singular value of a triangular matrix. BIT, 15:1–4, 1975. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Axelsson:1990:PCG

[AK90] Owe Axelsson and Lily Yu. Kolotilina, editors. Preconditioned Conjugate Gradient Methods: Proceedings of a Conference Held in Nijmegen, the Netherlands, June 19–21, 1989, volume 1457 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1990. ISBN 0-387-53515-2 (New York), 3-540-53515-2 (Berlin). 196 pp. LCCN QA3 .L28 no. 1457.

Aho:1988:APL

[AKW88] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. The AWK Programming Language. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-07981-X. x + 210 pp. LCCN QA76.73.A95 A35 1988. See also [?, ?, ?].

Anderson:1973:QCA

[AL73] P. Anderson and G. Loizou. On the quadratic convergence of an algorithm which diagonalizes a complex symmetric matrix. J. Inst. Maths Applics, 12:261– 271, 1973.

Aird:1975:CAU

[AL75] T. J. Aird and R. E. Lynch. Computable accurate upper and lower error bounds for approximate solutions of linear algebraic systems. ACM Transactions on Mathematical Software, 1:217– 231, 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Arnold:1984:GEA

[AL84] William F. Arnold and Alan J. Laub. Generalized eigenproblem algorithms and software for algebraic Riccati equations. Proc. IEEE, 72(12):1746–1754, 1984. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Arioli:1985:EAA

[AL85] M. Arioli and A. Laratta. Error analysis of an algorithm for solving an underdetermined linear system. Numer. Math., 46:255–268, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1986:EAA

[AL86] M. Arioli and A. Laratta. Error analysis of algorithms for computing the projection of a point onto a linear manifold. Linear Algebra and Appl., 82:1–26, 1986.

Albrecht:1977:BIB

[Alb77] J. Albrecht. Bemerkungen zu iterationsverfahren zur berechnung von A^{1/2} und A^{−1}. Z. Angew. Math. Mech, 57: T262–T263, 1977.

Albers:1991:PHC

[Alb91] Donald J. Albers. Paul Halmos by parts (interviews by Donald J. Albers). In Ewing and Gehring [EG91], pages 3– 32. ISBN 3-540-97509-8. LCCN QA300.5 .H35 1991.

Allen:1983:PEN

[All83] D. M. Allen. Parameter estimation for nonlinear models with emphasis on compartmental models. Biometrics, 39:629– 637, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Allwright:1988:PSM

[All88] J. C. Allwright. Positive semidefinite matrices: Characterization via conical hulls and least-squares solution of a matrix equation. SIAM Journal on Control and Optimization, 26(3):537–556, 1988.

Ashenhurst:1965:EEC

[AM65] R. L. Ashenhurst and N. Metropolis.

Error estimation in computer calculation. Amer. Math. Monthly, 72(2):47–58, 1965.

Althoen:1987:GJR

[AM87] Steven C. Althoen and Renate Mclaughlin. Gauss–Jordan reduction: A brief history. Amer. Math. Monthly, 94 (2):130–142, 1987.

Amodio:1994:BEA

[AM94] Pierluigi Amodio and Francesca Mazzia. Backward error analysis of cyclic reduction for the solution of tridiagonal systems. Math. Comp., 62(206):601–617, 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Physics:1965:SM

[Ame65] American Institute of Physics. Style Manual. American Institute of Physics, ????, revised second edition, 1965. ???? pp.

Ames:1977:NMP

[Ame77] William F. Ames. Numerical Methods for Partial Differential Equations.

Academic Press, New York, second edition, 1977. ISBN 0-12-056760-1. xiv + 365 pp.

AMS:1991:AVU

[Ame91] American Mathematical Society. AMS-LaTEX Version 1.1 User’s Guide. American Mathematical Society, Providence, RI, USA, 1991. ???? pp.

Ascher:1988:NSB

[AMR88] Uri M. Ascher, Robert M. M. Mattheij, and Robert D. Russell. Numerical Solution of Boundary Value Problems for Ordinary Differential Equations. Prentice-Hall, Englewood Cliffs, NJ, USA, 1988. ISBN 0-13-627266-5. ???? pp.

Amir-Moez:1974:SRP

[AMS74] A. R. Amir-Moéz and S. Smyrl. Square roots of a positive matrix and polar decompositions. Indust. Math., 24(1): 55–60, 1974.

Ahlberg:1963:CPS

[AN63] J. H. Ahlberg and E. N. Nilson. Convergence properties of the spline fit. J. Soc. Indust. Appl. Math., 11(1):95–104, 1963.

Anderson:1971:SAT

[And71] T. W. Anderson. The Statistical Analysis of Time Series. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1971. ISBN 0-471-02900-9. xiv + 704 pp.

Ando:1985:ATN

[And85] T. Ando. Approximation in trace norm by positive semidefinite matrices. Linear Algebra and Appl., 71:15–21, 1985.

Ando:1987:TPM

[And87] T. Ando. Totally positive matrices. Linear Algebra and Appl., 90:165– 219, 1987.

Anderson:1990:RNG

[And90a] S. L. Anderson. Random number generators on vector supercomputers and other advanced architectures. SIAM Review, 32:221–251, 1990.

Andrew:1990:ESV

[And90b] Alan L. Andrew. Eigenvalues and singular values of certain random matrices. J. Comp. Appl. Math., 30:165–171, 1990.

Anderson:1991:RTS

[And91a] Edward Anderson. Robust triangular solves for use in condition estimation. Technical Report CS-91-142, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, August 1991. 35 pp. LAPACK Working Note 36.

Anderson:1991:YM

[And91b] Sylvia Anderson. Yes, M’Lady.

Smith Gryphon, London, 1991. ???? pp.

Anholt:1994:DSA

[Anh94] Robert R. H. Anholt. Dazzle ’em With Style: The Art of Oral Scientific Presentation. W.H. Freeman, San Francisco, CA, USA and New York, NY, USA, 1994. ISBN 0-7167-2583-5. xiii + 200 pp.

Anonymous:1904:CMP

[Ano04] Anonymous, editor. The Collected Mathematical Papers of James Joseph Sylvester, volume 1 (1837–1853). Cambridge University Press, ????, 1904. xii + 650 pp.

Anonymous:1922:CC

[Ano22] Anonymous. Le Commandant Cholesky. Bulletin Géodésique (Toulouse), pages 159–161, 1922. CODEN BGDQAG. ISSN 0007-4632. Translation by Richard W. Cottle (“Major Cholesky”) appears in [Mai84, Appendix] and in NA-Digest, Volume 90, Issue 7, 1990.

Anonymous:1924:GFN

[Ano24] Anonymous. Gleanings far and near. Mathematical Gazette, 22(170):95, 1924. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:1946:MOA

[Ano46] Anonymous, editor. A Manual of Operation for the Automatic Sequence Controlled Calculator. Harvard University Press, Cambridge, MA, USA, 1946. ISBN 0-262-01084-4. Reprinted, with new foreword and introduction, Volume 8 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

Anonymous:1948:PSL

[Ano48] Anonymous, editor. Proceedings of a Symposium on Large-Scale Digital Calculating Machinery, volume 16 of The Annals of the Computation Laboratory of Harvard University. Harvard University Press, Cambridge, MA, USA, 1948. ISBN 0-262-08152-0. Reprinted, with a new introduction by William Aspray, Volume 7 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

Anonymous:1976:BGG

[Ano76] Anonymous. The Best of George Gershwin. Chappell and Co. Ltd., International Music Publications, ????, 1976. ???? pp.

Anonymous:1978:CDJ

[Ano78a] Anonymous, editor. The Contribution of Dr. J. H. Wilkinson to Numerical Analysis, Symposium Proceedings Series No. 19. The Institute of Mathematics and its Applications, 1978.

Anonymous:1978:LAA

[Ano78b] Anonymous. Leroy Anderson (almost complete) 25 Great Melodies as originally composed for Piano Solo. Belwin-Mills, ????, 1978. ???? pp.

Anonymous:1978:NSP

[Ano78c] Anonymous. Next slide please. Nature, 272(5656):743, 1978. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1979:LFB

[Ano79] Anonymous. The Legal Fake Book.

Warner Bros, ????, revised edition, 1979. ISBN 0-89898-201-4. ???? pp.

Anonymous:1983:RWL

[Ano83] Anonymous. Richard Wolfe’s Legit Professional Fake Book. Columbia Pictures Publications, Florida, sixth edition, 1983. ISBN 0-89898-201-4. ???? pp.

Anonymous:1984:FSA

[Ano84] Anonymous. Frank sinatra—L.A. is my lady, with quincy jones and orchestra. Technical report, Warner Bros., ????, 1984. ???? pp.

Anonymous:1985:CCVa

[Ano85a] Anonymous. Carpenters Complete, Volume 1. Columbia Pictures Publications, ????, 1985. ISBN 0-89898-362-2. ???? pp.

Anonymous:1985:CCVb

[Ano85b] Anonymous. Carpenters Complete, Volume 2. Columbia Pictures Publications, ????, 1985. ISBN 0-89898-363-0. ???? pp.

Anonymous:1985:LRL

[Ano85c] Anonymous. Linda Ronstadt—Lush Life, What’s New. Warner Bros., ????, 1985. ???? pp.

Anonymous:1985:UFB

[Ano85d] Anonymous. The Ultimate Fake Book, Volume 2. Hal Leonard, ????, 1985. ???? pp.

Anonymous:1986:FAF

[Ano86a] Anonymous. FPV: A floatingpoint validation package. Release I. User’s guide. Technical Report NP 1201, NAG Ltd., Oxford, May 1986. ???? pp.

Anonymous:1986:LRN

[Ano86b] Anonymous. Linda Ronstadt with Nelson Riddle and his Orchestra, For Sentimental Reasons. Warner Bros., ????, 1986. ISBN 0-88188-284-4. ???? pp.

Anonymous:1987:JW

[Ano87] Anonymous. James Wilkinson (1919– 1986). Annals of the History of Computing, 9(2):205–210, 1987. From the introduction: “A series of lightly edited extracts from messages that were sent over various computer networks during the period October 5, 1986–February 13, 1987”.

Anonymous:1988:NRB

[Ano88] Anonymous. The New Real Book. Sher Music, P.O. Box 445, Petaluma, California 94953, 1988. ISBN 0-9614701-4-3. ???? pp.

Anonymous:198x:UBF

[Ano8x] Anonymous. The Ultimate Broadway Fake Book (updated). Hal Leonard, Winona, MN, 198x. ISBN 0-88188-284-4. ???? pp.

Anonymous:1991:EN

[Ano91] Anonymous. Editor’s note. SIAM J.

Matrix Anal. Appl., 12(3):??, 1991.

Anonymous:1993:KSL

[Ano93] Anonymous. Kent State launches electronic numerical analysis journal. SIAM News, 26(6):20, October 1993. ISSN 0036-1437.

Anonymous:19xx:PTL

[Anoxxa] Anonymous. 500 Plus Top Line Greats. Chappell, ????, 19xx. ???? pp.

Anonymous:19xx:ATE

[Anoxxb] Anonymous. All Time Enduring Standards, Volume 2 of the T.P.A. Library, 1930s thru 1950s. ????, 19xx. ???? pp.

Anonymous:19xx:OLM

[Anoxxc] Anonymous. The Original, Legal, Musicians’ Fake Book. Hansen House, 1860 West Avenue, Miami Beach, Florida 33139, 19xx. ISBN 0-8494-0001-5. ???? pp.

Anonymous:19xx:VJS

[Anoxxd] Anonymous. Vintage Jazz Standards, Book 1. ????, 19xx. ???? pp.

Anthony:1985:APS

[Ant85] G. T. Anthony. On an algorithm for Pythagorean sums and its connection with continued fractions. IMA Bulletin, 21(9/10):171–172, 1985.

Anton:1994:ELAa

[Ant94] Howard Anton. Elementary Linear Algebra. Wiley, New York, NY, USA; London, UK; Sydney, Australia, seventh edition, 1994. ISBN 0-471-58742-7. xvi + 570 + Appendix + Index pp.

Ahac:1986:SMF

[AO86] Alan A. Ahac and D. D. Olesky. A stable method for the LU factorization of M-matrices. SIAM J. Alg. Disc. Meth., 7 (3):368–378, 1986.

Anderson:1987:GRO

[AOU87] T. W. Anderson, I. Olkin, and L. G. Underhill. Generation of random orthogonal matrices. SIAM J. Sci. Stat. Comput., 8(4):625–629, 1987.

Andrews:1976:SVD

[AP76] H. C. Andrews and C. L. Patterson. Singular value decompositions and digital image processing. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-24: 26–53, 1976. CODEN IETABA. ISSN 0096-3518.

Allen:1981:AIP

[AP81] R. C. Allen and S. A. Pruess. An analysis of an inverse problem in ordinary differential equations. SIAM J. Sci. Stat. Comput., 2:176–185, 1981.

Anda:1994:FPR

[AP94] Andrew A. Anda and Haesun Park. Fast plane rotations with dynamic scaling. SIAM J. Matrix Anal. Appl., 15(1): 162–174, 1994.

Appel:1985:EPM

[App85] A. W. Appel. An efficient program for many-body simulation. SIAM J. Sci. Stat. Comput., 6:85–103, 1985.

Alexander:1988:ARL

[APP88] S. T. Alexander, C.-T. Pan, and R. J. Plemmons. Analysis of a recursive least squares hyperbolic rotation algorithm for signal processing. Linear Algebra and Appl., 98:3–40, 1988.

Alvarado:1993:HPS

[APS93] Fernando L. Alvarado, Alex Pothen, and Robert S. Schreiber. Highly parallel sparse triangular solution. In J. Alan George, John R. Gilbert, and Joseph W. H. Liu, editors, Graph Theory and Sparse Matrix Computations, volume 56 of IMA Volumes in Mathematics and its Applications, pages 141–158. Springer-Verlag, New York, 1993.

Arioli:1985:RBC

[AR85] Mario Arioli and Francesco Romani.

Relations between condition numbers and the convergence of the Jacobi method for real positive definite matrices. Numer. Math., 46:31–42, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1992:SCC

[AR92] Mario Arioli and Francesco Romani.

Stability, convergence, and conditioning of stationary iterative methods of the form x^{(i+1)} = P x^{(i)} + q for the solution of linear systems. IMA J. Numer. Anal., 12:21–30, 1992.

Anton:1994:ELAb

[AR94] Howard Anton and Chris Rorres. Elementary Linear Algebra: Applications Version. Wiley, New York, NY, USA; London, UK; Sydney, Australia, seventh edition, 1994. ISBN 0-471-30570-7. xxvi + 779 + Appendix + Index pp.

Arkless:1986:HBE

[Ark86] Jan Arkless. How to Boil an Egg: Simple Cookery for One. Elliot Right Way Books, Kingswood, Surrey, 1986. ???? pp.

Artzrouni:1983:TPM

[Art83] Marc Artzrouni. A theorem on products of matrices. Linear Algebra and Appl., 49:153–159, 1983.

Artzrouni:1986:CIP

[Art86] Marc Artzrouni. On the convergence of infinite products of matrices. Linear Algebra and Appl., 74:11–21, 1986.

Arun:1992:UCT

[Aru92] K. S. Arun. A unitarily constrained total least squares problem in signal processing. SIAM J. Matrix Anal. Appl., 13 (3):729–745, 1992.

Abramowitz:1964:HMF

[AS64] Milton Abramowitz and Irene A. Stegun, editors. Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables, volume 55 of Applied mathematics series. U. S. Department of Commerce, Washington, DC, USA, 1964. xiv + 1046 pp. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available

                                        online at http://www.convertit.com/
                                        Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Alefeld:1982:SRM

[AS82] G. Alefeld and N. Schneider. On square roots of M-matrices. Linear Algebra and Appl., 42:119–132, 1982.

Andrews:1983:CNC

[AS83] G. R. Andrews and F. B. Schneider.

Concepts and notations for concurrent programming. ACM Computing Surveys, 15:3–43, 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Auzinger:1985:AAR

[AS85] W. Auzinger and H. J. Stetter. Accurate arithmetic results for decimal data on non-decimal computers. Computing, 35:141–151, 1985. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Aberth:1992:PCU

[AS92] Oliver Aberth and Mark. J. Schaefer. Precise computation using range arithmetic, via C++. ACM Transactions on Mathematical Software, 18(4): 481–491, 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Alvarado:1993:OPS

[AS93] Fernando L. Alvarado and Robert S.

Schreiber. Optimal parallel solution of sparse triangular systems. SIAM J. Sci. Comput., 14(2):446–460, 1993.

Ash:1986:ML

[Ash86a] A. Ash. A matrix laboratory. Notices Amer. Math. Soc., 33:482–485, 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Asherman:1986:STC

[Ash86b] Allan Asherman. The Star Trek Compendium. Pocket Books, New York, NY, USA, 1986. ISBN 0-671-62726-0. ???? pp.

Ash:1991:TE

[Ash91] Russell Ash. The Top 10 of Everything 1992. Queen Anne Press, Macdonald, London, 1991. ???? pp.

Ash:1993:RVB

[Ash93] Robert B. Ash. Real Variables with Basic Metric Space Topology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ???? pp. Year unsure. Have only ph’copy from Nil.

Aslett:1982:DDV

[Asl82] Don Aslett. Do I Dust or Vacuum First? Writer’s Digest Books, Cincinnati, Ohio, 1982. ???? pp.

Aslett:1989:TSI

[Asl89] Don Aslett. Is There a Speech Inside You? Writer’s Digest Books, Cincinnati, Ohio, 1989. ISBN 0-89879-361-0. ???? pp.

Aslaksen:1993:TTT

[Asl93] Helmer Aslaksen. Ten TEX tricks for the mathematician. TUGboat, 14(4):423– 424, 1993. ISSN 0896-3207.

Asner:1970:TNH

[Asn70] Bernard A. Asner, Jr. On the total nonnegativity of the Hurwitz matrix. SIAM J. Appl. Math., 18(2):407– 414, 1970.

Asplund:1959:IMW

[Asp59] Edgar Asplund. Inverses of matrices {a_{ij}} which satisfy a_{ij} = 0 for j > i + p. Math. Scand, 7:57–60, 1959. CODEN MTSCAN. ISSN 0025-5521 (print), 1903-1807 (electronic).

Aspray:1981:WLP

[Asp81] William Aspray. Wilkinson lecture on Pilot ACE. Annals of the History of Computing, 3(3):281–282, 1981.

Aspray:1990:JNO

[Asp90] William Aspray. John von Neumann and the Origins of Modern Computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-01121-2. xvii + 376 pp.

Ando:1973:APO

[ASS73] T. Ando, T. Sekiguchi, and T. Suzuki. Approximation by positive operators. Math. Zeitschr., 131:273–282, 1973.

Atanasoff:1940:CMS

[Ata40] John V. Atanasoff. Computing machine for the solution of large systems of linear algebraic equations. Unpublished manuscript, Iowa State College, Ames, IA, USA, August 1940. Reprinted in [Ran75b, pp. 315–335].

Atkinson:1989:INA

[Atk89] Kendall E. Atkinson. An Introduction to Numerical Analysis. Wiley, New York, NY, USA; London, UK; Sydney, Australia, second edition, 1989. ISBN 0-471-50023-2. ???? pp.

Atkinson:1993:ENA

[Atk93] Kendall Atkinson. Elementary Numerical Analysis. Wiley, New York, NY, USA; London, UK; Sydney, Australia, second edition, 1993. ISBN 0-471-60010-5. xiii + 425 pp.

Atta:1957:EPE

[Att57] Susie E. Atta. Effect of propagated error on inverse of Hilbert matrix. J. Assoc. Comput. Mach., 4:36–40, 1957.

Auchmuty:1992:PEE

[Auc92] G. Auchmuty. A posteriori error estimates for linear equations. Numer. Math., 61:1–6, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Augarde:1984:OGW

[Aug84] Tony Augarde. The Oxford Guide to Word Games. Oxford University Press, ????, 1984. ISBN 0-19-282005-2. ???? pp.

Autonne:1902:GLR

[Aut02] L. Autonne. Sur les groupes linéaires, Réels et orthogonaux. Bulletin de la Société Mathématique de France, 30: 121–134, 1902. CODEN BSMFAA. ISSN 0037-9484 (print), 2102-622X (electronic).

Abelson:1985:SIC

[AwJS85] Harold Abelson and Gerald Jay Sussman with Julie Sussman. Structure and Interpretation of Computer Programs. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-01077-1. xx + 542 pp.

Axelsson:1994:ISM

[Axe94] Owe Axelsson. Iterative Solution Methods. Cambridge University Press, ????, 1994. ISBN 0-521-44524-8. xiii + 654 pp.

Araki:1981:IHS

[AY81] Huzihiro Araki and Shigeru Yamagami. An inequality for Hilbert–Schmidt norm. Commun. Math. Phys., 81:89–96, 1981. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

Alvarado:1990:PSM

[AYB90] F. L. Alvarado, D. C. Yu, and R. Betancourt. Partitioned sparse A^{−1} methods. IEEE Transactions on Power Systems, 5(2):452–458, 1990. CODEN ITPSEG. ISSN 0885-8950 (print), 1558-0679 (electronic).

Bultman:1981:UFB

[B^{+}81] Mary Bultman et al. The Ultimate Fake Book. Hal Leonard, Milwaukee, WI, 1981. ISBN 0-9607350-0-3. ???? pp.

Babbage:1837:MPC

[Bab37] Charles Babbage. On the mathematical powers of the calculating engine. Unpublished manuscript. Buxton MS7, Museum of the History of Science, Oxford. Reprinted in [Ran75b, pp. 19–54], December 1837.

Babuska:1969:NSM

[Bab69] Ivo Babuška. Numerical stability in mathematical analysis. In Proc. IFIP Congress, Information Processing 68, pages 11–23. North-Holland, Amsterdam, The Netherlands, 1969.

Babuska:1972:NSP

[Bab72] Ivo Babuška. Numerical stability in problems of linear algebra. SIAM J. Numer. Anal., 9(1):53–77, 1972.

Baggett:1994:ETV

[Bag94] Jeffrey S. Baggett. Exponential type versus spectral abscissa: The Hille– Phillips example. Manuscript, 1994.

Bailey:1987:DEP

[Bai87] Richard W. Bailey, editor. Dictionaries of English: Prospects for the Record of Our Language. Cambridge University Press, ????, 1987.

Bailey:1988:NRT

[Bai88a] D. H. Bailey. Numerical results on the transcendence of constants involving π, e, and Euler’s constant. Math. Comp., 50:275–281, 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1988:CDD

[Bai88b] David H. Bailey. The computation of π to 29,360,000 decimal digits using Borweins’ quartically convergent algorithm. Math. Comp., 50(181):283–296, 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1988:EHS

[Bai88c] David H. Bailey. Extra high speed matrix multiplication on the Cray-2. SIAM J. Sci. Stat. Comput., 9(3):603– 607, 1988.

Bailey:1989:FEH

[Bai89] D. H. Bailey. FFTs in external or hierarchical memory. 1989.

Bailey:1991:TWF

[Bai91] David H. Bailey. Twelve ways to fool the masses when giving performance results on parallel computers. Supercomputer Rev., pages 54–55, August 1991.

Also in Supercomputer, Sept. 1991, pp. 4–7.

Bailey:1993:AMT

[Bai93a] David H. Bailey. Algorithm 719:

Multiprecision translation and execution of FORTRAN programs. ACM Transactions on Mathematical Software, 19(3): 288–319, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bailey:1993:ATF

[Bai93b] David H. Bailey. Automatic translation of Fortran programs to multiprecision. Technical Report RNR-91-025, NASA Ames Research Center, Moffett Field, CA, USA, May 1993. ???? pp.

Bailey:1993:EPC

[Bai93c] David H. Bailey. Experience with parallel computers at NASA Ames. Int Journal of High Speed Computing, 5(1): 51–62, 1993. CODEN IHSCEZ. ISSN 0129-0533.

Bailey:1993:PHP

[Bai93d] David H. Bailey. A portable high performance multiprecision package. Technical Report RNR-90-022, NASA Ames Research Center, Moffett Field, CA, USA, May 1993. ???? pp.

Bailey:1993:RMS

[Bai93e] David H. Bailey. RISC microprocessors and scientific computing. Technical Report RNR-93-004, NASA Ames Research Center, Moffett Field, CA, USA, March 1993. ???? pp.

Bai:1994:CTM

[Bai94a] Zhaojun Bai. A collection of test matrices for large scale nonsymmetric eigenvalue problems (version 1.0). Manuscript, July 1994.

Bailey:1994:FBM

[Bai94b] David H. Bailey. A Fortran-90 based multiprecision system. Technical Report RNR-94-013, NASA Ames Research Center, Moffett Field, CA, USA, June 1994. 12 pp.

Baker:1961:GTW

[Bak61] C. Baker. A Guide to Technical Writing. Pitman Publishing Ltd., London, UK, 1961. ???? pp.

Baksalary:1984:NDP

[Bak84] J. K. Baksalary. Nonnegative definite and positive definite solutions to the matrix equation AXA^{∗} = B. Linear and Multilinear Algebra, 16:133–139, 1984. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Baker:1985:SC

[Bak85a] Jenny Baker. The Student’s Cookbook. Faber and Faber, London, 1985. ???? pp.

Baker:1985:PS

[Bak85b] Sheridan Baker. The Practical Stylist. Harper & Row, New York, NY, USA, sixth edition, 1985. ISBN 0-06-040439-6. ???? pp.

Balzer:1980:ACM

[Bal80] L. A. Balzer. Accelerated convergence of the matrix sign function method of solving Lyapunov, Riccati and other matrix equations. Internat. J. Control, 32 (6):1057–1078, 1980.

Balliett:1986:AMP

[Bal86] Whitney Balliett. American Musicians: 56 Portraits in Jazz. Oxford University Press, ????, 1986. ???? pp.

Bank:1976:MAB

[Ban76] Randolph E. Bank. Marching algorithms and block Gaussian elimination. In Sparse Matrix Computations [BR76], pages 293–307.

Barnes:1965:ASN

[Bar65] J. G. P. Barnes. An algorithm for solving non-linear equations based on the secant method. The Computer Journal, 8:66–72, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Bard:1974:NPE

[Bar74] Yonathan Bard. Nonlinear Parameter Estimation. Academic Press, New York, 1974. ISBN 0-12-078250-2. x + 341 pp.

Barnett:1975:IMC

[Bar75a] S. Barnett. Introduction to Mathematical Control Theory. Oxford University Press, ????, 1975. ???? pp.

Barnett:1975:SAC

[Bar75b] S. Barnett. Some applications of the comrade matrix. Internat. J. Control, 21: 849–855, 1975.

Barrass:1978:SMW

[Bar78] Robert Barrass. Scientists Must Write: A Guide to Better Writing for Scientists, Engineers and Students. Chapman and Hall, Ltd., London, UK, 1978. ISBN 0-412-15430-7. ???? pp.

Barraud:1979:CNS

[Bar79a] A. Y. Barraud. Comments on “The numerical solution of A^{T} Q + QA = −C . IEEE Trans. Automat. Control., AC-24 (4):671–672, 1979.

Barraud:1979:IAF

[Bar79b] A. Y. Barraud. Investigations autour de la fonction signe d’une matrice application a l’équation de Riccati. R.A.I.R.O. Automatique/Systems Analysis and Control, 13(4):335–368, 1979. CODEN RFAADN. ISSN 0035-3035, 0399-0524.

Barlow:1981:DAR

[Bar81a] Jesse L. Barlow. On the distribution of accumulated roundoff error in floating point arithmetic. In Proc. 5th IEEE Symposium on Computer Arithmetic, Ann Arbor, Michigan, pages 100– 105. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981.

Barlow:1981:PEA

[Bar81b] Jesse L. Barlow. Probabilistic Error Analysis of Floating Point and CRD Arithmetics. PhD thesis, Northwestern University, Evanston, IL, USA, June 1981. ???? pp.

Barrass:1982:SMW

[Bar82] Robert Barrass. Students Must Write: A Guide to Better Writing in Course Work and Examinations. Methuen, London, 1982. ISBN 0-416-33620-5. ???? pp.

Barrett:1984:TMB

[Bar84] Wayne W. Barrett. Toeplitz matrices with banded inverses. Linear Algebra and Appl., 57:131–145, 1984.

Barlow:1986:NMS

[Bar86a] J. L. Barlow. A note on monitoring the stability of triangular decomposition of sparse matrices. SIAM J. Sci. Stat. Comput., 7(1):166–168, 1986.

Barlow:1986:SPS

[Bar86b] Jesse L. Barlow. On the smallest positive singular value of a singular M-matrix with applications to ergodic Markov chains. SIAM J. Alg. Disc. Meth., 7(3):414–424, 1986.

Baron:1986:CLG

[Bar86c] N. S. Baron. Computer Languages:

A Guide. Penguin, ????, 1986. ???? pp.

Barlow:1988:EAI

[Bar88a] Jesse L. Barlow. Error analysis and implementation aspects of deferred correction for equality constrained least squares problems. SIAM J. Numer. Anal., 25(6):1340–1358, 1988.

Barlow:1988:DDL

[Bar88b] Jesse L. Barlow. On the discrete distribution of leading significant digits in finite precision arithmetic. Technical Report CS-88-35, Department of Computer Science, Pennsylvania State University, University Park, PA, USA, September 1988. 15 pp.

Barrett:1989:FMA

[Bar89] Geoff Barrett. Formal methods applied to a floating-point number system. IEEE Trans. Software Engrg., 15(5):611– 621, 1989.

Barnett:1990:MMA

[Bar90a] S. Barnett. Matrices: Methods and Applications. Oxford University Press, ????, 1990. ???? pp.

Barrlund:1990:PBP

[Bar90b] Anders Barrlund. Perturbation bounds on the polar decomposition. BIT, 30:101–113, 1990. CODEN BITTEL.

ISSN 0006-3835 (print), 1572-9125 (electronic).

Barlow:1991:EAP

[Bar91a] Jesse L. Barlow. Error analysis of a pairwise summation algorithm to compute the sample variance. Numer. Math., 58:583–590, 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrlund:1991:PBD

[Bar91b] Anders Barrlund. Perturbation bounds for the LDL^{H} and LU decompositions. BIT, 31:358–363, 1991. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Barry:1991:T

[Bar91c] John A. Barry. Technobbable. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-02333-4. ???? pp.

Bartels:1991:TTM

[Bar91d] Sven G. Bartels. Two topics in matrix analysis: Structured sensitivity for Vandermonde-like systems and a subgradient method for matrix norm estimation. M.Sc. Thesis, Department of Mathematics and Computer Science, University of Dundee, Dundee, Scotland, September 1991. ???? pp.

Barrlund:1992:HIC

[Bar92a] Anders Barrlund. How integrals can be used to derive matrix perturbation bounds. Report UMINF 92.11, Institute of Information Processing, University of Umeå, Umeå, Sweden, September 1992. 8 pp.

Bartlett:1992:BFQ

[Bar92b] John Bartlett. Bartlett’s Familiar Quotations. Little, Brown and Company, Boston, sixteenth edition, 1992. ISBN 0-316-08277-5. lvi + 1405 pp. Justin Kaplan, general editor.

Barlow:1993:EAU

[Bar93a] Jesse L. Barlow. Error analysis of update methods for the symmetric eigenvalue problem. SIAM J. Matrix Anal. Appl., 14(2):598–618, 1993.

Barlow:1993:EBC

[Bar93b] Jesse L. Barlow. Error bounds for the computation of null vectors with applications to Markov chains. SIAM J. Matrix Anal. Appl., 14(3):798–812, 1993.

Barlow:1993:LNP

[Bar93c] Jesse L. Barlow. Lecture notes on parallel solution of the symmetric eigenvalue problem. In J. Gilbert and D. Kershaw, editors, Numerical Analysis 1992. Proceedings of SERC Summer School in Lancaster, page ?? Oxford University Press, 1993.

Barnsley:1993:FE

[Bar93d] Michael F. Barnsley. Fractals Everywhere. Academic Press, Boston, MA, USA, second edition, 1993. ISBN 0-12-079061-0. xiv + 531 pp.

Barrlund:1994:NSM

[Bar94] Anders Barrlund. A note on some matrix perturbation bounds. Report UMINF 94.03, Department of Information Processing, University of Umeå, Sweden, 1994. ???? pp.

Basilevsky:1983:AMA

[Bas83] A. Basilevsky. Applied Matrix Algebra in the Statistical Sciences. North-Holland, Amsterdam, The Netherlands, 1983. ???? pp.

Bates:1983:DXU

[Bat83] Douglas Bates. The derivative of |X′X| and its uses. Technometrics, 25 (4):373–376, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/1267859.

Bauer:1963:OSMb

[Bau63a] F. L. Bauer. Optimal scaling of matrices and the importance of the minimal condition. In Cicely M. Popplewell, editor, Proc. IFIP Congress 1962, Information Processing 62, pages 198–201. North-Holland, Amsterdam, The Netherlands, 1963.

Bauer:1963:OSMa

[Bau63b] F. L. Bauer. Optimally scaled matrices. Numer. Math., 5:73–87, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1966:GLG

[Bau66] F. L. Bauer. Genauigkeitsfragen bei der Lösung linearer Gleichungssysteme. Z. Angew. Math. Mech, 46(7):409–421, 1966.

Bauer:1969:ROS

[Bau69] F. L. Bauer. Remarks on optimally scaled matrices. Numer. Math., 13:1–3, 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1974:CGR

[Bau74] F. L. Bauer. Computational graphs and rounding errors. SIAM J. Numer. Anal., 11(1):87–96, 1974.

Bauer:1975:PN

[Bau75] F. L. Bauer. Positivity and norms.

Comm. ACM, 18:9–13, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baxter:1991:CPF

[Bax91] B. J. C. Baxter. Conditionally positive functions and p-norm distance matrices. Constructive Approximation, 7:427– 440, 1991. ISSN 0176-4276 (print), 1432-0940 (electronic).

Booth:1949:PPC

[BB49] Andrew D. Booth and Kathleen H. V.

Britten. Principles and progress in the construction of high-speed digital computers. Quart. J. Mech. and Applied Math., 2:182–197, 1949.

Beckenbach:1961:I

[BB61] E. F. Beckenbach and R. Bellman. Inequalities. Springer-Verlag, Berlin, 1961. ???? pp.

Bjoerck:1971:IAC

[BB71] Å. Björck and C. Bowie. An iterative algorithm for computing the best estimate of an orthogonal matrix. SIAM J. Numer. Anal., 8(2):358–364, 1971.

Borwein:1984:AGM

[BB84] J. M. Borwein and P. B. Borwein. The arithmetic-geometric mean and fast computation of elementary functions. SIAM Review, 26(3):351–366, 1984.

Barlow:1985:RED

[BB85a] Jesse L. Barlow and E. H. Bareiss. On roundoff error distributions in floating point and logarithmic arithmetic. Computing, 34:325–347, 1985. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Barlow:1985:PEA

[BB85b] Jesse L. Barlow and E. H. Bareiss. Probabilistic error analysis of Gaussian elimination in floating point and logarithmic arithmetic. Computing, 34:349–364, 1985. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Borwein:1987:PAA

[BB87] Jonathan M. Borwein and Peter B. Borwein. Pi and the AGM: A Study in Analytic Number Theory and Computational Complexity. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1987. ISBN 0-471-83138-7. ???? pp.

Barzilai:1988:TPS

[BB88a] Jonathan Barzilai and Jonathan M. Borwein. Two-point step size gradient methods. IMA J. Numer. Anal., 8:141– 148, 1988.

Brassard:1988:ATP

[BB88b] Gilles Brassard and Paul Bratley. Algorithmics: Theory and Practice. Prentice-Hall, Englewood Cliffs, NJ, USA, 1988. ISBN 0-13-023243-2. xvi + 361 pp.

Borwein:1992:SOC

[BB92] Jonathan Borwein and Peter Borwein. Some observations on computer aided analysis. Notices Amer. Math. Soc., 39 (8):825–829, 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bennani:1994:SCE

[BB94] Maria Bennani and Thierry Braconnier. Stopping criteria for eigensolvers. Journal of Numerical Linear Algebra with Applications, page ??, 1994. ISSN 0129-3281. Submitted.

Balakrishnan:1995:EUO

[BB95] V. Balakrishnan and S. Boyd. Existence and uniqueness of optimal matrix scalings. SIAM J. Matrix Anal. Appl., 16 (1):29–39, 1995.

Borwein:1989:RME

[BBB89] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to Pi or how to compute one billion digits of Pi. Amer. Math. Monthly, 96(3):201–219, 1989.

Bailey:1991:NPB

[BBB^{+}91] D. H. Bailey, E. Barszcz, J. T.

Barton, D. S. Browning, R. L. Carter, L. Dagum, R. A. Fatoohi, P. O. Frederickson, T. A. Lasinski, R. S. Schreiber, H. D. Simon, V. Venkatakrishnan, and S. K. Weeratunga. The NAS parallel benchmarks. Int. J. Supercomputer Appl., 5 (3):63–73, 1991. CODEN IJSAE9. ISSN 0890-2720.

Barrett:1994:TSL

[BBC^{+}94] Richard Barrett, Michael Berry, Tony F. Chan, James Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-328-5. xiii + 112 pp.

Barrett:1994:ABI

[BBD^{+}94] Richard Barrett, Michael Berry, Jack Dongarra, Victor Eijkhout, and Charles Romine. Algorithmic bombardment for the iterative solution of linear systems: A poly-iterative approach. Technical Report CS-94-239, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, August 1994. ???? pp.

Bojanczyk:1991:SBA

[BBdH91] A. W. Bojanczyk, R. P. Brent, and F. R. de Hoog. On stability of Bareiss algorithm. 1991.

Bojanczyk:1995:SBR

[BBdHS95] A. W. Bojanczyk, R. P. Brent, F. R. de Hoog, and D. R. Sweet. On the stability of the Bareiss and related Toeplitz factorization algorithms. SIAM J. Matrix Anal. Appl., 16(1):40–57, 1995.

Bliss:1992:API

[BBG92] B. Bliss, M.-C. Brunet, and E. Gallopoulos. Automatic program instrumentation with applications in performance and error analysis. In E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors, Expert Systems for Scientific Computing, pages 235–260. North-Holland, Amsterdam, The Netherlands, 1992.

Boyd:1989:BMC

[BBK89] S. Boyd, V. Balakrishan, and P. Kabamba. A bisection method for computing the H_{∞} norm of a transfer matrix and related problems. Math. Control, Signals and Sys., 2:207–219, 1989.

Benzinger:1987:CCD

[BBP87] Harold E. Benzinger, Scott A. Burns, and Julian I. Palmore. Chaotic complex dynamics and Newton’s method. Physics Letters A, 119(9):441–446, 1987. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Bojanczyk:1987:NDC

[BBVdH87] A. W. Bojanczyk, R. P. Brent, P. Van Dooren, and F. R. de Hoog. A note on downdating the Cholesky factorization. SIAM J. Sci. Stat. Comput., 8: 210–221, 1987.

Brunet:1986:CTS

[BC86] Marie-Christine Brunet and Françoise Chatelin. CESTAC, a tool for a stochastic round-off error analysis in scientific computing. In R. Vichnevetsky and J. Vignes, editors, Numerical Mathematics and Applications, pages 11–20. Elsevier North Holland, Inc., New York, NY, USA, 1986.

Blish:1990:GFA

[BC90] S. L. Blish and J. H. Curry. On the geometry of factorization algorithms. In Allgower and Georg [AG90], pages 47– 60. ISBN 0-8218-1131-2. LCCN QA372 .C6374 1990.

Braconnier:1995:HNE

[BCD95] Thierry Braconnier, Françoise Chatelin, and J. C. Dunyach. Highly nonnormal eigenproblems in the aeronautical industry. Japan J. Indust. Appl. Math., 12(1): 123–136, 1995. CODEN JAPJI7. ISSN 0916-7005.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Biles:1987:UES

[BCJR87] A. Biles, F. Cost, G. Johnson, and K. Reek. Using expert systems in typographic design. IEEE Trans. Prof. Commun., PC-30:102–111, 1987.

Beavers:1973:CME

[BD73] A. N. Beavers, Jr. and E. D. Denman. A computational method for eigenvalues and eigenvectors of a matrix with real eigenvalues. Numer. Math., 21:389–396, 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beavers:1974:ASM

[BD74a] A. N. Beavers, Jr. and Eugene D. Denman. Asymptotic solutions to the matrix Riccati equation. Mathematical Biosciences, 20:339–344, 1974. CODEN MABIAR. ISSN 0025-5564 (print), 1879-3134 (electronic).

Beavers:1974:NSM

[BD74b] A. N. Beavers, Jr. and Eugene D. Denman. A new solution method for quadratic matrix equations. Mathematical Biosciences, 20:135–143, 1974. CO-DEN MABIAR. ISSN 0025-5564 (print), 1879-3134 (electronic).

Boggs:1976:SAP

[BD76] P. T. Boggs and J. E. Dennis, Jr. A stability analysis for perturbed nonlinear iterative methods. Math. Comp., 30:199– 215, 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyle:1985:MFP

[BD85] J. P. Boyle and R. L. Dykstra. A method for finding projections onto the intersection of convex sets in Hilbert spaces. In Advances in Order Restricted Inference, volume 37 of Lecture Notes in Statistics, pages 28–47. Springer-Verlag, Berlin, 1985.

Boyce:1986:EDE

[BD86] William E. Boyce and Richard C.

DiPrima. Elementary Differential Equations and Boundary Value Problems. Wiley, New York, NY, USA; London, UK; Sydney, Australia, fourth edition, 1986. ISBN 0-471-83824-1. ???? pp.

Baker:1987:FGQ

[BD87] C. T. H. Baker and M. S. Derakhshan.

Fast generation of quadrature rules with some special properties. In P. Keast and G. Fairweather, editors, Numerical Integration: Recent Developments, Software and Applications, pages 53–60. D. Reidel Publishing Company, Dordrecht, Holland, 1987.

Bai:1989:BIH

[BD89a] Zhaojun Bai and James W. Demmel.

On a block implementation of Hessenberg multishift QR iteration. Int Journal of High Speed Computing, 1(1):97– 112, 1989. CODEN IHSCEZ. ISSN 0129-0533.

Bischof:1989:PDL

[BD89b] C. H. Bischof and J. J. Dongarra.

A project for developing a linear algebra library for high-performance computers. In Parallel Supercomputing: Methods, Algorithms and Applications [Car89], pages 45–56.

Barlow:1990:CAE

[BD90] Jesse L. Barlow and James W. Demmel. Computing accurate eigensytems of scaled diagonally dominant matrices. SIAM J. Numer. Anal., 27(3):762–791, 1990.

Bai:1991:DAC

[BD91] Z. Bai and J. W. Demmel. On a direct algorithm for computing invariant subspaces with specified eigenvalues. Technical Report CS-91-139, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, November 1991. ???? pp. LAPACK Working Note 38.

Bai:1992:DPN

[BD92] Zhaojun Bai and James W. Demmel. Design of a parallel nonsymmetric eigenroutine toolbox (part I). Research Report 92-09, Department of Mathematics, University of Kentucky, Lexington, KY, USA, December 1992. 30 pp.

Bai:1993:CGS

[BD93a] Zhaojun Bai and James W. Demmel. Computing the generalized singular value decomposition. SIAM J. Matrix Anal. Appl., 14(6):1464–1486, 1993.

Bai:1993:DPN

[BD93b] Zhaojun Bai and James W. Demmel. Design of a parallel nonsymmetric eigenroutine toolbox, Part I. In Richard F. Sincovec, David E. Keyes, Michael R. Leuze, Linda R. Petzold, and Daniel A. Reed, editors, Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, Volume I, pages 391–398. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993.

Bai:1993:SDB

[BD93c] Zhaojun Bai and James W. Demmel. On swapping diagonal blocks in real Schur form. Linear Algebra and Appl., 186:73–95, 1993.

Bai:1994:DPN

[BD94] Zhaojun Bai and James W. Demmel.

Design of a parallel nonsymmetric eigenroutine toolbox, part II. Computer Science Division Report UCB/CSD-94-???, University of California, Berkeley, CA, USA, ???? 1994. ???? pp.

Brewer:1988:TAA

[BDCS88] O. Brewer, J. J. Dongarra, J. J. Du Croz, and D. C. Sorensen. Tools to aid in the analysis of memory access patterns for Fortran programs. Parallel Computing, 9: 25–35, 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bischof:1988:GJH

[BDDC88] C. H. Bischof, J. W. Demmel, J. J. Dongarra, and J. J. Du Croz. A. Greenbaum, S. J. Hammarling, D. C. Sorensen, Provisional contents. Technical report, LAPACK Working Note #5, Report ANL-88-38, Mathematics and Computer Science Division, Argonne National Laboratory, Illinois, ????, 1988. ???? pp.

Bai:1994:IFP

[BDG94a] Zhaojun Bai, James W. Demmel, and Ming Gu. Inverse free parallel spectral divide and conquer algorithms for nonsymmetric eigenproblems. Computer Science Division Report UCB/CSD-94-793, University of California, Berkeley, CA, USA, February 1994. 34 pp.

Browne:1994:NSR

[BDG^{+}94b] Shirley V. Browne, Jack J. Dongarra, Stan C. Green, Keith Moore, Thomas H. Rowan, and Reed C. Wade. Netlib services and resources. Report ORNL/TM-12680, Oak Ridge National Laboratory, Oak Ridge, TN, USA, April 1994. 42 pp.

Barnsley:1988:SFI

[BDM^{+}88] M. F. Barnsley, R. L. Devaney, B. B. Mandelbrot, H.-O. Peitgen, D. Saupe, and R. F. Voss. The Science of Fractal Images. Springer-Verlag, Berlin, 1988. ???? pp.

Bai:1989:CNE

[BDM89] Z. Bai, J. W. Demmel, and A. McKenney. On the conditioning of the nonsymmetric eigenproblem: Theory and software. Technical report, LAPACK Working Note #13, Department of Computer Science, University of Tennessee, Knoxville, ????, 1989. ???? pp.

Bai:1993:CCN

[BDM93] Zhaojun Bai, James W. Demmel, and Alan McKenney. On computing condition numbers for the nonsymmetric eigenproblem. ACM Transactions on Mathematical Software, 19(2):202– 223, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bunch:1989:SSA

[BDV89] James R. Bunch, James W. Demmel, and Charles F. Van Loan. The strong stability of algorithms for solving symmetric linear systems. SIAM J. Matrix Anal. Appl., 10(4):494–499, 1989.

Bjoerck:1973:ACV

[BE73a] Åke Björck and Tommy Elfving. Algorithms for confluent Vandermonde systems. Numer. Math., 21:130–137, 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bukhberger:1973:MIT

[BE73b] B. Bukhberger and G. A. Emel’yanenko.

Methods of inverting tridiagonal matrices. U.S.S.R. Computational Math. and Math. Phys., 13:10–20, 1973. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Beauwens:1979:FIM

[Bea79] R. Beauwens. Factorization iterative methods, M-operators and Hoperators. Numer. Math., 31:335–357, 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beckmann:1971:H

[Bec71] Petr Beckmann. A History of π. St.

Martin’s Press, New York, 1971. ???? pp. Pi.

Beer:1992:WST

[Bee92] David F. Beer, editor. Writing and Speaking in the Technology Professions: A Practical Guide. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-87942-284-X. ???? pp.

Beebe:1993:BPS

[Bee93] Nelson H. F. Beebe. Bibliography prettyprinting and syntax checking. TUGboat, 14(4):395–419, 1993. ISSN 0896-3207.

Beiler:1966:RTN

[Bei66] Albert H. Beiler. Recreations in the Theory of Numbers: The Queen of Mathematics Entertains. Dover, New York, NY, USA, 1966. ISBN 0-486-21096-0. ???? pp.

Bell:1931:RCH

[Bel31] E. T. Bell. Review of “Contributions to the History of Determinants, 1900-1920”, by Sir Thomas Muir. Amer. Math. Monthly, 38:161–164, 1931. Reprinted in [Ewi94].

Bellman:1968:SIS

[Bel68] R. Bellman. Some inequalities for the square root of a positive definite matrix. Linear Algebra and Appl., 1:321– 324, 1968.

Bellman:1970:IMA

[Bel70] Richard Bellman. Introduction to Matrix Analysis. McGraw-Hill, New York, second edition, 1970. ISBN 0-89871-346-3. xxiii + 403 pp. First edition (1960) reprinted by Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995.

Benoit:1924:NMR

[Ben24] Commandant Benoit. Note sur une méthode de résolution des équations normales provenant de l’application de la méthode des moindres carrés à un système d’équations linéaires en nombre inférieur à celui des inconnues. Application de la méthode à la résolution d’un système défini d’équations linéaires (Procédé de Commandant Cholesky). Bulletin Géodésique (Toulouse), 7(1):67– 77, 1924. CODEN BGDQAG. ISSN 0007-4632. Cited in [Ano22, Cottle’s translation].

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society, 78(4): 551–572, March 1938. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:

//links.jstor.org/sici?sici=0003-
049X%2819380331%2978%3A4%3C551%3ATLOAN%

3E2.0.CO%3B2-G. See comments about Benford’s biased rounding practices [?].

Bentley:1982:WEP

[Ben82] Jon Loius Bentley. Writing Efficient Programs. Prentice-Hall, Englewood Cliffs, NJ, USA, 1982. ???? pp.

Bentley:1985:PPC

[Ben85] J. L. Bentley. Programming pearls:

Confessions of a coder. Comm. ACM, 28: 671–679, 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1986:PP

[Ben86] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10331-1 (paperback). viii + 195 pp. LCCN QA76.6.B453 1986. Reprinted with corrections.

Bentley:1987:PPA

[Ben87a] J. L. Bentley. Programming pearls:

Abstract data types. Comm. ACM, 30: 284–290, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1987:PPP

[Ben87b] J. L. Bentley. Programming pearls:

Profilers. Comm. ACM, 30:587–592, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1988:MPP

[Ben88] Jon Louis Bentley. More Programming Pearls: Confessions of a Coder. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-11889-0. viii + 207 pp. LCCN QA76.6 .B452 1988. US$18.75.

Berkeley:1949:GBM

[Ber49] Edmund Callis Berkeley. Giant Brains or Machines That Think. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1949. xvi + 270 pp.

Bernstein:1965:CWM

[Ber65] Theodore M. Bernstein. The Careful Writer: A Modern Guide to English Usage. Atheneum, New York, 1965. ISBN 0-689-70555-7. xviii + 487 pp.

Bernstein:1971:AEA

[Ber71a] A. R. Bernstein. Almost eigenvectors for almost commuting matrices. SIAM J. Appl. Math., 21:232–235, 1971.

Bernstein:1971:MTH

[Ber71b] Theodore M. Bernstein. Miss Thistlebottom’s Hobgoblins: The Careful Writer’s Guide to the Taboos, Bugbears and Outmoded Rules of English Usage. Farrar, Straus and Giroux, New York, 1971. ISBN 0-671-50404-5. ???? pp. Reprinted by Simon and Schuster, New York, 1984.

Bergerson:1973:PA

[Ber73] H. W. Bergerson. Palindromes and Anagrams. Dover, New York, NY, USA, 1973. ???? pp.

Berry:1975:YVH

[Ber75] Cicely Berry. Your Voice and How to Use it Successfully. Harrap, London, 1975. ISBN 0-245-52886-5. ???? pp.

Bernstein:1977:DME

[Ber77a] Theodore M. Bernstein. Dos, Dont’s and Maybes of English Usage. Barnes and Noble, New York, NY, USA, 1977. ISBN 0-88029-944-4. 250 pp.

Bernstein:1977:P

[Ber77b] Theodore M. Bernstein. Punctuation. IEEE Trans. Prof. Commun., PC-20(1):38–44, 1977. Reprinted from [Ber65].

Berliner:1980:BCP

[Ber80] H. J. Berliner. Backgammon computer program beats world champion. Artificial Intelligence, 14:205–220, 1980. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Bernard:1990:RRL

[Ber90] André Bernard, editor. Rotten Rejections: A Literary Companion. Penguin, New York, 1990. ISBN 0-14-014786-1. ???? pp.

Berenbaum:1991:DSM

[Ber91] M. C. Berenbaum. Direct search methods in the optimisation of cancer chemotherapy. British J. Cancer, 61:101– 109, 1991. CODEN BJCAAI. ISSN 0007-0920 (print), 1532-1827 (electronic).

Berry:1992:LSS

[Ber92] Michael W. Berry. Large-scale sparse singular value computations. Int. J. Supercomputer Appl., 6(1):13–49, 1992. CO-DEN IJSAE9. ISSN 0890-2720.

Bevington:1969:DRE

[Bev69] Philip R. Bevington. Data Reduction and Error Analysis for the Physical Sciences. McGraw-Hill, New York, 1969. xv + 336 pp.

Bauer:1960:NET

[BF60] F. L. Bauer and C. T. Fike. Norms and exclusion theorems. Numer. Math., 2:137–141, 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1981:IBM

[BF81] Wayne W. Barrett and Philip J. Feinsilver. Inverses of banded matrices. Linear Algebra and Appl., 41:111–130, 1981.

Brandreth:1987:SC

[BF87] G. Brandreth and D. Francis. The Scrabble Companion. Stanley Paul, London, 1987. ???? pp.

Bailey:1988:SNA

[BF88] D. H. Bailey and H. R. P. Ferguson. A Strassen–Newton algorithm for highspeed parallelizable matrix inversion. In Proceedings of Supercomputing ’88, pages 419–424. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988.

Burden:1989:NA

[BF89] Richard L. Burden and J. Douglas Faires. Numerical Analysis. PWS-KENT, Boston, fourth edition, 1989. ISBN 0-534-91585-X. ???? pp.

Brown:1990:SAM

[BF90] P. J. Brown and W. A. Fuller. Statistical Analysis of Measurement Error Models and Applications, volume 112 of Contemporary Mathematics. American Mathematical Society, Providence, RI, USA, 1990. ???? pp.

Bull:1992:NPA

[BF92] J. M. Bull and T. L. Freeman. Numerical performance of an asynchronous Jacobi iteration. Numerical Analysis Report No. 210, University of Manchester, Manchester, England, February 1992. ???? pp.

Bini:1993:MAR

[BF93a] Dario Bini and Paola Favati. On a matrix algebra related to the discrete Hartley transform. SIAM J. Matrix Anal. Appl., 14(2):500–507, 1993.

Burden:1993:NA

[BF93b] Richard L. Burden and J. Douglas Faires. Numerical Analysis. PWS, Boston, fifth edition, 1993. ISBN 0-534-93219-3. xiv + 768 pp.

Baumann:1964:IA

[BFBS64] R. Baumann, M. Feliciano, F. L.

Bauer, and K. Samelson. Introduction to Algol. Prentice-Hall, Englewood Cliffs, NJ, USA, 1964. ???? pp.

Barrett:1988:SRM

[BFP88] Wayne W. Barrett, Rodney W. Forcade, and Andrew D. Pollington. On the spectral radius of a (0,1) matrix related to Mertens’ function. Linear Algebra and Appl., 107:151–159, 1988.

Businger:1965:LLS

[BG65] P. Businger and G. H. Golub. Linear least squares solutions by Householder transformations. Numer. Math., 7:269–276, 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 111–118], Contribution I/8.

Bjoerck:1967:IRLb

[BG67] Åke Björck and Gene H. Golub. Iterative refinement of linear least squares solutions by Householder transformation.

BIT, 7:322–337, 1967. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjoerck:1973:NMC

[BG73] Åke Björck and Gene H. Golub. Numerical methods for computing angles between linear subspaces. Math. Comp., 27 (123):579–594, 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barwell:1976:CAS

[BG76] Victor Barwell and Alan George. A comparison of algorithms for solving symmetric indefinite systems of linear equations. ACM Transactions on Mathematical Software, 2(3):242–251, 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bjoerck:1977:EMA

[BG77] Åke Björck and Gene H. Golub. Eigenproblems for matrices associated with periodic boundary conditions. SIAM Review, 19(1):5–16, 1977.

Birkhoff:1979:IDT

[BG79] Garrett Birkhoff and Surender Gulati. Isotropic distributions of test matrices. J. Appl. Math. Phys. (ZAMP), 30:148–158, 1979. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Boley:1987:SMI

[BG87] D. Boley and G. H. Golub. A survey of matrix inverse eigenvalue problems. Inverse Problems, 3:595–622, 1987. CO-DEN INPEEY. ISSN 0266-5611 (print), 1361-6420 (electronic).

Boller:1989:TNS

[BG89] P. F. Boller, Jr. and J. George. They Never Said It: A Book of Fake Quotes, Misquotes, and Misleading Attributions. Oxford University Press, ????, 1989. ???? pp.

Brankin:1990:CAB

[BG90] R. W. Brankin and I. Gladwell. Codes for almost block diagonal systems. International J. Computer and Mathematics with Applications, 19:1–6, 1990.

Bondeli:1994:CRS

[BG94] S. Bondeli and W. Gander. Cyclic reduction for special tridiagonal systems. SIAM J. Matrix Anal. Appl., 15(1):321– 330, 1994.

Brainerd:1990:PGF

[BGA90] Walter S. Brainerd, Charles H.

Goldberg, and Jeanne C. Adams. Programmer’s Guide to Fortran 90. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-000248-7. vii + 410 pp. LCCN QA76.73.F25 B735 1990. US$37.95. See [?].

Blackmore:1987:FRT

[bGBB87] Arranged by George Blackmore and Editor Cecil Bolton. Favourite Radio Themes. International Music Publications, ????, 1987. ISBN 0-86359-420-4. ???? pp.

Bunse-Gerstner:1993:NMS

[BGBM93] Angelika Bunse-Gerstner, Ralph Byers, and Volker Mehrmann. Numerical methods for simultaneous diagonalization. SIAM J. Matrix Anal. Appl., 14 (4):927–949, 1993.

Bunse-Gerstner:1990:NCA

[BGBMN90] A. Bunse-Gerstner, Ralph Byers, V. Mehrmann, and N. K. Nichols. Numerical computation of an analytic singular value decomposition of a matrix valued function. Technical report, Materialien LIX, Universität Bielefeld, FRG, ????, 1990. ???? pp. to appear in Numer. Math.

Brankin:1989:ARK

[BGD^{+}89] R. W. Brankin, I. Gladwell, J. R. Dormand, P. J. Prince, and W. L. Seward. Algorithm 670: A Runge–Kutta Nyström code. ACM Transactions on Mathematical Software, 15:31–40, 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Boyd:1994:LMI

[BGFB94] Stephen Boyd, Laurent El Ghaoui, Eric Feron, and Venkataramanan Balakrishnan. Linear Matrix Inequalities in System and Control Theory. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-334-X. ix + 193 pp.

Bunse-Gerstner:1988:SVD

[BGG88] Angelika Bunse-Gerstner and William B. Gragg. Singular value decompositions of complex symmetric matrices. J. Comp. Appl. Math., 21:41–54, 1988.

Baumert:1962:DHM

[BGH62] L. Baumert, S. W. Golomb, and M. Hall, Jr. Discovery of an Hadamard matrix of order 92. Bull. Amer. Math. Soc., 68:237–238, 1962.

Bunse-Gerstner:1992:SSP

[BGH92] A. Bunse-Gerstner and C. He. A Sturm sequence of polynomials for unitary Hessenberg matrices. 1992.

Boley:1994:FPF

[BGM^{+}94] Daniel Boley, Gene H. Golub, Samy Makar, Nirmal Saxena, and Edward J. McCluskey. Floating point fault tolerance with backward error assertions. IEEE Trans. Comput., 44(2):302–311, 1994.

Buzbee:1970:DMS

[BGN70] B. L. Buzbee, G. H. Golub, and C. W. Nielson. On direct methods for solving Poisson’s equations. SIAM J. Numer. Anal., 7(4):627–656, 1970.

Brankin:1987:CRS

[BGS87] R. W. Brankin, I. Gladwell, and L. F.

Shampine. Codes for reliable solution of special event location problems for ODEs. Numerical Analysis Report No. 139, University of Manchester, Manchester, England, 1987. ???? pp.

Bunch:1974:TFI

[BH74] James R. Bunch and John E.

Hopcroft. Triangular factorization and inversion by fast matrix multiplication. Math. Comp., 28(125):231–236, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bjoerck:1983:SMS

[BH83] Åke Björck and Sven Hammarling. A Schur method for the square root of a matrix. Linear Algebra and Appl., 52/53: 127–140, 1983.

Broom:1987:IUT

[BH87] M. A. Broom and T. R. Hopkins. The implementation and use of Toolpack/1 on a graphics workstation. Software— Practice and Experience, 17(9):561–573, 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barlow:1988:DSW

[BH88] Jesse L. Barlow and Susan L. Handy. The direct solution of weighted and equality constrained least-squares problems. SIAM J. Sci. Stat. Comput., 9(4):704– 716, 1988.

Bischof:1991:SPR

[BH91] C. H. Bischof and P. C. Hansen. Structure-preserving and rank-revealing QR-factorizations. SIAM J. Sci. Stat. Comput., 12:1332–1350, 1991.

Bartels:1992:SSV

[BH92] Sven G. Bartels and Desmond J. Higham. The structured sensitivity of Vandermonde-like systems. Numer. Math., 62:17–33, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braconnier:1995:CFV

[BH95a] Thierry Braconnier and Nicholas J. Higham. Computing the field of values and pseudospectra using the Lanczos method with continuation. Numerical Analysis Report No. 279, Manchester Centre for Computational Mathematics, Manchester, England, November 1995. 20 pp. Submitted to BIT.

Briggs:1995:DOM

[BH95b] William L. Briggs and Van Emden Henson. The DFT: An Owner’s Manual for the Discrete Fourier Transform. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-342-0. xv + 434 pp.

Bhatia:1993:FSO

[Bha93] Rajendra Bhatia. First and second order perturbation bounds for the operator absolute value. Manuscript, Indian Statistical Institute, 1993.

Bhatia:1994:MFT

[Bha94] Rajendra Bhatia. Matrix factorisations and their perturbations. Linear Algebra and Appl., 197/198:245–276, 1994.

Balle:1993:STM

[BHH93] Susanne M. Balle, Per Christian Hansen, and Nicholas J. Higham. A Strassen-type matrix inversion algorithm for the Connection Machine. Technical Report CNC/1993/028, Centre for Novel Computing, University of Manchester, Manchester, England, October 1993. 29 pp.

Bhatia:1991:NAB

[BHK91] R. Bhatia, R. A. Horn, and F. Kittaneh. Normal approximants to binormal operators. Linear Algebra and Appl., 147: 169–179, 1991.

Byers:1994:MSF

[BHM94] Ralph Byers, Chunyang He, and Volker Mehrmann. The matrix sign function method and the computation of invariant subspaces. Manuscript, February 1994.

Bai:1994:NBG

[BHR94] Z. Bai, D. Hu, and L. Reichel. A Newton basis GMRES implementation.

IMA J. Numer. Anal., 14(4):563–581, 1994.

Brent:1980:AIB

[BHY80] Richard P. Brent, Judith A. Hooper, and J. Michael Yohe. An AUGMENT interface for Brent’s multiple precision arithmetic package. ACM Transactions on Mathematical Software, 6(2):146–149, 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bar-Itzhack:1975:IOO

[BI75] Itzhack Y. Bar-Itzhack. Iterative optimal orthogonalization of the strapdown matrix. IEEE Trans. Aerospace Electronic Systems, AES-11(1):30–37, 1975.

Bar-Itzhack:1977:UCT

[BI77] Itzhack Y. Bar-Itzhack. A unidimensional convergence test for matrix iterative processes applied to strapdown navigation. International J. Numer. Methods Engineering, 11:115–130, 1977.

Barlow:1987:SGR

[BI87] Jesse L. Barlow and Ilse C. F. Ipsen. Scaled Givens rotations for the solution of linear least squares problems on systolic arrays. SIAM J. Sci. Stat. Comput., 8(5): 716–733, 1987.

Bierce:1911:DD

[Bie11] Ambrose Bierce. The Devil’s Dictionary. Neale Publishing Company, ????, 1911. ISBN 0-486-20487-1. ???? pp. Reprinted by Dover, New York, 1958.

Bierce:1937:WIR

[Bie37] Ambrose Bierce. Write it Right: A Little Blacklist of Literary Faults. The Union Library Association, New York, 1937. ???? pp. Reprinted in [Ber71b, p. 209–253].

Bar-Itzhack:1969:OTD

[BIF69] I. Y. Bar-Itzhack and K. A. Fegley. Orthogonalization techniques of a direction cosine matrix. IEEE Trans. Aerospace Electronic Systems, AES-5(5): 798–804, 1969.

Bar-Itzhack:1976:CIO

[BIM76] Itzhack Y. Bar-Itzhack and Jeffrey Meyer. On the convergence of iterative orthogonalization processes. IEEE Trans. Aerospace Electronic Systems, AES-12 (2):146–151, 1976.

Bar-Itzhack:1990:MPS

[BIM90] I. Y. Bar-Itzhack and F. L. Markley.

Minimal parameter solution of the orthogonal matrix differential equation. IEEE Trans. Automat. Control., AC-35: 314–317, 1990.

Bar-Itzhack:1976:SMO

[BIMF76] IItzhack Y. Bar-Itzhack, J. Meyer, and P. A. Fuhrmann. Strapdown matrix orthogonalization: The dual iterative algorithm. IEEE Trans. Aerospace Electronic Systems, AES-12(1):32–37, 1976.

Birkhoff:1975:THN

[Bir75] Garrett Birkhoff. Two Hadamard numbers for matrices. Comm. ACM, 18(1):25–29, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bischof:1988:AB

[Bis88a] C. H. Bischof. Adaptive blocking. Technical report, Preprint MCS-P39-1288, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1988. ???? pp.

Bischof:1988:QFA

[Bis88b] C. H. Bischof. QR Factorization Algorithms for Course-grained Distributed Systems. Ph.d. thesis, Cornell University, Ithaca, NY, USA, 1988. ???? pp.

Bischof:1989:BQF

[Bis89a] C. H. Bischof. A block QR factorization algorithm using restricted pivoting. Technical report, Preprint MCS-P94-0789, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1989. ???? pp.

Bischof:1989:CSV

[Bis89b] C. H. Bischof. Computing the singular value decomposition on a distributed system of vector processors. Parallel Computing, 11:171–186, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bischof:1990:FLA

[Bis90a] C. H. Bischof. Fundamental linear algebra computations on high-performance computers. Technical report, Preprint MCS-P150-0490, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1990. ???? pp.

Bischof:1990:ICEa

[Bis90b] Christian H. Bischof. Incremental condition estimation. SIAM J. Matrix Anal. Appl., 11(2):312–322, 1990.

Bischof:1991:PQF

[Bis91] C. H. Bischof. A parallel QR factorization algorithm with controlled local pivoting. SIAM J. Sci. Stat. Comput., 21: 36–57, 1991.

Bronlund:1974:FPM

[BJ74] O. E. Brønlund and T. L. Johnsen.

QR-factorization of partitioned matrices. Computer Methods in Applied Mechanics and Engineering, 3:153–172, 1974.

Bartels:1989:GDL

[BJ89] Richard Bartels and Barry Joe. On generating discrete linear l_{∞} test problems. SIAM J. Sci. Stat. Comput., 10(3): 550–561, 1989.

Barrett:1992:SPM

[BJ92] Wayne W. Barrett and Tyler J. Jarvis.

Spectral properties of a matrix of Redheffer. Linear Algebra and Appl., 162:673– 683, 1992.

Bjoerck:1967:IRLa

[Bjö67a] Åke Björck. Iterative refinement of linear least squares solutions I. BIT, 7: 257–278, 1967. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjoerck:1967:SLL

[Bjö67b] Åke Björck. Solving linear least squares problems by Gram–Schmidt orthogonalization. BIT, 7:1–21, 1967. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjoerck:1968:IRL

[Bjö68] Åke Björck. Iterative refinement of linear least squares solutions II. BIT, 8:8– 30, 1968. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjoerck:1978:CIR

[Bjö78] Åke Björck. Comment on the iterative refinement of least-squares solutions. J. Amer. Statist. Assoc., 73(361):161–166,

Bjoerck:1987:SAM

[Bjö87] Åke Björck. Stability analysis of the method of seminormal equations for linear least squares problems. Linear Algebra and Appl., 88/89:31–48, 1987.

Bjoerck:1988:CBE

[Bjö88] Å. Björck. Componentwise backward errors and condition estimates for linear least squares problems. 1988.

Bjoerck:1990:IRR

[Bjö90a] Åke Björck. Iterative refinement and reliable computing. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 249–266. Oxford University Press, 1990.

Bjoerck:1990:LSM

[Bjö90b] Åke Björck. Least squares methods. In P. G. Ciarlet and J. L. Lions, editors, Handbook of Numerical Analysis, volume I: Finite Difference Methods—Solution of Equations in R^{n}, page ?? Elsevier/North-Holland, Amsterdam, The Netherlands, 1990.

Bjoerck:1991:CWP

[Bjö91a] Åke Björck. Component-wise perturbation analysis and error bounds for linear least squares solutions. BIT, 31:238– 244, 1991. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjoerck:1991:EAL

[Bjö91b] Åke Björck. Error analysis of least squares algorithms. In G. H. Golub and P. M. Van Dooren, editors, Numerical Linear Algebra, Digital Signal Processing and Parallel Algorithms, volume F70 of NATO ASI Series, pages 41–73. Springer-Verlag, Berlin, 1991.

Bjoerck:1991:LSM

[Bjö91c] Åke Björck. Least squares methods.

Technical report, Department of Mathematics, Linköping University, Sweden, May 1991. ???? pp.

Bjoerck:1992:PSA

[Bjö92] Åke Björck. Pivoting and stability in the augmented system method. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 1–16. Longman Scientific and Technical, Harlow, Essex, UK, 1992.

Bjoerck:1994:NGS

[Bjö94] Åke Björck. Numerics of Gram– Schmidt orthogonalization. Linear Algebra and Appl., 197/198:297–316, 1994.

Bjoerck:1996:NML

[Bjö96] Åke Björck. Numerical Methods for Least Squares Problems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-360-9. ???? pp. In Press.

Butler:1985:NSQ

[BJW85] G. J. Butler, C. R. Johnson, and H. Wolkowicz. Nonegative solutions of a quadratic matrix equation arising from comparison theorems in ordinary differential equations. SIAM J. Alg. Disc. Meth., 6:47–53, 1985.

Browne:1969:ORF

[BK69] Michael Browne and Walter Kristof.

On the oblique rotation of a factor matrix to a specified pattern. Psychometrika, 34(2):237–248, 1969. CODEN PSMIA3.

ISSN 0033-3123 (print), 1860-0980 (electronic).

Bunch:1977:SSM

[BK77] James R. Bunch and Linda Kaufman. Some stable methods for calculating inertia and solving symmetric linear systems. Math. Comp., 31(137):163–179, 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baksalary:1979:ME

[BK79] J. K. Baksalary and R. Kala. The matrix equation AX − Y B = C. Linear Algebra and Appl., 25:41–43, 1979.

Baksalary:1980:ME

[BK80] J. K. Baksalary and R. Kala. The matrix equation AXB + CY D = E. Linear Algebra and Appl., 30:141–147, 1980.

Bentley:1986:GLTa

[BK86a] Jon L. Bentley and Brian W. Kernighan. GRAP—A language for typesetting displays. Technical report, Computer Science and Statistics, Proceedings of the 18th Symposium on the Interface, Amer. Statist. Assoc., Washington DC, ????, 1986. ???? pp.

Bentley:1986:GLTb

[BK86b] Jon L. Bentley and Brian W. Kernighan. GRAP—A language for typesetting graphs. Comm. ACM, 29:782–792, 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1988:TPI

[BK88] Jon L. Bentley and Brian W. Kernighan. Tools for printing indexes. Electronic Publishing—Origination, Dissemination, and Design, 1(1):3–17, 1988. CODEN EPODEU. ISSN 0894-3982.

Also Computer Science Technical Report No. 128, AT&T Bell Laboratories, Murray Hill, NJ, October 1986.

Bentley:1986:PPL

[BKM86] Jon Bentley, Donald E. Knuth, and M. Douglas McIlroy. Programming pearls: a literate program: a WEB program for common words. Comm. ACM, 29(6): 471–483, June 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/ft_
gateway.cfm?id=315653. See another

solution to the problem of printing the k most common words of a text file in [?]. Reprinted in [?, 151–177].

Boros:1994:FAS

[BKO94] T. Boros, T. Kailath, and V. Olshevsky. Fast algorithms for solving Vandermonde and Chebyshev–Vandermonde systems. Manuscript, 1994.

Bunch:1976:DSM

[BKP76] James R. Bunch, Linda Kaufman, and Beresford N. Parlett. Decomposition of a symmetric matrix. Numer. Math., 27:95–109, 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bailey:1993:MPM

[BKP93] David H. Bailey, Robert Krasny, and Richard Pelz. Multiple precision, multiple processor vortex sheet roll-up computation. In Richard F. Sincovec, David E. Keyes, Michael R. Leuze, Linda R. Petzold, and Daniel A. Reed, editors, Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, Volume I, pages 52–56. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993.

Belsley:1980:RDI

[BKW80] D. A. Belsley, E. Kuh, and R. E. Welsch. Regression Diagnostics: Identifying Influential Data and Sources of Collinearity. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1980. ???? pp.

Bjoerck:1990:WPC

[BKW90] Lennart Björck, Michael Knight, and Eleanor Wikborg. The Writing Process: Composition Writing for University Students. Studentlitteratur (Lund, Sweden) and Chartwell Bratt (Bromley, Kent, England), ????, second edition, 1990. ISBN 91-44-28222-2 and 0-86238-300-5. ???? pp.

Birkhoff:1977:SMA

[BL77] Garrett Birkhoff and Saunders Mac Lane. A Survey of Modern Algebra. MacMillan Publishing Company, New York, NY, USA, fourth edition, 1977. ISBN 0-02-310070-2. xi + 500 pp.

Bini:1980:SFA

[BL80] Dario Bini and Grazia Lotti. Stability of fast algorithms for matrix multiplication. Numer. Math., 36:63–72, 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Birkhoff:1984:NSE

[BL84] Garrett Birkhoff and Robert E. Lynch. Numerical Solution of Elliptic Problems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1984. ISBN 0-89871-197-5. xi + 319 pp.

Byrnes:1986:CCM

[BL86] C. I. Byrnes and A. Lindquist. Computational and Combinatorial Methods in Systems Theory. Elsevier North Holland, Inc., New York, NY, USA, 1986. ???? pp.

Belitskii:1988:MNT

[BL88] G. R. Belitskii and Yu. I. Lyubich. Matrix Norms and Their Applications, volume 36 of Operator Theory: Advances and Applications. Birkhäuser Verlag, Basel, Switzerland, 1988. ???? pp.

Bischof:1990:ABS

[BL90] C. H. Bischof and P. G. Lacroute. An adaptive blocking strategy for matrix factorizations. Technical report, Preprint MCS-P151-0490, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1990. ???? pp.

Bojanczyk:1991:CEA

[BL91] Adam W. Bojanczyk and Adam Lutoborski. Computation of the Euler angles of a symmetric 3 × 3 matrix. SIAM J. Matrix Anal. Appl., 12(1):41–48, 1991.

Barlow:1995:MCO

[BL95] J. B. Barlow and S.-M. Lu. Multifrontal computation with the orthogonal factors of sparse matrices. SIAM J. Matrix Anal. Appl., page ??, 1995. To appear.

Blanch:1964:NEC

[Bla64] G. Blanch. Numerical evaluation of continued fractions. SIAM Review, 6(4): 383–421, 1964.

Bligh:1972:WUL

[Bli72] Donald A. Bligh. What’s the Use of Lectures? Penguin, ????, 1972. ???? pp.

Bloch:1948:MC

[Blo48] Richard M. Bloch. Mark I calculator. In Anonymous [Ano48], pages 23–30. ISBN 0-262-08152-0.

Bischof:1990:ICEb

[BLP90] Christian H. Bischof, John G. Lewis, and Daniel J. Pierce. Incremental condition estimation for sparse matrices. SIAM J. Matrix Anal. Appl., 11(4):644–659, 1990.

Bevilacqua:1990:SCI

[BLR90] R. Bevilacqua, C. Lotti, and F. Romani. Storage compression of inverses of band matrices. Computers and Mathematics with Applications, 20:1–11, 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Bailey:1991:USA

[BLS91] David H. Bailey, King Lee, and Horst D. Simon. Using Strassen’s algorithm to accelerate the solution of linear systems. The Journal of Supercomputing, 4:357–371, 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Blue:1978:PFP

[Blu78] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. ACM Transactions on Mathematical Software, 4(1):15–23, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brent:1985:CSV

[BLV85] Richard P. Brent, Franklin T. Luk, and Charles F. Van Loan. Computation of the singular value decomposition using mesh-connected processors. J. VLSI and Computer Systems, 1:242–270, 1985.

Borodin:1975:CCA

[BM75] A. Borodin and I. Munro. The Computational Complexity of Algebraic and Numeric Problems. American Elsevier, New York, 1975. ???? pp.

Balfour:1979:PSF

[BM79] A. Balfour and D. H. Marwick. Programming in Standard Fortran 77. Heinemann, London, 1979. ISBN 0-435-77486-7. ???? pp.

Baker:1982:TIE

[BM82] C. T. H. Baker and G. F. Miller, editors. Treatment of Integral Equations by Numerical Methods. Academic Press, London, 1982. ???? pp.

Berry:1984:BC

[BM84] R. E. Berry and B. A. E. Meekings. A Book on C. MacMillan Publishing Company, New York, NY, USA, 1984. ???? pp.

Bradley:1986:MS

[BM86] I. Bradley and R. L. Meek. Matrices and Society. Penguin, ????, 1986. ???? pp.

Brankin:1992:FFC

[BM92] L. A. Brankin and A. M. Mumford.

File formats for computer graphics: Unraveling the confusion. Technical Report TR1/92, NAG Ltd., Oxford, June 1992. ???? pp.

Bhatia:1994:VUP

[BM94] Rajendra Bhatia and Kalyan Mukherjea. Variation of the unitary part of a matrix. SIAM J. Matrix Anal. Appl., 15(3): 1007–1014, 1994.

Barlow:1992:CMS

[BMO92] J. B. Barlow, M. M. Monahemi, and D. P. O’Leary. Constrained matrix Sylvester equations. SIAM J. Matrix Anal. Appl., 13(1):1–9, 1992.

Bowdler:1966:SRC

[BMPW66] H. J. Bowdler, R. S. Martin, G. Peters, and J. H. Wilkinson. Solution of real and complex systems of linear equations. Numer. Math., 8:217–234, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 93–110], Contribution I/7.

Bjoerstad:1992:EMM

[BMSV92] P. Bjørstad, F. Manne, T. Sørevik, and M. Vajteršic. Efficient matrix multiplication on SIMD computers. SIAM J. Matrix Anal. Appl., 13(1):386–401, 1992.

Bargmann:1946:SLS

[BMvN46] V. Bargmann, D. Montgomery, and J. von Neumann. Solution of linear systems of high order. Report prepared for Navy Bureau of Ordnance, ????, 1946. ???? pp. Reprinted in [Tau63, pp. 421– 477].

Bardsley:1986:FTM

[BMW86] W. G. Bardsley, P. B. McGinlay, and A. J. Wright. The F test for model discrimination with exponential functions. Biometrika, 73:501–508, 1986. ISSN 0006-3444 (print), 1464-3510 (electronic).

Bell:1971:CSR

[BN71] C. Gordon Bell and Allen Newell. Computer Structures: Readings and Examples. McGraw-Hill, New York, 1971.

ISBN 0-07-004357-4. xix + 668 pp. LCCN TK7888.3 .B4.

Byers:1987:ARP

[BN87a] Ralph Byers and S. G. Nash. Approaches to robust pole assignment. Technical report, Technical Report 87-101, Department of Operations Research and Applied Statistics, George Mason University, Virginia, ????, 1987. ???? pp.

Byers:1987:SVL

[BN87b] Ralph Byers and Stephen Nash. On the singular “vectors” of the Lyapunov operator. SIAM J. Alg. Disc. Meth., 8 (1):59–66, 1987.

Boas:1966:CLR

[Boa66] R. P. Boas. Capitalization of language references. Comm. ACM, 9(6):409, 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boas:1979:SN

[Boa79] R. P. Boas. The Skewes number.

In R. Honsberger, editor, Mathematical Plums, number 4 in Dolciani Mathematical Expositions, page ?? Mathematical Association of America, Washington, DC, USA, 1979.

Boas:1981:CWM

[Boa81] Ralph P. Boas. Can we make mathematics intelligible? Amer. Math. Monthly, 88:727–731, 1981.

Boas:19xx:CDC

[Boaxx] R. P. Boas. Convergence, divergence, and the computer. In ??, page ?? C??, 19xx.

Bodewig:1949:RRE

[Bod49] E. Bodewig. Review of “Rounding-Off Errors in Matrix Processes” by A. M.

Turing. Math. Rev., 10:405, 1949. CO-DEN MAREAR. ISSN 0025-5629.

Bodewig:1956:MC

[Bod56] E. Bodewig. Matrix Calculus. North-Holland (Amsterdam) and InterScience (New York), ????, 1956. ???? pp.

Bodanis:1986:SH

[Bod86] D. Bodanis. The Secret House. Simon and Schuster, New York, NY, USA, 1986. ???? pp.

Boggs:1974:NAC

[Bog74] P. T. Boggs. A new algorithm for the Chebyshev solution of overdetermined linear systems. Math. Comp., 28:203–217, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bohte:1975:BRE

[Boh75] Z. Bohte. Bounds for rounding errors in the Gaussian elimination for band systems. J. Inst. Maths Applics, 16:133–142, 1975.

Bohlender:1977:FPC

[Boh77] Gerd Bohlender. Floating-point computation of functions with maximum accuracy. IEEE Trans. Comput., C-26(7): 621–632, 1977.

Boisvert:1990:GAM

[Boi90] Ronald F. Boisvert. The guide to available mathematical software advisory system. In E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors, Intelligent Mathematical Software Systems, pages 167–177. North-Holland, Amsterdam, The Netherlands, 1990.

Bojanczyk:1984:CSL

[Boj84] A. Bojańczyk. Complexity of solving linear systems in different models of computation. SIAM J. Sci. Stat. Comput., 21:591–603, 1984.

Bollen:1984:NSD

[Bol84] Jo A. M. Bollen. Numerical stability of descent methods for solving linear equations. Numer. Math., 43:361–377, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Boley:1985:PRL

[Bol85] D. Boley. A perturbation result for linear control problems. SIAM J. Alg. Disc. Meth., 6(1):66–72, 1985.

Bollobas:1986:LM

[Bol86] Béla Bollobás, editor. Littlewood’s Miscellany. Cambridge University Press, ????, 1986. ISBN 0-521-33702-X. ???? pp.

Bolsky:1988:BST

[Bol88] Morris I. Bolsky. Better Scientific and Technical Writing. Prentice-Hall, Englewood Cliffs, NJ, USA, 1988. ISBN 0-13-074253-8. ???? pp.

Boman:1993:EKC

[Bom93] Erik Boman. Experiences on the KSR1 computer. NAS Systems Division Technical Report RNR-93-008, NASA Ames Research Center, Moffett Field, CA, USA, April 1993. ???? pp.

Bondi:1991:NAM

[Bon91] Christine Bondi, editor. New Applications of Mathematics. Penguin, London, 1991. ISBN 0-14-012491-8. x + 289 pp.

Booth:1984:CSW

[Boo84] Vernon Booth. Communicating in Science: Writing and Speaking. Cambridge University Press, ????, 1984. ISBN 0-521-27771-X. ???? pp.

Booth:1993:CSW

[Boo93] Vernon Booth. Communicating in Science: Writing a Scientific Paper and Speaking at Scientific Meetings. Cambridge University Press, ????, second edition, 1993. ISBN 0-521-42915-3. ???? pp.

Borovkov:1994:REE

[Bor94] K. A. Borovkov. Russian–English English–Russian Dictionary on Probability, Statistics, and Combinatorics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-316-1. viii + 154 pp.

Boswell:1986:LSJ

[Bos86] James Boswell, editor. The Life of Samuel Johnson. Penguin, London, 1986. ISBN 0-14-043116-0. ???? pp. Edited and abridged with an introduction and notes by Christopher Hibbert.

Bouldin:1973:OUP

[Bou73a] R. Bouldin. Operators with a unique positive near-approximant. Indiana Univ. Math. J., 23:421–427, 1973. CODEN IUMJAB. ISSN 0022-2518.

Bouldin:1973:PA

[Bou73b] R. Bouldin. Positive approximants. Trans. Amer. Math. Soc., 177:391–403, 1973.

Bourne:1978:US

[Bou78] S. R. Bourne. The UNIX shell. The Bell System Technical Journal, 57:1971– 1990, 1978. CODEN BSTJAN. ISSN 0005-8580.

Bowden:1953:OTM

[Bow53] B. V. Bowden. The organization of a typical machine. In B. V. Bowden, editor, Faster than Thought: A Symposium on Digital Computing Machines, pages 67– 77. Pitman Publishing Ltd., London, UK, 1953.

Bowler:1990:VH

[Bow90] Jane Bowler. The Vegetarian Handbook. The Vegetarian Society UK Ltd., ????, 1990. ???? pp.

Box:1965:NMC

[Box65] M. J. Box. A new method of constrained optimization and a comparison with other methods. The Computer Journal, 8:42–52, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Box:1966:CSC

[Box66] M. J. Box. A comparison of several current optimization methods, and the use of transformations in constrained problems. The Computer Journal, 9:67– 77, 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Boyer:1968:HM

[Boy68] Carl B. Boyer. A History of Mathematics. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1968. ISBN 0-691-02391-3. xv + 717 pp. Reprinted by Princeton University Press, Princeton, NJ, USA, 1985.

Boyd:1974:PMN

[Boy74] David W. Boyd. The power method for ℓ^{p} norms. Linear Algebra and Appl., 9:95–101, 1974.

Boyle:1994:AFS

[Boy94] Jeff Boyle. An application of Fourier series to the most significant digit problem. Amer. Math. Monthly, 101(11):879– 886, 1994.

Ballester:1967:CDA

[BP67] C. Ballester and V. Pereyra. On the construction of discrete approximations to linear differential expressions. Math. Comp., 21:297–302, 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bjoerck:1970:SVS

[BP70] Åke Björck and Victor Pereyra. Solution of Vandermonde systems of equations. Math. Comp., 24(112):893–903, 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunch:1971:DMS

[BP71] James R. Bunch and Beresford N. Parlett. Direct methods for solving symmetric indefinite systems of linear equations. SIAM J. Numer. Anal., 8(4):639– 655, 1971.

Berman:1979:NMM

[BP79] A. Berman and R. J. Plemmons. Nonnegative Matrices in the Mathematical Sciences. Academic Press, New York, 1979. ???? pp.

Baker:1981:NSN

[BP81] C. T. H. Baker and C. Phillips, editors. The Numerical Solution of Nonlinear Problems. Oxford University Press, ????, 1981. ???? pp.

Bjoerck:1992:LRO

[BP92a] Åke Björck and C. C. Paige. Loss and recapture of orthogonality in the modified Gram–Schmidt algorithm. SIAM J. Matrix Anal. Appl., 13(1):176–190, 1992.

Brawer:1992:LAP

[BP92b] Robert Brawer and Magnus Pirovino.

The linear algebra of the Pascal matrix. Linear Algebra and Appl., 174:13– 23, 1992.

Berman:1994:NMM

[BP94a] Abraham Berman and Robert J.

Plemmons. Nonnegative Matrices in the Mathematical Sciences, volume 9 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-321-8. xx + 340 pp. LCCN QA188 .B47 1994. URL http:

//www.loc.gov/catdir/enhancements/
fy0708/94037449-d.html;   http://
www.loc.gov/catdir/enhancements/
fy0708/94037449-t.html.  Revised

reprint of [BP79].

Bini:1994:PMC

[BP94b] Dario Bini and Victor Y. Pan.

Polynomial and Matrix Computations. Volume 1: Fundamental Algorithms. Birkhäuser, Boston, MA, USA, 1994. ISBN 0-8176-3786-9. xvi + 415 pp.

Bjoerck:1994:SAL

[BP94c] Åke Björck and C. C. Paige. Solution of augmented linear systems using orthogonal factorizations. BIT, 34:1–24, 1994. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjoerck:1992:ADL

[BPE92] Åke Björck, H. Park, and L. Eldén. Accurate downdating of least squares solutions. IMA Preprint Series # 947, Institute for Mathematics and its Applications, University of Minnesota, Minneapolis, MN, USA, April 1992. ???? pp.

Bischof:1990:SCD

[BPT90] C. H. Bischof, C.-T. Pan, and P. T. P. Tang. Stable Cholesky up- and downdating algorithms for systolic and SIMD architectures. Technical report, Preprint MCS-P167-0790, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1990. ???? pp.

Brookes:1967:EFL

[BR67] H. F. Brookes and H. Ross. English as a Foreign Language for Science Students, volume 1. Heinemann, London, 1967. ???? pp.

Brookes:1968:EFL

[BR68] H. F. Brookes and H. Ross. English as a Foreign Language for Science Students, volume 2. Heinemann, London, 1968. ???? pp.

Brown:1969:CB

[BR69] W. S. Brown and P. L. Richman. The choice of base. Comm. ACM, 12(10):560– 561, 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1971:IPD

[BR71a] F. L. Bauer and C. Reinsch. Inversion of positive definite matrices by the Gauss–Jordan method. In J. H. Wilkinson and C. Reinsch, editors, Linear Algebra, volume II of Handbook for Automatic Computation, pages 45–49. Springer-Ver-lag, Berlin, 1971. Contribution I/3.

Benschop:1971:MSE

[BR71b] N. F. Benschop and H. C. Ratz. A mean square estimate of the generated roundoff error in constant matrix iterative processes. J. Assoc. Comput. Mach., 18(1):48–62, 1971.

Bunch:1976:SMC

[BR76] J. R. Bunch and D. J. Rose. Sparse Matrix Computations. Academic Press, New York, 1976. ???? pp.

Bank:1977:MAE

[BR77] Randolph E. Bank and Donald J.

Rose. Marching algorithms for elliptic boundary value problems. I: The constant coefficient case. SIAM J. Numer. Anal., 14(5):792–829, 1977.

Barron:1987:TPT

[BR87] D. Barron and M. Rees. Text Processing and Typesetting with UNIX. Addison-Wesley, Wokingham, England, 1987. ???? pp.

Battle:1993:ZAR

[BR93] Mary Vroman Battle and Ernest A.

Rakow. Zen and the art of reporting differences in data that are not statistically significant. IEEE Trans. Prof. Commun., 36(2):75–80, 1993.

Bhatia:1996:HWS

[BR96] Rajendra Bhatia and Peter Rosenthal.

How and why to solve the operator equation AX − XB = Y. Bull. London Math. Soc., page ??, 1996. ISSN 0024-6093 (print), 1469-2120 (electronic). To appear.

Brandreth:1983:CPS

[Bra83] Gyles Brandreth. The Complete Public Speaker. Sheldon Press, London, 1983. ???? pp.

Braunwald:1987:ASF

[Bra87] Eugene Braunwald. On analysing scientific fraud. Nature, 325:215–216, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Braams:1993:IQ

[Bra93] Johannes Braams. International quotations. TEX and TUG News, 2(3):5–9, 1993. ISSN 1065-240X.

Braconnier:1995:IOB

[Bra95] Thierry Braconnier. Influence of orthogonality on the backward error and the stopping criterion for Krylov methods. Numerical analysis report, Manchester Centre for Computational Mathematics, Manchester, England, ???? 1995. ?? pp. In preparation.

Beaton:1976:ARS

[BRB76] Albert E. Beaton, Donald B. Rubin, and John L. Barone. The acceptability of regression solutions: Another look at computational accuracy. J. Amer. Statist. Assoc., 71(353):158–168, 1976.

Brent:1970:AMM

[Bre70a] R. P. Brent. Algorithms for matrix multiplication. Technical Report CS 157, Department of Computer Science, Stanford University, Stanford, CA, USA, March 1970. ii + 52 pp.

Brent:1970:EAA

[Bre70b] R. P. Brent. Error analysis of algorithms for matrix multiplication and triangular decomposition using Winograd’s identity. Numer. Math., 16:145–156, 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brent:1973:AMD

[Bre73a] R. P. Brent. Algorithms for Minimization Without Derivatives. Prentice-Hall, Englewood Cliffs, NJ, USA, 1973. ???? pp.

Brent:1973:PAV

[Bre73b] Richard P. Brent. On the precision attainable with various floating-point number systems. IEEE Trans. Comput., C-22(6):601–607, 1973.

Brent:1974:PEG

[Bre74] Richard P. Brent. The parallel evaluation of general arithmetic expressions. J. Assoc. Comput. Mach., 21(2):201–206, 1974.

Brent:1976:MPZ

[Bre76] Richard P. Brent. Multiple-precision zero-finding methods and the complexity of elementary function evaluation. In J. F. Traub, editor, Analytic Computational Complexity, pages 151–176. Academic Press, New York, 1976. ISBN 0-12-697560-4.

Brent:19xx:FMP

[Bre78a] Richard P. Brent. A Fortran multiple-precision arithmetic package. ACM Transactions on Mathematical Software, 4(1):57–70, 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brent:19xx:AMF

[Bre78b] Richard P. Brent. ALGORITHM 524: MP, A Fortran multiple-precision arithmetic package. ACM Transactions on Mathematical Software, 4(1):71–81, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brent:1991:PAT

[Bre91] Richard P. Brent. Parallel algorithm for Toeplitz systems. In G. H. Golub and P. M. Van Dooren, editors, Numerical Linear Algebra, Digital Signal Processing and Parallel Algorithms, volume F70 of NATO ASI Series, pages 75–92. Springer-Verlag, Berlin, 1991.

Briggs:1987:MT

[Bri87] W. L. Briggs. A Multigrid Tutorial. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987. ???? pp.

Briggs:1991:PCF

[Bri91] K. Briggs. A precise calculation of the Feigenbaum constants. Math. Comp., 57: 435–439, 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Britton:1992:CWM

[Bri92] J. L. Britton, editor. Collected Works of A. M. Turing: Pure Mathematics. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88059-3. xxii + 287 pp.

Bleher:1987:SFE

[BRK^{+}87] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. A study of a FORTRAN extension for engineering/scientific computation with access to ACRITH. Computing, 39:93–110, 1987. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Browne:1967:OPR

[Bro67] Michael W. Browne. On oblique Procrustes rotation. Psychometrika, 32(2): 125–132, 1967. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Brock:1968:OMD

[Bro68] John E. Brock. Optimal matrices describing linear systems. AIAA Journal, 6 (7):1292–1296, 1968. CODEN AIAJAH. ISSN 0001-1452.

Browne:1972:ORP

[Bro72] M. W. Browne. Orthogonal rotation to a partially specified target. British J. Math. Statist. Psychol., 25:115–120, 1972. CODEN BJMSAK. ISSN 0007-1102.

Broyden:1973:SCN

[Bro73] C. G. Broyden. Some conditionnumber bounds for the Gaussian elimination process. J. Inst. Maths Applics, 12: 273–286, 1973.

Brooks:1975:MMM

[Bro75] Frederick P. Brooks, Jr. The Mythical Man-Month: Essays on Software Engineering. Addison-Wesley, Reading, MA, USA, 1975. ISBN 0-201-00650-2. xi + 195 pp.

Brown:1978:LE

[Bro78] George Brown. Lecturing and Explaining. Methuen, London, 1978. ???? pp.

Brown:1979:WIC

[Bro79] P. J. Brown. Writing Interactive Compilers and Interpreters. John Wiley, New York, NY, USA, 1979. ???? pp.

Brown:1981:SRM

[Bro81] W. S. Brown. A simple but realistic model of floating-point arithmetic. ACM Transactions on Mathematical Software, 7(4):445–480, 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Broyden:1985:NSH

[Bro85] C. G. Broyden. On the numerical stability of Huang’s and related methods. Journal of Optimization Theory and Applications, 47(4):401–412, 1985. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Brown:1989:APG

[Bro89] G. A. Brown. The Airline Passenger’s Guerrilla Handbook. Blakes, Washington, DC, 1989. ???? pp.

Brocklehurst:1991:SB

[Bro91a] E. R. Brocklehurst. Survey of benchmarks. NPL Report DITC 192/91, National Physical Laboratory, Teddington, UK, November 1991. ???? pp.

Bronstein:1991:CAI

[Bro91b] Manuel Bronstein. Computer algebra and indefinite integrals. In Kenneth R. Meyer and Dieter S. Schmidt, editors, Computer Aided Proofs in Analysis, volume 28 of IMA Volumes in Mathematics and its Applications, pages 33–42. Springer-Verlag, New York, 1991.

Brown:1991:TCA

[Bro91c] Peter N. Brown. A theoretical comparison of the Arnoldi and GMRES algorithms. SIAM J. Sci. Stat. Comput., 12 (1):58–78, 1991.

Bleher:1985:AHA

[BRR85] J. H. Bleher, A. E. Roeder, and S. M.

Rump. ACRITH: High-accuracy arithmetic. An advanced tool for numerical computation. In Kai Hwang, editor, Proceedings of the 7th Symposium on Computer Arithmetic, pages 318–321. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985.

Brumfit:1982:EIC

[Bru82] Christopher J. Brumfit, editor. English for International Communication. Pergamon, New York, NY, USA, 1982. ISBN 0-08-028613-5. ???? pp.

Brunet:1989:CFL

[Bru89] Marie-Christine Brunet. Contribution à la Fiabilité de Logiciels Numériques et à L’analyse de Leur Comportement: Une Approche Statistique. PhD thesis, Université de Paris IX Dauphine, U.E.R. Mathématiques de la Décision, ????, January 1989. viii + 214 pp.

Brualdi:1992:IC

[Bru92] R. A. Brualdi. Introductory Combinatorics. North-Holland, Amsterdam, The Netherlands, second edition, 1992. ???? pp.

Bruaset:1995:SPI

[Bru95a] Are Magnus Bruaset. A Survey of Preconditioned Iterative Methods, volume 328 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, UK, 1995. ISBN 0-582-27654-3. xii + 162 pp.

Bruehl:1995:CTA

[Brü95b] Martin Brühl. A curve tracing algorithm for computing the pseudospectrum.

Manuscript. Submitted to BIT, October 1995.

Bryson:1987:PDT

[Bry87] Bill Bryson. The Penguin Dictionary of Troublesome Words. Penguin, London, second edition, 1987. ISBN 0-14-051200-4. ???? pp.

Bryson:1989:LCT

[Bry89] Bill Bryson. The Lost Continent: Travels In Small Town America. Abacus, London, 1989. ISBN 0-349-10056-X. ???? pp.

Bryson:1990:MTE

[Bry90] Bill Bryson. Mother Tongue: English and How it Got That Way. Avon Books, New York, NY, USA, 1990. ISBN 0-380-71543-0. ???? pp.

Bryson:1991:NHN

[Bry91] Bill Bryson. Neither Here Nor There: Travels in Europe. Minerva, London, 1991. ISBN 0-7493-9815-9. ???? pp.

Barnett:1968:SAL

[BS68] S. Barnett and C. Storey. Some applications of the Lyapunov matrix equation. J. Inst. Maths Applics, 4:33–42, 1968.

Bartels:1972:ASM

[BS72] R. H. Bartels and G. W. Stewart. Algorithm 432: Solution of the matrix equation AX +XB = C. Comm. ACM, 15(9): 820–826, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blevins:1974:CED

[BS74] M. M. Blevins and G. W. Stewart. Calculating the eigenvectors of diagonally dominant matrices. J. Assoc. Comput. Mach., 21:261–271, 1974.

Baur:1983:CPD

[BS83] W. Baur and V. Strassen. The complexity of partial derivatives. Theoret. Comput. Sci., 22:317–330, 1983. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berry:1989:OPA

[BS89a] Michael Berry and Ahmed H. Sameh.

An overview of parallel algorithms for the singular value and symmetric eigenvalue problems. J. Comp. Appl. Math., 27:191– 213, 1989.

Bischof:1989:USS

[BS89b] C. H. Bischof and G. M. Shroff. On updating signal subspaces. Technical report, Preprint MCS-P101-0989, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1989. ???? pp.

Beattie:1992:OMA

[BS92a] C. A. Beattie and S. W. Smith.

Optimal matrix approximants in structural identification. Journal of Optimization Theory and Applications, 74(1):23– 56, 1992. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Bischof:1992:DCM

[BS92b] Christian H. Bischof and Xiaobai Sun. A divide-and-conquer method for tridiagonalizing symmetric matrices with repeated eigenvalues. Preprint MCS-P286-0192, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, March 1992. ???? pp.

Bramley:1992:RPM

[BS92c] R. Bramley and A. Sameh. Row projection methods for large nonsymmetric linear systems. SIAM J. Sci. Stat. Comput., 13(1):168–193, 1992.

Baskette:1992:AE

[BSB92] Floyd K. Baskette, Jack Z. Sissors, and Brian S. Brooks. The Art of Editing. MacMillan Publishing Company, New York, NY, USA, fifth edition, 1992. ???? pp.

Bailey:1989:FPA

[BSBF89] David H. Bailey, Horst D. Simon, John T. Barton, and Martin J. Fouts. Floating point arithmetic in future supercomputers. Int. J. Supercomputer Appl., 3(3):86–90, 1989. CODEN IJSAE9. ISSN 0890-2720.

Barron:1960:SSL

[BSD60] D. W. Barron and H. P. F. Swinnerton-Dyer. Solution of simultaneous linear equations using a magnetictape store. The Computer Journal, 3(1): 28–33, 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Bauer:1961:AMN

[BSW61] F. L. Bauer, J. Stoer, and C. Witzgall. Absolute and monotonic norms. Numer. Math., 3:257–264, 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bickley:1964:MTM

[BT64] W. G. Bickley and R. S. H. G. Thompson. Matrices: Their Meaning and Manipulation. Van Nostrand, Princeton, New Jersey, 1964. ???? pp. My copy signed by Givens.

Bertsekas:1989:PDC

[BT89] Dimitri P. Bertsekas and John N. Tsitsiklis. Parallel and Distributed Computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1989. ???? pp.

Buchanan:1992:NMA

[BT92] James L. Buchanan and Peter R.

Turner. Numerical Methods and Analysis. McGraw-Hill, New York, 1992. ISBN 0-07-008717-2, 0-07-112922-7 (international paperback edition). xv + 751 pp.

Buchholz:1959:FFC

[Buc59] W. Buchholz. Fingers or fists? (The choice of decimal or binary representation). Comm. ACM, 2(12):3–11, 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Buckley:1977:SCS

[Buc77] A. Buckley. On the solution of certain skew symmetric linear systems. SIAM J. Numer. Anal., 14:566–570, 1977.

Buchholz:1981:OWE

[Buc81] W. Buchholz. Origin of the word byte.

Annals of the History of Computing, 3(1): 72, 1981.

Buckley:1994:AXF

[Buc94a] A. G. Buckley. Algorithm XXX: A Fortran 90 code for unconstrained nonlinear minimization. ACM Transactions on Mathematical Software, 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). To appear, 1994.

Buckley:1994:CFC

[Buc94b] A. G. Buckley. Conversion to Fortran 90: A case study. ACM Transactions on Mathematical Software, 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). To appear, 1994.

Budgett:1987:WED

[Bud87] H. Budgett. What every desktop publisher should know. Personal computer world, 10:146–150, 1987.

Buehler:1977:RCT

[Bue77] Mary Fran Buehler. Report construction: Tables. IEEE Trans. Prof. Commun., PC-20(1):29–32, 1977.

Bultheel:1980:EAI

[Bul80] A. Bultheel. Error analysis of incoming and outgoing schemes for the trigonometric moment problem. Technical report, Proc. of the Conference on Padé Approximation and its Applications, Lecture Notes in Mathematics 888, Springer-Verlag, Berlin, ????, 1980. ???? pp.

Bull:1991:AIL

[Bul91] J. M. Bull. Asynchronous iterations on local memory parallel computers. M.Sc. Thesis, University of Manchester, Manchester, England, October 1991. ???? pp.

Bunch:1971:ADP

[Bun71a] James R. Bunch. Analysis of the diagonal pivoting method. SIAM J. Numer. Anal., 8(4):656–680, 1971.

Bunch:1971:ESM

[Bun71b] James R. Bunch. Equilibration of symmetric matrices in the max-norm. J. Assoc. Comput. Mach., 18(4):566–572, 1971.

Bunch:1974:PPS

[Bun74] James R. Bunch. Partial pivoting strategies for symmetric matrices. SIAM J. Numer. Anal., 11(3):521–528, 1974.

Bunch:1976:BMS

[Bun76] James R. Bunch. Block methods for solving sparse linear systems. In J. R. Bunch and D. J. Rose, editors, Sparse Matrix Computations, pages 39–58. Academic Press, New York, 1976.

Bunch:1982:NSD

[Bun82] James R. Bunch. A note on the stable decomposition of skew-symmetric matrices. Math. Comp., 38(158):475–479, 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunch:1985:SMS

[Bun85] James R. Bunch. Stability of methods for solving Toeplitz systems of equations. SIAM J. Sci. Stat. Comput., 6(2): 349–364, 1985.

Bunch:1987:WSS

[Bun87] James R. Bunch. The weak and strong stability of algorithms in numerical linear algebra. Linear Algebra and Appl., 88/89:49–66, 1987.

Burrill:1974:CPS

[Bur74] Donald F. Burrill. Computational precision in statistics: Variances and covariances. Journal of Statistical Computation and Simulation, 3(2):105–115, 1974. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burton:1980:ENT

[Bur80] David M. Burton. Elementary Number Theory. Allyn and Bacon, Needham Heights, MA, USA, 1980. ISBN 0-205-06978-9. ???? pp.

Burleigh:1987:TFC

[Bur87] D. W. Burleigh. Three fortran 77 compilers. BYTE Magazine, 12:187–194, 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Burchfield:1989:UEL

[Bur89] Robert Burchfield. Unlocking the English Language. Faber and Faber, London, 1989. ISBN 0-571-14416-0. ???? pp.

Burrage:1990:ANI

[Bur90] K. Burrage. An adaptive numerical integration code for a chain of transputers. Parallel Computing, 16:305–312, 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Burrage:1995:PSM

[Bur95] Kevin Burrage. Parallel and Sequential Methods for Ordinary Differential Equations. Oxford University Press, ????, 1995. ISBN 0-19-853432-9. xvi + 446 pp.

Businger:1968:MWC

[Bus68] P. A. Businger. Matrices which can be optimally scaled. Numer. Math., 12:346–348, 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Businger:1969:RMH

[Bus69a] P. A. Businger. Reducing a matrix to Hessenberg form. Math. Comp., 23: 819–821, 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Businger:1969:EPB

[Bus69b] Peter A. Businger. Extremal properties of balanced tri-diagonal matrices. Math. Comp., 23:193–195, 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Businger:1971:MNS

[Bus71] P. A. Businger. Monitoring the numerical stability of Gaussian elimination. Numer. Math., 16:360–361, 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bush:1990:MCP

[Bus90] Stephen G. Bush. The mostcited physical-sciences publications in the 1945–1954 Science Citation Index. Part 2. Mathematics. Current Contents: Physical, Chemical & Earth Sciences, (42):381– 386, 1990. October 15.

Butcher:1981:CEC

[But81] Judith Butcher. Copy-Editing: The Cambridge Handbook. Cambridge University Press, ????, second edition, 1981. ISBN 0-521-25638-0. ???? pp.

Butcher:1987:NAO

[But87] J. C. Butcher. The Numerical Analysis of Ordinary Differential Equations: Runge–Kutta and General Linear Methods. Wiley, Chichester, UK, 1987. ISBN 0-471-91046-5. xv + 512 pp.

Butcher:1992:CEC

[But92a] Judith Butcher. Copy-Editing: The Cambridge Handbook for Editors, Authors and Publishers. Cambridge University Press, ????, third edition, 1992. ISBN 0-521-40074-0. ???? pp.

Butler:1992:IMS

[But92b] B. P. Butler. Interfacing Matlab and software written in other languages. NPL Report DITC 200/92, National Physical Laboratory, Teddington, UK, May 1992. ???? pp.

Buzan:1982:UYH

[Buz82] Tony Buzan. Use Your Head. Ariel Books, British Broadcasting Corporation, ????, 1982. ISBN 0-563-16552-9. 156 pp.

Buzbee:1992:PER

[Buz92] Bill Buzbee. Poisson’s equation revisited. Current Contents: Physical, Chemical & Earth Sciences, (36):8, 1992. No. 36, Sept. 7.

Bischof:1986:CSV

[BV86] C. H. Bischof and C. F. Van Loan. Computing the singular value decomposition on a ring of array processors. In Large Scale Eigenvalue Problems [CW86], pages 51–66.

Bischof:1987:WRP

[BV87] Christian H. Bischof and Charles F. Van Loan. The WY representation for products of Householder matrices. SIAM J. Sci. Stat. Comput., 8(1):s2–s13, 1987.

Barlow:1992:NDC

[BV92a] Jesse L. Barlow and Udaya B. Vemulapati. A note on deferred correction for equality constrained least squares problems. SIAM J. Numer. Anal., 29(1):249– 256, 1992.

Barlow:1992:RDM

[BV92b] Jesse L. Barlow and Udaya B. Vemulapati. Rank detection methods for sparse matrices. SIAM J. Matrix Anal. Appl., 13 (4):1279–1297, 1992.

Beam:1993:ASB

[BW93] Richard M. Beam and Robert F. Warming. The asymptotic spectra of banded Toeplitz and quasi-Toeplitz matrices. SIAM J. Sci. Comput., 14(4):971– 1006, 1993.

Byers:1984:LSC

[Bye84] Ralph Byers. A LINPACK-style condition estimator for the equation AX − XB^{T} = C. IEEE Trans. Automat. Control., AC-29(10):926–928, 1984.

Byers:1985:NCA

[Bye85] Ralph Byers. Numerical condition of the algebraic Riccati equation. In B. N. Datta, editor, Linear Algebra and its Role in Systems Theory, volume 47 of Contemporary Math., pages 35–49. American Mathematical Society, Providence, RI, USA, 1985.

Byers:1986:NSI

[Bye86] Ralph Byers. Numerical stability and instability in matrix sign function based algorithms. In Christopher I. Byrnes and Anders Lindquist, editors, Computational and Combinatorial Methods in Systems Theory, pages 185–200. Elsevier North Holland, Inc., New York, NY, USA, 1986.

Byers:1987:SAR

[Bye87] Ralph Byers. Solving the algebraic Riccati equation with the matrix sign function. Linear Algebra and Appl., 85: 267–279, 1987.

Byers:1988:BMM

[Bye88] Ralph Byers. A bisection method for measuring the distance of a stable matrix to the unstable matrices. SIAM J. Sci. Stat. Comput., 9:875–881, 1988.

Byrne:1987:HIT

[Byr87] George D. Byrne. How to improve technical presentations. SIAM News, 20: 10–11, January 1987. ISSN 0036-1437.

Byrne:1988:BTA

[Byr88] Robert Byrne, editor. 1,911 Best Things Anybody Ever Said. Fawcett Columbine, New York, 1988. ISBN 0-449-90285-4. xiv + 461 pp.

Barlow:1990:EAU

[BZ90] Jesse L. Barlow and Richard J. Zaccone. Error analysis in unnormalized floating point arithmetic. Technical Report CS-90-23, Department of Computer Science, Pennsylvania State University, University Park, PA, April 1990. ???? pp.

Breuer:1993:NML

[BZ93] Shlomo Breuer and Gideon Zwas. Numerical Mathematics—A Laboratory Approach. Cambridge University Press, ????, 1993. ISBN 0-521-44040-8. xiii + 267 pp.

Barlow:1993:SCA

[BZY93] Jesse L. Barlow, Hongyuan Zha, and Peter A. Yoon. Stable chasing algorithms for modifying complete and partial singular value decompositions. Technical Report CSE-93-19, Department of Computer Science, Pennsylvania State University, University Park, PA, December 1993. ???? pp.

Cadzow:1973:FAM

[Cad73] J. A. Cadzow. A finite algorithm for the minimum l_{∞} solution to a system of consistent linear equations. SIAM J. Numer. Anal., 10:607–617, 1973.

Cadzow:1974:EAP

[Cad74] J. A. Cadzow. An efficient algorithmic procedure for obtaining a minimum l_{∞}-norm solution to a system of consistent linear equations. SIAM J. Numer. Anal., 11:1151–1165, 1974.

Cadzow:1988:SEC

[Cad88] J. A. Cadzow. Signal enhancement— A composite property mapping algorithm. IEEE Trans. Acoustics, Speech, and Signal Processing, 36(1):49–62, 1988. CODEN IETABA. ISSN 0096-3518.

Caffney:1963:ATM

[Caf63] John Caffney. Another test matrix for determinants and inverses. Comm. ACM, 6(6):310, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cahn:1989:NSC

[Cah89] Sammy Cahn. The New Sammy Cahn Songbook. Warner Bros., ????, 1989. ???? pp.

Cain:1984:SSP

[Cai84] Bryan E. Cain. Some spectral properties of polar decompositions. Canadian Journal of Mathematics, 36(6):973–985, 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cajori:1925:ACM

[Caj25] Florian Cajori. American contributions to mathematical symbolism. Amer. Math. Monthly, 32:414–416, 1925. Reprinted in [Ewi94].

Cajori:1929:HMN

[Caj29] Florian Cajori. A History of Mathematical Notations. Volume II. Notations Mainly in Higher Mathematics.

The Open Court Publishing Company, Chicago, 1929. ???? pp.

Calahan:1986:BOL

[Cal86] D. A. Calahan. Block-oriented, localmemory-based linear equation solution on the Cray-2: uniprocessor algorithms. Technical report, Proceedings of 1986 International Conference on Parallel Processing, Washington DC, IEEE Computer Society Press, ????, 1986. ???? pp.

Caprani:1971:ILR

[Cap71] Ole Caprani. Implementation of a low round-off summation method. BIT, 11: 271–275, 1971. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Carey:1958:MSB

[Car58] G. V. Carey. Mind the Stop: A Brief Guide to Punctuation with a Note on Proof-Correction. Penguin, London, second edition, 1958. ISBN 0-14-051072-9. ???? pp. Reprinted 1976.

Carr:1959:EAF

[Car59] John W. Carr III. Error analysis in floating point arithmetic. Comm. ACM, 2(5):10–15, 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carey:1989:PSM

[Car89] G. F. Carey. Parallel Supercomputing: Methods, Algorithms and Applications. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1989. ???? pp.

Carter:1991:MFP

[Car91] Russell Carter. Y-MP floating point and Cholesky factorization. Int Journal of High Speed Computing, 3(3/4):215– 222, 1991. CODEN IHSCEZ. ISSN 0129-0533.

Cash:1979:SRA

[Cas79] J. R. Cash. Stable Recursions: With Applications to the Numerical Solution of Stiff Systems. Academic Press, London, 1979. ???? pp.

Cash:1989:BRK

[Cas89] J. R. Cash. A block 6(4) Runge– Kutta formula for nonstiff initial value problems. ACM Transactions on Mathematical Software, 15:15–28, 1989. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). Algorithm 669 description, pp. 29–30.

Casti:1991:VTT

[Cas91] J. L. Casti. VTEX typesetting package: A TEX for tomorrow. Notices Amer. Math. Soc., 38:105–109, 1991. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casti:1993:BFT

[Cas93] John L. Casti. Bitmap-free TEX. Notices Amer. Math. Soc., 40(5):464–469, 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cauchy:1841:EDP

[Cau41] A. Cauchy. Exercices d’Analyse et de Phys. Math., volume 2. ????, Paris, 1841. ???? pp. Cited in [SL54].

Causey:1964:CNM

[Cau64] R. L. Causey. On Closest Normal Matrices. PhD thesis, Stanford University, Stanford, CA, May 1964. ???? pp.

Calnan:1981:SAM

[CB81] James Calnan and Andras Barabas. Speaking at Medical Meetings: A Practical Guide. William Heinemann Medical, London, second edition, 1981. ISBN 0-433-05001-2. ???? pp.

Chatelin:1990:PRE

[CB90a] Françoise Chatelin and Marie-Christine Brunet. A probabilistic roundoff error propagation model. Application to the eigenvalue problem. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 139–160. Oxford University Press, 1990.

Cybenko:1990:HHA

[CB90b] G. Cybenko and M. Berry. Hyperbolic Householder algorithms for factoring structured matrices. SIAM J. Matrix Anal. Appl., 11:499–520, 1990.

Chaundy:1954:PMA

[CBB54] T. W. Chaundy, P. R. Barrett, and Charles Batey. The Printing of Mathematics: Aids for Authors and Editors and Rules for Compositors and Readers at the University Press, Oxford. Oxford University Press, ????, 1954. ???? pp.

Corless:1991:RGO

[CC91] Robert M. Corless and George F. Corliss. Rationale for guaranteed ODE defect control. Preprint MCS-P273-1191, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, December 1991. ???? pp.

Cody:1993:AFS

[CC93] W. J. Cody and Jerome T. Coonen. ALGORITHM 722: Functions to support the IEEE standard for binary floating-point arithmetic. ACM Transactions on Mathematical Software, 19(4): 443–451, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cooke:1982:BSC

[CCC82] D. Cooke, A. H. Craven, and G. M.

Clarke. Basic Statistical Computing. Edward Arnold, London, UK, 1982. ISBN 0-7131-3441-0. ???? pp.

Chaitin-Chatelin:1996:LFP

[CCF96] Françoise Chaitin-Chatelin and Valérie Frayssé. Lectures on Finite Precision Computations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-358-7. ???? pp. In Press.

Cody:1984:PRW

[CCG^{+}84] W. J. Cody, J. T. Coonen, D. M.

Gay, K. Hanson, D. Hough, W. Kahan, R. Karpinski, J. Palmer, F. N. Ris, and D. Stevenson. A proposed radixand word-length-independent standard for floating-point arithmetic. IEEE Micro, 4(4):86–100, 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chambers:1983:GMD

[CCKT83] John M. Chambers, William S.

Cleveland, Beat Kleiner, and Paul A. Tukey. Graphical Methods for Data Analysis. Wadsworth International Group (Belmont, CA) and Duxbury Press (Boston), ????, 1983. ???? pp.

Cline:1982:GLC

[CCV82] Alan K. Cline, Andrew R. Conn, and Charles F. Van Loan. Generalizing the LINPACK condition estimator. In J. P. Hennart, editor, Numerical Analysis, Mexico 1981, volume 909 of Lecture Notes in Mathematics, pages 73–83. Springer-Verlag, Berlin, 1982.

Carpenter:1986:MTA

[CD86] B. E. Carpenter and R. W. Doran, editors. A. M. Turing’s ACE Report of 1946 and Other Papers, volume 10 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-03114-0. vii + 141 pp.

Conte:1980:ENA

[CdB80] Samuel D. Conte and Carl de Boor. Elementary Numerical Analysis: An Algorithmic Approach. McGraw-Hill, Tokyo, third edition, 1980. ISBN 0-07-066228-2. xii + 432 pp.

Cosgrove:1991:AIP

[CDG91] J. D. F. Cosgrove, J. C. Díaz, and A. Griewank. Approximate inverse preconditionings for sparse linear systems. Preprint MCS-P268-1091, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, November 1991. ???? pp.

Crowder:1979:RCE

[CDM79] H. Crowder, R. S. Dembo, and J. M. Mulvey. On reporting computational experiments with mathematical software. ACM Transactions on Mathematical Software, 5:193–203, 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Choi:1994:DIS

[CDO^{+}94] Jaeyoung Choi, Jack J. Dongarra, Susan Ostrouchov, Antoine P. Petitet, David W. Walker, and R. Clint Whaley. The design and implementation of the ScaLAPACK LU, QR and Cholesky factorization routines. Report ORNL/TM-12470, Oak Ridge National Laboratory, Oak Ridge, TN, USA, September 1994. 26 pp. LAPACK Working Note 80.

Choi:1992:SSL

[CDPW92] Jaeyoung Choi, Jack J. Dongarra, Roldan Pozo, and David W. Walker. ScaLAPACK: A scalable linear algebra library for distributed memory concurrent computers. Technical Report CS-92-181, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, November 1992. 8 pp. LAPACK Working Note 55.

Choi:1994:DPD

[CDW94] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. The design of a parallel, dense linear algebra software library: Reduction to Hessenberg, tridiagonal, and bidiagonal form. In Jack. J. Dongarra and Bernard Tourancheau, editors, Environments and Tools For Parallel Scientific Computing, pages 98–111. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Corless:1991:NMC

[CEN91] Robert M. Corless, C. Essex, and M. A. H. Nerenberg. Numerical methods can suppress chaos. Physics Letters A, 157(1):27–36, 1991. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Carpraux:1993:SPN

[CES93] J. Carpraux, J. Erhel, and M. Sadkane. Spectral portrait for non Hermitian large sparse matrices. Technical Report 777, INRIA/IRISA, Unité de Recherche, Rennes, France, 1993. ???? pp.

Chan:1988:EWC

[CF88] Tony F. Chan and David E. Foulser. Effectively well-conditioned linear systems. SIAM J. Sci. Stat. Comput., 9(6): 963–969, 1988.

Chatelin:1991:SSS

[CF91] F. Chatelin and V. Frayssé. A statistical study of the stability of linear systems. Technical report, Report TR/PA/91/43, CERFACS, Toulouse, France, ????, 1991. ???? pp.

Chatelin:1992:ECT

[CF92] Françoise Chatelin and Valérie Frayssé. Elements of a condition theory for the computational analysis of algorithms. In R. Beauwens and P. de Groen, editors, Iterative Methods in Linear Algebra, pages 15–25. Elsevier North Holland, Inc., New York, NY, USA, 1992.

Chatelin:1993:QCEa

[CF93] Françoise Chatelin and Valérie Frayssé. Qualitative computing: Elements of a theory for finite precision computation. Lecture notes, CERFACS, Toulouse, France and THOMSON-CSF, Orsay, France, ????, June 1993. ???? pp.

Chatelin:1993:QCEb

[CFB93] F. Chatelin, V. Frayssé, and T. Braconnier. Qualitative computing: Elements of a theory for finite precision computation. Lecture notes, CERFACS, Toulouse, France, April 1993. ???? pp.

Chartres:1967:CEB

[CG67] Bruce A. Chartres and James C. Geuder. Computable error bounds for direct solution of linear equations. J. Assoc. Comput. Mach., 14(1):63–71, 1967.

Chu:1985:NEE

[CG85] Eleanor Chu and Alan George. A note on estimating the error in Gaussian elimination without pivoting. ACM SIGNUM Newsletter, 20(2):2–7, 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Char:1991:MVL

[CGG^{+}91] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton L. Leong, Michael B. Monagan, and Stephen M. Watt. Maple V Library Reference Manual. Springer-Verlag, Berlin, 1991. ISBN 3-540-97592-6. xxv + 698 pp.

Chan:1983:ACS

[CGL83] Tony F. Chan, Gene H. Golub, and Randall J. LeVeque. Algorithms for computing the sample variance: Analysis and recommendations. The American Statistician, 37(3):242–247, 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Concus:1985:BPC

[CGM85] P. Concus, G. H. Golub, and G. Meurant. Block preconditioning for the conjugate gradient method. SIAM J. Sci. Stat. Comput., 6(1):220–252, 1985.

Concus:1976:GCG

[CGO76] Paul Concus, Gene H. Golub, and Dianne P. O’Leary. A generalized conjugate gradient method for the numerical solution of elliptic partial differential equations. In Sparse Matrix Computations [BR76], pages 309–332. Reprinted in [Gol84, pp. 178–198].

Antonioordova:1990:VMU

[CGR90] Antonio Córdova, Walter Gautschi, and Stephan Ruscheweyh. Vandermonde matrices on the unit circle: Spectral properties and conditioning. Numer. Math., 57:577–591, 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvetti:1993:ACP

[CGR93] D. Calvetti, E. Gallopoulos, and L. Reichel. Accuracy control for parallel evaluation of matrix rational functions. Technical Report 1278, CSRD, University of Illinois, Urbana, 1993? ???? pp.

Curry:1983:IRF

[CGS83] J. H. Curry, L. Garnett, and D. Sullivan. On the iteration of a rational function: Computer experiments with newton’s method. Commun. Math. Phys., 91: 267–277, 1983. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

Conn:1992:LFP

[CGT92a] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. LANCELOT: A Fortran Package for Large-Scale Nonlinear Optimization (Release A). Springer-Verlag, Berlin, 1992. ISBN 0-387-55470-X. xviii + 330 pp.

Conn:1992:NEL

[CGT92b] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. Numerical experiments with the LANCELOT package (Release A) for large-scale nonlinear optimization. Report RAL-92-075, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, November 1992. ???? pp.

Chambers:1968:NGR

[CH68] J. M. Chambers and Agnes M.

Herzberg. A note on the game of refereeing. Applied Statistics, 17(3):260–263, 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1980:ELA

[CH80] M. G. Cox and S. J. Hammarling.

Evaluation of the language Ada for use in numerical computations. NPL Report DNACS 30/80, National Physical Laboratory, Teddington, UK, July 1980. ???? pp.

Ch:1985:DBN

[Ch85] E. Goles Ch. Dynamical behaviour of neural networks. SIAM J. Alg. Disc. Meth., 6(4):749–754, 1985.

Choudhury:1987:COS

[CH87] D. Choudhury and R. A. Horn. A complex orthogonal-symmetric analog of the polar decomposition. SIAM J. Alg. Disc. Meth., 8:219–225, 1987.

Chen:1988:IPP

[CH88a] Pehong Chen and Michael A. Harrison. Index preparation and processing. Software—Practice and Experience, 18(9):897–915, 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Coleman:1988:DAS

[CH88b] T. F. Coleman and L. A. Hulbert. A direct active set algorithm for large sparse quadratic programs with simple bounds. Technical report, Technical Report TR 88-926, Department of Computer Science, Cornell University, ????, 1988. ???? pp.

Cox:1989:OIA

[CH89] M. G. Cox and P. M. Harris. Overcoming an instability arising in a spline approximation algorithm by using an alternative form of a simple rational function. IMA Bulletin, 25(9):228–232, 1989.

Chan:1990:CTS

[CH90a] T. F. Chan and P. C. Hansen. Computing truncated singular value decomposition least squares solutions by rank revealing QR-factorizations. SIAM J. Sci. Stat. Comput., 11:519–530, 1990.

Coleman:1990:CTR

[CH90b] T. F. Coleman and C. Hempel. Computing a trust region step for a penalty function. SIAM J. Sci. Stat. Comput., 11:180–??, 1990.

Cox:1990:RNC

[CH90c] M. G. Cox and S. J. Hammarling, editors. Reliable Numerical Computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$45.00.

Croz:1990:PNL

[CH90d] J. J. Du Croz and S. J. Hammarling. Parallelism and the NAG library. Technical report, NAG Technical Report TR5/90, NAG Ltd., Oxford, ????, 1990. ???? pp.

Chan:1992:LAL

[CH92a] T. F. Chan and P. C. Hansen. A look-ahead Levinson algorithm for indefinite Toeplitz systems. SIAM J. Matrix Anal. Appl., 13(2):490–506, 1992.

Chan:1992:SAR

[CH92b] T. F. Chan and P. C. Hansen. Some applications of the rank revealing QR factorization. SIAM J. Sci. Stat. Comput., 13(3):727–741, 1992.

Croz:1992:SMM

[CH92c] Jeremy J. Du Croz and Nicholas J.

Higham. Stability of methods for matrix inversion. IMA J. Numer. Anal., 12:1–19, 1992.

Christiansen:1994:ECN

[CH94] Søren Christiansen and Per Christian Hansen. The effective condition number applied to error analysis of certain boundary collocation methods. J. Comp. Appl. Math., 54:15–36, 1994.

Chartres:1966:ACP

[Cha66] Bruce A. Chartres. Automatic controlled precision calculations. J. Assoc. Comput. Mach., 13(3):386–403, 1966.

Chargaff:1974:BTB

[Cha74a] Erwin Chargaff. Building the tower of babble. Nature, 248:776–779, 1974. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Chatterjee:1974:NIP

[Cha74b] G. Chatterjee. Negative integral powers of a bidiagonal matrix. Math. Comp., 28:713–714, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chan:1984:ECF

[Cha84] Tony F. Chan. On the existence and computation of LU-factorizations with small pivots. Math. Comp., 42(166):535– 547, 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chan:1987:RRQ

[Cha87] Tony F. Chan. Rank revealing QR factorizations. Linear Algebra and Appl., 88/89:67–82, 1987.

Chan:1988:OCP

[Cha88] Tony F. Chan. An optimal circulant preconditioner for Toeplitz systems. SIAM J. Sci. Stat. Comput., 9(4):766– 771, 1988.

Chambers:1990:CED

[Cha90] Chambers. Chambers English Dictionary. Chambers, Edinburgh, seventh edition, 1990. ISBN 0-550-10250-7. ???? pp.

Chambers:1991:CCD

[Cha91] Chambers. Chambers Concise Dictionary. Chambers, Edinburgh, 1991. ???? pp.

Chatelin:1992:AIN

[Cha92] Françoise Chatelin. About the influence of nonnormality on matrix computations. Talk given at Workshop on Linear Algebra, Markov Chains, and Queuing Models, Institute for Mathematics and its Applications, University of Minnesota, January 1992.

Chambers:1993:CD

[Cha93a] Chambers. The Chambers Dictionary. Chambers, Edinburgh, 1993. ISBN 0-550-10255-8. xviii + 2062 pp.

Chatelin:1993:EM

[Cha93b] Françoise Chatelin. Eigenvalues of Matrices. Pure and applied mathematics. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-93538-7. xviii + 382 pp.

LCCN QA188 .C44 1993. URL http:

//www.loc.gov/catdir/enhancements/
fy0707/93003430-d.html;   http://
www.loc.gov/catdir/toc/onix04/93003430.

html. With exercises by Mario Ahués and the author, Translated from the French and with additional material by Walter Ledermann.

Cheney:1982:IAT

[Che82a] E. W. Cheney. Introduction to Approximation Theory. Chelsea, New York, second edition edition, 1982. ???? pp.

Cherry:1982:WT

[Che82b] Lorinda L. Cherry. Writing tools.

IEEE Transactions on Communications, COM-30(1):100–105, 1982. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Cheney:1986:MAT

[Che86] E. W. Cheney. Multivariate Approximation Theory: Selected topics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1986. ???? pp.

Cohen:1983:CCP

[CHH83] Marty S. Cohen, T. E. Hull, and V. Carl Hamacher. CADAC: A controlled-precision decimal arithmetic unit. IEEE Trans. Comput., C-32(4):370– 377, 1983.

Ching:1993:MDL

[Chi93] Li Ching. The maximum determinant of an n × n lower Hessenberg (0, 1) matrix. Linear Algebra and Appl., 183:147– 153, 1993.

Chow:1969:CHM

[Cho69] T. S. Chow. A class of Hessenberg matrices with known eigenvalues and inverses. SIAM Review, 11(3):391–395, 1969.

Choi:1983:TTH

[Cho83] Man-Duen Choi. Tricks or treats with the Hilbert matrix. Amer. Math. Monthly, 90:301–312, 1983.

Canuto:1988:SMF

[CHQZ88] Claudio Canuto, M. Yousuff Hussaini, Alfio Quarteroni, and Thomas A. Zang. Spectral Methods in Fluid Dynamics. Springer-Verlag, Berlin, 1988. ISBN 3-540-52205-0. xv + 567 pp.

Chu:1986:DEA

[Chu86] M. T. Chu. A differential equation approach to the singular value decomposition of bidiagonal matrices. Linear Algebra and Appl., 80:71–79, 1986.

Chu:1987:SME

[Chu87] King-wah Eric Chu. The solution of the matrix equations AXB − CXD = E and (Y A−DZ, Y C −BZ) = (E, F). Linear Algebra and Appl., 93:93–105, 1987.

Chu:1989:SSL

[Chu89] K.-W. E. Chu. Symmetric solutions of linear matrix equations by matrix decompositions. Linear Algebra and Appl., 119:35–50, 1989.

Chu:1991:CJL

[Chu91a] M. T. Chu. A continuous Jacobi-like approach to the simultaneous reduction of real matrices. Linear Algebra and Appl., 147:75–96, 1991.

Chu:1991:LSA

[Chu91b] M. T. Chu. Least squares approximation by real normal matrices with specified spectrum. SIAM J. Matrix Anal. Appl., 12:115–127, 1991.

Chu:1995:CHM

[Chu95] Moody T. Chu. Constructing a Hermitian matrix from its diagonal entries and eigenvalues. SIAM J. Matrix Anal. Appl., 16(1):207–217, 1995.

Chandrasekaran:1991:RRQ

[CI91] S. Chandrasekaran and I. Ipsen. On rank-revealing QR factorisations. Research Report YALEU/DCS/RR-880, Department of Computer Science, Yale University, ????, December 1991. ???? pp.

Chandrasekaran:1994:BEE

[CI94] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. Backward errors for eigenvalue and singular value decompositions. Numer. Math., 68:215–223, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandrasekaran:1995:AQA

[CI95a] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. Analysis of a QR algorithm for computing singular values. SIAM J. Matrix Anal. Appl., 16(2):520–535, 1995.

Chandrasekaran:1995:SSC

[CI95b] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. On the sensitivity of solution components in linear systems of equations. SIAM J. Matrix Anal. Appl., 16(1):93–112, 1995.

Ciarlet:1989:INL

[Cia89] P. G. Ciarlet. Introduction to Numerical Linear Algebra and Optimisation. Cambridge University Press, ????, 1989. ???? pp.

Cipra:1988:CDP

[Cip88] Barry A. Cipra. Computer-drawn pictures stalk the wild trajectory. Science, 241:1162–1163, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Cipra:1989:MDH

[Cip89] Barry Cipra. Misteaks … and How to Find Them Before the Teacher Does . . .. Academic Press, New York, USA, second edition, 1989. ???? pp.

Clint:1970:EEE

[CJ70] M. Clint and A. Jennings. The evaluation of eigenvalues and eigenvectors of real symmetric matrices by simultaneous iteration. The Computer Journal, 13(1): 76–80, 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Carlson:1992:GEE

[CJLP92] David Carlson, Charles R. Johnson, David C. Lay, and A. Duane Porter. Gems of exposition in elementary linear algebra. College Mathematics Journal, 23 (4):299–303, 1992.

Carlson:1993:LAC

[CJLP93] David Carlson, Charles R. Johnson, David C. Lay, and A. Duane Porter. The Linear Algebra Curriculum Study Group recommendations for the first course in linear algebra. College Mathematics Journal, 24(1):41–46, 1993.

Carlson:1957:OPL

[CK57] B. C. Carlson and Joseph M. Keller. Orthogonalization procedures and the localization of Wannier functions. Physical Review, 105:102–103, 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Chen:1975:TPP

[CK75] Shyh-Ching Chen and David J. Kuck.

Time and parallel processor bounds for linear recurrence systems. IEEE Trans. Comput., C-24(7):701–717, 1975.

Campbell-Kelly:1980:PEE

[CK80a] Martin Campbell-Kelly. Programming the EDSAC: Early programming activity at the University of Cambridge. Annals of the History of Computing, 2(1): 7–36, 1980.

Campbell-Kelly:1980:PME

[CK80b] Martin Campbell-Kelly. Programming the Mark I: Early programming activity at the University of Manchester. Annals of the History of Computing, 2(2): 130–168, 1980.

Campbell-Kelly:1981:PPA

[CK81] Martin Campbell-Kelly. Programming the Pilot ACE: Early programming activity at the National Physical Laboratory. Annals of the History of Computing, 3(2):133–162, 1981.

Campbell-Kelly:1984:RAT

[CK84] Martin Campbell-Kelly. Review of “Alan Turing: The Enigma”, by Andrew Hodges. Annals of the History of Computing, 6(2):176–178, 1984.

Campbell-Kelly:1989:IBT

[CK89] Martin Campbell-Kelly. ICL: A Business and Technical History. Oxford University Press, ????, 1989. ISBN 0-19-853918-5. xii + 409 pp.

Coomes:1995:RCS

[CKP95] Brian A. Coomes, Hüseyin Koçak, and Kenneth J. Palmer. Rigorous computational shadowing of orbits of ordinary differential equations. Numer. Math., 69 (4):401–421, 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1978:PPB

[CKS78] S. C. Chen, D. J. Kuck, and A. H. Sameh. Practical parallel band triangular system solvers. ACM Transactions on Mathematical Software, 4(3):270–277, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Campbell-Kelly:1985:MSL

[CKW85] Martin Campbell-Kelly and Michael R. Williams. The Moore School Lectures: Theory and Techniques for Design of Electronic Digital Computers, volume 9 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-03109-4. xxxiv + 568 pp.

Cart:1972:EOT

[CL72] D. G. Cart and D. H. Lackowski. Estimation of orthogonal transformations in strapdown inertial systems. IEEE Trans. Automat. Control., AC-??:97–100, February 1972.

Cross:1974:SRC

[CL74] G. W. Cross and P. Lancaster. Square roots of complex matrices. Linear and Multilinear Algebra, 1:289–293, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Chan:1979:CSD

[CL79] Tony F. Chan and John Gregg Lewis.

Computing standard deivations: Accuracy. Comm. ACM, 22(9):526–531, 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ciarlet:1990:HNA

[CL90a] P. G. Ciarlet and J. L. Lions. Handbook of Numerical Analysis, volume I. Elsevier/North Holland, ????, 1990. ???? pp.

Coleman:1990:LSN

[CL90b] Thomas F. Coleman and Yuying Li.

Large-Scale Numerical Optimization. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-268-8. ???? pp.

Codenotti:1991:PCL

[CL91] Bruno Codenotti and Mauro Leoncini.

Parallel Complexity of Linear System Solution. World Scientific, Singapore, 1991. ISBN 981-02-0503-1. ix + 218 pp.

Coleman:1992:GQC

[CL92] Thomas F. Coleman and Yuying Li.

A global and quadratically convergent method for linear L_{∞} problems. SIAM Journal on Numerical Analysis, 29(4): 1166–1186, August 1992. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clark:1992:PTP

[Cla92] Malcolm Clark. A Plain TEX Primer.

Oxford University Press, ????, 1992. ISBN 0-19-853724-7. ???? pp.

Clenshaw:1955:NSC

[Cle55] C. W. Clenshaw. A note on the summation of Chebyshev series. M.T.A.C., 9 (51):118–120, 1955.

Clement:1959:CTD

[Cle59] Paul A. Clement. A class of triplediagonal matrices for test purposes. SIAM Review, 1(1):50–52, 1959.

Clenshaw:1978:NLA

[Cle78] C. W. Clenshaw. A new look at error analysis. In Anonymous [Ano78a], pages 54–75.

Cliff:1966:ORC

[Cli66] Norman Cliff. Orthogonal rotation to congruence. Psychometrika, 31(1):33–42, 1966. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Cline:1972:RCL

[Cli72] A. K. Cline. Rate of convergence of Lawson’s algorithm. Math. Comp., 26: 167–176, 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cline:1973:EMS

[Cli73] A. K. Cline. An elimination method for the solution of linear least squares problems. SIAM J. Numer. Anal., 10(2): 283–289, 1973.

Clinger:1990:HRF

[Cli90] William D. Clinger. How to read floating point numbers accurately. ACM SIG-PLAN Notices, 25(6):92–101, 1990. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cormen:1990:IA

[CLR90] Thomas H. Cormen, Charles E. Leiserson, and Ronald L. Rivest. Introduction to Algorithms. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-03141-8. xvii + 1028 pp.

Chazan:1969:CR

[CM69] D. Chazan and W. Miranker. Chaotic relaxation. Linear Algebra and Appl., 2: 199–22, 1969.

Campbell:1979:GIL

[CM79] S. L. Campbell and C. D. Meyer, Jr.

Generalized Inverses of Linear Transformations. Pitman Publishing Ltd., London, UK, 1979. ISBN 0-486-66693-X. xi + 272 pp. Reprinted by Dover, New York, 1991.

Crawford:1983:FPD

[CM83] C. R. Crawford and Y. S. Moon. Finding a positive definite linear combination of two Hermitian matrices. Linear Algebra and Appl., 51:37–48, 1983.

Concus:1986:CIB

[CM86] P. Concus and G. Meurant. On computing INV block preconditionings for the conjugate gradient method. Technical report, Technical Report LBL-20939, Lawrence Berkeley Laboratory, University of California, Berkeley, ????, 1986. ???? pp.

Charlton:1987:WHC

[CM87] James Charlton and Lisbeth Mark, editors. The Writer’s Home Companion. Penguin, ????, 1987. ???? pp.

Cox:1989:BEA

[CM89] C. L. Cox and W. F. Moss. Backward error analysis for a pole assignment algorithm. SIAM J. Matrix Anal. Appl., 10: 446–456, 1989.

Chen:1993:SMT

[CM93] Denise Chen and Cleve Moler. Symbolic Math Toolbox: User’s Guide. The MathWorks, Inc., Natick, MA, USA, 1993. ???? pp.

Croz:1990:FBM

[CMR90] J. J. Du Croz, P. Mayes, and G. Radicati. Factorizations of band matrices using level 3 BLAS. Technical report, LAPACK Working Note #21, Report CS-90-109, Department of Computer Science, University of Tennessee, Knoxville, ????, 1990. ???? pp.

Cline:1979:ECN

[CMSW79] A. K. Cline, C. B. Moler, G. W. Stewart, and J. H. Wilkinson. An estimate for the condition number of a matrix. SIAM J. Numer. Anal., 16(2):368– 375, 1979.

Chan:1994:CPT

[CNP94] Raymond H. Chan, James G. Nagy, and Robert J. Plemmons. Circulant preconditioned Toeplitz least squares iterations. SIAM J. Matrix Anal. Appl., 15 (1):80–97, 1994.

Croz:1981:SLF

[CNRT81] J. J. Du Croz, S. M. Nugent, J. K. Reid, and D. B. Taylor. Solving large full sets of linear equations in a paged virtual store. ACM Transactions on Mathematical Software, 7(4):527–536, 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Clenshaw:1984:BFP

[CO84] C. W. Clenshaw and F. W. J. Olver. Beyond floating point. J. Assoc. Comput. Mach., 31(2):319–328, 1984.

Cody:1973:SDN

[Cod73] William J. Cody, Jr. Static and dynamic numerical characteristics of floating-point arithmetic. IEEE Trans. Comput., C-22(6):598–601, 1973.

Cody:1981:APF

[Cod81] W. J. Cody. Analysis of proposals for the floating-point standard. Computer, 14:63–68, 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cody:1982:ITF

[Cod82] W. J. Cody. Implementation and testing of function software. In Paul C. Messina and Almerico Murli, editors, Problems and Methodologies in Mathematical Software Production, volume 142 of Lecture Notes in Computer Science, pages 24–47. Springer-Verlag, Berlin, 1982.

Cody:1987:SPS

[Cod87] W. J. Cody. SPECFUN: A portable special function package. In Arthur Wouk, editor, New Computing Environments: Microcomputers in Large-Scale Computing, pages 1–12. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987.

Cody:1988:AMS

[Cod88a] W. J. Cody. ALGORITHM 665 MACHAR: A subroutine to dynamically determine machine parameters. ACM Transactions on Mathematical Software, 14(4):303–311, 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1988:FPS

[Cod88b] W. J. Cody. Floating-point standards—Theory and practice. In Ramon E. Moore, editor, Reliability in Computing: The Role of Interval Methods in Scientific Computing, pages 99–107. Academic Press, Boston, MA, USA, 1988.

Cody:1993:ACP

[Cod93] W. J. Cody. Algorithm 714. CELE-FUNT: A portable test package for complex elementary functions. ACM Transactions on Mathematical Software, 19(1): 121, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cohen:1974:NPS

[Coh74a] A. M. Cohen. A note on pivot size in Gaussian elimination. Linear Algebra and Appl., 8:361–368, 1974.

Cohen:1974:FPM

[Coh74b] A. M. Cohen. On fractional powers of a matrix. 1974.

Cohen:1975:IPM

[Coh75] A. M. Cohen. The inverse of a Pascal matrix. Mathematical Gazette, 59(408): 111–112, 1975. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cohen:1994:PP

[Coh94] A. M. Cohen. Is the polynomial so perfidious? Numer. Math., 68:225–238, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Coker:1975:JI

[Cok75] Jerry Coker. The Jazz Idiom. Prentice-Hall, Englewood Cliffs, NJ, USA, 1975. ???? pp.

Colaianne:1977:AMA

[Col77] A. J. Colaianne. The aims and methods of annotated bibliography. Scholarly Publishing, 11(4):321–331, 1977. ISSN 0036-634X.

Coleman:1984:LSN

[Col84] T. F. Coleman. Large Sparse Numerical Optimization, volume 165 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1984. ???? pp.

Collins:1987:CCE

[Col87] Collins. Collins Cobuild English Language Dictionary. ????, London, 1987. ISBN 0-00-375021-3. ???? pp.

Coleman:1991:LSN

[Col91] T. F. Coleman. Large-scale numerical optimization: Introduction and overview. Technical Report CTC 91 TR85, Advanced Computing Research Institute, Cornell Theory Center, Cornell University, Ithaca, New York, 1991. ???? pp. To appear in the Encyclopedia of Computer Science and Technology, Marcel Dekker.

Company:1983:PPG

[Com83] International Paper Company. Pocket Pal: A Graphic Arts Production Handbook. International Paper Company, New York, thirteenth edition, 1983. ???? pp.

Computer:1988:ANM

[Com88] Apple Computer. Apple Numerics Manual. Addison-Wesley, Reading, MA, USA, second edition, 1988. ???? pp.

Computer:1994:IMP

[Com94] Apple Computer. Inside Macintosh: PowerPC Numerics. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-40728-0. ???? pp.

Cooper:1964:WTR

[Coo64] Bruce M. Cooper. Writing Technical Reports. Penguin, London, 1964. ISBN 0-14-020676-0. ???? pp. Reprinted 1986.

Cooper:1967:ECC

[Coo67] G. J. Cooper. The evaluation of the coefficients in a Chebyshev expansion. The Computer Journal, 10(1):94– 100, 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Coonen:1981:UDN

[Coo81] Jerome T. Coonen. Underflow and the denormalized numbers. Computer, 14:75–87, 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cooley:1990:HFG

[Coo90] James W. Cooley. How the FFT gained acceptance. In Stephen G. Nash, editor, A History of Scientific Computing, pages 133–140. Addison-Wesley, Reading, MA, USA, 1990.

Cooley:1994:LFD

[Coo94] James W. Cooley. Lanczos and the FFT: A discovery before its time. In J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors, Proceedings of the Cornelius Lanczos International Centenary Conference, pages 3–9. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Cornwall:1988:HHI

[Cor88] H. Cornwall. The Hacker’s Handbook III. Century Hutchinson, London, 1988. ???? pp.

Corless:1992:CFC

[Cor92a] Rob Corless. Continued fractions and chaos. Amer. Math. Monthly, 99(3): 203–215, 1992.

Corless:1992:DCN

[Cor92b] Robert M. Corless. Defectcontrolled numerical methods and shadowing for chaotic differential equations. Physica D, 60:323–334, 1992. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Corey:1994:MTT

[Cor94] Jim Corey. Mark Twain–Technical writer. IEEE Trans. Prof. Commun., 37 (2):108, 1994.

Cottle:1974:MSC

[Cot74] Richard W. Cottle. Manifestations of the Schur complement. Linear Algebra and Appl., 8:189–211, 1974.

Clenshaw:1989:LIA

[COT89] C. W. Clenshaw, F. W. J. Olver, and P. R. Turner. Level-index arithmetic: An introductory survey. In Peter R. Turner, editor, Numerical Analysis and Parallel Processing, Lancaster 1987, volume 1397 of Lecture Notes in Mathematics, pages 95–168. Springer-Verlag, Berlin, 1989. ISBN 3-540-51645-X.

Cowell:1977:PNS

[Cow77] Wayne R. Cowell, editor. Portability of Numerical Software, volume 57 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ???? pp.

Cowell:1984:SDM

[Cow84] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp.

Cowell:1989:DLT

[Cow89a] W. R. Cowell. DO loop transforming tools. In Pollicini [Pol89], pages 161– 180.

Cowell:1989:TW

[Cow89b] Wayne R. Cowell. Tool writing. In Pollicini [Pol89], pages 251–275.

Cox:1972:NES

[Cox72] M. G. Cox. The numerical evaluation of B-splines. J. Inst. Maths Applics, 10: 134–149, 1972.

Cox:1975:ASI

[Cox75] M. G. Cox. An algorithm for spline interpolation. J. Inst. Maths Applics, 15: 95–108, 1975.

Cox:1978:NES

[Cox78] M. G. Cox. The numerical evaluation of a spline from its B-spline representation. J. Inst. Maths Applics, 21:135–143, 1978.

Cox:1986:PCS

[Cox86] M. G. Cox. Piecewise Chebyshev series. IMA Bulletin, 22(11/12):163–166, 1986.

Cline:1976:SUL

[CP76] R. E. Cline and R. J. Plemmons. l_{2}- solutions to underdetermined linear systems. SIAM Review, 18(1):92–106, 1976.

Chang:1995:NPB

[CP95] Xiao-Wen Chang and C. C. Paige.

New perturbation bounds for the Cholesky factorization. Manuscript, February 1995.

Courant:1941:WM

[CR41] R. Courant and H. Robbins. What Is Mathematics? Oxford University Press, ????, 1941. ???? pp.

Curtis:1972:ASM

[CR72] A. R. Curtis and J. K. Reid. On the automatic scaling of matrices for Gaussian elimination. J. Inst. Maths Applics, 10:118–124, 1972.

Cohen:1976:ISF

[CR76] Jacques Cohen and Martin Roth. On the implementation of Strassen’s fast multiplication algorithm. Acta Informatica, 6:341–355, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cope:1979:BSL

[CR79] J. E. Cope and B. W. Rust. Bounds on solutions of linear systems with inaccurate data. SIAM J. Numer. Anal., 16 (6):950–963, 1979.

Cline:1983:SCE

[CR83] A. K. Cline and R. K. Rew. A set of counter-examples to three condition number estimators. SIAM J. Sci. Stat. Comput., 4(4):602–611, 1983.

Chan:1986:CNS

[CR86] Tony F. Chan and Diana C. Resasco.

On the condition of nearly singular matrices under rank-1 perturbations. Linear Algebra and Appl., 76:223–232, 1986.

Carroll:1988:TWR

[CR88] R. J. Carroll and D. Ruppert. Transformation and Weighting in Regression. Chapman and Hall, Ltd., London, UK, 1988. ???? pp.

Cameron:1991:LGE

[CR91] Debra Cameron and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-84-6. xxvii + 411 pp. LCCN QA76.76.T49 C35

1991. URL http://www.oreilly.com/
catalog/9780937175842.

Calvetti:1992:CVS

[CR92] D. Calvetti and L. Reichel. A Chebychev-Vandermonde solver. Linear Algebra and Appl., 172:219–229, 1992.

Calvetti:1993:FIV

[CR93] D. Calvetti and L. Reichel. Fast inversion of Vandermonde-like matrices involving orthogonal polynomials. BIT, 33: 473–484, 1993. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Choi:1994:SCD

[CR94] Man-Duen Choi and Peter Rosenthal. A survey of Chandler Davis. Linear Algebra and Appl., 208/209:3–18, 1994.

Cramer:1974:BSO

[Cra74] Elliot M. Cramer. On Browne’s solution for oblique Procrustes rotation. Psychometrika, 39(2):159–163, 1974. CO-DEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Crary:1979:VPN

[Cra79] Fred D. Crary. A versatile precompiler for nonstandard arithmetics. ACM Transactions on Mathematical Software, 5(2):204–217, 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Crawford:1986:APR

[Cra86] C. R. Crawford. ALGORITHM 646 PDFIND: A routine to find a positive definite linear combination of two real symmetric matrices. ACM Transactions on Mathematical Software, 12:278–282, 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Inc:1989:UMS

[Cra89] Cray Research, Inc., Minneapolis, MN, USA. UNICOS Math and Scientific Library Reference Manual, 1989. ???? pp.

Crick:1974:DHP

[Cri74] Francis Crick. The double helix:

A personal view. Nature, 248:766–769, 1974. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Crichton:1975:MOS

[Cri75] Michael Crichton. Medical obfuscation: Structure and function. The New England Journal of Medicine, 293(24): 1257–1259, 1975. CODEN NEJMAG. ISSN 0028-4793 (print), 1533-4406 (electronic).

Crick:1989:WMP

[Cri89] Francis Crick. What Mad Pursuit:

A Personal View of Scientific Discovery. Penguin, ????, 1989. ISBN 0-14-011973-6. ???? pp.

Crout:1941:SME

[Cro41] Prescott D. Crout. A short method for evaluating determinants and solving systems of linear equations with real or complex coefficients. Trans. Amer. Inst. Elec. Eng., 60:1235–1241, 1941. CODEN TAEEA5. ISSN 0096-3860.

Croarken:1990:ESC

[Cro90] Mary Croarken. Early Scientific Computing in Britain. Clarendon Press, New York, NY, USA, 1990. ISBN 0-19-853748-4. viii + 160 pp.

Cryer:1968:PSG

[Cry68] C. W. Cryer. Pivot size in Gaussian elimination. Numer. Math., 12:335–345, 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cryer:1973:FTP

[Cry73] Colin W. Cryer. The LU-factorization of totally positive matrices. Linear Algebra and Appl., 7:83–92, 1973.

Cryer:1976:SPT

[Cry76] Colin W. Cryer. Some properties of totally positive matrices. Linear Algebra and Appl., 15:1–25, 1976.

Crystal:1987:CEL

[Cry87] David Crystal. The Cambridge Encyclopedia of Language. Cambridge University Press, ????, 1987. ISBN 0-521-42443-7. ???? pp.

Crystal:1988:EL

[Cry88] David Crystal. The English Language. Penguin, London, 1988. ISBN 0-14-022730-X. ???? pp.

Coleman:1984:NCO

[CS84] Thomas F. Coleman and Danny C. Sorensen. A note on the computation of an orthonormal basis for the null space of a matrix. Math. Prog., 29:234–242, 1984. CODEN MHPGA4. ISSN 0025-5610.

Cao:1986:NIH

[CS86] Wei-Lu Cao and William J. Stewart.

A note on inverses of Hessenberg-like matrices. Linear Algebra and Appl., 76:233– 240, 1986.

Coschi:1987:WUG

[CS87] G. Coschi and J. B. Schueler.

WATFOR-77 user’s guide for the IBM PC with DOS. Technical report, WATCOM Publications, Waterloo, Ontario, Canada, 1987. ???? pp.

Chan:1989:TEC

[CS89a] R. H. Chan and G. Strang. Toeplitz equations by conjugate gradients with circulant preconditioner. SIAM J. Sci. Stat. Comput., 10:104–119, 1989.

Chen:1989:PBP

[CS89b] Chun-hui Chen and Ji-guang Sun.

Perturbation bounds for the polar factors. J. Comput. Math., 7:397–401, 1989. CO-DEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

Csanky:1976:FPM

[Csa76] L. Csanky. Fast parallel matrix inversion algorithms. SIAM J. Comput., 5(4): 618–623, 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chui:1983:ATI

[CSW83] C. K. Chui, L. L. Schumaker, and J. D. Ward, editors. Approximation Theory IV. Academic Press, New York, 1983. ???? pp.

Cooley:1965:AMC

[CT65] James W. Cooley and John W. Tukey.

An algorithm for the machine calculation of complex Fourier series. Math. Comp., 19(90):297–301, 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooley:1993:OPF

[CT93] James W. Cooley and John W. Tukey. On the origin and publication of the FFT paper. Current Contents: Physical, Chemical & Earth Sciences, (51–52):8–9, 1993.

Culver:1966:EUR

[Cul66] W. J. Culver. On the existence and uniqueness of the real logarithm of a matrix. Proc. Amer. Math. Soc., 17:1146– 1151, 1966.

Cullen:1985:LMB

[Cul85] M. R. Cullen. Linear Modes in Biology. Ellis Horwood, New York, NY, USA, 1985. ISBN 0-470-20206-8. ???? pp.

Cuppen:1984:UTP

[Cup84] J. J. M. Cuppen. On updating triangular products of Householder reflections. Numer. Math., 45:403–409, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cybenko:1986:CME

[CV86] George Cybenko and Charles F. Van Loan. Computing the minimum eigenvalue of a symmetric positive definite Toeplitz matrix. SIAM J. Sci. Stat. Comput., 7(1):123–131, 1986.

Coleman:1988:HMC

[CV88] Thomas F. Coleman and Charles F. Van Loan. Handbook for Matrix Computations, volume 4 of Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1988. ISBN 0-89871-227-0 (paperback), 1-61197-104-7 (e-book). vii + 264 pp. LCCN QA188 .C651 1988. US$34.00.

Charlier:1989:SAR

[CV89] J.-P. Charlier and P. Van Dooren. A systolic algorithm for Riccati and Lyapunov equations. Math. Control, Signals and Sys., 2:109–136, 1989.

Cody:1980:SME

[CW80] William J. Cody, Jr. and William Waite. Software Manual for the Elementary Functions. Prentice-Hall, Englewood Cliffs, NJ, USA, 1980. ISBN 0-13-822064-6. x + 269 pp.

Coppersmith:1982:ACM

[CW82] D. Coppersmith and S. Winograd.

On the asymptotic complexity of matrix multiplication. SIAM J. Comput., 11(3): 472–492, 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cullum:1985:LAL

[CW85] Jane K. Cullum and Ralph A.

Willoughby. Lanczos Algorithms for Large Symmetric Eigenvalue Computations, Volume I Theory. Birkhäuser, Boston, MA, USA, 1985. ISBN 0-8176-3058-9. xiv + 273 pp.

Cullum:1986:LSE

[CW86] J. Cullum and R. A. Willoughby.

Large Scale Eigenvalue Problems. Elsevier Science Publishers B.V., North-Holland, Amsterdam, 1986. ???? pp.

Coppersmith:1987:MMA

[CW87] Don Coppersmith and Shmuel Winograd. Matrix multiplication via arithmetic progressions. In Proceedings of the Nineteenth Annual ACM Symposium on Theory of Computing, pages 1–6. ACM Press, New York, NY, USA, 1987.

Cybenko:1978:EAS

[Cyb78] G. Cybenko. Error analysis of some signal processing algorithms. Ph.d. thesis, Princeton University, Princeton, NJ, USA, 1978. ???? pp.

Cybenko:1979:REP

[Cyb79] G. Cybenko. Round-off error propagation in Durbin’s, Levinson’s, and Trench’s algorithms. Technical report, Proc. International Conference on Acoustics, Speech and Signal Processing, Washington DC, 498-501, ????, 1979. ???? pp.

Cybenko:1980:NSL

[Cyb80] G. Cybenko. The numerical stability of the Levinson–Durbin algorithm for Toeplitz systems of equations. SIAM J. Sci. Stat. Comput., 1:303–319, 1980.

Cybenko:1981:EAA

[Cyb81] G. Cybenko. Error analysis of algorithms for Toeplitz matrices and linear prediction. Technical report, in Circuit Theory and Design, R. Boite and P. Dewilde, eds, North-Holland, 379-386, ????, 1981. ???? pp.

Cybenko:1983:GOT

[Cyb83] George Cybenko. A general orthogonalization technique with applications to time series analysis and signal processing. Math. Comp., 40(161):323–336, 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cybenko:1984:NSL

[Cyb84a] G. Cybenko. The numerical stability of the lattice algorithm for least squares linear prediction problems. BIT, 24:441– 455, 1984. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cybenko:1984:ETM

[Cyb84b] G. Cybenko. On the eigenstructure of Toeplitz matrices. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-32:918–921, 1984. CODEN IETABA. ISSN 0096-3518.

Cybenko:1986:CEM

[Cyb86] G. Cybenko. The conditioning of eigenvector methods for beamforming problems. 1986.

Daintith:1993:BT

[D^{+}93] John Daintith et al., editors. Bloomsbury Thesaurus. Bloomsbury, London, 1993. ISBN 0-7475-1226-4. xxx + 1569 pp. LCCN PE1591.B48 1993.

Dahl:1978:SLA

[Dah78] Grete Dahl. On scaling in linear algebraic systems. BIT, 18:363–365, 1978. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Dahlquist:1983:MMM

[Dah83] G. Dahlquist. On matrix majorants and minorants, with applications to differential equations. Linear Algebra and Appl., 52/53:199–216, 1983.

Duren:1988:CMA

[DAM88] Peter Duren, Richard A. Askey, and Uta C. Merzbach, editors. A Century of Mathematics in America, Part I. American Mathematical Society, Providence, RI, USA, 1988. ISBN 0-8218-0124-4. ???? pp.

Danloy:1976:CSH

[Dan76] B. Danloy. On the choice of signs for Householder’s matrices. J. Comp. Appl. Math., 2(1):67–69, 1976.

Dantzig:1984:RAO

[Dan84] George B. Dantzig. Reminiscences about the origins of linear programming. Memoirs of the American Mathematical Society, 48(298):1–11, 1984. CO-DEN MAMCAU. ISSN 0065-9266 (print), 1947-6221 (electronic).

Dantzig:1987:OSM

[Dan87] G. B. Dantzig. Origins of the simplex method. Technical report, Technical Report SOL 87-5, Department of Operations Research, Stanford University, ????, 1987. ???? pp.

Datta:1985:LAR

[Dat85] B. N. Datta, editor. Linear Algebra and its Role in Systems Theory, volume 47 of Contemporary Math. American Mathematical Society, Providence, RI, USA, 1985.

Datta:1987:AAE

[Dat87] B. N. Datta. An algorithm to assign eigenvalues in a Hessenberg matrix: Single input case. IEEE Trans. Automat. Control., AC-32(5):414–417, 1987.

Datta:1988:MEAa

[Dat88] Karabi Datta. The matrix equation XA − BX = R and its applications. Linear Algebra and Appl., 109:91–105, 1988.

Datta:1989:PLS

[Dat89] B. N. Datta. Parallel and large-scale matrix computations in control: Some ideas. Linear Algebra and Appl., 121:243– 264, 1989.

Davis:1972:FMD

[Dav72] P. J. Davis. Fidelity in mathematical discourse: Is one and one really two? Amer. Math. Monthly, 79:252–263, 1972.

Davis:1973:EFC

[Dav73] C. Davis. Explicit functional calculus. Linear Algebra and Appl., 6:193–199, 1973.

Davies:1975:VPP

[Dav75] Andrew Davies. A Very Peculiar Practice. Coronet (Hodder and Stoughton), London, 1975. ???? pp.

Davis:1979:CM

[Dav79] Philip J. Davis. Circulant Matrices.

Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1979. ISBN 0-471-05771-1. xv + 250 pp.

Davis:1981:NSQ

[Dav81] G. J. Davis. Numerical solution of a quadratic matrix equation. SIAM J. Sci. Stat. Comput., 2:164–175, 1981.

Davis:1983:AAC

[Dav83] G. J. Davis. Algorithm 598: An algorithm to compute solvents of the matrix equation AX^{2}+BX+C = 0. ACM Transactions on Mathematical Software, 9:246– 254, 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davies:1985:LCF

[Dav85] E. B. Davies. Lipschitz continuity of functions of operators in the Schatten classes. J. London Math. Soc., 37(2):148– 157, 1985.

Davenport:1986:SNM

[Dav86] J. H. Davenport. Symbolic and numeric manipulation of integrals. In Willard L. Miranker and Richard A. Toupin, editors, Accurate Scientific Computations, Proceedings, 1985, volume 235 of Lecture Notes in Computer Science, pages 168–180. Springer-Verlag, Berlin, 1986.

Davies:1988:ECO

[Dav88] Louise Davies. Easy Cooking for One or Two. Penguin, ????, 1988. ???? pp.

Davis:1992:WMT

[Dav92] Chandler Davis. Which is to be master, that’s all. The Mathematical Intelligencer, 14(2):51, 1992. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Davis:1993:WMIa

[Dav93a] Chandler Davis. Which is to be master—III. The Mathematical Intelligencer, 15(2):26, 1993. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Davis:1993:STC

[Dav93b] Philip J. Davis. Spirals from Theodorus to Chaos. A. K. Peters, Ltd., Wellesley, MA, USA, 1993. ISBN 1-56881-010-5. ???? pp. With contributions by Walter Gautschi and Arieh Iserles.

Dax:1982:PPS

[Dax82] Achiya Dax. Partial pivoting strategies for symmetric Gaussian elimination. Math. Prog., 22:288–303, 1982. CODEN MHPGA4. ISSN 0025-5610.

Dax:1990:CLS

[Dax90a] Achiya Dax. The convergence of linear stationary iterative processes for solving singular unstructured systems of linear equations. SIAM Review, 32(4):611– 635, 1990.

Dax:1990:LSA

[Dax90b] Achiya Dax. Line search acceleration of iterative methods. Linear Algebra and Appl., 130:43–63, 1990.

Dax:1990:HAL

[Dax90c] Achiya Dax. On hybrid acceleration of a linear stationary iterative process. Linear Algebra and Appl., 130:99– 110, 1990.

Day:1977:HWS

[Day77] Robert A. Day. How to write a scientific paper. IEEE Trans. Prof. Commun., PC-20(1):32–37, 1977.

Day:1983:TMM

[Day83a] R. A. Day. Typesetting mathematics on multi-access systems. Software —Practice and Experience, 13:131–138, 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Day:1983:HWP

[Day83b] Robert A. Day. How To Write and Publish a Scientific Paper. ISI Press, Philadelphia, second edition, 1983. ???? pp. First edition, 1979.

Day:1988:HWP

[Day88] Robert A. Day. How To Write and Publish a Scientific Paper. Cambridge University Press and Oryx Press (Phoenix, Arizona), ????, third edition, 1988. ISBN 0-521-36760-3. xi + 211 pp.

Day:1992:SEG

[Day92] Robert A. Day. Scientific English: A Guide for Scientists and Other Professionals. Oryx Press, Phoenix, Arizona, 1992. ISBN 0-89774-722-4. ???? pp.

Boor:1972:CS

[dB72] Carl de Boor. On calculating with Bsplines. J. Approximation Theory, 6:50– 62, 1972.

Denman:1976:MSF

[DB76] Eugene D. Denman and Alex N. Beavers, Jr. The matrix sign function and computations in systems. Applied Mathematics and Computation, 2:63–94, 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dahlquist:1974:NM

[DBA74] Germund Dahlquist, Åke Björck, and Ned Anderson. Numerical Methods. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1974. ISBN 0-13-627315-7. xviii + 573 pp. LCCN QA297 .D131 1969. Translated by Ned Anderson.

Boor:1986:CLS

[dBK86] C. de Boor and H.-O. Kreiss. On the condition of the linear systems associated with discretized BVPs of ODEs. SIAM J. Numer. Anal., 23(5):936–939, 1986.

Dongarra:1990:TAD

[DBKF90] J. J. Dongarra, O. Brewer, J. A. Kohl, and S. Finberg. A tool to aid in the design, implementation, and understanding of matrix algorithms for parallel processors. Journal of Parallel and Distributed Computing, 9:185–202, 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dongarra:1979:LUG

[DBMS79] J. J. Dongarra, J. R. Bunch, C. B.

Moler, and G. W. Stewart. LINPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X. ???? pp.

Boor:1977:BEA

[dBP77] Carl de Boor and Allan Pinkus.

Backward error analysis for totally positive linear systems. Numer. Math., 27:485–490, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dent:1977:MCA

[DC77] Warren T. Dent and David C. Cavander. More on computational accuracy in regression. J. Amer. Statist. Assoc., 72 (359):598–600, 1977.

Dressel:1987:ABE

[DC87] Susan Dressel and Joe Chew. Authenticity beats eloquence. IEEE Trans. Prof. Commun., PC-30:82–83, 1987.

Dongarra:19xx:SLB

[DCDH90a] J. J. Dongarra, J. J. Du Croz, I. S. Duff, and S. J. Hammarling. A set of Level 3 basic linear algebra subprograms. ACM Transactions on Mathematical Software, 16:1–17, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:19xx:ASL

[DCDH90b] J. J. Dongarra, J. J. Du Croz, I. S. Duff, and S. J. Hammarling. Algorithm 679. A set of Level 3 basic linear algebra subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 16:18– 28, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1995:PFI

[DCH^{+}95] Jack Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, Jerzy Waśniewski, and Adam Zemla. A proposal for a Fortran 90 interface for LA-PACK. Technical Report CS-95-301, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, July 1995. ???? pp. LAPACK Working Note 101.

Dongarra:19xx:ESF

[DCHH88a] Jack J. Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, and Richard J. Hanson. An extended set of Fortran basic linear algebra subprograms. ACM Transactions on Mathematical Software, 14(1):1–17, 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:19xx:AES

[DCHH88b] Jack J. Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, and Richard J. Hanson. Algorithm 656. An extended set of Fortran basic linear algebra subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 14(1):18–32, 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Datardina:1990:PSB

[DCHP90] S. P. Datardina, J. J. Du Croz, S. J. Hammarling, and M. W. Pont. A proposed specification of BLAS routines in C. Technical Report TR6/90, NAG Ltd., Oxford, December 1990. ???? pp.

Demmel:1988:GDS

[DCHS88] J. W. Demmel, J. J. Du Croz, S. J.

Hammarling, and D. C. Sorensen. Guidelines for the design of symmetric eigenroutines, SVD, and iterative refinement and condition estimation for linear systems. Technical report, LAPACK Working Note #4, Technical Memorandum 111, Mathematics and Computer Science Division, Argonne National Laboratory, ????, 1988. ???? pp.

Datta:1986:TCA

[DD86] B. N. Datta and K. Datta. Theoretical and computational aspects of some linear algebra problems in control theory. In C. I. Byrnes and A. Lindquist, editors, Computational and Combinatorial Methods in Systems Theory, pages 201–212. Elsevier North Holland, Inc., New York, NY, USA, 1986.

Datta:1988:MEAb

[DD88] B. N. Datta and K. Datta. The matrix equation XA = A^{T} X and an associated algorithm for solving the inertia and stability problems. Linear Algebra and Appl., 97:103–119, 1988.

Dayde:1989:ULB

[DD89] M. J. Daydé and I. S. Duff. Use of level 3 BLAS in LU factorization on the CRAY-2, the ETA-10P, and the IBM 3090/VF. Int. J. Supercomputer Appl., 3:40–70, 1989. CODEN IJSAE9. ISSN 0890-2720.

Dayde:1990:ULB

[DD90] M. J. Daydé and I. S. Duff. Use of level 3 BLAS in LU factorization in a multiprocessing environment on three vector multiprocessors: the ALLIANT FX/ 80. Technical report, the CRAY-2, and the IBM 3090 VF Report RAL-90-083, Rutherford Appleton Laboratory, Didcot, Oxon, ????, 1990. ???? pp.

Demmel:1987:PDL

[DDC^{+}87] J. W. Demmel, J. J. Dongarra, J. J. Du Croz, A. Greenbaum, S. J. Hammarling, and D. C. Sorensen. Prospectus for the development of a linear algebra library for high-performance computers. Technical Memorandum No. 97, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, September 1987. ???? pp.

Dongarra:1989:VPC

[DDGM89] J. J. Dongarra, I. S. Duff, P. W. Gaffney, and S. McKee. Vector and Parallel Computing. Ellis Horwood, New York, NY, USA, 1989. ???? pp.

Demmel:1992:DPH

[DDK92] James W. Demmel, J. J. Dongarra, and W. Kahan. On designing portable high performance numerical libraries. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 69–84. Longman Scientific and Technical, Harlow, Essex, UK, 1992.

Deift:1991:BSV

[DDLT91] Percy Deift, James W. Demmel, Luen-Chau Li, and Carlos Tomei. The bidiagonal singular value decomposition and Hamiltonian mechanics. SIAM J. Numer. Anal., 28:1463–1516, 1991.

Demmel:1994:CPB

[DDR94] James W. Demmel, Inderjit Dhillon, and Huan Ren. On the correctness of parallel bisection in floating point. Technical Report CS-94-228, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, March 1994. 38 pp. LAPACK Working Note 70.

Dongarra:1991:SLS

[DDSv91] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Solving Linear Systems on Vector and Shared Memory Computers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-270-X. x + 256 pp. LCCN QA184 .S65 1991.

Dongarra:1984:SMA

[DE84] J. J. Dongarra and S. C. Eisenstat.

Squeezing the most out of an algorithm in CRAY FORTRAN. ACM Transactions on Mathematical Software, 10:219– 230, 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dahlquist:1972:BEL

[DEG72] G. Dahlquist, S. C. Eisenstat, and G. H. Golub. Bounds for the error of linear systems of equations using the theory of moments. Journal of Mathematical Analysis and Applications, 37:151–166, 1972. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

Duff:1988:SSG

[DEGR88] I. S. Duff, A. M. Erisman, C. W.

Gear, and J. K. Reid. Sparsity structure and Gaussian elimination. ACM SIGNUM Newsletter, 23(2):2–8, 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Deif:1986:SAL

[Dei86] A. Deif. Sensitivity Analysis in Linear Systems. Springer-Verlag, Berlin, 1986. ???? pp.

Deif:1990:RPP

[Dei90] A. S. Deif. Realistic a priori and a posteriori error bounds for computed eigenvalues. IMA J. Numer. Anal., 9:323–329, 1990.

Dekker:1969:FZM

[Dek69] T. J. Dekker. Finding a zero by means of successive linear interpolation. In Bruno Dejon and Peter Henrici, editors, Constructive Aspects of the Fundamental Theorem of Algebra, pages 37–48. Wiley-Interscience, New York, NY, USA, 1969.

Dekker:1971:FPT

[Dek71] T. J. Dekker. A floating-point technique for extending the available precision. Numer. Math., 18:224–242, 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dekker:1982:PCM

[Dek82] T. J. Dekker. Program correctness and machine arithmetic. In Paul C. Messina and Almerico Murli, editors, Problems and Methodologies in Mathematical Software Production, volume 142 of Lecture Notes in Computer Science, pages 48–80. Springer-Verlag, Berlin, 1982.

Dongarra:1995:RCI

[DEK95] Jack Dongarra, Victor Eijkhout, and Ajay Kalhan. Reverse communication interface for linear algebra templates for iterative methods. Technical Report CS-95-291, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, May 1995. ???? pp. LAPACK Working Note 99.

Dackland:1992:PBM

[DEKV92] Krister Dackland, Erik Elmroth, Bo Kågström, and Charles F. Van Loan. Parallel block matrix factorizations on the shared-memory multiprocessor IBM 3090 VF/600J. Int. J. Supercomputer Appl., 6(1):69–97, 1992. CODEN IJSAE9. ISSN 0890-2720.

Demko:1977:IBM

[Dem77] Stephen Demko. Inverses of band matrices and local convergence of spline projections. SIAM J. Numer. Anal., 14 (4):616–619, 1977.

Demko:1979:BB

[Dem79] Stephen Demko. On bounding ∥A^{−1}∥_{∞} for banded A. Math. Comp., 33 (148):1283–1288, 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demmel:1983:CNE

[Dem83] James W. Demmel. The condition number of equivalence transformations that block diagonalize matrix pencils. SIAM J. Numer. Anal., 20(3):599– 610, 1983.

Demmel:1984:URN

[Dem84] James W. Demmel. Underflow and the reliability of numerical software. SIAM J. Sci. Stat. Comput., 5(4):887– 919, 1984.

Demmel:1987:SPS

[Dem87a] J. W. Demmel. The smallest perturbation of a submatrix which lowers the rank and constrained total least squares problems. SIAM J. Numer. Anal., 24: 199–206, 1987.

Demmel:1987:CTC

[Dem87b] James W. Demmel. A counterexample for two conjectures about stability. IEEE Trans. Automat. Control., AC-32 (4):340–342, 1987.

Demmel:1987:CND

[Dem87c] James W. Demmel. On condition numbers and the distance to the nearest ill-posed problem. Numer. Math., 51:251–289, 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:1987:EAA

[Dem87d] James W. Demmel. On error analysis in arithmetic with varying relative precision. In Mary Jane Irwin and Renato Stefanelli, editors, Proceedings of the Eighth Symposium on Computer Arithmetic, Como, Italy, pages 148–152. IEEE Computer Society, Washington, DC, 1987.

Demmel:1987:TMR

[Dem87e] James W. Demmel. Three methods for refining estimates of invariant subspaces. Computing, 38:43–57, 1987. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Demmel:1988:SSV

[Dem88a] J. W. Demmel. On structured singular values. Technical report, Technical Report No. 349, Computer Science Department, Courant Institute of Mathematical Sciences, New York, ????, 1988. ???? pp.

Demmel:1988:PNA

[Dem88b] J. W. Demmel. The probability that a numerical analysis problem is difficult. Math. Comp., 50:449–480, 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demeure:1989:FFV

[Dem89a] Cédric J. Demeure. Fast QR factorization of Vandermonde matrices. Linear Algebra and Appl., 122/3/4:165–194, 1989.

Demmel:1989:FPE

[Dem89b] James W. Demmel. On floating point errors in Cholesky. Technical Report CS-89-87, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October 1989. 6 pp. LAPACK Working Note 14.

Demeure:1990:FCV

[Dem90a] Cédric J. Demeure. QR factorization of confluent Vandermonde matrices. IEEE Trans. Acoustics, Speech, and Signal Processing, 38(10):1799–1802, 1990. CODEN IETABA. ISSN 0096-3518.

Demmel:1990:NDM

[Dem90b] J. W. Demmel. Nearest defective matrices and the geometry of illconditioning. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 35–55. Oxford University Press, 1990.

Demmel:1991:OIA

[Dem91] James W. Demmel. On the odor of IEEE arithmetic. NA Digest, 91(39): ??, 1991. (Response to a message “IEEE Arithmetic Stinks” in Volume 91, Issue 33). Electronic mail magazine: na.help@

na-net.ornl.gov.

Demmel:1992:III

[Dem92a] J. W. Demmel. The inherent inaccuracy of implicit tridiagonal QR. IMA Preprint Series # 963, Institute for Mathematics and its Applications, University of Minnesota, Minneapolis, MN, USA, April 1992. ???? pp.

Demmel:1992:SRP

[Dem92b] J. W. Demmel. Specifications for robust parallel prefix operations. Technical report, Thinking Machine Corp., Boston, 1992. ???? pp.

Demmel:1992:CDN

[Dem92c] James W. Demmel. The componentwise distance to the nearest singular matrix. SIAM J. Matrix Anal. Appl., 13(1): 10–19, 1992.

Demmel:1992:OPN

[Dem92d] James W. Demmel. Open problems in numerical linear algebra. IMA Preprint Series #961, Institute for Mathematics and its Applications, University of Minnesota, Minneapolis, MN, USA, April 1992. 21 pp. LAPACK Working Note 47.

Demmel:1992:SFP

[Dem92e] James W. Demmel. A specification for floating point parallel prefix. Technical Report CS-92-167, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, July 1992. 8 pp. LAPACK Working Note 49.

Demmel:1993:NLA

[Dem93a] James W. Demmel. Numerical Linear Algebra, volume 1 of Berkeley Lecture Notes in Mathematics. Mathematics Department, University of California at Berkeley, Berkeley, CA, USA, 1993. ???? pp.

Demmel:1993:TPN

[Dem93b] James W. Demmel. Trading off parallelism and numerical stability. In Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors, Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E, pages 49–68. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993.

Denman:1981:RRM

[Den81] Eugene D. Denman. Roots of real matrices. Linear Algebra and Appl., 36:133– 139, 1981.

Dennis:1984:UGN

[Den84] J. E. Dennis, Jr. A user’s guide to nonlinear optimization algorithms. Proc. IEEE, 72:1765–1776, 1984. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Duff:1986:DMS

[DER86] I. S. Duff, A. M. Erisman, and J. K.

Reid. Direct Methods for Sparse Matrices. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-853408-6 (hardcover), 0-19-853421-3 (paperback). xiii + 341 pp. LCCN QA188 .D841 1986. US$37.50.

Descloux:1963:BSN

[Des63a] J. Descloux. Bounds for the spectral norm of functions of matrices. Numer. Math., 15:185–190, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Descloux:1963:NRE

[Des63b] J. Descloux. Note on the roundoff errors in iterative processes. Math.

Comp., 17:18–27, 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deutsch:1983:NAM

[Deu83] F. Deutsch. Von Neumann’s alternating method: The rate of convergence. In C. K. Chui, L. L. Schumaker, and J. D. Ward, editors, Approximation Theory IV, pages 427–434. Academic Press, New York, 1983.

Drapkin:1987:PRC

[DF87] T. Drapkin and R. Forsyth. The Punter’s Revenge: Computers in the World of Gambling. Chapman and Hall, Ltd., London, UK, 1987. ???? pp.

Deuflhard:1990:FSM

[DFW90] P. Deuflhard, R. Freund, and A. Walter. Fast secant methods for the iterative solution of large nonsymmetric linear systems. Technical report, Preprint SC-90-5, Konrad-Zuse-Zentrum für Informationstechnik Berlin, ????, 1990. ???? pp.

Dongarra:1987:DMS

[DG87] Jack J. Dongarra and Eric Grosse. Distribution of mathematical software via electronic mail. Comm. ACM, 30(5):403– 407, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Demmel:1993:CAS

[DG93] James W. Demmel and William Gragg. On computing accurate singular values and eigenvalues of matrices with acyclic graphs. Linear Algebra and Appl., 185:203–217, 1993.

Dubrulle:1994:MQI

[DG94] Augustin A. Dubrulle and Gene H.

Golub. A multishift QR iteration without computation of the shifts. Numerical Algorithms, 7:173–181, 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

DeTurck:1989:YCH

[DGGW89] Dennis DeTurck, Herman Gluck, Carolyn Gordon, and David Webb. You can not hear the mass of a homology class. Comment. Math. Helvetici, 64:589–617, 1989. CODEN COMHAX. ISSN 0010-2571 (print), 1420-8946 (electronic).

Delsarte:1980:MMI

[DGK80] P. Delsarte, Y. Genin, and Y. Kamp. A method of matrix inverse triangular decomposition based on contiguous principal submatrices. Linear Algebra and Appl., 31:199–212, 1980.

Dongarra:1984:ILA

[DGK84] J. J. Dongarra, F. G. Gustavson, and A. Karp. Implementing linear algebra algorithms for dense matrices on a vector pipeline machine. SIAM Review, 26(1): 91–112, 1984.

Daniel:1976:RSA

[DGKS76] J. W. Daniel, W. B. Gragg, L. Kaufman, and G. W. Stewart. Reorthogonalization and stable algorithms for updating the Gram–Schmidt QR factorization. Math. Comp., 30(136):772– 795, 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duff:1989:SMT

[DGL89] Iain S. Duff, Roger G. Grimes, and John G. Lewis. Sparse matrix test problems. ACM Transactions on Mathematical Software, 15(1):1–14, 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dodson:19xx:SEF

[DGL91a] D. S. Dodson, R. G. Grimes, and J. G. Lewis. Sparse extensions to the FORTRAN basic linear algebra subprograms. ACM Transactions on Mathematical Software, 17(2):253–263, 1991. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dodson:19xx:AMI

[DGL91b] D. S. Dodson, R. G. Grimes, and J. G. Lewis. Algorithm 692: Model implementation and test package for the sparse basic linear algebra subprograms. ACM Transactions on Mathematical Software, 17(2):264–272, 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Duff:1992:UGH

[DGL92] Iain S. Duff, Roger G. Grimes, and John G. Lewis. Users’ guide for the Harwell–Boeing sparse matrix collection (release 1). Report RAL-92-086, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, December 1992. 84 pp.

Daubechies:1986:PNE

[DGM86] Ingrid Daubechies, A. Grossmann, and Y. Meyer. Painless nonorthogonal expansions. J. Math. Phys., 27(5):1271– 1283, 1986. CODEN JMAPAQ. ISSN 0022-2488.

Dubois:1979:AIM

[DGR79] P. F. Dubois, A. Greenbaum, and G. H. Rodrigue. Approximating the inverse of a matrix for use in iterative algorithms on vector processors. Computing, 22:257–268, 1979. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Duff:1991:FSS

[DGR^{+}91] I. S. Duff, N. I. M. Gould, J. K.

Reid, J. A. Scott, and K. Turner. The factorization of sparse symmetric indefinite matrices. IMA J. Numer. Anal., 11: 181–204, 1991.

Dongarra:1979:ULF

[DH79] J. J. Dongarra and A. R. Hinds. Unrolling loops in FORTRAN. Software —Practice and Experience, 9:216–226, 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dongarra:1986:IDL

[DH86] J. J. Dongarra and T. Hewitt. Implementing dense linear algebra algorithms using multitasking on the Cray X-MP-4 (or approaching the gigaflop). SIAM J. Sci. Stat. Comput., 7:347–350, 1986.

Dekker:1989:RGJ

[DH89] T. J. Dekker and W. Hoffmann. Rehabilitation of the Gauss–Jordan algorithm. Numer. Math., 54:591–599, 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dongarra:1990:ENS

[DH90] J. J. Dongarra and S. J. Hammarling. Evolution of numerical software for dense linear algebra. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 297–327. Oxford University Press, 1990.

Demmel:1992:SBA

[DH92] James W. Demmel and Nicholas J. Higham. Stability of block algorithms with fast level-3 BLAS. ACM Transactions on Mathematical Software, 18 (3):274–291, September 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demmel:1993:IEB

[DH93] James W. Demmel and Nicholas J. Higham. Improved error bounds for underdetermined system solvers. SIAM J. Matrix Anal. Appl., 14(1):1–14, January 1993.

Deuflhard:1995:NAF

[DH95] Peter Deuflhard and Andreas Hohmann. Numerical Analysis: A First Course in Scientific Computation. Walter de Gruyter, Berlin, 1995. ISBN 3-11-013882-4. xiii + 355 pp.

Dekker:1994:PAS

[DHP94] T. J. Dekker, W. Hoffmann, and K. Potma. Parallel algorithms for solving large linear systems. J. Comp. Appl. Math., 50:221–232, 1994.

Dongarra:1989:BRM

[DHS89] Jack J. Dongarra, Sven J. Hammarling, and Danny C. Sorensen. Block reduction of matrices to condensed forms for eigenvalue computations. J. Comp. Appl. Math., 27:215–227, 1989.

Demmel:1995:SBF

[DHS95] James W. Demmel, Nicholas J. Higham, and Robert S. Schreiber. Stability of block LU factorization. Numerical Linear Algebra with Applications, 2 (2):173–190, 1995.

Douglas:1994:GPL

[DHSS94] Craig C. Douglas, Michael Heroux, Gordon Slishman, and Roger M. Smith. GEMMW: A portable level 3 BLAS Winograd variant of Strassen’s matrix–matrix multiply algorithm. J. Comput. Phys., 110:1–10, 1994.

Demmel:1992:PNL

[DHvdV92] James W. Demmel, Michael T.

Heath, and Henk A. van der Vorst. Parallel numerical linear algebra. Technical Report UCB/CSD-92-703, Computer Science Division, University of California, Berkeley, ????, 1992. ???? pp. To appear in Acta Numerica, Cambridge University Press.

Demmel:1993:PNL

[DHvdV93] James W. Demmel, Michael T.

Heath, and Henk A. van der Vorst. Parallel numerical linear algebra. In Acta Numerica, pages 111–198. Cambridge University Press, 1993.

Dongarra:1992:NCC

[DHW92] J. J. Dongarra, S. J. Hammarling, and J. H. Wilkinson. Numerical considerations in computing invariant subspaces. SIAM J. Matrix Anal. Appl., 13(1):145– 161, 1992.

Delosme:1986:PSS

[DI86] J.-M. Delosme and I. C. F. Ipsen.

Parallel solution of symmetric positive definite systems with hyperbolic rotations. Linear Algebra and Appl., 77:75– 111, 1986.

Diamond:1978:SRI

[Dia78] Harold G. Diamond. Stability of rounded off inverses under iteration.

Math. Comp., 32(141):227–232, 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diaz:1989:MLS

[Día89] J. C. Díaz. Mathematics for Large Scale Computing. Marcel Dekker, Inc., New York, NY, USA, 1989. ???? pp.

Dickson:1992:DWT

[Dic92] Paul Dickson. Dickson’s Word Treasury: A Connoisseur’s Collection of Old and New, Weird and Wonderful, Useful and Outlandish Words. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1992. ISBN 0-471-55168-6. ???? pp.

Digest:1991:FGA

[Dig91] Reader’s Digest. Family Guide to Alternative Medicine. The Reader’s Digest Association Ltd., London, 1991. ???? pp.

Dillon:1981:ECE

[Dil81] J. T. Dillon. The emergence of the colon: An empirical correlate of scholarship. American Psychologist, 36(8):879– 884, 1981. CODEN AMPSAB. ISSN 0003-066x (print), 1935-990x (electronic).

Diller:1993:LLL

[Dil93] Antoni Diller. LaTEX Line by Line: Tips and Techniques for Document Processing. Wiley Professional Computing. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-93471-2 (softcover). xiii + 291 pp. LCCN Z253.4.L38 D55 1993. US$29.95. Optionally accompanied by disk with examples, ISBN 0-471-93797-5.

Dixon:1972:NO

[Dix72] L. C. W. Dixon. Nonlinear Optimisation. English Universities Press, London, 1972. ISBN 0-340-05241-4. ix + 214 pp.

Dixon:1973:S

[Dix73] Bernard Dixon. Sciwrite. Chemistry in Britain, 9(1):70–72, 1973. CODEN CHMBAY. ISSN 0009-3106.

Dixon:1976:OA

[Dix76] L. C. W. Dixon. Optimization in Action. Academic Press, ????, 1976. ???? pp.

Dixon:1983:EEE

[Dix83] John D. Dixon. Estimating extremal eigenvalues and condition numbers of matrices. SIAM J. Numer. Anal., 20(4):812– 814, 1983.

DePrima:1974:R

[DJ74] C. R. DePrima and C. R. Johnson.

The range of A^{−1}A^{∗} in GL(n, C). Linear Algebra and Appl., 9:209–222, 1974.

Jong:1977:TFD

[dJ77] Lieuwe Sytse de Jong. Towards a formal definition of numerical stability. Numer. Math., 28:211–219, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davis:1970:REP

[DK70] Chandler Davis and W. M. Kahan.

The rotation of eigenvectors by a perturbation. III. SIAM J. Numer. Anal., 7(1): 1–46, 1970.

Dax:1977:PTS

[DK77] Achiya Dax and S. Kaniel. Pivoting techniques for symmetric Gaussian elimination. Numer. Math., 28:221–241, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dax:1981:EMC

[DK81] A. Dax and S. Kaniel. The ELR method for computing the eigenvalues of a general matrix. SIAM J. Numer. Anal., 18:597–605, 1981.

Daniel:1983:CUN

[DK83] R. W. Daniel and B. Kouvaritakis. The choice and use of normal matrix approximations to transfer-function matrices of multivariable control systems. Internat. J. Control, 37:1121–1133, 1983.

Daniel:1984:ADL

[DK84] R. W. Daniel and B. Kouvaritakis. Analysis and design of linear multivariable feedback systems in the presence of additive perturbations. Internat. J. Control, 39:551–580, 1984.

Demmel:1985:IAS

[DK85] James W. Demmel and F. Krückeberg. An interval algorithm for solving systems of linear equations to prespecified accuracy. Computing, 34:117–129, 1985. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Demmel:1987:CSE

[DK87] James W. Demmel and Bo Kågström. Computing stable eigendecompositions of matrix pencils. Linear Algebra and Appl., 88/89:139–186, 1987.

Demmel:1988:ASI

[DK88] James W. Demmel and Bo Kågström. Accurate solutions of ill-posed problems in control theory. SIAM J. Matrix Anal. Appl., 9(1):126–145, 1988.

Demmel:1990:ASV

[DK90] James W. Demmel and W. Kahan.

Accurate singular values of bidiagonal matrices. SIAM J. Sci. Stat. Comput., 11(5):873–912, 1990.

Demmel:1993:GSDa

[DK93a] J. W. Demmel and B. Kågström.

The generalized Schur decomposition of an arbitrary pencil A − λB: Robust software with error bounds and applications. Part I: Theory and Algorithms. ACM Transactions on Mathematical Software, 19(2):160–174, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demmel:1993:GSDb

[DK93b] J. W. Demmel and B. Kågström.

The generalized Schur decomposition of an arbitrary pencil A − λB: Robust software with error bounds and applications. Part II: Software and Applications. ACM Transactions on Mathematical Software, 19(2):175–201, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1994:CCI

[DK94] Jack Dongarra and Michael Kolatis.

Call conversion interface (CCI) for LA-PACK/ESSL. Technical Report CS-94-250, Department of Computer Science, University of Tennessee, Knoxville, October 1994. 23 pp. LAPACK Working Note 82.

Dongarra:1986:SME

[DKH86] Jack J. Dongarra, Linda Kaufman, and Sven J. Hammarling. Squeezing the most out of eigenvalue solvers on highperformance computers. Linear Algebra and Appl., 77:113–136, 1986.

Davis:1982:NPD

[DKW82] Chandler Davis, W. M. Kahan, and H. F. Weinberger. Norm-preserving dilations and their applications to optimal error bounds. SIAM J. Numer. Anal., 19 (3):445–469, 1982.

Daubechies:1992:SMA

[DL92a] I. Daubechies and J. C. Lagarias. Sets of matrices all infinite products of which converge. Linear Algebra and Appl., 161:227–263, 1992.

Dieci:1992:BMC

[DL92b] L. Dieci and J. Lorenz. Block Mmatrices and computation of invariant tori. SIAM J. Sci. Stat. Comput., 13(4): 885–903, 1992.

Demmel:1993:FNA

[DL93] James W. Demmel and Xiaoye Li. Fast numerical algorithms via exception handling. LAPACK Working Note #59, Department of Computer Science, University of Tennessee, ????, 1993. ???? pp. To appear in Proceedings of 11th IEEE Symposium on Computer Arithmetic.

Demmel:1994:FNA

[DL94] James W. Demmel and Xiaoye Li. Faster numerical algorithms via exception handling. IEEE Trans. Comput., 43(8): 983–992, 1994.

Dongarra:1994:SML

[DLN^{+}94] Jack Dongarra, Andrew Lumsdaine, Xinhiu Niu, Roldan Pozo, and Karin Remington. A sparse matrix library in C++ for high performance architectures. Technical Report CS-94-236, Department of Computer Science, University of Tennessee, Knoxville, July 1994. 5 pp. LAPACK Working Note 74.

Millo:1979:SPP

[DLP79] R. A. De Millo, R. J. Lipton, and A. J. Perlis. Social processes and proofs of theorems and programs. Comm. ACM, 22:271–280, 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denman:1981:SDM

[DLR81] E. D. Denman and J. Leyva-Ramos.

Spectral decomposition of a matrix using the generalized sign matrix. Applied Mathematics and Computation, 8: 237–250, 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Denman:1985:CRR

[DLR85] E. D. Denman and J. Leyva-Ramos.

Computation of roots of real and complex matrices. Applied Mathematics and Computation, 16:213–228, 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Durham:1983:SCU

[DLS83] I. Durham, D. A. Lamb, and J. B.

Saxe. Spelling correction in user interfaces. Comm. ACM, 26:764–773, 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dorr:1971:REC

[DM71] Fred W. Dorr and Cleve B. Moler.

Roundoff error on the CDC 6600/7600 computers. ACM SIGNUM Newsletter, 8(2):24–26, 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Davis:1978:SME

[DM78] G. J. Davis and C. B. Moler. Sensitivity of matrix eigenvalues. International J. Numer. Methods Engineering, 12:1367–1373, 1978.

Dongarra:1984:EPS

[DM84] J. J. Dongarra and C. B. Moler. EISPACK—a package for solving matrix eigenvalue problems. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 68–87. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Davies:1986:DMC

[DM86] Roy O. Davies and J. J. Modi. A direct method for completing eigenproblem solutions on a parallel computer. Linear Algebra and Appl., 77:61–74, 1986.

Demmel:1989:TMG

[DM89] James W. Demmel and A. McKenney. A test matrix generation suite. Preprint MCS-P69-0389, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, March 1989. 16 pp. LAPACK Working Note 9.

Dixon:1990:ERE

[DM90] L. C. W. Dixon and D. J. Mills. The effect of rounding error on the variable metric method. Technical Report 229, Numerical Optimisation Centre, Hatfield Polytechnic, Hatfield, UK, April 1990. ???? pp.

Doyle:1991:LHL

[DM91] Terry Doyle and Paul Meara. Lingo! How to Learn a Language. BBC Books, ????, 1991. ???? pp.

Dongarra:1990:IRS

[DMdB90] J. J. Dongarra, P. Mayes, and G. Radicati di Brozolo. The IBM RISC system/6000 and linear algebra operations. Technical report, Report CS-90-122, Department of Computer Science, University of Tennessee, Knoxville, ????, 1990. ???? pp.

Duff:1992:PUL

[DMR92] Iain S. Duff, Michael Marrone, and Giuseppe Radicati. A proposal for user level sparse BLAS. Technical Report TR/PA/92/85, CERFACS, Toulouse, France, August 1992. ???? pp.

Demko:1984:DRI

[DMS84] Stephen Demko, William F. Moss, and Philip W. Smith. Decay rates for inverses of band matrices. Math. Comp., 43(168):491–499, 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dongarra:1983:IACb

[DMW83] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. SIAM J. Numer. Anal., 20:23– 45, 1983.

Daintith:1989:PDM

[DN89] John Daintith and R. D. Nelson, editors. The Penguin Dictionary of Mathematics. Penguin, London, 1989. ISBN 0-14-051119-9. ???? pp.

Duff:1987:ULP

[DNR87] I. S. Duff, J. Nocedal, and J. K.

Reid. The use of linear programming for the solution of sparse sets of nonlinear equations. SIAM J. Sci. Stat. Comput., 8:99–108, 1987.

Oliveira:1969:CSM

[dO69] G. N. de Oliveira. On a certain system of matrix equations. SIAM Review, 11: 378–381, 1969.

Dongarra:1990:LBF

[DO90] J. J. Dongarra and S. Ostrouchov. LAPACK block factorization algorithms on the intel iPSC/860. Technical report, LAPACK Working Note #24, Report CS-90-115, Department of Computer Science, University of Tennessee, Knoxville, ????, 1990. ???? pp.

Dongara:1982:ASF

[Don82] J. J. Dongara. Algorithm 589 SICEDR: A FORTRAN subroutine for improving the accuracy of computed matrix eigenvalues. ACM Transactions on Mathematical Software, 8:371–375, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1983:IACa

[Don83] J. J. Dongarra. Improving the accuracy of computed singular values. SIAM J. Sci. Stat. Comput., 4:712–719, 1983.

Don:1987:SSL

[Don87] F. J. Henk Don. On the symmetric solutions of a linear matrix equation. Linear Algebra and Appl., 93:1–7, 1987.

Dongarra:1988:LBE

[Don88a] J. J. Dongarra. The LINPACK benchmark: An explanation. In E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors, Supercomputing, First International Conference, Athens, Proceedings, volume 297 of Lecture Notes in Computer Science, pages 456–474. Springer-Verlag, Berlin, 1988.

Dongarra:1988:PC

[Don88b] J. J. Dongarra. Private communication, 1988. Technical report, ????, 1988. ???? pp.

Dongarra:1995:PVC

[Don95] Jack J. Dongarra. Performance of various computers using standard linear equations software. Technical Report CS-89-85, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, February 1995. 34 pp.

Dorr:1970:DSD

[Dor70] F. W. Dorr. The direct solution of the discrete poisson equation on a rectangle. SIAM Review, 12(2):248–263, 1970.

Dorr:1971:EIC

[Dor71] Fred W. Dorr. An example of illconditioning in the numerical solution of singular perturbation problems. Math. Comp., 25(114):271–283, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1959:REN

[Dou59] Jim Douglas, Jr. Round-off error in the numerical solution of the heat equation. J. Assoc. Comput. Mach., 6:48–58, 1959.

Drmac:1994:PCF

[DOV94] Zlatko Drmac̆, Matjaz̆ Omladic̆, and Kres̆imir Veselić. On the perturbation of the Cholesky factorization. SIAM J. Matrix Anal. Appl., 15(4):1319–1332, 1994.

Dowd:1993:HPC

[Dow93] Kevin Dowd. High Performance Computing. O’Reilly & Associates, Sebastopol, CA, 1993. ISBN 1-56592-032-5. xxv + 370 pp.

Doyle:1957:ISC

[Doy57] Thomas C. Doyle. Inversion of symmetric coefficient matrix of positivedefinite quadratic form. M.T.A.C., 11: 55–58, 1957.

Pillis:1980:HEY

[dP80] John de Pillis. How to embrace your spectrum for faster iterative results. Linear Algebra and Appl., 34:125–143, 1980.

Day:1988:GGE

[DP88] Jane M. Day and Brian Peterson. Growth in Gaussian elimination. Amer. Math. Monthly, 95(6):489–513, 1988.

Dongarra:1994:LDO

[DPW94] Jack Dongarra, Roldan Pozo, and David W. Walker. LAPACK++: A design overview of object-oriented extensions for high performance linear algebra. Technical Report CS-xx-yyy, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, ???? 1994. ???? pp. LAPACK Working Note ??

Dongarra:1995:LVH

[DPW95] Jack Dongarra, Roldan Pozo, and David W. Walker. LAPACK++ V1.0: High performance linear algebra users’ guide. Technical Report CS-95-290, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, May 1995. 31 pp. LAPACK Working Note 98.

Duff:1982:MSF

[DR82] Iain S. Duff and John K. Reid. MA27—A set of Fortran subroutines for solving sparse symmetric sets of linear equations. Technical Report AERE R10533, AERE Harwell Laboratory, London, July 1982. ISBN 0-7058-0985-4. ???? pp. Published by Her Majesty’s Stationary Office.

Duff:1983:MSI

[DR83] Iain S. Duff and John K. Reid. The multifrontal solution of indefinite sparse symmetric linear equations. ACM Transactions on Mathematical Software, 9(3): 302–325, 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davis:1984:MNI

[DR84] Philip J. Davis and Philip Rabinowitz.

Methods of Numerical Integration. Academic Press, London, second edition, 1984. ISBN 0-12-206360-0. xiv + 612 pp.

Dongarra:1991:NNA

[DR91] Jack J. Dongarra and Bill Rosener.

NA-NET: Numerical analysis NET. Technical Report CS-91-146, Department of Computer Science, University of Tennessee, Knoxville, September 1991. ???? pp.

Duff:1993:MFC

[DR93] Iain S. Duff and John K. Reid. MA48, a Fortran code for direct solution of sparse unsymmetric linear systems of equations. Report RAL-93-072, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, October 1993. ???? pp.

Deif:1994:ISP

[DR94] A. Deif and J. Rohn. On the invariance of the sign pattern of matrix eigenvectors under perturbation. Linear Algebra and Appl., 196:63–70, 1994.

Duff:1995:MFC

[DR95] Iain S. Duff and John K. Reid. MA47, a Fortran code for direct solution of indefinite sparse symmetric linear systems. Report RAL-95-001, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, January 1995. 63 pp.

Drazin:1992:NS

[Dra92] P. G. Drazin. Nonlinear Systems. Cambridge University Press, ????, 1992. ISBN 0-521-40668-4. ???? pp.

Duff:1979:DSS

[DRMN79] Iain S. Duff, John K. Reid, Neils Munskgaard, and Hans B. Nielsen. Direct solution of sets of linear equations whose matrix is sparse, symmetric and indefinite. J. Inst. Maths Applics, 23:235–250, 1979.

Dongarra:1993:SDU

[DRW93] Jack J. Dongarra, Thomas H. Rowan, and Reed C. Wade. Software distribution using Xnetlib. Report ORNL/TM-12318, Oak Ridge National Laboratory, Oak Ridge, TN, USA, June 1993. ???? pp.

Dixon:1975:TGO

[DS75] L. C. W. Dixon and G. P. Szegő. Towards Global Optimization. North-Holland, Amsterdam, The Netherlands, 1975. ???? pp.

Dennis:1979:LCS

[DS79a] J. E. Dennis, Jr. and R. B. Schnabel. Least change secant updates for quasi-Newton methods. SIAM Review, 21:443– 459, 1979.

Duff:1979:SMP

[DS79b] I. S. Duff and G. W. Stewart, editors. Sparse Matrix Proceedings 1978. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ???? pp.

Deuflhard:1980:RDP

[DS80] P. Deuflhard and W. Sautter. On rank-deficient pseudoinverses. Linear Algebra and Appl., 29:91–111, 1980.

Dennis:1983:NMU

[DS83] J. E. Dennis, Jr. and Robert B.

Schnabel. Numerical Methods for Unconstrained Optimization and Nonlinear Equations. Prentice-Hall, Englewood Cliffs, NJ, USA, 1983. ISBN 0-13-627216-9. xiii + 378 pp.

Dongarra:1984:LPS

[DS84] J. J. Dongarra and G. W. Stewart.

LINPACK—A package for solving linear systems. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 20–48. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Dongarra:1986:HPC

[DS86a] J. J. Dongarra and D. C. Sorensen.

High performance computers and algorithms from linear algebra. In Large Scale Eigenvalue Problems [CW86], pages 15– 36.

Dongarra:1986:LAH

[DS86b] J. J. Dongarra and D. C. Sorensen.

Linear algebra on high-performance computers. Applied Mathematics and Computation, 20:57–88, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dongarra:1987:FPA

[DS87] J. J. Dongarra and D. C. Sorensen. A fully parallel algorithm for the symmetric eigenvalue problem. SIAM J. Sci. Stat. Comput., 8:s139–s154, 1987.

Dongarra:1988:ADH

[DS88] J. J. Dongarra and D. C. Sorensen. Algorithm design for high-performance computers. In Paul and Almasi [PA88], pages 83–97.

Demeure:1989:FLS

[DS89] Cédric J. Demeure and Louis L. Scharf. Fast least squares solution of Vandermonde systems of equations. In Proceedings of the IEEE International Conference on Acoustics, Speech and Signal Processing, May 1989, Glasgow, volume 4, pages 2198–2201. ????, ????, 1989.

Dongarra:1992:PAN

[DS92] J. J. Dongarra and M. Sidani. A parallel algorithm for the non-symmetric eigenvalue problem. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 85– 102. Longman Scientific and Technical, Harlow, Essex, UK, 1992.

Dongarra:1993:PAN

[DS93a] Jack J. Dongarra and Majed Sidani. A parallel algorithm for the nonsymmetric eigenvalue problem. SIAM J. Sci. Comput., 14(3):542–569, 1993.

Duff:1993:MNF

[DS93b] Iain S. Duff and Jennifer A. Scott. MA42—A new frontal code for solving sparse unsymmetric systems. Report RAL-93-064, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, September 1993. ???? pp.

Douglas:1994:VMM

[DS94] Craig C. Douglas and Gordon Slishman. Variants of matrix-matrix multiplication for Fortran-90. ACM SIGNUM Newsletter, 29:4–6, 1994. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Don:1995:ASC

[DS95] Wai Sun Don and Alex Solomonoff.

Accuracy and speed in computing the Chebyshev collocation derivative. SIAM J. Sci. Comput., 16(6):1253–1268, 1995.

Dongarra:1988:TMP

[DSB88] J. J. Dongarra, D. C. Sorensen, and O. Brewer. Tools and methodology for programming parallel processors. Technical report, Preprint MC-P29-1288, Mathematics and Computer Science Division, Argonne National Laboratory, Illinois, ????, 1988. ???? pp.

Deif:1995:SEA

[DSH95] A. S. Deif, N. P. Seif, and S. A. Hussein. Sylvester’s equation: Accuracy and computational stability. J. Comp. Appl. Math., 61:1–11, 1995.

Davenport:1988:CAS

[DST88] J. H. Davenport, Y. Siret, and E. Tournier, editors. Computer Algebra: Systems and Algorithms for Algebraic Computation. Academic Press, London, 1988. ???? pp.

Dennis:1991:DSM

[DT91] J. E. Dennis, Jr. and Virginia Torczon. Direct search methods on parallel machines. SIAM J. Optimization, 1(4): 448–474, 1991.

Dongarra:1993:ETP

[DT93] Jack. J. Dongarra and Bernard Tourancheau, editors. Environments and Tools For Parallel Scientific Computing.

North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-89963-4.

Dongarra:1994:ETP

[DT94] Jack. J. Dongarra and Bernard Tourancheau, editors. Environments and Tools For Parallel Scientific Computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-343-9.

Dubrulle:1983:CNM

[Dub83] Augustin A. Dubrulle. A class of numerical methods for the computation of Pythagorean sums. IBM Journal of Research and Development, 27(6):582–589, 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Dubrulle:1991:BGR

[Dub91a] A. A. Dubrulle. Block Gauss reduction to Hessenberg form. SIAM J. Sci. Stat. Comput., 12:1245–1253, 1991.

Dubrulle:1991:MQA

[Dub91b] A. A. Dubrulle. The multishift QR algorithm: Is it worth the trouble? Report No. G320-3558x, IBM Corporation, Palo Alto Scientific Center, CA, August 1991. ???? pp.

Dubrulle:1994:FIM

[Dub94] Augustin A. Dubrulle. Frobenius iteration for the matrix polar decomposition. Report HPL-94-??, Hewlette Packard Computer Research Center, CA, December 1994. 9 pp.

Dudley:1977:POW

[Dud77] Hugh Dudley, editor. The Presentation of Original Work in Medicine and Biology. Churchill Livingstone, Edinburgh, 1977. ISBN 0-443-01583-X. ???? pp.

Duff:1977:SSM

[Duf77] I. S. Duff. A survey of sparse matrix research. Proc. IEEE, 65:500–535, 1977. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Duff:1981:MAS

[Duf81] I. S. Duff. ME28: A sparse unsymmetric linear equation solver for complex equations. ACM Transactions on Mathematical Software, 7(4):505–511, 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Duff:1984:RDS

[Duf84a] I. S. Duff. Research directions in sparse matrix computations. In Golub [Gol84], pages 83–139.

Duff:1984:SSM

[Duf84b] I. S. Duff. A survey of sparse matrix software. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 165–199. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Duff:1988:PSM

[Duf88] I. S. Duff. Parallelism in sparse matrices. In Paul and Almasi [PA88], pages 99–106.

Duff:1989:MSM

[Duf89a] I. S. Duff. Multiprocessing a sparse matrix code on the alliant FX/8. J. Comp. Appl. Math., 27:229–239, 1989.

Duff:1989:PAS

[Duf89b] I. S. Duff. Parallel algorithms for sparse matrix solution. In Parallel Computing: Methods, Algorithms and Applications [ES89], pages 73–82.

Dunham:1982:CBC

[Dun82] Charles B. Dunham. Choice of basis for Chebyshev approximation. ACM Transactions on Mathematical Software, 8(1):21–25, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dunham:1990:JTG

[Dun90] William Dunham. Journey Through Genius: The Great Theorems of Mathematics. Penguin, New York, 1990. ISBN 0-14-014739-X. xiii + 300 pp.

Demmel:1992:JMM

[DV92a] James W. Demmel and Kres̆imir Veselić. Jacobi’s method is more accurate than QR. SIAM J. Matrix Anal. Appl., 13(4):1204–1245, 1992.

Dongarra:1992:PVC

[DV92b] Jack J. Dongarra and Henk A. Van der Vorst. Performance of various computers using standard techniques for solving sparse linear equations. Technical Report CS-92-168, Department of Computer Science, University of Tennessee, ????, 1992. ???? pp.

Dongarra:1992:RCF

[DvdG92] J. J. Dongarra and R. A. van de Geijn. Reduction to condensed form for the eigenvalue problem on distributed memory architectures. Report ORNL/TM-12006, Oak Ridge National Laboratory, Tennessee, January 1992. ???? pp.

Dongarra:1992:LAS

[DvdGW92] J. J. Dongarra, R. A. van de Geijn, and D. W. Walker. A look at scalable dense linear algebra libraries. Report ORNL/TM-12126, Oak Ridge National Laboratory, Tennessee, July 1992. ???? pp.

Daniel:1980:FED

[DW80] C. Daniel and F. S. Wood. Fitting Equations to Data. Wiley, New York, NY, USA; London, UK; Sydney, Australia, second edition, 1980. ???? pp.

Dennis:1984:IQM

[DW84] J. E. Dennis, Jr. and Homer F.

Walker. Inaccuracy in quasi-Newton methods: Local improvement theorems. Mathematical Programming Study, 22:70– 85, 1984. CODEN MPSTDF. ISSN 0303-3929.

Wouk:1987:OMN

[DW87] John E. Dennis, Jr. and Daniel J.

Woods. Optimization on microcomputers: The Nelder–Mead simplex algorithm. In Arthur Wouk, editor, New Computing Environments: Microcomputers in Large-Scale Computing, pages 116–122. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987.

Dongarra:1993:DLA

[DW93] Jack J. Dongarra and David W.

Walker. The design of linear algebra libraries for high performance computers. Report ORNL/TM-12404, Oak Ridge National Laboratory, Tennessee, August 1993. ???? pp.

Dongarra:1995:SLL

[DW95] Jack J. Dongarra and David W.

Walker. Software libraries for linear algebra computations on high performance computers. SIAM Review, 37(2):151–180,

Dwyer:1951:LC

[Dwy51] Paul S. Dwyer. Linear Computations. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1951. xi + 344 pp.

Dwyer:1953:EMC

[Dwy53] Paul S. Dwyer. Errors of matrix computations. In L. J. Paige and Olga Taussky, editors, Simultaneous Linear Equations and the Determination of Eigenvalues, number 29 in Applied Mathematics Series, pages 49–58. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1953.

Dwyer:1967:SAM

[Dwy67] P. S. Dwyer. Some applications of matrix derivatives in multivariate analysis. J. Amer. Statist. Assoc., 62:607–625, 1967.

Dyson:1993:GGP

[Dys93] Freeman Dyson. George Green and physics. Physics world, 6(8):33– 38, August 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

                                        tronic). URL iopscience.iop.org/pwa/
                                        full/pwa-pdf/6/8/phwv6i8a28.pdf.

Edelman:1994:CCG

[EAS94] Alan Edelman, Tomaás A. Arias, and Steven T. Smith. Curvature in conjugate gradient eigenvalue computation with applications to materials and chemistry calculations. 1994.

Eberlein:1965:MNN

[Ebe65] P. J. Eberlein. On measures of nonnormality for matrices. Amer. Math. Monthly, 72:995–996, 1965.

Eberlein:1971:NMD

[Ebe71a] P. J. Eberlein. A note on the matrices denoted by B_{n}. SIAM J. Appl. Math., 20(1):87–92, 1971.

Eberlein:1971:DCS

[Ebe71b] P. J. Eberlein. On the diagonalization of complex symmetric matrices. J. Inst. Maths Applics, 7:377–383, 1971.

Ebel:1987:ASW

[EBR87] Hans F. Ebel, Claus Bliefert, and William E. Russey. The Art of Scientific Writing: From Student Reports to Professional Publications in Chemistry and Related Fields. VCH Publishers, New York, 1987. ISBN 0-89573-645-4. ???? pp.

Evans:1983:PST

[ED83] D. J. Evans and R. C. Dunbar. The parallel solution of triangular systems of equations. IEEE Trans. Comput., C-32 (2):201–204, 1983.

Edelman:1988:ECN

[Ede88] Alan Edelman. Eigenvalues and condition numbers of random matrices. SIAM J. Matrix Anal. Appl., 9(4):543– 560, 1988.

Edelman:1989:ECN

[Ede89] Alan Edelman. Eigenvalues and Condition Numbers of Random Matrices. PhD thesis, M.I.T., Cambridge, MA, May 1989. ???? pp. Numerical Analysis Report 89-7, Department of Mathematics.

Edelman:1991:DMS

[Ede91a] Alan Edelman. The distribution and moments of the smallest eigenvalue of a random matrix of Wishart type. Linear Algebra and Appl., 159:55–80, 1991.

Edelman:1991:FAL

[Ede91b] Alan Edelman. The first annual large dense linear system survey. ACM SIGNUM Newsletter, 26:6–12, October 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Edelman:1992:CPC

[Ede92a] Alan Edelman. The complete pivoting conjecture for Gaussian elimination is false. Mathematica Journal, 2:58–61, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Edelman:1992:DSC

[Ede92b] Alan Edelman. On the distribution of a scaled condition number. Math. Comp., 58(197):185–190, 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Edelman:1993:ERR

[Ede93a] Alan Edelman. Eigenvalue roulette and random test matrices. In Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors, Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E, pages 365–368. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993.

Edelman:1993:LDN

[Ede93b] Alan Edelman. Large dense numerical linear algebra in 1993: The parallel computing influence. Int. J. Supercomputer Appl., 7(2):113–128, 1993. CODEN IJSAE9. ISSN 0890-2720.

Edelman:1994:SDN

[Ede94a] Alan Edelman. Scalable dense numerical linear algebra in 1994: The multicomputer influence. In John G. Lewis, editor, Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, pages 344–348. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Edelman:1994:W

[Ede94b] Alan Edelman. When is x ∗ (1/x) ≠ 1? Manuscript, 1994.

Edelman:1995:PFD

[Ede95] Alan Edelman. The Pentium flaw:

A devilish recreational mathematical puzzle. Manuscript, July 7, 1995.

Edmondson:1985:CCF

[Edm85] Tod Edmondson, editor. Columbia’s Colossal Fake Book. Columbia Pictures Publications, Hialeah, FL 33014, 1985. ISBN 0-89898-342-8. ???? pp.

Edmonds:1984:WGB

[EE84] Allan Edmonds and John Ewing.

What’s a good book review? An editorial comment. Amer. Math. Monthly, 91 (5):318–319, 1984.

Ewing:1991:PHC

[EG91] John H. Ewing and F. W. Gehring, editors. Paul Halmos: celebrating 50 years of mathematics. Springer-Verlag, Berlin, 1991. ISBN 3-540-97509-8. viii + 320 pp. LCCN QA300.5 .H35 1991.

Erisman:1987:EOU

[EGL^{+}87] A. M. Erisman, R. G. Grimes, J. G.

Lewis, W. G. Poole, and H. D. Simon. Evaluation of orderings for unsymmetric sparse matrices. SIAM J. Sci. Stat. Comput., 8(4):600–624, 1987.

Evans:1979:COS

[EH79] D. J. Evans and M. Hatzopoulos. A comparison of optimal scaling and preconditioning. ACM SIGNUM Newsletter, 14:20–22, 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Elsner:1991:ACD

[EH91a] L. Elsner and C. He. An algorithm for computing the distance to uncontrollability. Systems and Control Letters, 17: 453–464, 1991.

Elsner:1991:PIT

[EH91b] L. Elsner and C. He. Perturbation and interlace theorems for the unitary eigenvalue problem. Linear Algebra and Appl., page ??, 1991. To appear.

Elsner:1992:EFS

[EH92] L. Elsner and C. He. An eigenvector formula for symmetric and unitary matrices. Manuscript, 1992.

Eisenstat:1988:MCA

[EHHR88] S. C. Eisenstat, Michael T. Heath, C. S. Henkel, and Charles H. Romine. Modified cyclic algorithms for solving triangular systems on distributed-memory multiprocessors. SIAM J. Sci. Stat. Comput., 9(3):589–600, 1988.

Elsner:1994:MCN

[EHM94] L. Elsner, C. He, and V. Mehrmann. Minimizing the condition number of a positive definite matrix by completion. Numer. Math., 69:17–23, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ehrlich:1970:CMI

[Ehr70] L. W. Ehrlich. Complex matrix inversion versus real. Comm. ACM, 13(9):

561–562, 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eisenstat:1993:RPT

[EI93] Stanley C. Eisenstat and Ilse C. F.

Ipsen. Relative perturbation techniques for singular value problems. Research Report YALEU/DCS/RR-942, Department of Computer Science, Yale University, ????, July 1993. ???? pp.

Einarsson:1979:BEN

[Ein79] B. Einarsson. Bibliography on the evaluation of numerical software. J. Comp. Appl. Math., 5:145–159, 1979.

Eirola:1993:ABE

[Eir93] Timo Eirola. Aspects of backward error analysis of numerical ODEs. J. Comp. Appl. Math., 45(1-2):65–73, 1993.

Eisenhart:1968:EUF

[Eis68] Churchill Eisenhart. Expression of the uncertainties of final results. Science, 160 (3833):1201–1204, 1968. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Eisenberg:1992:GTE

[Eis92] Anne Eisenberg. Guide to Technical Editing: Discussion, Dictionary, and Exercises. Oxford University Press, New York, 1992. ISBN 0-19-506306-6. ???? pp.

Edlefsen:1986:GPL

[EJ86] L. E. Edlefsen and S. D. Jones. GAUSS programming language manual. Technical report, Aptech Systems, Inc., 26250 196th Place South East, Kent, WA 98042, ????, 1986. ???? pp.

Egecloglu:1989:FAR

[EK89] Ömer Eg̃eclog̃lu and Çetin K. Koç. A fast algorithm for rational interpolation via orthogonal polynomials. Math. Comp., 53(187):249–264, 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Egecloglu:1992:PPC

[EK92] Ö. Eg̃eclog̃lu and Ç. K. Koç. Parallel prefix computation with few processors. Computers and Mathematics with Applications, 24(4):77–84, 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Edelman:1993:HMR

[EK93] Alan Edelman and Eric Kostlan. How many roots of a random polynomial are real? 1993.

Edelman:1994:RKM

[EK94] Alan Edelman and Eric Kostlan. The road from Kac’s matrix to Kac’s random polynomials. In John G. Lewis, editor, Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, pages 503–507. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Elsner:1993:ACP

[EKNX93] L. Elsner, I. Koltracht, M. Neumann, and D. Xiao. On accurate computations of the Perron root. SIAM J. Matrix Anal. Appl., 14(2):456–467, 1993.

Edelman:1994:HME

[EKS94] Alan Edelman, Eric Kostlan, and Michael Shub. How many eigenvalues of a random matrix are real? J. Amer. Math. Soc., 7(1):247–267, 1994.

Ewer:1969:CBS

[EL69] J. R. Ewer and G. Latorre. A Course in Basic Scientific English. Longman, Harlow, Essex, 1969. ISBN 0-582-52009-6. ???? pp.

Elden:1980:PTL

[Eld80] Lars Eldén. Perturbation theory for the least squares problem with linear equality constraints. SIAM J. Numer. Anal., 17(3):338–350, 1980.

Ellacott:1994:TMA

[Ell94] S. W. Ellacott. Techniques for the mathematical analysis of neural networks. J. Comp. Appl. Math., 50:283–297, 1994.

Elsner:1976:IIC

[Els76] Ludwig Elsner. Inverse iteration for calculating the spectral radius of a nonnegative irreducible matrix. Linear Algebra and Appl., 15:235–242, 1976.

Edelman:1994:PRC

[EM94] Alan Edelman and H. Murakami.

Polynomial roots from companion matrix eigenvalues. In John G. Lewis, editor, Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, pages 326– 330. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. Full version [EM95b].

Edelman:1995:CPC

[EM95a] Alan Edelman and Walter Mascarenhas. On the complete pivoting conjecture for a Hadamard matrix of order 12. Linear and Multilinear Algebra, 38(3): 181–188, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Edelman:1995:PRC

[EM95b] Alan Edelman and H. Murakami. Polynomial roots from companion matrix eigenvalues. Math. Comp., 64(210):763– 776, 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eiermann:1988:SSL

[EMN88] Michael Eiermann, Ivo Marek, and Wilhelm Niethammer. On the solution of singular linear systems of algebraic equations by semiiterative methods. Numer. Math., 53:265–283, 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Enis:1973:R

[Eni73] Peter Enis. On the relation E(X) = E{E(x|Y)}. Biometrika, 60(2):432–433, 1973. ISSN 0006-3444 (print), 1464-3510 (electronic).

Enright:1979:ERN

[Enr79] W. Enright. On the efficient and reliable numerical solution of large linear systems of ODE’s. IEEE Trans. Automat. Control., AC-24:905–908, 1979.

Enright:1989:NEC

[Enr89] W. H. Enright. A new error-control for initial value solvers. Applied Mathematics and Computation, 31:288–301, 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1979:GSF

[EO79] D. J. Evans and S. O. Okolie. A generalised sparse factorisation method for the solution of periodic tridiagonal systems. Computers and Mathematics with Applications, 5:211–216, 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Elsner:1987:MNM

[EP87a] L. Elsner and M. H. C. Paardekooper.

On measures of nonnormality of matrices. Linear Algebra and Appl., 92:107– 124, 1987.

Enright:19xx:TFP

[EP87b] W. H. Enright and J. D. Pryce.

Two FORTRAN packages for assessing initial value methods. ACM Transactions on Mathematical Software, 13(1):1– 27, 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Enright:19xx:ANS

[EP87c] W. H. Enright and J. D. Pryce. Algorithm 648. NSDTST and STDTST: Routines for assessing the performance of IV solvers. ACM Transactions on Mathematical Software, 13(1):28–34, 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Eisinburg:1994:EIC

[EP94a] Alfredo Eisinburg and Paolo Pugliese.

Exact inversion of a class of Vandermonde matrices. In John G. Lewis, editor, Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, pages 239–243. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Elden:1994:PAB

[EP94b] Lars Eldén and Haesun Park. Perturbation analysis for block downdating of a Cholesky decomposition. Numer. Math., 68(4):457–467, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ellis:1994:FP

[EPL94] T. M. R. Ellis, Ivor R. Phillips, and Thomas M. Lahey. Fortran 90 Programming. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54446-6. xviii + 825 pp.

Epton:1980:MSA

[Ept80] Michael A. Epton. Methods for the solution of AXD − BXC = E and its application in the numerical solution of implicit ordinary differential equations. BIT, 20:341–345, 1980. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Erisman:1974:MST

[ER74] A. M. Erisman and J. K. Reid. Monitoring the stability of the triangular factorization of a sparse matrix. Numer. Math., 22:183–186, 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Erhel:1993:EDP

[Erh93] J. Erhel. Experiments with data perturbations to study condition numbers and numerical stability. Computing, 51:29–44, 1993. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Ellacott:1988:CMG

[ES88] S. W. Ellacott and E. B. Saff. On Clenshaw’s method and a generalisation to Faber series. Numer. Math., 52:499–509, 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Evans:1989:PCM

[ES89] D. J. Evans and C. Sutti. Parallel Computing: Methods, Algorithms and Applications. Adam Hilger, Bristol, 1989. ???? pp.

Eldersveld:1992:BUL

[ES92] Samuel K. Eldersveld and Michael A.

Saunders. A block-LU update for largescale linear programming. SIAM J. Matrix Anal. Appl., 13(1):191–201, 1992.

Edelman:1993:SO

[ES93] Alan Edelman and G. W. Stewart.

Scaling for orthogonality. IEEE Trans. Signal Processing, 41(4):1676–1677, 1993.

Espelid:1978:FPS

[Esp78] Terje O. Espelid. On floating-point summation. Report No. 67, Department of Applied Mathematics, University of Bergen, Bergen, Norway, December 1978. ???? pp.

Eisenhart:1973:CHP

[ET73] Laurance C. Eisenhart and George E.

Trapp, Jr. Computation of Hermite polynomials. Math. Comp., 27(123):625–632, 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Enns:1990:SMI

[ETA90] M. K. Enns, W. F. Tinney, and F. L.

Alvarado. Sparse matrix inverse factors. IEEE Transactions on Power Systems, 5(2):466–472, 1990. CODEN ITPSEG. ISSN 0885-8950 (print), 1558-0679 (electronic).

Evard:1992:ME

[EU92] Jean-Claude Evard and Frank Uhlig.

On the matrix equation f(X) = A. Linear Algebra and Appl., 162–164:447–519,

Eichorn:1985:JPW

[EV85] Dorothy H. Eichorn and Gary R. VandenBos. Journal publication within the IPA. American Psychologist, 40(12): 1309–1316, 1985. CODEN AMPSAB. ISSN 0003-066x (print), 1935-990x (electronic).

Evans:1971:NSS

[Eva71] D. J. Evans. Numerical solution of the Sturm Liouville problem with periodic boundary conditions. In J. Ll. Morris, editor, Conference on Applications of Numerical Analysis, volume 228 of Lecture Notes in Mathematics, pages 277– 289. Springer-Verlag, Berlin, 1971.

Evans:1974:SNM

[Eva74] D. J. Evans, editor. Software for Numerical Mathematics. Academic Press, London, 1974. ???? pp.

Evans:1976:IJH

[Eva76] Christopher Evans. Interview with J. H. Wilkinson, page ?? Number 10 in Pioneers of Computing, 60-Minute Recordings of Interviews. Science Museum, London, 1976.

Evans:1981:MMH

[Eva81] Christopher Evans. The Making of the Micro: A History of the Computer. Victor Gollanz, London, 1981. ISBN 0-575-02913-7. 118 pp.

Evans:1982:PPS

[Eva82] David J. Evans, editor. Parallel Processing Systems. Cambridge University Press, ????, 1982. ISBN 0-521-24366-1. ???? pp.

Eves:1966:EMT

[Eve66] Howard Eves. Elementary Matrix Theory. Dover, New York, NY, USA, 1966. ???? pp.

Eves:1969:MC

[Eve69] Howard W. Eves. In Mathematical Circles. Prindle, Weber and Schmidt, Boston, MA, USA, 1969. ???? pp.

Eves:1988:RMC

[Eve88] Howard W. Eves. Return to Mathematical Circles. PWS-Kent Publishing Company, Boston, 1988. ???? pp.

Evans:1990:SS

[EW90] R. J. Evans and S. Wilson, editors.

Supercomputational Science. Plenum Press, New York, NY, USA; London, UK, 1990. ???? pp.

Ewing:1994:CMT

[Ewi94] John Ewing, editor. A Century of Mathematics Through the Eyes of the Monthly. Mathematical Association of America, Washington, DC, USA, 1994. ISBN 0-88385-459-7. xi + 323 pp.

Elden:1990:NAI

[EWK90] L. Eldén and L. Wittmeyer-Koch.

Numerical Analysis: An Introduction. Academic Press, ????, 1990. ???? pp.

Eckart:1936:AOM

[EY36] Carl Eckart and Gale Young. The approximation of one matrix by another of lower rank. Psychometrika, 1(3):211–218, 1936. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Faddeeva:1959:CML

[Fad59] V. N. Faddeeva. Computational Methods of Linear Algebra. Dover, New York, NY, USA, 1959. ISBN 0-486-60424-1. x + 252 pp.

Fairweather:1971:EEC

[Fai71] Graeme Fairweather. On the eigenvalues and eigenvectors of a class of Hessenberg matrices. SIAM Review, 13(2):220– 221, 1971.

Fam:1988:ECM

[Fam88] Adly T. Fam. Efficient complex matrix multiplication. IEEE Trans. Comput., C-37(7):877–879, 1988.

Fan:1951:MPI

[Fan51] K. Fan. Maximum properties and inequalities for the eigenvalues of completely continuous operators. Proceedings of the National Academy of Sciences, 37: 760–766, 1951.

Fan:1960:NM

[Fan60] Ky Fan. Note on M-matrices. Quarterly Journal of Mathematics, 11:43–49, 1960. CODEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Fan:1993:QCL

[Fan93] Michael K. H. Fan. A quadratically convergent local algorithm on minimizing the largest eigenvalue of a symmetric matrix. Linear Algebra and Appl., 188:231– 253, 1993.

Farebrother:1985:LUA

[Far85] R. W. Farebrother. Linear unbiased approximators of the disturbances in the standard linear model. Linear Algebra and Appl., 67:259–274, 1985.

Farebrother:1987:MLM

[Far87a] R. W. Farebrother. A memoir of the life of M. H. Doolittle. IMA Bulletin, 23 (6/7):102, 1987.

Farebrother:1987:TTA

[Far87b] R. W. Farebrother. Three theorems with applications to Euclidean distance matrices. Linear Algebra and Appl., 95: 11–16, 1987.

Farnon:1987:JM

[Far87c] Robert Farnon. Journey Into Melody.

Chappell Music Ltd., International Music Publications, ????, 1987. ISBN 0-86359-529-4. ???? pp.

Farebrother:1988:LLS

[Far88a] R. W. Farebrother. Linear Least Squares Computations. Marcel Dekker, Inc., New York, NY, USA, 1988. ISBN 0-8247-7661-5. xiii + 293 pp.

Farnum:1988:CSF

[Far88b] Charles Farnum. Compiler support for floating-point computation. Software —Practice and Experience, 18(7):701– 709, 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fateman:1982:HLL

[Fat82] Richard J. Fateman. High-level language implications of the proposed IEEE floating-point standard. ACM Transactions on Programming Languages and Systems, 4(2):239–257, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kornerup:1991:AMA

[FB91] Warren E. Ferguson, Jr. and Tom Brightman. Accurate and monotone approximations of some transcendental functions. In Peter Kornerup and David W. Matula, editors, Proc. 10th IEEE Symposium on Computer Arithmetic, pages 237–244. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991.

Friedman:1986:SJR

[FDR86] Sharon M. Friedman, Sharon Dunwoody, and Carol L. Rogers, editors. Scientists as Journalists: Reporting Science as News. American Association for the Advancement of Science, Washington, 1986. ???? pp.

Feilmeier:1982:PNA

[Fei82] M. Feilmeier. Parallel numerical algorithms. In David J. Evans, editor, Parallel Processing Systems, pages 285–338. Cambridge University Press, 1982.

Felippa:1975:SLE

[Fel75] C. A. Felippa. Solution of linear equations with skyline-stored symmetric matrix. Computers and Structures, 5:13–29, 1975. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ferguson:1994:ECS

[Fer94] Warren E. Ferguson, Jr. Exact computation of a sum or difference with applications to argument reduction. Submitted to 12th IEEE Symposium on Computer Arithmetic, 1994.

Knowles:1995:ECS

[Fer95] Warren E. Ferguson, Jr. Exact computation of a sum or difference with applications to argument reduction. In Simon Knowles and William H. McAllister, editors, Proc. 12th IEEE Symposium on Computer Arithmetic, Bath, England, pages 216–221. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995.

Feynman:1985:SYJ

[Fey85] Richard P. Feynman. “Surely You’re Joking, Mr. Feynman!”: Adventures of a Curious Character. Bantam, New York, 1985. ISBN 0-553-25649-1. xi + 322 pp.

Fowler:1931:KE

[FF31] H. W. Fowler and F. G. Fowler. The King’s English. Oxford University Press, ????, third edition, 1931. ISBN 0-19-881330-9. ???? pp. Reprinted 1990.

Feller:1951:NMT

[FF51] William Feller and George E. Forsythe.

New matrix transformations for obtaining characteristic vectors. Quarterly of Applied Mathematics, 8(4):325–331, 1951. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

Faddeev:1963:CML

[FF63] D. K. Faddeev and V. N. Faddeeva.

Computational Methods of Linear Algebra. W.H. Freeman, San Francisco, CA, USA and New York, NY, USA, 1963. ???? pp.

Faddeev:1964:NML

[FF64] D. K. Faddeev and V. N. Faddeeva.

Numerische Methoden der Linearen Algebra. Veb Deutscher Verlag der Wissenschaften, Berlin, 1964. ???? pp.

Farewell:1987:WSG

[FF87a] W. S. Farewell and W. P. Forever. A Wordstar Survivor’s Guide to WordPerfect. WordPerfect Publications, Orem, Utah, 1987. ???? pp.

Fincham:1987:PC

[FF87b] A. E. Fincham and B. Ford, editors. Parallel Computation. Oxford University Press, ????, 1987. ISBN 0-19-853680-1.

Friedman:1987:LL

[FF87c] D. P. Friedman and M. Felleisen. The Little LISPer. MIT Press, ????, trade edition, 1987. ???? pp.

Flavell:1992:DIT

[FF92] Linda M. Flavell and Roger H. Flavell. Dictionary of Idioms and their Origins. Kyle Cathie Ltd., London, 1992. ISBN 1-85626-067-4. ???? pp.

Feldstein:1982:LSF

[FG82] Alan Feldstein and Richard Goodman. Loss of significance in floating point subtraction and addition. IEEE Trans. Comput., C-31(4):328–335, 1982.

Feuer:1984:CAP

[FG84] A. R. Feuer and N. Gehani, editors. Comparing and Assessing Programming Languages: Ada, C, and Pascal. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984. ???? pp.

Flury:1986:ASO

[FG86] B. N. Flury and W. Gautschi. An algorithm for simultaneous orthogonal transformation of several positive definite symmetric matrices to nearly diagonal form. SIAM J. Sci. Stat. Comput., 7:169–184, 1986.

Fletcher:1971:CLB

[FGH71] R. Fletcher, J. A. Grant, and M. D. Hebden. The calculation of linear best L_{p} approximations. The Computer Journal, 14:276–279, 1971. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic).

Freeman:1983:PSP

[FGHK83] Daniel H. Freeman, Jr., María Elena González, David C. Hoaglin, and Beth A. Kilss. Presenting statistical papers. The American Statistician, 37(2):106– 110, 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Forsgren:1995:CMN

[FGM95] Anders Forsgren, Philip E. Gill, and Walter Murray. Computing modified Newton directions using a partial Cholesky factorization. SIAM J. Sci. Comput., 16(1):139–150, 1995.

Feldman:1990:FCC

[FGMS90] S. I. Feldman, David M. Gay, Mark W. Maimone, and N. L. Schryer. A Fortran to C converter. Computing Science Technical Report No. 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990. 26 pp.

Freund:1992:ISL

[FGN92] R. W. Freund, G. H. Golub, and N. M. Nachtigal. Iterative solution of linear systems. In Acta Numerica, pages 57– 100. Cambridge University Press, 1992.

Ferng:1991:ALM

[FGP91] William R. Ferng, Gene H. Golub, and Robert J. Plemmons. Adaptive Lanczos methods for recursive condition estimation. Numerical Algorithms, 1(1):1–19, 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Forsgren:1995:SSI

[FGS95] Anders Forsgren, Philip E. Gill, and Joseph R. Shinnerl. Stability of symmetric ill-conditioned systems arising in interior methods for constrained optimization. SIAM J. Matrix Anal. Appl., 16(4): ??, 1995. To appear.

Fan:1955:SMI

[FH55] Ky Fan and A. J. Hoffman. Some metric inequalities in the space of matrices. Proc. Amer. Math. Soc., 6:111–116, 1955.

Forsythe:1960:CJM

[FH60] G. E. Forsythe and P. Henrici. The cyclic Jacobi method for computing the principal values of a complex matrix. Trans. Amer. Math. Soc., 94:1–23, 1960.

Fletcher:1990:TRL

[FH90] R. Fletcher and J. A. J. Hall. Towards reliable linear programming. In Griffiths and Watson [GW90], pages 89–104. to appear in Proc. 1989 Dundee Conference.

Flowe:1993:NGV

[FH93] Randolph P. Flowe and Gary A. Harris. A note on generalized Vandermonde determinants. SIAM J. Matrix Anal. Appl., 4(4):1146–1151, 1993.

Fox:1967:ABE

[FHM67] L. Fox, P. Henrici, and C. B. Moler. Approximations and bounds for eigenvalues of elliptic operators. SIAM J. Numer. Anal., 4:89–102, 1967.

Fox:1978:PMS

[FHS78] P. A. Fox, A. D. Hall, and N. L. Schryer. The PORT mathematical subroutine library. ACM Transactions on Mathematical Software, 4(2):104–126, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fox:1948:NSA

[FHW48] L. Fox, H. D. Huskey, and J. H.

Wilkinson. Notes on the solution of algebraic linear simultaneous equations. Quart. J. Mech. and Applied Math., 1: 149–173, 1948.

Fiedler:1986:SMT

[Fie86] Miroslav Fiedler. Special Matrices and Their Applications in Numerical Mathematics. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1986. ISBN 90-247-2957-2. xi + 307 pp.

Fike:1967:MEP

[Fik67] C. T. Fike. Methods of evaluating polynomial approximations in function evaluation routines. Comm. ACM, 10(3): 175–178, 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Filipponi:1981:ACF

[Fil81] P. Filipponi. An algorithm for computing functions of triangular matrices. Computing, 26:67–71, 1981. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Finney:1994:BLA

[Fin94] David J. Finney. On biometric language and its abuses. Biometric Bulletin, pages 2–4, November 1994. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Fischer:1974:FSC

[Fis74] Patrick C. Fischer. Further schemes for combining matrix algorithms. In Jacques Loeckx, editor, Automata, Languages and Programming, volume 14 of Lecture Notes in Computer Science, pages 428–436. Springer-Verlag, Berlin, 1974.

Fischer:1983:CBD

[Fis83] J. Fischer. The convergence of the best discrete linear L_{p} approximation as p → 1. J. Approximation Theory, 39:374–385, 1983.

Fitch:1993:MGA

[Fit93] John Fitch. Mathematics goes automatic. Physics world, 6(6):48–52, June 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Fairgrieve:1991:KFM

[FJ91] Thomas F. Fairgrieve and Allan D. Jepson. O.K. Floquet multipliers. SIAM J. Numer. Anal., 28(5):1446–1462, 1991.

Fox:1977:CCH

[FK77] Kenneth Fox and Burton J. Krohn. Computation of cubic harmonics. J. Comput. Phys., 25:386–408, 1977.

Faddeev:1968:SLAa

[FKF68a] D. K. Faddeev, V. N. Kublanovskaja, and V. N. Faddeeva. Solution of linear algebraic systems with rectangular matrices. Proc. Steklov Inst. Math., 96:93– 111, 1968. CODEN TMISAF. ISSN 0081-5438.

Faddeev:1968:SLAb

[FKF68b] D. K. Faddeev, V. N. Kublanovskaja, and V. N. Faddeeva. Sur les systémes linéaires algébriques de matrices rectangulaires et mal-conditionées. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7 of Éditions Centre Nat. Recherche Sci., Paris, pages 161–170, 1968.

Ford:1986:SA

[FKR86] B. Ford, J. Kok, and M. W. Rogers, editors. Scientific Ada. Cambridge University Press, ????, 1986. ISBN 0-521-33258-3. 388 pp.

Fenner:1974:SNB

[FL74] T. I. Fenner and G. Loizou. Some new bounds on the condition numbers of optimally scaled matrices. J. Assoc. Comput. Mach., 21(3):514–524, 1974.

Flanders:1971:MMA

[Fla71] Harley Flanders. Manual for Monthly authors. Amer. Math. Monthly, 78:1–10, 1971.

Flesch:1946:APT

[Fle46] Rudolf Flesch. The Art of Plain Talk.

Harper and Brothers, New York, 1946. ???? pp.

Fletcher:1965:FME

[Fle65] R. Fletcher. Function minimization without evaluating derivatives—A review. The Computer Journal, 8:33–41, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Fletcher:1976:FSI

[Fle76] R. Fletcher. Factorizing symmetric indefinite matrices. Linear Algebra and Appl., 14:257–272, 1976.

Fletcher:1980:PMO

[Fle80] R. Fletcher. Practical Methods of Optimization, Volume 1, Unconstrained Optimization. Wiley, Chichester, UK, 1980. ???? pp.

Fletcher:1981:NPP

[Fle81a] R. Fletcher. A nonlinear programming problem in statistics (educational testing). SIAM J. Sci. Stat. Comput., 2: 257–267, 1981.

Fletcher:1981:PMO

[Fle81b] R. Fletcher. Practical Methods of Optimization, Volume 2, Constrained Optimization. Wiley, Chichester, UK, 1981. ???? pp.

Fletcher:1984:SCF

[Fle84] R. Fletcher. The self consistent field problem. IMA Bulletin, 20:72–76, 1984.

Fletcher:1985:EC

[Fle85a] R. Fletcher. Expected conditioning. IMA J. Numer. Anal., 5:247–273, 1985.

Fletcher:1985:SDM

[Fle85b] R. Fletcher. Semi-definite matrix constraints in optimization. SIAM Journal on Control and Optimization, 23:493– 513, 1985.

Fletcher:1986:CEQ

[Fle86] R. Fletcher. Cancellation errors in quasi-Newton methods. SIAM J. Sci. Stat. Comput., 7(4):1387–1399, 1986.

Fletcher:1987:PMO

[Fle87] R. Fletcher. Practical Methods of Optimization. Wiley, Chichester, UK, second edition, 1987. ISBN 0-471-91547-5. xiv + 436 pp.

Fletcher:1988:DPR

[Fle88] R. Fletcher. Degeneracy in the presence of roundoff errors. Linear Algebra and Appl., 106:149–183, 1988.

Fletcher:1989:NVR

[Fle89] R. Fletcher. A new variational result for quasi-Newton formulae. Technical report, Report NA/119, Department of Mathematics and Computer Science, University of Dundee, ????, 1989. ???? pp.

Fletcher:1990:LSM

[Fle90] R. Fletcher. Low storage methods for unconstrained optimization. In Allgower and Georg [AG90], pages 165–179. ISBN 0-8218-1131-2. LCCN QA372 .C6374 1990.

Fletcher:1993:RDQ

[Fle93] R. Fletcher. Resolving degeneracy in quadratic programming. Annals of Operations Research, 47:307–334, 1993. CO-DEN AOREEV. ISSN 0254-5330 (print), 1572-9338 (electronic).

Fletcher:1994:SED

[Fle94] R. Fletcher. Steepest edge, degeneracy and conditioning in LP. Numerical Analysis Report NA/154, Department of Mathematics and Computer Science, University of Dundee, Dundee, Scotland, November 1994. 21 pp.

Forsythe:1952:EGT

[FM52] George E. Forsythe and Theodore S.

Motzkin. An extension of Gauss’ transformation for improving the condition of systems of linear equations. Mathematical Tables and Other Aids to Computation, 6 (37):9–17, 1952.

Forsythe:1967:CSL

[FM67] George E. Forsythe and Cleve B.

Moler. Computer Solution of Linear Algebraic Systems. Prentice-Hall, Englewood Cliffs, NJ, USA, 1967. xi + 148 pp.

Fiacco:1968:NPS

[FM68a] A. V. Fiacco and G. P. McCormick.

Nonlinear Programming: Sequential Unconstrained Minimization Techniques.

Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1968. ???? pp. reprinted in SIAM Classics in Applied Mathematics series, 1990.

Fox:1968:CMS

[FM68b] L. Fox and D. F. Mayers. Computing Methods for Scientists and Engineers. Oxford University Press, ????, 1968. xii + 255 pp.

Fletcher:1984:SME

[FM84] R. Fletcher and S. P. J. Matthews. Stable modification of explicit LU factors for simplex updates. Math. Prog., 30: 267–284, 1984. CODEN MHPGA4. ISSN 0025-5610.

Fletcher:1985:SAU

[FM85] R. Fletcher and S. P. J. Matthews. A stable algorithm for updating triangular factors under a rank one change. Math. Comp., 45(172):471–485, 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Funderlic:1986:SSD

[FM86] R. E. Funderlic and C. D. Meyer, Jr. Sensitivity of the stationary distribution vector for an ergodic Markov chain. Linear Algebra and Appl., 76:1–17, 1986.

Francois:1991:SPM

[FM91] Philippe François and Jean-Michel Muller. The SCALP perturbation method. In Proceedings of IMACS ’91, 13th World Congress on Computation and Applied Mathematics, Dublin, Volume 1, pages 59–60, 1991.

Forsythe:1977:CMM

[FMM77] George E. (George Elmer) Forsythe, Michael A. Malcolm, and Cleve B.

Moler. Computer Methods for Mathematical Computations. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1977. ISBN 0-13-165332-6. xi + 259 pp. LCCN QA297 .F5681. US$16.95. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from

math.liu.se in pub/references.

Friedland:1987:FAN

[FNO87] S. Friedland, J. Nocedal, and M. L.

Overton. The formulation and analysis of numerical methods for inverse eigenvalue problems. SIAM J. Numer. Anal., 24:634–667, 1987.

Funderlic:1982:DGD

[FNP82] R. E. Funderlic, M. Neumann, and R. J. Plemmons. LU decompositions of generalized diagonally dominant matrices. Numer. Math., 40:57–69, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forsythe:1951:GGR

[For51] George E. Forsythe. Gauss to Gerling on relaxation. Mathematical Tables and Other Aids to Computation, 5:255– 258, 1951.

Forsythe:1953:SLA

[For53] George E. Forsythe. Solving linear algebraic equations can be interesting. Bull. Amer. Math. Soc., 59(4):299–329, 1953.

Forsythe:1957:GUO

[For57a] G. E. Forsythe. Generation and use of orthogonal polynomials for data-fitting with a digital computer. J. Soc. Indust. Appl. Math., 5(2):74–88, 1957.

Forsythe:1957:SST

[For57b] G. E. Forsythe. Suggestions to students on talking about mathematics papers. Amer. Math. Monthly, 64:16–18, 1957. Reprinted in [Ewi94].

Forsythe:1959:RNR

[For59a] George E. Forsythe. Reprint of a note on rounding-off errors. SIAM Review, 1 (1):66–67, 1959.

Forsythe:1959:SNS

[For59b] George E. Forsythe. Singularity and near singularity in numerical analysis. Amer. Math. Monthly, 65(4):229–240, 1959.

Forsythe:1960:ACP

[For60] George E. Forsythe. Algorithm 16: Crout with pivoting. Comm. ACM, 3 (9):507–508, September 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1967:TCM

[For67] George E. Forsythe. Today’s computational methods of linear algebra. SIAM Review, 9(3):489–515, 1967.

Forsythe:1968:WDT

[For68] George E. Forsythe. What to do till the computer scientist comes. American Mathematical Monthly, 75(5):454– 462, May 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1969:SQE

[For69a] George E. Forsythe. Solving a quadratic equation on a computer. In National Research Council’s Committee on Support of Research in the Mathematical Sciences, editor, The Mathematical Sciences: A Collection of Essays, pages 138–152. MIT Press, Cambridge, MA, USA, 1969.

Forsythe:1969:WSQ

[For69b] George E. Forsythe. What is a satisfactory quadratic equation solver? In Bruno Dejon and Peter Henrici, editors, Constructive Aspects of the Fundamental Theorem of Algebra, pages 53–61. Wiley-Interscience, New York, NY, USA, 1969.

Forsythe:1970:PCW

[For70] George E. Forsythe. Pitfalls in computation, or why a math book isn’t enough. Amer. Math. Monthly, 77:931–956, 1970.

Ford:1978:PET

[For78] Brian Ford. Parameterization of the environment for transportable numerical software. ACM Transactions on Mathematical Software, 4(2):100–103, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Forscher:1980:RR

[For80] Bernard K. Forscher. The role of the referee. Scholarly Publishing, 11(2):165– 169, 1980. ISSN 0036-634X.

Forbes:1992:PCS

[For92] A. B. Forbes. Position calibration software for the form assessment of complex workpieces. NPL Report DITC 203/92, National Physical Laboratory, Teddington, UK, July 1992. ???? pp.

Forum:1993:HPF

[For93] High Performance Fortran Forum.

High performance Fortran language specification. January 25, 1993, Version 1.0 Draft, 1993.

Fosdick:1979:DEP

[Fos79a] L. D. Fosdick. Detecting errors in programs. In Fosdick [Fos79b], pages 77– 87. ISBN 0-444-85330-8.

Fosdick:1979:PEN

[Fos79b] Lloyd D. Fosdick, editor. Performance Evaluation of Numerical Software. North-Holland, Amsterdam, The Netherlands, 1979. ISBN 0-444-85330-8. xi + 335 pp.

Foster:1986:RNS

[Fos86] Leslie V. Foster. Rank and null space calculations using matrix decomposition without column interchanges. Linear Algebra and Appl., 74:47–71, 1986.

Foster:1990:PLD

[Fos90] L. V. Foster. The probability of large diagonal elements in the QR factorization. SIAM J. Sci. Stat. Comput., 11: 531–544, 1990.

Foster:1991:MNE

[Fos91] Leslie Foster. Modifications of the normal equations method that are numerically stable. In G. H. Golub and P. M. Van Dooren, editors, Numerical Linear Algebra, Digital Signal Processing and Parallel Algorithms, volume F70 of NATO ASI Series, pages 501–512. Springer-Verlag, Berlin, 1991.

Foster:1994:GEP

[Fos94] Leslie V. Foster. Gaussian elimination with partial pivoting can fail in practice. SIAM J. Matrix Anal. Appl., 15(4):1354– 1362, 1994.

Foulds:1981:OTI

[Fou81] L. R. Foulds. Optimization Techniques: An Introduction. Springer-Verlag, Berlin, 1981. ???? pp.

Fowler:1968:DME

[Fow68] H. W. Fowler. A Dictionary of Modern English Usage. Oxford University Press, ????, second edition, 1968. ???? pp. Revised by Sir Ernest Gowers.

Fox:1964:INL

[Fox64] L. Fox. An Introduction to Numerical Linear Algebra. Oxford University Press, ????, 1964. xi + 328 pp.

Fox:1969:ALA

[Fox69] B. L. Fox. Accelerating LP algorithms. Comm. ACM, 12(7):384–385, 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:1971:HGM

[Fox71] L. Fox. How to get meaningless answers in scientific computation (and what to do about it). IMA Bulletin, 7(10):296– 302, 1971.

Fox:1978:AAJ

[Fox78] L. Fox. All about Jim Wilkinson, with a commemorative snippet on backward error analysis. In The Contribution of Dr. J. H. Wilkinson to Numerical Analysis, Symposium Proceedings Series No. 19, pages 1–20. The Institute of Mathematics and its Applications, Southend-On-Sea, Essex, UK, 1978.

Fox:1987:JHW

[Fox87a] L. Fox. James Hardy Wilkinson, 1919–1986. Biographical Memoirs of Fellows of the Royal Society, 33:671–708, 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Foxley:1987:MLT

[Fox87b] E. Foxley. Music—A language for typesetting music scores. Software —Practice and Experience, 17:485–502, 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fox:1990:ENA

[Fox90a] L. Fox. Early numerical analysis in the United Kingdom. In Stephen G. Nash, editor, A History of Scientific Computing, pages 280–300. Addison-Wes-ley, Reading, MA, USA, 1990.

Fox:1990:EJW

[Fox90b] L. Fox. Epilogue. Jim Wilkinson: Some after-dinner sentiments. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 329– 335. Oxford University Press, 1990.

Fiedler:1962:MNP

[FP62] Miroslav Fiedler and Vlastimil Pták. On matrices with non-positive offdiagonal elements and positive principal minors. Czechoslovak Math. J., 12(87): 382–400, 1962. CODEN CZMJAE. ISSN 0528-9181.

Fletcher:1963:RCD

[FP63] R. Fletcher and M. J. D. Powell. A rapidly convergent descent method for minimization. The Computer Journal, 6: 163–168, 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Fox:1968:CPN

[FP68] L. Fox and I. B. Parker. Chebyshev Polynomials in Numerical Analysis. Oxford University Press, ????, 1968. ???? pp.

Fischer:1974:EPU

[FP74a] Patrick C. Fischer and Robert L.

Probert. Efficient procedures for using matrix algorithms. In Jacques Loeckx, editor, Automata, Languages and Programming, volume 14 of Lecture Notes in Computer Science, pages 413–427. Springer-Verlag, Berlin, 1974.

Fletcher:1974:MF

[FP74b] R. Fletcher and M. J. D. Powell.

On the modification of LDL^{T} factorizations. Math. Comp., 28(128):1067–1087, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Funderlic:1981:DME

[FP81] R. E. Funderlic and R. J. Plemmons. LU decomposition of M-matrices by elimination without pivoting. Linear Algebra and Appl., 41:99–110, 1981.

Funderlic:1986:ULF

[FP86a] R. E. Funderlic and R. J. Plemmons.

Updating LU factorizations for computing stationary distributions. SIAM J. Alg. Disc. Meth., 7:30–42, 1986.

Funderlic:1986:UFC

[FP86b] R. E. Funderlic and R. J. Plemmons.

Updating LU factorizations for computing stationary distributions. SIAM J. Alg. Disc. Meth., 7(1):30–42, 1986.

Fernando:1992:ASV

[FP92a] K. Vince Fernando and Beresford N.

Parlett. Accurate singular values and differential qd algorithms. Technical Report TR2/92, NAG Ltd., Oxford, July 1992. ???? pp. To appear in Numer. Math.

Freeman:1992:PNA

[FP92b] T. L. Freeman and C. Phillips. Parallel Numerical Algorithms. Prentice-Hall, Englewood Cliffs, NJ, USA, 1992. ???? pp.

Farouki:1987:NCB

[FR87] R. T. Farouki and V. T. Rajan. On the numerical condition of Bernstein polynomials. Technical report, Report RC 12626, Manufacturing Research Department, IBM Thomas J. Watson Research Center, Yorktown Heights, New York, ????, 1987. ???? pp.

Frank:1958:CEC

[Fra58] Werner L. Frank. Computing eigenvalues of complex matrices by determinant evaluation and by methods of Danilewski and Wielandt. J. Soc. Indust. Appl. Math., 6(4):378–392, 1958.

Franca:1989:ACS

[Fra89] L. P. Franca. An algorithm to compute the square root of a 3 × 3 positive definite matrix. Computers and Mathematics with Applications, 18(5):459–466, 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Franzblau:1992:USE

[Fra92a] Deborah S. Franzblau. An undergraduate seminar emphasizing oral presentation of research mathematics. PRIMUS (Problems Resources and Issues in Mathematics Undergraduate Studies), 2(1):16–32, 1992. ISSN 1051-1970 (print), 1935-4053 (electronic). Problems, Resources, and Issues in Mathematics Undergraduate Studies, Rose-Hulman Institute of Technology, Terre Haute, Indiana 47803.

Fraysse:1992:RCS

[Fra92b] V. Frayssé. Reliability of Computer Solutions. PhD thesis, L’Institut National Polytechnique de Toulouse, Toulouse, France, July 1992. ???? pp. CERFACS Report TH/PA/92/11.

Freeman:1979:MCA

[Fre79] T. L. Freeman. A method for computing all the zeros of a polynomial with real coefficients. BIT, 19:321–333, 1979. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Freeman:1980:SVS

[Fre80] T. L. Freeman. Solution of Vandermonde systems of equations: An alternative view. Numerical Analysis Report No. 45, University of Manchester, Manchester, England, 1980. ???? pp.

Frenkel:1987:PCB

[Fre87a] K. A. Frenkel. Profiles in computing:

Brian K. Reid. Comm. ACM, 30:820–823, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frenkel:1987:PCD

[Fre87b] K. A. Frenkel. Profiles in computing: Donald E. Knuth. Comm. ACM, 30: 816–819, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freund:1990:CGT

[Fre90] R. W. Freund. Conjugate gradient type methods for linear systems with complex symmetric coefficient matrices. Technical report, ????, 1990. ???? pp. to appear in SIAM J. Sci. Stat. Comput.

Freund:1993:TFQ

[Fre93] Roland W. Freund. A transpose-free quasi-minimal residual algorithm for non-Hermitian linear systems. SIAM J. Sci. Comput., 14(2):470–482, 1993.

Fried:1984:EPI

[Fri84] S. S. Fried. Evaluating 8087 performance on the IBM PC. BYTE Magazine, 9:197–208, 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fried:1985:PC

[Fri85] S. S. Fried. The 8087/80287 performance curve. Technical report, BYTE, ????, 1985. ???? pp.

Friedland:1991:RMS

[Fri91] Shmuel Friedland. Revisiting matrix squaring. Linear Algebra and Appl., 154-156:59–63, 1991.

Froeberg:1953:SLS

[Frö53] Carl-Erik Fröberg. Solutions of linear systems on a relay machine. In L. J. Paige and Olga Taussky, editors, Simultaneous Linear Equations and the Determination of Eigenvalues, number 29 in Applied Mathematics Series, pages 39– 42. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1953.

Froberg:1969:INA

[Fro69] C.-E. Froberg. Introduction to Numerical Analysis. Addison-Wesley, Reading, MA, USA, second edition edition, 1969. ???? pp.

Froeberg:1990:BCC

[Frö90] Carl-Erik Fröberg. BIT—A child of the computer. In Stephen G. Nash, editor, A History of Scientific Computing, pages 228–233. Addison-Wesley, Reading, MA, USA, 1990.

Forsythe:1955:BCM

[FS55] G. E. Forsythe and E. G. Straus. On best conditioned matrices. Proc. Amer. Math. Soc., 6:340–345, 1955.

Ford:1976:DSN

[FS76] B. Ford and D. K. Sayers. Developing a single numerical algorithms library for different machine ranges. ACM Transactions on Mathematical Software, 2(2): 115–131, 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Friedland:1980:GPN

[FS80] Shmuel Friedland and Hans Schneider.

The growth of powers of a nonnegative matrix. SIAM J. Alg. Disc. Meth., 1(2): 185–200, 1980.

Fletcher:1983:ADJ

[FS83] R. Fletcher and D. C. Sorensen. An algorithmic derivation of the Jordan canonical form. Amer. Math. Monthly, 90(1): 12–16, 1983.

Fritsch:1973:STE

[FSC73] F. N. Fritsch, R. E. Shafer, and W. P.

Crowley. Solution of the transcendental equation we^{w} = χ. Comm. ACM, 16(2): 123–124, 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldstein:1986:OUS

[FT86] Alan Feldstein and Peter Turner.

Overflow, underflow, and severe loss of significance in floating-point addition and subtraction. IMA J. Numer. Anal., 6: 241–251, 1986.

Fan:1988:FNR

[FT88] M. K. H. Fan and A. L. Tits. mform numerical range and the computation of the structured singular value.

IEEE Trans. Automat. Control., AC-33: 284–289, 1988.

Fan:19xx:MCD

[FTxx] M. K. H. Fan and A. L. Tits. A method for computing the distance of a stable matrix to the set of unstable matrices. Manuscript, 19xx.

FroeseFischer:1969:PST

[FU69] C. Froese Fischer and R. A. Usmani. Properties of some tridiagonal matrices and their applications to boundary value problems. SIAM J. Numer. Anal., 6:127– 142, 1969.

Fuller:1987:MEM

[Ful87] W. A. Fuller. Measurement Error Models. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1987. ???? pp.

Funaro:1988:CIC

[Fun88] D. Funaro. Computing the inverse of the Chebyshev collocation derivative. SIAM J. Sci. Stat. Comput., 9:1050–1057, 1988.

Feingold:1962:BDD

[FV62] David G. Feingold and Richard S. Varga. Block diagonally dominant matrices and generalizations of the Gerschgorin circle theorem. Pacific J. Math., 12:1241– 1250, 1962.

Foley:1990:CGP

[FvDFH90] James D. Foley, Andries van Dam, Steven K. Feiner, and John F. Hughes. Computer Graphics: Principles and Practice. Addison-Wesley, Reading, MA, USA, second edition, 1990. ???? pp.

Forsythe:1960:FDM

[FW60] George E. Forsythe and Wolfgang R.

Wasow. Finite-Difference Methods for Partial Differential Equations. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1960. x + 444 pp.

Gustafson:1994:MPS

[GA94] John L. Gustafson and Srinivas Aluru.

Massively parallel searching for better algorithms or, how to do a cross product with five multiplications. Scientific Programming, page ??, 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). Submitted.

Gabriel:1979:MMD

[Gab79] R. Gabriel. Matrizen mit maximaler diagonale bei unitärer Similarität. Journal für die reine und angewandte Mathematik, 307/308:31–52, 1979.

Gabriel:1987:NMC

[Gab87] R. Gabriel. The normal ΔHmatrices with connection to some Jacobilike methods. Linear Algebra and Appl., 91:181–194, 1987.

Gehringer:1988:SCP

[GAG88] E. F. Gehringer, J. Abullarade, and M. H. Gulyn. A survey of commercial parallel processors. ACM SIGARCH Computer Architecture News, 16:75–107, 1988. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Godunov:1993:GAN

[GAKK93] S. K. Godunov, A. G. Antonov, O. P. Kiriljuk, and V. I. Kostin. Guaranteed Accuracy in Numerical Linear Algebra. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-2352-1. 535 pp.

Galant:1971:ICT

[Gal71] David Galant. An implementation of Christoffel’s theorem in the theory of orthogonal polynomials. Math. Comp., 25 (113):111–113, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gallup:1991:I

[Gal91] Hutchinson Gallup. INFO 92. Hutchinson, London, 1991. ???? pp.

Gallian:1994:CAA

[Gal94] Joseph A. Gallian. Contemporary Abstract Algebra. D. C. Heath, Lexington, MA, USA, third edition, 1994. ISBN 0-669-33907-5. xix + 525 + 52 pages back matter pp.

Gantmacher:1959:TMa

[Gan59a] F. R. Gantmacher. The Theory of Matrices, volume one. Chelsea, New York, 1959. ISBN 0-8284-0131-4. x + 374 pp.

Gantmacher:1959:TMb

[Gan59b] F. R. Gantmacher. The Theory of Matrices, volume two. Chelsea, New York, 1959. ISBN 0-8284-0133-0. ix + 276 pp.

Gander:1985:HIM

[Gan85] W. Gander. On Halley’s iteration method. Amer. Math. Monthly, 92:131– 134, 1985.

Gander:1990:APD

[Gan90] Walter Gander. Algorithms for the polar decomposition. SIAM J. Sci. Stat. Comput., 11(6):1102–1115, 1990.

Gardner:1961:MMP

[Gar61] Martin Gardner. More Mathematical Puzzles and Diversions. Penguin, New York, 1961. ISBN 0-14-020748-1. 187 pp.

Garfield:1964:SCI

[Gar64] Eugene Garfield. “Science Citation Index”—A new dimension in indexing. Science, 144(3619):649–654, 1964. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Garfield:1972:CAT

[Gar72] Eugene Garfield. Citation analysis as a tool in journal evaluation. Science, 178: 471–479, 1972. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Garfield:1973:HCWa

[Gar73a] Eugene Garfield. Highly cited works in mathematics. Part 1. “Pure” mathematics. Current Contents: Physical, Chemical & Earth Sciences, (47):504– 508, November 21, 1973.

Garfield:1973:HCWb

[Gar73b] Eugene Garfield. Highly cited works in mathematics. Part 2. “Applied” mathematics. Current Contents: Physical, Chemical & Earth Sciences, (48):509– 513, November 28, 1973.

Garfield:1973:WSJ

[Gar73c] Eugene Garfield. What scientific journals can tell us about scientific journals. IEEE Trans. Prof. Commun., PC-16(4):200–203, 1973.

Gardner:1975:MC

[Gar75] Martin Gardner. Mathematical Carnival. Penguin, ????, 1975. ISBN 0-14-022041-0. x + 274 pp.

Garfield:1976:SJS

[Gar76a] Eugene Garfield. Significant journals of science. Nature, 264:609–615, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Garner:1976:SSR

[Gar76b] Harvey L. Garner. A survey of some recent contributions to computer arithmetic. IEEE Trans. Comput., C-25(12): 1277–1282, 1976.

Garfield:1977:HCA

[Gar77] Eugene Garfield. Highly cited articles. 36. Physics, chemistry and mathematics papers published in the 1940s. Current Contents: Physical, Chemical & Earth Sciences, (10):??, March 7, 1977.

Garfield:1979:CIT

[Gar79] Eugene Garfield. Citation Indexing: Its Theory and Application in Science, Technology, and Humanities. John Wiley, New York, NY, USA, 1979. ???? pp.

Garver:1980:PPP

[Gar80] Robert V. Garver. Presenting the peer paper. IEEE Trans. Prof. Commun., PC-23(1):18–22, 1980.

Garfield:1982:PMM

[Gar82a] Eugene Garfield. The 200 “pure” mathematicians most cited in 1978 and 1979, including a list of most-cited publications for the top 100. Current Contents: Physical, Chemical & Earth Sciences, (36):666–675, September 6, 1982.

Garfield:1982:JCS

[Gar82b] Eugene Garfield. Journal citation studies. 36. Pure and applied mathematics journals: What they cite and vice versa. Current Contents: Physical, Chemical & Earth Sciences, (5 (or 15??)): 5–10, March 8, 1982.

Garfield:1984:MCP

[Gar84] Eugene Garfield. The 100 most-cited papers ever and how we select citation classics. Current Contents: Physical, Chemical & Earth Sciences, (23):??, June 4, 1984.

Garfield:1985:ASO

[Gar85] Eugene Garfield. The Awards of Science and Other Essays. Essays of an Information Scientist: 1984. ISI Press, Philadelphia, 1985. ???? pp.

Garfield:1990:JEA

[Gar90a] Eugene Garfield. Journal editors awaken to the impact of citation errors. How we control them at ISI. Current Contents: Physical, Chemical & Earth Sciences, (41):5–13, October 8, 1990.

Garfield:1990:MCPa

[Gar90b] Eugene Garfield. The most-cited papers of all time, SCI 1945–1988. Part 1A. The SCI top 100—will the Lowry method ever be obliterated? Current Contents: Physical, Chemical & Earth Sciences, (7): 3–14, February 12, 1990.

Garfield:1990:MCPb

[Gar90c] Eugene Garfield. The most-cited papers of all time, SCI 1945–1988. Part 1B. Superstars new to the SCI top 100. Current Contents: Physical, Chemical & Earth Sciences, (8):3–13, February 19, 1990.

Garfield:1990:MCPc

[Gar90d] Eugene Garfield. The most-cited papers of all time, SCI 1945–1988. Part 3.

Another 100 from the Citation Classics Hall of Fame. Current Contents: Physical, Chemical & Earth Sciences, (34):3– 13, August 20, 1990.

Garfield:1990:MCPd

[Gar90e] Eugene Garfield. The most-cited physical-sciences publications in the 1945–1954 Science Citation Index. Part 2. Twenty Citation Classics in mathematics. Current Contents: Physical, Chemical & Earth Sciences, (42):376–381, October 15, 1990.

Garfield:1991:HUS

[Gar91a] Eugene Garfield. How to use the science citation index (SCI). In SCI Science Citation Index 1991. Guide and List of Source Publications, pages 26–33. ISI Press, Philadelphia, 1991. Reprinted from Current Contents 9, pp. 5–14, February 28 1983.

Garfield:1991:MCP

[Gar91b] Eugene Garfield. The most-cited papers of all time, SCI 1945–1988. Part 4. The papers ranked 301–400. Current Contents: Physical, Chemical & Earth Sciences, 31(21):5–16, 1991. No. 21, May 27.

Garfield:1994:IF

[Gar94a] Eugene Garfield. The impact factor. Current Contents: Physical, Chemical & Earth Sciences, 34(25):3–7, 1994. June 20.

Garfield:1994:UIF

[Gar94b] Eugene Garfield. Using the impact factor. Current Contents: Physical, Chemical & Earth Sciences, 34(29):3–5, 1994. July 18.

Gastinel:1983:LNA

[Gas83] Noel Gastinel. Linear Numerical Analysis. Kershaw Publishing, London, 1983. ISBN 0-901665-16-9. ix + 341 pp. First published in English by Academic Press, London, 1970.

Gautschi:1953:ABPa

[Gau53a] Walter Gautschi. The asymptotic behaviour of powers of matrices. Duke Math. J., 20:127–140, 1953. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Gautschi:1953:ABPb

[Gau53b] Walter Gautschi. The asymptotic behaviour of powers of matrices. II. Duke Math. J., 20:375–379, 1953. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Gautschi:1962:IVC

[Gau62] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. Numer. Math., 4:117–123, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1963:IVC

[Gau63] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. II. Numer. Math., 5:425–430, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1967:CAT

[Gau67] Walter Gautschi. Computational aspects of three-term recurrence relations. SIAM Review, 9(1):24–82, 1967.

Gautschi:1969:CMA

[Gau69] Walter Gautschi. On the condition of a matrix arising in the numerical inversion of the Laplace transform. Math. Comp., 23(105):109–118, 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1975:NEI

[Gau75a] Walter Gautschi. Norm estimates for inverses of Vandermonde matrices. Numer. Math., 23:337–347, 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1975:OCV

[Gau75b] Walter Gautschi. Optimally conditioned Vandermonde matrices. Numer. Math., 24:1–12, 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1978:IVC

[Gau78a] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices III. Numer. Math., 29:445–450, 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1978:QNC

[Gau78b] Walter Gautschi. Questions of numerical condition related to polynomials. pages 45–72, 1978.

Gautschi:1983:CVL

[Gau83a] Walter Gautschi. The condition of Vandermonde-like matrices involving orthogonal polynomials. Linear Algebra and Appl., 52/53:293–300, 1983.

Gautschi:1983:HHC

[Gau83b] Walter Gautschi. How and how not to check Gaussian quadrature formulae. BIT, 23:209–216, 1983. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gautschi:1990:HSV

[Gau90] Walter Gautschi. How (un)stable are Vandermonde systems? In R. Wong, editor, Asymptotic and Computational Analysis, volume 124 of Lecture Notes in Pure and Applied Mathematics, pages 193–210. Marcel Dekker, Inc., New York, NY, USA, 1990.

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gauss:1995:TCO

[Gau95] Carl Friedrich Gauss. Theory of the Combination of Observations Least Subject to Errors. Part One, Part Two, Supplement. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-347-1. xi + 241 pp. Translated from the Latin by G. W. Stewart.

Gay:1979:SCP

[Gay79] D. M. Gay. Some convergence properties of Broyden’s method. SIAM J. Numer. Anal., 16:623–630, 1979.

Gay:1984:TRA

[Gay84] D. M. Gay. A trust-region approach to linearly constrained optimization. In Griffiths [Gri84], pages 72–105.

Gay:1985:EMD

[Gay85] David M. Gay. Electronic mail distribution of linear programming test problems. Mathematical Programming Society COAL Newsletter, pages 10–12, December 1985.

Gay:1987:ULL

[Gay87] D. M. Gay. Using a large library on a small machine. In Wouk [Wou87], pages 80–91.

Gay:1988:ILS

[Gay88] David M. Gay. Interval least squares—A diagnostic tool. In Ramon E. Moore, editor, Reliability in Computing: The Role of Interval Methods in Scientific Computing, pages 183–205. Academic Press, Boston, MA, USA, 1988.

Gay:1990:CRB

[Gay90] David M. Gay. Correctly rounded binary-decimal and decimal-binary conversions. Numerical Analysis Manuscript 90-10, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1990. 16 pp.

Gay:1991:MMB

[Gay91] D. M. Gay. Massive memory buys little speed for complete, in-core sparse Cholesky factorizations on some scalar computers. Linear Algebra and Appl., 152:291–314, 1991.

Golomb:1963:SHM

[GB63] S. W. Golomb and L. D. Baumert.

The search for Hadamard matrices. Amer. Math. Monthly, 70:12–17, 1963.

Greenacre:1986:EAL

[GB86] Michael J. Greenacre and Michael W.

Browne. An efficient alternating leastsquares algorithm to perform multidimensional unfolding. Psychometrika, 51 (2):241–250, 1986. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Gover:1989:AMT

[GB89] M. J. C. Gover and S. Barnett. Applications of Matrix Theory. Oxford University Press, ????, 1989. ???? pp.

Gal:1991:AEM

[GB91] Shmuel Gal and Boris Bachelis. An accurate elementary mathematical library for the IEEE floating point standard. ACM Transactions on Mathematical Software, 17(1):26–45, 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Garbow:1977:MER

[GBDM77] B. S. Garbow, J. M. Boyle, J. J.

Dongarra, and C. B. Moler. Matrix Eigensystem Routines—EISPACK Guide Extension, volume 51 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1977. ???? pp.

Griewank:1991:ADA

[GC91] Andreas Griewank and George F.

Corliss, editors. Automatic Differentiation of Algorithms: Theory, Implementation, and Application. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-284-X.

Greenbaum:1989:EQQ

[GD89] A. Greenbaum and J. J. Dongarra. Experiments with QR/QL methods for the symmetric tridiagonal eigenproblem. Technical report, LAPACK Working Note #17, Department of Computer Science, University of Tennessee, Knoxville, ????, 1989. ???? pp.

Gu:1994:ECS

[GDD94] Ming Gu, James W. Demmel, and Inderjit Dhillon. Efficient computation of the singular value decomposition with applications to least squares problems. Technical Report CS-94-257, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October 1994. 19 pp. LAPACK Working Note 88.

Gu:1993:RPT

[GE93] Ming Gu and Stanley C. Eisenstat. Relative perturbation theory for eigenproblems. Research Report YALEU/DCS/RR-934, Department of Computer Science, Yale University, ????, February 1993. ???? pp.

Gu:1994:SEA

[GE94] Ming Gu and Stanley C. Eisenstat. A stable and efficient algorithm for the rank-one modification of the symmetric eigenproblem. SIAM J. Matrix Anal. Appl., 15(4):1266–1276, 1994.

Gu:1995:DCAa

[GE95a] Ming Gu and Stanley C. Eisenstat. A divide-and-conquer algorithm for the bidiagonal SVD. SIAM J. Matrix Anal. Appl., 16(1):79–92, 1995.

Gu:1995:DCAb

[GE95b] Ming Gu and Stanley C. Eisenstat. A divide-and-conquer algorithm for the symmetric tridiagonal eigenproblem. SIAM J. Matrix Anal. Appl., 16(1):172– 191, 1995.

Gu:1995:DSV

[GE95c] Ming Gu and Stanley C. Eisenstat.

Downdating the singular value decomposition. SIAM J. Matrix Anal. Appl., 16 (3):793–810, 1995.

Gear:1969:SST

[Gea69] C. W. Gear. A simple set of test matrices for eigenvalue programs. Math. Comp., 23(105):119–125, 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gear:1977:SCB

[Gea77] C. W. Gear. Simulation: Conflicts between real-time and software. In Rice [Ric77], page ?? ISBN 0-12-587260-7.

Gear:1981:NSO

[Gea81] C. W. Gear. Numerical solution of ordinary differential equations: Is there anything left to do? SIAM Review, 23: 10–24, 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gear:1991:ISM

[Gea91] C. William Gear. Inside SIAM’s mysterious journal publication process. SIAM News, 24(2):6, March 1991. ISSN 0036-1437.

Gehani:1989:AAI

[Geh89] Narain Gehani, editor. ADA: An Advanced Introduction. Prentice-Hall, Englewood Cliffs, NJ, USA, second edition, 1989. ISBN 0-13-004359-1. xvi + 280 pp.

Geist:1991:RGM

[Gei91] George A. Geist. Reduction of a general matrix to tridiagonal form. SIAM J. Matrix Anal. Appl., 12(2):362–373, 1991.

Guggenheimer:1995:SEC

[GEJ95] Heinrich W. Guggenheimer, Alan S. Edelman, and Charles R. Johnson. A simple estimate of the condition number of a linear system. College Mathematics Journal, 26(1):2–5, 1995.

Geman:1980:LTN

[Gem80] Stuart Geman. A limit theorem for the norm of random matrices. Ann. Probab., 8(2):252–261, April 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL http://projecteuclid.
org/euclid.aop/1176994775.

Geman:1986:SRL

[Gem86] Stuart Geman. The spectral radius of large random matrices. Ann. Probab., 14(4):1318–1328, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.
aop/1176992372.

Gentleman:1969:EAG

[Gen69] W. M. Gentleman. An error analysis of Goertzel’s (Watt’s) method for computing Fourier coefficients. The Computer Journal, 12:160–165, 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gentleman:1973:LSC

[Gen73] W. Morven Gentleman. Least squares computations by Givens transformations without square roots. J. Inst. Maths Applics, 12:329–336, 1973.

Gentleman:1974:RPQ

[Gen74] W. Morven Gentleman. Regression problems and the QR decomposition. IMA Bulletin, 10(5/6):195–197, 1974.

Gentleman:1975:EAD

[Gen75] W. Morven Gentleman. Error analysis of QR decompositions by Givens transformations. Linear Algebra and Appl., 10:189–197, 1975.

Gentleman:1978:SCR

[Gen78] W. M. Gentleman. Some complexity results for matrix computations on parallel processors. J. Assoc. Comput. Mach., 25:112–115, 1978.

George:1972:BES

[Geo72] A. George. On block elimination for sparse linear systems. SIAM J. Numer. Anal., 11:585–603, 1972.

George:1978:ARR

[Geo78] James E. George. Algorithms to reveal the representation of characters, integers, and floating-point numbers. ACM Transactions on Mathematical Software, 1(3):210–216, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Geurts:1982:CTC

[Geu82] A. J. Geurts. A contribution to the theory of condition. Numer. Math., 39:85–96, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gohberg:1974:CEP

[GF74] I. C. Gohberg and I. A. Feld́man. Convolution Equations and Projection Methods for Their Solution, volume 41 of Translations of Mathematical Monographs. American Mathematical Society, Providence, RI, USA, 1974. ISBN 0-8218-1591-1. ???? pp.

Goodman:1975:REP

[GF75] R. Goodman and A. Feldstein. Round-off error in products. Computing, 15:263–273, 1975. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Goodman:1985:REF

[GFB85] R. H. Goodman, A. Feldstein, and J. Bustoz. Relative error in floating-point multiplication. Computing, 35:127–139, 1985. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Gander:1989:CEP

[GGM89] W. Gander, G. H. Golub, and U. Von Matt. A constrained eigenvalue problem. Linear Algebra and Appl., 114/115:815–839, 1989.

Graves:1943:RYS

[GH43] Robert Graves and Alan Hodge. The Reader Over Your Shoulder: A Handbook for Writers of English Prose. Jonathan Cape, London, 1943. ???? pp.

Goldstine:1959:PDN

[GH59] H. H. Goldstine and L. P. Horwitz. A procedure for the diagonalization of normal matrices. J. Assoc. Comput. Mach., 6:176–195, 1959.

Graves:1970:UAE

[GH70] Robert Graves and Alan Hodge. The Use and Abuse of the English Language. Paragon House, New York, second edition, 1970. ISBN 1-55778-317-9. ???? pp. Originally entitled The Reader Over Your Shoulder. Paragon edition 1990.

Gander:1993:SPS

[GH93] Walter Gander and Jiřı Hřebıček.

Solving Problems in Scientific Computing Using Maple and Matlab. Springer-Ver-lag, Berlin, 1993. ISBN 3-540-57329-1. xiii + 268 pp.

George:1986:PCF

[GHL86] Alan George, Michael T. Heath, and Joseph Liu. Parallel Cholesky factorization on a shared-memory multiprocessor. Linear Algebra and Appl., 77:165– 187, 1986.

Gill:1986:UGL

[GHM^{+}86] P. E. Gill, S. J. Hammarling, W. Murray, M. A. Saunders, and M. H. Wright. User’s guide for LSSOL (version 1.0): A Fortran package for constrained linear least-squares and convex quadratic programming. Technical report, Technical Report SOL 86-1, Department of Operations Research, Stanford University, ????, 1986. ???? pp.

Goult:1974:CML

[GHMP74] R. J. Goult, R. F. Hoskins, J. A.

Milner, and M. J. Pratt. Computational Methods in Linear Algebra. Stanley Thornes, London, 1974. ISBN 0-85950-007-1. vii + 204 pp.

Gallivan:1990:PAM

[GHN^{+}90] K. A. Gallivan, M. T. Heath, E. Ng, J. M. Ortega, B. W. Peyton, R. J.

Plemmons, C. H. Romine, A. H. Sameh, and R. G. Voigt. Parallel Algorithms for Matrix Computations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ???? pp.

Giangardella:1967:SCV

[GHR67] J. J. Giangardella, J. F. Hudson, and R. S. Roper. Spelling correction by vector representation using a digital computer. IEEE Trans. Engineering Writing and Speech, EWS-10:57–62, 1967.

Glunt:1993:MCD

[GHR93] W. Glunt, T. L. Hayden, and M. Raydan. Molecular conformations from distance matrices. Journal of Computational Chemistry, 14(1):114– 120, January 1993. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Golub:1987:GEY

[GHS87] G. H. Golub, A. Hoffman, and G. W. Stewart. A generalization of the Eckart–Young–Mirsky matrix approximation theorem. Linear Algebra and Appl., 88/89:317–328, 1987.

Gautschi:1988:LBC

[GI88] Walter Gautschi and Gabriele Inglese. Lower bounds for the condition number of Vandermonde matrices. Numer. Math., 52:241–250, 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Giddins:1986:RNJ

[Gid86] Gary Giddins. Rhythm-a-ning: Jazz Tradition and Innovation in the ’80s. Oxford University Press, ????, 1986. ???? pp.

Giddins:1987:CB

[Gid87] Gary Giddins. Celebrating Bird.

Beech Tree Books, New York, 1987. ???? pp.

Gill:1951:PSS

[Gil51] S. Gill. A process for the step-by-step integration of differential equations in an automatic digital computing machine. Proc. Cambridge Phil. Soc., 47: 96–108, 1951. CODEN PCPSA4. ISSN 0008-1981.

Gill:1974:RDN

[Gil74] P. E. Gill. Recent developments in numerically stable methods for linear programming. IMA Bulletin, 10:180–186, 1974.

Gilvary:1986:EEM

[Gil86] B. Gilvary. Element by Element Methods for Heat Conduction Problems. Ph.d. thesis, University of Manchester, England, ????, 1986. ???? pp.

Gillman:1987:WMW

[Gil87] Leonard Gillman. Writing Mathematics Well: A Manual for Authors. Mathematical Association of America, Washington, DC, USA, 1987. ISBN 0-88385-443-0. ???? pp.

Gillespie:1991:NVT

[Gil91a] T. A. Gillespie. Noncommutative variations on theorems of Marcel Riesz and others. In John H. Ewing and F. W. Gehring, editors, Paul Halmos: Celebrating 50 Years of Mathematics, pages 221– 236. Springer-Verlag, Berlin, 1991.

Gillman:1991:PHE

[Gil91b] Leonard Gillman. Paul Halmos’s expository writing. In John H. Ewing and F. W. Gehring, editors, Paul Halmos: Celebrating 50 Years of Mathematics, pages 33–48. Springer-Verlag, Berlin, 1991.

Givens:1954:NCC

[Giv54] Wallace J. Givens. Numerical computation of the characteristic values of a real symmetric matrix. Technical Report ORNL-1574, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1954. 107 pp.

Givens:1958:CPU

[Giv58] Wallace Givens. Computation of plane unitary rotations transforming a general matrix to triangular form. J. Soc. Indust. Appl. Math., 6(1):26–50, 1958.

Gruenberger:1965:PCS

[GJ65] Fred Gruenberger and George Jaffray. Problems for Computer Solution. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1965. xvi + 401 pp.

Garey:1979:CIG

[GJ79] Michael R. Garey and David S. Johnson. Computers and Intractability: A Guide to the Theory of NP-Completeness. W.H. Freeman, San Francisco, CA, USA and New York, NY, USA, 1979. ISBN 0-7167-1045-5. ???? pp.

Gallivan:1987:UBL

[GJM87] Kyle Gallivan, William Jalby, and Ulrike Meier. The use of BLAS3 in linear algebra on a parallel processor with a hierarchical memory. SIAM J. Sci. Stat. Comput., 8(6):1079–1084, 1987.

Gallivan:1988:IHM

[GJMS88] Kyle Gallivan, William Jalby, Ulrike Meier, and Ahmed Sameh. Impact of hierarchical memory systems on linear algebra algorithm design. Int. J. Supercomputer Appl., 2(1):12–48, 1988. CO-DEN IJSAE9. ISSN 0890-2720.

Grone:1987:NM

[GJSW87] R. Grone, C. R. Johnson, E. M. Sá, and H. Wolkowicz. Normal matrices. Linear Algebra and Appl., 87:213–225, 1987.

Gantmakher:1937:MCN

[GK37] F. Gantmakher and M. Krein. Sur les matrices complètement non Négatives et oscillatoires. Compositio Mathematica, 4:445–476, 1937. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

Golub:1965:CSV

[GK65] G. H. Golub and W. Kahan. Calculating the singular values and pseudo-inverse of a matrix. SIAM J. Numer. Anal., 2: 205–224, 1965.

Gregory:1969:CMT

[GK69] Robert T. Gregory and David L. Karney. A Collection of Matrices for Testing Computational Algorithms. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1969. ISBN 0-88275-649-4. ix + 154 pp. Reprinted with corrections by Robert E. Krieger, Huntington, New York, 1978.

Golub:1983:CGQ

[GK83] G. H. Golub and J. Kautsky. Calculation of Gauss quadratures with multiple free and fixed knots. Numer. Math., 41:147–163, 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gohberg:1989:EAC

[GK89] I. Gohberg and I. Koltracht. Error analysis for Cauchy and Vandermonde matrices. 1989.

Gohberg:1990:EAT

[GK90a] I. Gohberg and I. Koltracht. Error analysis for triangular factorization of Cauchy and Vandermonde matrices. Manuscript, 1990.

Gohberg:1990:SPS

[GK90b] I. Gohberg and I. Koltracht. On the inversion of Cauchy matrices. In M. A. Kaashoek, J. H. van Schuppen, and A. C. M. Ran, editors, Signal Processing, Scattering and Operator Theory, and Numerical Methods. Proceedings of the International Symposium MTNS-89, Volume III, pages 381–392. Birkhäuser, Boston, MA, USA, 1990.

Gohberg:1992:SCN

[GK92] I. Gohberg and I. Koltracht. Structured condition numbers for linear matrix structures. IMA Preprint Series # 1030, Institute for Mathematics and its Applications, University of Minnesota, Minneapolis, MN, USA, September 1992. ???? pp.

Gohberg:1993:CNF

[GK93a] I. Gohberg and I. Koltracht. Condition numbers for functions of matrices. Applied Numerical Mathematics, 12:107– 117, 1993.

Gohberg:1993:MCS

[GK93b] I. Gohberg and I. Koltracht. Mixed, componentwise, and structured condition numbers. SIAM J. Matrix Anal. Appl., 14(3):688–704, 1993.

Godunov:1990:SPM

[GKK90] S. K. Godunov, O. P. Kirilyuk, and V. I. Kostin. Spectral portraits of matrices. Preprint 3, Inst. of Math., Sib. Branch of USSR Acad. Sci., 1990. ???? pp. In Russian. Cited in [Trexx].

Gudmundsson:1995:SSS

[GKL95] Thorkell Gudmundsson, Charles S.

Kenney, and Alan J. Laub. Small-sample statistical estimates for matrix norms. SIAM J. Matrix Anal. Appl., 16(3):776– 792, 1995.

Gohberg:1994:GEP

[GKO94] I. Gohberg, T. Kailath, and V. Olshevsky. Gaussian elimination with partial pivoting for structured matrices. 1994.

Graham:1989:CM

[GKP89] Ronald L. Graham, Donald E.

Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14236-8. xiii + 625 pp. LCCN QA39.2 .G7331 1989.

Golub:1976:RDL

[GKS76] Gene H. Golub, Virginia Klema, and G. W. Stewart. Rank degeneracy and least squares problems. Technical Report TR-456, Department of Computer Science, University of Maryland, College Park, June 1976. ???? pp.

Gries:1980:CFN

[GL80] David Gries and Gary Levin. Computing Fibonacci numbers (and similarly defined functions) in log time. Information Processing Letters, 11(2):68–69, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

George:1981:CSL

[GL81a] Alan George and Joseph W-H Liu. Computer Solution of Large Sparse Positive Definite Systems. Prentice-Hall, Englewood Cliffs, NJ, USA, 1981. ISBN 0-13-165274-5. xii + 324 pp.

Grimes:1981:CNE

[GL81b] Roger G. Grimes and John G. Lewis. Condition number estimation for sparse matrices. SIAM J. Sci. Stat. Comput., 2 (4):384–388, 1981.

Gardiner:1986:GMS

[GL86] Judith D. Gardiner and Alan J. Laub. A generalization of the matrixsign-function solution for algebraic Riccati equations. Internat. J. Control, 44 (3):823–832, 1986.

Gardiner:1988:SAR

[GL88] Judith D. Gardiner and Alan J. Laub. Solving the algebraic Riccati equation on a hypercube multiprocessor. In G. Fox, editor, Hypercube Concurrent Computers and Applications, Vol. II, pages 1562– 1568. ACM Press, New York, NY, USA, 1988.

Gardiner:1991:PAA

[GL91a] Judith D. Gardiner and Alan J. Laub. Parallel algorithms for algebraic Riccati equations. Internat. J. Control, 54(6):1317–1333, 1991.

Goldstein:1991:LAQ

[GL91b] Jerome A. Goldstein and Mel Levy. Linear algebra and quantum chemistry. Amer. Math. Monthly, 98(8):710–718,

Galligani:1994:EAN

[GL94] E. Galligani and A. Laratta. Error analysis of null space algorithm for linear equality constrained least squares problems. Computing, 52:161–176, 1994. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Ghavimi:1995:BES

[GL95a] Ali R. Ghavimi and Alan J. Laub.

Backward error, sensitivity, and refinement of computed solutions of algebraic Riccati equations. Numerical Linear Algebra with Applications, 2(1):29–49, 1995.

Ghavimi:1995:CAN

[GL95b] Ali R. Ghavimi and Alan J. Laub.

Computation of approximate null vectors of Sylvester and Lyapunov operators. IEEE Trans. Automat. Control., 40(2): 387–391, 1995.

Ghavimi:1995:RBD

[GL95c] Ali R. Ghavimi and Alan J. Laub.

Residual bounds for discrete-time Lyapunov equations. IEEE Transactions on Automatic Control, 40(7):1244–1249, 1995.

Gladwell:1982:POD

[Gla82] I. Gladwell, editor. Proceedings of a one-day colloquium on numerical linear algebra and its applications. University of Manchester, Manchester, England, July 1982. Numerical Analysis Report No. 78.

Glassey:1993:NCU

[Gla93] Robert Glassey. Numerical Computation Using C. Academic Press, New York, USA, 1993. ISBN 0-12-286155-8. vii + 283 pp.

Glanz:1995:MLF

[Gla95] James Glanz. Mathematical logic flushes out the bugs in chip designs. Science, 267:332–333, 1995. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). 20 January.

Gardiner:1992:SSM

[GLAM92] Judith D. Gardiner, Alan J. Laub, James J. Amato, and Cleve B. Moler. Solution of the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 18(2): 223–231, 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gleser:1986:SNR

[Gle86] Leon J. Gleser. Some notes on refereeing. The American Statistician, 40(4): 310–312, 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gliedman:1989:TTU

[Gli89] John Gliedman. Tips and Techniques for Using Low-Cost and Public Domain Software. McGraw-Hill, New York, 1989. ???? pp.

Gahinet:1990:SSD

[GLKH90] Pascal M. Gahinet, Alan J. Laub, Charles S. Kenney, and Gary A. Hewer. Sensitivity of the stable discrete-time Lyapunov equation. IEEE Trans. Automat. Control., AC-35(11):1209–1217, 1990.

Golub:1981:BLM

[GLO81] G. H. Golub, F. T. Luk, and M. L. Overton. A block Lanczos method for computing the singular values and corresponding singular vectors of a matrix.

ACM Transactions on Mathematical Software, 7:149–169, 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Geist:1989:SGR

[GLW89] G. A. Geist, A. Lu, and E. L.

Wachpress. Stabilized Gaussian reduction of an arbitrary matrix to tridiagonal form. Technical report, Report ORNL/TM-11089, Oak Ridge National Laboratory, Tennessee, ????, 1989. ???? pp.

Gill:1973:NSF

[GM73] P. E. Gill and W. Murray. A numerically stable form of the simplex algorithm. Linear Algebra and Appl., 7:99– 138, 1973.

Gentleman:1974:MAR

[GM74a] W. Morven Gentleman and Scott B.

Marovich. More on algorithms that reveal properties of floating point arithmetic units. Comm. ACM, 17(5):276–277, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gill:1974:NMC

[GM74b] P. E. Gill and W. Murray, editors.

Numerical Methods for Constrained Optimization. Academic Press, New York, USA, 1974.

Gill:1974:NTM

[GM74c] Philip E. Gill and Walter Murray.

Newton-type methods for unconstrained and linearly constrained optimization. Math. Prog., 7:311–350, 1974. CODEN MHPGA4. ISSN 0025-5610.

Goodman:1977:TNE

[GM77] J. T. Goodman and C. B. Moler.

Three numerical experiments with Gaussian elimination. Technical report, LIN-PACK Working Note #8, Technical Memorandum No. 311, Applied Mathematics Division, Argonne National Laboratory, ????, 1977. ???? pp.

Golub:1986:UQF

[GM86] G. H. Golub and C. D. Meyer, Jr. Using the QR factorization and group inversion to compute, differentiate, and estimate the sensitivity of stationary probabilities for Markov chains. SIAM J. Alg. Disc. Meth., 7:273–281, 1986.

Gill:1992:PIS

[GMPS92] Philip E. Gill, Walter Murray, Dulce B. Ponceleón, and Michael A. Saunders. Preconditioners for indefinite systems arising in optimization. SIAM J. Matrix Anal. Appl., 13(1):292–311, 1992.

Gill:1985:PRB

[GMS^{+}85] P. E. Gill, W. Murray, M. A. Saunders, G. W. Stewart, and M. H. Wright. Properties of a representation of a basis for the null space. Math. Prog., 33: 172–186, 1985. CODEN MHPGA4. ISSN 0025-5610.

Gill:1986:PNB

[GMS^{+}86] P. E. Gill, W. Murray, M. A. Saunders, J. A. Tomlin, and M. H. Wright. On projected Newton barrier methods for linear programming and an equivalence to Karmarkar’s projective method. Math. Prog., 36:183–209, 1986. CODEN MH-PGA4. ISSN 0025-5610.

Gill:1988:IPM

[GMS88] P. E. Gill, W. Murray, and M. A. Saunders. Interior-point methods for linear programming: A challenge to the simplex method. Technical report, Technical Report SOL 88-14, Department of Operations Research, Stanford University, ????, 1988. ???? pp.

Gilbert:1992:SMM

[GMS92] J. R. Gilbert, C. B. Moler, and R. S.

Schreiber. Sparse matrices in MATLAB: Design and implementation. SIAM J. Matrix Anal. Appl., 13(1):333–356, 1992.

Goossens:1993:LC

[GMS93] Michael Goossens, Frank Mittelbach, and Alexander Samarin. The LaTEX Companion. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-54199-8. xxx + 528 pp.

Gill:1984:SMM

[GMSW84] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Sparse matrix methods in optimization. SIAM J. Sci. Stat. Comput., 5 (3):562–589, 1984.

Gill:1986:CNA

[GMSW86a] P. E. Gill, W. Murray, M. A.

Saunders, and M. H. Wright. Considerations of numerical analysis in a sequential quadratic programming method. In Hennart [Hen86], pages 46–62.

Gill:1986:NNA

[GMSW86b] P. E. Gill, W. Murray, M. A.

Saunders, and M. H. Wright. A note on nonlinear approaches to linear programming. Technical report, Technical Report SOL 86-7, Department of Operations Research, Stanford University, ????, 1986. ???? pp.

Gill:1986:UGN

[GMSW86c] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H.

Wright. User’s guide for NPSOL (version 4.0): A Fortran package for nonlinear programming. Technical Report SOL 86-2, Department of Operations Research, Stanford University, Stanford, CA, USA, January 1986. 53 pp.

Gill:1987:CNP

[GMSW87a] P. E. Gill, W. Murray, M. A. Saunders, and M. H. Wright. Constrained nonlinear programming. Technical report, Technical Report SOL 87-13, Department of Operations Research, Stanford University, ????, 1987. ???? pp.

Gill:1987:MFG

[GMSW87b] P. E. Gill, W. Murray, M. A. Saunders, and M. H. Wright. Maintaining LU factors of a general sparse matrix. Linear Algebra and Appl., 88/89:239–270, 1987.

Gill:1988:RDC

[GMSW88] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Recent developments in constrained optimzation. J. Comp. Appl. Math., 22:257–270, 1988.

Gill:1990:SCM

[GMSW90] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. A Schur-complement method for sparse quadratic programming. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 113– 138. Oxford University Press, 1990.

Gill:1991:ICM

[GMSW91] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Inertia-controlling methods for general quadratic programming. SIAM Review, 33(1):1–36, 1991.

Gill:1981:PO

[GMW81] Philip E. Gill, Walter Murray, and Margaret H. Wright. Practical Optimization. Academic Press, London, 1981. ISBN 0-12-283952-8. xvi + 401 pp.

Gill:1991:NLA

[GMW91] Philip E. Gill, Walter Murray, and Margaret H. Wright. Numerical Linear Algebra and Optimization, volume 1. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-12649-4. xvii + 426 pp.

Gunzburger:1982:SGE

[GN82] M. D. Gunzburger and R. A. Nicolaides. Stability of Gaussian elimination without pivoting on tridiagonal Toeplitz matrices. Linear Algebra and Appl., 45: 21–28, 1982.

Golub:1979:HSM

[GNV79] G. H. Golub, S. Nash, and C. F.

Van Loan. A Hessenberg–Schur method for the problem AX + XB = C. IEEE Trans. Automat. Control., AC-24(6):909– 913, 1979.

Golub:1978:NAP

[GO78] G. H. Golub and J. Oliger. Numerical analysis: Proceedings of symposia in applied mathematics. Technical report, Vol. 22, American Mathematical Society, Providence, RI, ????, 1978. ???? pp.

Golub:1988:CIC

[GO88] G. H. Golub and M. L. Overton.

The convergence of inexact Chebyshev and Richardson iterative methods for solving linear systems. Numer. Math., 53:571–593, 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1989:SHC

[GO89] Gene H. Golub and Dianne P. O’Leary. Some history of the conjugate gradient and Lanczos algorithms: 1948– 1976. SIAM Review, 31(1):50–102, 1989.

Gohberg:1994:FIC

[GO94] I. Gohberg and V. Olshevsky. Fast inversion of Chebyshev–Vandermonde matrices. Numer. Math., 67(1):71–92, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1996:SCD

[GO96] Gene Howard Golub and James M. Ortega. Scientific Computing and Differential Equations: an Introduction to Numerical Methods. Academic Press, New York, NY, USA, fourth edition, 1996. ISBN 0-12-289255-0. xi + 337 pp. LCCN

                                        QA371 1996. URL http://www.gbv.de/
                                        dms/bowker/toc/9780122892554.pdf.

Goedecker:1994:RAF

[Goe94] S. Goedecker. Remark on algorithms to find roots of polynomials. SIAM J. Sci. Comput., 15(5):1059–1063, 1994.

Golub:1962:BRE

[Gol62] Gene H. Golub. Bounds for the roundoff errors in the Richardson second order method. BIT, 2:212–223, 1962. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Golub:1965:NMS

[Gol65] G. H. Golub. Numerical methods for solving linear least squares problems. Numer. Math., 7:206–216, 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldberg:1967:BED

[Gol67] I. Bennett Goldberg. 27 bits are not enough for 8-digit accuracy. Comm. ACM, 10(2):105–106, February 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Golub:1968:LSS

[Gol68] G. H. Golub. Least squares, singular values and matrix approximations. Aplikace Matematiky, 13:44–51, 1968.

Golub:1969:MDS

[Gol69] Gene H. Golub. Matrix decompositions and statistical calculations. In Roy C. Milton and John A. Nelder, editors, Statistical Computation, pages 365– 397. Academic Press, New York, 1969.

Goldstine:1972:CPN

[Gol72] Herman H. Goldstine. The Computer:

From Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-02367-0. xii + 378 pp. 1993 printing with new preface.

Golub:1973:SMM

[Gol73a] G. H. Golub. Some modified matrix eigenvalue problems. SIAM Review, 15: 318–334, 1973.

Golub:1973:TNA

[Gol73b] G. H. Golub. Some uses of the Lanczos algorithm in numerical linear algebra. In J. J. H. Miller, editor, Topics in Numerical Analysis, pages 173–184. Academic Press, New York, USA, 1973.

Goldstine:1977:HNA

[Gol77] Herman H. Goldstine. A History of Numerical Analysis From the 16th Through the 19th Century. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-90277-5. xiv + 348 pp.

Goldstein:1983:MBP

[Gol83] Rebecca Goldstein. The Mind Body Problem. Laurel (Dell Publishing), New York, 1983. ???? pp.

Golub:1984:SNA

[Gol84] G. H. Golub, editor. Studies in Numerical Analysis, volume 24 of Studies in Mathematics. Mathematical Association of America, Washington, DC, 1984. ???? pp.

Golub:1990:PRJ

[Gol90] G. H. Golub. Prologue: Reflections on Jim Wilkinson. In Cox and Hammarling [CH90c], pages 1–5. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$45.00.

Goldberg:1991:WEC

[Gol91a] David Goldberg. What every computer scientist should know about floating-point arithmetic. ACM Computing Surveys, 23(1):5–48, 1991. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Goldberg:1991:MTA

[Gol91b] Jack L. Goldberg. Matrix Theory With Applications. McGraw-Hill, New York, 1991. ISBN 0-07-557200-1. ???? pp.

Goodnight:1979:TSO

[Goo79] James H. Goodnight. A tutorial on the SWEEP operator. The American Statistician, 33(3):149–158, 1979. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Good:1992:IRA

[Goo92] I. J. Good. Introductory remarks for the article in Biometrika 66 (1979), “A. M. turing’s statistical work in World War II”. In J. L. Britton, editor, Collected Works of A. M. Turing: Pure Mathematics, pages 211–223. North-Holland, Amsterdam, The Netherlands, 1992.

Gordon:1984:TVH

[Gor84] Karen Elizabeth Gordon. The Transitive Vampire: A Handbook of Grammar for the Innocent, the Eager, and the Doomed. Times Books, New York, revised and expanded edition, 1984. ISBN 0-8129-1101-6. ???? pp.

Gordon:1993:WOW

[Gor93a] Ian Gordon. Wobbling one’s way along in words. New scientist, pages 47– 48, 1993. CODEN NWSCAL. ISSN 0262-4079, 0028-6664. 1 May.

Gordon:1993:NWT

[Gor93b] Karen Elizabeth Gordon. The New Well-Tempered Sentence: A Punctuation Handbook for the Innocent, the Eager, and the Doomed. Ticknor and Fields, New York, revised and expanded edition, 1993. ISBN 0-395-62883-0. ???? pp.

Gottlieb:1992:APS

[Got92] Allan Gottlieb. Architectures for parallel supercomputing. Technical report, Ultracomputer Research Laboratory, New York University, ????, December 1992. ???? pp.

Gould:1972:OP

[Gou72] Calvin R. Gould. The overhead projector. IEEE Trans. Prof. Commun., PC-15(1):2–6, 1972.

Gould:1973:VAH

[Gou73] Calvin R. Gould. Visual aids—how to make them positively legible. IEEE Trans. Prof. Commun., PC-16(2):35–38, 1973.

Gould:1986:ADS

[Gou86a] N. I. M. Gould. On the accurate determination of search directions for simple differentiable penalty functions. IMA J. Numer. Anal., 6:357–372, 1986.

Gourlay:1986:LMP

[Gou86b] J. S. Gourlay. A language for music printing. Comm. ACM, 29:388–401, 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gould:1991:GGE

[Gou91a] N. I. M. Gould. On growth in Gaussian elimination with complete pivoting. SIAM J. Matrix Anal. Appl., 12(2):354– 361, 1991.

Gouvea:1991:LAS

[Gou91b] F. Q. Gouvêa. A look at pari and gp. Notices Amer. Math. Soc., 38:903– 904, 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gower:1958:NIM

[Gow58] J. C. Gower. A note on an iterative method for root extraction. The Computer Journal, 1:142–143, 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gowers:1973:CPW

[Gow73] Sir Ernest Gowers. The Complete Plain Words. Penguin, London, second edition, 1973. ???? pp. Revised by Sir Bruce Fraser.

Gower:1975:GPA

[Gow75] J. C. Gower. Generalized Procrustes analysis. Psychometrika, 40(1):33–51, 1975. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Gower:1984:MAO

[Gow84] J. C. Gower. Multivariate analysis: Ordination, multidimensional scaling and allied topics. In E. H. Lloyd, editor, Statistics, volume VI of Handbook of Applicable Mathematics, pages 727–781. Wiley, Chichester, UK, 1984.

Gower:1985:DSC

[Gow85] J. C. Gower. The development of statistical computing at rothamsted. Technical report, Rothamsted Report for 1985, Part 2, ????, 1985. ???? pp.

Gowers:1986:CPW

[Gow86] Sir Ernest Gowers. The Complete Plain Words. Penguin, London, third edition, 1986. ISBN 0-14-051199-7. ???? pp. Revised by Sidney Greenbaum and Janet Whitcut.

Godfrey:1959:TWA

[GP59] J. W. Godfrey and G. Parr. The Technical Writer: An Aid to the Presentation and Production of Technical Literature. Chapman and Hall, Ltd., London, UK, 1959. ???? pp.

Galimberti:1970:NDS

[GP70] G. Galimberti and V. Pereyra. Numerical differentiation and the solution of multidimensional Vandermonde systems. Math. Comp., 24(110):357–364, 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Galimberti:1971:SCV

[GP71] G. Galimberti and V. Pereyra. Solving confluent Vandermonde systems of Hermite type. Numer. Math., 18:44–60, 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gibbs:1974:TP

[GP74] N. E. Gibbs and W. G. Poole, Jr. Tridiagonalization by permutations. Comm. ACM, 17:20–24, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Govaerts:1990:BEO

[GP90] W. Govaerts and J. D. Pryce. Block elimination with one iterative refinement solves bordered linear systems accurately. BIT, 30:490–507, 1990. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gasca:1992:TPN

[GP92] M. Gasca and J. M. Peña. Total positivity and Neville elimination. Linear Algebra and Appl., 165:25–44, 1992.

Gasca:1993:TPF

[GP93] M. Gasca and J. M. Peña. Total positivity, QR factorization, and Neville elimination. SIAM J. Matrix Anal. Appl., 4 (4):1132–1140, 1993.

Gallivan:1990:PAD

[GPS90] K. A. Gallivan, R. J. Plemmons, and A. H. Sameh. Parallel algorithms for dense linear algebra computations. SIAM Review, 32:54–135, 1990.

Golub:1970:SVD

[GR70] G. H. Golub and C. Reinsch. Singular value decomposition and least squares solutions. Numer. Math., 14:403–420, 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenspan:1988:NM

[GR88] D. Greenspan and Rózsa. Numerical Methods. Colloquia Mathematica Societatis János Bolyai 50, Miskolc, 1986. North-Holland, Amsterdam, The Netherlands, 1988. ???? pp.

Greenbaum:1989:OPG

[GR89] A. Greenbaum and G. Rodrigue. Optimal preconditioners of a given sparsity pattern. BIT, 29:610–634, 1989. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gragg:1990:DCM

[GR90] W. B. Gragg and L. Reichel. A divide and conquer method for unitary and orthogonal eigenproblems. Numer. Math., 57:695–718, 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Graybill:1969:IMA

[Gra69] F. A. Graybill. Introduction to Matrices with Applications in Statistics. Wadsworth Pub. Co., Pacific Grove, CA, USA, 1969. ???? pp.

Gray:1970:YWT

[Gra70] Dwight E. Gray. So You Have to Write a Technical Report. Information Resources Press, Washington, DC, 1970. ???? pp.

Grad:1971:MB

[Gra71] J. Grad. Matrix balancing. The Computer Journal, page ??, 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gray:1983:TT

[Gra83] Bill Gray. Tips on Type. Van Nostrand Reinhold Co., New York, NY, USA, 1983. ???? pp.

Graetzer:1985:MTS

[Grä85] George Grätzer. Math into TEX: A Simple Introduction to AMS-LaTEX. Birkhäuser Boston Inc., Cambridge, MA, USA, 1985. ISBN 0-8176-3637-4. xxix + 294 pp.

Gragg:1986:QAU

[Gra86] W. B. Gragg. The QR algorithm for unitary Hessenberg matrices. J. Comp. Appl. Math., 16:1–8, 1986.

Graham:1987:NMA

[Gra87] A. Graham. Nonnegative Matrices and Applicable Topics in Linear Algebra. Ellis Horwood, New York, NY, USA, 1987. ???? pp.

Grandine:1990:NMP

[Gra90] T. A. Grandine. The Numerical Methods Programming Projects Book. Oxford University Press, ????, 1990. ???? pp.

Granville:1992:ZBB

[Gra92] Andrew Granville. Zaphod Beeblebrox’s brain and the fifty-ninth row of Pascal’s triangle. Amer. Math. Monthly, 99:318–331, 1992.

Grassmann:1993:REC

[Gra93a] Winfried K. Grassmann. Rounding errors in certain algorithms involving Markov chains. ACM Transactions on Mathematical Software, 19(4):496– 508, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Graetzer:1993:A

[Grä93b] George Grätzer. AMS-LaTEX. Notices Amer. Math. Soc., 40(2):148–150, 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Graetzer:1994:ATI

[Grä94] George Grätzer. Advances in TEX:

IV. Header and footer control in LaTEX. Notices Amer. Math. Soc., 41(7):772–777, 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grcar:1989:OCM

[Grc89] J. F. Grcar. Operator coefficient methods for linear equations. Report SAND89-8691, Sandia National Laboratories, Albuquerque, New Mexico, November 1989. ???? pp.

Grcar:1990:MSL

[Grc90] J. F. Grcar. Matrix stretching for linear equations. Report SAND90-8723, Sandia National Laboratories, Albuquerque, New Mexico, 1990. ???? pp.

Grcar:1992:CWE

[Grc92] J. F. Grcar. Component-wise error bounds for linear equations. Report SAND92-8548, Sandia National Laboratories, Albuquerque, New Mexico, June 1992. ???? pp.

Green:1952:OAO

[Gre52] Bert F. Green. The orthogonal approximation of an oblique structure in factor analysis. Psychometrika, 17(4): 429–440, 1952. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Greenspan:1955:MMI

[Gre55] Donald Greenspan. Methods of matrix inversion. Amer. Math. Monthly, 62 (5):303–318, 1955.

Gregory:1972:CFP

[Gre72] James Gregory. A comparison of floating point summation methods. Comm. ACM, 15(9):838, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greaser:1979:IER

[Gre79] Constance U. Greaser. Improving the effectiveness of research writing. Scholarly Publishing, 11(1):61–71, 1979. ISSN 0036-634X.

Greig:1980:O

[Gre80] D. M. Greig. Optimisation. Longman, London, 1980. ???? pp.

Greenstadt:1985:QQN

[Gre85] J. Greenstadt. A quasi-quasi-Newton method for generating quasi-Choleski factors. SIAM J. Sci. Stat. Comput., 6:425– 437, 1985.

Green:1987:PS

[Gre87] J. Green. Putting on the style. Personal computer world, 10:144–148, 1987.

Greenbaum:1989:BSP

[Gre89] A. Greenbaum. Behavior of slightly perturbed Lanczos and conjugategradient recurrences. Linear Algebra and Appl., 113:7–63, 1989.

Gregory:1992:IPP

[Gre92] Martin W. Gregory. The infectiousness of pompous prose. Nature, 360:11– 12, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Greenbaum:1993:NFM

[Gre93] Anne Greenbaum. Norms of functions of matrices. Technical Report 645, Courant Institute of Mathematical Sciences, New York University, New York, August 1993. ???? pp.

Greenbaum:1994:LCG

[Gre94] Anne Greenbaum. The Lanczos and conjugate gradient algorithms in finite precision arithmetic. In J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors, Proceedings of the Cornelius Lanczos International Centenary Conference, pages 49–60. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Griffiths:1984:NAD

[Gri84] D. F. Griffiths, editor. Numerical Analysis, Dundee 1983, volume 1066 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1984.

Griewank:1988:AD

[Gri88] A. Griewank. On automatic differentiation. Technical report, Preprint ANL/MCS-P10-1088, Mathematics and Computer Science Division, Argonne National Laboratory, Illinois, ????, 1988. ???? pp.

Grigson:1989:CPS

[Gri89] Lionel Grigson. A Charlie Parker Study Album. Novello, London, 1989. ISBN 0-85360-140-2. ???? pp.

Gropp:1987:SNL

[Gro87] W. Gropp. A system for numerical linear algebra. In Wouk [Wou87], pages 26–38.

Grossman:1989:SCA

[Gro89] R. Grossman, editor. Symbolic Computation: Applications to Scientific Computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1989. ???? pp.

Grosse:1990:CAA

[Gro90] E. Grosse. A catalogue of algorithms for approximation. In Algorithms for Approximation II [MC90], pages 479–514.

Grone:1994:BMM

[Gro94] Robert Grone. A biography of Marvin Marcus. Linear Algebra and Appl., 201: 1–20, 1994.

Grosse:1995:BNR

[Gro95a] Eric Grosse. BibNet: A new resource for netlib users. SIAM News, 28(3):4, March 1995. ISSN 0036-1437.

Grosse:1995:RM

[Gro95b] Eric Grosse. Repository mirroring. ACM Transactions on Mathematical Software, 21(1):89–97, 1995. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gruvaeus:1970:GAP

[Gru70] Gunnar T. Gruvaeus. A general approach to Procrustes pattern rotation. Psychometrika, 35(4):493–505, 1970. CO-DEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Grenander:1958:TFT

[GS58] U. Grenander and G. Szegő. Toeplitz forms and their applications. University of California Press, Berkeley, Berkeley, CA, 1958. ???? pp.

Gentleman:1966:FFT

[GS66] W. M. Gentleman and G. Sande. Fast Fourier transforms—for fun and profit. In Fall Joint Computer Conference, volume 29 of AFIPS Conference Proceedings, pages 563–578. Spartan Books, New York, NY, USA, 1966.

Golub:1971:ACA

[GS71] G. H. Golub and L. B. Smith. Algorithm 414: Chebyshev approximation of continuous functions by a Chebyshev system of functions [E2]. Comm. ACM, 14: 737–746, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Golub:1974:SAN

[GS74] G. H. Golub and G. P. H. Styan. Some aspects of numerical computations for linear models. Technical report, Interface— Proceedings of Computer Science and Statistics, 7th Annual Symposium on the Interface 1973, ????, 1974. ???? pp.

Gaffney:1976:ABE

[GS76a] Matthew P. Gaffney and Lynn Arthur Steen. Annotated Bibliography of Expository Writing in the Mathematical Sciences. Mathematical Association of America, Washington, DC, USA, 1976. ISBN 0-88385-422-8. ???? pp.

Gragg:1976:SVS

[GS76b] W. B. Gragg and G. W. Stewart. A stable variant of the secant method for solving nonlinear equations. SIAM J. Numer. Anal., 13:889–903, 1976.

Gay:1978:SSN

[GS78] D. M. Gay and R. B. Schnabel. Solving systems of nonlinear equations by Broyden’s method with projected updates.

In Mangasarian et al. [MMR78b], pages 245–281. ISBN 0-12-486660-3. LCCN T57.8 .N67 1977.

Goldberg:1983:MNM

[GS83] Moshe Goldberg and E. G. Straus. Multiplicativity of l_{p} norms for matrices. Linear Algebra and Appl., 52/53:351–360, 1983.

Graham:1984:AHM

[GS84] R. L. Graham and N. J. A. Sloane. Anti-Hadamard matrices. Linear Algebra and Appl., 62:113–137, 1984.

Gear:1990:DOM

[GS90a] C. W. Gear and R. D. Skeel. The development of ODE methods: A symbiosis between hardware and numerical analysis. In Stephen G. Nash, editor, A History of Scientific Computing, pages 88–105. Addison-Wesley, Reading, Massachusetts, 1990.

Gluchowska:1990:SLL

[GS90b] J. Gluchowska and A. Smoktunowicz. Solving the linear least squares problem with very high relative accuracy. Computing, 45:345–354, 1990. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Gopen:1990:SSW

[GS90c] George D. Gopen and Judith A. Swan. The science of scientific writing. American Scientist, 78:550–558, 1990.

Greenbaum:1992:PBF

[GS92] A. Greenbaum and Z. Strakos. Predicting the behavior of finite precision Lanczos and conjugate gradient computations. SIAM J. Matrix Anal. Appl., 13 (1):121–137, 1992.

Gladwell:1988:LSE

[GSB88] I. Gladwell, L. F. Shampine, and R. W. Brankin. Locating special events when solving ODEs. Applied Mathematics Letters, 1:153–156, 1988. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Gill:1996:SCF

[GSS96] Philip E. Gill, Michael A. Saunders, and Joseph R. Shinnerl. On the stability of Cholesky factorization for symmetric quasi-definite systems. SIAM J. Matrix Anal. Appl., 17:??, 1996. To appear.

Greenbaum:1993:DPM

[GT93] Anne Greenbaum and Lloyd N. Trefethen. Do the pseudospectra of a matrix determine its behavior? Technical Report TR 93-1371, Department of Computer Science, Cornell University, Ithaca, New York, August 1993. ???? pp.

Golub:1977:BLM

[GU77] G. H. Golub and R. Underwood.

The block Lanczos method for computing eigenvalues. In Rice [Ric77], pages 361– 377. ISBN 0-12-587260-7.

Gulliksson:1992:EAA

[Gul92] M. Gulliksson. Error analysis of an algorithm for constrained and weighted linear least squares. Report UMINF 92.11, Department of Information Processing, University of Umeå, Sweden, November 1992. ???? pp.

Gulliksson:1994:IRC

[Gul94] Mårten Gulliksson. Iterative refinement for constrained and weighted linear least squares. BIT, 34:239–253, 1994. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gulliksson:1995:BEA

[Gul95] Mårten Gulliksson. Backward error analysis for the constrained and weighted linear least squares problem when using the weighted QR factorization. SIAM J. Matrix Anal. Appl., 16(2):675–687, 1995.

Gurwitz:1992:TCE

[Gur92] Chaya Gurwitz. A test for cancellation errors in quasi-Newton methods. SIAM J. Sci. Stat. Comput., 18(2):134– 140, 1992.

Gurari:1993:TLD

[Gur93] Eitan M. Gurari. TEX and LaTEX: Drawing and Literate Programming. Mc-Graw-Hill, New York, 1993. ISBN 0-07-911616-7. ???? pp.

Gustafson:1970:CEC

[Gus70] Sven-Åke Gustafson. Control and estimation of computational errors in the evaluation of interpolation formulae and quadrature rules. Math. Comp., 24(112): 847–854, 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gustafson:1988:RAL

[Gus88] J. L. Gustafson. Reevaluating Amdahl’s law. Comm. ACM, 31:532–533, 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Golub:1974:CBS

[GV74] G. H. Golub and J. M. Varah. On a characterization of the best ℓ_{2}-scaling of a matrix. SIAM J. Numer. Anal., 11(3): 472–479, 1974.

Golub:1979:UPD

[GV79] Gene H. Golub and Charles F. Van Loan. Unsymmetric positive definite linear systems. Linear Algebra and Appl., 28:85–97, 1979.

Golub:1980:ATL

[GV80] G. H. Golub and C. F. Van Loan. An analysis of the total least squares problem. SIAM J. Numer. Anal., 17:883–893, 1980.

Golub:1983:MC

[GV83] G. H. Golub and C. F. Van Loan. Matrix Computations. The Johns Hopkins University Press, Baltimore, MD, USA, 1983. ???? pp.

Golub:1989:MC

[GV89] Gene H. Golub and Charles F. Van Loan. Matrix Computations, volume 3 of Johns Hopkins Series in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, second edition, 1989. ISBN 0-8018-3772-3 (hardcover), 0-8018-3739-1 (paperback). xix + 642 pp. LCCN QA188 .G65 1989. US$14.50.

Goldstine:1951:NIM

[GvN51] Herman H. Goldstine and John von Neumann. Numerical inverting of matrices of high order II. Proc. Amer. Math. Soc., 2:188–202, 1951. Reprinted in [Tau63, pp. 558–572].

Golub:1966:NIR

[GW66] G. H. Golub and J. H. Wilkinson. Note on the iterative refinement of least squares solution. Numer. Math., 9:139–148, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gourlay:1973:CMM

[GW73] A. R. Gourlay and G. A. Watson. Computational Methods for Matrix Eigenproblems. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1973. ISBN 0-471-31915-5. ???? pp.

Golub:1976:ICE

[GW76] G. H. Golub and J. H. Wilkinson. Illconditioned eigensystems and the computation of the Jordan canonical form. SIAM Review, 18(4):578–619, 1976.

Griffiths:1986:NAD

[GW86] D. F. Griffiths and G. A. Watson. Numerical analysis, dundee 1985. Technical report, Pitman Research Notes in Mathematics 140, Longman Scientific and Technical, ????, 1986. ???? pp.

Griffiths:1990:NAP

[GW90] D. F. Griffiths and G. A. Watson, editors. Numerical Analysis 1989, Proceedings of the 13th Dundee Conference, volume 228 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, UK, 1990.

Griffiths:1992:NAP

[GW92a] D. F. Griffiths and G. A. Watson, editors. Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, UK, 1992.

Gulliksson:1992:MQD

[GW92b] Mårten Gulliksson and Per-Åke Wedin. Modifying the QR-decomposition to constrained and weighted linear least squares. SIAM J. Matrix Anal. Appl., 13 (4):1298–1313, 1992.

Griffiths:1994:NAP

[GW94] D. F. Griffiths and G. A. Watson, editors. Numerical Analysis 1993, Proceedings of the 15th Dundee Conference, volume 303 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, UK, 1994.

Gardiner:1992:AFS

[GWL^{+}92] Judith D. Gardiner, Matthew R.

Wette, Alan J. Laub, James J. Amato, and Cleve B. Moler. Algorithm 705: A FORTRAN-77 software package for solving the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 18(2): 232–238, 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gordon:1992:OCH

[GWW92] Carolyn Gordon, David L. Webb, and Scott Wolpert. One cannot hear the shape of a drum. Bull. Amer. Math. Soc., 27(1):134–138, 1992.

Golub:1993:PAC

[GZ93] G. H. Golub and Hongyuan Zha. Perturbation analysis of the canonical correlations of matrix pairs. Technical Report CS-93-2, Department of Computer Science, Pennsylvania State University, University Park, PA, February 1993. ???? pp.

Hull:1985:NT

[HAC^{+}85] T. E. Hull, A. Abraham, M. S. Cohen, A. F. X. Curley, C. B. Hall, D. A. Penny, and J. T. M. Sawchuk. Numerical TURING. ACM SIGNUM Newsletter, 20(3):26–34, 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hacon:1993:JMS

[Hac93] Derek Hacon. Jacobi’s method for skew-symmetric matrices. SIAM J. Matrix Anal. Appl., 14(3):619–628, 1993.

Hager:1984:CE

[Hag84] William W. Hager. Condition estimates. SIAM J. Sci. Stat. Comput., 5(2): 311–316, 1984.

Hager:1988:ANL

[Hag88] William W. Hager. Applied Numerical Linear Algebra. Prentice-Hall, Englewood Cliffs, NJ, USA, 1988. ISBN 0-13-041369-0. viii + 424 pp.

Hager:1989:UIM

[Hag89] William W. Hager. Updating the inverse of a matrix. SIAM Review, 31(2): 221–239, 1989.

Hall:1967:CT

[Hal67a] Marshall Hall, Jr. Combinatorial Theory. Blaisdell, Waltham, MA, USA, 1967. x + 310 pp.

Halmos:1967:HSP

[Hal67b] Paul R. Halmos. A Hilbert Space Problem Book. Springer-Verlag, Berlin, 1967. ???? pp.

Halmos:1970:HWM

[Hal70] Paul R. Halmos. How to write mathematics. L’Enseignement Mathématique, 16:123–152, 1970. Reprinted in [SHSD73] and [SG83a].

Halmos:1972:PAO

[Hal72] P. R. Halmos. Positive approximants of operators. Indiana Univ. Math. J., 21: 951–960, 1972. CODEN IUMJAB. ISSN 0022-2518.

Halmos:1974:NST

[Hal74a] P. R. Halmos. Naive Set Theory.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ???? pp.

Halmos:1974:SAN

[Hal74b] P. R. Halmos. Spectral approximants of normal operators. Proc. Edinburgh Math. Soc., 19:51–58, 1974. CO-DEN PEMSA3. ISSN 0013-0915 (print), 1464-3839 (electronic).

Halmos:1974:FDV

[Hal74c] Paul R. Halmos. Finite-Dimensional Vector Spaces. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ???? pp.

Halmos:1974:HTM

[Hal74d] Paul R. Halmos. How to talk mathematics. Notices Amer. Math. Soc., 21(3): 155–158, April 1974. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted in [SG83a].

Halmos:1975:WP

[Hal75] Paul R. Halmos. What to publish.

Amer. Math. Monthly, 82(1):14–17, 1975. Reprinted in [SG83a].

Haley:1977:SMM

[Hal77] Stephen B. Haley. Solution of modified matrix equations. SIAM J. Numer. Anal., 24(4):946–951, 1977.

Haley:1980:SBM

[Hal80] Stephen B. Haley. Solution of band matrix equations by projectionrecurrence. Linear Algebra and Appl., 32: 33–48, 1980.

Halmos:1982:HSP

[Hal82a] Paul R. Halmos. A Hilbert Space Problem Book. Springer-Verlag, Berlin, second edition, 1982. ???? pp.

Halmos:1982:TIG

[Hal82b] Paul R. Halmos. Think it gooder. The Mathematical Intelligencer, 4(1):20– 21, 1982. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Halmos:1985:WMA

[Hal85] Paul R. Halmos. I Want to be a Mathematician: An Automathography in Three Parts. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-88385-445-7. ???? pp.

Halmos:1986:EA

[Hal86] P. R. Halmos. Extensions and approximations. Linear and Multilinear Algebra, 19:297–298, 1986. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Halmos:1987:PM

[Hal87] Paul R. Halmos. I Have A Photographic Memory. American Mathematical Society, Providence, RI, USA, 1987. ???? pp.

Halmos:1988:SBA

[Hal88] Paul R. Halmos. Some books of auld lang syne. In Peter Duren, Richard A. Askey, and Uta C. Merzbach, editors, A Century of Mathematics in America, Part I, pages 131–174. American Mathematical Society, Providence, RI, USA, 1988.

Hald:1990:HPS

[Hal90a] Anders Hald. A History of Probability and Statistics and Their Applications Before 1750. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1990. ISBN 0-471-50230-8. xiii + 586 pp.

Halmos:1990:PMS

[Hal90b] P. R. Halmos. Has progress in mathematics slowed down? Amer. Math. Monthly, 97(7):561–588, 1990.

Halmos:1991:BPG

[Hal91a] P. R. Halmos. Bad products of good matrices. Linear and Multilinear Algebra, 29:1–20, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Halmos:1991:PMY

[Hal91b] Paul R. Halmos. Problems for Mathematicians, Young and Old. Number 12 in The Dolcani Mathematical Expositions. Mathematical Association of America, Washington, DC, 1991. ISBN 0-88385-320-5. ???? pp.

Halmos:1994:WT

[Hal94] Paul R. Halmos. What is teaching?

Amer. Math. Monthly, 101(9):848–854, 1994.

Halfhill:1995:TBP

[Hal95] Tom R. Halfhill. The truth behind the Pentium bug. BYTE Magazine, pages 163–164, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hamming:1973:NMS

[Ham73] R. W. Hamming. Numerical Methods for Scientists and Engineers. McGraw-Hill, New York, second edition, 1973. ISBN 0-07-025887-2. ix + 721 pp.

Hammarling:1974:NMG

[Ham74] Sven Hammarling. A note on modifications to the Givens plane rotation. J. Inst. Maths Applics, 13:215–218, 1974.

Hammarling:1982:NMS

[Ham82a] S. J. Hammarling. Newton’s method for solving the algebraic Riccati equation. NPL Report DITC 12/82, National Physical Laboratory, Teddington, UK, September 1982. ???? pp.

Hammarling:1982:NSS

[Ham82b] S. J. Hammarling. Numerical solution of the stable, non-negative definite Lyapunov equation. IMA J. Numer. Anal., 2:303–323, 1982.

Hammarling:1982:SNU

[Ham82c] S. J. Hammarling. Some notes on the use of orthogonal similarity transformations in control. NPL Report DITC 8/82, National Physical Laboratory, Teddington, UK, August 1982. ???? pp.

Hammarling:1985:SVD

[Ham85] S. J. Hammarling. The singular value decomposition in multivariate statistics. ACM SIGNUM Newsletter, 20: 2–25, 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hamada:1987:NRN

[Ham87] Hozumi Hamada. A new real number representation and its operation. In Mary Jane Irwin and Renato Stefanelli, editors, Proceedings of the Eighth Symposium on Computer Arithmetic, Como, Italy, pages 153–157. IEEE Computer Society, Washington, DC, 1987.

Hammarling:1994:PLW

[Ham94] Sven J. Hammarling. Parallel library work at NAG. In Jack. J. Dongarra and Bernard Tourancheau, editors, Environments and Tools For Parallel Scientific Computing, pages 172–182. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Hansen:1969:TIA

[Han69] Eldon Hansen, editor. Topics in Interval Analysis. Oxford University Press, ????, 1969.

Hansen:1987:TSM

[Han87] P. C. Hansen. The truncated SVD as a method for regularization. BIT, 27: 534–553, 1987. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hansen:1989:RGT

[Han89] P. C. Hansen. Regularization, GSVD and truncated GSVD. BIT, 29:491–504, 1989. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hansen:1990:DPC

[Han90a] P. C. Hansen. The discrete Picard condition for discrete ill-posed problems. BIT, 30:658–672, 1990. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hansen:1990:RBS

[Han90b] P. C. Hansen. Relations between SVD and GSVD of discrete regularization problems in standard and general form. Linear Algebra and Appl., 141:165–176, 1990.

Hansen:1990:TSV

[Han90c] P. C. Hansen. Truncated singular value decomposition solutions to discrete ill-posed problems with ill-determined numerical rank. SIAM J. Sci. Stat. Comput., 11:503–518, 1990.

Hansen:1992:RTM

[Han92] Per Christian Hansen. Regularization tools. A Matlab package for analysis and solution of discrete ill-posed problems. Report UNIC-92-03, UNI·C, Technical University of Denmark, DK-2800 Lyngby, Denmark, June 1992. ???? pp.

Hansen:1993:TMR

[Han93] Per Christian Hansen. Test matrices for regularization methods. Report UNIC-93-09, UNI·C, Technical University of Denmark, DK-2800 Lyngby, Denmark, August 1993. ???? pp.

Hansen:1995:TMR

[Han95] Per Christian Hansen. Test matrices for regularization methods. SIAM J. Sci. Comput., 16(2):506–512, 1995.

Hardy:1909:I

[Har09] G. H. Hardy. The integral ∫_{0}^{∞} \frac{sin x}{x} dx. Mathematical Gazette, page ??, 1909. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). reprinted vol. 55, 1971, pp. 152–158.

Hardy:1940:MA

[Har40] G. H. Hardy. A Mathematician’s Apology. Cambridge University Press, ????, 1940. ???? pp. 1967 edition with foreward by C. P. Snow.

Hartree:1958:NA

[Har58] Douglas R. Hartree. Numerical Analysis. Oxford University Press, ????, second edition, 1958. xvi + 302 pp.

Harlow:1965:NFD

[Har65] F. H. Harlow. Numerical fluid dynamics. Amer. Math. Monthly, 72:84–91, 1965.

Harrison:1966:IS

[Har66] G. B. Harrison, editor. Introducing Shakespeare. Penguin, ????, third edition, 1966. ISBN 0-14-013613-4. ???? pp.

Hardy:1967:CPM

[Har67] G. H. Hardy. A Course of Pure Mathematics. Cambridge University Press, ????, tenth edition, 1967. ISBN 0-521-09227-2. xii + 509 pp.

Harter:1972:OWF

[Har72] Richard Harter. The optimality of Winograd’s formula. Comm. ACM, 15 (5):352, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harman:1975:HP

[Har75] Eleanor Harman. Hints on proofreading. Scholarly Publishing, 6(2):151–157, 1975. ISSN 0036-634X.

Hartree:1976:EAC

[Har76] Edward F. Hartree. Ethics for authors: A case history of acrosin. Perspectives in Biology and Medicine, 20:82–91, 1976. CODEN PBMEA8. ISSN 0031-5982 (print), 1529-8795 (electronic).

Hartfiel:1980:CSS

[Har80] D. J. Hartfiel. Concerning the solution set of Ax = b where P ≤ A ≤ Q and p ≤ b ≤ q. Numer. Math., 35:355–359, 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hartley:1981:EWI

[Har81a] James Hartley. Eighty ways of improving instructional text. IEEE Trans. Prof. Commun., PC-24:17–27, 1981.

Hartmanis:1981:OAD

[Har81b] Juris Hartmanis. Observations about the development of theoretical computer science. Annals of the History of Computing, 3(1):42–51, 1981.

Hart:1983:HRC

[Har83] Horace Hart. Hart’s Rules for Compositors and Readers at the University Press Oxford. Oxford University Press, ????, thirty-ninth edition, 1983. ISBN 0-19-212983-X. ???? pp. First edition 1893.

Hartley:1984:RCT

[Har84a] James Hartley. The role of colleagues and text-editing programs in improving text. IEEE Trans. Prof. Commun., PC-27:42–44, 1984.

Hartree:1984:CMR

[Har84b] Douglas R. Hartree. Calculating Machines: Recent and Prospective Developments and Their Impact on Mathematical Physics, and Calculating Instruments and Machines, volume 6 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-08147-4. ???? pp.

Hartley:1985:DIT

[Har85] James Hartley. Designing Instructional Text. Kogan Page, London, second edition, 1985. ISBN 0-85038-943-7. ???? pp.

Harrod:1986:DTI

[Har86a] W. J. Harrod. LU-decompositions of tridiagonal irreducible H-matrices. SIAM J. Alg. Disc. Meth., 7:180–187, 1986.

Hartnell:1986:DPB

[Har86b] T. Hartnell. Desktop Publishing:

The Book. Interface Publications, London, 1986. ???? pp.

HarperCollins:1991:CED

[Har91a] HarperCollins. Collins English Dictionary. HarperCollins College Publishers, New York, NY, USA, third edition, 1991. ISBN 0-00-433286-5. ???? pp.

Harris:1991:YWP

[Har91b] Sidney Harris. “You Want Proof?

I’ll Give You Proof!” More Cartoons from Sidney Harris. W.H. Freeman, San Francisco, CA, USA and New York, NY, USA, 1991. ISBN 0-7167-2159-7. ???? pp.

Harris:1995:EC

[Har95] Robert Harris. The Enigma Connection. Random House, New York, NY, USA, 1995. ???? pp.

Hastings:1955:ADC

[Has55] Cecil Hastings, Jr. Approximations for Digital Computers. Princeton University Press, Princeton, NJ, USA, 1955. viii + 201 pp.

Hausner:1967:UPD

[Hau67] A. Hausner. Uniqueness of the polar decomposition. Amer. Math. Monthly, 74: 303–304, 1967.

Hauptman:1991:CUP

[Hau91] Don Hauptman. Cruel and Unusual Puns. Dell Publishing, New York, 1991. ISBN 0-440-20850-5. ???? pp.

Hay:1988:MMP

[Hay88] Cynthia Hay, editor. Mathematics from Manuscript to Print 1300–1600. Clarendon Press, New York, NY, USA, 1988.

Haykin:1989:STS

[Hay89] S. Haykin, editor. Selected Topics in Signal Processing. Prentice-Hall, Englewood Cliffs, NJ, USA, 1989. ???? pp.

Householder:1959:CME

[HB59] Alston S. Householder and Friedrich L. Bauer. On certain methods for expanding the characteristic polynomial. Numer. Math., 1:29–37, 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Howard:1986:TP

[HB86] V. A. Howard and J. H. Barton. Thinking on Paper. Quill/William Morrow, New York, 1986. ISBN 0-688-07758-7. ???? pp.

Householder:1963:SVI

[HC63] A. S. Householder and J. A. Carpenter. The singular values of involutory and of idempotent matrices. Numer. Math., 5:234–237, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Holt:1988:TPL

[HC88] R. C. Holt and J. R. Cordy. The Turing programming language. Comm. ACM, 31(12):1410–1423, 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harper:1992:IM

[HC92] David Harper and Rachid Chalabi. An introduction to Maple. Report, Computer Algebra Support Project, Astronomy Unit, School of Mathematical Sciences, Queen Mary and Westfield College, University of London, London, E1 4NS, November 1992. ???? pp. Third edition. Available by anonymous ftp

from maths.qmw.ac.uk (138.37.80.15)
in pub/mapledoc.

Hull:1988:EHS

[HCSW88] T. E. Hull, M. S. Cohen, J. T. M.

Sawchuk, and D. B. Wortman. Exception handling in scientific computing. ACM Transactions on Mathematical Software, 14(3):201–217, 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hennell:1980:PAN

[HD80] M. A. Hennell and L. M. Delves.

Production and Assessment of Numerical Software. Academic Press, New York, USA, 1980. ???? pp.

He:1990:ELS

[He90] C. He. On estimating the least singular value of a matrix. Linear Algebra and Appl., 128:133–138, 1990.

Hearon:1967:PFM

[Hea67] John Z. Hearon. Polar factorization of a matrix. J. Res. Nat. Bur. Standards, 71B(2-3):65–67, 1967.

Hearon:1977:NS

[Hea77] John Z. Hearon. Nonsingular solutions of T A − BT = C. Linear Alg. and its Appl., 16:57–63, 1977.

Heath:1984:NML

[Hea84] Michael T. Heath. Numerical methods for large sparse linear least squares problems. SIAM J. Sci. Stat. Comput., 5 (3):497–513, 1984.

Hein:1992:G

[Hei92a] Piet Hein. Grooks. Number 85 in Borgen’s Pocketbooks. Narayana Press, Gylling, second edition, 1992. ISBN 87-418-1079-1. 53 pp.

Hein:1992:GI

[Hei92b] Piet Hein. Grooks II. Number 84 in Borgen’s Pocketbooks. Narayana Press, Gylling, second edition, 1992. ISBN 87-418-1090-2. 53 pp.

Helton:1968:LM

[Hel68] Burrell W. Helton. Logarithms of matrices. Proc. Amer. Math. Soc., 19:733– 738, 1968.

Heller:1974:DTA

[Hel74] Don Heller. A determinant theorem with applications to parallel algorithms. SIAM J. Numer. Anal., 11(3):559–568, 1974.

Heller:1978:SPA

[Hel78] Don Heller. A survey of parallel algorithms in numerical linear algebra. SIAM Review, 20(4):740–777, 1978.

Hemmerle:1967:SCD

[Hem67] William J. Hemmerle. Statistical Computations on a Digital Computer.

Blaisdell, Waltham, MA, USA, 1967. x + 230 pp.

Henrici:1962:BII

[Hen62a] Peter Henrici. Bounds for iterates, inverses, spectral variation and fields of values of non-normal matrices. Numer. Math., 4:24–40, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henrici:1962:DVM

[Hen62b] Peter Henrici. Discrete Variable Methods in Ordinary Differential Equations. John Wiley, New York, NY, USA, 1962. xi + 407 pp.

Henrici:1963:EPD

[Hen63] Peter Henrici. Error Propagation for Difference Methods. John Wiley, New York, NY, USA, 1963. vi + 73 pp.

Henrici:1964:ENA

[Hen64] Peter Henrici. Elements of Numerical Analysis. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1964. xv + 328 pp.

Henrici:1966:TPM

[Hen66] Peter Henrici. Test of probabilistic models for the propagation of roundoff errors. Comm. ACM, 9(6):409–410, 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henon:1976:TDM

[Hén76] M. Hénon. A two-dimensional mapping with a strange attractor. Commun. Math. Phys., 50:69–77, 1976. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

Henrici:1980:MPR

[Hen80] Peter Henrici. A model for the propagation of rounding error in floating arithmetic. In Karl L. E. Nickel, editor, Interval Mathematics 1980, pages 49–73. Academic Press, New York, 1980.

Hennart:1982:NAM

[Hen82] J. P. Hennart, editor. Numerical Analysis, Mexico 1981, volume 909 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1982.

Hennart:1986:NAM

[Hen86] J. P. Hennart, editor. Numerical Analysis, Mexico 1984, volume 1230 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1986.

Henry:1992:IDR

[Hen92] G. Henry. Increasing data reuse in the unsymmetric QR algorithm. Technical Report CTC92TR100, Cornell Theory Center, Ithaca, New York, July 1992. ???? pp.

Herbert:1965:STE

[Her65] A. J. Herbert. The Structure of Technical English. Longman, Harlow, Essex, 1965. ISBN 0-582-52523-3. ???? pp.

Herman:1988:MR

[Her88] Eugene A. Herman. Mathematica— A review. Notices Amer. Math. Soc., 35:1334–1344, November 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

HP:1982:HAF

[Hew82] Hewlett–Packard, Portable Computer Division, Corvallis, OR, USA. HP-15C Advanced Functions Handbook, 1982. 221 pp. Part number 00015-90011 Rev. C.

HP:1993:HSU

[Hew93] Hewlett–Packard, Corvallis Division, Corvallis, OR, USA. HP 48G Series User’s Guide, 1993. ???? pp. Part number 00048-90126, Edition 3.

Heyman:1987:FCD

[Hey87] Daniel P. Heyman. Further comparisons of direct methods for computing stationary distributions of Markov chains. SIAM J. Alg. Disc. Meth., 8(2):226–232, 1987.

Hull:1994:ICE

[HFT94] T. E. Hull, Thomas F. Fairgrieve, and Ping Tak Peter Tang. Implementing complex elementary functions using exception handling. ACM Transactions on Mathematical Software, 20(2): 215–244, 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ha:1980:NDF

[HG80] T. T. Ha and J. A. Gibson. A note on the determinant of a functional confluent Vandermonde matrix and controllability. Linear Algebra and Appl., 30:69–75, 1980.

Heath:1990:EEI

[HGD90] M. T. Heath, G. A. Geist, and J. B.

Drake. Early experience with the Intel iPSC/860 at Oak Ridge National Laboratory. Report ORNL/TM-11655, Oak Ridge National Laboratory, Oak Ridge, TN, USA, September 1990. 26 pp.

Halme:1988:GED

[HH88] H. Halme and J. Heinänen. GNU Emacs as a dynamically extensible programming environment. Software—Practice and Experience, 18:999–1009, 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Higham:1989:LGF

[HH89] Nicholas J. Higham and Desmond J. Higham. Large growth factors in Gaussian elimination with pivoting. SIAM J. Matrix Anal. Appl., 10(2):155–164, April 1989.

Hake:1991:IMO

[HH91a] J.-Fr. Hake and W. Homberg. The impact of memory organization on the performance of matrix calculations. Parallel Computing, 17:311–327, 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Haemmerlin:1991:NM

[HH91b] Günther Hämmerlin and Karl-Heinz Hoffmann. Numerical Mathematics. Springer-Verlag, New York, 1991. ISBN 0-387-97494-6. xi + 422 pp.

Higham:1992:BEC

[HH92a] Desmond J. Higham and Nicholas J. Higham. Backward error and condition of structured linear systems. SIAM J. Matrix Anal. Appl., 13(1):162–175, January 1992.

Higham:1992:CPT

[HH92b] Desmond J. Higham and Nicholas J. Higham. Componentwise perturbation theory for linear systems with multiple right-hand sides. Linear Algebra and Appl., 174:111–129, 1992.

Harmon:1993:BCT

[HH93] Joseph E. Harmon and David R. Hamrin. Bibliography on communicating technical research information. IEEE Trans. Prof. Commun., 36(1):2–6, 1993.

Hammarling:1993:CHW

[HHK93] Sven Hammarling, Nicholas J.

Higham, and Bo Kågström. CERFACS hosts workshop on reliability of computations. SIAM News, 26:4, July 1993. ISSN 0036-1437.

Hammer:1993:NTV

[HHKR93] Rolf Hammer, Matthias Hocks, Ulrich Kulisch, and Dietmar Ratz. Numerical Toolbox for Verified Computing I. Basic Numerical Problems: Theory, Algorithms, and Pascal-XSC Programs. Springer-Verlag, Berlin, 1993. ISBN 3-540-57118-3. xiii + 337 pp.

Higham:1983:MCN

[Hig83a] N. J. Higham. Matrix condition numbers. M.Sc. Thesis, University of Manchester, Manchester, England, October 1983. iv + 86 pp.

Higham:1983:UBC

[Hig83b] Nicholas J. Higham. Upper bounds for the condition number of a triangular matrix. Numerical Analysis Report No. 86, University of Manchester, Manchester, England, May 1983. ???? pp.

Higham:1985:NPN

[Hig85] Nicholas J. Higham. Nearness Problems in Numerical Linear Algebra. PhD thesis, University of Manchester, Manchester, England, July 1985. 173 pp.

Higham:1986:CPD

[Hig86a] Nicholas J. Higham. Computing the polar decomposition—with applications. SIAM J. Sci. Stat. Comput., 7(4):1160– 1174, October 1986.

Higham:1986:EAC

[Hig86b] Nicholas J. Higham. Efficient algorithms for computing the condition number of a tridiagonal matrix. SIAM J. Sci. Stat. Comput., 7(1):150–165, 1986.

Higham:1986:MCB

[Hig86c] Nicholas J. Higham. Matrix computations in Basic on a microcomputer. IMA Bulletin, 22:13–20, 1986.

Higham:1986:NMM

[Hig86d] Nicholas J. Higham. Newton’s method for the matrix square root. Math. Comp., 46(174):537–549, April 1986. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Higham:1987:CRS

[Hig87a] Nicholas J. Higham. Computing real square roots of a real matrix. Linear Algebra and Appl., 88/89:405–430, 1987.

Higham:1987:EAB

[Hig87b] Nicholas J. Higham. Error analysis of the Björck-Pereyra algorithms for solving Vandermonde systems. Numer. Math., 50(5):613–632, 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:1987:SCN

[Hig87c] Nicholas J. Higham. A survey of condition number estimation for triangular matrices. SIAM Review, 29(4):575–596, December 1987.

Higham:1988:CNS

[Hig88a] Nicholas J. Higham. Computing a nearest symmetric positive semidefinite matrix. Linear Algebra and Appl., 103: 103–118, 1988.

Higham:1988:FSV

[Hig88b] Nicholas J. Higham. Fast solution of Vandermonde-like systems involving orthogonal polynomials. IMA J. Numer. Anal., 8:473–486, 1988.

Higham:1988:FCE

[Hig88c] Nicholas J. Higham. FORTRAN codes for estimating the one-norm of a real or complex matrix, with applications to condition estimation (Algorithm 674). ACM Transactions on Mathematical Software, 14(4):381–396, December 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Higham:1988:SPP

[Hig88d] Nicholas J. Higham. The symmetric Procrustes problem. BIT, 28:133–143, 1988. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:1989:AST

[Hig89a] Nicholas J. Higham. The accuracy of solutions to triangular systems. SIAM J. Numer. Anal., 26(5):1252–1265, October 1989.

Higham:1989:CTM

[Hig89b] Nicholas J. Higham. A collection of test matrices in MATLAB. Numerical Analysis Report No. 172, University of Manchester, Manchester, England, July 1989. ???? pp.

Higham:1989:MCP

[Hig89c] Nicholas J. Higham. Matrix computations on a PC. SIAM News, 22:16, January 1989. ISSN 0036-1437.

Higham:1989:MNP

[Hig89d] Nicholas J. Higham. Matrix nearness problems and applications. In M. J. C.

Gover and S. Barnett, editors, Applications of Matrix Theory, pages 1–27. Oxford University Press, 1989.

Higham:1990:ACD

[Hig90a] Nicholas J. Higham. Analysis of the Cholesky decomposition of a semi-definite matrix. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 161–185. Oxford University Press, 1990.

Higham:1990:BEG

[Hig90b] Nicholas J. Higham. Bounding the error in Gaussian elimination for tridiagonal systems. SIAM J. Matrix Anal. Appl., 11(4):521–530, October 1990.

Higham:1990:CEB

[Hig90c] Nicholas J. Higham. Computing error bounds for regression problems. In Philip J. Brown and Wayne A. Fuller, editors, Statistical Analysis of Measurement Error Models and Applications, Contemporary Mathematics 112, pages 195–208. American Mathematical Society, Providence, RI, USA, 1990.

Higham:1990:EMN

[Hig90d] Nicholas J. Higham. Experience with a matrix norm estimator. SIAM J. Sci. Stat. Comput., 11(4):804–809, July 1990.

Higham:1990:EFM

[Hig90e] Nicholas J. Higham. Exploiting fast matrix multiplication within the level 3 BLAS. ACM Transactions on Mathematical Software, 16(4):352–368, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Higham:1990:HAG

[Hig90f] Nicholas J. Higham. How accurate is Gaussian elimination? In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1989, Proceedings of the 13th Dundee Conference, volume 228 of Pitman Research Notes in Mathematics, pages 137–154. Longman Scientific and Technical, Harlow, Essex, UK, 1990.

Higham:1990:FMM

[Hig90g] Nicholas J. Higham. Is fast matrix multiplication of practical use? SIAM News, 23:12, November 1990. ISSN 0036-1437.

Higham:1990:IRE

[Hig90h] Nicholas J. Higham. Iterative refinement enhances the stability of QR factorization methods for solving linear equations. Numerical Analysis Report No. 182, University of Manchester, Manchester, England, April 1990. ???? pp.

Higham:1990:RGH

[Hig90i] Nicholas J. Higham. Review of “G. H.

Golub and C. F. Van Loan, Matrix Computations, Second Edition, Johns Hopkins University Press, Baltimore, Maryland, 1989”. Linear Algebra and Appl., 141:289–292, 1990.

Higham:1990:RMJ

[Hig90j] Nicholas J. Higham. Review of “M.

J. C. Gover and S. Barnett, eds., Applications of Matrix Theory, Oxford University Press, 1989”. Mathematical Gazette, 74(468):202, 1990. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Higham:1990:SAA

[Hig90k] Nicholas J. Higham. Stability analysis of algorithms for solving confluent Vandermonde-like systems. SIAM J. Matrix Anal. Appl., 11(1):23–41, January 1990.

Higham:1991:RA

[Hig91a] Desmond J. Higham. Remark on Algorithm 669. ACM Transactions on Mathematical Software, 17(3):424–426, 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Higham:1991:ACT

[Hig91b] Nicholas J. Higham. Algorithm 694: A collection of test matrices in MAT-LAB. ACM Transactions on Mathematical Software, 17(3):289–305, September 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Higham:1991:IRE

[Hig91c] Nicholas J. Higham. Iterative refinement enhances the stability of QR factorization methods for solving linear equations. BIT, 31:447–468, 1991. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:1991:SLE

[Hig91d] Nicholas J. Higham. Solving linear equations. In Christine Bondi, editor, New Applications of Mathematics, pages 33–56. Penguin, London, 1991.

Higham:1991:TMP

[Hig91e] Nicholas J. Higham. Three measures of precision in floating point arithmetic. NA Digest, 91(16):??, 1991. Electronic

                                        mail magazine: na.help@na-net.ornl.
                                        gov.

Higham:1992:MCG

[Hig92a] Desmond J. Higham. More commandments of good writing. Manuscript, Department of Mathematics and Computer Science, University of Dundee, November 1992.

Higham:1992:EMN

[Hig92b] Nicholas J. Higham. Estimating the matrix p-norm. Numer. Math., 62: 539–555, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:1992:IWP

[Hig92c] Nicholas J. Higham. IMA workshop participants Fête Golub on 60th birthday. SIAM News, 25:3, May 1992. ISSN 0036-1437.

Higham:1992:JAF

[Hig92d] Nicholas J. Higham. The joy of anonymous FTP. IMA Numerical Analysis Newsletter, 17(1):61–64, October 1992.

Higham:1992:LRM

[Hig92e] Nicholas J. Higham. LAPACK released in March. SIAM News, 25:20, May 1992. ISSN 0036-1437.

Higham:1992:SMM

[Hig92f] Nicholas J. Higham. Stability of a method for multiplying complex matrices with three real matrix multiplications. SIAM J. Matrix Anal. Appl., 13(3):681– 687, July 1992.

Higham:1993:AFP

[Hig93a] Nicholas J. Higham. The accuracy of floating point summation. SIAM J. Sci. Comput., 14(4):783–799, July 1993.

Higham:1993:HWM

[Hig93b] Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-314-5. xii + 241 pp.

Higham:1993:ODS

[Hig93c] Nicholas J. Higham. Optimization by direct search in matrix computations. SIAM J. Matrix Anal. Appl., 14(2):317– 333, April 1993.

Higham:1993:PTB

[Hig93d] Nicholas J. Higham. Perturbation theory and backward error for AX − XB = C. BIT, 33:124–136, 1993. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:1993:TMT

[Hig93e] Nicholas J. Higham. The Test Matrix Toolbox for Matlab. Numerical Analysis Report No. 237, Manchester Centre for Computational Mathematics, Manchester, England, December 1993. 76 pp.

Higham:1994:BVT

[Hig94a] Nicholas J. Higham. BibTEX: A versatile tool for LaTEX users. SIAM News, 27:10, 11, 19, January 1994. ISSN 0036-1437.

Higham:1994:HWM

[Hig94b] Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. Nippon Hyoron Sha, Tokyo, 1994. ISBN 4-535-78211-3. x + 235 pp. Japanese edition of [Hig93b], translated by Shoji Okumura and Takemitsu Hasegawa.

Higham:1994:MSD

[Hig94c] Nicholas J. Higham. The matrix sign decomposition and its relation to the polar decomposition. Linear Algebra and Appl., 212/213:3–20, 1994.

Higham:1994:SCP

[Hig94d] Nicholas J. Higham. A survey of componentwise perturbation theory in numerical linear algebra. In Walter Gautschi, editor, Mathematics of Computation 1943–1993: A Half Century of Computational Mathematics, volume 48 of Proceedings of Symposia in Applied Mathematics, pages 49–77. American Mathematical Society, Providence, RI, USA, 1994.

Higham:1994:WDM

[Hig94e] Nicholas J. Higham. Which dictionary for the mathematical scientist? IMA Bulletin, 30(5/6):81–88, 1994.

Higham:1995:CNT

[Hig95a] Desmond J. Higham. Condition numbers and their condition numbers. Linear Algebra and Appl., 214:193–213, 1995.

Higham:1995:ASN

[Hig95b] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-355-2 (paperback). xxviii + 688 pp. LCCN QA297.H53 1996. US$39.00. Typeset with LaTEX2e.

Higham:1995:IRL

[Hig95c] Nicholas J. Higham. Iterative refinement and LAPACK. Numerical Analysis Report No. 277, Manchester Centre for Computational Mathematics, Manchester, England, September 1995. 17 pp. Submitted to IMA J. Numer. Anal.

Higham:1995:RHA

[Hig95d] Nicholas J. Higham. Review of “Howard Anton and Chris Rorres, Elementary Linear Algebra: Applications Version, Seventh edition, Wiley, New York, 1994.”. IMA Bulletin, 31(7/8):122– 123, 1995.

Higham:1995:SPT

[Hig95e] Nicholas J. Higham. Stability of parallel triangular system solvers. SIAM J. Sci. Comput., 16(2):400–413, March 1995.

Higham:1995:SDP

[Hig95f] Nicholas J. Higham. Stability of the diagonal pivoting method with partial pivoting. Numerical Analysis Report No. 265, Manchester Centre for Computational Mathematics, Manchester, England, July 1995. 17 pp. To appear in SIAM J. Matrix Anal. Appl.

Higham:1995:TMT

[Hig95g] Nicholas J. Higham. The Test Matrix Toolbox for Matlab (version 3.0). Numerical Analysis Report No. 276, Manchester Centre for Computational Mathematics, Manchester, England, September 1995. 70 pp.

Higham:1996:ASN

[Hig96a] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-355-2. xxviii + 688 pp.

Higham:1996:CCL

[Hig96b] Nicholas J. Higham. Commentary on C. Lanczos, “Introduction” to “Tables of Chebyshev Polynomials S_{n}(x) and C_{n}(x)”, Nat. Bur. Stand., Appl. Math. Series 9, 1952, v–xxvi. In M. Chu et al., editors, The Lanczos Collected Papers with Commentaries, page ?? ????, ????, 1996. To appear.

Higham:1996:RAN

[Hig96c] Nicholas J. Higham. Review of “Acta Numerica 1994 and 1995, Cambridge University Press”. IMA Bulletin, 1996. To appear.

Higham:1996:RAB

[Hig96d] Nicholas J. Higham. Review of “Are Magnus Bruaset. A Survey of Preconditioned Iterative Methods, volume 328 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Essex, UK, 1995.”. IMA Bulletin, 1996. To appear.

Hilbert:1894:BTL

[Hil94] David Hilbert. Ein Beitrag zur Theorie des Legendre’schen Polynoms. Acta Mathematica, 18:155–159, 1894. CODEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Hill:1929:CAA

[Hil29] L. S. Hill. Cryptography in an algebraic alphabet. Amer. Math. Monthly, 36: 306–312, 1929.

Hildebrand:1974:INA

[Hil74] F. B. Hildebrand. Introduction to Numerical Analysis. McGraw-Hill, New York, second edition, 1974. ISBN 0-486-65363-3. xiii + 669 pp. Reprinted by Dover, New York, 1987.

Hills:1987:PP

[Hil87] Philip J. Hills, editor. Publish or Perish. Peter Francis Publishers, Berrycroft, Cambs., 1987. ISBN 1-870167-00-7. ???? pp.

Hill:1988:ECM

[Hil88] D. R. Hill. Experiments in Computational Matrix Algebra. Random House, New York, NY, USA, 1988. ???? pp.

Hill:1989:UPB

[Hil89] R. J. Hill. Using personal bibliographic data bases to keep up with the engineering literature. IEEE Trans. Prof. Commun., PC-32:189–193, 1989.

Hill:1995:SDP

[Hil95] Theodore P. Hill. The significant-digit phenomenon. Amer. Math. Monthly, 102 (4):322–327, 1995.

Hinch:1977:SB

[Hin77] D. Hinch. The Scrabble Book. Pan Books, ????, 1977. ???? pp.

Hirschman:1965:SRC

[Hir65] I. I. Hirschman, Jr. Studies in Real and Complex Analysis. Mathematical Association of America and Prentice-Hall, ????, 1965. ???? pp.

Hockney:1981:PCA

[HJ81] R. W. Hockney and C. R. Jesshope. Parallel Computers: Architecture, Programming and Algorithms. Adam Hilger, Bristol, 1981. ISBN 0-85274-752-7. xiii + 423 pp.

Horn:1985:MA

[HJ85] Roger A. Horn and Charles R. Johnson. Matrix Analysis. Cambridge University Press, ????, 1985. ISBN 0-521-30586-1. xiii + 561 pp.

Hockney:1988:PCA

[HJ88] Roger W. Hockney and Chris R.

Jesshope. Parallel Computers: Architecture, Programming, and Algorithms. Adam Hilger Ltd., Bristol, UK, second edition, 1988. ISBN 0-85274-811-6 (hardcover), 0-85274-812-4 (paperback). xv + 625 pp. LCCN QA76.5 .H57 1988.

Horn:1991:TMA

[HJ91] Roger A. Horn and Charles R. Johnson. Topics in Matrix Analysis. Cambridge University Press, ????, 1991. ISBN 0-521-30587-X. viii + 607 pp.

Hall:1965:CAC

[HK65] M. Hall, Jr. and D. E. Knuth. Combinatorial analysis and computers. Amer. Math. Monthly, 72:21–28, 1965.

Hopcroft:1971:MNM

[HK71] J. E. Hopcroft and L. R. Kerr. On minimizing the number of multiplications necessary for matrix multiplication. SIAM J. Appl. Math., 20(1):30–36, 1971.

Hanson:1987:ATA

[HK87] R. J. Hanson and F. T. Krogh. Algorithm 653. translation of algorithm 539: PC-BLAS, basic linear algebra subprograms for FORTRAN usage with the IN-TEL 8087, 80287 numeric data processor. ACM Transactions on Mathematical Software, 13(3):311–317, 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hewer:1988:SSL

[HK88] Gary Hewer and Charles Kenney. The sensitivity of the stable Lyapunov equation. SIAM Journal on Control and Optimization, 26(2):321–344, 1988.

Higham:1993:CEA

[HK93a] Nicholas J. Higham and Philip A. Knight. Componentwise error analysis for stationary iterative methods. In Carl D. Meyer and Robert J. Plemmons, editors, Linear Algebra, Markov Chains, and Queueing Models, volume 48 of IMA Volumes in Mathematics and its Applications, pages 29–46. Springer-Verlag, New York, 1993.

Higham:1993:FPB

[HK93b] Nicholas J. Higham and Philip A. Knight. Finite precision behavior of stationary iteration for solving singular systems. Linear Algebra and Appl., 192:165– 186, 1993.

Higham:1995:MPF

[HK95] Nicholas J. Higham and Philip A. Knight. Matrix powers in finite precision arithmetic. SIAM J. Matrix Anal. Appl., 16(2):343–358, April 1995.

Hoffman:1986:NFL

[HL86] W. Hoffman and W. M. Lioen. NUMVEC Fortran library manual, chapter: Simultaneous equations. Technical report, Report NM-R8614, Centre for Mathematics and Computer Science, Amsterdam, ????, 1986. ???? pp.

Hardy:1952:I

[HLP52] G. H. Hardy, J. E. Littlewood, and G. Pólya. Inequalities. Cambridge University Press, ????, second edition, 1952. xii + 324 pp.

Heath:1986:CSV

[HLPW86] M. T. Heath, Alan J. Laub, C. C.

Paige, and R. C. Ward. Computing the singular value decomposition of a product of two matrices. SIAM J. Sci. Stat. Comput., 7(4):1147–1159, 1986.

Hughes:1983:EEI

[HLW83] T. J. R. Hughes, I. Levit, and J. Winget. Element-by-element implicit algorithms for heat conduction. J. Eng. Mech., 109:576–585, 1983. CODEN JEN-MDT. ISSN 0733-9399 (print), 1943-7889 (electronic).

Horn:1990:ACS

[HM90a] R. A. Horn and R. Mathias. An analog of the Cauchy–Schwarz inequality for Hadamard products and unitarily invariant norms. SIAM J. Matrix Anal. Appl., 11:481–498, 1990.

Horn:1990:CSI

[HM90b] R. A. Horn and R. Mathias.

Cauchy–Schwarz inequalities associated with positive semidefinite matrices. Linear Algebra and Appl., 142:63–82, 1990.

Hoskins:1977:NSa

[HMW77a] W. D. Hoskins, D. S. Meek, and D. J. Walton. The numerical solution of A′Q + QA = −C. IEEE Trans. Automat. Control., AC-22:882–883, 1977.

Hoskins:1977:NSb

[HMW77b] W. D. Hoskins, D. S. Meek, and D. J. Walton. The numerical solution of Ẋ = A_{1}X + XA_{2} + D, X(0) = C. IEEE Trans. Automat. Control., AC-22: 881–882, 1977.

Hoskins:1977:NSM

[HMW77c] W. D. Hoskins, D. S. Meek, and D. J. Walton. The numerical solution of the matrix equation XA + AY = F. BIT, 17:184–190, 1977. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hanson:1981:AMB

[HN81] Richard J. Hanson and Michael J. Norris. Analysis of measurements based on the singular value decomposition. SIAM J. Sci. Stat. Comput., 2(3):363– 373, 1981.

Hodges:1983:ATE

[Hod83] Andrew Hodges. Alan Turing: The Enigma. Burnett Books, London, 1983. ISBN 0-09-911641-3. xix + 586 pp. 1992 edition with preface, Vintage, London.

Hodel:1992:RAL

[Hod92] A. Scottedward Hodel. Recent applications of the Lyapunov equation in control theory. In R. Beauwens and P. de Groen, editors, Iterative Methods in Linear Algebra, pages 217–227. Elsevier North Holland, Inc., New York, NY, USA, 1992.

Hoenig:1986:STT

[Hoe86] A. Hoenig. Scientific typesetting with TEX. SIAM News, pages 14–16, January 1986. ISSN 0036-1437.

Hoenig:1987:RCT

[Hoe87] A. Hoenig. Review of Computers and Typesetting by D. E. Knuth. BYTE Magazine, 12:67–78, 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hofstadter:1985:MMM

[Hof85a] Douglas R. Hofstadter. Metafont, metamathematics, and metaphysics: Comments on Donald Knuth’s article “the concept of a meta-font”. In Metamagical Themas: Questing for the Essence of Mind and Pattern, pages 483–491. Bantam Books, New York, NY, USA, 1985.

Hofstadter:1985:MTQ

[Hof85b] Douglas R. Hofstadter. Metamagical Themas: Questing for the Essence of Mind and Pattern. Bantam Books, New York, NY, USA, 1985. ISBN 0-553-34279-7. xxviii + 852 pp.

Hofstadter:1985:RAT

[Hof85c] Douglas R. Hofstadter. Review of “Alan Turing: The Enigma”. In Metamagical Themas: Questing for the Essence of Mind and Pattern, pages 483–491. Bantam Books, New York, NY, USA, 1985.

Hoffmann:1987:SLS

[Hof87] W. Hoffmann. Solving linear systems on a vector computer. J. Comp. Appl. Math., 18:353–367, 1987.

Hoffmann:1989:PAR

[Hof89a] Christoph M. Hoffmann. The problems of accuracy and robustness in geometric computation. Computer, pages 31–41, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoffmann:1989:IAG

[Hof89b] W. Hoffmann. Iterative algorithms for Gram–Schmidt orthogonalization. Computing, 41:335–348, 1989. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Holland:1973:BB

[Hol73] Tim Holland. Beginning Backgammon. W.H. Allen, London, 1973. ???? pp.

Holmes:1974:TBW

[Hol74a] Olive Holmes. Thesis to book: What to get rid of. Scholarly Publishing, 6(1): 40–50, 1974. ISSN 0036-634X.

Holmes:1974:BAN

[Hol74b] R. B. Holmes. Best approximation by normal operators. J. Approximation Theory, 12:412–417, 1974. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Holmes:1975:TBW

[Hol75] Olive Holmes. Thesis to book: What to do with what is left. Scholarly Publishing, 6(2):165–176, 1975. ISSN 0036-634X.

Holmes:1980:PS

[Hol80] Olive Holmes. The pinyin syndrome. Scholarly Publishing, 11(3):221– 227, 1980. ISSN 0036-634X.

Holbrook:1992:SVN

[Hol92] J. A. Holbrook. Spectral variation of normal matrices. Linear Algebra and Appl., 174:131–144, 1992.

Homer:1992:BVC

[Hom92] Steve Homer. Battling on with veteran computers. New scientist, pages 32– 35, 1992. CODEN NWSCAL. ISSN 0262-4079, 0028-6664. 14 November.

Honey:1991:DAM

[Hon91] John Honey. Does Accent Matter?

Faber and Faber, London, 1991. ???? pp.

Hoover:1980:EST

[Hoo80] H. Hoover. Essentials for the Scientific and Technical Writer. Dover, New York, NY, USA, 1980. ???? pp.

Hopper:1981:FB

[Hop81] Grace Murray Hopper. The first bug.

Annals of the History of Computing, 3(3): 285–286, 1981.

Hopkins:1982:SCP

[Hop82] T. R. Hopkins. On the sensitivity of the coefficients of padè approximants with respect to their defining power series coefficients. J. Comp. Appl. Math., 8 (2):105–109, 1982.

Horn:1954:EMP

[Hor54] A. Horn. On the eigenvalues of a matrix with prescribed singular values. Proc. Amer. Math. Soc., 5:4–7, 1954.

Horning:1979:NPR

[Hor79] Jim Horning. Note on program reliability. ACM Sigsoft, 4(4):6, 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). Cited in [Tro84a].

Hotelling:1943:SNM

[Hot43] Harold Hotelling. Some new methods in matrix calculation. Annals of mathematical statistics, 14(1):1–34, 1943. ISSN 0003-4851.

Householder:1950:SNM

[Hou50] A. S. Householder. Some numerical methods for solving systems of linear equations. Amer. Math. Monthly, 57:453– 459, 1950.

Householder:1953:PNA

[Hou53] Alston S. Householder. Principles of Numerical Analysis. McGraw-Hill, New York, 1953. ISBN 0-486-61116-7. x + 274 pp. Reprinted by Dover, New York, 1974.

Householder:1958:GER

[Hou58a] Alston S. Householder. Generated error in rotational tridiagonalization. J. Assoc. Comput. Mach., 5:335–338, 1958.

Householder:1958:UTN

[Hou58b] Alston S. Householder. Unitary triangularization of a nonsymmetric matrix. J. Assoc. Comput. Mach., 5:339– 342, 1958.

Householder:1964:TMN

[Hou64] Alston S. Householder. The Theory of Matrices in Numerical Analysis. Blaisdell, New York, 1964. ISBN 0-486-61781-5. xi + 257 pp. Reprinted by Dover, New York, 1975.

Hough:1981:API

[Hou81] David Hough. Applications of the proposed IEEE 754 standard for floatingpoint arithmetic. Computer, 14:70–74, 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

HoughtonMifflin:1982:AHD

[Hou82] Houghton Mifflin. The American Heritage Dictionary. Houghton-Mifflin, Boston, second college edition, 1982. ???? pp. 1985 printing.

Hough:1989:RS

[Hou89] David Hough. Random story. NA Digest, 89(1):??, 1989. Electronic mail mag-

azine: na.help@na-net.ornl.gov.

Howland:1983:FWA

[How83a] J. L. Howland. Further ways to approximate the exponential of a matrix. Linear and Multilinear Algebra, 14:121– 129, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Howland:1983:SMS

[How83b] James Lucien Howland. The sign matrix and the separation of matrix eigenvalues. Linear Algebra and Appl., 49: 221–232, 1983.

Houp:1977:RTI

[HP77] Kenneth W. Houp and Thomas E.

Pearsall. Reporting Technical Information. Glencino Publishing (Encino, CA 91316) and Collier Macmillan (London), ????, third edition, 1977. ???? pp.

Harrod:1984:CSD

[HP84] W. J. Harrod and R. J. Plemmons.

Comparison of some direct methods for computing stationary distributions of Markov chains. SIAM J. Sci. Stat. Comput., 5(2):453–469, 1984.

Hinrichsen:1986:SRL

[HP86a] D. Hinrichsen and A. J. Pritchard.

Stability radii of linear systems. Systems and Control Letters, 7:1–10, 1986.

Hoenig:1986:TTE

[HP86b] A. Hoenig and M. Pfeffer. TEX text editors for the IBM PC. TUGboat, 7:51– 54, 1986. ISSN 0896-3207.

Hopkins:1988:NMP

[HP88] Tim Hopkins and Chris Phillips. Numerical Methods in Practice: Using the NAG Library. Addison-Wesley, Wokingham, England, 1988. ISBN 0-201-19248-9. ???? pp.

Hennessy:1990:CAQ

[HP90] John L. Hennessy and David A. Patterson. Computer Architecture: A Quantitative Approach. Morgan Kaufmann, San Mateo, CA, USA, 1990. ISBN 1-55860-188-0. xxviii + 594 + appendices pp.

Hong:1991:RRQ

[HP91] Y. P. Hong and C.-T. Pan. Rankrevealing QR factorizations and the singular value decomposition. Math. Comp., 197:213–232, 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hodel:1992:PSL

[HP92a] A. Scottedward Hodel and Kameshwar Poolla. Parallel solution of large Lyapunov equations. SIAM J. Matrix Anal. Appl., 13(4):1189–1203, 1992.

Hong:1992:LBS

[HP92b] Y. P. Hong and C.-T. Pan. A lower bound for the smallest singular value. Linear Algebra and Appl., 172:27– 32, 1992.

Higham:1993:PSV

[HP93] Nicholas J. Higham and Pythagoras Papadimitriou. Parallel singular value decomposition via the polar decomposition. Numerical Analysis Report No. 239, Manchester Centre for Computational Mathematics, Manchester, England, October 1993. ???? pp.

Higham:1994:NPA

[HP94a] Nicholas J. Higham and Pythagoras Papadimitriou. A new parallel algorithm for computing the singular value decomposition. In John G. Lewis, editor, Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, pages 80–84. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-336-6.

Higham:1994:PAC

[HP94b] Nicholas J. Higham and Pythagoras Papadimitriou. A parallel algorithm for computing the polar decomposition. Parallel Computing, 20(8):1161–1173, August 1994.

Higham:1994:SPI

[HP94c] Nicholas J. Higham and Alex Pothen.

Stability of the partitioned inverse method for parallel solution of sparse triangular systems. SIAM J. Sci. Comput., 15(1):139–148, January 1994.

Houstis:1988:SFI

[HPP88] E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors. Supercomputing, First International Conference, Athens, Proceedings, volume 297 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1988.

Henderson:1983:HKP

[HPS83] Harold V. Henderson, Friedrich Pukelsheim, and Shayle R. Searle. On the history of the Kronecker product. Linear and Multilinear Algebra, 14:113–120, 1983. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hansen:1990:PES

[HPW90] Eldon R. Hansen, Merrell L.

Patrick, and Richard L. C. Wang. Polynomial evaluation with scaling. ACM Transactions on Mathematical Software, 16(1): 86–93, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Henwood:1979:EMN

[HR80] Mervyn R. Henwood and Ivan Rival. Eponymy in mathematical nomenclature: What’s in a name, and what should be? The Mathematical Intelligencer, 2(4):205–206, 1979/1980. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Heath:1988:PST

[HR88] Michael T. Heath and Charles H. Romine. Parallel solution of triangular systems on distributed-memory multiprocessors. SIAM J. Sci. Stat. Comput., 9(3): 558–588, 1988.

Huitfeldt:1990:NAN

[HR90] J. Huitfeldt and A. Ruhe. A new algorithm for numerical path following applied to an example from hydrodynamical flow. SIAM J. Sci. Stat. Comput., 11(6): 1181–1192, 1990.

Hu:1992:KSM

[HR92] D. Y. Hu and L. Reichel. Krylov subspace methods for the Sylvester equation. Linear Algebra and Appl., 172:283–313, 1992.

Houstis:1990:IMS

[HRV90] E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors. Intelligent Mathematical Software Systems. North-Holland, Amsterdam, The Netherlands, 1990. ISBN 0-444-88682-6.

Houstis:1992:ESS

[HRV92] E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors. Expert Systems for Scientific Computing. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89226-5.

Horn:1959:EUP

[HS59] A. Horn and R. Steinberg. Eigenvalues of the unitary part of a matrix. Pacific J. Math., 9:541–550, 1959.

Hull:1966:TPM

[HS66] T. E. Hull and J. R. Swenson. Tests of probabilistic models for propagation of roundoff errors. Comm. ACM, 9(2):108– 113, 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henderson:1981:DIS

[HS81a] H. V. Henderson and S. R. Searle. On deriving the inverse of a sum of matrices. SIAM Review, 23(1):53–60, 1981.

Henderson:1981:VPM

[HS81b] Harold V. Henderson and S. R.

Searle. The vec-permutation matrix, the vec operator and Kronecker products: A review. Linear and Multilinear Algebra, 9: 271–288, 1981. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Howard:1983:MSR

[HS83] K. Howard and J. A. Sharp. The Management of a Student Research Project. Gower, Aldershot, 1983. ???? pp.

Higham:1987:NLA

[HS87] Nicholas J. Higham and G. W. Stewart. Numerical linear algebra in statistical computing. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 41–57. Oxford University Press, 1987.

Higham:1990:FPD

[HS90] Nicholas J. Higham and Robert S.

Schreiber. Fast polar decomposition of an arbitrary matrix. SIAM J. Sci. Stat. Comput., 11(4):648–655, July 1990.

Hinsley:1993:CIS

[HS93a] F. H. Hinsley and Alan Stripp, editors. Codebreakers: The Inside Story of Bletchley Park. Oxford University Press, ????, 1993. ISBN 0-19-820479-5. xxi + 321 pp.

Huo:1993:EMP

[HS93b] Yan Huo and Robert Schreiber. Efficient, massively parallel eigenvalue computation. Int. J. Supercomputer Appl., 7(4):292–303, 1993. CODEN IJSAE9. ISSN 0890-2720.

Hochbruck:1995:PKS

[HS95] Marlis Hochbruck and Gerhard Starke. Preconditioned Krylov subspace methods for Lyapunov matrix equations. SIAM J. Matrix Anal. Appl., 16(1):156–171, 1995.

Heaton:1987:LDC

[HT87] J. B. Heaton and N. D. Turton. Longman Dictionary of Common Errors. Longman, Harlow, Essex, 1987. ISBN 0-582-96410-5. ???? pp.

Higham:1993:SO

[HT93] Desmond J. Higham and Lloyd N. Trefethen. Stiffness of ODEs. BIT, 33: 285–303, 1993. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Huang:1975:JTM

[Hua75] C. P. Huang. A Jacobi-type method for triangularizing an arbitrary matrix. SIAM J. Numer. Anal., 12:566–570, 1975.

Hua:1990:SSL

[Hua90] D. Hua. On the symmetric solutions of linear matrix equations. Linear Algebra and Appl., 131:1–7, 1990.

Huff:1954:HLS

[Huf54] Darrell Huff. How to Lie with Statistics. Penguin, ????, 1954. ISBN 0-14-021300-7. 124 pp.

Hull:1967:SOM

[Hul67] T. E. Hull. A search for optimum methods for the numerical integration of ordinary differential equations. SIAM Review, 9:647–654, 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1979:CNS

[Hul79] T. E. Hull. Correctness of numerical software. In Lloyd D. Fosdick, editor, Performance Evaluation of Numerical Software, pages 3–15. North-Holland, Amsterdam, The Netherlands, 1979.

Hull:1982:PCE

[Hul82] T. E. Hull. Precision control, exception handling and A choice of numerical algorithms. In G. A. Watson, editor, Numerical Analysis Proceedings, Dundee 1981, volume 912 of Lecture Notes in Mathematics, pages 169–178. Springer-Verlag, Berlin, 1982.

Hume:1988:TTG

[Hum88] A. Hume. A tale of two greps.

Software—Practice and Experience, 18: 1063–1072, 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hurford:1990:MMO

[Hur90] Peter Hurford. Making Music on the Organ. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised edition, 1990. ???? pp.

Henry:1994:PQA

[HV94] Greg Henry and Robert Van de Geijn. Parallelizing the QR algorithm for the unsymmetric algebraic eigenvalue problem: Myths and reality. Technical Report CS-94-244, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, August 1994. ???? pp.

Hamacher:1990:CO

[HVZ90] V. Carl Hamacher, Zvonko G. Vranesic, and Safwat G. Zaky. Computer Organization. McGraw-Hill, New York, third edition, 1990. ISBN 0-07-100742-3. xx + 617 pp.

Hetherington:1975:TTF

[HW75] J. H. Hetherington and F. D. C. Willard. Two-, three-, and four-atom exchange effects in bcc^{3}He. Phys. Rev. Lett, 35(21):1442–1443, 1975. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Hall:1976:MNM

[HW76a] G. Hall and J. M. Watt, editors. Modern Numerical Methods for Ordinary Differential Equations. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1976. ISBN 0-19-853348-9. 336 pp.

Hammarling:1976:PBL

[HW76b] S. J. Hammarling and J. H. Wilkinson. The practical behaviour of linear iterative methods with particular reference to S.O.R. Report NAC 69, National Physical Laboratory, Teddington, UK, September 1976. 19 pp.

Hartnell:1976:PFM

[HW76c] B. L. Hartnell and H. C. Williams, editors. Proc. Fifth Manitoba Conference on Numerical Math., Winnipeg 1975. Utilitas Math. Publ., Winnipeg, Manitoba, Canada, 1976. ???? pp.

Hedayat:1978:HMT

[HW78a] A. Hedayat and W. D. Wallis.

Hadamard matrices and their applications. Annals of Statistics, 6(6):1184– 1238, 1978. ISSN 0090-5364.

Hoskins:1978:FMC

[HW78b] W. D. Hoskins and D. J. Walton.

A faster method of computing the square root of a matrix. IEEE Trans. Automat. Control., AC-23:494–495, 1978.

Hodge:1979:ABR

[HW79a] James O. Hodge and James L. W.

West III. Academic book reviewing: Some problems and suggestions. Scholarly Publishing, 11(1):35–41, 1979. ISSN 0036-634X.

Hoskins:1979:FMS

[HW79b] W. D. Hoskins and D. J. Walton. A faster, more stable method for computing the pth roots of positive definite matrices. Linear Algebra and Appl., 26:139– 163, 1979.

Hammarling:1980:LSA

[HW80] S. J. Hammarling and J. H. Wilkinson. On linear systems arising from finite difference approximations to elliptic differential equations. Report DNACS 34/80, National Physical Laboratory, Teddington, UK, October 1980. ???? pp.

Hayden:1988:AMP

[HW88] T. L. Hayden and J. Wells. Approximation by matrices positive semidefinite on a subspace. Linear Algebra and Appl., 109:115–130, 1988.

Hairer:1991:SOD

[HW91a] E. Hairer and G. Wanner. Solving Ordinary Differential Equations II. Springer-Verlag, Berlin, 1991. ISBN 3-540-53775-9. xv + 601 pp.

Hubbard:1991:DED

[HW91b] John H. Hubbard and Beverly H. West. Differential Equations. A Dynamical Systems Approach. Part I: Ordinary Differential Equations. Springer-Verlag, Berlin, 1991. ISBN 3-540-97286-2. xix + 348 pp.

He:1995:ACN

[HW95] Chunyang He and G. A. Watson. An algorithm for computing the numerical radius. Manuscript, 1995.

Hwang:1995:WAL

[Hwa95] Andrew D. Hwang. Writing in the age of LaTEX. Notices Amer. Math. Soc., 42(8):878–882, 1995. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harper:1991:GCA

[HWH91] David Harper, Chris Wooff, and David Hodgkinson. A Guide to Computer Algebra Systems. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-92910-7. ???? pp.

Hageman:1981:AIM

[HY81] L. A. Hageman and D. Young. Applied Iterative Methods. Academic Press, New York, USA, 1981. ???? pp.

Hammel:1988:NOC

[HYG88] Stephen M. Hammel, James A.

Yorke, and Celso Grebogi. Numerical orbits of chaotic processes represent true orbits. Bull. Amer. Math. Soc., 19(2):465– 469, 1988.

Hyman:1957:EEG

[Hym57] M. A. Hyman. Eigenvalues and eigenvectors of general matrices. Presented at the 12th National Meeting of the Association for Computing Machinery, Houston, Texas, 1957. Cited in [Wil63b].

Hyman:1986:DHP

[Hym86] Dick Hyman. Dick Hyman’s Professional Chord Changes and Substitutions for 100 Tunes Every Musician Should Know. Ekay Music, Inc, 223 Katonah Ave, Katonah, NY 10536, 1986. ISBN 0-943748-14-3. ???? pp.

IBM:1988:ESS

[IBM88] IBM. Engineering and Scientific Subroutine Library, Guide and Reference, Release 3. IBM Corporation, fourth edition (program number 5668-863) edition, 1988. ??–?? pp.

IBM:1993:AVR

[IBM93] IBM Corporation, San Jose, CA, USA. AIX Version 3.2 for RISC System/6000: Optimization and Tuning Guide for Fortran, C, and C++, December 1993. viii + 305 pp. Publication No. SC09-1705-00.

Group:1981:PSB

[IEE81] IEEE Computer Society Microprocessor Standards Committee, Floating-Point Working Group. A proposed standard for binary floating-point arithmetic, Draft 8.0 of IEEE Task P754 (with introductory comments by David Stevenson). Computer, 14:51–62, 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

IEEE:1985:ISB

[IEE85] IEEE Standard for Binary Floating-Point Arithmetic, ANSI/IEEE Standard 754-1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ???? pp. Reprinted in SIGPLAN Notices, 22(2):9– 25, 1987.

IEEE:1987:RSF

[IEE87] A Radix-Independent Standard for Floating-Point Arithmetic, IEEE Standard 854-1987. IEEE Computer Society, New York, 1987. ???? pp.

Saussy:1984:PDC

[III84] G. S. Saussy III. The Penguin Dictionary of Curious and Interesting Words. Penguin, ????, 1984. ???? pp.

Ikebe:1987:MTC

[IIM87] Yasuhiko Ikebe, Toshiyuki Inagaki, and Sadaaki Miyamoto. The monotonicity theorem, Cauchy’s interlace theorem, and the Courant–Fischer theorem. Amer. Math. Monthly, 94(4):352–354, 1987.

Isaacson:1966:ANM

[IK66] Eugene Isaacson and Herbert Bishop Keller. Analysis of Numerical Methods. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1966. ISBN 0-486-68029-0. xv + 541 pp. Reprinted by Dover, New York, 1994.

Ikebe:1979:IHM

[Ike79] Yasuhiko Ikebe. On inverses of Hessenberg matrices. Linear Algebra and Appl., 24:93–97, 1979.

Committee:1993:RGL

[ILA93] ILAS Education Committee. Report on graduate linear algebra courses. Manuscript from the International Linear Algebra Society, November 1993. URL

= http://gauss.technion.ac.il/iic/
GRAD-ED.SYLLABI.

Ipsen:1994:USM

[IM94] Ilse C. F. Ipsen and Carl D. Meyer.

Uniform stability of Markov chains. SIAM J. Matrix Anal. Appl., 15(4):1061– 1074, 1994.

Incertis:1983:ENF

[Inc83] F. Incertis. An extension of a new formulation of the algebraic Riccati equation problem. IEEE Trans. Automat. Control., AC-28(2):235–238, 1983.

Incertis:1984:SSF

[Inc84] F. Incertis. A skew-symmetric formulation of the algebraic Riccati equation problem. IEEE Trans. Automat. Control., AC-29(5):467–470, 1984.

Incertis:1985:FMC

[Inc85] F. Incertis. A faster method of computing matrix Pythagorean sums. IEEE Trans. Automat. Control., AC-30(3):273– 275, 1985.

Ince:1992:CWM

[Inc92] D. C. Ince, editor. Collected Works of A. M. Turing: Mechanical Intelligence. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88058-5. xix + 227 pp.

Ingleby:1995:LAW

[Ing95] N. Bruce Ingleby. Linear algebra in weather forecasting: Combining observations and forecasts. Presented at the IMA Conference on Linear Algebra and Its Applications, University of Manchester, July 1995.

Investigators:1993:IRT

[Inv93] The GUSTO Investigators. An international randomized trial comparing four thrombolytic strategies for acute myocardial infarction. The New England Journal of Medicine, 329(10):673–682, 1993. CO-DEN NEJMAG. ISSN 0028-4793 (print), 1533-4406 (electronic).

Iohvidov:1982:HTM

[Ioh82] I. S. Iohvidov. Hankel and Toeplitz Matrices and Forms: Algebraic Theory. Birkhäuser Boston Inc., Cambridge, MA, USA, 1982. ???? pp.

Ip:1987:LCS

[Ip87a] C. M. Ip. On least-change secant updates in factored form. SIAM J. Numer. Anal., 24:1126–1132, 1987.

Iserles:1987:SAN

[IP87b] A. Iserles and M. J. D. Powell, editors. The State of the Art in Numerical Analysis. Oxford University Press, ????, 1987. ISBN 0-19-853614-3.

Ipsen:1994:HWC

[Ips94a] Ilse C. F. Ipsen. Helmut Wielandt’s contributions to the numerical solution of complex eigenvalue problems. In B. Huppert and Schneider H, editors, Helmut Wielandt, Mathematische Werke, Mathematical Works, Volume II: Matrix Theory and Analysis, page ?? Walter de Gruyter, Berlin, 1994. To appear.

Ipsen:1994:HII

[Ips94b] Ilse C. F. Ipsen. A history of inverse iteration. In B. Huppert and Schneider H, editors, Helmut Wielandt, Mathematische Werke, Mathematical Works, Volume II: Matrix Theory and Analysis, page ?? Walter de Gruyter, Berlin, 1994. To appear.

Iri:1991:HAD

[Iri91] Masao Iri. History of automatic differentiation and rounding error estimation. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, pages 3–16. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991.

Iserles:1992:AN

[Ise92] A. Iserles, editor. Acta Numerica 1992.

Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41026-6. 407 pp. LCCN QA297.A21 1992.

Ibbett:1989:AHP

[IT89] R. N. Ibbett and N. P. Topham. Architecture of High Performance Computers, Volume I: Uniprocessors and Vector Processors. MacMillan Publishing Company, New York, NY, USA, 1989. ???? pp.

Iri:1988:ACP

[ITH88] M. Iri, T. Tsuchiya, and M. Hoshi. Automatic computation of partial derivatives and rounding error estimates with applications to large-scale systems of nonlinear equations. J. Comp. Appl. Math., 24:365–392, 1988.

Iverson:1962:PL

[Ive62] Kenneth E. Iverson. A Programming Language. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1962. ???? pp.

Jacobson:1953:LAA

[Jac53] N. Jacobson. Lectures in Abstract Algebra, volume 2. Van Nostrand, Princeton, N.J., 1953. ???? pp.

Jacobi:1976:WAW

[Jac76] Ernst Jacobi. Writing at Work: DOs, DON’Ts and HOW TOs. Hayden, New Jersey, 1976. ???? pp.

Jacobs:1977:SAN

[Jac77] David A. H. Jacobs, editor. The State of the Art in Numerical Analysis. Academic Press, London, 1977. ISBN 0-12-378650-9.

Jacobs:1978:NSN

[Jac78] D. A. H. Jacobs, editor. Numerical Software—Needs and Availability. Academic Press, London, 1978.

Jackson:1986:BHS

[Jac86] Donald D. Jackson. A brief history of scholarly publishing. In Scherr [Sch86a], pages 197–?? ISBN 0-930376-42-0. LCCN Q167.J682 1986.

JaJa:1992:IPA

[JáJ92] Joseph JáJá. An Introduction to Parallel Algorithms. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54856-9. x + 566 pp.

James:1974:PIC

[Jam74] R. D. James. Proc. International Congress of Mathematicians, Vancouver 1974. Canadian Mathematical Congress, 1974. ???? pp.

Jackson:1991:GRR

[JBNP91] R. H. F. Jackson, P. T. Boggs, S. G.

Nash, and S. Powell. Guidelines for reporting results of computational experiments. Report of the ad hoc committee. Math. Prog., 49:413–425, 1991. CODEN MHPGA4. ISSN 0025-5610.

Jacoby:1970:BB

[JC70] Oswald Jacoby and John. R. Crawford. The Backgammon Book. Pan Books, ????, 1970. ???? pp.

Jungong:1995:ERP

[JE95] Xue Jungong and Jian Erxiong. Entrywise relative perturbation theory for nonsingular M-matrices and applications. BIT, 35:417–427, 1995. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jeffreys:1960:MFL

[Jef60] Alan E. Jeffreys. Michael Faraday: A List of His Lectures and Published Writings. Chapman and Hall, Ltd., London, UK, 1960. xxviii + 86 pp.

Jennings:1977:MAM

[Jen77a] A. Jennings. Matrices, ancient and modern. IMA Bulletin, 13:117–123, 1977.

Jennings:1977:MCE

[Jen77b] A. Jennings. Matrix Computation for Engineers and Scientists. Wiley, Chichester, UK, 1977. ???? pp.

Jennings:1982:BSV

[Jen82] A. Jennings. Bounds for the singular values of a matrix. IMA J. Numer. Anal., 2:459–474, 1982.

Jennings:1983:BSE

[Jen83] A. Jennings. On bounds for symmetric eigenvalue problems. In Kågström and Ruhe [KR83], pages 195–205.

Jessup:1993:CAD

[Jes93] E. R. Jessup. A case against a divide and conquer approach to the nonsymmetric eigenvalue problem. Applied Numerical Mathematics, 12:403–420, 1993.

Jones:1993:HGG

[JHL93] Simon L. Peyton Jones, John Hughes, and John Launchbury. How to give a good research talk. ACM SIG-PLAN Notices, 28(11):??, 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jessup:1992:IAI

[JI92] Elizabeth R. Jessup and Ilse C. F. Ipsen. Improving the accuracy of inverse iteration. SIAM J. Sci. Stat. Comput., 13 (2):550–572, 1992.

Jennings:1992:MC

[JM92] Alan Jennings and K. J. McKeown. Matrix Computation. Wiley, Chichester, UK, second edition, 1992. ISBN 0-471-93527-1. ???? pp. £19.95.

Johnson:1994:AIM

[JMV94] Jeffrey Johnson, Sean McKee, and Alfred Vella, editors. Artificial Intelligence in Mathematics. Oxford University Press, ????, 1994. ISBN 0-19-853686-0.

Jennings:1974:DEA

[JO74] L. S. Jennings and M. R. Osborne. A direct error analysis for least squares. Numer. Math., 22:325–332, 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:1978:NDF

[Joh78] Charles R. Johnson. Numerical determination of the field of values of a general complex matrix. SIAM J. Numer. Anal., 15(3):595–602, 1978.

Johnson:1982:IM

[Joh82] C. R. Johnson. Inverse M-matrices.

Linear Algebra and Appl., 47:195–216, 1982.

Jones:1870:LLM

[Jon70] Henry Bence Jones. The Life and Letters of Michael Faraday. Longmans, Green and Co., London, second, revised edition, 1870. ???? pp. In Two Volumes.

Jones:1962:ANA

[Jon62] Terence G. Jones. An algorithm for the numerical application of a linear operator. J. Assoc. Comput. Mach., 9:440– 449, 1962.

Jones:1974:RWR

[Jon74] Robert Jones. Rights, wrongs and referees. New scientist, 61(890):758–759, 1974. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Jonckheere:1984:NBS

[Jon84] Edmond A. Jonckheere. New bound on the sensitivity of the solution of the Lyapunov equation. Linear Algebra and Appl., 60:57–64, 1984.

Jordan:1968:EEG

[Jor68] T. L. Jordan. Experiments on error growth associated with some linear leastsquares procedures. Math. Comp., 22: 579–588, 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Joseph:1991:CPN

[Jos91] George Gheverghese Joseph. The Crest of the Peacock: Non-European Roots of Mathematics. Penguin, ????, 1991. ISBN 0-14-012529-9. xv + 371 pp.

Jalby:1990:SAI

[JP90] W. Jalby and B. Philippe. Stability analysis and improvement of the block gram–schmidt algorithm. Technical report, Publication Interne No. 509, Institute de Recherche en Informatique et Systèmes Aléatoires, Rennes, ????, 1990. ???? pp.

Jalby:1991:SAI

[JP91] William Jalby and Bernard Philippe. Stability analysis and improvement of the block Gram–Schmidt algorithm. SIAM J. Sci. Stat. Comput., 12(5):1058–1073, 1991.

Jones:1993:BKF

[JP93] Mark T. Jones and Merrell L. Patrick. Bunch–Kaufman factorization for real symmetric indefinite banded matrices. SIAM J. Matrix Anal. Appl., 14(2):553– 559, 1993.

Jones:1994:FSI

[JP94] Mark T. Jones and Merrell L. Patrick.

Factoring symmetric indefinite matrices on high-performance architectures. SIAM J. Matrix Anal. Appl., 15(1):273–283, 1994.

James:1975:CCS

[JR75] K. R. James and W. Riha. Convergence criteria for successive overrelaxation. SIAM J. Numer. Anal., 12(2):137– 143, 1975.

Johnson:1982:NA

[JR82] Lee W. Johnson and R. Dean Riess.

Numerical Analysis. Addison-Wesley, Reading, MA, USA, second edition, 1982. ISBN 0-201-10392-3. xii + 563 pp.

Jackson:1977:VMC

[JS77] L. W. Jackson and A. E. Sedgwick.

Vandermonde matrices, coordinate transformations and Adams’ methods. Report TR77-5, Department of Computer Science, University of Alberta, Canada, 1977. ???? pp.

Jones:1983:DEM

[JS83] D. S. Jones and B. D. Sleeman. Differential Equations and Mathematical Biology. George Allen and Unwin, London, 1983. ???? pp.

Jessup:1992:PAC

[JS92] E. R. Jessup and D. C. Sorensen. A parallel algorithm for computing the singular value decomposition of a matrix. Technical Report CU-CS-623-92, University of Colorado at Boulder, ????, October 1992. ???? pp. To appear in SIAM J. Matrix Anal. Appl.

Jankowski:1983:NFP

[JSW83] M. Jankowski, A. Smoktunowicz, and H. Woźniakowski. A note on floatingpoint summation of very many terms. J. Information Processing and Cybernetics-EIK, 19(9):435–440, 1983. CODEN JI-CYE5. ISSN 0863-0593.

Jones:1974:NSE

[JT74] William B. Jones and W. J. Thron. Numerical stability in evaluating continued fractions. Math. Comp., 28(127):795– 810, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jenkins:1975:PTP

[JT75] M. A. Jenkins and J. F. Traub. Principles for testing polynomial zerofinding programs. ACM Transactions on Mathematical Software, 1(1):26–34, 1975. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Judd:1990:CPG

[Jud90] Karen Judd. Copyediting: A Practical Guide. Crisp Publications, Los Altos, California, second edition, 1990. ISBN 1-56052-143-0. ???? pp.

Jungong:1995:CSE

[Jun95] Xue Jungong. Computing the smallest eigenvalue of an M-matrix. Submitted to SIAM J. Matrix Anal. Appl., 1995.

Justus:1972:TMT

[Jus72] P. E. Justus. There is more to typesetting than setting type. IEEE Trans. Prof. Commun., PC-15:13–16, 1972.

Jensen:1974:PUM

[JW74] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ???? pp.

Jankowski:1977:IRI

[JW77] M. Jankowski and H. Woźniakowski.

Iterative refinement implies numerical stability. BIT, 17:303–311, 1977. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jankowski:1985:ASC

[JW85a] M. Jankowski and H. Woźniakowski.

The accurate solution of certain continuous problems using only single precision arithmetic. BIT, 25:635–651, 1985. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jensen:1985:PUM

[JW85b] Kathleen Jensen and Niklaus Wirth.

Pascal User Manual and Report. Revised for the ISO Pascal Standard. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1985. ???? pp. Revised by Andrew B. Mickel and James F. Milner.

Jansen:1986:HAA

[JW86] Paul Jansen and Peter Weidner. Highaccuracy arithmetic software—some tests of the ACRITH problem-solving routines. ACM Transactions on Mathematical Software, 12(1):62–70, 1986. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kac:1966:COH

[Kac66] Marc Kac. Can one hear the shape of a drum? Amer. Math. Monthly, 73(4, Part II):1–23, 1966. (4, Part II), Slaught Papers No. 11.

Kaagstroem:1977:BPB

[Kåg77a] B. Kågström. Bounds and perturbation bounds for the matrix exponential. BIT, 17:39–57, 1977. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Kaagstroem:1977:NCM

[Kåg77b] B. Kågström. Numerical computation of matrix functions. Technical report, Report UMINF-58.77, Department of Information Processing, University of Umeå, Sweden, ????, 1977. ???? pp.

Kaagstroem:1984:GSV

[Kåg84] B. Kågström. The generalized singular value decomposition and the general (A − λB)-problem. BIT, 24:568–583, 1984. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Kaagstroem:1993:DMR

[Kåg93] Bo Kågström. A direct method for reordering eigenvalues in the generalized real Schur form of a regular matrix pair (A, B). In Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors, Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E, pages 195–218. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993.

Kaagstroem:1994:PAG

[Kåg94] Bo Kågström. A perturbation analysis of the generalized Sylvester equation (AR − LB, DR − LE) = (C, F). SIAM J. Matrix Anal. Appl., 15(4):1045–1060, 1994.

Kahan:1965:FRR

[Kah65] W. Kahan. Further remarks on reducing truncation errors. Comm. ACM, 8(1):40, 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahan:1966:AES

[Kah66a] W. Kahan. Accurate eigenvalues of a symmetric tri-diagonal matrix. Technical Report No. CS41, Department of Computer Science, Stanford University, ????, 1966. ???? pp. Revised June 1968.

Kahan:1966:NLA

[Kah66b] W. M. Kahan. Numerical linear algebra. Canadian Math. Bulletin, 9(??): 757–801, ???? 1966. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kahan:1970:ACM

[Kah70] W. M. Kahan. Almost-commuting matrices are near commuting matrices. 1970.

Kahan:1971:P

[Kah71] W. Kahan. A problem. ACM SIGNUM Newsletter, 6(3):6, 1971. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1972:CCC

[Kah72a] W. Kahan. Conserving confluence curbs ill-condition. Technical Report 6, Computer Science Department, University of California, Berkeley, August 1972. ???? pp.

Kahan:1972:SEA

[Kah72b] W. Kahan. A survey of error analysis. In Proc. IFIP Congress, Ljubljana, Information Processing 71, pages 1214– 1239. North-Holland, Amsterdam, The Netherlands, 1972.

Kahan:1973:EMR

[Kah73a] W. Kahan. Every n × n matrix Z with real spectrum satisfies ∥Z − Z^{∗}∥ ≤ ∥Z + Z^{∗}∥(log_{2} n + 0.038). Proc. Amer. Math. Soc., 39:235–241, 1973.

Kahan:1973:IAL

[Kah73b] W. Kahan. Implementation of algorithms (lecture notes by W. S. Haugeland and D. Hough). Technical Report 20, Department of Computer Science, University of California, Berkeley, CA, USA, 1973. ???? pp.

Kahan:1975:SNH

[Kah75] W. Kahan. Spectra of nearly Hermitian matrices. Proc. Amer. Math. Soc., 48:11–17, 1975.

Kahan:1977:NSC

[Kah77] W. Kahan. And now for something completely different: The Texas Instruments SR-52. Memorandum No. UCB/ERL M77/23, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, April 1977. ???? pp.

Kahan:1979:PCK

[Kah79] William M. Kahan. Personal calculator has key to solve any equation f(x) = 0. Hewlett–Packard Journal, pages 20–26, December 1979.

Kahan:1980:IAO

[Kah80a] W. Kahan. Interval arithmetic options in the proposed IEEE floating point arithmetic standard. In Karl L. E. Nickel, editor, Interval Mathematics 1980, pages 99–128. Academic Press, New York, 1980.

Kahan:1980:SPI

[Kah80b] W. Kahan. Software \sqrt{x} for the proposed IEEE floating-point standard. Manuscript, August 1980.

Kahan:1980:HCE

[Kah80c] William M. Kahan. Handheld calculator evaluates integrals. Hewlett– Packard Journal, pages 23–32, August 1980.

Kahan:1981:WDW

[Kah81] W. Kahan. Why do we need a floating-point arithmetic standard? Technical report, University of California, Berkeley, CA, USA, February 1981. 41 pp.

Kahan:1983:MWS

[Kah83a] W. Kahan. Mathematics written in sand—the HP-15C, Intel 8087, etc. Technical report, Statistical Computing Section, Proceedings of the American Statistical Association, Toronto 1983, ????, 1983. ???? pp.

Kahan:1983:PIS

[Kah83b] W. Kahan. The proposed IEEE Standard p754 for floating-point arithmetic: What good is it? Manuscript, 1983.

Kahaner:1984:SIQ

[Kah84] D. Kahaner. Sources of information on quadrature software. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 134– 164. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Kahan:1986:DVB

[Kah86a] W. Kahan. Decimal versus binary floating-point arithmetic. Manuscript, November 1986.

Kahan:1986:RAF

[Kah86b] W. Kahan. Rational arithmetic in floating-point. Technical Report PAM-343, Center for Pure and Applied Mathematics, University of California, Berkeley, September 1986. 8 pp.

Kahan:1986:RPE

[Kah86c] W. Kahan. Roundoff in polynomial evaluation. Manuscript, 1986.

Kahan:1986:SRC

[Kah86d] W. Kahan. To solve a real cubic equation. Technical Report PAM-352, Center for Pure and Applied Mathematics, University of California, Berkeley, CA, USA, November 1986. 20 pp.

Kahan:1987:BCC

[Kah87a] W. Kahan. Branch cuts for complex elementary functions or much ado about nothing’s sign bit. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 165–211. Oxford University Press, 1987.

Kahan:1987:DPI

[Kah87b] W. Kahan. Doubled-precision IEEE standard 754 floating-point arithmetic. Manuscript, February 1987.

Kahan:1987:HSM

[Kah87c] W. Kahan. How should Max and Min be defined? Manuscript, August

Kahan:1987:PCF

[Kah87d] W. Kahan. Presubstitution, and continued fractions. Manuscript, March 1987.

Kahan:1988:CPA

[Kah88] W. Kahan. A computer program with almost no significance. Manuscript, November 1988.

Kahan:1989:CAA

[Kah89a] W. Kahan. Calculating area and angle of a needle-like triangle. Manuscript, January 1989.

Kahan:1989:OFH

[Kah89b] W. Kahan. The occasional futility of higher-precision arithmetic. Manuscript, September 1989.

Kahan:1989:SRN

[Kah89c] W. Kahan. A singularity removable numerically when arithmetic is clean. Manuscript, May 1989.

Kahan:1990:BPA

[Kah90a] W. Kahan. Better to prescribe arithmetic than describe it. Technical report, Turing Award Lecture presented at ACM Conference, Washington, ????, 1990. ???? pp.

Kahan:1990:HCA

[Kah90b] W. Kahan. How Cray’s arithmetic hurts scientific computation (and what might be done about it). Manuscript prepared for the Cray User Group meeting in Toronto, June 1990.

Kahan:1990:POC

[Kah90c] W. Kahan. Paradoxes in our concepts of computational accuracy. Manuscript, October 1990.

Kahan:1990:TCC

[Kah90d] W. Kahan. Twenty challenges for computerized symbolic algebra systems. Manuscript, August 1990.

Kahan:1991:ARL

[Kah91a] W. Kahan. Analysis and refutation of the LCAS. ACM SIGNUM Newsletter, 26(3):2–15, 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1991:CMA

[Kah91b] W. Kahan. Contracted multiplyadds. Manuscript, September 1991.

Kahan:1991:VTC

[Kah91c] W. Kahan. Validating/testing/confirming/checking/correcting numerical software. Manuscript, November 1991.

Kahan:1992:FE

[Kah92a] W. Kahan. Floating-point exceptionhandling. Manuscript, July 1992.

Kahan:1992:NLI

[Kah92b] W. Kahan. Notes on Laguerre’s iteration. Manuscript, December 1992.

Kahan:1992:RCM

[Kah92c] W. Kahan. Roundoff in complex multiplication. Manuscript, 1992.

Kahan:1992:WN

[Kah92d] W. Kahan. What numbers?

Manuscript, September 1992.

Kahan:1993:AAS

[Kah93a] W. Kahan. Analysis and application of simply compensated summation. Manuscript, September 1993.

Kahan:1993:TIE

[Kah93b] W. Kahan. A tridiagonal inverse eigenproblem and the travails of a wouldbe portable programmer. Manuscript, June 1993.

Kahan:1994:CBV

[Kah94a] W. Kahan. Computer benchmarks versus accuracy. Draft manuscript, June 1994.

Kahan:1994:AS

[Kah94b] W. Kahan. On the advantages of the 8087’s stack. Manuscript, 1994.

Kahan:1994:W

[Kah94c] W. Kahan. When is x ∗ [1/x] < x/x = 1? Manuscript, January 1994.

Kahan:1995:LNS

[Kah95] W. Kahan. Lecture notes on the status of IEEE Standard 754 for binary floating-point arithmetic. Manuscript, May 1995.

Karlin:1968:TP

[Kar68] Samuel Karlin. Total Positivity, volume 1. Stanford University Press, Stanford, CA, USA, 1968. ???? pp.

Karasalo:1974:CTD

[Kar74] Ilkka Karasalo. A criterion for truncation of the QR-decomposition algorithm for the singular linear least squares problem. BIT, 14:156–166, 1974. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Karpinski:1985:PFP

[Kar85] Richard Karpinski. Paranoia: A floating-point benchmark. BYTE Magazine, 10(2):223–235, February 1985. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Karlson:1991:SSR

[Kar91] Rune Karlson. A study of some roundoff effects of the GMRES-method. Report LiTH-MAT-R-1990-11, Department of Mathematics, Linköping University, Sweden, 1991. ???? pp.

Kasenally:1995:GGM

[Kas95] Ebrahim M. Kasenally. GMBACK: A generalised minimum backward error algorithm for nonsymmetric linear systems. SIAM J. Sci. Comput., 16(3):698– 719, 1995.

Kato:1960:EIM

[Kat60] Tosio Kato. Estimation of iterated matrices, with application to the von Neumann condition. Numer. Math., 2:22– 29, 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kato:1976:PTL

[Kat76] Tosio Kato. Perturbation Theory for Linear Operators. Springer-Verlag, Berlin, second edition, 1976. ISBN 3-540-97588-5 (??invalid checksum??). xxi + 619 pp.

Katz:1985:ESP

[Kat85] Michael J. Katz. Elements of the Scientific Paper: A Step-by-Step Guide for Students and Professionals. Yale University Press, New Haven, CT, USA, 1985. ISBN 0-300-03532-2. ???? pp.

Kaufman:1978:FCB

[Kau78] Roger Emanuel Kaufman. A FOR-TRAN Coloring Book. MIT Press, Cambridge, MA, USA, 1978. ISBN 0-262-61026-4. ???? pp.

Kaufman:1983:MMQ

[Kau83] Linda Kaufman. Matrix methods for queuing problems. SIAM J. Sci. Stat. Comput., 4(3):525–552, 1983.

Kaufman:1987:GHT

[Kau87] Linda Kaufman. The generalized Householder transformation and sparse matrices. Linear Algebra and Appl., 90: 221–234, 1987.

Kelly:1992:XTF

[KB92] Nickolas J. Kelly and Christian H.

Bischof. XBibTEX and friends. TUGboat, 13(4):443–446, 1992. ISSN 0896-3207.

Khalfan:1990:TES

[KBS90] H. Khalfan, R. H. Byrd, and R. B.

Schnabel. A theoretical and experimental study of the symmetric rank one update. Technical report, Report CU-CS-489-90, Department of Computer Science, University of Colorado, Boulder, ????, 1990. ???? pp.

Koc:1994:CMS

[KBS94] Çetin K. Koç, Bertan Bakkaloğlu, and Leang S. Shieh. Computation of the matrix sign function using continued fraction expansion. IEEE Trans. Automat. Control., AC-39(8):1644–1647, 1994.

Kuki:1973:SSA

[KC73] H. Kuki and W. J. Cody. A statistical study of the accuracy of floating point number systems. Comm. ACM, 16(4): 223–230, 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kernighan:1975:STM

[KC75] Brian W. Kernighan and Lorinda L.

Cherry. A system for typesetting mathematics. Comm. ACM, 18(3):151–157, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahan:1982:NOS

[KC82] W. Kahan and Jerome T. Coonen. The near orthogonality of syntax, semantics, and diagnostics in numerical programming environments. In J. K. Reid, editor, The Relationship Between Numerical Computation and Programming Languages, pages 103–115. North-Holland, Amsterdam, The Netherlands, 1982. ISBN 0-444-86377-X.

Kincaid:1991:NA

[KC91] D. Kincaid and W. Cheney. Numerical Analysis. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1991. ???? pp.

Kopka:1993:GLD

[KD93] Helmut Kopka and Patrick W. Daly. A Guide to LaTEX: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Wokingham, England, 1993. ISBN 0-201-56889-6. ???? pp.

Kopka:1995:GLD

[KD95] Helmut Kopka and Patrick W. Daly. A Guide to LaTEX2ε: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Wokingham, England, second edition, 1995. ISBN 0-201-42777-X. x + 554 pp.

Kautsky:1982:CWI

[KE82] J. Kautsky and S. Elhay. Calculation of the weights of interpolatory quadratures. Numer. Math., 40:407–422, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Keener:1993:PFT

[Kee93] James P. Keener. The Perron– Frobenius theorem and the ranking of football teams. SIAM Review, 35(1):80– 93, 1993.

Keller:1962:FML

[Kel62] Joseph B. Keller. Factorization of matrices by least-squares. Biometrika, 49(1-2):239–242, 1962. ISSN 0006-3444 (print), 1464-3510 (electronic).

Keller:1965:SSS

[Kel65] Herbert B. Keller. On the solution of singular and semidefinite linear systems by iteration. SIAM J. Numer. Anal., 2 (2):281–290, 1965.

Keller:1975:CUO

[Kel75] J. B. Keller. Closest unitary, orthogonal and Hermitian operators to a given operator. Mathematics Magazine, 48: 192–197, 1975. CODEN MAMGA8. ISSN 0025-570X.

Kelly:1985:ICD

[Kel85] Kevin Kelly. Information as a Communicable Disease, page ?? In Essays of an Information Scientist: 1984 [Gar85], 1985.

Kelly:1988:SCT

[Kel88] Kevin Kelly, editor. SIGNAL: Communication Tools for the Information Age. Harmony Books, Crown Publishers, 225 Park Ave. South, New York, NY 10003, 1988. ISBN 0-517-57084-X. ???? pp.

Kelley:1991:WM

[Kel91] J. L. Kelley. Writing mathematics. In John H. Ewing and F. W. Gehring, editors, Paul Halmos: Celebrating 50 Years of Mathematics, pages 91–96. Springer-Verlag, Berlin, 1991.

Keller:1993:NMT

[Kel93] Herbert B. Keller. Numerical Methods for Two-Point Boundary-Value Problems. Dover, New York, NY, USA, 1993. ISBN 0-486-66925-4. ???? pp. Corrected and expanded republication of work first published by Blaisdell, Waltham, MA, 1964.

Kemperman:1982:HMI

[Kem82] J. H. B. Kemperman. A Hurwitz matrix is totally positive. SIAM J. Math. Anal., 13(2):331–341, 1982.

Kenny:1982:HPS

[Ken82] Peter Kenny. A Handbook of Public Speaking for Scientists and Engineers. Adam Hilger, Bristol, 1982. ISBN 0-85274-553-2. ???? pp.

Kenney:1985:EFM

[Ken85] Alice P. Kenney. The epidemiology of footnote-in-mouth disease. Scholarly Publishing, ??:263–267, 1985. ISSN 0036-634X.

KSR:1991:KFP

[Ken91a] Kendall Square Research Corporation, Waltham, MA. KSR Fortran Programming, 1991. ???? pp.

KSR:1991:KPP

[Ken91b] Kendall Square Research Corporation, Waltham, MA. KSR Parallel Programming, 1991. ???? pp.

KSR:1993:KBL

[Ken93a] Kendall Square Research Corporation, Waltham, MA. KSRlib/BLAS Library, Version 1.0, Installation Guide and Release Notes, 1993. ???? pp.

KSR:1993:KLL

[Ken93b] Kendall Square Research Corporation, Waltham, MA. KSRlib/LAPACK Library, Version 1.0b BETA, Installation Guide and Release Notes, 1993. ???? pp.

Kershaw:1969:EIT

[Ker69] D. Kershaw. The explicit inverses of two commonly occurring matrices. Math. Comp., 23(105):189–191, 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kernighan:1975:RPR

[Ker75] Brian W. Kernighan. RATFOR—A preprocessor for a rational Fortran. Software—Practice and Experience, 5:395– 406, 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kerkut:1983:CTP

[Ker83] G. A. Kerkut. Choosing a title for a paper. Comp. Biochem. Physiol., 47A(1): 1, 1983. CODEN CPCPAI. ISSN 0010-406x (print), 1879-3029 (electronic).

Kernighan:1984:WPM

[Ker84] Brian W. Kernighan. Why Pascal is not my favorite programming language. In A. R. Feuer and N. Gehani, editors, Comparing and Assessing Programming Languages: Ada, C, and Pascal, pages 170–186. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Kerner:1989:LSC

[Ker89] W. Kerner. Large-scale complex eigenvalue problems. J. Comput. Phys., 85:1–85, 1989.

Kerr:1990:FCT

[Ker90] Thomas H. Kerr. Fallacies in computational testing of matrix positive definiteness/semidefiniteness. IEEE Trans. Aerospace Electronic Systems, 26(2):415– 421, 1990.

Kershaw:1992:VTG

[Ker92] D. Kershaw. Variations on the theme of Gaussian Elimination. Journal of the Operational Research Society, 43(8):821– 827, 1992.

Kerrigan:1993:MF

[Ker93] James F. Kerrigan. Migrating to Fortran 90. O’Reilly & Associates, Sebastopol, CA, 1993. ISBN 1-56592-049-X. xxvi + 361 pp.

Kahan:1963:ACC

[KF63a] W. Kahan and I. Farkas. Algorithm 167: Calculation of confluent divided differences. Comm. ACM, 6(4):164–165, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahan:1963:ANIa

[KF63b] W. Kahan and I. Farkas. Algorithm 168: Newton interpolation with backward divided differences. Comm. ACM, 6(4): 165, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahan:1963:ANIb

[KF63c] W. Kahan and I. Farkas. Algorithm 169: Newton interpolation with forward divided differences. Comm. ACM, 6(4): 165, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Keast:1987:NIR

[KF87] P. Keast and G. Fairweather. Numerical Integration: Recent Developments, Software and Applications. D. Reidel Publishing Company, Dordrecht, Holland, 1987. ???? pp.

Karp:1990:MPP

[KF90] Alan H. Karp and Horace P. Flatt.

Measuring parallel processor performance. Comm. ACM, 33(5):539–543, 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kennedy:1980:SC

[KG80] William J. Kennedy, Jr. and James E.

Gentle. Statistical Computing. Marcel Dekker, Inc., New York, NY, USA, 1980. ISBN 0-8247-6898-1. xi + 591 pp.

Kautsky:1983:CJM

[KG83] J. Kautsky and G. H. Golub. On the calculation of Jacobi matrices. Linear Algebra and Appl., 52/53:439–455, 1983.

Kenney:1990:SAD

[KH90a] Charles Kenney and Gary Hewer.

The sensitivity of the algebraic and differential Riccati equations. SIAM Journal on Control and Optimization, 28(1): 50–69, 1990.

Kincaid:1990:IML

[KH90b] D. R. Kincaid and L. J. Hayes. Iterative Methods for Large Linear Systems. Academic Press, New York, USA, 1990. ???? pp.

Kenney:1995:TNB

[KH95] Charles Kenney and Gary Hewer.

Trace norm bounds for stable Lyapunov operators. Linear Algebra and Appl., 22: 1–18, 1995.

Khalil:1972:EBT

[Kha72] H. M. Khalil. The eigenproblem of block tridiagonal matrices. Comm. ACM, 15(9):839, 1972. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Kielbasinski:1973:SAC

[Kie73] Andrzej Kiełbasiński. Summation algorithm with corrections and some of its applications. Math. Stos., 1:22–41, 1973. (In Polish, cited in [JSW83] and [JW85a]).

Kielbasinski:1981:IRL

[Kie81] Andrzej Kiełbasiński. Iterative refinement for linear systems in variableprecision arithmetic. BIT, 21:97–103, 1981. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Kielbasinski:1987:NRE

[Kie87] Andrzej Kiełbasiński. A note on rounding-error analysis of Cholesky factorization. Linear Algebra and Appl., 88/89:487–494, 1987.

King:1967:MWNa

[Kin67a] Lester S. King. Medical writing number 1: Monotony. J. Amer. Medical Assoc., 201(6):362–363, 1967.

King:1967:MWNb

[Kin67b] Lester S. King. Medical writing number 3: Prepositions. J. Amer. Medical Assoc., 201(10):765–766, 1967.

King:1967:MWNc

[Kin67c] Lester S. King. Medical writing number 5: Verbs. J. Amer. Medical Assoc., 202(1):45–46, 1967.

King:1967:MWNd

[Kin67d] Lester S. King. Medical writing number 7: The opening sentence. J. Amer. Medical Assoc., 202(6):535–536,

Kirkman:1992:GSW

[Kir92] John Kirkman. Good Style: Writing for Science and Technology. E & FN Spon (Chapman and Hall), London, 1992. ISBN 0-419-17190-8. ???? pp.

Kittel:1971:ISS

[Kit71] Charles Kittel. Introduction to Solid State Physics. Wiley, New York, NY, USA; London, UK; Sydney, Australia, fourth edition, 1971. ISBN 0-471-49021-0. xv + 766 pp.

Kittaneh:1985:LFN

[Kit85] Fuad Kittaneh. On lipschitz functions of normal operators. Proc. Amer. Math. Soc., 94(3):416–418, 1985.

Kittaneh:1995:SVC

[Kit95] Fuad Kittaneh. Singular values of companion matrices and bounds on zeros of polynomials. SIAM J. Matrix Anal. Appl., 16(1):333–340, 1995.

Kemeny:1980:BP

[KK80] J. G. Kemeny and T. E. Kurtz. BA-SIC Programming. John Wiley, New York, NY, USA, third edition, 1980. ???? pp.

Klatte:1993:CXC

[KKL^{+}93] R. Klatte, U. W. Kulisch, C. Lawo, M. Rauch, and A. Wiethoff. C-XSC: A C++ Class Library for Extended Scientific Computing. Springer-Verlag, Berlin, 1993. ISBN 0-387-56328-8. ???? pp.

Kailath:1979:DRM

[KKM79] T. Kailath, S.-Y. Kung, and M. Morf. Displacement ranks of matrices and linear equations. Journal of Mathematical Analysis and Applications, 68:

395–407, 1979. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Klatte:1992:PXL

[KKN^{+}92] R. Klatte, U. W. Kulisch, M. Neaga, D. Ratz, and Ch. Ullrich. PASCAL-XSC—Language Reference With Examples. Springer-Verlag, Berlin, 1992. ???? pp.

Kaneko:1973:LRE

[KL73] Toyohisa Kaneko and Bede Liu. On local roundoff errors in floating-point arithmetic. J. Assoc. Comput. Mach., 20(3): 391–398, 1973.

Klema:1980:SVD

[KL80] Virginia C. Klema and Alan J. Laub. The singular value decomposition: Its computation and some applications. IEEE Trans. Automat. Control., AC-25 (2):164–176, 1980.

Kahan:1985:AIA

[KL85] W. Kahan and E. LeBlanc. Anomalies in the IBM ACRITH package. In Kai Hwang, editor, Proceedings of the 7th Symposium on Computer Arithmetic, pages 322–331. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985.

Kenney:1988:CSR

[KL88] Charles Kenney and Alan J. Laub. Controllability and stability radii for companion form systems. Math. Control, Signals and Sys., 1:239–256, 1988.

Kaagstroem:1989:LBR

[KL89a] B. Kågström and P. Ling. Level 2 and 3 BLAS routines for IBM 3090 VF:

Implementations and experiences. In Vector and Parallel Computing [DDGM89], pages 215–228.

Kenney:1989:CEM

[KL89b] Charles Kenney and Alan J. Laub.

Condition estimates for matrix functions. SIAM J. Matrix Anal. Appl., 10(2):191– 209, 1989.

Kenney:1989:PEE

[KL89c] Charles Kenney and Alan J. Laub.

Padé error estimates for the logarithm of a matrix. Internat. J. Control, 50(3):707– 730, 1989.

Kenney:1991:PDM

[KL91a] Charles Kenney and Alan J. Laub.

Polar decomposition and matrix sign function condition estimates. SIAM J. Sci. Stat. Comput., 12:488–504, 1991.

Kenney:1991:RIM

[KL91b] Charles Kenney and Alan J. Laub.

Rational iterative methods for the matrix sign function. SIAM J. Matrix Anal. Appl., 12(2):273–291, 1991.

Kenney:1992:SNM

[KL92] Charles Kenney and Alan J. Laub. On scaling Newton’s method for polar decomposition and the matrix sign function. SIAM J. Matrix Anal. Appl., 13(3):688– 706, 1992.

Kenney:1993:HTI

[KL93] Charles S. Kenney and Alan J. Laub.

A hyperbolic tangent identity and the geometry of Padé sign function iterations. Preprint, Department of Electrical and Computer Engineering, University of California, Santa Barbara, July 1993.

Kenney:1994:SSS

[KL94] Charles S. Kenney and Alan J. Laub. Small-sample statistical condition estimates for general matrix functions. SIAM J. Stat. Comput., 15(1):36–61, 1994.

Kenney:1995:MSF

[KL95] Charles S. Kenney and Alan J. Laub. The matrix sign function. IEEE Transactions on Automatic Control, 40(8):1330– 1348, 1995.

Klare:1963:MR

[Kla63] George R. Klare. The Measurement of Readability. Iowa State University Press, Ames, Iowa, 1963. ???? pp.

Kenney:1989:PNS

[KLJ89] Charles Kenney, Alan J. Laub, and E. A. Jonckheere. Positive and negative solutions of dual Riccati equations by matrix sign function iteration. Systems and Control Letters, 13:109–116, 1989.

Kenney:1992:MSAa

[KLP92a] Charles S. Kenney, Alan J. Laub, and Philip M. Papadopoulos. Matrix sign algorithms for Riccati equations. IMA J. of Math. Control and Information, 9:331– 344, 1992.

Kenney:1992:MSAb

[KLP92b] Charles S. Kenney, Alan J. Laub, and Philip M. Papadopoulos. Matrix sign algorithms for Riccati equations. In Proceedings of IMA Conference on Control: Modelling, Computation, Information, pages 1–10. The Institute of Mathematics and its Applications, Southend-On-Sea, Essex, UK, 1992.

Kenney:1993:NSA

[KLP93] Charles Kenney, Alan J. Laub, and P. M. Papadopoulos. A Newton-squaring algorithm for computing the negative invariant subspace of a matrix. IEEE Transactions on Automatic Control, 38 (8):1284–1289, 1993.

Knuth:1989:MW

[KLR89] Donald E. Knuth, Tracy Larrabee, and Paul M. Roberts. Mathematical Writing. MAA Notes Number 14. Mathematical Association of America, Washington, DC, USA, 1989. ISBN 0-88385-063-X. ???? pp. Also Report STAN-CS-88-1193, Department of Computer Science, Stanford University, January 1988.

Koelbel:1994:HPF

[KLS^{+}94] Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele Jr., and Mary E. Zosel. The High Performance Fortran Handbook. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-11185-3. xiv + 329 pp.

Kaagstroem:1991:HPG

[KLV91] Bo Kågström, Per Ling, and Charles F. Van Loan. High performance GEMM-based level-3 BLAS: Sample routines for double precision data. In M. Durand and F. El Dabaghi, editors, High Performance Computing II, pages 269– 281. North-Holland, Amsterdam, The Netherlands, 1991.

Kenney:1989:SES

[KLW89] Charles Kenney, Alan J. Laub, and M. Wette. A stability-enhancing scaling procedure for Schur-Riccati solvers. Math. Control, Signals and Sys., 12:241– 250, 1989.

Kenney:1990:EBN

[KLW90] Charles Kenney, Alan J. Laub, and M. Wette. Error bounds for Newton refinement of solutions to algebraic Riccati equations. Math. Control, Signals and Sys., 3:211–224, 1990.

Khatri:1976:HND

[KM76] C. G. Khatri and S. K. Mitra. Hermitian and nonnegative definite solutions of linear matrix equations. SIAM J. Appl. Math., 31:579–585, 1976.

Kulisch:1981:CAT

[KM81] Ulrich W. Kulisch and Willard L. Miranker. Computer Arithmetic in Theory and in Practice. Academic Press, New York, 1981. ISBN 0-12-428650-X. xiii + 249 pp.

Kulisch:1983:NAS

[KM83] Ulrich W. Kulisch and Willard L. Miranker, editors. A New Approach to Scientific Computation. Academic Press, New York, 1983. ISBN 0-12-428660-7.

Krishnakumar:1986:EST

[KM86a] A. S. Krishnakumar and M. Morf. Eigenvalues of a symmetric tridiagonal matrix: A divide-and-conquer approach. Numer. Math., 48:349–368, 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kulisch:1986:ADC

[KM86b] Ulrich W. Kulisch and Willard L. Miranker. The arithmetic of the digital computer: A new approach. SIAM Review, 28(1):1–40, 1986.

Kahaner:1989:NMS

[KMN89] David K. Kahaner, Cleve B. Moler, and Stephen G. Nash. Numerical Methods and Software. Prentice-Hall, Englewood Cliffs, NJ, USA, 1989. ISBN 0-13-627258-4. xii + 495 pp.

Knight:1993:EAS

[Kni93] Philip A. Knight. Error Analysis of Stationary Iteration and Associated Problems. PhD thesis, University of Manchester, Manchester, England, September 1993. 135 pp.

Knight:1995:FRM

[Kni95] Philip A. Knight. Fast rectangular matrix multiplication and QR decomposition. Linear Algebra and Appl., 221:69– 81, 1995.

Knuth:1962:EPC

[Knu62] Donald E. Knuth. Evaluation of polynomials by computer. Comm. ACM, 5 (12):595–599, December 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [?].

Knuth:1971:ESF

[Knu71] Donald E. Knuth. An empirical study of FORTRAN programs. Software—Practice and Experience, 1:105– 133, 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knuth:1974:SPG

[Knu74a] Donald E. Knuth. Structured programming with go to statements. ACM Computing Surveys, 6(4):261–301, 1974. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Knuth:1974:SN

[Knu74b] Donald E. Knuth. Surreal Numbers. Addison-Wesley, Reading, MA, USA, 1974. ISBN 0-201-03812-9. 119 pp.

Knuth:1979:MT

[Knu79] Donald E. Knuth. Mathematical typography. Bull. (New Series) AMS, 1(2): 337–372, 1979. ISSN 0273-0979 (print), 1088-9485 (electronic).

Knuth:1973:ACPa

[Knu81a] Donald E. Knuth. The Art of Computer Programming. Addison-Wes-ley, Reading, MA, USA, 1973–1981. ???? pp. Three volumes.

Knuth:1981:ACP

[Knu81b] Donald E. Knuth. The Art of Computer Programming, Volume 2, Seminumerical Algorithms. Addison-Wesley, Reading, MA, USA, second edition, 1981. ISBN 0-201-03822-6. xiii + 688 pp.

Knuth:1984:CS

[Knu84a] Donald E. Knuth. The complexity of songs. Comm. ACM, 27(4):344–346, April 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See erratum in [?].

Knuth:1984:LP

[Knu84b] Donald E. Knuth. Literate programming. The Computer Journal, 27: 97–111, 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Knuth:1985:ATM

[Knu85a] Donald E. Knuth. Algorithmic thinking and mathematical thinking. Amer. Math. Monthly, 92(3):170–181, 1985.

Knuth:1985:RF

[Knu85b] Donald E. Knuth. Recipes and fractions. TUGboat, 6:36–38, 1985. ISSN 0896-3207.

Knuth:1986:TP

[Knu86a] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986.

Knuth:1986:MP

[Knu86b] Donald E. Knuth. : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:1986:TB

[Knu86c] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13448-9. ix + 483 pp. LCCN Z253.4.T47 K58 1986.

Knuth:1986:CMT

[Knu86d] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:1986:LMM

[Knu86e] Donald E. Knuth. The Book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13444-6. xi + 361 pp. LCCN Z250.8.M46 K58 1986.

Knuth:1987:FDH

[Knu87a] Donald E. Knuth. Fonts for digital halftones. TUGboat, 8:135–160, 1987. ISSN 0896-3207.

Knuth:1987:MJ

[Knu87b] Donald E. Knuth. Macros for Jill. TUGboat, 8:309–314, 1987. ISSN 0896-3207.

Knuth:1989:ET

[Knu89a] Donald E. Knuth. The errors of TEX. Software—Practice and Experience, 19:607–685, 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Knuth:1989:TCM

[Knu89b] Donald E. Knuth. Typesetting Concrete Mathematics. TUGboat, 10:31–36, 1989. ISSN 0896-3207.

Knuth:1991:BTI

[Knu91] Donald E. Knuth. 3:16 Bible Texts Illuminated. A-R Editions, 801 Deming Way, Madison, Wisconsin, 1991. ISBN 0-89579-252-4. ???? pp.

Knuth:1992:TNN

[Knu92] Donald E. Knuth. Two notes on notation. Amer. Math. Monthly, 99(5):403– 422, 1992.

Knuth:1993:AP

[Knu93] Donald E. Knuth. Artistic programming. Current Contents: Physical, Chemical & Earth Sciences, (34):8, August 23, 1993.

Knuth:1973:ACPb

[Knuta] Donald E. Knuth. The Art of Computer Programming, Volume 1, Fundamental Algorithms. Addison-Wesley, Reading, MA, USA, second edition, 1973. ISBN 0-201-03821-8. xxi + 634 pp.

Karatsuba:1963:MMN

[KO63] A. Karatsuba and Yu. Ofman. Multiplication of multidigit numbers on automata. Soviet Physics—Doklady, 7 (7):595–596, January 1963. CODEN SPHDA9. ISSN 0038-5689.

Kowalik:1968:MUO

[KO68] J. Kowalik and M. R. Osborne. Methods for Unconstrained Optimization Problems. American Elsevier, New York, 1968. ???? pp.

King:1991:FLS

[KO91] K. N. King and A. J. Offutt. A Fortran language system for mutation-based software testing. Software—Practice and Experience, 21:685–718, 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kodak:1979:LSA

[Kod79] Kodak Limited. Let’s stamp out awful lecture slides. Kodak publication S-22(H), ????, April 1979. ???? pp.

Kogbetliantz:1955:SLE

[Kog55] E. G. Kogbetliantz. Solution of linear equations by diagonalization of coefficients matrix. Quarterly of Applied Mathematics, 13:123–132, 1955. CODEN QA-MAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Kogge:1974:PSR

[Kog74] P. M. Kogge. Parallel solution of recurrence problems. IBM Journal of Research and Development, 18(2):138–148, 1974. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kolata:1978:GDE

[Kol78] G. B. Kolata. Geodesy: Dealing with an enormous computer task. Science, 200: 421–466, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Kolotilina:1995:TSB

[Kol95] L. Yu. Kolotilina. Two-sided bounds for the inverse of an H-matrix. Linear Algebra and Appl., 225:117–123, 1995.

Korem:1988:PTP

[Kor88a] D. Korem. Powers: Testing the Psychic and Supernatural. Intervarsity Press, Downers Grove, Illinois, 1988. ???? pp.

Koerner:1988:FA

[Kör88b] T. W. Körner. Fourier Analysis. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-38991-7. xii + 591 pp.

Kostlan:1992:SGM

[Kos92] Eric Kostlan. On the spectra of Gaussian matrices. Linear Algebra and Appl., 162-164:385–388, 1992.

Kostin:1993:LAA

[Kos93] V. I. Kostin. Linear algebra algorithms with guaranteed accuracy. Technical Report TR/PA/93/05, CERFACS, Toulouse, France, 1993. ???? pp.

Kovarik:1970:SIM

[Kov70] Zdislav Kovarik. Some iterative methods for improving orthonormality. SIAM J. Numer. Anal., 7(3):386–389, 1970.

Kovarik:1976:CAS

[Kov76] Z. V. Kovarik. Compatibility of approximate solutions of inaccurate linear equations. Linear Algebra and Appl., 15: 217–225, 1976.

Kovarik:1977:MCS

[Kov77] Z. V. Kovarik. Minimal compatible solutions of linear equations. Linear Algebra and Appl., 17:95–106, 1977.

Kozen:1992:DAA

[Koz92] D. C. Kozen. The Design and Analysis of Algorithms. Springer-Verlag, Berlin, 1992. ???? pp.

Kernighan:1974:PSE

[KP74] Brian W. Kernighan and P. J.

Plauger. Programming style: Examples and counterexamples. ACM Computing Surveys, 6:303–319, 1974. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Kernighan:1976:ST

[KP76] Brian W. Kernighan and P. J.

Plauger. Software Tools. Addison-Wes-ley, Reading, MA, USA, 1976. ISBN 0-201-03669-X. ???? pp.

Kahan:1977:CYC

[KP77] W. Kahan and B. N. Parlett. Can you count on your calculator? Memorandum No. UCB/ERL M77/21, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, April 1977. ???? pp.

Kernighan:1978:EPS

[KP78] Brian W. Kernighan and P. J.

Plauger. The Elements of Programming Style. McGraw-Hill, New York, second edition, 1978. ???? pp.

Kahan:1979:PFP

[KP79] W. Kahan and J. Palmer. On a proposed floating-point standard. ACM SIGNUM Newsletter, 14:13–21, October 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knuth:1981:BPL

[KP81] D. E. Knuth and M. F. Plass. Breaking paragraphs into lines. Software— Practice and Experience, 11:1119–1184, 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kaagstroem:1992:DSM

[KP92] Bo Kågström and Peter Poromaa. Distributed and shared memory block algorithms for the triangular Sylvester equation with sep^{−1} estimators. SIAM J. Matrix Anal. Appl., 13(1):90–101, 1992.

Kaagstroem:1993:LSA

[KP93] Bo Kågström and Peter Poromaa. LAPACK-style algorithms and software for solving the generalized Sylvester equation and estimating the separation between regular matrix pairs. Report UMINF 93.23, Institute of Information Processing, University of Umeå, Umeå, Sweden, December 1993. 35 pp. LAPACK Working Note 75.

Kaagstroem:1994:CES

[KP94] Bo Kågström and Peter Poromaa. Computing eigenspaces with specified eigenvalues of a regular matrix pair (A, B) and condition estimation: Theory, algorithms and software. Report UMINF 94.04, Institute of Information Processing, University of Umeå, Umeå, Sweden, September 1994. 65 pp. LAPACK Working Note 87.

Kahan:1982:RBA

[KPJ82] W. Kahan, B. N. Parlett, and E. Jiang. Residual bounds on approximate eigensystems of nonnormal matrices. SIAM J. Numer. Anal., 19(3):470– 484, 1982.

Kuck:1977:ARM

[KPS77] David J. Kuck, Parker, Douglass S., Jr., and Ahmed H. Sameh. Analysis of rounding methods in floating-point arithmetic. IEEE Trans. Comput., C-26(7): 643–650, 1977.

Keller:1965:CSM

[KR65] H. B. Keller and E. L. Reiss. Computers in solid mechanics: A case history. Amer. Math. Monthly, 72:92–98, 1965.

Kleinman:1978:EBS

[KR78] David L. Kleinman and P. Krishna Rao. Extensions to the Bartels–Stewart algorithm for linear matrix equations. IEEE Trans. Automat. Control., AC-23 (1):85–87, 1978.

Kaagstroem:1980:AJA

[KR80a] B. Kågström and A. Ruhe. ALGO-RITHM 560 JNF: An algorithm for numerical computation of the Jordan normal form of a complex matrix. ACM Transactions on Mathematical Software, 6:437–443, 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaagstroem:1980:ANC

[KR80b] B. Kågström and A. Ruhe. An algorithm for numerical computation of the Jordan normal form of a complex matrix. ACM Transactions on Mathematical Software, 6:398–419, 1980. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaagstroem:1983:MPP

[KR83] B. Kågström and A. Ruhe, editors. Matrix Pencils, Proceedings, 1982, volume 973 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1983.

Krantz:1993:HTM

[Kra93] Steven G. Krantz. How To Teach Mathematics: A Personal Perspective. American Mathematical Society, Providence, RI, USA, 1993. ISBN 0-8218-0197-X. ???? pp.

Kreisselmeier:1972:SBM

[Kre72] G. Kreisselmeier. A solution of the bilinear matrix equation AY + Y B = −Q. SIAM J. Appl. Math., 23(3):334– 338, 1972.

Kreczmar:1976:MRS

[Kre76] Antoni Kreczmar. On memory requirements of Strassen’s algorithms. In Mazurkiewicz [Maz76], pages 404–407. ISBN 0-387-07854-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4 S989 1976. “The Symposium is being organized by the Computation Centre of the Polish Academy of Sciences in cooperation with the University of Gdansk.”.

Krishnaiah:1980:MAV

[Kri80] P. R. Krishnaiah, editor. Multivariate Analysis—V. North-Holland, Amsterdam, The Netherlands, 1980.

Kronsjoe:1987:ATC

[Kro87] L. Kronsjö. Algorithms: Their Complexity and Efficiency. John Wiley, New York, NY, USA, second edition, 1987. ???? pp.

Krol:1992:WIU

[Kro92] Ed Krol. The Whole Internet User’s Guide & Catalog. O’Reilly & Associates, Sebastopol, CA, USA, 1992. ISBN 1-56592-025-2. ???? pp.

Krol:1994:WIU

[Kro94] Ed Krol. The Whole Internet User’s Guide & Catalog. O’Reilly & Associates, Sebastopol, CA, USA, second edition, 1994. ISBN 1-56592-063-5. xxv + 543 pp.

Kielbasinski:1988:NLA

[KS88] Andrzej Kiełbasiński and Hubert Schwetlick. Numerische Lineare Algebra: Eine Computerorientierte Einführung. VEB Deutscher, Berlin, 1988. ISBN 3-87144-999-7. 472 pp.

Kielbasinski:1992:NAL

[KS92] Andrzej Kiełbasiński and Hubert Schwetlick. Numeryczna Algebra Liniowa: Wprowadzenie do Obliczeń Zautomatyzowanych. Wydawnictwa Naukowo-Techniczne, Warszawa, 1992. ISBN 83-204-1260-9. 502 pp.

Kreyszig:1976:TD

[KT76] E. Kreyszig and J. Todd. QR in two dimensions. Elemente der Mathematik, 31:109–114, 1976. CODEN ELMMAF. ISSN 0013-6018 (print), 1420-8962 (electronic).

Koshy:1985:URP

[KT85] M. Koshy and R. P. Tewarson. On the use of restricted pseudo-inverses for the unified derivation of quasi-Newton updates. IMA J. Numer. Anal., 5:141–151, 1985.

Kahan:1991:APL

[KT91] W. Kahan and J. W. Thomas. Augmenting a programming language with complex arithmetic. Manuscript, November 1991.

Kac:1968:MLR

[KU68] Marc Kac and Stanislaw Ulam. Mathematics and Logic: Retrospect and Prospects. Penguin, ????, 1968. ISBN 0-14-021253-1. 204 pp.

Kublanovskaja:1962:SAS

[Kub62] V. N. Kublanovskaja. On some algorithms for the solution of the complete eigenvalue problem. U.S.S.R. Computational Math. and Math. Phys., 1:637–657, 1962. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Kublanovskaja:1963:CIP

[Kub63] V. N. Kublanovskaja. On certain iteration processes for the symmetrisation of a matrix. U.S.S.R. Computational Math. and Math. Phys., 2:859–868, 1963. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Kublanovskaja:1965:PIO

[Kub65] V. N. Kublanovskaja. A process for improving the orthogonalization of a vector system. U.S.S.R. Computational Math. and Math. Phys., 5:215–220, 1965. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Kublanovskaya:1970:ASG

[Kub70] V. N. Kublanovskaya. On an approach to the solution of the generalized latent value problem for lambda-matrices. SIAM J. Numer. Anal., 7:532–537, 1970.

Kubota:1991:PFP

[Kub91] Koichi Kubota. PADRE2, a Fortran precompiler yielding error estimates and second derivatives. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, pages 251–262. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991.

Kuperman:1971:ALA

[Kup71] I. B. Kuperman. Approximate Linear Algebraic Equations. Van Nostrand Reinhold Co., New York, NY, USA, 1971. ISBN 0-442-04546-8. xi + 225 pp.

Kurtz:1992:SCA

[Kur92] David C. Kurtz. A sufficient condition for all the roots of a polynomial to be real. Amer. Math. Monthly, 99:259– 263, 1992.

Kuyper:1991:BSA

[Kuy91] Barbara J. Kuyper. Bringing up scientists in the art of critiquing research. BioScience, 41(4):248–250, 1991. CO-DEN BISNAS. ISSN 0006-3568 (print), 1525-3244 (electronic).

Kaagstroem:1991:GBL

[KV91] B. Kågström and C. F. Van Loan.

GEMM-based level-3 BLAS. Technical report, Report CTC91TR47, Cornell Theory Center, Cornell University, ????, 1991. ???? pp.

Kaashoek:1990:SPS

[KvSR90] M. A. Kaashoek, J. H. van Schuppen, and A. C. M. Ran, editors. Signal Processing, Scattering and Operator Theory, and Numerical Methods. Proceedings of the International Symposium MTNS-89, Volume III. Birkhäuser Verlag, Basel, Switzerland, 1990. ???? pp.

Kress:1974:NM

[KvW74] R. Kress, H. L. De vries, and R. Wegmann. On nonnormal matrices. Linear Algebra and Appl., 8:109–120, 1974.

Kahaner:1983:MSB

[KW83] D. K. Kahaner and W. L. Wyman. Mathematical software in basic: Evaluation of definite integrals. IEEE Micro, 3:42–46, 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaagstroem:1989:GSM

[KW89] Bo Kågström and Lars Westin. Generalized Schur methods with condition estimators for solving the generalized Sylvester equation. IEEE Trans. Automat. Control., AC-34(7):745–751, 1989.

Kuczynski:1992:ELE

[KW92] J. Kuczyński and H. Woźniakowski. Estimating the largest eigenvalue by the power and Lanczos algorithms with a random start. SIAM J. Matrix Anal. Appl., 13(4):1094–1122, 1992.

Kincaid:1988:BRI

[KY88] D. R. Kincaid and D. M. Young. A brief review of the ITPACK project. J. Comp. Appl. Math., 24:121–127, 1988.

Kwong:1991:NIP

[KZ91] M. K. Kwong and A. Zettl. Norm inequalities for the powers of a matrix. Amer. Math. Monthly, 98:533–538, 1991.

Lambuth:1964:GBW

[L^{+}64] David Lambuth et al. The Golden Book on Writing. Penguin, ????, 1964. ISBN 0-14-046263-5. ???? pp.

Laasonen:1958:ISM

[Laa58] P. Laasonen. On the iterative solution of the matrix equation AX^{2} − I = 0. M.T.A.C., 12:109–116, 1958.

Laboratory:1957:MCM

[Lab57] National Physical Laboratory. Modern Computing Methods. Number 16 in Notes on Applied Science. Her Majesty’s Stationery Office, London, 1957. vi + 129 pp.

Laderman:1976:NAM

[Lad76] J. D. Laderman. A noncommutative algorithm for multiplying 3 × 3 matrices using 23 multiplications. Bull. Amer. Math. Soc., 82(1):126–128, 1976.

Lagerloef:1974:IRR

[Lag74] R. O. E. Lagerlöf. Interpolation with rounded ramp functions. Comm. ACM, 17:476–479, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lammers:1986:PAW

[Lam86a] S. Lammers. Programmers at Work.

Microsoft Press, Bellevue, WA, USA, 1986. ???? pp.

Lamport:1986:LDP

[Lam86b] Leslie Lamport. LaTEX: a Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-15790-X. xiv + 242 pp. LCCN Z253.4.L38 L35 1986.

Lamport:1987:DPV

[Lam87a] Leslie Lamport. Document production: Visual or logical? Notices Amer. Math. Soc., 34:621–624, 1987. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lamport:1987:MIP

[Lam87b] Leslie Lamport. Makeindex: An index processor for LaTEX. Documentation distributed with MakeIndex program, February 17, 1987.

Lambert:1991:NMO

[Lam91] J. D. Lambert. Numerical Methods for Ordinary Differential Systems: The Initial Value Problem. John Wiley, New York, NY, USA, 1991. ISBN 0-471-92990-5. ???? pp.

Lamport:1994:LDP

[Lam94] Leslie Lamport. LaTEX: a Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-52983-1. xvi + 272 pp. LCCN Z253.4.L38L35 1994.

Lamport:1995:HWP

[Lam95] Leslie Lamport. How to write a proof. Amer. Math. Monthly, 102(7):600– 608, 1995.

Lanczos:1938:TIE

[Lan38] C. Lanczos. Trigonometric interpolation of empirical and analytical functions. J. Math. Phys., 17:123–199, 1938. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lanczos:1956:AA

[Lan56] Cornelius Lanczos. Applied Analysis. Prentice-Hall, Englewood Cliffs, NJ, USA, 1956. ISBN 0-486-65656-X. xx + 539 pp. Reprinted by Dover, New York, 1988.

Lancaster:1965:HM

[Lan65] H. O. Lancaster. The Helmert matrices. Amer. Math. Monthly, 72:4–12, 1965.

Lancaster:1969:TM

[Lan69] P. Lancaster. Theory of Matrices.

Academic Press, New York, USA, 1969. ???? pp.

Lancaster:1970:ESL

[Lan70] Peter Lancaster. Explicit solutions of linear matrix equations. SIAM Review, 12(4):544–566, 1970.

Landau:1989:DAC

[Lan89a] Sidney I. Landau. Dictionaries: The Art and Craft of Lexicography. Cambridge University Press, ????, 1989. ISBN 0-521-36725-5. ???? pp.

Langston:1989:UMT

[Lan89b] P. S. Langston. Unix music tools at Bellcore. Software—Practice and Experience, 20:S1/47–S1/61, 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lanham:1992:RP

[Lan92] Richard A. Lanham. Revising Prose.

MacMillan Publishing Company, New York, NY, USA, third edition, 1992. ISBN 0-02-367445-8. ???? pp.

Lang:1993:PAR

[Lan93] Bruno Lang. A parallel algorithm for reducing symmetric banded matrices to tridiagonal form. SIAM J. Matrix Anal. Appl., 14(6):1320–1338, 1993.

LaQuey:1994:ICB

[LaQ94] Tracey LaQuey. The Internet Companion: A Beginner’s Guide to Global Networking. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-40766-3. viii + 262 pp.

Larson:1966:LSE

[Lar66] Harold J. Larson. Least squares estimation of the components of a symmetric matrix. Technometrics, 8(2):360–362, 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Large:1983:FLB

[Lar83] J. A. Large. The Foreign-Language Barrier: Problems in Scientific Communication. Andre Deutsch, London, 1983. ???? pp.

Lajosaszlo:1991:UBM

[Lás91] Lajos László. Upper bounds for matrix diagonals. Linear and Multilinear Algebra, 29:283–301, 1991. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lajosaszlo:1994:ALB

[Lás94] Lajos László. An attainable lower bound for the best normal approximation. SIAM J. Matrix Anal. Appl., 15(3):1035– 1043, 1994.

Peteraeuchli:1961:JEA

[Läu61] Peter Läuchli. Jordan-Elimination und Ausgleichung nach kleinsten Quadraten. Numer. Math., 3:226–240, 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Laurie:1975:PIR

[Lau75] D. P. Laurie. Propagation of initial rounding error in Romberg-like quadrature. BIT, 15:277–282, 1975. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Laub:1979:SMS

[Lau79] Alan J. Laub. A Schur method for solving algebraic Riccati equations. IEEE Trans. Automat. Control., AC-24(6):913– 921, 1979.

Laub:1985:NLA

[Lau85] Alan J. Laub. Numerical linear algebra aspects of control design computations. IEEE Trans. Automat. Control., AC-30(2):97–108, 1985.

Laurie:1988:WGQ

[Lau88a] D. P. Laurie. What good is a quadrature estimate? Technical report, in Numerical Integration III, International Series of Numerical Mathematics 85, H. Braß and G. Hämmerlin, eds, Birkhäuser, ????, 1988. ???? pp.

Laurie:1988:NAI

[Lau88b] Dirk P. Laurie. A numerical approach to the inverse Toeplitz eigenproblem. SIAM J. Sci. Stat. Comput., 9(2): 401–405, 1988.

Lavington:1980:EBC

[Lav80] Simon Lavington. Early British Computers: The Story of Vintage Computers and the People Who Built Them. Manchester University Press, Manchester, UK, 1980. ISBN 0-7190-0803-4. 139 pp.

Lawden:1972:PM

[Law72] G. H. Lawden. Pascal matrices.

Mathematical Gazette, 56(398):325–327, 1972. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lax:1965:NSP

[Lax65] P. D. Lax. Numerical solution of partial differential equations. American Mathematical Monthly, 72:74–84, 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lay:1994:LAA

[Lay94] David C. Lay. Linear Algebra and Its Applications. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-84556-3. 624 pp.

Lazou:1988:STU

[Laz88] C. Lazou. Supercomputers and their Use. Oxford University Press, ????, revised edition, 1988. ???? pp.

Li:1989:NMS

[LC89] Guangye Li and Thomas F. Coleman. A new method for solving triangular systems on distributed-memory messagepassing multiprocessors. SIAM J. Sci. Stat. Comput., 10(2):382–396, 1989.

Le:1985:EDF

[Le85a] D. Le. An efficient derivative-free method for solving nonlinear equations. ACM Transactions on Mathematical Software, 11:250–262, 1985. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Le:1985:TNR

[Le85b] D. Le. Three new rapidly convergent algorithms for finding a zero of a function. SIAM J. Sci. Stat. Comput., 6:193–208, 1985.

Ledson:1968:AGB

[Led68] Sidney Ledson. The Awful Grammar Book. Wolfe Publishing Ltd., London, 1968. ???? pp.

Lederer:1987:AEA

[Led87] Richard Lederer. Anguished English: An Anthology of Accidental Assaults Upon Our Language. Wyrick & Company, Charleston, South Carolina, 1987. ISBN 0-941711-04-8. ???? pp. Also available in a Dell edition.

Lederer:1990:CE

[Led90] Richard Lederer. Crazy English.

Pocket Books, New York, NY, USA, 1990. ISBN 0-671-68907-X. ???? pp.

Lederer:1991:ML

[Led91] Richard Lederer. The Miracle of Language. Pocket Books, New York, NY, USA, 1991. ISBN 0-671-70940-2. ???? pp.

Lees:1981:SPM

[Lee81] Gene Lees. Songwriters and Poets Modern Rhyming Dictionary. Omnibus Press, London, 1981. ???? pp.

Lees:1988:MMA

[Lee88] Gene Lees. Meet Me at Jim & Andy’s—Jazz Musicians and their World. Oxford University Press, ????, 1988. ???? pp.

Lee:1995:PUB

[Lee95] Steven L. Lee. A practical upper bound for departure from normality. SIAM J. Matrix Anal. Appl., 16(2):462– 468, 1995.

Lefkovitch:1978:CCQ

[Lef78] L. P. Lefkovitch. Consensus coordinates from qualitative and quantitative attributes. Biometrical J., 20:679–691, 1978. CODEN BIJODN. ISSN 0323-3847.

Lehmer:1943:TMP

[Leh43] D. H. Lehmer. Tables to many places of decimals. Mathematical Tables and Other Aids to Computation, 1(1):30–31, 1943.

Lehmer:1960:MM

[Leh60] D. H. Lehmer. Mahler’s matrices. J. Australian Math. Soc., 1:385–395, 1960. CODEN JAUMAX. ISSN 0004-9735.

Lehtonen:1992:TUL

[Leh92] Ari Lehtonen. Typesetting using LAMS-TEX macros. Euromath Bulletin, 1:59–71, 1992.

Lehman:1994:PCQ

[Leh94a] Eugene Lehman. Pocket calculating the quartic. Notices Amer. Math. Soc., 41 (7):777–779, 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lehoucq:1994:CEU

[Leh94b] R. B. Lehoucq. The computation of elementary unitary matrices. Technical Report CS-94-233, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, May 1994. 9 pp. LAPACK Working Note 72.

Lei:1989:CAI

[Lei89] L. Lei. Convergence of asynchronous iteration with arbitrary splitting form. Linear Algebra and Appl., 113:119–127, 1989.

Lemisch:1961:BFA

[Lem61] L. Jesse Lemisch. Benjamin Franklin: The Autobiography and Other Writings. Signet Classic (Penguin), New York, 1961. ISBN 0-451-52497-7. ???? pp.

Lemeire:1975:BCN

[Lem75a] Frans Lemeire. Bounds for condition numbers of triangular and trapezoid matrices. BIT, 15:58–64, 1975. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lemeire:1975:CEI

[Lem75b] Frans Lemeire. Computation of equivalent inherent rounding errors in the solution of a set of linear equations. BIT, 15:65–71, 1975. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lemeire:1976:EMO

[Lem76] Frans Lemeire. Equilibration of matrices to optimize backward numerical stability. BIT, 16:143–145, 1976. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Leon:1993:LAA

[Leo93] Steven J. Leon. Linear Algebra with Applications. Prentice-Hall, Englewood Cliffs, NJ, USA, fourth edition, 1993. ISBN 0-02-369831-4. xiv + 505 pp.

Leron:1983:SMP

[Ler83] Uri Leron. Structuring mathematical proofs. Amer. Math. Monthly, 90(3):174– 185, 1983.

Levine:1958:VMS

[Lev58] J. Levine. Variable matrix substitution in algebraic cryptography. Amer. Math. Monthly, 65(3):170–179, 1958.

Levinger:1980:SRM

[Lev80] Bernard W. Levinger. The square root of a 2 × 2 matrix. Mathematics Magazine, 53(4):222–224, 1980. CODEN MAMGA8. ISSN 0025-570X.

Levison:1983:EMF

[Lev83] M. Levison. Editing mathematical formulae. Software—Practice and Experience, 13:189–195, 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levine:1989:JPB

[Lev89] Mark Levine. The Jazz Piano Book. Sher Music, P.O. Box 445, Petaluma, California 94953, 1989. ISBN 0-9614701-5-1. ???? pp.

Lewin:1978:UGW

[Lew78] Ronald Lewin. Ultra Goes to War: The Secret Story. Hutchinson, London, 1978. ISBN 0-09-134420-4. 398 pp.

Lewis:1982:ITM

[Lew82] Jerry W. Lewis. Inversion of tridiagonal matrices. Numer. Math., 38:333–345, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lewkowicz:1992:WCR

[Lew92] I. Lewkowicz. When are the complex and real stability radii equal? IEEE Trans. Automat. Control., AC-37(6):880– 883, 1992.

Ladner:1980:PPC

[LF80] R. E. Ladner and M. J. Fischer. Parallel prefix computations. J. Assoc. Comput. Mach., 27(4):831–838, 1980.

Lawson:1974:SLS

[LH74] Charles L. Lawson and Richard J. Hanson. Solving Least Squares Problems. Prentice-Hall, Englewood Cliffs, NJ, USA, 1974. ISBN 0-13-822585-0. xii + 340 pp.

Lawson:1979:BLA

[LHKK79] C. L. Lawson, R. J. Hanson, D. R.

Kincaid, and F. T. Krogh. Basic linear algebra subprograms for Fortran usage. ACM Transactions on Mathematical Software, 5(3):308–323, 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Laub:1987:CSB

[LHPW87] Alan J. Laub, M. T. Heath, C. C.

Paige, and R. C. Ward. Computation of system balancing transformations and other applications of simultaneous diagonalization algorithms. IEEE Trans. Automat. Control., AC-32:115–122, 1987.

Li:1988:MSE

[Li88] C. K. Li. Matrices with some extremal properties. Linear Algebra and Appl., 101:255–267, 1988.

Li:1991:GCM

[Li91] Y. Li. A globally convergent method for L_{p} problems. Technical report, Technical Report 91-1212, Department of Computer Science, Cornell University, Ithaca, NY, ????, 1991. ???? pp.

Li:1995:NPB

[Li95] Ren-Cang Li. New perturbation bounds for the unitary polar factor. SIAM J. Matrix Anal. Appl., 16(1):327– 332, 1995.

LIA:1993:ITI

[LIA93] Information Technology—Language Independent Arithmetic—Part I: Integer and Floating Point Arithmetic, Draft International Standard (Version 4.1), ISO/IEC DIS 10967-1:1993. August 1993.

Linz:1970:AFP

[Lin70] Peter Linz. Accurate floating-point summation. Comm. ACM, 13(6):361– 362, 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindsey:1972:AFT

[Lin72] Charles H. Lindsey. ALGOL 68 with fewer tears. The Computer Journal, 15: 176–188, 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Linnainmaa:1974:ASK

[Lin74] Seppo Linnainmaa. Analysis of some known methods of improving the accuracy of floating-point sums. BIT, 14:167– 202, 1974. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Linnainmaa:1975:TAS

[Lin75] Seppo Linnainmaa. Towards accurate statistical estimation of rounding errors in floating-point computations. BIT, 15: 165–173, 1975. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Linnainmaa:1976:TEA

[Lin76] Seppo Linnainmaa. Taylor expansion of the accumulated rounding error. BIT, 16:146–160, 1976. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Linnainmaa:1981:SDP

[Lin81] Seppo Linnainmaa. Software for doubled-precision floating-point computations. ACM Transactions on Mathematical Software, 7(3):272–283, 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lindley:1984:R

[Lin84a] Dennis V. Lindley. Refereeing. The Mathematical Intelligencer, 6(2):56–60, 1984. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Lindsay:1984:GSW

[Lin84b] David Lindsay. A Guide To Scientific Writing. Longman Cheshire, Melbourne, Australia, 1984. ISBN 0-582-71132-0. ???? pp.

Linz:1988:CNA

[Lin88] P. Linz. A critique of numerical analysis. Bull. Amer. Math. Soc., 19:407–416, 1988.

Linzer:1992:SSM

[Lin92a] Elliot Linzer. On the stability of solution methods for band Toeplitz systems. Linear Algebra and Appl., 170:1–32, 1992.

Linzer:1992:STB

[Lin92b] Elliot Linzer. On the stability of transform-based circular deconvolution. SIAM J. Numer. Anal., 29(5):1482–1492, 1992.

Littlewood:1978:MAW

[Lit78] J. E. Littlewood. The mathematician’s art of work. The Mathematical Intelligencer, 1(2):112–118, 1978. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). Also in [Bol86].

Liu:1987:PPS

[Liu87] Joseph W. H. Liu. A partial pivoting strategy for sparse symmetric matrix decomposition. ACM Transactions on Mathematical Software, 13(2):173– 182, 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lichtenstein:1993:BCD

[LJ93] Woody Lichtenstein and S. Lennart Johnson. Block-cyclic dense linear algebra. SIAM J. Matrix Anal. Appl., 14(6): 1259–1288, 1993.

Lyness:1976:CNA

[LK76] J. N. Lyness and J. J. Kaganove. Comments on the nature of automatic quadrature routines. ACM Transactions on Mathematical Software, 2(1):65–81, 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lorayne:1974:MB

[LL74] H. Lorayne and J. Lucas. The Memory Book. W.H. Allen & Co. Ltd., ????, 1974. ???? pp.

Lambert:1986:HFI

[LL86] Jill Lambert and Peter A. Lambert. How To Find Information in Science and Technology. Clive Bingley, London, 1986. ISBN 0-85157-394-0. ???? pp.

Li:1993:AST

[LL93] Kuiyuan Li and Tien-Yien Li. An algorithm for symmetric tridiagonal eigenproblems: Divide and conquer with homotopy continuation. SIAM J. Sci. Stat. Comput., 12(3):735–751, 1993.

Lloyd:1984:S

[Llo84] E. H. Lloyd, editor. Statistics, volume VI of Handbook of Applicable Mathematics. Wiley, Chichester, UK, 1984. ???? pp.

Lyness:1966:VMS

[LM66] J. N. Lyness and C. B. Moler. Van der Monde systems and numerical differentiation. Numer. Math., 8:458–464, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Levy:1985:TAG

[LM85] A. V. Levy and A. Montalvo. The tunneling algorithm for the global minimization of functions. SIAM J. Sci. Stat. Comput., 6:15–29, 1985.

Lazer:1990:LAP

[LM90] A. C. Lazer and P. J. McKenna.

Large-amplitude periodic oscillations in suspension bridges: Some new connections with nonlinear analysis. SIAM Review, 32(4):537–578, 1990.

Levine:1962:CIM

[LN62] J. Levine and H. M. Nahikian. On the construction of involutary matrices. Amer. Math. Monthly, 69:267–272, 1962.

Lee:1972:SQM

[LN72] C. C. Lee and H. P. Niu. Solution of quadratic matrix equations by leastsquares method. Quarterly of Applied Mathematics, pages 345–350, 1972. CO-DEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic). October.

Lei:1992:FAE

[LN92] Li Lei and T. Nakamura. A fast algorithm for evaluating the matrix polynomial I + A + ⋯ + A^{N−1}. IEEE Trans. Circuits and Systems, AC-32(39 (or 4??)): 299–300, 1992.

Li:1982:FSF

[LNOP82] M. R. Li, B. Nour-Omid, and B. N.

Parlett. A fast solver free of fill-in for finite-element problems. SIAM J. Numer. Anal., 19(6):1233–1242, 1982.

Leuprecht:1982:PAR

[LO82] H. Leuprecht and W. Oberaigner. Parallel algorithms for the rounding exact summation of floating point numbers. Computing, 28:89–104, 1982. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Lock:1985:HDI

[Loc85] Stephen Lock, editor. How To Do It. British Medical Association, London, second edition, 1985. ISBN 0-7279-0186-9. ???? pp.

Lodge:1975:CP

[Lod75] David Lodge. Changing Places. Penguin, ????, 1975. ???? pp.

Lodge:1984:SW

[Lod84] David Lodge. Small World. Penguin, ????, 1984. ???? pp.

Loeckx:1974:ALP

[Loe74] Jacques Loeckx, editor. Automata, Languages and Programming, volume 14 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1974.

Loizou:1969:NJC

[Loi69] Georghios Loizou. Nonnormality and Jordan condition numbers of matrices. J. Assoc. Comput. Mach., 16(4):580–584, 1969.

Longley:1967:ALS

[Lon67] J. W. Longley. An appraisal of least squares programs for the electronic computer from the point of view of the user. J. Amer. Statist. Assoc., 62:819–841, 1967.

Longley:1984:LSC

[Lon84] James W. Longley. Least Squares Computations Using Orthogonalization Methods. Marcel Dekker, Inc., New York, NY, USA, 1984. ISBN 0-8247-7232-6. xi + 308 pp.

Longman:1991:LDE

[Lon91] Longman. Longman Dictionary of the English Language. Longman, Harlow, Essex, UK, new edition, 1991. ISBN 0-582-07038-4. ???? pp.

Longyear:1993:MHS

[Lon93] Marie M. Longyear, editor. The McGraw-Hill Style Manual: A Concise Guide for Writers and Editors. McGraw-Hill, New York, 1993. ISBN 0-07-038684-6. ???? pp.

Longman:1995:LDC

[Lon95] Longman. Longman Dictionary of Contemporary English. Longman, Harlow, Essex, UK, third edition, 1995. ISBN 0-582-23750-5. xxii + 1668 pp.

Lootsma:1972:NMN

[Loo72] F. A. Lootsma. Numerical Methods for Non-linear Optimization. Academic Press, New York, USA, 1972. ???? pp.

Lorayne:1974:HDS

[Lor74] H. Lorayne. How to Develop a Super-Power Memory. Signet, New American Library, Inc., ????, 1974. ???? pp.

Lotkin:1955:STM

[Lot55] Mark Lotkin. A set of test matrices.

M.T.A.C., 9:153–161, 1955.

Peroetstedt:1983:PBL

[Löt83] Per Lötstedt. Perturbation bounds for the linear least squares problem subject to linear inequality constraints. BIT, 23: 500–519, 1983. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Peroetstedt:1984:SML

[Löt84] Per Lötstedt. Solving the minimal least squares problem subject to bounds on the variables. BIT, 24:206– 224, 1984. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lowan:1960:PRE

[Low60] Arnold N. Lowan. On the propagation of round-off errors in the numerical integration of the heat equation. Math. Comp., 14:139–146, 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lowman:1984:MTT

[Low84] Joseph Lowman. Mastering the Techniques of Teaching. Jossey-Bass Publishers, San Francisco and London, 1984. ???? pp.

Lasdon:1987:ODE

[LPBW87] L. S. Lasdon, John Plummer, B. Buehler, and A. D. Waren. Optimal design of efficient acoustic antenna arrays. Math. Prog., 39:131–155, 1987. CODEN MHPGA4. ISSN 0025-5610.

Lena:1993:RRP

[LPP93] G. Di Lena, R. I. Peluso, and G. Piazza. Results on the relative perturbation of the singular values of a matrix. BIT, 33:647–653, 1993. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lo:1987:MAS

[LPS87] S. Lo, B. Philippe, and A. Sameh. A multiprocessor algorithm for the symmetric tridiagonal eigenvalue problem. SIAM J. Sci. Stat. Comput., 8:s155–s165, 1987.

Laderman:1992:PAA

[LPS92] Julian Laderman, Victor Pan, and Xuan-He Sha. On practical algorithms for accelerated matrix multiplication. Linear Algebra and Appl., 162-164:557–588, 1992.

Larson:1983:ASR

[LPW83] John L. Larson, Mary E. Pasternak, and John A. Wisniewski. Algorithm 594: Software for relative error analysis. ACM Transactions on Mathematical Software, 9(1):125–130, 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Luk:1987:FUO

[LQ87] Franklin T. Luk and Sanzheng Qiao.

A fast but unstable orthogonal triangularization technique for Toeplitz matrices. Linear Algebra and Appl., 88/89:495–506, 1987.

Lau:1981:CBA

[LR81] K. K. Lau and W. O. J. Riha. Characterization of best approximations in normed linear spaces of matrices by elements of finite-dimensional linear subspaces. Linear Algebra and Appl., 35:109– 120, 1981.

Lancaster:1983:ME

[LR83] P. Lancaster and P. Rozsa. On the matrix equation AX + X^{∗}A^{∗} = C. SIAM J. Alg. Disc. Meth., 4(4):432–436, 1983.

LaQuey:1993:ICB

[LR93] Tracey LaQuey and Jeanne C. Ryer.

The Internet Companion: A Beginner’s Guide to Global Networking. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62224-6. ???? pp.

Larson:1978:ECE

[LS78] John L. Larson and Ahmed H. Sameh. Efficient calculation of the effects of roundoff errors. ACM Transactions on Mathematical Software, 4(3):228–236, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). Errata 5(3):372, 1979.

Larson:1980:ARE

[LS80] John L. Larson and Ahmed H. Sameh. Algorithms for roundoff error analysis— A relative error approach. Computing, 24:275–297, 1980. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Laub:1982:LAA

[LS82] Alan J. Laub and G. R. Shiflett. A linear algebra approach to the analysis of rigid body displacement from initial and final position data. Journal of Applied Mechanics, 49:213–216, 1982. CODEN JAMCAV. ISSN 0021-8936 (print), 1528-9036 (electronic).

Laub:1983:LAA

[LS83] Alan J. Laub and G. R. Shiflett. A linear algebra approach to the analysis of rigid body velocity from position and velocity data. Journal of Dynamic Systems, Measurement, and Control, 105:92– 95, 1983.

Lenferink:1990:GNR

[LS90] H. W. J. Lenferink and M. N. Spijker. A generalization of the numerical range of a matrix. Linear Algebra and Appl., 140: 251–266, 1990.

Lenferink:1991:USR

[LS91] H. W. J. Lenferink and M. N. Spijker. On the use of stability regions in the numerical analysis of initial value problems. Math. Comp., 57(195):221–237, 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehoucq:1994:DTI

[LS94] Richard Lehoucq and Danny C.

Sorensen. Deflation techniques for an implicitly restarted Arnoldi iteration. Technical Report TR94-13, Department of Computational and Applied Mathematics, Rice University, Houston, TX, USA, February 1994. 39 pp. Revised February 1995.

Lehoucq:1995:AII

[LSV95] R. B. Lehoucq, D. C. Sorensen, and P. Vu. ARPACK: An implementation of the implicitly re-started Arnoldi iteration that computes some of the eigenvalues and eigenvectors of a large sparse matrix, 1995. Available from netlib@ornl.gov under the directory scalapack.

Lancaster:1985:TM

[LT85] P. Lancaster and M. Tismenetsky. The Theory of Matrices. Academic Press, New York, USA, second edition, 1985. ???? pp.

LeVeque:1988:FAS

[LT88] R. J. LeVeque and L. N. Trefethen.

Fourier analysis of the SOR iteration. IMA J. Numer. Anal., 8:273–279, 1988.

Lederman:1992:PER

[LTT92] S. Lederman, A. Tsao, and T. Turnbull. A parallelizable eigensolver for real diagonalizable matrices with real eigenvalues. Technical Report TR-91-042, Supercomputing Research Center, Bowie, MD, March 1992. ???? pp.

Lu:1994:FSC

[Lu94] Hao Lu. Fast solution of confluent Vandermonde linear systems. SIAM J. Matrix Anal. Appl., 15(4):1277–1289, 1994.

Lu:1995:FAC

[Lu95] Hao Lu. Fast algorithms for confluent Vandermonde linear systems and generalized Trummer’s problem. SIAM J. Matrix Anal. Appl., 16(2):655–674, 1995.

Lu:1996:SVL

[Lu96] Hao Lu. Solution of Vandermonde-like systems and confluent Vandermonde-like systems. SIAM J. Matrix Anal. Appl., 17: ??, 1996. To appear.

Luey:1987:HAA

[Lue87] Beth Luey. Handbook for Academic Authors. Cambridge University Press, ????, 1987. ISBN 0-521-33682-1. ???? pp.

Lui:1996:CPC

[Lui96] S. H. Lui. Computation of pseudospectra by continuation. SIAM J. Sci. Comput., page ??, 1996. To appear.

Luk:1980:CSV

[Luk80] F. T. Luk. Computing the singularvalue decomposition on the ILLIAC IV. ACM Transactions on Mathematical Software, 6:524–539, 1980. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Luk:1983:ORP

[Luk83] Franklin T. Luk. Orthogonal rotation to a partially specified target. SIAM J. Sci. Stat. Comput., 4(2):223–228, 1983.

Luk:1984:OPR

[Luk84] F. T. Luk. Oblique Procrustes rotations in factor analysis. SIAM J. Sci. Stat. Comput., 5:764–770, 1984.

Luk:1985:MMF

[Luk85] F. T. Luk. On the minres method of factor analysis. SIAM J. Sci. Stat. Comput., 6:562–572, 1985.

Luk:1986:TPA

[Luk86] Franklin T. Luk. A triangular processor array for computing singular values. Linear Algebra and Appl., 77:259– 273, 1986.

Lustig:1989:AAS

[Lus89] I. J. Lustig. An analysis of an available set of linear programming test problems. Computers and Operations Research, 16:173–184, 1989. CODEN CMORAP. ISSN 0305-0548 (print), 1873-765X (electronic).

Latterner:1986:TAM

[LW86] D. C. Latterner and W. B. Woolf. TEX at Mathematical Reviews. TUGboat, 10 (4):639–654, 1986. ISSN 0896-3207.

Lay-Yong:1989:MSL

[LYK89] Lam Lay-Yong and Shen Kangshen.

Methods of solving linear equations in traditional China. Historia Mathematica, 16(2):107–122, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0315086089900335.

Lynn:1964:REM

[Lyn64] M. Stuart Lynn. On the round-off error in the method of successive overrelaxation. Math. Comp., 18(85):36–49, 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1969:EIC

[Lyn69] J. N. Lyness. The effect of inadequate convergence criteria in automatic routines. The Computer Journal, 12:279– 281, 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). See also letter and response in 13 (1970), p. 121.

Lyness:1982:RAP

[Lyn82] J. N. Lyness. Remarks about performance profiles. In Paul C. Messina and Almerico Murli, editors, Problems and Methodologies in Mathematical Software Production, volume 142 of Lecture Notes in Computer Science, pages 186– 197. Springer-Verlag, Berlin, 1982.

Lyness:1983:WUA

[Lyn83] J. N. Lyness. When not to use an automatic quadrature routine. SIAM Review, 25(1):63–87, 1983.

Lyness:1987:SQR

[Lyn87] J. N. Lyness. Some quadrature rules for finite trigonometric and related integrals. In P. Keast and G. Fairweather, editors, Numerical Integration: Recent Developments, Software and Applications, pages 17–33. D. Reidel Publishing Company, Dordrecht, Holland, 1987.

Lin:1991:PAC

[LZ91] Chih-Chang Lin and Earl Zmijewski. A parallel algorithm for computing the eigenvalues of an unsymmetric matrix on an SIMD mesh of processors. Report TRCS 91-15, Department of Computer Science, University of California, Santa Barbara, July 1991. ???? pp.

Li:1992:HDA

[LZ92] T. Y. Li and Zhonggang Zeng.

Homotopy-determinant algorithm for solving nonsymmetric eigenvalue problems. Math. Comp., 59(200):483–502, 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1994:LIS

[LZ94] T. Y. Li and Zhonggang Zeng. The Laguerre iteration in solving the symmetric tridiagonal eigenproblem, revisited. SIAM J. Sci. Comput., 15(5):1145–1173, 1994.

Li:1992:SEP

[LZC92] T.-Y. Li, Z. Zeng, and L. Cong.

Solving eigenvalue problems of real nonsymmetric matrices with real homotopies. SIAM J. Numer. Anal., 29(1):229–248, 1992.

Li:1991:PHA

[LZS91] T.-Y. Li, H. Zhang, and X.-H. Sun.

Parallel homotopy algorithm for the symmetric tridiagonal eigenvalue problem. SIAM J. Sci. Stat. Comput., 12(3):469– 487, 1991.

Meijer:1987:OCP

[MA87] H. Meijer and S. G. Akl. Optimal computation of prefix sums on a binary tree of processors. International Journal of Parallel Programming, 16(2):127–136, 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

MacDuffee:1943:VM

[Mac43] C. C. MacDuffee. Vectors and Matrices, volume 7 of The Carus Mathematical Monographs. Mathematical Association of America, Washington, DC, USA, 1943. ???? pp.

MacGregor:1977:GSC

[Mac77a] A. J. MacGregor. Graphics simplified: Charts and graphs. Scholarly Publishing, 8(2):151–164, 1977. ISSN 0036-634X.

MacGregor:1977:GSM

[Mac77b] A. J. MacGregor. Graphics simplified: Media specifications and artists’ aids. Scholarly Publishing, 8(4):367–376, 1977. ISSN 0036-634X.

MacGregor:1977:GSP

[Mac77c] A. J. MacGregor. Graphics simplified: Preparing charts and graphs. Scholarly Publishing, 8(3):257–274, 1977. ISSN 0036-634X.

Mackay:1977:HQE

[Mac77d] A. L. Mackay. The Harvest of a Quiet Eye: A Selection of Scientific Quotations. The Institute of Physics, Bristol and London, 1977. ???? pp.

MacLeod:1988:DGF

[Mac88] A. J. MacLeod. The distribution of the growth factor in Gaussian elimination with partial pivoting. Technical report, Technical Report, Department of Mathematics and Statistics, Paisley College of Technology, Paisley, Scotland, ????, 1988. ???? pp.

Macleod:1989:ISB

[Mac89a] A. J. Macleod. Instability in the solution of banded Toeplitz systems. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-37(9):1449–1450, 1989. CODEN IETABA. ISSN 0096-3518.

Macleod:1989:SSG

[Mac89b] Allan J. Macleod. Some statistics on Gaussian elimination with partial pivoting. ACM SIGNUM Newsletter, 24(2/3):

10–14, 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

MacQueen:1990:UNE

[Mac90] Donald S. MacQueen. Using Numbers in English: A Reference Guide for Swedish Speakers Including Basic Terminology for Describing Graphs. Studentlitteratur (Lund, Sweden) and Chartwell Bratt (Bromley, Kent, England), ????, 1990. ISBN 91-44-31921-5, 0-86238-264-5. ???? pp.

Mackey:1995:HJM

[Mac95] Niloufer Mackey. Hamilton and Jacobi meet again: Quaternions and the eigenvalue problem. SIAM J. Matrix Anal. Appl., 16(2):421–435, 1995.

Maddox:1974:IYJ

[Mad74] John Maddox (?). It’s your journal.

Nature, 252:337, 1974. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Maddox:1983:MSI

[Mad83] John Maddox. Must science be impenetrable? Nature, 305(6):477–478, 1983. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Maddox:1987:WEJ

[Mad87] John Maddox. What is an electronic journal? Nature, 328:471, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Maddox:1988:HCA

[Mad88] John Maddox. How not to catch attention. Nature, 332:391, 1988. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Maddox:1990:SSS

[Mad90] John Maddox. Saying sorry, but not sorry. Nature, 343:589, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Magula:1960:CFM

[Mag60] D. A. Magula. A conjugate factor method for the solution of a cubic. Math. Comp., 14:281–283, 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Maindonald:1984:SC

[Mai84] J. H. Maindonald. Statistical Computation. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1984. ISBN 0-471-86452-8. xviii + 370 pp.

Makhoul:1991:TDP

[Mak91] John Makhoul. Toeplitz determinants and positive semidefiniteness. IEEE Trans. Signal Processing, 39(3):743–746, 1991.

Malcolm:1971:AFP

[Mal71] Michael A. Malcolm. On accurate floating-point summation. Comm. ACM, 14(11):731–736, November 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malcolm:1972:ARP

[Mal72] Michael A. Malcolm. Algorithms to reveal properties of floating-point arithmetic. Comm. ACM, 15(11):949–951, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mallon:1989:SWF

[Mal89] Thomas Mallon. Stolen Words: Forays Into the Origins and Ravages of Plagiarism. Penguin, ????, 1989. ISBN 0-14-014440-4. xiv + 300 pp.

Malyshev:1992:GASa

[Mal92a] Alexander N. Malyshev. Guaranteed accuracy in spectral problems of linear algebra, I. Siberian Advances in Mathematics, 2(1):144–197, 1992. ISSN 1055-1344 (print), 1934-8126 (electronic).

Malyshev:1992:GASb

[Mal92b] Alexander N. Malyshev. Guaranteed accuracy in spectral problems of linear algebra, II. Siberian Advances in Mathematics, 2(2):153–204, 1992. ISSN 1055-1344 (print), 1934-8126 (electronic).

Malyshev:1993:PAS

[Mal93] Alexander N. Malyshev. Parallel algorithms for solving some spectral problems of linear algebra. Linear Algebra and Appl., 188:489–520, 1993.

Manheim:1975:SR

[Man75] Frank T. Manheim. The scientific referee. IEEE Trans. Prof. Commun., PC-18(3):190–195, 1975.

Manteuffel:1979:SIC

[Man79] T. A. Manteuffel. Shifted incomplete Cholesky factorization. In Duff and Stewart [DS79b], pages 41–61.

Mansbridge:1980:PTD

[Man80] Ronald Mansbridge. ‘my publishers are terrible … ’. Scholarly Publishing, 11 (2):133–140, 1980. ISSN 0036-634X.

Manteuffel:1981:IAA

[Man81] Thomas A. Manteuffel. An interval analysis approach to rank determination in linear least squares problems. SIAM J. Sci. Stat. Comput., 2(3):335–348, 1981.

Mandel:1982:USV

[Man82] J. Mandel. Use of the singular value decomposition in regression analysis. The American Statistician, 36:15–24, 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Manser:1988:GBW

[Man88] Martin Manser. The Guinness Book of Words. Guinness, Enfield, Middlesex, 1988. ???? pp.

Manser:1990:BGW

[Man90] Martin H. Manser. Bloomsbury Good Word Guide. Bloomsbury, London, second edition, 1990. ISBN 0-7475-0875-5. ???? pp.

Mangasarian:1994:NP

[Man94] Olvi L. Mangasarian. Nonlinear Programming. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-341-2. xv + 220 pp. Corrected republication of work first published in 1969 by McGraw-Hill, New York.

Mao:1986:OOS

[Mao86] Jianqin Mao. Optimal orthonormalization of the strapdown matrix by using singular value decomposition. Computers and Mathematics with Applications, 12 (Part 1) (3):353–362, March 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/089812218690194X.

Marquardt:1963:ALS

[Mar63] D. W. Marquardt. An algorithm for least-squares estimation of nonlinear parameters. J. Soc. Indust. Appl. Math., 11 (2):431–441, 1963.

Marcus:1989:UES

[Mar89] Marvin Marcus. A unified exposition of some classical matrix theorems. Linear and Multilinear Algebra, 25:137–147, 1989. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Markoff:1994:CFC

[Mar94] John Markoff. Circuit flaw causes Pentium chip to miscalculate, Intel admits. New York Times, page ??, 1994. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. 24 November.

Mascarenhas:1991:CJM

[Mas91] W. F. Mascarenhas. On the convergence of the Jacobi method for arbitrary orderings. PhD thesis, M.I.T., Cambridge, MA, January 1991. ???? pp. Numerical Analysis Report 91-2, Department of Mathematics.

Mascarenhas:1994:NJB

[Mas94] Walter F. Mascarenhas. A note on Jacobi being more accurate than QR. SIAM J. Matrix Anal. Appl., 15(1):215– 218, 1994.

Matula:1968:C

[Mat68] David W. Matula. In-and-out conversions. Comm. ACM, 11(1):47–50, January 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Gol67, ?].

Mathis:1969:CSU

[Mat69] R. F. Mathis. Completion of a symmetric unitary matrix. SIAM Review, 11: 261–263, 1969.

Matula:1970:FFP

[Mat70] David W. Matula. A formalization of floating-point numeric base conversion. IEEE Trans. Comput., C-19(8):681–692, 1970.

Mattheij:1984:SBD

[Mat84a] R. M. M. Mattheij. Stability of block LU-decompositions of matrices arising from BVP. SIAM J. Alg. Disc. Meth., 5(3):314–331, 1984.

Mattheij:1984:SLD

[Mat84b] R. M. M. Mattheij. The stability of LU-decompositions of block tridiagonal matrices. Bulletin of the Australian Mathematical Society, 29:177–205, 1984. CO-DEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (electronic).

Mathar:1985:BEF

[Mat85] Rudolf Mathar. The best Euclidian fit to a given distance matrix in prescribed dimensions. Linear Algebra and Appl., 67:1–6, 1985.

Mathias:1990:CMM

[Mat90a] R. Mathias. Concavity of monotone matrix functions of finite order. Linear and Multilinear Algebra, 27:129–138, 1990. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Mathias:1990:TTS

[Mat90b] R. Mathias. Two theorems on singular values and eigenvalues. Amer. Math. Monthly, 97:47–50, 1990.

Mathias:1990:SNN

[Mat90c] Roy Mathias. The spectral norm of a nonnegative matrix. Linear Algebra and Appl., 139:269–284, 1990.

Matthews:1990:SPG

[Mat90d] Diane L. Matthews. The scientific poster: Guidelines for effective visual communication. Technical Communication, 37(3):225–232, 1990. ISSN 0049-3155 (print), 1938-369x (electronic).

Mathias:1991:AMV

[Mat91a] R. Mathias. Approximating matrix valued functions. 1991.

Mathias:1991:EFD

[Mat91b] R. Mathias. Evaluating the Frechet derivative of the matrix exponential. 1991.

Mathias:1992:AAO

[Mat92a] R. Mathias. Analysis of algorithms for orthogonalizing products of unitary matrices. 1992.

Mathias:1992:MPD

[Mat92b] Roy Mathias. Matrices with positive definite Hermitian part: Inequalities and linear systems. SIAM J. Matrix Anal. Appl., 13(2):640–654, 1992.

MathWorks:1992:MUG

[Mat92c] The MathWorks, Inc., Natick, MA, USA. MATLAB User’s Guide, 1992. ???? pp.

Matstoms:1992:SQF

[Mat92d] P. Matstoms. Sparse QR factorization in MATLAB. Report LiTH-MAT-R-1992-05, Department of Mathematics, Linköping University, Sweden, March 1992. ???? pp.

Mathias:1993:FAE

[Mat93a] Roy Mathias. Fast accurate eigensystem computations by Jacobi methods. Manuscript. Institute for Mathematics and its Applications, University of Minnesota, April 1993.

Mathias:1993:PBP

[Mat93b] Roy Mathias. Perturbation bounds for the polar decomposition. SIAM J. Matrix Anal. Appl., 14(2):588–597, 1993.

Mathias:1995:AEC

[Mat95a] Roy Mathias. Accurate eigensystem computations by Jacobi methods. SIAM J. Matrix Anal. Appl., 16(3):977–1003, 1995.

Mathias:1995:AAO

[Mat95b] Roy Mathias. Analysis of algorithms for orthogonalizing products of unitary matrices. Numerical Linear Algebra with Applications, 1995. To appear.

Mathias:1995:CEM

[Mat95c] Roy Mathias. Condition estimation for matrix functions via the Schur decomposition. SIAM J. Matrix Anal. Appl., 16 (2):565–578, 1995.

Mathias:1995:IPP

[Mat95d] Roy Mathias. The instability of parallel prefix matrix multiplication. SIAM J. Sci. Comput., 16(4):956–973, 1995.

Maurer:1987:RRD

[Mau87] P. M. Maurer. A rose is a rose . . .. Amer. Math. Monthly, 94:631–645, 1987.

Maurer:1991:AUS

[Mau91] Stephen B. Maurer. Advice for undergraduates on special aspects of writing mathematics. PRIMUS (Problems Resources and Issues in Mathematics Undergraduate Studies), 1(1):9–28, March 1991. ISSN 1051-1970 (print), 1935-4053 (electronic).

Mazurkiewicz:1976:MFC

[Maz76] Antoni W. Mazurkiewicz, editor.

Mathematical foundations of computer science, 1976: proceedings, 5th Symposium, Gdansk, September 6–10, 1976, volume 45 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07854-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xi + 606 pp. LCCN QA267.5.S4 S989 1976. “The Symposium is being organized by the Computation Centre of the Polish Academy of Sciences in cooperation with the University of Gdansk.”.

Melloy:1993:CEI

[MB93] Brian J. Melloy and G. Kemble Bennett. Computing the exponential of an intensity matrix. J. Comp. Appl. Math., 46:405–413, 1993.

Meyer:1977:PCI

[MBI77] J. Meyer and I. Y. Bar-Itzhack. Practical comparison of iterative matrix orthogonalisation algorithms. IEEE Trans. Aerospace Electronic Systems, 13:230– 235, 1977.

Mason:1990:AAI

[MC90] J. C. Mason and M. G. Cox. Algorithms for Approximation II. Chapman and Hall, Ltd., London, UK, 1990. ???? pp.

McCarthy:1967:P

[McC67] Charles A. McCarthy. c_{p}. Israel J.

Math, 5:249–271, 1967.

McConnell:1978:CTW

[McC78] James V. McConnell. Confession of a textbook writer. American Psychologist, 33:159–169, 1978. CODEN AMPSAB. ISSN 0003-066x (print), 1935-990x (electronic).

McCorduck:1979:MWT

[McC79] Pamela McCorduck. Machines Who Think. W.H. Freeman, San Francisco, CA, USA and New York, NY, USA, 1979. ???? pp.

McCracken:1984:EHF

[McC84] Daniel D. McCracken. The early history of FORTRAN publications. Annals of the History of Computing, 6(1):33–34, 1984.

McClure:1987:RFU

[McC87] Glenda M. McClure. Readability formulas: Useful or useless? IEEE Trans. Prof. Commun., PC-30:12–15, 1987.

McIlroy:1982:DSL

[McI82] M. Douglas McIlroy. Development of a spelling list. IEEE Transactions on Communications, COM-30(1 pt 1):91–99, January 1982. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

McKeeman:1962:ACE

[McK62] William Marshall McKeeman. Algorithm 135: Crout with equilibration and iteration. Comm. ACM, 5:553–555, 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKelvey:1984:STE

[McK84] J. P. McKelvey. Simple transcendental expressions for the roots of cubic equations. American Journal of Physics, 52(3):269–270, 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

McCracken:1964:NMF

[MD64] Daniel D. McCracken and William S.

Dorn. Numerical Methods and Fortran Programming: With Applications in Science and Engineering. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1964. xii + 457 pp.

Moor:1992:TLL

[MD92] Bart De Moor and Johan David. Total linear least squares and the algebraic Riccati equation. Systems and Control Letters, 18:329–337, 1992.

Meek:1980:ISM

[Mee80] D. Meek. The inverses of some matrices deviating slightly from a symmetric, tridiagonal, Toeplitz form. SIAM J. Numer. Anal., 17(1):39–43, 1980.

Meinguet:1969:ES

[Mei69] Jean Meinguet. On the estimation of significance. In Elden Hansen, editor, Topics in Interval Analysis, pages 47–64. Oxford University Press, 1969.

Meinguet:1983:REA

[Mei83] Jean Meinguet. Refined error analyses of Cholesky factorization. SIAM J. Numer. Anal., 20(6):1243–1250, 1983.

Mendelsohn:1956:SEP

[Men56] N. S. Mendelsohn. Some elementary properties of ill conditioned matrices and linear equations. Amer. Math. Monthly, 63(5):285–295, 1956.

Merriam-Webster:1983:WNN

[Mer83] Merriam-Webster. Webster’s Ninth New Collegiate Dictionary. Merriam-Webster, Springfield, Massachusetts, 1983. ISBN 0-87779-509-6. 1563 pp. LCCN PE1628 .W5638 1983. 1991 printing.

Mermin:1988:WWL

[Mer88] N. David Mermin. What’s wrong with this Lagrangean? Physics Today, page 9, April 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Reprinted with postscript in [Mer90].

Mermin:1990:BAW

[Mer90] N. David Mermin. Boojums All the Way Through: Communicating Science in a Prosaic Age. Cambridge University Press, ????, 1990. ISBN 0-521-38880-5. ???? pp.

Anonymous:1991:MCH

[Mer91] The Merriam-Webster Concise Handbook for Writers. Merriam-Webster, Springfield, Massachusetts, 1991. ISBN 0-87779-602-5. ix + 310 pp. LCCN PE1408.M517 1991.

Merriam-Webster:1993:MWC

[Mer93] Merriam-Webster. Merriam-Webster’s Collegiate Dictionary. Merriam-Webster, Springfield, Massachusetts, tenth edition, 1993. ISBN 0-87779-708-0. 38a + 1559 pp. LCCN PE1628 .M36 1993.

Metropolis:1977:MSA

[Met77] N. Metropolis. Methods of significance arithmetic. In David A. H. Jacobs, editor, The State of the Art in Numerical Analysis, pages 179–192. Academic Press, London, 1977.

Meurant:1992:RIS

[Meu92] Gérard Meurant. A review on the inverse of symmetric tridiagonal and block tridiagonal matrices. SIAM J. Matrix Anal. Appl., 13(3):707–728, 1992.

Meyer:1975:RGG

[Mey75] C. D. Meyer, Jr. The role of the group generalized inverse in the theory of finite Markov chains. SIAM Review, 17 (3):443–464, 1975.

Meyer:1989:SCU

[Mey89] Carl D. Meyer. Stochastic complementation, uncoupling Markov chains, and the theory of nearly reducible systems. SIAM Review, 31(2):240–272, 1989.

Macdonald:1982:WWC

[MFGK82] Nina H. Macdonald, Lawrence T.

Frase, Patricia S. Gingrich, and Stacey A. Keenan. The Writer’s Workbench: Computer aids for text analysis. IEEE Transactions on Communications, COM-30(1): 105–110, 1982. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Moor:1989:GSV

[MG89] B. L. R. De Moor and G. H. Golub.

Generalized singular value decompositions: A proposal for a standardized nomenclature. 1989.

Moor:1991:RSV

[MG91] B. L. R. De Moor and G. H. Golub.

The restricted singular value decomposition: Properties and applications. SIAM J. Matrix Anal. Appl., 12:401–425, 1991.

More:1981:FST

[MGH81a] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. Fortran subroutines for testing unconstrained optimization software. ACM Transactions on Mathematical Software, 7:136–140, 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1981:TUO

[MGH81b] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. Testing unconstrained optimization software. ACM Transactions on Mathematical Software, 7:17–41, 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Moonen:1993:LAL

[MGM93] Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors. Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-2151-0.

Meyer:1957:MIL

[MH57] H. I. Meyer and B. J. Hollingsworth. A method of inverting large matrices of special form. M.T.A.C., 11:94–97, 1957.

Meek:1980:GGP

[MH80] B. L. Meek and P. M. Heath, editors. Guide to Good Programming Practice. Ellis Horwood, New York, NY, USA, 1980. ???? pp.

Martin:1987:SDP

[MH87] J. M. Martin and G. A. Hewer. Smallest destabilizing perturbations for linear systems. Internat. J. Control, 45(5): 1495–1504, 1987.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Matsui:1981:OUF

[MI81] Shouichi Matsui and Masao Iri.

An overflow/underflow-free floating-point representation of numbers. Journal of Information Processing, 4(3):123–133, 1981. CODEN JOSHA4. ISSN 0387-6101.

Michie:1980:TOC

[Mic80] Donald Michie. Turing and the origins of the computer. New scientist, 85(195): 580–583, 1980. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Michaelson:1990:HWP

[Mic90] Herbert B. Michaelson. How to Write and Publish Engineering Papers and Reports. Oryx Press, Phoenix, Arizona, third edition, 1990. ISBN 0-89774-650-3. ???? pp.

Michener:1992:JMW

[Mic92a] James A. Michener. James A. Michener’s Writer’s Handbook: Explorations in Writing and Publishing. Random House, New York, NY, USA, 1992. ISBN 0-679-74126-7. ???? pp.

Michener:1992:WMH

[Mic92b] James A. Michener. The World is My Home: A Memoir. Mandarin, London, 1992. ISBN 0-7493-1303-X. 618 pp.

Mikhlin:1991:EAN

[Mik91] Solomon G. Mikhlin. Error Analysis in Numerical Processes. Wiley, Chichester, UK, 1991. ISBN 0-471-92133-5. 283 pp.

Miller:1943:GMM

[Mil43] Glenn Miller. Glenn Miller’s Method for Orchestral Arranging. Chappell, London, 1943. ???? pp.

Miller:1971:LZC

[Mil71a] John J. H. Miller. On the location of zeros of certain classes of polynomials with applications to numerical analysis. J. Inst. Maths Applics, 8:397–406, 1971.

Miller:1971:NIG

[Mil71b] W. Miller. A note on the instability of Gaussian elimination. BIT, 11:422– 424, 1971. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Miller:1975:TNA

[Mil75a] J. J. H. Miller, editor. Topics in Numerical Analysis II. Academic Press, London, 1975. ???? pp.

Miller:1975:CCN

[Mil75b] Webb Miller. Computational complexity and numerical stability. SIAM Journal on Computing, 4(2):97–107, 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1975:CSN

[Mil75c] Webb Miller. Computer search for numerical instability. J. Assoc. Comput. Mach., 22(4):512–521, 1975.

Miller:1975:SRA

[Mil75d] Webb Miller. Software for roundoff analysis. ACM Transactions on Mathematical Software, 1(2):108–128, 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Miller:1976:GTR

[Mil76] Webb Miller. Graph transformations for roundoff analysis. SIAM Journal on Computing, 5(2):204–216, 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1978:PAR

[Mil78] Webb Miller. Performing armchair roundoff analyses of statistical algorithms. Commun. Statist.-Simula. Computa., B7(3):243–255, 1978.

Miller:1984:ENS

[Mil84] Webb Miller. The Engineering of Numerical Software. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984. ISBN 0-13-279043-2. viii + 167 pp.

Miles:1987:DDP

[Mil87a] John Miles. Design for Desktop Publishing: A Guide to Layout and Typography on the Personal Computer. Gordon Fraser, London, 1987. ???? pp.

Miller:1987:SEM

[Mil87b] K. S. Miller. Some Eclectic Matrix Theory. Robert E. Krieger Publishing Co., Krieger Drive, Malabar, FL 32950, 1987. ???? pp.

Milenkovic:1988:VIG

[Mil88a] Victor J. Milenkovic. Verifiable implementations of geometric algorithms using finite precision arithmetic. Artificial Intelligence, 37:377–401, 1988. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Miller:1988:ISM

[Mil88b] D. F. Miller. The iterative solution of the matrix equation XA + BX + C = 0. Linear Algebra and Appl., 105:131–137, 1988.

Minc:1974:SLA

[Min74] Henryk Minc. Six letters from Alexander C. Aitken. Linear and Multilinear Algebra, 2:1–12, 1974. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Minc:1988:NM

[Min88] Henryk Minc. Nonnegative Matrices. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1988. ISBN 0-471-83966-3. xiii + 206 pp.

Miola:1990:CTS

[Mio90] A. M. Miola, editor. Computing Tools for Scientific Problem Solving. Academic Press, London, 1990.

Mirsky:1956:NNM

[Mir56a] L. Mirsky. A note on normal matrices. Amer. Math. Monthly, 63:479, 1956.

Mirsky:1956:SM

[Mir56b] L. Mirsky. The spread of a matrix. Mathematika, 3:127–130, 1956. CODEN MTKAAB. ISSN 0025-5793.

Mirsky:1958:MMN

[Mir58] L. Mirsky. On the minimization of matrix norms. Amer. Math. Monthly, 65: 106–107, 1958.

Mirsky:1959:DEO

[Mir59] L. Mirsky. Diagonal elements of orthogonal matrices. Amer. Math. Monthly, 66:19–22, 1959.

Mirsky:1960:SGF

[Mir60] L. Mirsky. Symmetric gauge functions and unitarily invariant norms. Quarterly Journal of Mathematics, 11:50–59, 1960. CODEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Mirsky:1961:ILA

[Mir61] L. Mirsky. An Introduction to Linear Algebra. Oxford University Press, ????, 1961. ISBN 0-486-66434-1. viii + 440 pp. Reprinted by Dover, New York, 1990.

Mitchell:1948:IMO

[Mit48] Herbert F. Mitchell, Jr. Inversion of a matrix of order 38. M.T.A.C., 3:161–166, 1948.

Mitchell:1987:NWT

[Mit87] Joan P. Mitchell. The New Writer:

Techniques for Writing Well with a Computer. Microsoft Press, Bellevue, WA, USA, 1987. ISBN 1-55615-029-6. ???? pp.

Matula:1985:FPR

[MK85] David W. Matula and Peter Kornerup. Finite precision rational arithmetic: Slash number systems. IEEE Trans. Comput., C-34(1):3–18, 1985.

Montoye:1985:PAS

[ML85] Robert K. Montoye and Duncan H.

Lawrie. A practical algorithm for the solution of triangular systems on a parallel processing system. IEEE Trans. Comput., C-31(11):1076–1082, 1985.

Moler:1987:PMUa

[MLB87a] C. B. Moler, J. N. Little, and S. Bangert. PC-Matlab User’s Guide. The MathWorks, Inc., Natick, Massachusetts, 1987. ???? pp.

Moler:1987:PMUb

[MLB87b] C. B. Moler, J. N. Little, and S. Bangert. Pro-Matlab User’s Guide. The MathWorks, Inc., Natick, Massachusetts, 1987. ???? pp.

Marcus:1964:SMT

[MM64] Marvin Marcus and Henryk Minc. A Survey of Matrix Theory and Matrix Inequalities. Prindle, Weber and Schmidt, Boston, MA, USA, 1964. ISBN 0-486-67102-X. ???? pp. Reprinted by Dover, New York, 1992.

Marcus:1965:ILA

[MM65] Marvin Marcus and Henrykl Minc. Introduction to Linear Algebra. MacMillan Publishing Company, New York, NY, USA, 1965. ???? pp.

Hueller-Merbach:1970:REL

[MM70] H. Müller-Merbach. On Round-Off Errors in Linear Programming, volume 37 of Lecture Notes in Operations Research and Mathematical Systems. Springer-Verlag, Berlin, 1970. 48 pp.

Messina:1982:PMM

[MM82] Paul C. Messina and Almerico Murli, editors. Problems and Methodologies in Mathematical Software Production, volume 142 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-11603-6. ???? pp.

Moler:1983:RSR

[MM83a] Cleve B. Moler and Donald Morrison. Replacing square roots by Pythagorean sums. IBM Journal of Research and Development, 27(6):577–581, 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Moler:1983:SVA

[MM83b] Cleve B. Moler and Donald Morrison. Singular value analysis of cryptograms. Amer. Math. Monthly, 90:78– 87, 1983.

Mukherjee:1988:SPP

[MM88] B. N. Mukherjee and S. S. Maiti.

On some properties of positive definite Toeplitz matrices and their possible applications. Linear Algebra and Appl., 102: 211–240, 1988.

Mangasarian:1978:NP

[MMR78a] O. L. Mangasarian, R. R. Meyer, and S. M. Robinson. Nonlinear Programming 3. Academic Press, New York, USA, 1978. ???? pp.

Mangasarian:1978:NPP

[MMR78b] Olvi L. Mangasarian, Robert R.

Meyer, and Stephen M. Robinson, editors. Nonlinear programming 3: proceedings of the Special Interest Group on Mathematical Programming symposium / conducted by the Computer Sciences Department at the University of Wisconsin-Madison, July 11–13, 1977. Academic Press, New York, USA, 1978. ISBN 0-12-486660-3. LCCN T57.8 .N67 1977.

Masser:1979:SRS

[MN79] D. W. Masser and M. Neumann. On the square roots of strictly quasiaccretive complex matrices. Linear Algebra and Appl., 28:135–140, 1979.

McCullagh:1989:GLM

[MN89] P. McCullagh and J. A. Nelder. Generalized Linear Models. Chapman and Hall, Ltd., London, UK, second edition, 1989. ISBN 0-412-31760-5. xix + 511 pp.

McCurdy:1984:ACD

[MNP84] A. McCurdy, K. C. Ng, and B. N. Parlett. Accurate computation of divided differences of the exponential function. Math. Comp., 43(168):501–528, 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Marshall:1979:ITM

[MO79] Albert W. Marshall and Ingram Olkin. Inequalities: Theory of Majorization and Its Applications. Academic Press, New York, USA, 1979. ISBN 0-12-473750-1. ???? pp.

Marsaglia:1984:GCM

[MO84] George Marsaglia and Ingram Olkin. Generating correlation matrices. SIAM J. Sci. Stat. Comput., 5(2):470–475, 1984.

Modi:1988:PAM

[Mod88] J. J. Modi. Parallel Algorithms and Matrix Computation. Oxford University Press, ????, 1988. ???? pp.

Moharir:1985:ESG

[Moh85] P. S. Moharir. Extending the scope of Golub’s method beyond complex multiplication. IEEE Trans. Comput., C-34 (5):484–487, 1985.

Oleoeller:1965:NQD

[Møl65a] Ole Møller. Note on quasi doubleprecision. BIT, 5:251–255, 1965. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Oleoeller:1965:QDP

[Møl65b] Ole Møller. Quasi double-precision in floating point addition. BIT, 5:37–50, 1965. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic). See also [Møl65a] for remarks on this article.

Moler:1967:IRF

[Mol67] Cleve B. Moler. Iterative refinement in floating point. J. Assoc. Comput. Mach., 14(2):316–321, 1967.

Moler:1974:CRS

[Mol74] Cleve B. Moler. Cramer’s rule on 2by-2 systems. ACM SIGNUM Newsletter, 9(4):13–14, 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1978:TRP

[Mol78] Cleve B. Moler. Three research problems in numerical linear algebra. In G. H. Golub and J. Oliger, editors, Numerical Analysis, volume 22 of Proceedings of Symposia in Applied Mathematics, pages 1–18. American Mathematical Society, Providence, RI, USA, 1978.

Moler:1982:DML

[Mol82a] C. B. Moler. Demonstration of a matrix laboratory. In Hennart [Hen82], pages 84–98.

Moler:1982:MUG

[Mol82b] C. B. Moler. MATLAB users’ guide.

Technical report, Technical Report CS81-1 (revised), Department of Computer Science, University of New Mexico, Albuquerque, ????, 1982. ???? pp.

Moler:1990:CCW

[Mol90a] Cleve B. Moler. Cleve’s corner: The world’s simplest impossible problem. The MathWorks Newsletter, 4(2):6–7, 1990.

Moler:1990:TND

[Mol90b] Cleve B. Moler. Technical note: Double-rounding and implications for numeric computations. The MathWorks Newsletter, 4:6, 1990.

Moler:1993:MMM

[Mol93] Cleve B. Moler. MATLAB’s magical mystery tour. The MathWorks Newsletter, 7(1):8–9, 1993.

Moler:1995:TTN

[Mol95] Cleve B. Moler. A tale of two numbers. SIAM News, 28:1, 16, January 1995. ISSN 0036-1437. Also in Matlab News and Notes, Winter 1995, 10–12.

Moler:1972:MCF

[Mol2a] Cleve B. Moler. Matrix computations with Fortran and paging. Comm. ACM, 15(4):268–270, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1972:ALE

[Mol2b] Cleve B. Moler. Algorithm 423: Linear equation solver. Comm. ACM, 15 (4):274, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Monro:1989:CCF

[Mon89] Donald M. Monro. A Crash Course in Fortran 77. Edward Arnold, London, UK, 1989. ISBN 0-7131-3582-4. ???? pp.

Moon:1936:SBI

[Moo36] Parry Moon. The Scientific Basis of Illuminating Engineering. McGraw-Hill, New York, 1936. xii + 608 pp.

Moore:1966:IA

[Moo66] Ramon E. Moore. Interval Analysis. Prentice-Hall, Englewood Cliffs, NJ, USA, 1966. xi + 145 pp.

Moore:1975:MES

[Moo75] Ramon E. Moore, editor. Mathematical Elements of Scientific Computing. Holt, Rinehart and Winston, New York, 1975. ISBN 0-03-088125-0. ???? pp.

Moore:1979:MAI

[Moo79] Ramon E. Moore. Methods and Applications of Interval Analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-161-4. xi + 190 pp.

Moore:1988:RCR

[Moo88] Ramon E. Moore, editor. Reliability in Computing: the Role of Interval Methods in Scientific Computing, volume 19 of Perspectives in computing. Academic Press, New York, NY, USA, 1988. ISBN 0-12-505630-3. xv + 428 pp. LCCN QA76.9.E94 R45 1988.

Moore:1993:SW

[Moo93] P. Moore. Sky watchers. IMA Bulletin, 29(1/2):1–8, 1993.

Morton:1980:ESM

[Mor80] Patrick Morton. On the eigenvectors of Schur’s matrix. Journal of Number Theory, 12:122–127, 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

More:1990:CNM

[Mor90] J. Moré. A collection of nonlinear model problems. In Allgower and Georg [AG90], pages 723–762. ISBN 0-8218-1131-2. LCCN QA372 .C6374 1990.

Moses:1971:ASG

[Mos71a] Joel Moses. Algebraic simplification: A guide for the perplexed. Comm. ACM, 14(8):527–537, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moses:1971:SIS

[Mos71b] Joel Moses. Symbolic integration: The stormy decade. Comm. ACM, 14: 548–560, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mosier:1986:RNP

[Mos86] R. G. Mosier. Root neighborhoods of a polynomial. Math. Comp., 47(175): 265–273, 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Malcolm:1974:FMS

[MP74] Michael A. Malcolm and John Palmer. A fast method for solving a class of tridiagonal linear systems. Comm. ACM, 17(1):14–17, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:1977:CPM

[MP77] Carl D. Meyer, Jr. and R. J. Plemmons. Convergent powers of a matrix with applications to iterative methods for singular linear systems. SIAM J. Numer. Anal., 14(4):699–705, 1977.

Markus:1983:CJS

[MP83] A. S. Markus and E. È. Parilis. The change of the Jordan structure of a matrix under small perturbations. Linear Algebra and Appl., 54:139–152, 1983.

Martin:1966:IRS

[MPW66] R. S. Martin, G. Peters, and J. H.

Wilkinson. Iterative refinement of the solution of a positive definite system of equations. Numer. Math., 8:203–216, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 31–44], Contribution I/2.

Metcalf:1987:FE

[MR87] M. Metcalf and J. K. Reid. Fortran 8X Explained. Oxford University Press, ????, 1987. ???? pp.

Mayes:1989:BCF

[MR89] P. Mayes and G. Radicati. Banded Cholesky factorization using level 3 BLAS. Technical report, LAPACK Working Note #12, Mathematics and Computer Science Division, Argonne National Laboratory, Illinois, ????, 1989. ???? pp.

Metcalf:1990:FE

[MR90] Michael Metcalf and John Ker Reid.

Fortran 90 Explained. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853772-7 (paperback). xiv + 294 pp. LCCN QA76.73.F28 M48 1990. US$22.45. See [?].

Madsen:1976:MMD

[MRK76] Niel K. Madsen, Garry H. Rodrigue, and Jack I. Karush. Matrix multiplication by diagonals on a vector/parallel processor. Information Processing Letters, 5 (2):41–45, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meyers:1964:W

[MS64] Norman G. Meyers and James Serrin. H = W. Proceedings of the National Academy of Sciences of the United States of America, 51:1055–1056, 1964. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

McCarthy:1974:OCM

[MS74a] Charles McCarthy and Gilbert Strang. Optimal conditioning of matrices. SIAM J. Numer. Anal., 10(2):370– 388, 1974.

Merle:1974:CED

[MS74b] G. Merle and H. Späth. Computational experiences with discrete L_{p}- approximation. Computing, 12:315–321, 1974. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Miller:1976:AGF

[MS76] W. Miller and D. L. Spooner. Automatic generation of floating-point test data. IEEE Transactions on Software Engineering, SE-2(3):223–226, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702369.

Miller:19xx:SRA

[MS78a] Webb Miller and David Spooner. Software for roundoff analysis, II. ACM Transactions on Mathematical Software, 4(4):369–387, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Miller:19xx:ASR

[MS78b] Webb Miller and David Spooner. Algorithm 532: Software for roundoff analysis. ACM Transactions on Mathematical Software, 4(4):388–390, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Moler:1978:HFA

[MS78c] Cleve B. Moler and G. W. Stewart. On the Householder-Fox algorithm for decomposing a projection. J. Comput. Phys., 28:82–91, 1978.

Moler:1979:EMF

[MS79a] C. B. Moler and G. W. Stewart. An efficient matrix factorization for digital image processing. Technical report, Report LA-7637, Los Alamos Scientific Laboratory, ????, 1979. ???? pp.

More:1979:UDN

[MS79b] Jorge J. Moré and Danny C.

Sorensen. On the use of directions of negative curvature in a modified Newton method. Math. Prog., 16:1–20, 1979. CO-DEN MHPGA4. ISSN 0025-5610.

More:1983:CTR

[MS83] J. J. Moré and D. C. Sorensen. Computing a trust region step. SIAM J. Sci. Stat. Comput., 4:553–572, 1983.

More:1984:NM

[MS84] J. J. Moré and D. C. Sorensen. Newton’s method. In Golub [Gol84], pages 29–82.

Mangasarian:1986:VCN

[MS86] O. L. Mangasarian and T.-H. Shiau. A variable-complexity norm maximization problem. SIAM J. Alg. Disc. Meth., 7: 455–461, 1986.

McComb:1988:ESS

[MS88] J. McComb and S. Schmidt. Engineering and scientific subroutine library for the IBM 3090 vector facility. IBM Systems Journal, 27:404–415, 1988. CODEN IBMSA7. ISSN 0018-8670.

Miller:1989:HNS

[MS89] Casey Miller and Kate Swift. The Handbook of Non-Sexist Writing for Writers, Editors and Speakers. The Women’s Press, London, 1989. ISBN 0-7043-3878-5. ???? pp. Second British edition, revised and updated.

Meyer:1991:CAP

[MS91] Kenneth R. Meyer and Dieter S. Schmidt, editors. Computer Aided Proofs in Analysis, volume 28 of IMA Volumes in Mathematics and its Applications. Springer-Verlag, New York, 1991. ISBN 0-387-97426-1.

Mathias:1993:BQA

[MS93a] R. Mathias and G. W. Stewart. A block QR algorithm and the singular value decomposition. Linear Algebra and Appl., 188:91–100, 1993.

Messina:1993:SST

[MS93b] Paul Messina and Thomas Sterling, editors. System Software and Tools for High Performance Computing Environments. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-326-9. ???? pp.

More:1984:MP

[MSGH84] J. J. Moré, D. C. Sorensen, B. S. Garbow, and K. E. Hillstrom. The MIN-PACK project. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 88–111. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

McAllister:1984:RRR

[MSS84] D. F. McAllister, G. W. Stewart, and W. J. Stewart. On a Rayleigh–Ritz refinement technique for nearly uncoupled stochastic matrices. Linear Algebra and Appl., 60:1–25, 1984.

Miranker:1986:ASC

[MT86] Willard L. Miranker and Richard A.

Toupin, editors. Accurate Scientific Computations, Proceedings, 1985, volume 235 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1986.

Miranda:1993:TIS

[MT93] H. Miranda and Robert C. Thompson. A trace inequality with a subtracted term. Linear Algebra and Appl., 182:165– 172, 1993.

Marques:1994:SPC

[MT94] Osni A. Marques and Vincent Toumazou. Spectral portrait computation by a Lanczos method (normal equation version). Technical Report TR/PA/95/02, CERFACS, Toulouse, France, December 1994. ???? pp.

Marques:1995:SPC

[MT95] Osni A. Marques and Vincent Toumazou. Spectral portrait computation by a Lanczos method (augmented matrix version). Technical Report TR/PA/95/05, CERFACS, Toulouse, France, February 1995. ???? pp.

Muirhead:1982:AMS

[Mui82] Robb J. Muirhead. Aspects of Multivariate Statistical Theory. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1982. ISBN 0-471-09442-0. xix + 673 pp.

Hueller:1983:REA

[Mül83] K. H. Müller. Rounding error analysis of Horner’s scheme. Computing, 30:285–303, 1983. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1989:AO

[Mul89] Jean-Michel Muller. Arithmétique des Ordinateurs. Masson, Masson, France, 1989. ISBN 2-225-81689-1. 214 pp. In French. Cited in [FM91].

Murray:1972:NMU

[Mur72] W. Murray, editor. Numerical Methods for Unconstrained Optimization. Academic Press, ????, 1972.

Murdock:1976:NAS

[Mur76] Phoebe J. Murdock. New alphabets and symbols for typesetting mathematics. Scholarly Publishing, 8(1):44–53, 1976. ISSN 0036-634X.

Murray:1993:MB

[Mur93] J. D. Murray. Mathematical Biology. Springer-Verlag, Berlin, second edition, 1993. ISBN 3-540-57204-X. xiv + 767 pp.

Moler:1978:NDW

[MV78] Cleve B. Moler and Charles F. Van Loan. Nineteen dubious ways to compute the exponential of a matrix. SIAM Review, 20(4):801–836, 1978.

Miller:1980:SRA

[MW80] Webb Miller and Celia Wrathall. Software for Roundoff Analysis of Matrix Algorithms. Academic Press, New York, 1980. ISBN 0-12-497250-0. x + 151 pp.

Mohamed:1986:NA

[MW86] J. L. Mohamed and J. E. Walsh, editors. Numerical Algorithms. Oxford University Press, ????, 1986. ???? pp.

More:1993:OSG

[MW93] Jorge J. Moré and Stephen J. Wright.

Optimization Software Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-322-6. xii + 154 pp.

Moor:1991:TGO

[MZ91] B. L. R. De Moor and H. Zha. A tree of generalizations of the ordinary singular value decomposition. Linear Algebra and Appl., 147:469–500, 1991.

Milnazzo:1987:RGC

[MZB87] F. Milnazzo, C. Zala, and I. Barrodale. On the rate of growth of condition numbers for convolution matrices. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-35:471–475, 1987. CODEN IETABA. ISSN 0096-3518.

Anonymous:19xx:D

[NA xx] NA Digest. Electronic mail maga-

zine: na.help@na-net.ornl.gov, 19xx.

Nabben:1991:GHM

[Nab91] Reinhard Nabben. On generalized H-matrices. Preprint 91-044, Universität Bielefeld, Germany, 1991. ???? pp.

Nabben:1992:CMW

[Nab92] Reinhard Nabben. On a class of matrices which arise in the numerical solution of Euler equations. Numer. Math., 63:411–431, 1992. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Nabben:19xx:NAG

[Nab9x] Reinhard Nabben. A new application for generalized M-matrices. In Numerical Linear Algebra, pages 179–192, 199x.

NAG:1990:NFL

[NAG90] NAG Ltd., Oxford. NAG Fortran Library Manual, Mark 14, 1990. ???? pp.

NAG:1992:NFR

[NAG92] NAG Ltd., Oxford. NAGWare FTN90 Reference Manual, 1992. ???? pp.

Namias:1985:SDR

[Nam85] V. Namias. Simple derivation of the roots of a cubic equation. American Journal of Physics, 53(8):775, 1985. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Nashed:1976:GIA

[Nas76] M. Zuhair Nashed, editor. Generalized Inverses and Applications. Academic Press, New York, 1976. ISBN 0-12-514250-1.

Nash:1979:CNM

[Nas79] J. C. Nash. Compact Numerical Methods for Computers: Linear Algebra and Function Minimisation. Adam Hilger (Bristol) and John Wiley (New York), ????, 1979. ???? pp.

Nash:1984:ESP

[Nas84] J. C. Nash. Effective Scientific Problem Solving with Small Computers. Reston Publishing Company, Virginia, 1984. ???? pp.

Nash:1985:DIV

[Nas85] J. C. Nash. Design and implementation of a very small linear algebra program package. Comm. ACM, 28:89–94, 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nash:1990:CNM

[Nas90a] J. C. Nash. Compact Numerical Methods for Computers: Linear Algebra and Function Minimisation. Adam Hilger, Bristol, second edition, 1990. ISBN 0-85274-319-X. ???? pp.

Nash:1990:HSC

[Nas90b] Stephen. G. Nash, editor. A History of Scientific Computing, ACM Press History Series. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50814-1.

Neill:1982:BNB

[Nei82] A. R. Neill. Bob Neill’s Book of Typewriter Art (With Special Computer Programme). The Weavers Press, 4 Weavers Cottages, Goudhurst, Kent, 1982. ISBN 0-946017-01-8. ???? pp.

Neill:1984:BNS

[Nei84] A. R. Neill. Bob Neill’s Second Book of Typewriter Art. The Weavers Press, 4 Weavers Cottages, Goudhurst, Kent, 1984. ISBN 0-946017-02-6. ???? pp.

Neumaier:1974:REV

[Neu74] A. Neumaier. Rundungsfehleranalyse einiger Verfahren zur Summation endlicher Summen. Z. Angew. Math. Mech, 54:39–51, 1974.

Neumaier:1985:IPR

[Neu85] A. Neumaier. Inner product rounding error analysis in the presence of underflow. Computing, 34:365–373, 1985. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumaier:1986:CMM

[Neu86] A. Neumaier. On the comparison of H-matrices with M-matrices. Linear Algebra and Appl., 83:135–141, 1986.

Neufeldt:1988:WNW

[Neu88] Victoria Neufeldt, editor. Webster’s New World Dictionary of American English. Simon and Schuster, New York, NY, USA, third college edition, 1988. ISBN 0-13-949280-1. xxvi + 1574 pp. LCCN PE1628 .W5633 1988.

Neumaier:1992:HNB

[Neu92] A. Neumaier. Hybrid norms and bounds for overdetermined linear systems. Manuscript, 1992.

Newcomb:1881:NFU

[New81] Simon Newcomb. Note on the frequency of use of the different digits in natural numbers. Amer. J. Math., 4: 39–40, 1881. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic). Cited in [Rai76].

Newman:1962:MC

[New62] Morris Newman. Matrix computations. In John Todd, editor, Survey of Numerical Analysis, pages 222–254. Mc-Graw-Hill, New York, 1962.

Newbery:1973:EAF

[New73] A. C. R. Newbery. Error analysis for Fourier series evaluation. Math. Comp., 27(123):639–644, 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1974:EAP

[New74] A. C. R. Newbery. Error analysis for polynomial evaluation. Math. Comp., 28 (127):789–793, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ng:1984:CCM

[Ng84] K. C. Ng. Contributions to the computation of the matrix exponential. Technical Report PAM-212, Center for Pure and Applied Mathematics, University of California, Berkeley, February 1984. ???? pp.

Nickel:1970:KBU

[Nic70] K. Nickel. Das Kahan-Babus̆kasche Summierungsverfahren in Triplex-ALGOL 60. Z. Angew. Math. Mech, 50:369–373, 1970.

Nickel:1975:IM

[Nic75] Karl Nickel, editor. Interval Mathematics, volume 29 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1975.

Nickel:1977:IA

[Nic77] Karl Nickel. Interval-analysis. In David A. H. Jacobs, editor, The State of the Art in Numerical Analysis, pages 193– 225. Academic Press, London, 1977.

Nickel:1980:IM

[Nic80] Karl L. E. Nickel, editor. Interval Mathematics 1980. Academic Press, ????, 1980. ISBN 0-12-518850-1.

Nickel:1985:IM

[Nic85] Karl Nickel, editor. Interval Mathematics 1985, volume 212 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1985.

Nix:1981:ESE

[Nix81] Robert P. Nix. Experience with a space efficient way to store a dictionary. Comm. ACM, 24(5):297–298, 1981. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nelder:1965:SMF

[NM65] J. A. Nelder and R. Mead. A simplex method for function minimization. The Computer Journal, 7:308–313, 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Newby:1990:CCC

[NM90] Peter Newby and John Meade. The Complete COUNTDOWN Companion. Pelham Books, Stephen Greene Press, ????, 1990. ISBN 0-7207-1935-6. ???? pp.

Noble:1969:ALA

[Nob69] B. Noble. Applied Linear Algebra. Prentice-Hall, Englewood Cliffs, NJ, USA, 1969. ???? pp.

Nocedal:1992:TAU

[Noc92] J. Nocedal. Theory of algorithms for unconstrained optimization. In Acta Numerica, pages 199–242. Cambridge University Press, 1992.

Noda:1971:NCM

[Nod71] Takashi Noda. Note on the computation of the maximal eigenvalue of a non-negative irreducible matrix. Numer. Math., 17:382–386, 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nour-Omid:1984:HIS

[NOP84] N. Nour-Omid and B. N. Parlett. How to implement the spectral transformation. Technical report, Technical Report PAM-224, Center for Pure and Applied Mathematics, University of California, Berkeley, ????, 1984. ???? pp.

Norman:1991:D

[Nor91] David Norman. Dinosaur! Boxtree, London, ????, 1991. ???? pp.

Notay:1993:CRC

[Not93] Yvan Notay. On the convergence rate of the conjugate gradients in presence of rounding errors. Numer. Math., 65:301–317, 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neumann:1978:CNM

[NP78] M. Neumann and R. J. Plemmons.

Convergent nonnegative matrices and iterative methods for consistent linear systems. Numer. Math., 31:265–279, 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neumann:1984:BEA

[NP84] M. Neumann and R. J. Plemmons.

Backward error analysis for linear systems associated with inverses of Hmatrices. BIT, 24:102–112, 1984. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Nickel:1972:TCN

[NR72] K. Nickel and K. Ritter. Termination criterion and numerical convergence. SIAM J. Numer. Anal., 9(2):277–283, 1972.

Nashed:1976:ABG

[NR76] M. Zuhair Nashed and L. B. Rall. Annotated bibliography on generalized inverses and applications. In M. Zuhair Nashed, editor, Generalized Inverses and Applications, pages 771–1041. Academic Press, New York, 1976.

Nachtigal:1992:HFN

[NRT92a] N. M. Nachtigal, S. C. Reddy, and L. N. Trefethen. How fast are nonsymmetric matrix iterations? SIAM J. Matrix Anal. Appl., 13(3):778–795, 1992.

Nachtigal:1992:HGA

[NRT92b] N. M. Nachtigal, L. Reichel, and L. N. Trefethen. A hybrid GMRES algorithm for nonsymmetric linear systems. SIAM J. Matrix Anal. Appl., 13(3):796– 825, 1992.

Nash:1990:ASD

[NS90] S. G. Nash and A. Sofer. Assessing a search direction within a truncated-Newton method. to appear in Operations Research Letters, April 1990, 1990.

Newman:1958:EMI

[NT58] Morris Newman and John Todd. The evaluation of matrix inversion programs. J. Soc. Indust. Appl. Math., 6(4):466–476, 1958.

OCarroll:1973:IRC

[O’C73] M. J. O’Carroll. Inconsistencies and S.O.R. convergence for the discrete Neumann problem. J. Inst. Maths Applics, 11:343–350, 1973.

OConnor:1978:ESB

[O’C78] Maeve O’Connor. Editing Scientific Books and Journals. Pitman Medical, Tunbridge Wells, Kent, 1978. ISBN 0-272-79517-8. ???? pp.

OConnor:1986:HCS

[O’C86] Maeve O’Connor. How to Copyedit Scientific Books and Journals. ISI Press, Philadelphia, 1986. ISBN 0-89495-064-9. ???? pp.

OConnor:1991:WSS

[O’C91] Maeve O’Connor. Writing Successfully in Science. Chapman and Hall, Ltd., London, UK, 1991. ISBN 0-412-44630-8. xi + 229 pp.

OCinneide:1993:EPT

[O’C93] Colm Art O’Cinneide. Entrywise perturbation theory and error analysis for Markov chains. Numer. Math., 65: 109–120, 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ocker:1975:PHE

[Ock75] W. A. Ocker. A program to hyphenate English words. IEEE Trans. Prof. Commun., PC-18:78–84, 1975.

Oden:1990:HCF

[Ode90] J. Tinsley Oden. Historical comments on finite elements. In Stephen G. Nash, editor, A History of Scientific Computing, pages 152–166. Addison-Wesley, Reading, MA, USA, 1990.

Oettli:1965:SSL

[Oet65] W. Oettli. On the solution set of a linear system with inaccurate coefficients. SIAM J. Numer. Anal., 2(1):115– 118, 1965.

Ogryczak:1988:SMA

[Ogr88] W. Ogryczak. The simplex method is not always well behaved. Linear Algebra and Appl., 109:41–57, 1988.

OLeary:1980:EMC

[O’L80] Dianne Prost O’Leary. Estimating matrix condition numbers. SIAM J. Sci. Stat. Comput., 1(2):205–209, 1980.

OLeary:1990:BSP

[O’L90] Dianne P. O’Leary. On bounds for scaled projections and pseudoinverses. Linear Algebra and Appl., 132:115–117, 1990.

OLeary:1993:IMF

[O’L93a] Dianne P. O’Leary. Iterative methods for finding the stationary vector for Markov chains. In Carl D. Meyer and Robert J. Plemmons, editors, Linear Algebra, Markov Chains, and Queueing Models, volume 48 of IMA Volumes in Mathematics and its Applications, pages 125–136. Springer-Verlag, New York, 1993.

OLeary:1993:WBN

[O’L93b] Dianne P. O’Leary. Why Broyden’s nonsymmetric method terminates on linear equations. Technical Report CS-TR 3045, Department of Computer Science, University of Maryland, ????, March 1993. ???? pp.

Olagnon:1993:FFF

[Ola93] Michel Olagnon. F90 FAN’s: Fortran 90 frequently asked about news. December 7, 1993. Available by anonymous ftp

from molene.ifremer.fr, 1993.

Oliver:1975:SME

[Oli75] J. Oliver. Stable methods for evaluating the points cos(iπ/n). J. Inst. Maths Applics, 16:247–257, 1975.

Oliver:1977:EAM

[Oli77] J. Oliver. An error analysis of the modified Clenshaw method for evaluating Chebyshev and Fourier series. J. Inst. Maths Applics, 20:379–391, 1977.

Oliver:1979:REP

[Oli79] J. Oliver. Rounding error propagation in polynomial evaluation schemes. J. Comp. Appl. Math., 5(2):85–97, 1979.

Olver:1978:NAE

[Olv78] F. W. J. Olver. A new approach to error arithmetic. SIAM J. Numer. Anal., 15(2):368–393, 1978.

Olver:1983:EAC

[Olv83a] F. W. J. Olver. Error analysis of complex arithmetic. In H. Werner, L. Wuytack, E. Ng., and H. J. Bünger, editors, Computational Aspects of Complex Analysis, volume 102 of NATO Advanced Study Institute Series C, pages 279–292. D. Reidel, Dordrecht, Holland, 1983.

Olver:1983:EBA

[Olv83b] F. W. J. Olver. Error bounds for arithmetic operations on computers without guard digits. IMA J. Numer. Anal., 3:153–160, 1983.

Olver:1986:EBP

[Olv86] F. W. J. Olver. Error bounds for polynomial evaluation and complex arithmetic. IMA J. Numer. Anal., 6:373–379, 1986.

Olschowka:1993:NPS

[ON93] Markus Olschowka and Arnold Neumaier. A new pivoting strategy for Gaussian elimination. Manuscript, 1993.

Oettli:1964:CAS

[OP64] W. Oettli and W. Prager. Compatibility of approximate solution of linear equations with given error bounds for coefficients and right-hand sides. Numer. Math., 6:405–409, 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

OLeary:1983:DIC

[OP83] D. P. O’Leary and S. Peleg. Digital image compression by outer product expansion. IEEE Transactions on Communications, COM-31:441–444, 1983. CO-DEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Oettli:1965:ASL

[OPW65] W. Oettli, W. Prager, and J. H. Wilkinson. Admissible solutions of linear systems with not sharply defined coefficients. SIAM J. Numer. Anal., 2(2): 291–299, 1965.

Or:1994:DTT

[Or94] A. C. Or. The dynamics of a Tippe top. SIAM J. Appl. Math., 54(3):597–609, 1994.

Orr:1974:JC

[Orr74] Carol Orr. Jellied constellations. Scholarly Publishing, ??:66–68, 1974. ISSN 0036-634X.

Orszagh:1969:WBE

[Ors69] Ladislas Országh. Wanted: Better English dictionaries. English Language Teaching, 23(3):216–222, 1969. CODEN ELNTA7. ISSN 0013-8290.

Ortega:1963:EAH

[Ort63] James M. Ortega. An error analysis of Householder’s method for the symmetric eigenvalue problem. Numer. Math., 5:211–225, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ortega:1972:NAS

[Ort72] James M. Ortega. Numerical Analysis: A Second Course. Academic Press, New York, 1972. ISBN 0-89871-250-5. xiii + 201 pp. Reprinted by Society for Industrial and Applied Mathematics, Philadelphia, 1990.

Ortega:1987:MTS

[Ort87] James M. Ortega. Matrix Theory: A Second Course. Plenum Press, New York, NY, USA; London, UK, 1987. ISBN 0-306-42433-9. xii + 262 pp.

Ortega:1988:IPV

[Ort88] James M. Ortega. Introduction to Parallel and Vector Solution of Linear Systems. Plenum Press, New York, NY, USA; London, UK, 1988. ISBN 0-306-42862-8. ???? pp.

OLeary:1990:CEE

[OS90] D. P. O’Leary and G. W. Stewart.

Computing the eigenvalues and eigenvectors of symmetric arrowhead matrices. J. Comput. Phys., 90:497–505, 1990.

Osborne:1960:PCM

[Osb60] E. E. Osborne. On pre-conditioning of matrices. J. Assoc. Comput. Mach., 7: 338–345, 1960.

Osborne:1965:DMS

[Osb65] M. R. Osborne. Direct methods for the solution of finite-difference approximations to separable partial differential equations. The Computer Journal, 8(2):

150–156, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Osborne:1978:IIN

[Osb78] M. R. Osborne. Inverse iteration, Newton’s method, and non-linear eigenvalue problems. pages 21–53, 1978.

Osborne:1985:FAO

[Osb85] M. R. Osborne. Finite Algorithms in Optimization and Data Analysis. Wiley, Chichester, UK, 1985. ???? pp.

Ostrowski:1954:SOP

[Ost54] A. M. Ostrowski. On the spectrum of a one-parametric family of matrices. Journal für die reine und angewandte Mathematik, 193(3/4):143–160, 1954.

Ostrowski:1973:SEE

[Ost73] A. M. Ostrowski. Solution of Equations in Euclidean and Banach Spaces. Academic Press, New York, 1973. ISBN 0-12-530260-6. xx + 412 pp. Third edition of Solution of Equations and Systems of Equations.

Osterweil:1983:TES

[Ost83] L. J. Osterweil. TOOLPACK—an experimental software development environment research project. IEEE Trans. Software Engrg., SE-9:673–685, 1983.

OLeary:1979:ELE

[OSV79] D. P. O’Leary, G. W. Stewart, and J. S. Vandergraft. Estimating the largest eigenvalue of a positive definite matrix. Math. Comp., 33:1289–1292, 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ortega:1985:SPD

[OV85] J. M. Ortega and R. G. Voigt. Solution of partial differential equations on vector and parallel computers. SIAM Review, 27(2):149–240, 1985.

Overton:1983:QCM

[Ove83] M. L. Overton. A quadratically convergent method for minimizing a sum of Euclidean norms. Math. Prog., 27:34–63, 1983. CODEN MHPGA4. ISSN 0025-5610.

Ortega:1990:BPV

[OVR90] J. M. Ortega, R. G. Voigt, and C. H.

Romine. A bibliography on parallel and vector numerical algorithms. In Parallel Algorithms for Matrix Computations, pages 125–197. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-260-2.

OConnor:1977:WSP

[OW77] Maeve O’Connor and F. Peter Woodford. Writing Scientific Papers in English. Pitman Medical, Tunbridge Wells, Kent, 1977. ISBN 0-272-79515-1. ???? pp.

Olver:1982:EPE

[OW82] F. W. J. Olver and J. H. Wilkinson. A posteriori error bounds for Gaussian elimination. IMA J. Numer. Anal., 2:377–406, 1982.

Osborne:1985:ATA

[OW85] M. R. Osborne and G. A. Watson. An analysis of the total approximation problem in separable norms, and an algorithm for the total l_{1} problem. SIAM J. Sci. Stat. Comput., 6:410–424, 1985.

OLeary:1990:PQF

[OW90] D. P. O’Leary and P. Whitman. Parallel QR factorization by Householder and modified Gram–Schmidt algorithms. Parallel Computing, 16:99–112, 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Press:1989:COE

[Oxf89a] Oxford University Press. The Compact Oxford English Dictionary. ????, second edition, 1989. ???? pp.

Press:1989:OAL

[Oxf89b] Oxford University Press. Oxford Advanced Learner’s Dictionary of Current English. ????, fourth edition, 1989. ISBN 0-19-431141-4. ???? pp.

Press:1989:OED

[Oxf89c] Oxford University Press. The Oxford English Dictionary. ????, second edition, 1989. ???? pp.

Press:1990:COD

[Oxf90] Oxford University Press. The Concise Oxford Dictionary of Current English. ????, eighth edition, 1990. ISBN 0-19-861200-1. ???? pp.

Press:1993:NSO

[Oxf93] Oxford University Press. The New Shorter Oxford English Dictionary. ????, 1993. ISBN 0-19-861134-X. ???? pp.

Paul:1988:PSC

[PA88] George Paul and George S. Almasi, editors. Parallel Systems and Computation: Proceedings of the 1986 IBM Europe Institute–Seminar on Parallel Computing, Oberlech, Austria, August 11–15, 1986. Elsevier North Holland, Inc., New York, NY, USA, 1988. x + 294 pp.

Pothen:1992:FRA

[PA92] A. Pothen and F. L. Alvarado. A fast reordering algorithm for parallel sparse triangular solution. SIAM J. Sci. Stat. Comput., 13(2):645–653, 1992.

Paardekooper:1971:EAS

[Paa71] M. H. C. Paardekooper. An eigenvalue algorithm for skew-symmetric matrices. Numer. Math., 17:189–202, 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paige:1973:EAM

[Pai73] C. C. Paige. An error analysis of a method for solving matrix equations. Math. Comp., 27(122):355–359, 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paige:1979:FNS

[Pai79] C. C. Paige. Fast numerically stable computations for generalized linear least squares problems. SIAM J. Numer. Anal., 16:165–171, 1979.

Paige:1980:EAS

[Pai80] C. C. Paige. Error analysis of some techniques for updating orthogonal decompositions. Math. Comp., 34:465–471, 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paige:1981:PNA

[Pai81] C. C. Paige. Properties of numerical algorithms related to computing controllability. IEEE Trans. Automat. Control., AC-26:130–138, 1981.

Paige:1985:CMR

[Pai85a] C. C. Paige. Covariance matrix representation in linear filtering. In B. N. Datta, editor, Linear Algebra and its Role in Systems Theory, volume 47 of Contemporary Math., pages 309–321. American Mathematical Society, Providence, RI, USA, 1985.

Paige:1985:GLM

[Pai85b] C. C. Paige. The general linear model and the generalized singular value decomposition. Linear Algebra and Appl., 70: 269–284, 1985.

Paige:1986:CGS

[Pai86] C. C. Paige. Computing the generalized singular value decomposition. SIAM J. Sci. Stat. Comput., 7(4):1126–1146, 1986.

Paige:1987:SAG

[Pai87] C. C. Paige. Some aspects of generalized QR factorizations. To appear in Cox & Hammarling., 1987.

Paige:1994:KSP

[Pai94] C. C. Paige. Krylov subspace processes, Krylov subspace methods, and iteration polynomials. In J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors, Proceedings of the Cornelius Lanczos International Centenary Conference, pages 83–92. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Palais:1986:ITA

[Pal86] Richard S. Palais. An introduction to TEX and AMS-TEX. Notices Amer. Math. Soc., 33:303–308, 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pan:1978:SAO

[Pan78] V. Pan. Strassen algorithm is not optimal. Trilinear technique of aggregating, uniting and canceling for constructing fast algorithms for matrix multiplication. In Proc. 19th Annual Symposium on the Foundations of Computer Science, pages 166–176. Ann Arbor, MI, USA, Ann Arbor, MI, USA, 1978.

Pan:1984:HMM

[Pan84a] V. Pan. How to Multiply Matrices Faster, volume 179 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1984. ???? pp.

Pan:1984:TTA

[Pan84b] V. Ya. Pan. The techniques of trilinear aggregating and the recent progress in the asymptotic acceleration of matrix operations. Theoret. Comput. Sci., 33:117– 138, 1984. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pan:1984:HCW

[Pan84c] Victor Pan. How can we speed up matrix multiplication? SIAM Review, 26 (3):393–415, 1984.

Pan:1990:CDS

[Pan90] V. Pan. On computations with dense structured matrices. Math. Comp., 55: 179–190, 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pan:1992:CCM

[Pan92] Victor Pan. Complexity of computations with matrices and polynomials. SIAM Review, 34(2):225–262, 1992.

Paprzycki:1992:CGE

[Pap92] M. Paprzycki. Comparisons of Gaussian elimination algorithms on a Cray Y-MP. Linear Algebra and Appl., 172:57– 69, 1992.

Papadimitriou:1993:KNA

[Pap93a] Pythagoras Papadimitriou. The KSR1—A numerical analyst’s perspective. Numerical Analysis Report No. 242, University of Manchester, Manchester, England, December 1993. ???? pp.

Papadimitriou:1993:PSS

[Pap93b] Pythagoras Papadimitriou. Parallel Solution of SVD-Related Problems, With Applications. PhD thesis, University of Manchester, Manchester, England, October 1993. ???? pp.

Parlett:1964:LMA

[Par64] Beresford N. Parlett. Laguerre’s method applied to the matrix eigenvalue problem. Math. Comp., 18:464–485, 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1965:MEP

[Par65] Beresford N. Parlett. Matrix eigenvalue problems. Amer. Math. Monthly, 72(2):59–66, February 1965.

Parlett:1967:A

[Par67] B. N. Parlett. The LU and QR algorithms. In Ralston and Wilf [RW67], pages 116–130.

Parlett:1971:AAR

[Par71] Beresford N. Parlett. Analysis of algorithms for reflections in bisectors. SIAM Review, 13(2):197–208, 1971.

Partridge:1973:UAG

[Par73] Eric Partridge. Usage and Abusage: A Guide to Good English. Penguin, London, 1973. ISBN 0-14-051024-9. ???? pp.

Parlett:1974:CFT

[Par74] B. N. Parlett. Computation of functions of triangular matrices. Memorandum ERL-M481, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, November 1974. ???? pp.

Parlett:1976:RAE

[Par76] B. N. Parlett. A recurrence among the elements of functions of triangular matrices. Linear Algebra and Appl., 14:117– 121, 1976.

Parlett:1978:PNA

[Par78] B. N. Parlett. Progress in numerical analysis. SIAM Review, 20:443–456, 1978.

Parlett:1980:SEP

[Par80] Beresford N. Parlett. The Symmetric Eigenvalue Problem. Series in Computational Mathematics. Prentice-Hall, Englewood Cliffs, NJ, USA, 1980. ISBN 0-13-880047-2. xix + 348 pp. LCCN QA188 .P3. Prentice-Hall Series in Computational Mathematics.

Parlett:1982:WFT

[Par82] B. N. Parlett. Winograd’s Fourier transform via circulants. Linear Algebra and Appl., 45:137–155, 1982.

Parlett:1984:SSE

[Par84] B. N. Parlett. The software scene in the extraction of eigenvalues from sparse matrices. SIAM J. Sci. Stat. Comput., 5 (3):590–604, 1984.

Parkinson:1985:BCG

[Par85] C. L. Parkinson. Breakthroughs:

A Chronology of Great Achievements in Science and Mathematics 1200–1930. Mensell, London, 1985. ???? pp.

Parter:1986:DSV

[Par86] Seymour V. Parter. On the distribution of the singular values of Toeplitz matrices. Linear Algebra and Appl., 80:115– 130, 1986.

Parberry:1990:GNR

[Par90a] Ian Parberry. A guide for new referees in theoretical computer science. Manuscript, Department of Computer Sciences, University of North Texas, Den-

ton, Texas (ian@ponder.csci.unt.edu),

April 1990.

Park:1990:PAU

[Par90b] H. Park. A parallel algorithm for the unbalanced orthogonal Procrustes problem. Parallel Computing, 17:913–923, 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Parlett:1990:CJH

[Par90c] Beresford N. Parlett. The contribution of J. H. Wilkinson to numerical analysis. In Stephen G. Nash, editor, A History of Scientific Computing, pages 17– 30. Addison-Wesley, Reading, MA, USA, 1990.

Parlett:1990:MLA

[Par90d] Beresford N. Parlett. Misconvergence in the Lanczos algorithm. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 7–24. Oxford University Press, 1990.

Park:1992:MED

[Par92a] H. Park. On multiple error detection in matrix triangularizations using checksum methods. Journal of Parallel and Distributed Computing, 14:90–97, 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Parlett:1992:RTF

[Par92b] B. N. Parlett. Reduction to tridiagonal form and minimal realizations. SIAM J. Matrix Anal. Appl., 13(2):567– 593, 1992.

Parlett:1992:SBI

[Par92c] B. N. Parlett. Some basic information on information-based complexity theory. Bulletin of the American Mathematical Society, 26(1):3–27, 1992. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Partington:1992:ODQ

[Par92d] Angela Partington. The Oxford Dictionary of Quotations. Oxford University Press, ????, fourth edition, 1992. ISBN 0-19-866185-1. xvi + 1061 pp.

Parberry:1993:HPP

[Par93] Ian Parberry. How to present a paper in theoretical computer science: A speaker’s guide for students. Manuscript, Department of Computer Sciences, University of North Texas, Denton, Texas

(ian@ponder.csci.unt.edu), July 1993.

Parlett:1994:DWF

[Par94] Beresford N. Parlett. Do we fully understand the symmetric Lanczos algorithm yet? In J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors, Proceedings of the Cornelius Lanczos International Centenary Conference, pages 93–107. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Pascal:1975:TE

[Pas75] Naomi B. Pascal. Two enchirdia.

Scholarly Publishing, 6(4):363–369, 1975. ISSN 0036-634X.

Pascal:1985:FRA

[Pas85] Naomi B. Pascal. Freedom, responsibility, and the agile editor. Scholarly Publishing, ??(??):255–261, 1985. ISSN 0036-634X.

Patashnik:1988:BI

[Pat88] Oren Patashnik. BibTEXing. Documentation for general BibTEX users. File btxdoc.tex in the standard distribution, February 8, 1988.

Paxson:1991:PTI

[Pax91] Vern Paxson. A program for testing IEEE decimal–binary conversion.

                                        Manuscript. URL = ftp://ftp.ee.lbl.
                                        gov/testbase-report.ps.Z, May 1991.

PC-SIG:1987:PSL

[PC-87] PC-SIG, Inc. The PC-SIG Library. PC-SIG, Sunnyvale, California, fourth edition, 1987. ???? pp.

Puschmann:1983:CPR

[PC83] Heinrich Puschmann and Joaquín Cortés. The coordinex problem and its relation to the conjecture of Wilkinson. Numer. Math., 42:291–297, 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pasupathy:1992:GTM

[PD92] J. Pasupathy and R. A. Damodar. The Gaussian toeplitz matrix. Linear Algebra and Appl., 171:133–147, 1992.

Piessens:1983:QSP

[PdDKÜK83] Robert Piessens, Elise de Doncker-Kapenga, Christoph W. Überhuber, and David K. Kahaner. QUADPACK: A Subroutine Package for Automatic Integration. Springer-Verlag, Berlin, 1983. ISBN 3-540-12553-1. viii + 301 pp.

Park:1995:DRR

[PE95] Haesun Park and Lars Eldén. Downdating the rank-revealing URV decomposition. SIAM J. Matrix Anal. Appl., 16 (1):138–155, 1995.

Pechenik:1987:SGW

[Pec87] Jan A. Pechenik. A Short Guide To Writing About Biology. HarperCollins, New York, 1987. ISBN 0-673-39232-5. ???? pp.

Pei:1962:TMI

[Pei62] M. L. Pei. A test matrix for inversion procedures. Comm. ACM, 5(10):508, 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pemberton:1969:HFM

[Pem69] John E. Pemberton. How To Find Out in Mathematics. Pergamon, New York, NY, USA, second edition, 1969. ???? pp.

Postlethwaite:1981:PGP

[PEM81] I. Postlethwaite, J. M. Edmunds, and A. G. J. MacFarlane. Principal gains and principal phases in the analysis of linear multivariable feedback systems. IEEE Trans. Automat. Control., AC-26(1):32– 46, 1981.

Pena:1994:PSL

[Peñ94] J. M. Peña. Pivoting strategies leading to small bounds of the errors for certain linear systems. IMA J. Numer. Anal., page ??, 1994. Submitted.

Pena:1995:MWI

[Peñ95] J. M. Peña. M-matrices whose inverses are totally positive. Linear Algebra and Appl., 221:189–193, 1995.

Perlis:1952:TM

[Per52] S. Perlis. Theory of Matrices. Addison-Wesley, Reading, MA, USA, 1952. ???? pp.

Pereyra:1969:SGS

[Per69] V. Pereyra. Stability of general systems of linear equations. Aequationes Mathematicae, 2:194–206, 1969. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic).

Perry:1991:FAT

[Per91] Carol Rosenblum Perry. The Fine Art of Technical Writing. Blue Heron Publishing, Hillsboro, Oregon, 1991. ISBN 0-936085-24-X. ???? pp.

Petard:1966:BDP

[Pét66] H. Pétard. A brief dictionary of phrases used in mathematical writing. Amer. Math. Monthly, 73:196–197, 1966. Reprinted in [Ewi94].

Petrovsky:1968:PIC

[Pet68] I. G. Petrovsky. Proc. International Congress of Mathematicians, Moscow 1966. Mir Publishers, Moscow, 1968. ???? pp.

Peterson:1980:CPD

[Pet80a] James L. Peterson. Computer programs for detecting and correcting spelling errors. Comm. ACM, 23:676–687, 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peterson:1980:CPS

[Pet80b] James L. Peterson. Computer Programs for Spelling Correction: An Experiment in Program Design. Number 96 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1980. ???? pp.

Petersen:1981:ET

[Pet81] Karl Petersen. Ergodic Theory. Cambridge University Press, Cambridge, UK, 1981. ISBN 0-521-23632-0. xi + 329 pp.

Petzold:1982:DAE

[Pet82] L. Petzold. Differential/algebraic equations are not ODE’s. SIAM J. Sci. Stat. Comput., 3:367–384, 1982.

Peterson:1986:NUT

[Pet86] James L. Peterson. A note on undetected typing errors. Comm. ACM, 29(7): 633–637, 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peterson:1991:SNM

[Pet91] Ivars Peterson. Searching for new mathematics. SIAM Review, 33:37–42, 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Press:1986:NRA

[PFTV86] William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes: The Art of Scientific Computing (FORTRAN Version). Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-38330-7. xx + 702 pp. LCCN QA297 .N866 1989.

Press:1988:NRC

[PFTV88] William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes in C: The Art of Scientific Computing. Cambridge University Press, ????, 1988. ???? pp.

Parkinson:1972:IEV

[PH72] J. M. Parkinson and D. Hutchinson. An investigation into the efficiency of variants on the simplex method. In Numerical Methods for Non-linear Optimization [Loo72], pages 115–135.

Pfeffer:1985:AMP

[PH85] M. Pfeffer and A. Hoenig. Assembling a moderately-priced, high-performance clone of the IBM PC for running TEX. TUGboat, 6:140–145, 1985. ISSN 0896-3207.

Pfeffer:1986:SII

[PH86] M. Pfeffer and A. Hoenig. The sperry IT: An IBM AT compatible. TUGboat, 7: 104–105, 1986. ISSN 0896-3207.

Patterson:1994:COD

[PH94] David A. Patterson and John L. Hennessy. Computer Organization and Design: The Hardware/Software Interface. Morgan Kaufmann, San Mateo, CA, USA, 1994. ISBN 1-55860-282-8. xxiv + 648 + appendices pp.

Phillips:1977:NNA

[Phi77] J. Phillips. Nearest normal approximation for certain operators. Proc. Amer. Math. Soc., 67:236–240, 1977.

Philippe:1987:AIN

[Phi87] Bernard Philippe. An algorithm to improve nearly orthonormal sets of vectors on a vector processor. SIAM J. Alg. Disc. Meth., 8(3):396–403, 1987.

Pichat:1972:CDS

[Pic72] M. Pichat. Correction d’une somme en arithmétique à virgule flottante. Numer. Math., 19:400–406, 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pickover:1990:CPC

[Pic90] C. A. Pickover. Computers, Pattern, Chaos and Beauty: Graphics from an Unseen World. St. Martin’s Press, New York, 1990. ???? pp.

Pickover:1991:MBS

[Pic91] C. A. Pickover. Mathematics and beauty: Several short classroom experiments. Notices Amer. Math. Soc., 38: 192–196, 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pinkus:1985:SEP

[Pin85] A. Pinkus. Some extremal problems for strictly totally positive matrices. Linear Algebra and Appl., 64:141–156, 1985.

Piranian:1982:SIB

[Pir82] George Piranian. Say it better. The Mathematical Intelligencer, 4(1):17–19, 1982. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Pitcher:1990:SES

[Pit90] E. J. Pitcher, editor. Science and Engineering on Supercomputers. Computational Mechanics Publications (Southampton) and Springer-Verlag (Berlin), ????, 1990. ISBN 1-85312-115-0.

Peitgen:1992:FCPa

[PJS92a] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Part One: Introduction to Fractals and Chaos. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97041-X. xiv + 450 pp.

Peitgen:1992:FCPb

[PJS92b] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Part Two: Complex Systems and Mandelbrot Set. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97722-8. xii + 500 pp.

Pandey:1990:PAM

[PKL90] Pradeep Pandey, Charles Kenney, and Alan J. Laub. A parallel algorithm for the matrix sign function. Int Journal of High Speed Computing, 2(2):181–191, 1990. CODEN IHSCEZ. ISSN 0129-0533.

Pandey:1991:GMC

[PKPL91] P. Pandey, C. Kenney, A. Packard, and A. J. Laub. A gradient method for computing the optimal H_{∞} norm. IEEE Trans. Automat. Control., AC-36: 887–890, 1991.

Parlett:1982:MSD

[PL82] B. N. Parlett and T. L. Landis. Methods for scaling to doubly stochastic form. Linear Algebra and Appl., 48:53–79, 1982.

Parlett:1993:FIT

[PL93] Beresford N. Parlett and Jian Le. Forward instability of tridiagonal QR. SIAM J. Matrix Anal. Appl., 14(1):279–316, 1993.

Plauger:1988:PFP

[Pla88] P. J. Plauger. Properties of floatingpoint arithmetic. Computer Language Magazine, 5(3):17–22, 1988. CODEN COMLEF. ISSN 0749-2839.

Plauger:1993:PPE

[Pla93] P. J. Plauger. Programming on Purpose: Essays on Software Design. Prentice-Hall, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-721374-3. viii + 236 pp.

Plemmons:1974:LLS

[Ple74] Robert J. Plemmons. Linear least squares by elimination and MGS. J. Assoc. Comput. Mach., 21(4):581–585, 1974.

Plemmons:1976:RSD

[Ple76] R. J. Plemmons. Regular splittings and the discrete Neumann problem. Numer. Math., 25:153–161, 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plemmons:1986:PBI

[Ple86] R. J. Plemmons. A parallel block iterative scheme applied to computations in structural analysis. SIAM J. Alg. Disc. Meth., 7:337–347, 1986.

Prata:1987:USV

[PM87] S. Prata and D. Martin. UNIX System V Bible. The Waite Group, Howard W. Sams, ????, 1987. ???? pp.

Park:1988:RNG

[PM88] Stephen K. Park and Keith W. Miller.

Random number generators: Good ones are hard to find. Comm. ACM, 31 (10):1192–1201, October 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0001-0782/
63042.html.

Puschmann:1985:ZUV

[PN85] Heinrich Puschmann and Marcelo Nordio. Zwei Unzalässige Verstärkungen der Vermutung von Wilkinson. Linear Algebra and Appl., 72:167–176, 1985. In German.

Parlett:1989:TBB

[PNO89] Beresford N. Parlett and Bahram Nour-Omid. Towards a black box Lanczos program. Computer Physics Communications, 53(1–3):169–179, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL https://www.math.utah.
                                        edu/pub/bibnet/subjects/acc-stab-
                                        num-alg.bib;   http://www.sciencedirect.
                                        com/science/article/pii/0010465589901586.

Polya:1957:HSI

[Pol57] George Polya. How to Solve It: A New Aspect of Mathematical Method. Doubleday, New York, second edition, 1957. xxi + 253 pp.

Pollock:1985:TPM

[Pol85] D. S. G. Pollock. Tensor products and matrix differential calculus. Linear Algebra and Appl., 67:169–193, 1985.

Polman:1987:IBF

[Pol87] Ben Polman. Incomplete blockwise factorizations of (block) H-matrices. Linear Algebra and Appl., 90:119–132, 1987.

Pollicini:1989:UTS

[Pol89] A. A. Pollicini, editor. Using Toolpack Software Tools. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1989. ???? pp.

Potter:1976:OL

[Pot76a] Simeon Potter. Our Language. Penguin, London, 1976. ???? pp.

Potts:1976:SSR

[Pot76b] R. B. Potts. Symmetric square roots of the finite identity matrix. Utilitas Mathematica, 9:73–86, 1976.

Poundstone:1983:BS

[Pou83] W. Poundstone. Big Secrets. Quill, New York, 1983. ???? pp.

Poundstone:1986:BS

[Pou86] W. Poundstone. Bigger Secrets.

Houghton Mifflin, Boston, MA, USA, 1986. ???? pp.

Powell:1967:MEP

[Pow67] M. J. D. Powell. On the maximum errors of polynomial approximations defined by interpolation and by least squares criteria. The Computer Journal, 9:404–407, 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Powell:1969:TRO

[Pow69] M. J. D. Powell. A theorem on rank one modifications to a matrix and its inverse. The Computer Journal, 12:288– 290, 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Powell:1970:SNM

[Pow70] M. J. D. Powell. A survey of numerical methods for unconstrained optimization. SIAM Review, 12(1):79–97, 1970.

Powell:1975:VUM

[Pow75] M. J. D. Powell. A view of unconstrained minimization algorithms that do not require derivatives. ACM Transactions on Mathematical Software, 1(2):97– 107, 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Powell:1976:VUO

[Pow76] M. J. D. Powell. A view of unconstrained optimization. In Optimization in Action [Dix76], pages 117–152.

Powell:1981:ATM

[Pow81] M. J. D. Powell. Approximation Theory and Methods. Cambridge University Press, Cambridge, UK, 1981. ISBN 0-521-29514-9. ???? pp.

Powell:1986:HBB

[Pow86a] M. J. D. Powell. How bad are the BFGS and DFP methods when the objective function is quadratic? Math. Prog., 34:34–47, 1986. CODEN MHPGA4. ISSN 0025-5610.

Power:1986:CDH

[Pow86b] Stephen Power. The Cholesky decomposition in Hilbert space. IMA Bulletin, 22(11/12):186–187, 1986.

Powell:1987:EGB

[Pow87a] M. J. D. Powell. On error growth in the Bartels–Golub and Fletcher– Matthews algorithms for updating matrix factorizations. Linear Algebra and Appl., 88/89:597–622, 1987.

Powell:1987:UCD

[Pow87b] M. J. D. Powell. Updating conjugate directions by the BFGS formula. Math. Prog., 38:29–46, 1987. CODEN MHPGA4. ISSN 0025-5610.

Pan:1989:LSM

[PP89] C.-T. Pan and R. J. Plemmons. Least squares modifications with inverse factorizations: parallel implications. J. Comp. Appl. Math., 27:109–127, 1989.

Pierce:1992:FAC

[PP92] Daniel J. Pierce and Robert J. Plemmons. Fast adaptive condition estimation. SIAM J. Matrix Anal. Appl., 13(1): 274–291, 1992.

Phillips:1994:HGP

[PP94] Estelle M. Phillips and D. S. Pugh.

How to Get a PhD: A Handbook for Students and Their Supervisors. Open University Press, Buckingham, second edition, 1994. ISBN 0-335-19214-9. xiv + 203 pp.

Peyton:1992:PCG

[PPY92] B. W. Peyton, A. Pothen, and X. Yuan. Partitioning a chordal graph into transitive subgraphs for parallel sparse triangular solution. Technical Report CS-92-55, Department of Computer Science, University of Waterloo, Waterloo, Ontario, N2L 3G1, Canada, December 1992. ???? pp. To appear in Linear Algebra and Appl.

Peyton:1993:CTA

[PPY93] B. W. Peyton, A. Pothen, and X. Yuan. A clique tree algorithm for partitioning a chordal graph into transitive subgraphs. Technical Report CS-93-27, Department of Computer Science, University of Waterloo, Waterloo, Ontario, N2L 3G1, Canada, May 1993. ???? pp.

Powell:1968:AHT

[PR68] M. J. D. Powell and J. K. Reid. On applying Householder transformations to linear least squares problems. Technical Report T.P. 322, Mathematics Branch, Theoretical Physics Division, Atomic Energy Research Establishment, Harwell, UK, February 1968. 20 pp.

Parlett:1969:BMC

[PR69a] B. N. Parlett and C. Reinsch. Balancing a matrix for calculation of eigenvalues and eigenvectors. Numer. Math., 13:293–304, 1969. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Powell:1969:AHT

[PR69b] M. J. D. Powell and J. K. Reid. On applying Householder transformations to linear least squares problems. In Proc. IFIP Congress 1968, pages 122– 126. North-Holland, Amsterdam, The Netherlands, 1969.

Parlett:1981:TPL

[PR81] B. N. Parlett and J. K. Reid. Tracking the progress of the Lanczos algorithm for large symmetric eigenproblems. IMA J. Numer. Anal., 1:135–155, 1981.

Pan:1985:EPS

[PR85] V. Pan and J. Reif. Efficient parallel solution of linear systems. Technical report, Proceedings of the Seventeenth Annual ACM Symposium on Theory of Computing, ????, 1985. ???? pp.

Pan:1989:FEP

[PR89] V. Pan and J. Reif. Fast and efficient parallel solution of dense linear systems. Computers and Mathematics with Applications, 17(11):1481–1491, 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Poljak:1993:CRN

[PR93] Svatopluk Poljak and Jiří Rohn. Checking robust nonsingularity is NPhard. Math. Control, Signals and Sys., 6:1–9, 1993.

Prenter:1975:SVM

[Pre75] P. M. Prenter. Splines and Variational Methods. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1975. ???? pp.

Prentice-Hall:1978:PAG

[Pre78] Prentice-Hall, Inc. Prentice-Hall Author’s Guide. Prentice-Hall, Englewood Cliffs, NJ, USA, fifth edition, 1978. ISBN 0-13-695015-9. x + 115 pp. LCCN PN147.P68 1975.

Previn:1991:NMC

[Pre91] Andre Previn. No Minor Chords: My days in Hollywood. Bantam Books, New York, NY, USA, 1991. ISBN 0-553-40525-X. ???? pp.

Prentice:1993:PBS

[Pre93] John K. Prentice. A performance benchmark study of Fortran 90 compilers. Fortran Journal, 5(3):??, 1993. ISSN 1060-0221. May/June.

Priest:1991:AAP

[Pri91a] Douglas M. Priest. Algorithms for arbitrary precision floating point arithmetic. In Peter Kornerup and David W. Matula, editors, Proc. 10th IEEE Symposium on Computer Arithmetic, pages 132–143. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991.

Priestley:1991:PHW

[Pri91b] W. M. Priestley. Paul Halmos:

Words more than numbers. In John H. Ewing and F. W. Gehring, editors, Paul Halmos: Celebrating 50 Years of Mathematics, pages 49–69. Springer-Verlag, Berlin, 1991.

Priest:1992:PFP

[Pri92] Douglas M. Priest. On Properties of Floating Point Arithmetics: Numerical Stability and the Cost of Accurate Computations. PhD thesis, Mathematics Department, University of California, Berkeley, CA, USA, November 1992. 126

                                        pp. URL = ftp://ftp.icsi.berkeley.
                                        edu/pub/theory/priest-thesis.ps.Z.

Pryce:1981:REA

[Pry81] J. D. Pryce. Round-off error analysis with fewer tears. IMA Bulletin, 17:40–47, 1981.

Pryce:1984:NMR

[Pry84] J. D. Pryce. A new measure of relative error for vectors. SIAM J. Numer. Anal., 21(1):202–215, 1984.

Pryce:1985:MEA

[Pry85] J. D. Pryce. Multiplicative error analysis of matrix transformation algorithms. IMA J. Numer. Anal., 5:437–445, 1985.

Pearcy:1979:ACM

[PS79] C. Pearcy and A. Shields. Almost commuting matrices. J. Functional Analysis, 33:332–338, 1979.

Paige:1981:TGS

[PS81] C. C. Paige and M. A. Saunders. Towards a generalized singular value decomposition. SIAM J. Numer. Anal., 18(3): 398–405, 1981.

Pan:1991:INI

[PS91] Victor Pan and Robert Schreiber. An improved Newton iteration for the generalized inverse of a matrix, with applications. SIAM J. Sci. Stat. Comput., 12(5): 1109–1130, 1991.

Pan:1992:FPC

[PS92] Victor Pan and Robert Schreiber. A fast, preconditioned conjugate gradient Toeplitz solver. Computers and Mathematics with Applications, 24(7):17–24, 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Pan:1994:BIP

[PS94] C.-T. Pan and Kermit Sigmon. A bottom-up inductive proof of the singular value decomposition. SIAM J. Matrix Anal. Appl., 15(1):59–61, 1994.

Parlett:1982:ELE

[PSS82] B. N. Parlett, H. Simon, and L. M.

Stringer. On estimating the largest eigenvalue with the Lanczos algorithm. Math. Comp., 38(157):153–165, 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Press:1992:NRC

[PTVF92a] William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical Recipes in C: The Art of Scientific Computing. Cambridge University Press, ????, second edition, 1992. ISBN 0-521-43108-5. ???? pp.

Press:1992:NRF

[PTVF92b] William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical Recipes in FOR-TRAN: The Art of Scientific Computing. Cambridge University Press, ????, second edition, 1992. ISBN 0-521-43064-X. xxvi + 963 pp.

Puglisi:1992:MHM

[Pug92] Chiara Puglisi. Modification of the Householder method based on the compact WY representation. SIAM J. Sci. Stat. Comput., 13(3):723–726, 1992.

Pulay:1966:IMD

[Pul66] P. Pulay. An iterative method for the determination of the square root of a positive definite matrix. Z. Angew. Math. Mech, 46:151, 1966.

Putnam:1957:SRN

[Put57] C. R. Putnam. On square roots of normal operators. Proc. Amer. Math. Soc., 8:768–769, 1957.

Putnam:19xx:SRL

[Putxx] C. R. Putnam. On square roots and logarithms of self-adjoint operators. Technical report, Proc. Glasgow...?, ????, 19xx. ???? pp.

Porte:1974:ESE

[PV74a] M. La Porte and J. Vignes. Etude statistique des erreurs dans l’arithmétique des ordinateurs; application au contrôle des Résultats d’algorithmes numériques. Numer. Math., 23:63–72, 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Porte:1974:MND

[PV74b] M. La Porte and J. Vignes. Méthode numérique de Détection de la singularité d’une matrice. Numer. Math., 23:73–81, 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paige:1981:SDH

[PV81] C. C. Paige and C. F. Van Loan. A Schur decomposition for Hamiltonian matrices. Linear Algebra and Appl., 41:11– 32, 1981.

Paige:1986:QCK

[PV86] C. C. Paige and P. Van Dooren. On the quadratic convergence of Kogbetliantz’s algorithm for computing the singular value decomposition. Linear Algebra and Appl., 77:301–313, 1986.

Pitsianis:1994:VMS

[PV94] Nikos Pitsianis and Charles F. Van Loan. Vectorization of multiple small matrix problems. Technical Report CTC 94 TR172, Advanced Computing Research Institute, Cornell Theory Center, Cornell University, Ithaca, New York, March 1994. ???? pp.

Peters:1969:EBS

[PW69] G. Peters and J. H. Wilkinson. Eigenvalues of Ax = λBx with band symmetric A and B. The Computer Journal, 12: 398–404, 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Peters:1970:LSP

[PW70] G. Peters and J. H. Wilkinson.

The least squares problem and pseudoinverses. The Computer Journal, 13(3): 309–316, 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Peters:1971:PPA

[PW71] G. Peters and J. H. Wilkinson. Practical problems arising in the solution of polynomial equations. J. Inst. Maths Applics, 8:16–35, 1971.

Peters:1974:ACE

[PW74] G. Peters and J. H. Wilkinson. Accuracy of computed eigensystems and invariant subspaces. In B. K. P. Scaife, editor, Studies in Numerical Analysis: Papers in Honour of Cornelius Lanczos, pages 115–135. Academic Press, London, 1974.

Peters:1975:SGJ

[PW75] G. Peters and J. H. Wilkinson. On the stability of Gauss–Jordan elimination with pivoting. Comm. ACM, 18(1):20–24, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parlett:1977:ICC

[PW77] B. N. Parlett and Y. Wang. The influence of the compiler on the cost of mathematical software—in particular on the cost of triangular factorization. ACM Transactions on Mathematical Software, 1:35–46, 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Page:1978:IRM

[PW78] E. S. Page and L. B. Wilson. Information Representation and Manipulation in a Computer. Cambridge University Press, ????, second edition, 1978. ???? pp.

Page:1979:ICC

[PW79a] E. S. Page and L. B. Wilson. An Introduction to Computational Combinatorics. Cambridge University Press, ????, 1979. ???? pp.

Peters:1979:III

[PW79b] G. Peters and J. H. Wilkinson. Inverse iteration, ill-conditioned equations and Newton’s method. SIAM Review, 21: 339–360, 1979.

Perry:1985:DCH

[PW85] Tekla S. Perry and P. Wallich. Design case history: the Commodore 64. IEEE Spectrum, 22(3):48–58, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Paige:1994:HGC

[PW94] C. C. Paige and M. Wei. History and generality of the CS decomposition. Linear Algebra and Appl., 208/209:303–326, 1994.

Pyke:1985:RP

[Pyk85] D. A. Pyke. Referee a paper. In Stephen Lock, editor, How To Do It, pages 215–219. British Medical Association, London, second edition, 1985.

Quarterman:1994:MCC

[QCM94] John S. Quarterman and Smoot Carl-Mitchell. The E-Mail Companion: Communicate Effectively via the Internet and Other Global Networks. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-40658-6. xvii + 318 pp.

Qi:1984:SSE

[Qi84] Liqun Qi. Some simple estimates for singular values of a matrix. Linear Algebra and Appl., 56:105–119, 1984.

Quirk:1990:EU

[QS90] Randolph Quirk and Gabriele Stein.

English In Use. Longman, Harlow, Essex, 1990. ISBN 0-582-06613-1. ???? pp.

Quinn:1990:REL

[QT90] Thomas Quinn and Scott Tremaine.

Roundoff error in long-term planetary orbit integrations. Astron. J., 99(3):1016– 1023, 1990. CODEN ANJOAA. ISSN 0004-6256 (print), 1538-3881 (electronic).

Quinn:1991:TMY

[QTD91] Thomas R. Quinn, Scott Tremaine, and Martin Duncan. A three million year integration of the earth’s orbit. Astron. J., 101(6):2287–2305, 1991. CODEN AN-JOAA. ISSN 0004-6256 (print), 1538-3881 (electronic).

Quirk:1968:UE

[Qui68] Randolph Quirk. The Use of English.

Longman, Harlow, Essex, second edition, 1968. ISBN 0-582-32681-8. ???? pp.

Quinn:1983:EPR

[Qui83] Kevin Quinn. Ever had problems rounding off figures? This stock exchange has. Wall Street Journal, page ??, 1983. 8 November.

Quinn:1987:DEA

[Qui87] M. J. Quinn. Designing Efficient Algorithms for Parallel Computers. Mc-Graw-Hill, New York, 1987. ???? pp.

Quinlan:1994:REL

[Qui94] Gerald D. Quinlan. Round-off error in long-term orbital integrations using multistep methods. Celestial Mechanics and Dynamical Astronomy, 58:339–351, 1994.

Rademacher:1948:AEP

[Rad48] Hans A. Rademacher. On the accumulation of errors in processes of integration on high-speed calculating machines. In Anonymous [Ano48], pages 176–187. ISBN 0-262-08152-0. Reprinted, with a new introduction by William Aspray, Volume 7 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

Radjavi:1962:UEA

[Rad62] H. Radjavi. On unitary equivalence of arbitrary matrices. Trans. Amer. Math. Soc., 104:363–373, 1962.

Raimi:1969:DFS

[Rai69a] R. A. Raimi. On the distribution of first significant figures. Amer. Math. Monthly, 76(4):342–348, 1969.

Raimi:1969:PDF

[Rai69b] Ralph A. Raimi. The peculiar distribution of first digits. Scientific American, 221(6):109–120, 1969. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Raimi:1976:FDP

[Rai76] Ralph A. Raimi. The first digit problem. Amer. Math. Monthly, 83:521–538, 1976.

Rall:1965:EDC

[Ral65] Louis B. Rall, editor. Error In Digital Computation, volume 2. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1965.

Rall:1981:ADT

[Ral81] Louis B. Rall. Automatic Differentiation: Techniques and Applications, volume 120 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1981. viii + 165 pp.

Rall:1991:TMC

[Ral91] Louis B. Rall. Tools for mathematical computation. In Kenneth R. Meyer and Dieter S. Schmidt, editors, Computer Aided Proofs in Analysis, volume 28 of IMA Volumes in Mathematics and its Applications, pages 217–228. Springer-Ver-lag, New York, 1991.

Ramos:1971:REA

[Ram71] George U. Ramos. Roundoff error analysis of the fast Fourier transform. Math. Comp., 25(116):757–768, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Randell:1975:ODCa

[Ran75a] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Springer-Verlag, Berlin, 1975. ISBN 3-540-07114-8.

Randell:1975:ODCb

[Ran75b] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Springer-Verlag, Berlin, third edition, 1975. ISBN 3-540-11319-3. xvi + 580 pp.

RandomHouse:1988:RHC

[Ran88] Random House. The Random House College Dictionary. Random House, New York, NY, USA, revised edition, 1988. ???? pp.

House:1987:RHD

[87] . The Random House Dictionary of the English Language. Random House, New York, NY, USA, second edition, 1987. ???? pp.

Rao:1980:MAR

[Rao80] C. R. Rao. Matrix approximations and reduction of dimensionality in multivariate statistical analysis. In P. R. Krishnaiah, editor, Multivariate Analysis—V, pages 3–22. North-Holland, Amsterdam, The Netherlands, 1980.

Raskin:1986:QS

[Ras86] R. Raskin. The quest for style. IEEE Trans. Prof. Commun., PC-29:10– 18, 1986. Reviews the five style analyzers Electric Webster, Grammatik II, PC-Style, Punctutation and Style, and Rightwriter.

Rath:1982:FGR

[Rat82] Wolfgang Rath. Fast Givens rotations for orthogonal similarity transformations. Numer. Math., 40:47–56, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rathbone:1985:CTI

[Rat85] Robert R. Rathbone. Communicating Technical Information: A New Guide to Current Uses and Abuses in Scientific and Engineering Writing. Addison-Wes-ley, Reading, MA, USA, second edition, 1985. ISBN 0-201-06365-4. ???? pp.

Raymond:1991:NHD

[Ray91] Eric Raymond, editor. The New Hacker’s Dictionary. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-68069-6. ???? pp. Also available as the on-line hacker Jargon File, version 2.9.10, 01 JUL 1992.

Roebuck:1978:STR

[RB78] Pamela A. Roebuck and Stephen Barnett. A survey of Toeplitz and related matrices. Int. J. Systems Sci., 9(8):921– 934, 1978.

Rice:1985:SEP

[RB85] John R. Rice and R. E. Boisvert. Solving Elliptic Problems using ELLPACK. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ???? pp.

Robinson:1989:JSM

[RB89] Donald W. Robinson and Wayne W.

Barrett. The Jordan 1-structure of a matrix of Redheffer. Linear Algebra and Appl., 112:57–73, 1989.

Redhead:1993:MC

[Red93] Brian Redhead. Manchester: A Celebration. Andre Deutsch, London, 1993. ISBN 0-233-98816-5. ???? pp.

Regis:1987:WGE

[Reg87] Ed Regis. Who Got Einstein’s Office?

Penguin, ????, 1987. ???? pp.

Reid:1971:NSG

[Rei71] J. K. Reid. A note on the stability of Gaussian elimination. J. Inst. Maths Applics, 8:374–375, 1971.

Reid:1986:HC

[Rei86] Constance Reid. Hilbert-Courant. Springer-Verlag, Berlin, 1986. ???? pp.

Reid:1987:SM

[Rei87] J. K. Reid. Sparse matrices. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 59–85. Oxford University Press, 1987.

Reichel:1990:NIA

[Rei90] Lothar Reichel. Newton interpolation at Leja points. BIT, 30:332–346, 1990. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Reichel:1991:FQD

[Rei91] Lothar Reichel. Fast QR decomposition of Vandermonde-like matrices and polynomial least squares approximation. SIAM J. Matrix Anal. Appl., 12(3):552– 564, 1991.

Reid:1992:AF

[Rei92] John K. Reid. The advantages of Fortran 90. Computing, 48(3-4):219–238, 1992. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Reid:1993:SBK

[Rei93] Constance Reid. The Search for E. T. Bell, Also Known as John Taine. Mathematical Association of America, Washington, DC, USA, 1993. ISBN 0-88385-508-9. x + 372 pp.

Rejewski:1981:HPM

[Rej81] Marian Rejewski. How Polish mathematicians deciphered the Enigma. Annals of the History of Computing, 3(3): 213–234, 1981.

Rennie:1972:CWT

[Ren72] B. C. Rennie. Chains and whips in the teaching of mathematics. Mathematical Gazette, 56(398):271–274, 1972. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rigal:1967:CGS

[RG67] J. L. Rigal and J. Gaches. On the compatibility of a given solution with the data of a linear system. J. Assoc. Comput. Mach., 14(3):543–548, 1967.

Rothberg:1992:PIH

[RG92a] Edward Rothberg and Anoop Gupta.

Parallel ICCG on a hierarchical memory multiprocessor—Addressing the triangular solve bottleneck. Parallel Computing, 18:719–741, 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ruggiero:1992:PAS

[RG92b] V. Ruggiero and E. Galligani. A parallel algorithm for solving block tridiagonal linear systems. Computers and Mathematics with Applications, 24(4):15–21, 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Rheinboldt:1966:MIC

[Rhe66] W. C. Rheinboldt. On measures of ill-conditioning for nonlinear equations. Math. Comp., 30(133):104–111, 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rheingold:1985:TT

[Rhe85] H. Rheingold. Tools for Thought. Simon & Schuster, New York, 1985. ???? pp.

Richards:1955:AOD

[Ric55] R. K. Richards. Arithmetic Operations in Digital Computers. Van Nostrand, Princeton, NJ, 1955. iv + 397 pp.

Rice:1965:CPR

[Ric65] John R. Rice. On the conditioning of polynomial and rational forms. Numer. Math., 7:426–435, 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rice:1966:EGS

[Ric66a] John R. Rice. Experiments on Gram–Schmidt orthogonalization. Math. Comp., 20:325–328, 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1966:TC

[Ric66b] John R. Rice. A theory of condition. SIAM J. Numer. Anal., 3(2):287– 310, 1966.

Rice:1971:MS

[Ric71] John R. Rice, editor. Mathematical Software. Academic Press, New York, 1971. ???? pp.

Rice:1975:MAQ

[Ric75] John R. Rice. A metalgorithm for adaptive quadrature. J. Assoc. Comput. Mach., 22(1):61–82, 1975.

Rice:1977:MSI

[Ric77] John R. Rice, editor. Mathematical Software III. Academic Press, New York, 1977. ISBN 0-12-587260-7. ix + 388 pp.

Rice:1981:MCM

[Ric81] John R. Rice. Matrix Computations and Mathematical Software. McGraw-Hill, New York, 1981. ???? pp.

Rice:1983:NMS

[Ric83] John R. Rice. Numerical Methods, Software, and Analysis. McGraw-Hill International, Tokyo, 1983. ISBN 0-07-066507-9. ???? pp.

Rice:1988:MASa

[Ric88a] John R. Rice, editor. Mathematical Aspects of Scientific Software. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ???? pp.

Rice:1988:MASb

[Ric88b] John R. Rice. Mathematical aspects of scientific software. In Mathematical Aspects of Scientific Software [Ric88a], pages 1–39.

Rice:1992:NMS

[Ric92] John R. Rice. Numerical Methods, Software, and Analysis. Academic Press, San Diego, second edition, 1992. ISBN 0-12-587755-2. ???? pp.

Riesel:1956:NLL

[Rie56] Hans Riesel. A note on large linear systems. M.T.A.C., 10:226–227, 1956.

Rinehart:1955:EDM

[Rin55] R. F. Rinehart. The equivalence of definitions of a matric function. Amer. Math. Monthly, 62:395–414, 1955.

Rinehart:1956:DMF

[Rin56] R. F. Rinehart. The derivative of a matric function. Proc. Amer. Math. Soc., 7:2–5, 1956.

Rinehart:1960:SMS

[Rin60] R. F. Rinehart. Skew matrices as square roots. Amer. Math. Monthly, 67: 157–161, 1960.

Rinehart:1964:ERU

[Rin64] R. F. Rinehart. The exponential representation of unitary matrices. Amer. Math. Monthly, 37??:111–112, 1964.

Rivlin:1969:IAF

[Riv69] Theodore J. Rivlin. An Introduction to the Approximation of Functions. Blaisdell, New York, 1969. ISBN 0-486-64069-8. viii + 150 pp. Reprinted by Dover, New York, 1981.

Rivlin:1990:CPA

[Riv90] T. J. Rivlin. Chebyshev Polynomials: From Approximation Theory to Algebra and Number Theory. Wiley, New York, NY, USA; London, UK; Sydney, Australia, second edition, 1990. ???? pp.

Reyment:1993:AFA

[RJ93] Richard A. Reyment and K. G. Jöreskog. Applied Factor Analysis in the Natural Sciences. Cambridge University Press, ????, 1993. ISBN 0-521-41242-0. xii + 371 pp.

Ritchie:1978:CPL

[RJLK78] D. M. Ritchie, S. C. Johnson, M. E. Lesk, and B. W. Kernighan. The C programming language. The Bell System Technical Journal, 57:1991–2019, 1978. CODEN BSTJAN. ISSN 0005-8580.

Reiser:1975:EDF

[RK75] John F. Reiser and Donald E. Knuth. Evading the drift in floating-point addition. Information Processing Letters, 3 (3):84–87, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohn:1995:CEC

[RK95] Jiří Rohn and Vladik Kreinovich.

Computing exact componentwise bounds on solutions of linear systems with interval data is NP-hard. SIAM J. Matrix Anal. Appl., 16(2):415–420, 1995.

Richtmyer:1967:DMI

[RM67] Robert D. Richtmyer and K. W. Morton. Difference Methods for Initial-Value Problems. Interscience, New York, second edition, 1967. xiv + 405 pp.

Rao:1971:GIM

[RM71] C. R. Rao and S. K. Mitra. Generalized Inverse of Matrices and its Applications. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1971. ???? pp.

Ramsey:1991:LPT

[RM91] N. Ramsey and C. Marceau. Literate programming on a team project. Software—Practice and Experience, 21:677– 683, 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Romine:1988:PST

[RO88] Charles H. Romine and James M. Ortega. Parallel solution of triangular systems of equations. Parallel Computing, 6(1):109–114, 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Reichel:1991:CVS

[RO91] Lothar Reichel and Gerhard Opfer.

Chebyshev–Vandermonde systems. Math. Comp., 57(196):703–721, 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Roberts:1980:LMR

[Rob80] J. D. Roberts. Linear model reduction and solution of the algebraic Riccati equation by use of the sign function. Internat. J. Control, 32(4):677–687, 1980. First issued as report CUED/B-Control/TR13, Department of Engineering, University of Cambridge, 1971.

Roberts:1988:PEU

[Rob88] Philip Davies Roberts. Plain English: A User’s Guide. Penguin, ????, 1988. ISBN 0-14-008407-X. ???? pp.

Roberts:1990:IVP

[Rob90] Yves Roberts. The Impact of Vector and Parallel Architectures on the Gaussian Elimination Algorithm. Manchester University Press (Manchester) and Halsted Press (New York), ????, 1990. ???? pp.

Roche:1981:SME

[Roc81] J. R. Roche. On the sensitivity of the matrix exponential problem. R.A.I.R.O. Numerical Analysis, 15:249–255, 1981. CODEN RFANDS. ISSN 0399-0516.

Rodale:1978:SF

[Rod78] Jerome Irving Rodale. The Synonym Finder. Rodale Press, Emmaus, Philadelphia, 1978. ISBN 0-87857-236-8. ???? pp. Completely revised by Laurence Urdang and Nancy LaRoche.

Rodin:1992:SPE

[Rod92] Ervin Y. Rodin. Speed of publication—an editorial. Computers and Mathematics with Applications, 24(4):1– 2, 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Roesler:1986:RHE

[Roe86] Friedrich Roesler. Riemann’s hypothesis as an eigenvalue problem. Linear Algebra and Appl., 81:153–198, 1986.

Rogers:1976:PSN

[Rog76] D. D. Rogers. On proximinal sets of normal operators. Proc. Amer. Math. Soc., 61:44–48, 1976.

Rohn:1989:NCN

[Roh89a] Jiří Rohn. New condition numbers for matrices and linear systems. Computing, 41:167–169, 1989. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1989:SLI

[Roh89b] Jiří Rohn. Systems of linear interval equations. Linear Algebra and Appl., 126: 39–78, 1989.

Rohn:1990:NUD

[Roh90] Jiří Rohn. Nonsingularity under data rounding. Linear Algebra and Appl., 139: 171–174, 1990.

Rohn:1992:AFS

[Roh92] J. Rohn. An algorithm for finding a singular matrix in an interval matrix. J. Numer. Lin. Alg. Appl., 1(1):43–47, 1992.

Rohn:1993:ESL

[Roh93] J. Rohn. Enclosing solutions of linear interval equations is NP-hard. 1993.

Rohn:1994:PTL

[Roh94] J. Rohn. A perturbation theorem for linear equations. Commentationes Mathematicae Universitatis Carolinea, 35:213– 214, 1994. CODEN CMUCAA. ISSN 0010-2628 (print), 1213-7243 (electronic).

Rohn:1995:NHR

[Roh95] Jiří Rohn. NP-hardness results for some linear and quadratic problems. Technical Report No. 619, Institute of Computer Science, Academy of Sciences of the Czech Republic, Prague, January 1995. 11 pp.

Rojas:1994:WIC

[Roj94] Raúl Rojas. Who invented the computer? The debate from the viewpoint of computer architecture. In Walter Gautschi, editor, Mathematics of Computation 1943–1993: A Half Century of Computational Mathematics, volume 48 of Proceedings of Symposia in Applied Mathematics, pages 361–365. American Mathematical Society, Providence, RI, USA, 1994.

Roland:1967:MWNa

[Rol67a] Charles G. Roland. Medical writing number 4: Nouns and Adjectives. J. Amer. Medical Assoc., 201(12):961–962, 1967.

Roland:1967:MWNb

[Rol67b] Charles G. Roland. Medical writing number 6: Rewriting. J. Amer. Medical Assoc., 202(3):219–220, 1967.

Roland:1975:TAMc

[Rol75a] Charles G. Roland. Thoughts about medical writing XXIX. Authorship— who, and where? Anesthesia and Analgesia … Current Researches, 54(3):366– 367, 1975. CODEN AACRAT. ISSN 0003-2999.

Roland:1975:TAMa

[Rol75b] Charles G. Roland. Thoughts about medical writing XXVII. Adam’s rib and the anatomy of medical prose. Anesthesia and Analgesia … Current Researches, 54(1):139–141, 1975. CODEN AACRAT. ISSN 0003-2999.

Roland:1975:TAMb

[Rol75c] Charles G. Roland. Thoughts about medical writing XXVIII. What hue are you? Anesthesia and Analgesia … Current Researches, 54(2):230– 231, 1975. CODEN AACRAT. ISSN 0003-2999.

Roland:1975:TAMd

[Rol75d] Charles G. Roland. Thoughts about medical writing XXX. Crooked thinking. Anesthesia and Analgesia … Current Researches, 54(4):540–541, 1975. CODEN AACRAT. ISSN 0003-2999.

Roland:1975:TAMe

[Rol75e] Charles G. Roland. Thoughts about medical writing XXXI. The letter-to-the-editor. Anesthesia and Analgesia … Current Researches, 54(5):626–627, 1975. CODEN AACRAT. ISSN 0003-2999.

Roland:1975:TAMf

[Rol75f] Charles G. Roland. Thoughts about medical writing XXXII. So you want to publish a book. Anesthesia and Analgesia … Current Researches, 54(6):814– 815, 1975. CODEN AACRAT. ISSN 0003-2999.

Roland:1976:TAMa

[Rol76a] Charles G. Roland. Thoughts about medical writing XXXIII. Parallelism is a virtue. Anesthesia and Analgesia … Current Researches, 55(1):139–140, 1976. CODEN AACRAT. ISSN 0003-

Roland:1976:TAMb

[Rol76b] Charles G. Roland. Thoughts about medical writing XXXIV. The publication process. Anesthesia and Analgesia … Current Researches, 55(2):261– 262, 1976. CODEN AACRAT. ISSN 0003-2999.

Roland:1976:TAMc

[Rol76c] Charles G. Roland. Thoughts about medical writing XXXV. “Let’s call it hebetization. Anesthesia and Analgesia … Current Researches, 55(3):366– 367, 1976. CODEN AACRAT. ISSN 0003-2999.

Roland:1976:TAMd

[Rol76d] Charles G. Roland. Thoughts about medical writing XXXVI. Writing an editorial. Anesthesia and Analgesia … Current Researches, 55(4):534– 535, 1976. CODEN AACRAT. ISSN 0003-2999.

Roland:1976:TAMe

[Rol76e] Charles G. Roland. Thoughts about medical writing XXXVII. Verify your references. Anesthesia and Analgesia … Current Researches, 55(5):717–718, 1976. CODEN AACRAT. ISSN 0003-2999.

Ross:1965:RTE

[Ros65] D. R. Ross. Reducing truncation errors using cascading accumulators. Comm. ACM, 8(1):32–33, 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenthal:1969:ACM

[Ros69] P. Rosenthal. Are almost commuting matrices near commuting matrices? Amer. Math. Monthly, 76:925–926, 1969.

Rosen:1972:PSL

[Ros72] S. Rosen. Programming systems and languages 1965–1975. Comm. ACM, 15: 591–600, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roskies:1975:PP

[Ros75] Ethel Roskies. Publish and perish. American Psychologist, 30:1165– 1168, 1975. CODEN AMPSAB. ISSN 0003-066x (print), 1935-990x (electronic).

Anonymous:1990:RCI

[Ros90] Rosemary Conley’s Inch Loss Plan: A Complete Diet and Exercise Program that will Transform Your Shape in 28 Days. Arrow Books, London, 1990. ISBN 0-09-974120-2. 192 pp.

Rothblum:1984:ESO

[Rot84] U. G. Rothblum. Explicit solutions for optimization problems on the intersections of the unit ball of the l_{1} and l_{i}nfty norms with a hyperplane. SIAM J. Alg. Disc. Meth., 5:619–632, 1984.

Rowan:1990:FSA

[Row90] Thomas Harvey Rowan. Functional Stability Analysis of Numerical Algorithms. PhD thesis, University of Texas at Austin, Austin, TX, USA, May 1990. xii + 205 pp.

Rozsa:1987:IBM

[Róz87] P. Rózsa. On the inverse of band matrices. Integral Equations and Operator Theory, 10:82–95, 1987. ISSN 0378-620X (print), 1420-8989 (electronic).

Rozsa:1988:BMS

[Róz88] P. Rózsa. Band matrices and semiseparable matrices. In Numerical Methods [GR88], pages 229–237.

Robers:1973:ADL

[RR73] P. D. Robers and S. S. Robers. Algorithm 458: Discrete linear L_{1} approximation by interval linear programming. Comm. ACM, 16:629–631, 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ralston:1978:FCN

[RR78] A. Ralston and P. Rabinowitz. A First Course In Numerical Analysis. McGraw-Hill, New York, NY, USA, second edition, 1978. ???? pp.

Rack:1982:NSE

[RR82] Heinz-Joachim Rack and Manfred Reimer. The numerical stability of evaluation schemes for polynomials based on the Lagrange interpolation form. BIT, 22: 101–107, 1982. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Robinson:1981:ASC

[RS81] P. Robinson and D. Singer. Another spelling correction program. Comm. ACM, 24:296–297, 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rader:1986:HHT

[RS86] C. M. Rader and A. O. Steinhardt. Hyperbolic Householder transformations. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-34:1589–1602, 1986. CODEN IETABA. ISSN 0096-3518.

Rialan:1988:FAC

[RS88a] C. P. Rialan and L. L. Scharf. Fast algorithms for computing QR and Cholesky factors of Toeplitz operators. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-36:1740–1748, 1988. CO-DEN IETABA. ISSN 0096-3518.

Robertazzi:1988:BOF

[RS88b] T. G. Robertazzi and S. C. Schwartz.

Best “ordering” for floating-point addition. ACM Transactions on Mathematical Software, 14(1):101–110, 1988. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Routh:1977:PSM

[RSD77] M. W. Routh, P. A. Swartz, and M. B. Denton. Performance of the super modified simplex. Analytical Chemistry, 49(9):1422–1428, 1977. CODEN AN-CHAM. ISSN 0003-2700 (print), 1520-6882 (electronic).

Reddy:1992:SML

[RT92a] Satish C. Reddy and Lloyd N. Trefethen. Stability of the method of lines. Numer. Math., 62:235–267, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reichel:1992:EPE

[RT92b] Lothar Reichel and Lloyd N. Trefethen. Eigenvalues and pseudoeigenvalues of Toeplitz matrices. Linear Algebra and Appl., 162–164:153–185, 1992.

Reddy:1994:PCD

[RT94] Satish C. Reddy and Lloyd N. Trefethen. Pseudospectra of the convectiondiffusion operator. SIAM J. Appl. Math., 54(6):1634–1649, 1994.

Reddy:1993:PCD

[RTP93] Satish C. Reddy, Lloyd N. Trefethen, and D. Pathria. Pseudospectrum of the convection–diffusion operator. Technical Report CTC 93 TR126, Advanced Computing Research Institute, Cornell Theory Center, Cornell University, Ithaca, New York, 1993. ???? pp.

Rubinstein:1988:DTI

[Rub88] Richard Rubinstein. Digital Typography: An Introduction to Type and Composition for Computer System Design. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17633-5. ???? pp.

Ruhe:1968:QCG

[Ruh68] A. Ruhe. On the quadratic convergence of a generalization of the Jacobi method to arbitrary matrices. BIT, 8: 210–231, 1968. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1969:NMA

[Ruh69] A. Ruhe. The norm of a matrix after a similarity transformation. BIT, 9:53– 58, 1969. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1970:AND

[Ruh70] A. Ruhe. An algorithm for numerical determination of the structure of a general matrix. BIT, 10:196–216, 1970. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1973:ANE

[Ruh73] A. Ruhe. Algorithms for the nonlinear eigenvalue problem. SIAM J. Numer. Anal., 10:674–689, 1973.

Ruhe:1975:CES

[Ruh75] A. Ruhe. On the closeness of eigenvalues and singular values of almost normal matrices. Linear Algebra and Appl., 11:87–94, 1975.

Ruhe:1983:NAG

[Ruh83] Axel Ruhe. Numerical aspects of Gram–Schmidt orthogonalization of vectors. Linear Algebra and Appl., 52/53: 591–601, 1983.

Ruhe:1987:CNM

[Ruh87] Axel Ruhe. Closest normal matrix finally found! BIT, 27:585–598, 1987. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ruhe:1994:RKA

[Ruh94] Axel Ruhe. The rational Krylov algorithm for large nonsymmetric eigenvalues— Mapping the resolvent norms (pseudospectrum). Manuscript, July 1994.

Rump:1991:CAI

[Rum91] S. M. Rump. A class of arbitrarily ill conditioned floating-point matrices. SIAM J. Matrix Anal. Appl., 12:645–653, 1991.

Russell:1972:BL

[Rus72] Ross Russell. Bird Lives! Quartet Books, London, 1972. ???? pp.

Rutherford:1947:SCD

[Rut47] D. E. Rutherford. Some continuant determinants arising in physics and chemistry. Proc. Royal Soc. Edin., 62A:229– 236, 1947.

Rutherford:1952:SCD

[Rut52] D. E. Rutherford. Some continuant determinants arising in physics and chemistry—II. Proc. Royal Soc. Edin., 63A:232–241, 1952.

Rutishauser:1958:SEP

[Rut58] Heinz Rutishauser. Solution of eigenvalue problems with the LRtransformation. In Further Contributions to the Solution of Simultaneous Linear Equations and the Determination of Eigenvalues, number 49 in Applied Mathematics Series, pages 47–81. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1958.

Rutishauser:1960:KKV

[Rut60] H. Rutishauser. Über eine kubisch konvergente variante der LRtransformation. Z. Angew. Math. Mech, 40:49–54, 1960.

Rutishauser:1966:BEO

[Rut66] H. Rutishauser. Bestimmung der eigenwerte orthogonaler matrizen. Numer. Math., 9:104–108, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1968:TM

[Rut68] H. Rutishauser. On test matrices. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7 of Éditions Centre Nat. Recherche Sci., Paris, pages 349–365, 1968.

Rutishauer:1990:LNM

[Rut90] H. Rutishauer. Lectures on Numerical Mathematics. Birkhäuser Verlag, Basel, Switzerland, 1990. ???? pp. Edited by M. Gutknecht, with the assistance of P. Henrici, P. Läuchli, H.-R. Schwarz; translated by W. Gautschi.

Ralston:1967:MMD

[RW67] A. Ralston and H. S. Wilf, editors. Mathematical Methods for Digital Computers, volume II. Wiley, New York, NY, USA; London, UK; Sydney, Australia,

Rudolfer:1992:TE

[RW92] Stephan M. Rudolfer and Peter C.

Watson. Table errata. Math. Comp., 59(206):727, 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ryder:1974:PV

[Ryd74] B. G. Ryder. The PFORT verifier. Software—Practice and Experience, 4:359–377, 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stegun:1956:PC

[SA56] Irene A. Stegun and Milton Abramowitz.

Pitfalls in computation. J. Soc. Indust. Appl. Math., 4(4):207–219, 1956.

Shampine:1973:NCI

[SA73] Lawrence F. Shampine and Richard C.

Allen, Jr. Numerical Computing: An Introduction. W. B. Saunders, Philadelphia, PA, USA, 1973. ISBN 0-7216-8150-6. vii + 258 pp.

Swartzlander:1975:SLN

[SA75] Earl E. Swartzlander, Jr. and Aristides G. Alexopolous. The sign/logarithm number system. IEEE Trans. Comput., C-24(12):1238–1242, 1975.

Saad:1980:VAM

[Saa80] Y. Saad. Variations on Arnoldi’s method for computing eigenelements of large unsymmetric matrices. Linear Algebra and Appl., 34:269–295, 1980.

Saad:1984:CAT

[Saa84] Youcef Saad. Chebyshev acceleration techniques for solving nonsymmetric eigenvalue problems. Math. Comp., 42(166):567–588, 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1992:NML

[Saa92] Youcef Saad. Numerical Methods for Large Eigenvalue Problems. Manchester University Press (Manchester) and Halsted Press (New York), ????, 1992. ISBN 0-7190-3386-1 (UK), 0-470-21820-7 (US). ???? pp.

Safire:1990:FLG

[Saf90] William Safire. Fumblerules: A Lighthearted Guide to Grammar and Good Usage. Dell Publishing, New York, 1990. ISBN 0-440-21010-0. ???? pp.

Salzer:1960:NSQ

[Sal60] Herbert E. Salzer. A note on the solution of quartic equations. Math. Comp., 14:279–281, 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1981:RHN

[Sal81] Herbert E. Salzer. Review of “A history of numerical analysis from the 16th through the 19th century” by herman H. goldstine. Annals of the History of Computing, 3(3):289–294, 1981.

Salton:1989:ATP

[Sal89] G. Salton. Automatic Text Processing: The Transformation, Analysis, and Retrieval of Information by Computer. Addison-Wesley, Reading, MA, USA, 1989. ???? pp.

Salford:1993:FRM

[Sal93] Salford Software Ltd. and the Numerical Algorithms Group Ltd., Salford and Oxford, UK. FTN90 Reference Manual, second edition, 1993. ???? pp.

Samuel:1983:FGT

[Sam83] A. L. Samuel. First grade TEX: A beginner’s TEX manual. Technical report, Report No. STAN-CS-83-985, Department of Computer Science, Stanford University, ????, 1983. ???? pp. (available from TUG as reprint).

Sassoon:1993:CT

[Sas93] Rosemary Sassoon, editor. Computers and Typography. Intellect Books, London, 1993. ISBN 1-871516-23-4. ???? pp.

Saunders:1972:LSL

[Sau72] M. A. Saunders. Large-scale linear programming using the Cholesky factorization. Report CS 252, Department of Computer Science, Stanford University, ????, January 1972. ???? pp.

Sautter:1978:EAG

[Sau78] Werner Sautter. Error analysis of Gauss elimination for the best least squares problem. Numer. Math., 30: 165–184, 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saunders:1994:SSR

[Sau94] Michael A. Saunders. Solution of sparse rectangular systems using LSQR and CRAIG. Technical Report SOL 94-4, Department of Operations Research, Stanford University, ????, December 1994. ???? pp.

Savory:1967:LS

[Sav67] Theodore H. Savory. The Language of Science. Andre Deutsch, London, 1967. ???? pp.

Sawyer:1943:MD

[Saw43] W. W. Sawyer. Mathematician’s Delight. Penguin, ????, 1943. ???? pp.

Sawyer:1955:PM

[Saw55] W. W. Sawyer. Prelude to Mathematics. Penguin, ????, 1955. 214 pp.

Samelson:1953:ORR

[SB53] Klaus Samelson and Friedrich L. Bauer. Optimale Rechengenauigkeit bei Rechenanlagen mit gleitendem Komma. Z. Angew. Math. Phys, 4:312–316, 1953. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Sameh:1977:STS

[SB77] Ahmed H. Sameh and Richard P. Brent. Solving triangular systems on a parallel computer. SIAM J. Numer. Anal., 14(6):1101–1113, 1977.

Serbin:1980:ACM

[SB80a] S. M. Serbin and S. A. Blalock. An algorithm for computing the matrix cosine. SIAM J. Sci. Stat. Comput., 1:198–204, 1980.

Stoer:1980:INA

[SB80b] J. Stoer and R. Bulirsch. Introduction to Numerical Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-90420-4. ix + 609 pp.

Sidi:1988:CSA

[SB88] A. Sidi and J. Bridger. Convergence and stability analyses for some vector extrapolation methods in the presence of defective iteration matrices. J. Comp. Appl. Math., 22:35–61, 1988.

Shroff:1992:ACE

[SB92] Gautam M. Shroff and Christian H.

Bischof. Adaptive condition estimation for rank-one updates of QR factorizations. SIAM J. Matrix Anal. Appl., 13 (4):1264–1278, 1992.

Sharangpani:1994:SAF

[SB94] H. P. Sharangpani and M. L. Barton.

Statistical analysis of floating point flaw in the Pentium processor (1994). Technical report, Intel Corporation, Santa Clara, CA, USA, November 1994. 31 pp.

Sun:1995:BKR

[SB95] Xiaobai Sun and Christian H. Bischof.

A basis-kernel representation of orthogonal matrices. SIAM J. Matrix Anal. Appl., 16(4):1184–1196, 1995.

Smith:1976:MER

[SBD^{+}76] B. T. Smith, J. M. Boyle, J. J.

Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines—EISPACK Guide, volume 6 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1976. ???? pp.

Smith:1993:MSE

[SBKZ93] Suzanne Weaver Smith, John R.

Baker, Mohamed Kaouk, and David C. Zimmerman. Mode shape expansion for visualization and model correlation. In L. Meirovitch, editor, Dynamics and Control of Large Structures: Proceedings of the Ninth VPI&SU Symposium held in Blacksburg, Virginia, pages 385–396. ????, 1993.

Schoenemann:1970:FOM

[SC70] P. H. Schönemann and R. M. Carroll. Fitting one matrix to another under choice of a central dilation and a rigid motion. Psychometrika, 35(2):245–255, 1970. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Shaw:1990:SNR

[SC90] G. J. Shaw and P. I. Crumpton. Some novel results on Newton’s method for matrix inversion. Technical report, Report No. 90/21, Oxford University Computing Laboratory, ????, 1990. ???? pp.

Scarborough:1950:NMA

[Sca50] James B. Scarborough. Numerical Mathematical Analysis. The Johns Hopkins University Press, Baltimore, MD, USA, second edition, 1950. xviii + 511 pp.

Schulz:1933:IBR

[Sch33] Günther Schulz. Iterative Berechnung der reziproken Matrix. Z. Angew. Math. Mech, 13:57–59, 1933.

Schaeffer:1965:MAI

[Sch65] J. J. Schäffer. More about invertible operators without roots. Proc. Amer. Math. Soc., pages 213–219, April 1965.

Schoenemann:1966:GSO

[Sch66] Peter H. Schönemann. A generalized solution of the orthogonal Procrustes problem. Psychometrika, 31(1):1– 10, 1966. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Scheid:1968:NA

[Sch68a] F. Scheid. Numerical Analysis. Schaum’s Outline Series. McGraw-Hill, New York, NY, USA, 1968. ???? pp.

Schoenemann:1968:TSO

[Sch68b] P. H. Schönemann. On two-sided orthogonal Procrustes problems. Psychometrika, 33:19–33, 1968. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Schoenhage:1973:FSO

[Sch73] Arnold Schönhage. Fast schmidt orthogonalization and unitary transformations of large matrices. In Traub [Tra73], pages 283–291. ISBN 0-12-697550-7. LCCN QA76.6 .S918 1973.

Schneider:1974:TCI

[Sch74] Ben Ross Schneider, Jr. Travels in Computerland, or Incompatibilities and Interfaces. Addison-Wesley, Reading, MA, USA, 1974. ???? pp.

Schneider:1977:OTT

[Sch77] Hans Schneider. Olga Taussky-Todd’s influence on matrix theory and matrix theorists: A discursive personal tribute. Linear and Multilinear Algebra, 5:197– 224, 1977. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schaefer:1979:UML

[Sch79] Barbara Kirsch Schaefer. Using the Mathematical Literature: A Practical Guide. Marcel Dekker, Inc., New York, NY, USA, 1979. ISBN 0-8247-6675-X. ix + 141 pp.

Schryer:1981:TCF

[Sch81a] N. L. Schryer. A test of a computer’s floating-point arithmetic unit. Computing Science Technical Report No. 89, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981. ???? pp.

Schultz:1981:EPS

[Sch81b] Martin H. Schultz, editor. Elliptic Problem Solvers. Academic Press, New York, 1981. ISBN 0-12-632620-7.

Schwarz:1981:EYC

[Sch81c] H. R. Schwarz. The early years of computing in Switzerland. Annals of the History of Computing, 3(2):121–132, 1981.

Schulz:1982:SC

[Sch82] Charles M. Schulz. Snoopy Classics. W H Smith (Ravette Books), Horsham, West Sussex, England, 1982. ISBN 1-85304-032-0. ???? pp.

Schelin:1983:CFA

[Sch83a] Charles W. Schelin. Calculator function approximation. Amer. Math. Monthly, 90(5):317–325, 1983.

Schnabel:1983:QNM

[Sch83b] R. B. Schnabel. Quasi-Newton methods using multiple secant updates. Technical report, Report CU-CS-247-83, Department of Computer Science, University of Colorado, Boulder, ????, 1983. ???? pp.

Schreiber:1983:CGI

[Sch83c] R. Schreiber. Computing generalized inverses and eigenvalues of symmetric matrices using systolic arrays. Technical report, Proceedings of the Sixth International Conference on Computer Methods in Science and Engineering, ????, 1983. ???? pp.

Schryer:1984:DCF

[Sch84] N. L. Schryer. Determination of correct floating-point model parameters. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 360–366. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Scherr:1986:JIR

[Sch86a] George H. Scherr, editor. The Journal of Irreproducible Results: Selected Papers. Dorset Press, New York, NY, third edition, 1986. ISBN 0-930376-42-0. 308 pp. LCCN Q167.J682 1986.

Schmidt:1986:MEL

[Sch86b] R. O. Schmidt. Multiple emitter location and signal parameter estimation. IEEE Trans. Antennas Propagation, AP-34(3):276–280, 1986.

Schreiber:1986:SES

[Sch86c] Robert S. Schreiber. Solving eigenvalue and singular value problems on an undersized systolic array. SIAM J. Sci. Stat. Comput., 7(2):441–451, 1986.

Schwartz:1986:HPM

[Sch86d] David Louis Schwartz. How to be a published mathematician without trying harder than necessary. In Scherr [Sch86a], pages 205–?? ISBN 0-930376-42-0. LCCN Q167.J682 1986.

Schur:1987:DCW

[Sch87] N. W. Schur. A Dictionary of Challenging Words. Penguin, ????, 1987. ???? pp.

Schreiber:1988:BAP

[Sch88a] Robert S. Schreiber. Block algorithms for parallel machines. In M. H. Schultz, editor, Numerical Algorithms for Modern Parallel Computer Architectures, number 13 in IMA Volumes In Mathematics and Its Applications, pages 197– 207. Springer-Verlag, Berlin, 1988.

Schultz:1988:NAM

[Sch88b] M. H. Schultz, editor. Numerical Algorithms for Modern Parallel Computer Architectures, number 13 in IMA Volumes In Mathematics and Its Applications. Springer-Verlag, Berlin, 1988.

Schreiber:1989:SLA

[Sch89a] R. Schreiber. Systolic linear algebra machines: A survey. In Haykin [Hay89], pages 73–92.

Schreiber:1989:HRS

[Sch89b] Robert S. Schreiber. Hough’s random story explained. NA Digest, 89(3): ??, 1989. Electronic mail magazine: na.

help@na-net.ornl.gov.

Schreiber:1990:ACM

[Sch90a] R. S. Schreiber. An assessment of the Connection Machine. RIACS Technical Report TR 90.40, NASA Ames Research Center, Moffett Field, CA, USA, June 1990. ???? pp.

Schreiber:1990:SSD

[Sch90b] R. S. Schreiber. Scalability of sparse direct solvers. RIACS Technical Report TR 92.13, NASA Ames Research Center, Moffett Field, CA, USA, May 1990. ???? pp. To appear in Sparse Matrix Computations: Graph Theory Issues and Algorithms, J. A. George , J. R. Gilbert and J. W. H. Liu, eds, IMA Volumes in Mathematics and its Applications, Springer-Verlag.

Schultz:1990:TWT

[Sch90c] Dodi Schultz, editor. Tools of the Writer’s Trade: Writers Tell All About the Equipment and Services They Find The Best. HarperCollins, New York, 1990. ISBN 0-06-272003-1. ???? pp.

Schmitt:1991:AAM

[Sch91] Bernhard A. Schmitt. An algebraic approximation for the matrix exponential in singularly perturbed boundary value problems. SIAM J. Numer. Anal., 27:51– 66, 1991.

Schmitt:1992:PBM

[Sch92a] Bernhard A. Schmitt. Perturbation bounds for matrix square roots and Pythagorean sums. Linear Algebra and Appl., 174:215–227, 1992.

Schofield:1992:OMB

[Sch92b] Jack Schofield. The OED on a metallic beer mat. The Guardian, page ??, 1992. ISSN 0261-3077 (print), 1756-3224 (electronic). 25 June.

Schlick:1993:MCF

[Sch93] Tamar Schlick. Modified Cholesky factorizations for sparse preconditioners. SIAM J. Sci. Comput., 14(2):424–445, 1993.

Scott:1981:SSS

[Sco81] D. S. Scott. Solving sparse symmetric generalized eigenvalue problems without factorization. SIAM J. Numer. Anal., 18 (1):102–110, 1981.

Sweetgall:1986:WJE

[SD86] Robert Sweetgall and John Dignam.

The Walker’s Journal: Experiencing America on Foot. Creative Walking, Inc., 175 Elkton Road, Newark, DE 19711, 1986. ISBN 0-939041-02-2. ???? pp.

Schreiber:1990:ABN

[SD90] R. S. Schreiber and J. J. Dongarra. Automatic blocking of nested loops. Technical report, Report CS-90-108, Department of Computer Science, University of Tennessee, Knoxville, ????, 1990. ???? pp.

Sledd:1962:DDC

[SE62] James Sledd and Wilma R. Ebbit. Dictionaries and That Dictionary: A Casebook on the Aims of Lexicographers and the Targets of Reviewers. Scott, Foresman and Company, Glenview, IL, USA, 1962. ???? pp.

Strang:1987:TCE

[SE87] G. Strang and A. Edelman. The Toeplitz-circulant eigenvalue problem Ax = λCx. Technical report, in Oakland Conference on PDEs, L. Bragg and J. Dettman, eds, Longman, ????, 1987. ???? pp.

Schnabel:1990:NMC

[SE90] Robert B. Schnabel and Elizabeth Eskow. A new modified Cholesky factorization. SIAM J. Sci. Stat. Comput., 11(6): 1136–1158, 1990.

Seager:1989:SM

[Sea89] M. K. Seager. A SLAP for the masses. In Parallel Supercomputing: Methods, Algorithms and Applications [Car89], pages 135–155.

Seber:1977:LRA

[Seb77] G. A. F. Seber. Linear Regression Analysis. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1977. ISBN 0-471-01967-4. xvii + 465 pp.

Sedgewick:1978:IQP

[Sed78] R. Sedgewick. Implementing Quicksort programs. Comm. ACM, 21:847–857, 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sedgewick:1988:A

[Sed88] Robert Sedgewick. Algorithms. Addison-Wesley, Reading, MA, USA, second edition, 1988. ISBN 0-201-06673-4. xii + 657 pp. LCCN QA76.6 .S435 1988. US$34.95.

Sedgewick:1990:AC

[Sed90] Robert Sedgewick. Algorithms in C. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51425-7. xii + 657 pp.

Schmidt:1986:MSD

[SF86] R. O. Schmidt and R. E. Franks. Multiple source DF signal processing: An experimental system. IEEE Trans. Antennas Propagation, AP-34(3):281–290, 1986.

Stewart:1987:ISL

[SF87] Walter W. Stewart and Ned Feder.

The integrity of the scientific literature. Nature, 325:207–214, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Sarason:1983:PRH

[SG83a] Donald E. Sarason and Leonard Gillman, editors. P. R. Halmos. Selecta: Expository Writing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-90756-4. ???? pp.

Sarason:1983:PHSb

[SG83b] Donald E. Sarason and Leonard Gillman, editors. Paul Halmos. Selecta: Expository Writing. Springer-Verlag, Berlin, 1983. ISBN 3-540-90756-4. xix + 304 pp.

Sarason:1983:PHSa

[SG83c] Donald E. Sarason and Leonard Gillman, editors. Paul Halmos. Selecta: Research Contributions. Springer-Verlag, Berlin, 1983. ???? pp.

Shortland:1991:CSH

[SG91] Michael Shortland and Jane Gregory. Communicating Science: A Handbook. Longman, Harlow, Essex, UK, 1991. ISBN 0-582-05709-4. ???? pp.

Simoncini:1992:IMN

[SG92] V. Simoncini and E. Gallopoulos. An iterative method for nonsymmetric systems with multiple right-hand sides. Report 1242, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, July 1992. ???? pp.

Shampine:1991:RSS

[SGB91] L. F. Shampine, I. Gladwell, and R. W. Brankin. Reliable solution of special event location problems for ODEs. ACM Transactions on Mathematical Software, 17:11–25, 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Skillin:1974:WT

[SGo74] Marjorie E. Skillin, Robert M. Gay, and other authorities. Words Into Type. Prentice-Hall, Englewood Cliffs, NJ, USA, third edition, 1974. ???? pp.

Singer:1983:ARE

[SH83] M. A. Singer and S. J. Hammarling.

The algebraic Riccati equation: A summary review of some available results. NPL Report DITC 23/83, National Physical Laboratory, Teddington, UK, January 1983. ???? pp.

Suffridge:1993:AHP

[SH93] T. J. Suffridge and T. L. Hayden.

Approximation by a Hermitian positive semidefinite Toeplitz matrix. SIAM J. Matrix Anal. Appl., 14(3):721–734, 1993.

Shamir:1979:HSS

[Sha79] A. Shamir. How to share a secret.

Comm. ACM, 22:612–613, 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shampine:1982:CMA

[Sha82a] L. F. Shampine. Conditioning of matrices arising in the solution of stiff ODE’s. Technical report, SAND 82-0906, Sandia National Laboratories, Albuquerque, NM, ????, 1982. ???? pp.

Shampine:1982:IRM

[Sha82b] L. F. Shampine. Implementation of Rosenbrock methods. ACM Transactions on Mathematical Software, 8:93– 113, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Shapiro:1982:OSM

[Sha82c] Alexander Shapiro. Optimally scaled matrices, necessary and sufficient conditions. Numer. Math., 39:239–245, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shapiro:1985:OBD

[Sha85] Alexander Shapiro. Optimal block diagonal l_{2}-scaling of matrices. SIAM J. Numer. Anal., 22(1):81–94, 1985.

Shampine:1987:RNR

[Sha87] L. F. Shampine. Review of “numerical recipes: The art of scientific computing”. Amer. Math. Monthly, 94:889–892, 1987.

Shapiro:1991:UBN

[Sha91] Alexander Shapiro. Upper bounds for nearly optimal diagonal scaling of matrices. Linear and Multilinear Algebra, 29: 147, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Shampine:1994:NSO

[Sha94] Lawrence F. Shampine. Numerical Solution of Ordinary Differential Equations. Chapman and Hall, New York, 1994. ISBN 0-412-05151-6. x + 484 pp.

Shea:1993:PIP

[She93] Michael Shea. Personal Impact: Presence, Paralanguage and the Art of Good Communication. Sinclair-Stevenson, London, 1993. ISBN 1-85619-257-1. ???? pp.

Shewchuk:1994:ICG

[She94] Jonathan Richard Shewchuk. An introduction to the conjugate gradient method without the agonizing pain. Report CMU-CS-94-125, School of Computer Science, Carnegie Mellon University, Pittsburgh, March 1994. ???? pp.

Shisha:1967:I

[Shi67] I. Shisha. Inequalities. Academic Press, New York, 1967. ???? pp.

Shisha:1970:II

[Shi70] I. Shisha. Inequalities—II. Academic Press, New York, 1970. ???? pp.

Shisha:1972:II

[Shi72] I. Shisha. Inequalities—III. Academic Press, New York, 1972. ???? pp.

Shrager:1984:OSC

[Shr84] R. I. Shrager. Optical spectra from chemical titration: An analysis by SVD. SIAM J. Alg. Disc. Meth., 5(3):351–358, 1984.

Shrager:1985:RRR

[Shr85] R. I. Shrager. A rapid robust rootfinder. Math. Comp., 44:151–165, 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steenrod:1973:HWMb

[SHSD73] Norman E. Steenrod, Paul R. Halmos, Menahem M. Schiffer, and Jean A. Dieudonné, editors. How to Write Mathematics. American Mathematical Society, Providence, RI, USA, 1973. 64 pp. LCCN QA41 .H6.

Scott:1986:PBJ

[SHW86] David S. Scott, Michael T. Heath, and Robert C. Ward. Parallel block Jacobi eigenvalue algorithms using systolic arrays. Linear Algebra and Appl., 77:345– 355, 1986.

Siebenmann:1986:TWU

[Sie86] L. Siebenmann. Toward wider use of TEX typesetting: A pre-TEX manifesto. Notices Amer. Math. Soc., 33:597–607, 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Siegman:1987:SBP

[Sie87] T. Siegman. Scientific book publishing using TEX. TUGboat, 8:8–11, 1987. ISSN 0896-3207.

Sigler:1986:HSM

[Sig86] Stephen M. Sigler. The History of Statistics: The Measurement of Uncertainty Before 1900. Harvard University Press, Cambridge, MA, USA, 1986. ISBN 0-674-40340-1. xvi + 410 pp.

Sigmon:1992:MP

[Sig92] Kermit Sigmon. MATLAB primer. Technical report, Department of Mathematics, University of Florida, Gainesville, 1992. ???? pp. Available in TEX source code form by email from sigmon@math.

ufl.edu.

Sigmon:1994:MP

[Sig94] Kermit Sigmon. MATLAB Primer. CRC Press, Boca Raton, FL, USA, fourth edition, 1994. ISBN 0-8493-9440-6. 101 pp.

Simpson:1987:NOP

[Sim87] J. A. Simpson. The new OED project: A year’s work in lexicography. University Computing, 9:2–7, 1987. CODEN UN-COET. ISSN 0265-4385.

Sameh:1977:PAS

[SK77] A. H. Sameh and D. J. Kuck. A parallel QR algorithm for symmetric tridiagonal matrices. IEEE Trans. Comput., C-26 (2):147–153, 1977.

Skeel:1993:ENC

[SK93] Robert D. Skeel and Jerry B. Keiper. Elementary Numerical Computing with Mathematica. McGraw-Hill, New York, 1993. ISBN 0-07-057820-6. xiv + 434 pp.

Skeel:1979:SNS

[Ske79] Robert D. Skeel. Scaling for numerical stability in Gaussian elimination. J. Assoc. Comput. Mach., 26(3):494–526, 1979.

Skeel:1980:IRI

[Ske80] Robert D. Skeel. Iterative refinement implies numerical stability for Gaussian elimination. Math. Comp., 35(151):817– 832, 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1981:EER

[Ske81] Robert D. Skeel. Effect of equilibration on residual size for partial pivoting. SIAM J. Numer. Anal., 18(3):449–454, 1981.

Skeel:1989:SNB

[Ske89] Robert D. Skeel. Safety in numbers:

The boundless errors of numerical computation. Working Document 89-3, Department of Computer Science, University of Illinois, Urbana, IL, USA, 1989. 9 pp.

Savage:1954:TIF

[SL54] I. Richard Savage and Eugene Lukacs.

Tables of inverse of finite segments of the Hilbert matrix. In Olga Taussky, editor, Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, number 39 in Applied Mathematics Series, pages 105–108. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1954.

Slater:1987:PS

[Sla87] Robert Slater. Portraits in Silicon. The MIT Press, Cambridge, Massachusetts, 1987. ???? pp.

Slapnicar:1992:ASE

[Sla92] Ivan Slapnic̆ar. Accurate Symmetric Eigenreduction by a Jacobi Method. PhD thesis, Fernuniversität Hagen, Germany, ????, 1992. ???? pp.

Smale:1985:EAA

[Sma85] S. Smale. On the efficiency of algorithms of analysis. Bull. Amer. Math. Soc., 13:87–121, 1985.

Smale:1990:SRF

[Sma90] Steve Smale. Some remarks on the foundations of numerical analysis. SIAM Review, 32(2):211–220, 1990.

Smith:1962:ACD

[Smi62] Robert L. Smith. Algorithm 116: Complex division. Comm. ACM, 5(8): 435, 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1965:ASO

[Smi65] Francis J. Smith. An algorithm for summing orthogonal polynomial series and their derivatives with applications to curve-fitting and interpolation. Math. Comp., 19:33–36, 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1968:IPE

[Smi68a] R. A. Smith. Infinite product expansions for matrix n-th roots. J. Australian Math. Soc., 8:242–249, 1968. CO-DEN JAUMAX. ISSN 0004-9735.

Smith:1968:ME

[Smi68b] R. A. Smith. Matrix equation XA+ BX = C. SIAM J. Appl. Math., 16(1): 198–201, 1968.

Smith:1975:SAP

[Smi75] Jon M. Smith. Scientific Analysis on the Pocket Calculator. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1975. ISBN 0-471-79997-1. xii + 380 pp.

Smith:1981:DE

[Smi81] Jonathan D. H. Smith. In defense of eponymy. The Mathematical Intelligencer, 3(2):89–90, 1981. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Smith:1990:TR

[Smi90] Alan Jay Smith. The task of the referee. IEEE Computer, 23(4):65–71, 1990.

Smith:1991:AFP

[Smi91] David M. Smith. Algorithm 693:

A FORTRAN package for floating-point multiple-precision arithmetic. ACM Transactions on Mathematical Software, 17(2):273–283, 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smoktunowicz:1995:NSC

[Smo95] Alicja Smoktunowicz. A note on the strong componentwise stability of algorithms for solving symmetric linear systems. Demonstratio Mathematica, 28(2), 1995. CODEN DEMADO. ISSN 0420-1213.

Starnes:1946:EDC

[SN46] De Witt T. Starnes and Gertrude E.

Noyes. The English Dictionary from Cawdrey to Johnson 1604–1755. University of North Carolina Press, Chapel Hill, NC, USA, 1946. ISBN 90-272-4544-4. ???? pp. New edition with an introduction and a select bibliography by Gabriele Stein, John Benjamins Publishing Company, Amsterdam/Philadelphia, 1991.

Starke:1991:S

[SN91] Gerhard Starke and Wilhelm Niethammer. SOR for AX − XB = C. Linear Algebra and Appl., 154-156:355–375, 1991.

Snyder:1955:ISS

[Sny55] James N. Snyder. On the improvement of the solutions to a set of simultaneous linear equations using the ILLIAC. Mathematical Tables and Other Aids to Computation, 9:177–184, 1955.

Snyder:1966:CMN

[Sny66] Martin Avery Snyder. Chebyshev Methods in Numerical Approximation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1966. x + 114 pp.

Soederkvist:1993:PAO

[Söd93] Inge Söderkvist. Perturbation analysis of the orthogonal Procrustes problem. BIT, 33:687–694, 1993. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Society:1986:TWI

[SoIPUG86] Boston Computer Society and PC Technical Group of IBM PC Users Group. Technical wordprocessors for the IBM PC and compatibles. Notices Amer. Math. Soc., 33:8–37, 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Society:1987:TWI

[SoIPUG87] Boston Computer Society and PC Technical Group of IBM PC Users Group. Technical wordprocessors for the IBM PC and compatibles, Part IIB: Reviews. Notices Amer. Math. Soc., 34:462–491, 1987. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Solodovnikov:1980:ESE

[Sol80] V. I. Solodovnikov. Extension of Strassen’s estimate to the solution of arbitrary systems of linear equations. U.S.S.R. Computational Math. and Math. Phys., 19:21–33, 1980. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Sorensen:1982:NMM

[Sor82] D. C. Sorensen. Newton’s method with a model trust region modification. SIAM J. Numer. Anal., 19:409–426, 1982.

Sorensen:1985:APP

[Sor85] D. C. Sorensen. Analysis of pairwise pivoting in Gaussian elimination. IEEE Trans. Comput., C-34:274–278, 1985.

Sorensen:1990:NAD

[Sor90] D. C. Sorensen. Numerical aspects of divide and conquer methods for eigenvalue problems. In Griffiths and Watson [GW90], pages 229–242.

Sorensen:1992:IAP

[Sor92] D. C. Sorensen. Implicit application of polynomial filters in a k-step Arnoldi method. SIAM J. Matrix Anal. Appl., 13 (1):357–385, 1992.

Schreiber:1988:BRT

[SP88] Robert S. Schreiber and Beresford N.

Parlett. Block reflectors: Theory and computation. SIAM J. Numer. Anal., 25 (1):189–205, 1988.

Spellucci:1980:ABA

[Spe80] P. Spellucci. An approach to backward analysis for linear and nonlinear iterative methods. Computing, 25:269–282, 1980. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Spedicato:1991:CAS

[Spe91] E. Spedicato, editor. Computer Algorithms for Solving Linear Algebraic Equations: The State of the Art, volume 77 of NATO ASI Series F. Springer-Verlag, Berlin, 1991.

Spiess:1976:UZD

[Spi76] J. Spieß. Untersuchungen des Zeitgewinns durch neue Algorithmen zur Matrix-Multiplikation. Computing, 17: 23–36, 1976. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Spivak:1990:JTG

[Spi90] Michael D. Spivak. The Joy of TEX: A Gourmet Guide to Typesetting with the AMS-TEX Macro Package. American Mathematical Society, Providence, RI, USA, second edition, 1990. ???? pp.

Spivak:1991:LT

[Spi91] Michael D. Spivak. LAMS-TEX. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1991. ???? pp.

Squire:1987:LSC

[Squ87] William Squire. Landen’s solution of the cubic. American Journal of Physics, 55(4):374–375, 1987. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Schonfelder:1980:ECP

[SR80] J. L. Schonfelder and M. Razaz. Error control with polynomial approximations. IMA J. Numer. Anal., 1:105–114, 1980.

Smits:1990:LAC

[SR90] J. G. M. M. Smits and J. J. M. Rijpkema. A linear algebra course with PC-Matlab: Some experiences. Technical report, Preprint, Department of Mathematics and Computer Science, Eindhoven University of Technology, The Netherlands, ????, August 1990. ???? pp.

Shampine:1995:SMO

[SR95] Lawrence F. Shampine and Mark W.

Reichelt. The Matlab ODE suite. Manuscript, 1995.

Strakos:1994:NSG

[SRD94] Z. Strakos̆, Miroslav Rozloz̆ník, and Jitka Drkos̆ová. Numerical stability of GMRES. Technical Report 579, Institute of Computer Science, Academy of Sciences of the Czech Republic, ????, January 1994. ???? pp.

Schwarz:1973:NAS

[SRS73] H. R. Schwarz, H. Rutishauser, and E. Stiefel. Numerical Analysis of Symmetric Matrices. Prentice-Hall, Englewood Cliffs, NJ, USA, 1973. ISBN 0-13-626556-1. ???? pp.

Schneider:1962:CTS

[SS62] Hans Schneider and W. Gilbert Strang.

Comparison theorems for supremum norms. Numer. Math., 4:15–20, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Soederstroem:1974:NPI

[SS74] T. Söderström and G. W. Stewart. On the numerical properties of an iterative method for computing the Moore– Penrose generalized inverse. SIAM J. Numer. Anal., 11:61–74, 1974.

Singer:1976:RSP

[SS76] Burton Singer and Seymour Spilerman. The representation of social processes by Markov models. Amer. J. Sociology, 82 (1):1–54, 1976.

Smoktunowicz:1984:BCI

[SS84] Alicja Smoktunowicz and Jolanta Sokolnicka. Binary cascades iterative refinement in doubled-mantissa arithmetics. BIT, 24:123–127, 1984. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Shroff:1989:CCJ

[SS89] Gautam M. Shroff and Robert S. Schreiber. On the convergence of the cyclic Jacobi method for parallel block orderings. SIAM J. Matrix Anal. Appl., 10 (3):326–346, 1989.

Schreiber:1990:TTC

[SS90a] R. S. Schreiber and H. D. Simon. Towards the tetraflop in CFD. RIACS Technical Report TR 92.12, NASA Ames Research Center, Moffett Field, CA, USA, May 1990. ???? pp. To appear in Parallel CFD—Implementations and Results Using Parallel Computers, H. D. Simon, ed., MIT Press, 1992.

Stewart:1990:MPT

[SS90b] Gilbert Wright (Gilbert W.) Stewart and Ji-guang Sun. Matrix Perturbation Theory. Computer science and scientific computing. Academic Press, New York, NY, USA, 1990. ISBN 0-12-670230-6 (hardcover). xv + 365 pp. LCCN QA871

.S775 1990. URL http://catalog.
hathitrust.org/api/volumes/oclc/
21227976.html;   http://www.loc.
gov/catdir/description/els032/90033378.
htm;   http://www.loc.gov/catdir/
description/els032/90033378.html;
 http://www.loc.gov/catdir/toc/els031/
90033378.htm;   http://www.loc.gov/
catdir/toc/els031/90033378.html.

Shougen:1991:ASP

[SS91] Wang Shougen and Zhao Shuqin. An algorithm for Ax = λBx with symmetric and positive-definite A and B^{∗}. SIAM J. Matrix Anal. Appl., 12(4):654–660, 1991.

Sanz-Serna:1992:SIH

[SS92] J. M. Sanz-Serna. Symplectic integrators for Hamiltonian problems: An overview. In Acta Numerica, pages 243–

286. Cambridge University Press, 1992.

Sanz-Serna:1994:NHP

[SSC94] J. M. Sanz-Serna and M. P. Calvo.

Numerical Hamiltonian Problems. Chapman and Hall, London, 1994. ISBN 0-412-54290-0. xii + 207 pp.

Sanz-Serna:1993:SCS

[SSL93] J. M. Sanz-Serna and S. Larsson.

Shadows, chaos, and saddles. Applied Numerical Mathematics, 13:181–190, 1993.

Santoro:1987:GCV

[SSSU87] Nicola Santoro, Jeffrey B. Sidney, Stuart J. Sidney, and Jorge Urrutia. Geometric containment and vector dominance. Theoret. Comput. Sci., 53:345– 352, 1987. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Squire:1975:CQM

[ST75] W. Squire and G. E. Trapp. Chebychev quadrature methods for inverting Laplace transforms. Technical report, Technical Report, Dept. of Statistics and Computer Science, West Virginia University, ????, 1975. ???? pp.

Shapiro:1980:APT

[ST80] H. Shapiro and O. Taussky. Alternative proofs of a theorem of Moyls and Marcus on the numerical range of a square matrix. Linear and Multilinear Algebra, 8:337–340, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). field of values.

Schauer:1986:SLS

[ST86] U. Schauer and R. A. Toupin. Solving large sparse linear systems with guaranteed accuracy. In Willard L. Miranker and Richard A. Toupin, editors, Accurate Scientific Computations, Proceedings, 1985, volume 235 of Lecture Notes in Computer Science, pages 142–167. Springer-Verlag, Berlin, 1986.

Sorensen:1991:OEC

[ST91] D. C. Sorensen and P. T. P. Tang. On the orthogonality of eigenvectors computed by divide-and-conquer techniques. SIAM J. Numer. Anal., 28(6):1752–1775, 1991.

Stainton:1977:BA

[Sta77a] Elsie Myers Stainton. A bag for authors. Scholarly Publishing, 8(4):335–345, 1977. ISSN 0036-634X.

Stainton:1977:BE

[Sta77b] Elsie Myers Stainton. A bag for editors. Scholarly Publishing, 8(2):111–119, 1977. ISSN 0036-634X.

Stafney:1978:FMT

[Sta78] James D. Stafney. Functions of a matrix and their norms. Linear Algebra and Appl., 20:87–94, 1978. Correction in Linear Alg. and Appl., 39 (1981), pp. 259– 260.

Stainton:1980:UD

[Sta80] Elsie Myers Stainton. The uses of dictionaries. Scholarly Publishing, 11(3): 229–241, 1980. ISSN 0036-634X.

Stallman:1986:GEM

[Sta86] Richard Stallman. GNU Emacs Manual, Emacs Version 18. Free Software Foundation, 675 Mass Ave, Cambridge, MA 02139, USA, Tel: (617) 876-3296, USA, fifth edition, October 1986. ???? pp.

Startz:1988:IPC

[Sta88] R. Startz. 8087/80287/80387 for the IBM PC & Compatibles. Brady, New York, third edition, 1988. ???? pp.

Stainton:1991:FAC

[Sta91] Elsie Myers Stainton. The Fine Art of Copyediting. Columbia University Press, New York, NY, USA, 1991. ISBN 0-231-06961-8. ???? pp.

Stewart:1967:MDM

[Ste67] G. W. Stewart. A modification of Davidon’s minimization method to accept difference approximations of derivatives. J. Assoc. Comput. Mach., 14:72–83, 1967.

Stewart:1969:SIF

[Ste69] G. W. Stewart. Some iterations for factoring a polynomial. Numer. Math., 13:458–471, 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1971:EAA

[Ste71a] G. W. Stewart. Error analysis of the algorithm for shifting the zeros of a polynomial by synthetic division. Math. Comp., 25(113):135–139, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1971:EBA

[Ste71b] G. W. Stewart. Error bounds for approximate invariant subspaces of closed linear operators. SIAM J. Numer. Anal., 8(4):796–808, 1971.

Steenrod:1973:HWMa

[Ste73a] Norman E. Steenrod. How to write mathematics. In Steenrod et al. [SHSD73], pages 1–17. LCCN QA41 .H6.

Stetter:1973:ADM

[Ste73b] Hans J. Stetter. Analysis of Discretization Methods for Ordinary Differential Equations. Springer-Verlag, Berlin, 1973. ISBN 3-540-06008-1. xvi + 388 pp.

Stewart:1973:CDM

[Ste73c] G. W. Stewart. Conjugate direction methods for solving systems of linear equations. Numer. Math., 21:285–297, 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1973:EPB

[Ste73d] G. W. Stewart. Error and perturbation bounds for subspaces associated with certain eigenvalue problems. SIAM Review, 15(4):727–764, 1973.

Stewart:1973:IMC

[Ste73e] G. W. Stewart. Introduction to Matrix Computations. Academic Press, New York, 1973. ISBN 0-12-670350-7. xiii + 441 pp.

Sterbenz:1974:FPC

[Ste74a] Pat H. Sterbenz. Floating-Point Computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1974. ISBN 0-13-322495-3. xiv + 316 pp.

Stewart:1974:MPE

[Ste74b] G. W. Stewart. Modifying pivot elements in Gaussian elimination. Math. Comp., 28(126):537–542, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1975:CMC

[Ste75a] G. W. Stewart. The convergence of the method of conjugate gradients at isolated extreme points of the spectrum. Numer. Math., 24:85–93, 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1975:IPT

[Ste75b] G. W. Stewart. An inverse perturbation theorem for the linear least squares problem. ACM SIGNUM Newsletter, 10: 39–40, 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1975:MSI

[Ste75c] G. W. Stewart. Methods of simultaneous iteration for calculating eigenvectors of matrices. In Miller [Mil75a], pages 185–196.

Stewart:1975:CMM

[Ste75d] Ian Stewart. Concepts of Modern Mathematics. Penguin, ????, 1975. ISBN 0-14-021849-1. viii + 315 pp.

Stewart:1977:PPI

[Ste77a] G. W. Stewart. On the perturbation of pseudo-inverses, projections and linear least squares problems. SIAM Review, 19 (4):634–662, 1977.

Stewart:1977:PBF

[Ste77b] G. W. Stewart. Perturbation bounds for the QR factorization of a matrix. SIAM J. Numer. Anal., 14(3):509–518, 1977.

Stewart:1977:RDL

[Ste77c] G. W. Stewart. Research, development, and LINPACK. In John R. Rice, editor, Mathematical Software III, pages 1–14. Academic Press, New York, 1977.

Stewart:1978:CNT

[Ste78] W. J. Stewart. A comparison of numerical techniques in Markov modeling. Comm. ACM, 21(2):144–152, 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stewart:1979:ERE

[Ste79a] G. W. Stewart. The effects of rounding error on an algorithm for downdating a Cholesky factorization. J. Inst. Maths Applics, 23:203–213, 1979.

Stewart:1979:NPS

[Ste79b] G. W. Stewart. A note on the perturbation of singular values. Linear Algebra and Appl., 28:213–216, 1979.

Stern:1980:WIF

[Ste80a] Nancy Stern. Who “invented” the first electronic digital computer? Annals of the History of Computing, 2(4): 375–376, 1980.

Stewart:1980:EGR

[Ste80b] G. W. Stewart. The efficient generation of random orthogonal matrices with an application to condition estimators. SIAM J. Numer. Anal., 17(3):403–409, 1980.

Sternberg:1981:HCS

[Ste81a] David Sternberg. How to Complete and Survive a Doctoral Dissertation. St. Martin’s Press, New York, 1981. ISBN 0-312-39606-6. 231 pp.

Stewart:1981:CDH

[Ste81b] G. W. Stewart. Constrained definite hessians tend to be well conditioned. Math. Prog., 21:235–238, 1981. CODEN MHPGA4. ISSN 0025-5610.

Stewart:1981:MCH

[Ste81c] G. W. Stewart. Matrix calculations on hand-held calculators. ACM SIGNUM Newsletter, 16:10–13, 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1981:IDN

[Ste81d] G. W. Stewart. On the implicit deflation of nearly singular systems of linear equations. SIAM Journal on Scientific and Statistical Computing, 2(2):136– 140, June 1981. CODEN SIJCD4. ISSN 0196-5204.

Stewart:1983:CEB

[Ste83a] G. W. Stewart. Computable error bounds for aggregated Markov chains. J. Assoc. Comput. Mach., 30:271–285, 1983.

Stewart:1983:MCG

[Ste83b] G. W. Stewart. A method for computing the generalized singular value decomposition. In Kågström and Ruhe [KR83], pages 207–220.

Stewart:1984:ABS

[Ste84a] G. W. Stewart. On the asymptotic behavior of scaled singular value and QR decompositions. Math. Comp., 43:483– 489, 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1984:IPN

[Ste84b] G. W. Stewart. On the invariance of perturbed null vectors under column scaling. Numer. Math., 44:61–65, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1984:RD

[Ste84c] G. W. Stewart. Rank degeneracy. SIAM J. Sci. Stat. Comput., 5:403–413, 1984.

Stewart:1985:JLA

[Ste85a] G. W. Stewart. A Jacobi-like algorithm for computing the Schur decomposition of a non-Hermitian matrix. SIAM J. Sci. Stat. Comput., 6(4):853–864, 1985.

Stewart:1985:NCD

[Ste85b] G. W. Stewart. A note on complex division. ACM Transactions on Mathematical Software, 11(3):238–241, 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stewart:1987:CLS

[Ste87a] G. W. Stewart. Collinearity and least squares regression, (with discussion). Statistical Science, 2:68–100, 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Stewart:1987:ISC

[Ste87b] G. W. Stewart. Invariant subspaces and capital punishment (a participatory paper). Technical report, Technical Report CS-TR 1923, Department of Computer Science, University of Maryland, ????, 1987. ???? pp.

Stewart:1989:DGP

[Ste89] Ian Stewart. Does God Play Dice?

The Mathematics of Chaos. Basil Blackwell, Oxford, 1989. ???? pp.

Sterrett:1990:UWT

[Ste90a] Andrew Sterrett, editor. Using Writing to Teach Mathematics. MAA Notes Number 16. Mathematical Association of America, Washington, DC, USA, 1990. ISBN 0-88385-066-4. ???? pp.

Stewart:1990:CMC

[Ste90b] G. W. Stewart. Communication and matrix computations on large message passing systems. Parallel Computing, 16: 27–40, 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Stewart:1990:ICC

[Ste90c] G. W. Stewart. Incremental condition calculation and column selection. Technical report, Technical Report CS-TR 2495, Department of Computer Science, University of Maryland, ????, 1990. ???? pp.

Stewart:1990:IMS

[Ste90d] G. W. Stewart. An iterative method for solving linear equalities. In Cox and Hammarling [CH90c], pages 241–247. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$45.00.

Stewart:1990:PTL

[Ste90e] G. W. Stewart. Perturbation theory and least squares with errors in the variables. In Statistical Analysis of Measurement Error Models and Applications [BF90], pages 171–181.

Stewart:1990:PTS

[Ste90f] G. W. Stewart. Perturbation theory for the singular value decomposition. Technical Report CS-TR 2539, Department of Computer Science, University of Maryland, College Park, MD, USA, 1990. ???? pp.

Stewart:1990:SPT

[Ste90g] G. W. Stewart. Stochastic perturbation theory. SIAM Review, 32(4):579–610, 1990.

Stewart:1990:UAS

[Ste90h] G. W. Stewart. An updating algorithm for subspace tracking. Technical report, Technical Report CS-TR 2494, Department of Computer Science, University of Maryland, ????, 1990. ???? pp. revised January 1991.

Stewart:1991:FFT

[Ste91a] G. W. Stewart. FTP—file transfer program. ACM SIGNUM Newsletter, 26 (4):2–3, 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1991:FTR

[Ste91b] G. W. Stewart. FTPing technical reports. NA Digest, 91(39):??, 1991. Electronic mail magazine: na.help@na-net.

                                        ornl.gov.

Stewart:1991:LLS

[Ste91c] G. W. Stewart. Lanczos and linear systems. Technical report, Technical Report CS-TR 2641, Department of Computer Science, University of Maryland, ????, 1991. ???? pp.

Stewart:1991:ARR

[Ste91d] G. W. Stewart. On an algorithm for refining a rank-revealing URV factorization and a perturbation theorem for singular values. Technical Report CS-TR 2626, Department of Computer Science, University of Maryland, College Park, MD, USA, 1991. ???? pp.

Stewart:1991:TSR

[Ste91e] G. W. Stewart. Two simple residual bounds for the eigenvalues of a Hermitian matrix. SIAM J. Matrix Anal. Appl., 12: 205–208, 1991.

Stewart:1992:EAQ

[Ste92a] G. W. Stewart. Error analysis of QR updating with exponential windowing. Math. Comp., 59(199):135–140, 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1992:NGS

[Ste92b] G. W. Stewart. Note on a generalized Sylvester equation. IMA Preprint Series #985, Institute for Mathematics and its Applications, University of Minnesota, Minneapolis, MN, USA, May 1992. 3 pp.

Stewart:1993:GEP

[Ste93a] G. W. Stewart. Gaussian elimination, perturbation theory, and Markov chains. In Carl D. Meyer and Robert J. Plemmons, editors, Linear Algebra, Markov Chains, and Queueing Models, volume 48 of IMA Volumes in Mathematics and its Applications, pages 59–69. Springer-Verlag, New York, 1993.

Stewart:1993:EHS

[Ste93b] G. W. Stewart. On the early history of the singular value decomposition. SIAM Review, 35(4):551–566, 1993.

Stewart:1993:PLC

[Ste93c] G. W. Stewart. On the perturbation of LU, Cholesky, and QR factorizations. SIAM J. Matrix Anal. Appl., 4(4):1141– 1145, 1993.

Stewart:1993:URR

[Ste93d] G. W. Stewart. Updating a rankrevealing ULV decomposition. SIAM J. Matrix Anal. Appl., 14(2):494–499, 1993.

Stewart:1994:GSG

[Ste94a] G. W. Stewart. Gauss, statistics, and Gaussian elimination. Technical Report TR-3307, Department of Computer Science, University of Maryland, College Park, MD, USA, August 1994. 14 pp.

Stewart:1994:GQD

[Ste94b] G. W. Stewart. On graded QR decompositions of products of matrices. Technical Report TR-3263, Department of Computer Science, University of Maryland, College Park, MD, USA, April 1994. 14 pp.

Stewart:1994:MCS

[Ste94c] G. W. Stewart. On Markov chains with sluggish transients. Technical Report TR-3306, Department of Computer Science, University of Maryland, College Park, MD, USA, June 1994. 8 pp.

Stewart:1994:CMI

[Ste94d] G. W. Stewart. On the convergence of multipoint iterations. Numer. Math., 68:143–147, 1994. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1994:SSU

[Ste94e] G. W. Stewart. On the stability of sequential updates and downdates. Technical Report TR-3238, Department of Computer Science, University of Maryland, College Park, MD, USA, March 1994. 17 pp.

Stewart:1994:PTR

[Ste94f] G. W. Stewart. Perturbation theory for rectangular matrix pencils. Linear Algebra and Appl., 208/209:297–301, 1994.

Stickel:1991:SEU

[Sti91] Eberhard U. Stickel. Separating eigenvalues using the matrix sign function. Linear Algebra and Appl., 148:75–88, 1991.

Stone:1962:BPR

[Sto62] Betty Jane Stone. Best possible ratios of certain matrix norms. Numer. Math., 4:114–116, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoneham:1975:PJO

[Sto75a] Harry Stoneham. Play Jazz Organ.

Wise Publications, London, 1975. ISBN 0-86001-060-0. ???? pp.

Stoutemyer:1975:AOU

[Sto75b] D. R. Stoutemyer. Analytical optimization using computer algebraic manipulation. ACM Transactions on Mathematical Software, 1:147–164, 1975. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stoutemyer:1977:AEA

[Sto77] David R. Stoutemyer. Automatic error analysis using computer algebraic manipulation. ACM Transactions on Mathematical Software, 3(1):26–43, 1977. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stoneham:1980:HSS

[Sto80] Harry Stoneham. The Harry Stoneham Style, volume 19 of Organ Adventure Series. Hal Leonard, ????, 1980. ???? pp.

Stoutemyer:1991:CMC

[Sto91] David R. Stoutemyer. Crimes and misdemeanours in the computer algebra trade. Notices Amer. Math. Soc., 38(7): 778–785, 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stpiczynski:1993:EAT

[Stp93] Przemysław Stpiczyński. Error analysis of two parallel algorithms for solving linear recurrence systems. Parallel Computing, 19:917–923, 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Strassen:1969:GEO

[Str69] V. Strassen. Gaussian elimination is not optimal. Numer. Math., 13:354–356, 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stroem:1972:MNL

[Str72] Torsten Ström. Minimization of norms and logarithmic norms by diagonal similarities. Computing, 10:1–7, 1972. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Stroem:1975:LN

[Str75] Torsten Ström. On logarithmic norms.

SIAM J. Numer. Anal., 12(5):741–753, 1975.

Strang:1986:IAM

[Str86a] Gilbert Strang. Introduction to Applied Mathematics. Wellesley-Cambridge Press, Wellesley, MA, USA, 1986. ISBN 0-9614088-0-4. ix + 758 pp.

Strang:1986:PTM

[Str86b] Gilbert Strang. A proposal for Toeplitz matrix calculations. Studies in Applied Mathematics, 74:171–176, 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic).

Strang:1987:KAP

[Str87] G. Strang. Karmarkar’s algorithm and its place in applied mathematics. The Mathematical Intelligencer, 9:4–10, 1987. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Strang:1988:LAA

[Str88] Gilbert Strang. Linear Algebra and Its Applications. Harcourt, Brace, Jovanovich, San Diego, CA, USA, third edition, 1988. ISBN 0-15-551005-3. xii + 505 pp.

Strang:1989:PLA

[Str89] G. Strang. Patterns in linear algebra.

Amer. Math. Monthly, 96:105–117, 1989.

Strakos:1991:RCR

[Str91] Z. Strakos̆. On the real convergence rate of the conjugate gradient method. Linear Algebra and Appl., 154-156:535– 549, 1991.

Strang:1993:ILA

[Str93] Gilbert Strang. Introduction to Linear Algebra. Wellesley-Cambridge Press, Wellesley, MA, USA, 1993. ISBN 0-9614088-5-5. viii + 472 pp.

Stummel:1980:REA

[Stu80] F. Stummel. Rounding error analysis of elementary numerical algorithms. Computing, Suppl. 2:169–195, 1980. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Stummel:1981:PTE

[Stu81] F. Stummel. Perturbation theory for evaluation algorithms of arithmetic expressions. Math. Comp., 37(156):435– 473, 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stummel:1982:OEE

[Stu82] F. Stummel. Optimal error estimates for Gaussian elimination in floating-point arithmetic. Z. Angew. Math. Mech, 62: T355–T357, 1982.

Stummel:1985:SOE

[Stu85a] F. Stummel. Strict optimal error estimates for Gaussian elimination. Z. Angew. Math. Mech, 65:T405–T407, 1985.

Stummel:1985:FEAa

[Stu85b] Friedrich Stummel. Forward error analysis of Gaussian elimination, Part I: Error and residual estimates. Numer. Math., 46:365–395, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stummel:1985:FEAb

[Stu85c] Friedrich Stummel. Forward error analysis of Gaussian elimination, Part II: Stability theorems. Numer. Math., 46:397–415, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shieh:1984:MSF

[STW84] L. S. Shieh, Y. T. Tsay, and C. T.

Wang. Matrix sector functions and their applications to system theory. IEE Proc., 131(5):171–181, 1984.

Shieh:1983:SPM

[STY83] L. S. Shieh, Y. T. Tsay, and R. E.

Yates. Some properties of matrix sign fucntions derived from continued fractions. IEE Proc., 130(3):111–118, 1983.

Saari:1984:NMC

[SU84] Donald G. Saari and John B. Urenko.

Newton’s method, circle maps, and chaotic motion. Amer. Math. Monthly, 91(1):3–17, 1984.

Summers:1987:LDC

[Sum87] Della Summers, editor. Longman Dictionary of Contemporary English. Longman, Harlow, Essex, UK, second edition, 1987. ISBN 0-582-84222-0, 0-582-84223-9. F53 + 1229 + B29 pp. LCCN PE1628.L58 1987.

Sun:1991:PBC

[Sun91] Ji-guang Sun. Perturbation bounds for the Cholesky and QR factorizations. BIT, 31:341–352, 1991. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Sun:1992:CPB

[Sun92a] Ji-guang Sun. Componentwise perturbation bounds for some matrix decompositions. BIT, 32(4):702–714, 1992. CO-DEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Sun:1992:REP

[Sun92b] Ji-guang Sun. Rounding-error and perturbation bounds for the Cholesky and LDL^{T} factorizations. Linear Algebra and Appl., 173:77–97, 1992.

Sun:1992:SFU

[Sun92c] Sun Microsystems, Inc., Mountain View, CA, USA. SPARCompiler FOR-TRAN 2.0.1: User’s Guide, October 1992. ???? pp. Part No. 800-6552-11, Revision A.

Sun:1992:SFN

[Sun92d] Sun Microsystems, Inc., Mountain View, CA, USA. SPARCompiler FOR-TRAN: Numerical Computation Guide, October 1992. ???? pp. Part No. 800-7097-11, Revision A.

Sun:1995:NBP

[Sun95a] Ji-guang Sun. A note on backward perturbations for the Hermitian eigenvalue problem. BIT, 35:385–393, 1995. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Sun:1995:PBQ

[Sun95b] Ji-guang Sun. On perturbation bounds for the QR factorization. Linear Algebra and Appl., 215:95–111, 1995.

Sun:1995:OBP

[Sun95c] Ji-guang Sun. Optimal backward perturbation bounds for the linear LS problem with multiple right-hand sides. IMA J. Numer. Anal., page ??, 1995. Submitted.

Saff:1977:PRA

[SV77] E. B. Saff and R. S. Varga, editors.

Padé and Rational Approximation: Theory and Applications. Academic Press, New York, 1977.

Sahasrabudhe:1979:ESV

[SV79] S. C. Sahasrabudhe and P. M. Vaidya.

Estimation of singular values of an image matrix. IEEE Trans. Acoustics, Speech, and Signal Processing, ASSP-27:434–436, 1979. CODEN IETABA. ISSN 0096-3518.

Schreiber:1989:SEW

[SV89] Robert S. Schreiber and Charles F.

Van Loan. A storage efficient WY representation for products of Householder transformations. SIAM J. Sci. Stat. Comput., 10:53–57, 1989.

Stoer:1962:TDM

[SW62] J. Stoer and C. Witzgall. Transformations by diagonal matrices in a normed space. Numer. Math., 4:158–171, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strunk:1979:ES

[SW79] William Strunk, Jr. and E. B. White.

The Elements of Style. MacMillan Publishing Company, New York, NY, USA, third edition, 1979. ISBN 0-02-418200-1. xvii + 92 pp.

Scott:1981:SSD

[SW81] D. S. Scott and R. C. Ward. Solving symmetric-definite quadratic λ-matrix problems without factorization. SIAM J. Sci. Stat. Comput., 3(1):58–67, 1981.

Sorensen:1982:ATF

[SW82] D. C. Sorensen and R. J.-B. Wets, editors. Algorithms and Theory in Filtering and Control. Number 18 in Mathematical Programming Study. North-Holland, Amsterdam, The Netherlands, 1982. ???? pp.

Shampine:1984:SOD

[SW84] L. F. Shampine and H. A. Watts. Software for ordinary differential equations. In W. R. Cowell, editor, Sources and Development of Mathematical Software, pages 112–133. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Spaeth:1987:OLA

[SW87] H. Späth and G. A. Watson. On orthogonal linear ℓ_{1} approximation. Numer. Math., 51:531–543, 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Simon:1988:SP

[SW88] B. Simon and R. M. Wilson. Supercalculators on the PC. Notices Amer. Math. Soc., 35:978–1001, 1988. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shepherd:1989:MCW

[SW89] David Shepherd and Greg Wilson. Making chips that work. New scientist, pages 61–64, 1989. CODEN NWSCAL. ISSN 0262-4079, 0028-6664. 13 May.

Steele:1990:HPF

[SW90] Guy L. Steele, Jr. and Jon L. White. How to print floating-point numbers accurately. ACM SIGPLAN Notices, 25(6):112–126, 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soederkvist:1991:PCN

[SW91] Inge Söderkvist and Per-Åke Wedin.

A practical condition number for a configuration of points defining a rigid body movement. Report UMINF-91.29, Institute of Information Processing, University of Umeå, Sweden, December 1991. ???? pp.

Soederkvist:1994:CNA

[SW94] Inge Söderkvist and Per-Åke Wedin.

On condition numbers and algorithms for determining a rigid body movement. BIT, 34:424–436, 1994. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Swales:1971:WSE

[Swa71] John Swales. Writing Scientific English. Thomas Nelson, London, 1971. ???? pp.

Swann:1972:DSM

[Swa72] W. H. Swann. Direct search methods. In W. Murray, editor, Numerical Methods for Unconstrained Optimization, pages 13–28. Academic Press, London, 1972.

Swann:1974:COD

[Swa74] W. H. Swann. Constrained optimization by direct search. In P. E. Gill and W. Murray, editors, Numerical Methods for Constrained Optimization, pages 191– 217. Academic Press, London, 1974.

Swanson:1979:MTC

[Swa79] Ellen Swanson. Mathematics Into Type: Copy Editing and Proofreading of Mathematics for Editorial Assistants and Authors. American Mathematical Society, Providence, RI, USA, revised edition, 1979. ISBN 0-8218-0053-1. ???? pp.

Swan:1980:PEU

[Swa80a] Michael Swan. Practical English Usage. Oxford University Press, ????, 1980. ISBN 0-19-431185-6. ???? pp.

Swartzlander:1980:CA

[Swa80b] Earl E. Swartzlander, Jr., editor. Computer Arithmetic, volume 21 of Benchmark Papers in Electrical Engineering and Computer Science. Dowden, Hutchinson and Ross, Stroudsburg, PA, USA, 1980. ???? pp.

Swales:1988:EES

[Swa88] John Swales. Episodes in ESP: A Source and Reference Book on the Development of English for Science and Technology. Prentice Hall International, Hemel Hempstead, Hampshire, 1988. ISBN 0-13-283383-2. ???? pp.

Swarztrauber:1993:PAC

[Swa93] Paul N. Swarztrauber. A parallel algorithm for computing the eigenvalues of a symmetric tridiagonal matrix. Math. Comp., 60(202):651–668, 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1976:SNO

[SWD76] L. F. Shampine, H. A. Watts, and S. M. Davenport. Solving nonstiff ordinary differential equations—the state of the art. SIAM Review, 18:376–411, 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sweeney:1965:AFP

[Swe65] D. W. Sweeney. An analysis of floating-point addition. IBM Systems Journal, 4:31–42, 1965. CODEN IBMSA7. ISSN 0018-8670. Reprinted in [Swa80b, pp. 317–328].

Sweet:1984:FTO

[Swe84] D. R. Sweet. Fast Toeplitz orthogonalization. Numer. Math., 43:1–21, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Szidarovszky:1978:PPN

[SY78] Ferenc Szidarovszky and Sidney Yakowitz. Principles and Procedures of Numerical Analysis. Plenum, New York, 1978. ISBN 0-306-40087-1. xii + 331 pp.

Sylvester:1840:NE

[Syl40] J. J. Sylvester. Note on elimination. Philosophical Magazine, 17:379–380, 1840. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Ano04, p. 58].

Sylvester:1850:AAN

[Syl50] J. J. Sylvester. Additions to the articles, “On a New Class of Theorems,” and “On Pascal’s Theorem”. Philosophical Magazine, 37:363–370, 1850. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Ano04, pp. 1451–151].

Sylvester:1851:ECT

[Syl51a] J. J. Sylvester. Explanation of the coincidence of a theorem given by Mr Sylvester in the December number of this journal, with one stated by Professor Donkin in the June number of the same. Philosophical Magazine, (Fourth Series) 1:44–46, 1851. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Ano04, pp. 217– 218].

Sylvester:1851:RBM

[Syl51b] J. J. Sylvester. On the relation between the minor determinants of linearly equivalent quadratic functions. Philosophical Magazine, (Fourth Series) 1:295– 305, 1851. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Ano04, pp. 241– 250].

Sylvester:1884:LM

[Syl84] J. J. Sylvester. Sur l’equation en matrices px = xq. Comptes Rendus de l’Académie des Sciences, pages 67–71, 115–116, 1884.

Scherer:1980:SNR

[SZ80] R. Scherer and K. Zeller. Shorthand notation for rounding errors. Computing, Suppl. 2:165–168, 1980. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Stewart:1991:EGM

[SZ91a] G. W. Stewart and G. Zhang. Eigenvalues of graded matrices and the condition numbers of a multiple eigenvalue. Numer. Math., 58:703–712, 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1991:DMS

[SZ91b] G. W. Stewart and G. Zhang. On a direct method for the solution of nearly uncoupled Markov chains. Numer. Math., 59:1–11, 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Turnbull:1932:ITC

[TA32] H. W. Turnbull and A. C. Aitken. An Introduction to the Theory of Canonical Matrices. Blackie, Glasgow, Scotland, 1932. xiii + 200 pp. Reprinted with appendix, 1952.

Tal-Ezer:1991:HDP

[Tal91] Hillel Tal-Ezer. High degree polynomial interpolation in Newton form.

SIAM J. Sci. Stat. Comput., 12(3):648– 667, 1991.

Tanabe:1975:NTE

[Tan75] K. Tanabe. Neumann-type expansion of reflexive generalized inverses of a matrix and the hyperpower iterative method. Linear Algebra and Appl., 10: 163–175, 1975.

Tang:1989:TDI

[Tan89] Ping Tak Peter Tang. Table-driven implementation of the exponential function in IEEE floating-point arithmetic. ACM Transactions on Mathematical Software, 15(2):144–157, 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1990:AET

[Tan90a] Ping Tak Peter Tang. Accurate and efficient testing of the exponential and logarithm functions. ACM Transactions on Mathematical Software, 16(3): 185–200, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1990:TDI

[Tan90b] Ping Tak Peter Tang. Table-driven implementation of the logarithm function in IEEE floating-point arithmetic. ACM Transactions on Mathematical Software, 16(3):378–400, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1991:TLA

[Tan91] Ping Tak Peter Tang. Table-lookup algorithms for elementary functions and their error analysis. In Peter Kornerup and David W. Matula, editors, Proc. 10th IEEE Symposium on Computer Arithmetic, pages 232–236. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991.

Tang:1992:TDI

[Tan92] Ping Tak Peter Tang. Table-driven implementation of the expm1 function in IEEE floating-point arithmetic. ACM Transactions on Mathematical Software, 18(2):211–222, 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1994:DCE

[Tan94] Ping Tak Peter Tang. Dynamic condition estimation and Rayleigh–Ritz approximation. SIAM J. Matrix Anal. Appl., 15(1):331–346, 1994.

Tao:1984:CSM

[Tao84a] Pham Dinh Tao. Convergence of a subgradient method for computing the bound norm of matrices. Linear Algebra and Appl., 62:163–182, 1984. In French.

Tao:1984:SMC

[Tao84b] Pham Dinh Tao. Some methods for computing the maximum of quadratic form on the unit ball of the maximum norm. Numer. Math., 45:377–401, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). In French.

Tarutz:1992:TEP

[Tar92] Judith A. Tarutz. Technical Editing: The Practical Guide for Editors and Writers. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56356-8. ???? pp.

Taussky:1949:RCC

[Tau49] Olga Taussky. A remark concerning the characteristic roots of the finite segments of the Hilbert matrix. Quarterly Journal of Mathematics, 20:80–83, 1949. CODEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Taussky:1950:NNA

[Tau50] Olga Taussky. Notes on numerical analysis—2. note on the condition of matrices. M.T.A.C., 4:111–112, 1950.

Taub:1963:JNC

[Tau63] A. H. Taub, editor. John von Neumann Collected Works, volume V, Design of Computers, Theory of Automata and Numerical Analysis. Pergamon, New York, NY, USA, 1963. ix + 784 pp.

Taussky:1967:PDM

[Tau67] O. Taussky. Positive-definite matrices. In Inequalities [Shi67], pages 309– 319.

Taussky:1970:PDM

[Tau70] O. Taussky. Positive-definite matrices and their role in the study of the characteristic roots of general matrices. [Shi70], pages 389–400.

Taussky:1972:RSM

[Tau72] O. Taussky. The role of symmetric matrices in the study of general matrices. Linear Algebra and Appl., 5:147–154, 1972.

Taussky:1988:HBT

[Tau88] Olga Taussky. How I became a torchbearer for matrix theory. Amer. Math. Monthly, 95(9):801–812, November 1988.

Taylor:1982:IEA

[Tay82] John R. Taylor. An Introduction to Error Analysis: The Study of Uncertainties in Physical Measurements. University Science Books, Mill Valley, CA, 1982. ISBN 0-935702-07-5. xiii + 270 pp.

Berge:1984:ORM

[tBK84] Jos M. ten Berge and Dirk K. Knol. Orthogonal rotations to maximal agreement for two or more matrices of different column orders. Psychometrika, 49(1):49– 55, 1984. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Thisted:1987:DSW

[TE87] R. Thisted and B. Efron. Did Shakespeare write a newly-discovered poem? Biometrika, 74:445–455, 1987. ISSN 0006-3444 (print), 1464-3510 (electronic).

Teagle:1991:LC

[Tea91] F. Teagle. The LaTEX cookbook. Technical Report ETR 7/91, Engineering Applications Support Environment, ????, 1991. ???? pp.

Temperton:1975:ASC

[Tem75] Clive Temperton. Algorithms for the solution of cyclic tridiagonal systems. J. Comput. Phys., 19:317–323, 1975.

Terras:1985:HAS

[Ter85] Audrey Terras. Harmonic Analysis on Symmetric Spaces and Applications I. Springer-Verlag, Berlin, 1985. ISBN 0-387-96159-3. ???? pp.

Tang:1981:BDV

[TG81] W. P. Tang and G. H. Golub. The block decomposition of a Vandermonde matrix and its applications. BIT, 21:505– 517, 1981. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Thacher:1961:ACP

[Tha61] Henry C. Thacher, Jr. Algorithm 43:

Crout with pivoting II. Comm. ACM, 4 (4):176–177, 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

MathWorks:1993:MST

[The93] The MathWorks, Inc. Maple Symbolic Toolbox. The MathWorks, Inc., Natick, Massachusetts, May 1993. ???? pp. Beta test edition.

Thimbleby:1986:EPU

[Thi86] H. Thimbleby. Experiences of ‘literate programming’ using cweb (a variant of WEB). The Computer Journal, 29: 201–211, 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Thisted:1988:ESC

[Thi88] Ronald A. Thisted. Elements of Statistical Computing: Numerical Computation. Chapman and Hall, Ltd., London, UK, 1988. ISBN 0-412-01371-1. xx + 427 pp.

Thompson:1942:GFC

[Tho42] D’Arcy Wentworth Thompson. On Growth and Form. The Complete Revised Edition. Cambridge University Press, ????, 1942. ISBN 0-486-67135-6. viii + 1116 pp. Reprinted by Dover, New York, 1992.

Thompson:1983:AVR

[Tho83] Robert C. Thompson. Author vs.

referee: A case history for middle level mathematicians. Amer. Math. Monthly, 90(10):661–668, 1983.

Thompson:1988:IVW

[Tho88] T. Thompson. The Intel 80387 vs. the Weitek 1167. BYTE Magazine, 13: 205, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:HLQ

[Tho92] R. C. Thompson. High, low and quantitative roads in linear algebra. Linear Algebra and Appl., 162-164:23–64, 1992.

Tienari:1970:SMR

[Tie70] Martti Tienari. A statistical model of roundoff error for varying length floatingpoint arithmetic. BIT, 10:355–365, 1970. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Turk:1989:EWI

[TK89] Christopher Turk and John Kirkman. Effective Writing: Improving Scientific, Technical and Business Communication. E & FN Spon (Chapman and Hall), London, second edition, 1989. ISBN 0-419-14660-1. ???? pp.

Ting:1981:CPB

[TL81] B. Y. Ting and Y. L. Luke. Conversion of polynomials between different polynomial bases. IMA J. Numer. Anal., 1:229– 234, 1981.

Taussky:1962:EFM

[TM62] Olga Taussky and Marvin Marcus. Eigenvalues of finite matrices. In John Todd, editor, Survey of Numerical Analysis, pages 279–313. McGraw-Hill, New York, 1962.

Todd:1949:CCM

[Tod49] John Todd. The condition of certain matrices I. Quart. J. Mech. and Applied Math., 2(4):469–472, 1949.

Todd:1950:CCM

[Tod50] John Todd. The condition of a certain matrix. Proc. Cambridge Phil. Soc., 46(1):116–118, 1950. CODEN PCPSA4. ISSN 0008-1981.

Todd:1953:EIM

[Tod53] John Todd. Experiments on the inversion of a 16×16 matrix. In L. J. Paige and Olga Taussky, editors, Simultaneous Linear Equations and the Determination of Eigenvalues, number 29 in Applied Mathematics Series, pages 113–115. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1953.

Todd:1954:CCM

[Tod54a] John Todd. The condition of certain matrices II. Archiv der Mathematik (Basel), 5:249–257, 1954.

Todd:1954:CFS

[Tod54b] John Todd. The condition of the finite segments of the Hilbert matrix. In Olga Taussky, editor, Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, number 39 in Applied Mathematics Series, pages 109–116. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1954.

Todd:1961:CPC

[Tod61] John Todd. Computational problems concerning the Hilbert matrix. J. Res. Nat. Bur. Standards Sect. B, 65(1):19–22, 1961.

Todd:1962:SNA

[Tod62] John Todd, editor. Survey of Numerical Analysis. McGraw-Hill, New York, 1962. xvi + 589 pp.

Todd:1965:SIG

[Tod65] J. Todd. On smallest isolated Gerschgorin disks for eigenvalues. Numer. Math., 7:170–175, 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Todd:1968:CN

[Tod68] J. Todd. On condition numbers. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7 of Éditions Centre Nat. Recherche Sci., Paris, pages 141–159, 1968.

Todd:1977:BNM

[Tod77] John Todd. Basic Numerical Mathematics, Vol. 2: Numerical Algebra. Birkhäuser (Basel) and Academic Press (New York), ????, 1977. ISBN 0-12-692402-3. 216 pp.

Togai:1986:ASV

[Tog86] M. Togai. An application of the singular value decompositon to manipulability and sensitivity of industrial robots. SIAM J. Alg. Disc. Meth., 7(2):315–320, 1986.

Tompa:1988:FM

[Tom88] Martin Tompa. Figures of merit. Research Report RC 14211 (#63576), IBM Thomas J. Watson Research Center, Yorktown Heights, New York, November 1988. ???? pp.

Tommasini:1992:NAS

[Tom92] T. Tommasini. A new algorithm for special Vandermonde matrices. Numerical Algorithms, 2:299–306, 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tong:1992:CSP

[Ton92] Charles H. Tong. A comparative study of preconditioned Lanczos methods for nonsymmetric linear systems. Report SAND91-8240, Sandia National Laboratories, Albuquerque, New Mexico, January 1992. ???? pp.

Tornheim:1965:MTP

[Tor65] L. Tornheim. Maximum third pivot for Gaussian reduction. Technical report, Calif. Research Corp., Richmond, CA, 1965. ???? pp. Cited in [Cry68].

Torczon:1989:MDS

[Tor89] Virginia J. Torczon. Multi-Directional Search: A Direct Search Algorithm for Parallel Machines. PhD thesis, Rice University, Houston, TX, USA, May 1989. vii + 85 pp.

Torczon:1991:CMS

[Tor91] Virginia J. Torczon. On the convergence of the multidirectional search algorithm. SIAM J. Optimization, 1(1):123– 145, 1991.

Torczon:1992:PDS

[Tor92] Virginia J. Torczon. PDS: Direct search methods for unconstrained optimization on either sequential or parallel machines. Report TR92-9, Department of Mathematical Sciences, Rice University, Houston, TX, USA, March 1992. ???? pp. To appear in ACM Transactions on Mathematical Software.

Torczon:1993:CPS

[Tor93] Virginia J. Torczon. On the convergence of pattern search algorithms. Report TR93-10, Department of Mathematical Sciences, Rice University, Houston, TX, USA, June 1993. ???? pp. Submitted to SIAM J. Optimization.

Traub:1966:APU

[Tra66] J. F. Traub. Associated polynomials and uniform methods for the solution of linear problems. SIAM Review, 8(3):277– 301, 1966.

Traub:1973:CSP

[Tra73] J. F. Traub, editor. Complexity of Sequential and Parallel Numerical Algorithms. Academic Press, New York, USA, 1973. ISBN 0-12-697550-7. ix + 300 pp. LCCN QA76.6 .S918 1973.

Trefethen:1985:TMG

[Tre85] Lloyd N. Trefethen. Three mysteries of Gaussian elimination. ACM SIGNUM Newsletter, 20:2–5, 1985. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Trefethen:1990:ATN

[Tre90] Lloyd N. Trefethen. Approximation theory and numerical linear algebra. In J. C. Mason and M. G. Cox, editors, Algorithms for Approximation II, pages 336– 360. Chapman and Hall, London, 1990.

Trefethen:1992:DNA

[Tre92a] Lloyd N. Trefethen. The definition of numerical analysis. SIAM News, 25: 6, 22, November 1992. ISSN 0036-1437. Reprinted in IMA Bulletin, 29 (3/4), pp. 47–49, 1993.

Trefethen:1992:PM

[Tre92b] Lloyd N. Trefethen. Pseudospectra of matrices. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 234– 266. Longman Scientific and Technical, Harlow, Essex, UK, 1992.

Trefethen:19xx:SPB

[Trexx] Lloyd N. Trefethen. Spectra and Pseudospectra: The Behavior of Non-Normal Matrices and Operators. ????, 19xx. ???? pp. Book in preparation.

Tropp:1984:FA

[Tro84a] Henry S. Tropp. FORTRAN anecdotes. Annals of the History of Computing, 6(1):59–64, 1984.

Tropp:1984:OTE

[Tro84b] Henry S. Tropp. Origin of the term bit. Annals of the History of Computing, 6(2):152–155, 1984.

Trudgill:1985:CA

[Tru85] P. Trudgill. Coping with America.

Basil Blackwell, Oxford, second edition, 1985. ???? pp.

Trzeciak:1993:WMP

[Trz93] Jerzy Trzeciak. Writing Mathematical Papers in English: A Practical Guide. Gdańsk Teachers’ Press, P.O. Box 59, 80-876 Gdańsk 52, Poland, 1993. ISBN 83-85694-02-1. 48 pp.

Trapp:1975:SNV

[TS75] G. E. Trapp, Jr. and W. Squire.

Solving nonlinear Vandermonde systems. The Computer Journal, 18(4):373–374, 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Tredennick:1980:SGS

[TS80] Nick Tredennick and Brion Shimamoto. On systematic generation of scientific papers (or how to write a crummy paper). Research Report RC 8425 (#36568), IBM Thomas J. Watson Research Center, Yorktown Heights, New York, August 1980. ???? pp.

Trefethen:1990:ACS

[TS90] Lloyd N. Trefethen and Robert S. Schreiber. Average-case stability of Gaussian elimination. SIAM J. Matrix Anal. Appl., 11(3):335–360, 1990.

Tsao:1974:SPB

[Tsa74a] N.-K. Tsao. Some a posteriori bounds in floating-point computations. J. Assoc. Comput. Mach., 21:6–17, 1974.

Tsao:1974:DSD

[Tsa74b] Nai-kuan Tsao. On the distributions of significant digits and roundoff errors. Comm. ACM, 17(5):269–271, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tsao:1975:NIH

[Tsa75] Nai-kuan Tsao. A note on implementing the Householder transformation. SIAM J. Numer. Anal., 12(1):53– 58, 1975.

Tsao:1981:ECA

[Tsa81a] Nai-kuan Tsao. Error complexity analysis of algorithms for matrix multiplication and matrix chain product. IEEE Trans. Comput., C-30(10):758–771, 1981.

Tsao:1981:NIB

[Tsa81b] Nai-kuan Tsao. The numerical instability of Bini’s algorithm. Information Processing Letters, 12(1):17–19, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsao:1983:SAE

[Tsa83] Nai-kuan Tsao. A simple approach to the error analysis of division-free numerical algorithms. IEEE Trans. Comput., C-32(4):343–351, 1983.

Tsao:1991:AST

[Tsa91] Nai-kuan Tsao. On the accuracy of solving triangular systems in parallel. Applied Numerical Mathematics, 7:207–215, 1991.

Tsao:1992:AST

[Tsa92] Nai-kuan Tsao. On the accuracy of solving triangular systems in parallel—II. Applied Numerical Mathematics, 9:73–89, 1992.

Tsing:1990:NUC

[Tsi90] N.-K. Tsing. On the norms used in computing the structured singular value. IEEE Trans. Automat. Control., AC-35: 201–203, 1990.

Tsay:1986:FMC

[TST86] Yih T. Tsay, Leang S. Shieh, and Jason S. H. Tsai. A fast method for computing the principal nth roots of complex matrices. Linear Algebra and Appl., 76: 205–221, 1986.

Tsai:1988:FSA

[TSY88] J. S. H. Tsai, L. S. Shieh, and R. E.

Yates. Fast and stable algorithms for computing the principal nth root of a complex matrix and the matrix sector function. Computers and Mathematics with Applications, 15(11):903–913, 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Trefethen:1987:IPS

[TT87] Lloyd N. Trefethen and Manfred R. Trummer. An instability phenomenon in spectral methods. SIAM J. Numer. Anal., 24(5):1008–1023, 1987.

Taussky:1991:ALA

[TT91] Olga Taussky and John Todd. Another look at a matrix of Mark Kac. Linear Algebra and Appl., 150:341–360, 1991.

Toh:1994:PPP

[TT94] Kim-Chuan Toh and Lloyd N. Trefethen. Pseudozeros of polynomials and pseudospectra of companion matrices. Numer. Math., 68(3):403–425, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Toh:1996:CPA

[TT96] Kim-Chuan Toh and Lloyd N. Trefethen. Calculation of pseudospectra by the Arnoldi iteration. SIAM J. Sci. Comput., page ??, 1996. To appear.

Trefethen:1992:PLN

[TTR92] L. N. Trefethen, A. E. Trefethen, and S. C. Reddy. Pseudospectra of the linear Navier–Stokes evolution operator and instability of plane Poiseuille and Couette flows: Preliminary report. Technical Report TR 92-1291, Department of Computer Science, Cornell University, Ithaca, New York, June 1992. ???? pp.

Trefethen:1992:NDH

[TTRD92] L. N. Trefethen, A. E. Trefethen, S. C. Reddy, and T. A. Driscoll. A new direction in hydrodynamic stability: Beyond eigenvalues. Technical Report CTC 92 TR115, Advanced Computing Research Institute, Cornell Theory Center, Cornell University, Ithaca, New York, December 1992. ???? pp.

Trefethen:1993:HSE

[TTRD93] Lloyd N. Trefethen, Anne E. Trefethen, Satish C. Reddy, and Tobin A. Driscoll. Hydrodynamic stability without eigenvalues. Science, 261:578–584, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Tufte:1983:VDQ

[Tuf83] Edward R. Tufte. The Visual Display of Quantitative Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1983. ISBN 0-9613921-0-X, 0-9613921-4-2. 197 pp. LCCN K27.S8 T84. US$34.00.

Tufte:1988:PG

[Tuf88] Edward R. Tufte. The power of graphics. PC/Computing, pages 88–93, 1988. CODEN PCMPEI. ISSN 0899-1847 (print), 1081-8642 (electronic).

Tufte:1990:EI

[Tuf90] Edward R. Tufte. Envisioning Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1990. 126 pp.

Tulloch:1991:ODN

[Tul91] Sara Tulloch. The Oxford Dictionary of New Words. Oxford University Press, ????, 1991. ???? pp.

Turnbull:1929:TDM

[Tur29] H. W. Turnbull. The Theory of Determinants, Matrices, and Invariants. Blackie, Glasgow, Scotland, 1929. xvi + 338 pp.

Turing:1936:CNA

[Tur36] A. M. Turing. On computable numbers, with an application to the Entscheidungsproblem. Proc. London Math. Soc., 42:230–265, 1936.

Turing:1945:PDM

[Tur45] A. M. Turing. Proposal for development in the Mathematics Division of an Automatic Computing Engine (ACE). Report E.882, Executive Committee, National Physical Laboratory, Teddington, UK, 1945. ???? pp. Reprinted in [CD86, pp. 20-105] and [Inc92, pp. 1–86].

Turing:1948:REM

[Tur48] A. M. Turing. Rounding-off errors in matrix processes. Quart. J. Mech. and Applied Math., 1:287–308, 1948. Reprinted in [Bri92] with summary and notes (including corrections).

Turing:1950:CMI

[Tur50] A. M. Turing. Computing machinery and intelligence. MIND: A Quarterly Review of Pyschology and Philosophy, 59 (236):433–460, 1950. ISSN 0026-4423 (print), 1460-2113 (electronic). Reprinted in [Inc92].

Turing:1959:AMT

[Tur59] Sara Turing. Alan M. Turing. W. Heffer and Sons, Cambridge, 1959. xiv + 157 pp.

Turner:1974:ETW

[Tur74] Barry T. Turner. Effective Technical Writing and Speaking. Business Books, London, second edition, 1974. ISBN 0-220-66227-4. ???? pp.

Turner:1982:DLS

[Tur82] Peter R. Turner. The distribution of leading significant digits. IMA J. Numer. Anal., 2:407–412, 1982.

Turner:1984:FRD

[Tur84] Peter R. Turner. Further revelations on L.S.D. IMA J. Numer. Anal., 4:225– 231, 1984.

Turk:1985:ESC

[Tur85] Christopher Turk. Effective Speaking: Communicating in Speech. E & FN Spon (Chapman and Hall), London, 1985. ISBN 0-419-13030-6. ???? pp.

Turabian:1987:MWT

[Tur87] Kate L. Turabian. A Manual for Writers of Term Papers, Theses, and Dissertations. The University of Chicago Press, Chicago and London, fifth edition, 1987. ISBN 0-226-81625-7. ???? pp.

Turner:1989:NAP

[Tur89] Peter R. Turner. Numerical Analysis and Parallel Processing, Lancaster 1987, volume 1397 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1989. ???? pp.

Turner:1991:CPK

[Tur91a] Kathryn Turner. Computing projections for the Karmakar algorithm. Linear Algebra and Appl., 152:141–154, 1991.

Turner:1991:WRR

[Tur91b] Peter R. Turner. Will the “real”

real arithmetic please stand up? Notices Amer. Math. Soc., 38(4):298–304, 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trefethen:1991:TRP

[TW91a] L. N. Trefethen and J. A. C. Weideman. Two results on polynomial interpolation in equally spaced points. J. Approximation Theory, 65:247–260, 1991.

Trew:1991:PPP

[TW91b] Arthur Trew and Greg Wilson, editors. Past, Present, Parallel: A Survey of Available Parallel Computing Systems. Springer-Verlag, Berlin, 1991. ???? pp.

Traub:1992:PIB

[TW92a] J. F. Traub and H. Woźniakowski. Perspectives on information-based complexity. Bulletin of the American Mathematical Society, 26(1):29–52, 1992. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic). Reply to [Par92c].

Turner:1992:EHA

[TW92b] K. Turner and H. F. Walker. Efficient high accuracy solutions with GMRES(m). SIAM J. Sci. Stat. Comput., 12(3):815–825, 1992.

Tyrtyshnikov:1991:CTM

[Tyr91] Evgenij E. Tyrtyshnikov. Cauchy-Toeplitz matrices and some applications. Linear Algebra and Appl., 149:1–18, 1991.

Tyrtyshnikov:1994:HBH

[Tyr94] Evgenij E. Tyrtyshnikov. How bad are Hankel matrices? Numer. Math., 67 (2):261–269, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Taussky:1959:STB

[TZ59] O. Taussky and H. Zassenhaus. On the similarity transformation between a matrix and its transpose. Pacific J. Math., 9:893–896, 1959.

Uhlig:1981:EPD

[Uhl81] Frank Uhlig. Explicit polar decomposition and a near-characteristic polynomial: The 2 × 2 case. Linear Algebra and Appl., 38:239–249, 1981.

Uhlig:1985:FVC

[Uhl85] F. Uhlig. The field of values of a complex matrix, an explicit description in the 2 × 2 case. SIAM J. Alg. Disc. Meth., 6: 541–545, 1985.

Uhlig:1992:CPW

[Uhl92a] F. Uhlig. Are the coefficients of a polynomial well-conditioned functions of its roots? Numer. Math., 61:383–393, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Uhlig:1992:RTM

[Uhl92b] F. Uhlig. Review of Topics in Matrix Analysis by Roger A. Horn and Charles R. Johnson. Linear Algebra and Appl., 174:243–246, 1992.

Ulam:1976:AM

[Ula76] S. M. Ulam. Adventures of a Mathematician. Charles Scribner’s Sons, New York, 1976. ???? pp.

Underhill:1980:UYD

[Und80] Adrian Underhill. Use Your Dictionary: A Practice Book for Users of Oxford Advanced Learner’s Dictionary of Current English and Oxford Student’s Dictionary of Current English. Oxford University Press, ????, 1980. ISBN 0-19-431104-X. ???? pp.

Press:1969:MS

[Uni69] University of Chicago Press, Chicago, IL, USA. A Manual of Style, twelfth edition, 1969. ix + 546 pp.

Press:1982:CMS

[Uni82] University of Chicago Press. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, thirteenth edition, 1982. ix + 737 pp.

USGAO:1992:PMD

[Uni92] Patriot missile defense: Software problem led to system failure at Dhahran, Saudi Arabia. Report GAO/IMTEC-92-26, Information Management and Technology Division, United States General Accounting Office, Washington, DC, February 1992. 16 pp.

Urabe:1968:RED

[Ura68] Minoru Urabe. Roundoff error distribution in fixed-point multiplication and a remark about the rounding rule. SIAM J. Numer. Anal., 5(2):202–210, 1968.

Uspensky:1948:TE

[Usp48] J. V. Uspensky. Theory of Equations. McGraw-Hill, New York, 1948. vii + 353 pp.

Ullrich:1989:ANA

[UvG89] Ch. Ullrich and J. Wolff von Gudenberg. Accurate Numerical Algorithms: A Collection of Research Papers. Springer-Verlag, Berlin, 1989. ISBN 3-540-51477-5. ix + 234 pp.

Vignes:1986:ESM

[VA86] J. Vignes and R. Alt. An efficient stochastic method for round-off error analysis. In Willard L. Miranker and Richard A. Toupin, editors, Accurate Scientific Computations, Proceedings, 1985, volume 235 of Lecture Notes in Computer Science, pages 183–205. Springer-Verlag, Berlin, 1986.

Loan:1975:SME

[Van75] C. F. Van Loan. A study of the matrix exponential. Technical report, Numerical Analysis Report No. 10, University of Manchester, England, ????, 1975. ???? pp.

Loan:1976:GSV

[Van76a] C. F. Van Loan. Generalizing the singular value decomposition. SIAM J. Numer. Anal., 13:76–83, 1976.

Loan:1976:LLS

[Van76b] C. F. Van Loan. Lectures in least squares. Technical report, Technical Report TR 76-279, Department of Computer Science, Cornell University, ????, 1976. ???? pp.

Vel:1977:NTG

[Van77a] H. Van de Vel. Numerical treatment of a generalized Vandermonde system of equations. Linear Algebra and Appl., 17: 149–179, 1977.

Loan:1977:LAP

[Van77b] C. F. Van Loan. On the limitation and application of Padé approximation to the matrix exponential. In E. B. Saff and R. S. Varga, editors, Padé and Rational Approximation: Theory and Applications, pages 439–448. Academic Press, New York, USA, 1977.

Loan:1977:SME

[Van77c] C. F. Van Loan. The sensitivity of the matrix exponential. SIAM J. Numer. Anal., 14:971–981, 1977.

Veldhuizen:1977:NPP

[van77d] M. van Veldhuizen. A note on partial pivoting and Gaussian elimination.

Numer. Math., 29:1–10, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Loan:1978:CII

[Van78] C. F. Van Loan. Computing integrals involving the matrix exponential. IEEE Trans. Automat. Control., AC-23: 395–404, 1978.

Loan:1979:NEM

[Van79] C. F. Van Loan. A note on the evaluation of matrix polynomials. IEEE Trans. Automat. Control., AC-24:320–321, 1979.

VanDooren:1981:GEA

[Van81] P. M. Van Dooren. A generalized eigenvalue approach for solving Riccati equations. SIAM J. Sci. Stat. Comput., 2 (2):121–135, 1981.

Loan:1982:STC

[Van82a] C. F. Van Loan. Some thoughts on condition estimation for invariant subspaces and eigenvalues. In Gladwell [Gla82], pages 4.1–4.12. Numerical Analysis Report No. 78.

Loan:1982:UHD

[Van82b] C. F. Van Loan. Using the Hessenberg decomposition in control theory. In D. C. Sorensen and R. J.-B. Wets, editors, Algorithms and Theory in Filtering and Control, number 18 in Mathematical Programming Study, pages 102–111. North-Holland, Amsterdam, The Netherlands, 1982.

Loan:1983:GSA

[Van83a] C. F. Van Loan. A generalized SVD analysis of some weighting methods for equality constrained least squares. In Kågström and Ruhe [KR83], pages 245– 262.

Vandergraft:1983:INC

[Van83b] James S. Vandergraft. Introduction to Numerical Computations. Academic, New York, second edition, 1983. ISBN 0-12-711356-8. xiv + 372 pp.

Loan:1984:SMA

[Van84a] C. F. Van Loan. A symplectic method for approximating all the eigenvalues of a Hamiltonian matrix. Linear Algebra and Appl., 61:233–251, 1984.

Loan:1984:ASM

[Van84b] Charles F. Van Loan. Analysis of some matrix problems using the CS decomposition. Technical Report TR 84-603, Department of Computer Science, Cornell University, Ithaca, New York, March 1984. ???? pp.

Loan:1985:HNS

[Van85a] C. F. Van Loan. How near is a stable matrix to an unstable matrix? In B. N. Datta, editor, Linear Algebra and its Role in Systems Theory, volume 47 of Contemporary Math., pages 465–478. American Mathematical Society, Providence, RI, USA, 1985.

Loan:1985:CCG

[Van85b] Charles F. Van Loan. Computing the CS and the generalized singular value decompositions. Numer. Math., 46:479–491, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Loan:1985:MWE

[Van85c] Charles F. Van Loan. On the method of weighting for equalityconstrained least-squares problems. SIAM J. Numer. Anal., 22(5):851–864, 1985.

Loan:1986:BJM

[Van86a] Charles F. Van Loan. The block Jacobi method for computing the singular value decomposition. In C. I. Byrnes and A. Lindquist, editors, Computational and Combinatorial Methods in Systems Theory, pages 245–255. Elsevier North Holland, Inc., New York, NY, USA, 1986.

Vliet:1986:TPD

[van86b] J. C. van Vliet, editor. Text Processing and Document Manipulation. Cambridge University Press, ????, 1986.

VanHuffel:1987:ATL

[Van87a] S. Van Huffel. Analysis of The Total Least Squares Problem and its Use in Parameter Esimation. Ph.d. thesis, Katholieke Universiteit Leuven, Belgium, 1987. ???? pp.

Loan:1987:UME

[Van87b] C. F. Van Loan. A unitary method for the ESPRIT direction-of-arrival estimation algorithm. Technical report, Technical Report 87-862, Department of Computer Science, Cornell University, ????, 1987. ???? pp.

Loan:1987:ECE

[Van87c] Charles F. Van Loan. On estimating the condition of eigenvalues and eigenvectors. Linear Algebra and Appl., 88/89: 715–732, 1987.

VanWyk:1987:LPP

[Van87d] C. J. Van Wyk. Literate programming: Printing common words. Comm. ACM, 30:594–599, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loan:1988:BQF

[Van88] C. F. Van Loan. A block QR factorization scheme for loosely coupled systems of array processors. In Schultz [Sch88b], pages 217–232.

Loan:1989:MCS

[Van89] C. F. Van Loan. Matrix computations and signal processing. In Haykin [Hay89], pages 48–72.

Loan:1990:SMC

[Van90] C. F. Van Loan. A survey of matrix computations. Technical Report CTC 90 TR26, Advanced Computing Research Institute, Cornell Theory Center, Cornell University, Ithaca, New York, October 1990. ???? pp.

VanDooren:1991:SLA

[Van91] Paul M. Van Dooren. Structured linear algebra problems in digital signal processing. In G. H. Golub and P. M. Van Dooren, editors, Numerical Linear Algebra, Digital Signal Processing and Parallel Algorithms, volume F70 of NATO ASI Series, pages 361–384. Springer-Ver-lag, Berlin, 1991.

Loan:1992:CFF

[Van92] Charles F. Van Loan. Computational Frameworks for the Fast Fourier Transform. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-285-8 (paperback), 1-61197-099-7 (e-book). xiii + 273 pp. LCCN QA403.5 .V35 1992.

Vanderbei:1995:SQD

[Van95] Robert J. Vanderbei. Symmetric quasi-definite matrices. SIAM J. Optimization, 5(1):100–113, 1995.

Varga:1962:MIA

[Var62] Richard S. Varga. Matrix Iterative Analysis. Prentice-Hall, Englewood Cliffs, NJ, USA, 1962. ???? pp.

Varah:1972:SBT

[Var72] J. M. Varah. On the solution of blocktridiagonal systems arising from certain finite-difference equations. Math. Comp., 26(120):859–868, 1972. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1975:LBS

[Var75] J. M. Varah. A lower bound for the smallest singular value of a matrix. Linear Algebra and Appl., 11:3–5, 1975.

Varga:1976:DDA

[Var76a] Richard S. Varga. On diagonal dominance arguments for bounding ∥A^{−1}∥_{∞}. Linear Algebra and Appl., 14:211–217, 1976.

Varga:1976:RTD

[Var76b] Richard S. Varga. On recurring theorems on diagonal dominance. Linear Algebra and Appl., 13:1–9, 1976.

Varah:1979:STM

[Var79] J. M. Varah. On the separation of two matrices. SIAM J. Numer. Anal., 16(2): 216–222, 1979.

Varah:1986:GFM

[Var86a] J. M. Varah. A generalization of the Frank matrix. SIAM J. Sci. Stat. Comput., 7(3):835–839, 1986.

Varian:1986:PTM

[Var86b] H. R. Varian. PCTEX and MicroTEX. BYTE Magazine, pages 267–272, April 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Varah:1990:EPV

[Var90a] J. M. Varah. Errors and perturbations in Vandermonde systems. Manuscript, 1990.

Varah:1990:IGF

[Var90b] James M. Varah. The influence of George Forsythe and his students. In Stephen G. Nash, editor, A History of Scientific Computing, pages 31–40. Addison-Wesley, Reading, MA, USA, 1990.

Varga:1990:RUM

[Var90c] Richard S. Varga. Reminiscences on the University of Michigan, summer schools, the Gatlinburg symposia and numerische mathematik. In Stephen G. Nash, editor, A History of Scientific Computing, pages 206–210. Addison-Wes-ley, Reading, MA, USA, 1990.

Varga:1990:SCM

[Var90d] Richard S. Varga. Scientific Computation on Mathematical Problems and Conjectures. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-257-2. vi + 122 pp.

Varah:1991:BEE

[Var91] J. M. Varah. Backward error estimates for Toeplitz and Vandermonde systems. 1991.

Varah:1993:PM

[Var93] J. M. Varah. The Prolate matrix. Linear Algebra and Appl., 187:269–278, 1993.

Vassilevski:1990:SWA

[Vas90] P. S. Vassilevski. On some ways of approximating inverses of banded matrices in connection with deriving preconditioners based on incomplete block factorizations. Computing, 43:277–296, 1990. CO-DEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Vavasis:1989:GEP

[Vav89] S. A. Vavasis. Gaussian elimination with partial pivoting is P-complete. SIAM Journal on Discrete Mathematics, 2:413–423, 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vavasis:1991:NOC

[Vav91] Stephen A. Vavasis. Nonlinear Optimization: Complexity Issues. Oxford University Press, ????, 1991. ISBN 0-19-507208-1. ???? pp.

Vavasis:1992:SNA

[Vav92] S. A. Vavasis. Stable numerical algorithms for equilibrium systems. Technical Report TR 92-1280, Department of Computer Science, Cornell University, Ithaca, NY, April 1992. ???? pp.

Varga:1981:FM

[VC81] Richard S. Varga and Da-Yong Cai. On the LU factorization of M-matrices. Numer. Math., 38:179–192, 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Geijn:1993:DSS

[vdG93] Robert A. van de Geijn. Deferred shifting schemes for parallel QR methods. SIAM J. Matrix Anal. Appl., 14(1):180– 194, 1993.

Sluis:1969:CNE

[vdS69] A. van der Sluis. Condition numbers and equilibration of matrices. Numer. Math., 14:14–23, 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sluis:1970:CEP

[vdS70a] A. van der Sluis. Condition, equilibration and pivoting in linear algebraic systems. Numer. Math., 15:74–86, 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sluis:1970:SSL

[vdS70b] A. van der Sluis. Stability of solutions of linear algebraic systems. Numer. Math., 14:246–251, 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sluis:1975:PEN

[vdS75a] A. van der Sluis. Perturbations of eigenvalues of non-normal matrices. Comm. ACM, 18(1):30–36, 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sluis:1975:SSL

[vdS75b] A. van der Sluis. Stability of the solutions of linear least squares problems. Numer. Math., 23:241–254, 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sluis:1987:CBR

[vdSvdV87] A. van der Sluis and H. A. van der Vorst. The convergence behaviour of Ritz values in the presence of close eigenvalues. Linear Algebra and Appl., 88/89:651–694,

vanderVorst:1990:CBP

[vdV90] H. A. van der Vorst. The convergence behaviour of preconditioned CG and CG-S in the presence of rounding errors. In Owe Axelsson and Lily Yu. Kolotilina, editors, Preconditioned Conjugate Gradient Methods, volume 1457 of Lecture Notes in Mathematics, pages 126–136. Springer-Verlag, Berlin, 1990.

vanderVorst:1992:BCF

[vdV92] H. A. van der Vorst. Bi-CGSTAB: A fast and smoothly converging variant of Bi-CG for the solution of nonsymmetric linear systems. SIAM J. Sci. Stat. Comput., 12(2):631–644, 1992.

Vorst:1989:VLR

[vdVD89] H. A. van der Vorst and K. Dekker. Vectorization of linear recurrence relations. SIAM J. Sci. Stat. Comput., 10 (1):27–35, 1989.

Waerden:1975:SAE

[vdW75] B. L. van der Waerden. Science Awakening I: Egyptian, Babylonian and Greek Mathematics. Kluwer Academic, Dordrecht, Netherlands, 1975. ???? pp. Translated by A. Dresden.

Verzuh:1949:SSL

[Ver49] Frank M. Verzuh. The solution of simultaneous linear equations with the aid of the 602 calculating punch. M.T.A.C., 3:453–462, 1949.

Vetter:1975:VSS

[Vet75] W. J. Vetter. Vector structures and solutions of linear matrix equations. Linear Algebra and Appl., 10:181–188, 1975.

Valvi:1987:AID

[VG87] F. N. Valvi and V. S. Geroyannis.

Analytic inverses and determinants for a class of matrices. IMA J. Numer. Anal., 7:123–128, 1987.

Vigilans:1952:CHG

[‘Vi52] ‘Vigilans’. Chamber of Horrors: A Glossary of Official Jargon Both English and American. Andre Deutsch, London, 1952. ???? pp. With an introduction by Eric Partridge.

Vitenko:1968:OAA

[Vit68] I. V. Viten’ko. Optimum algorithms for adding and multiplying on computers with a floating point. U.S.S.R. Computational Math. and Math. Phys., 8(5): 183–195, 1968. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Vitasek:1969:NSS

[Vit69] Emil Vitasek. The numerical stability in solution of differential equations. In J. Ll. Morris, editor, Conference on the Numerical Solution of Differential Equations, volume 109 of Lecture Notes in Mathematics, pages 87–111. Springer-Verlag, Berlin, 1969.

Leunen:1978:HS

[vL78] Mary-Claire van Leunen. A Handbook for Scholars. Alfred A. Knopf, New York, NY, USA, 1978. ???? pp.

Leunen:1992:HS

[vL92] Mary-Claire van Leunen. A Handbook for Scholars. Oxford University Press, New York, revised edition, 1992. ISBN 0-19-506954-4. xi + 348 pp. First edition Knopf, 1978.

Matt:1994:RES

[vMS94] Urs von Matt and G. W. Stewart. Rounding errors in solving block Hessenberg systems. Technical Report TR-3345, Department of Computer Science, University of Maryland, College Park, MD, USA, September 1994. ???? pp.

Neumann:1947:NIM

[vNG47] John von Neumann and Herman H. Goldstine. Numerical inverting of matrices of high order. Bull. Amer. Math. Soc., 53:1021–1099, 1947. Reprinted in [Tau63, pp. 479–557].

VanLoan:1993:AKP

[VP93] C. F. Van Loan and N. Pitsianis. Approximation with Kronecker products. In Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors, Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E, pages 293–314. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993.

Vrscay:1986:JSM

[Vrs86] E. R. Vrscay. Julia sets and Mandelbrot-like sets associated with higher order schröder rational iteration functions: A computer assisted study. Math. Comp., 46:151–169, 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Venkaiah:1988:CMS

[VS88a] V. Ch. Venkaiah and S. K. Sen. Computing a matrix symmetrizer exactly using modified multiple modulus residue arithmetic. J. Comp. Appl. Math., 21:27– 40, 1988.

Verde-Star:1988:IGV

[VS88b] Luis Verde-Star. Inverses of generalized Vandermonde matrices. Journal of Mathematical Analysis and Applications, 131:341–353, 1988. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Veselic:1993:FPP

[VS93] Kres̆imar Veselić and Ivan Slapnic̆ar.

Floating-point perturbations of Hermitian matrices. Linear Algebra and Appl., 195:81–116, 1993.

VanHuffel:1987:SSU

[VV87] S. Van Huffel and J. Vandewalle. Subset selection using the total least squares approach in collinearity problems with errors in the variables. Linear Algebra and Appl., 88/89:695–714, 1987.

VanHuffel:1991:TLS

[VV91] Sabine Van Huffel and Joos Vandewalle. The Total Least Squares Problem: Computational Aspects and Analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-275-0. xiii + 300 pp.

Wijngaarden:1966:NAI

[vW66] A. van Wijngaarden. Numerical analysis as an independent science. BIT, 6: 66–81, 1966. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Waugh:1967:FPM

[WA67] F. V. Waugh and M. E. Abel. On fractional powers of a matrix. J. Amer. Statist. Assoc., 62:1018–1021, 1967.

Wachpress:1988:ISL

[Wac88] Eugene L. Wachpress. Iterative solution of the Lyapunov matrix equation.

Applied Mathematics Letters, 1(1):87–90, 1988. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Wachpress:1992:AIS

[Wac92] Eugene L. Wachpress. ADI iterative solution of Lyapunov equations. In R. Beauwens and P. de Groen, editors, Iterative Methods in Linear Algebra, pages 229–231. Elsevier North Holland, Inc., New York, NY, USA, 1992.

Wagon:1991:MA

[Wag91] Stan Wagon. Mathematica In Action. W.H. Freeman, San Francisco, CA, USA and New York, NY, USA, 1991. ISBN 0-7167-2202-X. xv + 419 pp.

Wahba:1965:PLS

[Wah65] G. Wahba. Problem 65-1: A least squares estimate of satellite attitude. SIAM Review, 7:409, 1965. Solutions in vol. 8, 1966, 384–386.

Wainwright:1985:WPW

[Wai85] A. Wainwright. Wainwright on the Pennine Way. Michael Joseph, London, 1985. ???? pp.

Walker:1971:BS

[Wal71] R. J. Walker. Binary summation. Comm. ACM, 14(6):417, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wallis:1982:PP

[Wal82] P. J. L. Wallis. Portable Programming. MacMillan Publishing Company, New York, NY, USA, 1982. ???? pp.

Wallis:1990:IFP

[Wal90] Peter J. L. Wallis, editor. Improving Floating-Point Programming. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1990. ISBN 0-471-92437-7. xvi + 191 pp.

Wallis:1993:HM

[Wal93] W. D. Wallis. Hadamard matrices.

In Richard A. Brualdi, Shmuel Friedland, and Victor Klee, editors, Combinatorial and Graph-Theoretical Problems in Linear Algebra, volume 50 of IMA Volumes in Mathematics and its Applications, pages 235–243. Springer-Verlag, New York, 1993. ISBN 0-387-94086-3.

Wand:1984:WL

[Wan84] Mitchell Wand. What is LISP?

Amer. Math. Monthly, 91(1):32–42, 1984.

Wang:1988:IBU

[Wan88] Paul S. Wang. An Introduction to Berkeley Unix. Wadsworth Pub. Co., Pacific Grove, CA, USA, 1988. ???? pp.

Wang:1991:GMM

[Wan91] Johnson J. H. Wang. Generalized Moment Methods in Electromagnetics: Formulation and Computer Solution of Integral Equations. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-51443-8. xiii + 553 pp.

Ware:1960:SCT

[War60] Willis H. Ware, editor. Soviet computer technology—1959. Comm. ACM, 3(3):131–166, 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ward:1976:QAH

[War76] Robert C. Ward. The QR algorithm and Hyman’s method on vector computers. Math. Comp., 30(133):132–142, 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ward:1977:SRE

[War77] R. C. Ward. Statistical roundoff error analysis of a Padé algorithm for computing the matrix exponential. In E. B. Saff and R. S. Varga, editors, Padé and Rational Approximation: Theory and Applications, pages 449–460. Academic Press, New York, USA, 1977.

Warburton:1989:DD

[War89] Y. L. Warburton. From dictionary to database. The Stag (Journal of the Society of Authors), 4:??, 1989.

Watson:1977:TMD

[Wat77] G. A. Watson. On two methods for discrete L_{p}-approximation. Computing, 18:263–266, 1977. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Watson:1980:ATN

[Wat80a] G. A. Watson. Approximation Theory and Numerical Methods. Wiley, Chichester, UK, 1980. ???? pp.

Watson:1980:DHP

[Wat80b] James D. Watson. The Double Helix: A Personal Account of the Discovery of the Structure of DNA. W.W. Norton, New York, NY, USA, 1980. ISBN 0-393-95075-1. ???? pp. Text, commentary, reviews and original papers, edited by Gunther S. Stent.

Watkins:1982:UQA

[Wat82] D. S. Watkins. Understanding the QR algorithm. SIAM Review, 24:427– 440, 1982.

Waterman:1983:PTP

[Wat83a] Fanny Waterman. On Piano Teaching and Performing. Faber Music, London, 1983. ???? pp.

Watson:1983:TAP

[Wat83b] G. A. Watson. The total approximation problem. In C. K. Chui, L. L. Schumaker, and J. D. Ward, editors, Approximation Theory IV, pages 723–728. Academic Press, New York, USA, 1983.

Wathen:1987:REB

[Wat87] A. J. Wathen. Realistic eigenvalue bounds for the Galerkin mass matrix. IMA J. Numer. Anal., 7:449–457, 1987.

Watson:1988:SPSa

[Wat88a] G. A. Watson. The smallest perturbation of a submatrix which lowers the rank of the matrix. IMA J. Numer. Anal., 8:295–303, 1988.

Watson:1988:SPSb

[Wat88b] G. A. Watson. The smallest perturbations of submatrices for rankdeficiency. 1988.

Waterhouse:1989:NS

[Wat89] Keith Waterhouse. On Newspaper Style. Viking, London, 1989. ISBN 0-670-82626-X. ???? pp.

Watson:1990:CSS

[Wat90] G. A. Watson. Computing the structured singular value, and some related problems. In Griffiths and Watson [GW90], pages 258–275.

Waterhouse:1991:EOE

[Wat91a] Keith Waterhouse. English our English (and How to Sing It). Viking, London, 1991. ISBN 0-670-83269-3. ???? pp.

Watkins:1991:FMC

[Wat91b] David S. Watkins. Fundamentals of Matrix Computations. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-54601-1. ???? pp.

Watson:1991:AOS

[Wat91c] G. A. Watson. An algorithm for optimal ℓ_{2} scaling of matrices. IMA J. Numer. Anal., 11:481–492, 1991.

Watson:1991:GCM

[Wat91d] G. A. Watson. On a general class of matrix nearness problems. Constructive Approximation, 7:299–314, 1991. ISSN 0176-4276 (print), 1432-0940 (electronic).

Watkins:1992:SSP

[Wat92a] D. S. Watkins. Shifting strategies for the parallel QR algorithm. May 1992.

Watson:1992:AMT

[Wat92b] G. A. Watson. Algorithms for minimum trace factor analysis. SIAM J. Matrix Anal. Appl., 13(4):1039–1053, 1992.

Watson:1992:CSSa

[Wat92c] G. A. Watson. Characterization of the subdifferential of some matrix norms. Linear Algebra and Appl., 170: 33–45, 1992.

Watson:1992:CSSb

[Wat92d] G. A. Watson. Computing the structured singular value. SIAM J. Matrix Anal. Appl., 13(4):1064–1066, 1992.

Watson:1992:EHO

[Wat92e] G. A. Watson. Estimating Hadamard operator norms, with application to triangular truncation. 1992.

Watson:1992:LBA

[Wat92f] G. A. Watson. Linear best approximation using a class of polyhedral norms. Numerical Algorithms, 2:321–326, 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Watson:1992:MAP

[Wat92g] G. A. Watson. On matrix approximation problems with Ky Fan k norms. 1992.

Watkins:1993:SPE

[Wat93a] David S. Watkins. Some perspectives on the eigenvalue problem. SIAM Review, 35(3):430–471, 1993.

Watson:1993:SGO

[Wat93b] G. A. Watson. Solving generalizations of orthogonal Procrustes problems. In R. P. Agarwal, editor, Contributions in Numerical Mathematics, volume 2, pages 413–426. World Scientific, Singapore, 1993.

Watkins:1994:NLM

[Wat94] A. J. Watkins. A note on Laguerre’s method for finding the roots of polynomials. IMA Bulletin, 30(5/6):91, 1994.

Watkins:1995:FST

[Wat95] David S. Watkins. Forward stability and transmission of shifts in the QR algorithm. SIAM J. Matrix Anal. Appl., 16 (2):469–487, 1995.

Watson:1996:CNR

[Wat96] G. A. Watson. Computing the numerical radius. Linear Algebra and Appl., page ??, 1996. To appear.

Watson:1953:MSN

[WC53] J. D. Watson and F. H. C. Crick. Molecular structure of nucleic acids: A structure for deoxyribose nucleic acid. Nature, 171(4356):737–738, 1953. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). April 25.

Williams:1989:EBC

[WCK89] M. R. Williams and Martin Campbell-Kelly, editors. The Early British Computer Conferences, volume 14 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-23136-0. xvi + 508 pp.

Watkins:1991:CAE

[WE91] D. S. Watkins and L. Elsner. Chasing algorithms for the eigenvalue problem. SIAM J. Matrix Anal. Appl., 12(2):374– 384, 1991.

Weber:1973:RWS

[Web73] Robert L. Weber. A Random Walk In Science. The Institute of Physics, Bristol and London, 1973. ISBN 0-85498-027-X. xvii + 206 pp. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Weber:1982:MRW

[Web82] Robert L. Weber. More Random Walks In Science. The Institute of Physics, Bristol and London, 1982. ISBN 0-85498-040-7. xv + 208 pp. LCCN Q167.M67 1982.

Weber:1992:SS

[Web92a] Robert L. Weber. Science with a Smile. Institute of Physics Publishing, Bristol and Philadelphia, 1992. ISBN 0-7503-0211-9. 452 pp. LCCN Q167.S34 1992.

Webster:1992:RMS

[Web92b] Michael Webster. Review of “Mathematica: A System for Doing Mathematics by Computer, Second Edition, by Stephen Wolfram”. SIAM Review, 34(3): 519–522, 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wedderburn:1934:LM

[Wed34] J. H. M. Wedderburn. Lectures On Matrices, volume 17 of American Mathematical Society Colloquium Publications. American Mathematical Society, Providence, RI, USA, 1934. vii + 205 pp.

Wedin:1972:PBC

[Wed72] Per-Åke Wedin. Perturbation bounds in connection with singular value decomposition. BIT, 12:90–111, 1972. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Wedin:1973:ARD

[Wed73a] P.-Å. Wedin. On the almost rank deficient case of the least squares problem. BIT, 13:344–354, 1973. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Wedin:1973:PTP

[Wed73b] Per-Åke Wedin. Perturbation theory for pseudo-inverses. BIT, 13:217–232, 1973. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Wedin:1978:PPI

[Wed78] P.-Å. Wedin. Perturbations of the pseudo-inverse and the linear least squares problem analysed with suitable matrix decompositions. Technical report, Report UMINF 69.78, Department of Information Processing, University of Umeå, Sweden, ????, 1978. ???? pp.

Wedin:1985:PTC

[Wed85] Per-Åke Wedin. Perturbation theory and condition numbers for generalized and constrained linear least squares problems. Report UMINF 125.85, Institute of Information Processing, University of Umeå, Umeå, Sweden, May 1985. ???? pp.

Weeks:1989:VRI

[Wee89] D. Weeks. Vectorizing a robust inner product algorithm. Technical report, Proceedings of 1989 International Conference on Supercomputing, Crete, Greece, ACM Publications, ????, 1989. ???? pp.

Weiner:1985:COE

[Wei85] Edmund Weiner. Computerizing the Oxford English Dictionary. Scholarly Publishing, ??:239–253, 1985. ISSN 0036-634X.

Wei:1990:PLS

[Wei90] Musheng Wei. Perturbation of the least squares problem. Linear Algebra and Appl., 141:177–182, 1990.

Wendroff:1966:TNA

[Wen66] Burton Wendroff. Theoretical Numerical Analysis. Academic Press, New York, USA, 1966. xi + 239 pp.

Werner:1984:PIL

[Wer84] Wilhelm Werner. Polynomial interpolation: Lagrange versus Newton. Math. Comp., 43(167):205–217, 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Westlake:1968:HNM

[Wes68] Joan R. Westlake. A Handbook of Numerical Matrix Inversion and Solution of Linear Equations. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1968. ???? pp.

Wexelblat:1981:HPL

[Wex81] R. L. Wexelblat. History of Programming Languages. Academic Press, New York, 1981. ???? pp.

Waser:1982:IAD

[WF82] Shlomo Waser and Michael J. Flynn.

Introduction to Arithmetic for Digital Systems Designers. Holt, Rinehart and Winston, New York, 1982. ISBN 0-03-060571-7. xvii + 308 pp.

Wood:1989:CEU

[WFF89] Frederick T. Wood, Roger H.

Flavell, and Linda M. Flavell. Current English Usage. MacMillan Publishing Company, New York, NY, USA, 1989. ISBN 0-333-27840-2. ???? pp.

Ward:1978:ECS

[WG78] R. C. Ward and L. J. Gray. Eigensystem computation for skew-symmetric matrices and a class of symmetric matrices. ACM Transactions on Mathematical Software, 4:278–285, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wright:1916:AOM

[WH16] Thomas Wallace Wright and John Fillmore Hayworth. The Adjustment of Observations by the Method of Least Squares with Applications to Geodetic Work. Constable, London, second edition, 1916. ix + 298 pp.

Whitemore:1987:BCS

[Whi87] Hugh Whitemore. Breaking the Code: The Story of Alan Turing. 1987. ???? pp.

Wichmann:1989:TFS

[Wic89] B. A. Wichmann. Towards a formal specification of floating point. The Computer Journal, 32:432–436, 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wichmann:1992:NUF

[Wic92] B. A. Wichmann. A note on the use of floating point in critical systems. The Computer Journal, 35(1):41– 44, 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wichmann:1993:PSC

[Wic93] B. A. Wichmann. PostScript—A study in conformity. NPL Report DITC 214/93, National Physical Laboratory, England, April 1993. ???? pp.

Widom:1965:TM

[Wid65] H. Widom. Toeplitz matrices. In Studies in Real and Complex Analysis [Hir65], pages 179–209.

Williamson:1942:GPR

[Wil42] J. Williamson. A generalization of the polar representation of nonsingular matrices. Bulletin of the American Mathematical Society, 48:856–863, 1942. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Wilkinson:1948:ACE

[Wil48a] J. H. Wilkinson. The Automatic Computing Engine at the National Physical Laboratory. Proc. Roy. Soc. London Ser. A, 195:285–286, 1948.

Wilkinson:1948:PRA

[Wil48b] J. H. Wilkinson. Progress report on the Automatic Computing Engine. Report MA/17/1024, Mathematics Division, Department of Scientific and Industrial Research, National Physical Laboratory, Teddington, UK, April 1948. 127 pp.

Wilkinson:1954:LAP

[Wil54a] J. H. Wilkinson. Linear algebra on the Pilot ACE. In Automatic Digital Computation, page ?? Her Majesty’s Stationery Office, London, 1954. Reprinted in [WCK89, pp. 337-344].

Wilkinson:1954:PA

[Wil54b] J. H. Wilkinson. The Pilot ACE. In Automatic Digital Computation, pages 5– 14. Her Majesty’s Stationery Office, London, 1954. Reprinted in [BN71, pp. 193– 199] and [WCK89, pp. 219–228].

Wilkinson:1955:ECM

[Wil55a] J. H. Wilkinson. Electronic computing machines and their uses. Journal of Scientific Instruments, 32:409–415, 1955. CODEN JSINAY. ISSN 0368-4253, 0950-7671.

Wilkinson:1955:UIM

[Wil55b] J. H. Wilkinson. The use of iterative methods for finding the latent roots and vectors of matrices. Mathematical Tables and Other Aids to Computation, 9:184– 191, 1955.

Wilkinson:1958:CEC

[Wil58] J. H. Wilkinson. The calculation of the eigenvectors of codiagonal matrices. The Computer Journal, 1:90–96, 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilkinson:1959:SRM

[Wil59] J. H. Wilkinson. Stability of the reduction of a matrix to almost triangular and triangular forms by elementary similarity transformations. J. Assoc. Comput. Mach., 6:336–359, 1959.

Wilf:1960:ADM

[Wil60a] H. S. Wilf. Almost diagonal matrices. Amer. Math. Monthly, 67:431–434, 1960.

Wilkinson:1960:EAF

[Wil60b] J. H. Wilkinson. Error analysis of floating-point computation. Numer. Math., 2:319–340, 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1960:SMS

[Wil60c] J. H. Wilkinson. Symposium on methods for solving linear systems. In Proc. International Conference on Information Processing, Unesco, Paris, pages 108–109. Butterworth, London, June 1960.

Wilkinson:1961:EAD

[Wil61a] J. H. Wilkinson. Error analysis of direct methods of matrix inversion. J. Assoc. Comput. Mach., 8:281–330, 1961.

Wilkinson:1961:REB

[Wil61b] J. H. Wilkinson. Rigorous error bounds for computed eigensystems. The Computer Journal, 4:230–241, 1961. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilkinson:1962:EAE

[Wil62] J. H. Wilkinson. Error analysis of eigenvalue techniques based on orthogonal transformations. J. Soc. Indust. Appl. Math., 10(1):162–195, 1962.

Wilkinson:1963:PRF

[Wil63a] J. H. Wilkinson. Plane-rotations in floating-point arithmetic. In Experimental Arithmetic, High Speed Computing and Mathematics, volume 15 of Proceedings of Symposia in Applied Mathematics, pages 185–198. American Mathematical Society, Providence, RI, USA, 1963.

Wilkinson:1963:REA

[Wil63b] J. H. Wilkinson. Rounding Errors in Algebraic Processes, volume 32 of Notes on Applied Science. Her Majesty’s Stationery Office, London, 1963. ISBN 0-486-67999-3. vi + 161 pp. Also published by Prentice-Hall, Englewood Cliffs, NJ, USA. Reprinted by Dover, New York, 1994.

Wilkinson:1965:AEP

[Wil65a] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Oxford University Press, ????, 1965. ISBN 0-19-853403-5 (hardback), 0-19-853418-3 (paperback). xviii + 662 pp.

Wilkinson:1965:CLP

[Wil65b] J. H. Wilkinson. The calculation of Lamé polynomials. The Computer Journal, 8:273–277, 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilkinson:1965:EAT

[Wil65c] J. H. Wilkinson. Error analysis of transformations based on the use of matrices of the form I − 2ww^{H}. In Louis B. Rall, editor, Error In Digital Computation, volume 2, pages 77–101. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1965.

Williams:1965:MFB

[Wil65d] L. Pearce Williams. Michael Faraday: A Biography. Chapman and Hall, Ltd., London, UK, 1965. ???? pp.

Wilkinson:1967:BZW

[Wil67a] J. H. Wilkinson. Bledy Zaokragleń w Procesach Algebraicznych. PWW, Warszawa, 1967. ???? pp. Polish translation of [Wil63b].

Wilkinson:1967:SIC

[Wil67b] J. H. Wilkinson. The solution of illconditioned linear equations. In Ralston and Wilf [RW67], pages 65–93.

Wilkinson:1968:ADM

[Wil68a] J. H. Wilkinson. Almost diagonal matrices with multiple or close eigenvalues. Linear Algebra and Appl., 1:1–12, 1968.

Wilkinson:1968:PEA

[Wil68b] J. H. Wilkinson. A priori error analysis of algebraic processes. In I. G. Petrovsky, editor, Proc. International Congress of Mathematicians, Moscow 1966, pages 629–640. Mir Publishers, Moscow, 1968.

Wilkinson:1968:SEA

[Wil68c] James Hardy Wilkinson. A survey of error analysis of matrix algorithms. Aplikace Matematiky, 13:93–102, 1968.

Wilkinson:1969:R

[Wil69] J. H. Wilkinson. Rundungsfehler.

Springer-Verlag, Berlin, 1969. ???? pp. German translation of [Wil63b].

Wilkinson:1970:AEP

[Wil70] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Nauka, Moscow, Russia, 1970. 564 pp. Russian translation of [Wil65a].

Wilkinson:1971:MEA

[Wil71a] J. H. Wilkinson. Modern error analysis. SIAM Review, 13(4):548–568, 1971.

Wilkinson:1971:SCN

[Wil71b] J. H. Wilkinson. Some comments from a numerical analyst. J. Assoc. Comput. Mach., 18(2):137–147, 1971. Reprinted in [ACM87].

Williams:1971:SCMa

[Wil71c] L. Pearce Williams, editor. The Selected Correspondence of Michael Faraday: Volume 1, 1812–1848. Cambridge University Press, Cambridge, 1971. ISBN 0-521-07908-X. ???? pp.

Williams:1971:SCMb

[Wil71d] L. Pearce Williams, editor. The Selected Correspondence of Michael Faraday: Volume 2, 1849–1866. Cambridge University Press, Cambridge, 1971. ISBN 0-521-07913-6. ???? pp.

Wilder:1972:APS

[Wil72a] Alec Wilder. American Popular Song: The Great Innovators, 1900–1950. Oxford University Press, ????, 1972. ???? pp. 1990 reprint with new foreword by Gene Lees.

Wilkinson:1972:IIT

[Wil72b] J. H. Wilkinson. Inverse iteration in theory and in practice. Symposia Mathematica, X:361–379, 1972. ISSN 0082-0725.

Wilkinson:1972:NMV

[Wil72c] J. H. Wilkinson. Note on matrices with a very ill-conditioned eigenproblem. Numer. Math., 19:176–178, 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1974:CEA

[Wil74a] J. H. Wilkinson. The classical error analyses for the solution of linear systems. IMA Bulletin, 10(5/6):175–180, 1974.

Wilkinson:1974:NLA

[Wil74b] J. H. Wilkinson. Numerical linear algebra on digital computers. IMA Bulletin, 10(9/10):354–356, 1974.

Wilkes:1975:ECD

[Wil75a] M. V. Wilkes. Early computer developments at Cambridge: The EDSAC. The Radio and Electronic Engineer, 45 (7):332–335, 1975.

Wilkinson:1975:IS

[Wil75b] J. H. Wilkinson. Invariant subspaces. In Proc. International Congress of Mathematicians, Vancouver 1974 [Jam74], pages 443–448.

Wilkinson:1975:PAA

[Wil75c] J. H. Wilkinson. The Pilot ACE at the National Physical Labotatory. The Radio and Electronic Engineer, 45(7): 336–340, 1975.

Williams:1975:ECA

[Wil75d] F. C. Williams. Early computers at Manchester university. The Radio and Electronic Engineer, 45(7):327–331, 1975.

Wilkinson:1976:LDE

[Wil76] J. H. Wilkinson. Linear differential equations and Ax = λBx. [HW76b], pages 137–158.

Wilkinson:1977:DSG

[Wil77a] J. H. Wilkinson. The differential system Bẋ = Ax and the generalized eigenvalue problem Au = λBu. Technical report, NPL Report NAC 73, National Physical Laboratory, England, ????, 1977. ???? pp.

Wilkinson:1977:SRA

[Wil77b] J. H. Wilkinson. Some recent advances in numerical linear algebra. In David A. H. Jacobs, editor, The State of the Art in Numerical Analysis, pages 3– 53. Academic Press, London, 1977.

Wilkinson:1977:USP

[Wil77c] J. H. Wilkinson. The use of the single-precision residual in the solution of linear systems. Unpublished manuscript, 1977.

Wilkinson:1978:PNL

[Wil78a] J. H. Wilkinson. A problem in numerical linear algebra. In The Contribution of Dr. J. H. Wilkinson to Numerical Analysis, Symposium Proceedings Series No. 19, pages 76–91. The Institute of Mathematics and its Applications, Southend-On-Sea, Essex, UK, 1978.

Wilkinson:1978:SVD

[Wil78b] J. H. Wilkinson. Singular-value decomposition—basic aspects. In D. A. H. Jacobs, editor, Numerical Software—Needs and Availability, pages 109–135. Academic Press, London, 1978.

Wilkinson:1979:KCF

[Wil79] J. H. Wilkinson. Kronecker’s canonical form and the QZ algorithm. Linear Algebra and Appl., 28:285–303, 1979.

Wilkinson:1980:TWA

[Wil80] J. H. Wilkinson. Turing’s work at the National Physical Laboratory and the construction of Pilot ACE, DEUCE, and ACE. In N. Metropolis, J. Howlett, and Gian-Carlo Rota, editors, A History of Computing in the Twentieth Century: A Collection of Essays, pages 101–114. Academic Press, New York, USA, 1980.

Wilcox:1984:HWR

[Wil84a] Alan D. Wilcox. How to write a recommendation. IEEE Trans. Prof. Commun., PC-27(4):211–214, 1984.

Wilkinson:1984:NMQ

[Wil84b] J. H. Wilkinson. On neighbouring matrices with quadratic elementary divisors. Numer. Math., 44:1–21, 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1984:PP

[Wil84c] James H. Wilkinson. The perfidious polynomial. In G. H. Golub, editor, Studies in Numerical Analysis, volume 24 of Studies in Mathematics, pages 1–28. Mathematical Association of America, Washington, DC, 1984.

Wilkinson:1985:SAE

[Wil85] J. H. Wilkinson. The state of the art in error analysis. In NAG Newsletter 2/85, pages 5–28. Numerical Algorithms Group, Oxford, UK, November 1985.

Wilf:1986:TNR

[Wil86a] Herbert S. Wilf. TEX: A non-review.

Amer. Math. Monthly, 93:309–315, 1986.

Wilkinson:1986:EAR

[Wil86b] J. H. Wilkinson. Error analysis revisited. IMA Bulletin, 22(11/12):192–200, 1986.

Wilkinson:1986:SEI

[Wil86c] J. H. Wilkinson. Sensitivity of eigenvalues II. Utilitas Mathematica, 30:243– 286, 1986.

Wilkinson:1987:EP

[Wil87a] J. H. Wilkinson. Eigenvalue problems. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 1–39. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987.

Wilkinson:1987:TF

[Wil87b] J. H. Wilkinson. On a theorem of Feingold. Linear Algebra and Appl., 88/89:13–30, 1987.

Williams:1988:SW

[Wil88a] N. Williams. Style wars. Personal computer world, 10:140–143, 1988.

Wilson:1988:FPS

[Wil88b] P. Wilson. Floating-point survival kit. BYTE Magazine, 13:217–226, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilf:1989:CAU

[Wil89] Herbert S. Wilf. Combinatorial Algorithms: An Update. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1989. ISBN 0-89871-231-9. 45 pp.

Wille:1995:ASF

[Wil95] David R. Willé. Advanced Scientific Fortran. Wiley, Chichester, UK, 1995. ISBN 0-471-95383-0. xiii + 234 pp.

Winograd:1968:NAI

[Win68] S. Winograd. A new algorithm for inner product. IEEE Trans. Comput., C-18:693–694, 1968.

Winograd:1971:MM

[Win71] S. Winograd. On multiplication of 2 × 2 matrices. Linear Algebra and Appl., 4:381–388, 1971.

Walden:1995:OBP

[WKS95] Bertil Waldén, Rune Karlson, and Ji-guang Sun. Optimal backward perturbation bounds for the linear least squares problem. Numerical Linear Algebra with Applications, 2(3):271–286, 1995.

Woelfle:1992:NGB

[Woe92] Robert M. Woelfle, editor. A New Guide for Better Technical Presentations: Applying Proven Techniques with Modern Tools. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-87942-283-1. ???? pp.

Wolfe:1964:RTE

[Wol64] Jack M. Wolfe. Reducing truncation errors by programming. Comm. ACM, 7(6):355–356, 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wolfe:1965:ESL

[Wol65] Philip Wolfe. Error in the solution of linear programming problems. In Louis B. Rall, editor, Error In Digital Computation, volume 2, pages 271–284. Wiley, New York, NY, USA; London, UK; Sydney, Australia, 1965.

Wolper:1985:ART

[Wol85] Roy S. Wolper. On academic reviewing: Ten common errors. Scholarly Publishing, ??:269–275, 1985. ISSN 0036-634X.

Wolfe:1989:OSS

[Wol89] M. J. Wolfe. Optimizing Supercompilers for Supercomputers. Pitman Publishing Ltd., London, UK, 1989. ???? pp.

Wolfram:1991:MSD

[Wol91] Stephen Wolfram. Mathematica: A System for Doing Mathematics by Computer. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-51507-5. xxii + 961 pp.

Wong:1988:UJF

[Won88] compiled by Herb Wong. The Ultimate Jazz Fake Book. Hal Leonard, ????, 1988. ???? pp.

Wong:1990:ACA

[Won90] R. Wong, editor. Asymptotic and Computational Analysis, volume 124 of Lecture Notes in Pure and Applied Mathematics. Marcel Dekker, Inc., New York, NY, USA, 1990.

Woodger:1958:HPU

[Woo58] Michael Woodger. The history and present use of digital computers at the National Physical Laboratory. Process Control and Automation, pages 437–442, November 1958. CODEN PCAUA3. ISSN 0552-234X. Reprinted in [CD86, pp. 125-140].

Wood:1962:CEU

[Woo62] F. T. Wood. Current English Usage: A Concise Dictionary. Macmillan and Co. Ltd., London, 1962. ???? pp.

Woodford:1967:STT

[Woo67] F. Peter Woodford. Sounder thinking through clearer writing. Science, 156 (3776):743–745, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Woods:1985:IAS

[Woo85] D. J. Woods. An Interactive Approach for Solving Multi-objective Optimization Problems. Ph.d. thesis, Rice University, Houston, Texas, 1985. ???? pp.

Woodgate:1989:NCL

[Woo89] K. G. Woodgate. A new constrained Lyapunov-type problem. Linear Algebra and Appl., 121:697–702, 1989.

Woodwark:1992:HRP

[Woo92] John R. Woodwark. How to Run a Paper Mill: Writing Technical Papers and Getting them Published. Information Geometers, Winchester, 1992. ISBN 1-874728-00-3. xv + 111 pp.

Worland:1989:MF

[Wor89] Peter B. Worland. Modern Fortran 77. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1989. ISBN 0-15-561156-9. ???? pp.

Wouk:1965:IRL

[Wou65] Arthur Wouk. Integral representation of the logarithm of matrices and operators. Journal of Mathematical Analysis and Applications, 11:131–138, 1965. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Wouk:1987:NCE

[Wou87] A. Wouk, editor. New Computing Environments: Microcomputers in Large-Scale Computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987.

Wozniakowski:1974:REA

[Woź74] H. Woźniakowski. Rounding error analysis for the evaluation of a polynomial and some of its derivatives. SIAM J. Numer. Anal., (4):780–787, 1974.

Wozniakowski:1977:NSS

[Woź77a] H. Woźniakowski. Numerical stability for solving nonlinear equations. Numer. Math., 27:373–390, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1977:NSC

[Woź77b] H. Woźniakowski. Numerical stability of the Chebyshev method for the solution of large linear systems. Numer. Math., 28:191–209, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1978:REA

[Woź78] H. Woźniakowski. Roundoff-error analysis of iterations for large linear systems. Numer. Math., 30:301–314, 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1980:REA

[Woź80] H. Woźniakowski. Roundoff-error analysis of a new class of conjugategradient algorithms. Linear Algebra and Appl., 29:507–529, 1980.

Waser:1992:WEM

[WPG92] Nickolas M. Waser, Mary V. Price, and Richard K. Grosberg. Writing an effective manuscript review. BioScience, 42(8):621–623, 1992. CODEN BISNAS. ISSN 0006-3568 (print), 1525-3244 (electronic).

Troupe:1989:MA

[wQT89] Miles Davis with Quincy Troupe. Miles—The Autobiography. Pan Books, London, 1989. ???? pp.

Whittaker:1944:COT

[WR44] Sir Edmund Whittaker and G. Robinson. The Calculus of Observations: A Treatise on Numerical Mathematics. Blackie & Son, London and Glasgow, fourth edition, 1944. xvi + 397 pp.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. viii + 439 pp. LCCN QA251 .W67.

Wright:1982:APM

[Wri82] K. Wright. Asymptotic properties of matrices associated with the quadrature method for integral equations. In Baker and Miller [BM82], page ??

Wright:1990:PAQ

[Wri90] K. Wright. Parallel algorithms for QR decomposition on a shared memory multiprocessor. Technical report, Parallel Processing Memorandum PPM/007, Computing Laboratory, University of Newcastle upon Tyne, ????, 1990. ???? pp.

Wright:1992:IMC

[Wri92] Margaret H. Wright. Interior methods for constrained optimization. In Acta Numerica, pages 341–407. Cambridge University Press, 1992.

Wright:1993:CPW

[Wri93] Stephen J. Wright. A collection of problems for which Gaussian elimination with partial pivoting is unstable. SIAM J. Sci. Stat. Comput., 14(1):231–238, 1993.

Wright:1995:SLE

[Wri95] Stephen J. Wright. Stability of linear equations solvers in interior-point methods. SIAM J. Matrix Anal. Appl., 16(4): 1287–1307, 1995.

Wallis:1972:CRS

[WSW72] W. D. Wallis, Anne Penfold Street, and Jennifer Seberry Wallis. Combinatorics: Room Squares, Sum-Free Sets, Hadamard Matrices, volume 292 of Lecture Notes in Mathematics. Springer-Ver-lag, Berlin, 1972. ISBN 3-540-06035-9. 508 pp.

Wegert:1994:BNP

[WT94] Elias Wegert and Lloyd N. Trefethen.

From the Buffon needle problem to the Kreiss matrix theorem. Amer. Math. Monthly, 101(2):132–139, 1994.

Wundheiler:1948:NEA

[Wun48] Alexander W. Wundheiler. The necessity of error analysis in numerical computation. In Anonymous [Ano48], pages 83–90. ISBN 0-262-08152-0. Reprinted, with a new introduction by William Aspray, Volume 7 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

Wilkes:1951:PPE

[WWG51] Maurice V. Wilkes, David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer. With Special Reference to the EDSAC and the Use of a Library of Subroutines. Addison-Wesley, Reading, MA, USA, 1951. ???? pp. Reprinted by Tomash Publishers, Los Angeles, 1982, Volume 1 in the Charles Babbage Institute Reprint Series for the History of Computing, with a new Introduction by Martin Campbell-Kelly.

Weiss:1986:ACN

[WWWS86] N. Weiss, G. W. Wasilkowski, H. Woźniakowski, and M. Shub. Average condition number for solving linear equations. Linear Algebra and Appl., 83: 79–102, 1986.

Wigner:1963:ICD

[WY63] E. P. Wigner and M. M. Yanase. Information contents of distributions. Proceedings of the National Academy of Sciences, 49:910–918, 1963.

Wigner:1964:PSN

[WY64] E. P. Wigner and M. M. Yanase. On the positive semidefinite nature of a certain matrix expression. Canadian Journal of Mathematics, 16:397–406, 1964. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yalamov:1991:BSG

[Yal91] P. Y. Yalamov. On the backward stability of Gauss–Jordan elimination. Computing, 47:193–197, 1991. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Yalamov:1994:REG

[Yal94] P. Y. Yalamov. Roundoff errors and graphs. Manuscript, 1994.

Yau:1993:RSM

[Yau93] Ya Yan Lu Shing-Tung Yau. Reducing the symmetric matrix eigenvalue problem to matrix multiplications. SIAM J. Sci. Comput., 14(1):121–136, 1993.

Yeung:1995:PAG

[YC95] Man-Chung Yeung and Tony. F.

Chan. Probabilistic analysis of Gaussian elimination without pivoting. CAM Report 95-29, Department of Mathematics, University of California, Los Angeles, USA, June 1995. 32 pp.

Young:1972:SNM

[YG72] David M. Young and Robert T. Gregory. A Survey of Numerical Mathematics, volume 1. Addison-Wesley, Reading, MA, USA, 1972. ISBN 0-486-65691-8. x + 492 + appendices pp. Reprinted by Dover, New York, 1988.

Young:1973:SNM

[YG73] David M. Young and Robert T. Gregory. A Survey of Numerical Mathematics, volume 2. Addison-Wesley, Reading, MA, USA, 1973. ISBN 0-486-65692-6. xi + 1099 + appendices: numbering continues from Vol 1. pp. Reprinted by Dover, New York, 1988.

Yamamoto:1979:IBM

[YI79] Tetsuro Yamamoto and Yasuhiko Ikebe. Inversion of band matrices. Linear Algebra and Appl., 24:105–111, 1979.

Yip:1986:NSS

[Yip86] E. L. Yip. A note on the stability of solving a rank-p modification of a linear system by the Sherman–Morrison– Woodbury formula. SIAM J. Sci. Stat. Comput., 7(3):507–513, 1986.

Yohe:1973:RFP

[Yoh73] J. Michael Yohe. Roundings in floating-point arithmetic. IEEE Trans. Comput., C-22(6):577–586, 1973.

Yohe:1979:INA

[Yoh79a] J. M. Yohe. Implementing nonstandard arithmetics. SIAM Review, 21(1): 34–56, 1979.

Yohe:1979:SIA

[Yoh79b] J. M. Yohe. Software for interval arithmetic: A reasonably portable package. ACM Transactions on Mathematical Software, 5(1):50–63, 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Yohe:1980:PSI

[Yoh80] J. M. Yohe. Portable software for interval arithmetic. Computing, Suppl. 2:211–229, 1980. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Young:1954:IMS

[You54] D. M. Young. Iterative methods for solving partial difference equations of elliptic type. Trans. Amer. Math. Soc., 76: 92–111, 1954.

Young:1971:ISL

[You71] David M. Young. Iterative Solution of Large Linear Systems. Academic Press, New York, USA, 1971. ISBN 0-12-773050-8. xxiv + 570 pp.

Young:1979:NPM

[You79] N. J. Young. Norms of powers of matrices with constrained spectra. Linear Algebra and Appl., 23:227–244, 1979.

Young:1988:IHS

[You88] N. Young. An Introduction to Hilbert Space. Cambridge University Press, ????, 1988. ???? pp.

Young:1989:HOI

[You89a] David M. Young. A historical overview of iterative methods. Computer Physics Communications, 53 (1–3):1–17, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465589901458.

Young:1989:TWH

[You89b] Matt Young. The Technical Writer’s Handbook: Writing with Style and Clarity. University Science Books, Mill Valley, CA, 1989. ISBN 0-935702-60-1. ???? pp.

Young:1990:HRI

[You90] David M. Young. A historical review of iterative methods. In Stephen G. Nash, editor, A History of Scientific Computing, pages 180–194. Addison-Wesley, Reading, MA, USA, 1990.

Yakowitz:1989:INC

[YS89] Sidney Yakowitz and Ferenc Szidarovszky. An Introduction to Numerical Computations. MacMillan Publishing Company, New York, NY, USA, 1989. ISBN 0-02-430821-8. xiv + 462 pp.

Yuval:1978:SPS

[Yuv78] Gideon Yuval. A simple proof of Strassen’s result. Information Processing Letters, 7(6):285–286, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zawilski:1991:NSC

[Zaw91] Adam T. Zawilski. Numerical stability of the cyclic Richardson iteration. Numer. Math., 60:251–290, 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zha:1992:TWC

[Zha92] H. Zha. A two-way chasing scheme for reducing a symmetric arrowhead matrix to tridiagonal form. J. Numer. Lin. Alg. Appl., 1(1):49–57, 1992.

Zha:1993:CPA

[Zha93a] Hongyuan Zha. A componentwise perturbation analysis of the QR decomposition. SIAM J. Matrix Anal. Appl., 4 (4):1124–1131, 1993.

Zha:1993:P

[Zha93b] Hongyuan Zha. Problem 10312. Amer. Math. Monthly, 100(5):499, 1993.

Zielke:1974:TMK

[Zie74] G. Zielke. Testmatrizen mit maximaler konditionszahl. Computing, 13:33– 54, 1974. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Zielke:1975:TFP

[Zie75] G. Zielke. Testmatrizen mit freien parametern. Computing, 15:87–103, 1975. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Zielke:1978:IRG

[Zie78] G. Zielke. Iterative refinement of generalized matrix inverses now practicable. ACM SIGNUM Newsletter, 13:9–10, 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zielinski:1979:ICM

[Zie79] Adam Zielinski. On inversion of complex matrices. International J. Numer. Methods Engineering, 14:1563–1566, 1979.

Zietak:1983:PMS

[Ziȩ83] K. Ziȩtak. The properties of the minimax solution of a non-linear matrix equation XY = A. IMA J. Numer. Anal., 3: 229–244, 1983.

Zielke:1984:SGM

[Zie84] G. Zielke. A survey of generalized matrix inverses. Computational Mathematics, 13:499–526, 1984. Banach Center Publications, PWN-Polish Scientific Publishers, Warsaw.

Zietak:1985:PCI

[Ziȩ85] K. Ziȩtak. On a particular case of the inconsistent linear matrix equation AX + Y B = C. Linear Algebra and Appl., 66: 249–258, 1985.

Zielke:1986:RTM

[Zie86] G. Zielke. Report on test matrices for generalized inverses. Computing, 36:105–162, 1986. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Zietak:1987:PSL

[Ziȩ87] K. Ziȩtak. Properties of the ℓ_{1}- solutions of the linear matrix equation AX + Y B = C. IMA J. Numer. Anal., 7: 223–233, 1987.

Zielke:1988:SRM

[Zie88] Gerhard Zielke. Some remarks on matrix norms, condition numbers, and error estimates for linear equations. Linear Algebra and Appl., 110:29–41, 1988.

Zinsser:1983:WWP

[Zin83] William Zinsser. Writing With a Word Processor. Harper & Row, New York, NY, USA, 1983. ISBN 0-06-091060-7. ???? pp.

Zinsser:1988:WWI

[Zin88a] William Zinsser. On Writing Well: An Informal Guide to Writing Nonfiction. Harper & Row, New York, NY, USA, third edition, 1988. ???? pp.

Zinsser:1988:WL

[Zin88b] William Zinsser. Writing to Learn. Harper & Row, New York, NY, USA, 1988. ISBN 0-06-091576-5. ???? pp.

Zinsser:1990:WWI

[Zin90] William Zinsser. On Writing Well: An Informal Guide to Writing Nonfiction. HarperCollins, New York, fourth edition, 1990. ISBN 0-06-272027-9. ???? pp.

Ziv:1982:RDE

[Ziv82] Abraham Ziv. Relative distance—An error measure in round-off error analysis. Math. Comp., 39(160):563–569, 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ziv:1983:SME

[Ziv83] Abraham Ziv. A stable method for the evaluation of a polynomial and of a rational function of one variable. Numer. Math., 41:309–319, 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ziv:1991:FEE

[Ziv91] Abraham Ziv. Fast evaluation of elementary mathematical functions with correctly rounded last bit. ACM Transactions on Mathematical Software, 17(3): 410–423, 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ziv:1995:CAE

[Ziv95] Abraham Ziv. Converting approximate error bounds into exact ones. Math. Comp., 64(209):265–277, 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zlatev:1982:UIR

[Zla82] Zahari Zlatev. Use of iterative refinement in the solution of sparse linear systems. SIAM J. Numer. Anal., 19(2):381– 399, 1982.

Zwillinger:1992:HI

[Zwi92] Daniel Zwillinger. Handbook of Integration. Jones and Bartlett, Boston, MA, USA, 1992. ISBN 0-86720-293-9. xv + 367 pp.

Zlatev:1986:CNE

[ZWS86] Zahari Zlatev, Jerzy Wasniewski, and Kjeld Schaumburg. Condition number estimators in a sparse matrix software. SIAM J. Sci. Stat. Comput., 7(4): 1175–1189, 1986.